A Complete Bibliography of Publications in Journal of Computational Chemistry: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

12 June 2023 Version 1.07

Title word cross-reference

(Φ, Ψ) [WD04]. (φ, ψ) [TTB01b]. −2 [ZZZ^{+}06]. −4 [ZZZ^{+}06]. 1 [FROD08]. 1/r^{p} [SG01]. 12 [KGL07]. $139.00 [Sta00]. 14 [LMGR06, ZGXX06]. 16 [HXD08]. 2 [DMC05, GdAcV^{+}07, LW04b, LXZ06, LW06, PF06, SFC04, WDZS07, YNW05, hYDN^{+}08, ZTS09]. 2 ≤ N ≤ 372 [Pul05]. $3 [Lip00]. 3 [BAH^{+}02, FROD08, GDPCPU07, GdSuM^{+}07, HP05, mJlZyL^{+}08, KSN01, LYZ^{+}08, MP03b, PF06, SHBD05, SFC04, TJM^{+}03, WLL07a, WDZS07, XLL^{+}02, ZLY07]. 3_{10} [JS07b]. 3d [Hol05, LD05a, Wu06, SZT08]. 4 [FHF^{+}01, LLXS02, QB05, WDX^{+}02]. 4, 7 [ZWS^{+}02]. 4d [ABYM08]. 5 [ZXL^{+}04]. 5d [HZ09]. 6 [Han01, LLXS02]. 7 [LMO09]. 8 [KS05a]. 9 [UM03].

[n] [VKP^{+}08]. ^{##} [BRV^{+}07]. ^{+} [FKŠ^{+}09, GLRL02, GPSP06, GWL07, HDO^{+}02, ITS06, KT08, LMO09, Mck07b, PV07, Sha02, WWT08, ZWY^{+}09, ZL05, dRLMS00]. ^{+2} [PNG08]. ^{−} [Bac09, Dib05, HTN03, LYZ^{+}08, Mas01a]. ^{1} [AGI^{+}07, IS03, LDT^{+}02a, XZ04, ZX04, ZLLS04b]. ^{125} [HWFN01]. ^{13} [FHF^{+}01, VBS09, ZLD09]. ^{15} [LKA01]. ^{19} [FO04]. ^{1}Σ [BAL^{+}01]. ^{1}Σ_{g}^{+} [BAL^{+}01]. ^{2} [BAL^{+}01, Gog08, PRSMM03, ZWS^{+}02]. ^{2+} [AS06, BTP09, BL00, FHRR07, GPSP06, HLB09, KRLD09, Kri08, ROG00, TFZRG01]. ^{2−} [IvSV06, JD09]. ^{2}Π [Gog08]. ^{3} [BPC01, LDT^{+}02b]. ^{3+} [BTP09, BL00, Kri09b]. ^{5} [WC08]. ^{57} [HLLN06]. ^{α} [VBS09, CJW^{+}09]. ^{•} [Mui05]. ^{ISQ} [Bac09]. ^{MT CP} 1 [BRDC02]. ^{n} [ZZZ^{+}06]. ^{p+} [GPSP06]. _{0} [DLD^{+}02]. _{1} [BPC01, DLD^{+}02, WLPF05, XWXC08, YKK09, ZZW09, XWXC08]. _{103} [NSO^{+}07]. _{11} [XB08]. _{12} [WZZ^{+}09, XB08]. _{141} [GYCZ04]. _{18} [ZZZ^{+}06]. _{2} [Bac09, BAL^{+}01, BL00, BBI^{+}09, CPJ00, CCCJ09, CTFC08, DLD^{+}02, DRAS04, DRAS05, Dib05, DMN05, Don08, FJ08, GCCVB00, GBDP05, GPSP06, HYR06, HK07, Hua09b, IN08, IV04, IvSV06, IS03, JPF^{+}00, mJlZsLyL07, KZY09, KS05a, KKJH08, KBL08, KS05c, LDMR01, LC07, LWK08, LLXS02, LW04a, LZZC09, LAT05, LF02, LDT^{+}02b, LWY^{+}09, LMO09, MR02, Mas01a, Mas01b, MGLL03, McD08, Mck07b, MY08a, NA06, NyHN06, NHN06, Owe05, PGNG03, PGRRNG03, PC00, PRSMM03, PLC08, RAGLL09a, RAGLL09b, RD00, RDM^{+}08, SLL^{+}04a, SLL^{+}04b, SRE08, STC^{+}08, TJM^{+}03, TD08, UCT^{+}03, UTT^{+}04, VS02, WLLS05, WDWS06, WDS06, WLL07a, WDZS07, WLL^{+}07b, WD08, XFF06, YTY07, YLW^{+}08, YLWL09, ZY01, ZZL04, ZLLS04b, ZWL^{+}05, ZLLS06a, ZLLS06b, ZKZ^{+}07, ZXY03, ZX09, dOMSL01, dRLMS00]. _{20} [Ber03, CCB04, Ell07, GB02, Var09, WLPF05, WL09b, ZZZ^{+}06]. _{2}^{+} [dOMSL01]. _{2}^{−} [LMCD09]. _{2}^{2+} [GHLK^{+}02]. _{2}^{2−} [GYMN07]. _{2A} [YKK09]. _{3} [BPC01, CPJ00, DRAS04, DMN05, FJ08, GBDP05, GD06, HLLS05, HYA02, HDO^{+}02, HTN03, ITS06, mJlZsLyL07, mJlZyL^{+}08, KZY09, KT08, KKJH08, KSTC01, LDMR01, LD05a, LMK01, LDC^{+}07, LW04a, LWW^{+}06, LF02, LWY^{+}09, MR02, Mas01b, MGLL03, Mck07b, Mui05, NA06, ON07, PGNG03, PGRRNG03, RD00, SAS05, SLL^{+}04a, SM06, SCP08, TJM^{+}03, TK08, TDK07, UTM^{+}02, UTT^{+}04, VDM06, WDS06, WLL^{+}07b, WDX^{+}02, XDS06a, XFF06, YLW^{+}08, YLWL09, ZJM^{+}07, ZZL04, ZLLS06a, ZZW^{+}07, ZZL^{+}08, ZXYF09, ZZL^{+}09, ZKZ^{+}07, ZWS^{+}02, ZX08]. _{3−n} [mJlZyL^{+}08, KSN01, WLL07a, WDZS07]. _{36} [CS01]. _{4} [DRAS04, DMN05, IN08, IvSV06, JD09, LDMR01, LMK01, Mck07a, PHH^{+}08, TYN05, TK08, UNM^{+}01, UTM^{+}02]. _{4−n} [FHF^{+}01, WDX^{+}02, XLL^{+}02]. _{5} [GD06, LYZ^{+}08, SLL^{+}04b, ZZL04, XWXC08]. _{50} [ZZvRSC08]. _{54} [GZ07]. _{6} [BSJ01, GHLK^{+}02, Han01, KRLD09, LDMR01, LZZC09, LAT05, LYZ^{+}08, RPNJ07, SLL^{+}04b, TYN05, XB08]. _{6−n} [Han01]. _{60} [CJS^{+}03, Ell07, KJP^{+}07, Owe05]. _{62} [HK07, STC^{+}08]. _{64} [CTFC08]. _{64}^{4−} [CTFC08]. _{7} [HYA02, LZZC09]. _{8} [GZ07, Mck07a, Mck07b, PLC08, WLL01]. 80 [KSN01]. _{9} [Mck07a, Mck07b, PLC08]. _{90} [NSO^{+}07]. _{A} [Sha02, KKS04, MK02, ZPL07]. _{cam} [ZAT07]. _{carbene} [HA04]. _{h} [WLPF05]. _{i} [OS08]. _{m} [LYZ^{+}08]. _{N} [BP01, BSB05, HTN03, MG00, Owe05, RC04, YT04, DRAS05, FHF^{+}01, Han01, HXD08, mJlZyL^{+}08, KS05a, KSN01, LLXS02, LMGR06, LMO09, Mas01a, Mas01b, QB05, TJM^{+}03, UM03, WLL07a, WDZS07, WDX^{+}02, XLL^{+}02, ZY01, ZXL^{+}04, ZXY03, ZGXX06]. _{S} [WLL01]. x [KGL07, RAGLL09a]. A [Dib05]. α [CMD^{+}04, CPML08b, FKU^{+}05, IGNH03, KLS02, LZZC09, LDY^{+}08, LLL07, MCF07, MML^{+}06, PCO^{+}07b, PCO^{+}07a, RY09, SWBM08, SBG^{+}09a, TFZRG01, Van08, WCF04, ZWPR^{+}04]. α_{3} [MH09]. β [AGO^{+}02, APG05, BTP09, BCP04, BSP06a, CSJ01, CJW^{+}09, CLA^{+}00, ESM06, GTC06, GAS04, HL08, IGNH03, KF03, LLW02, MCF07, MML^{+}06, MS03, MK02, PP08a, PGC05, PSF^{+}08, PMM05, RMP01, SBG^{+}09a, SDM02, TFZRG01, TK08, ZALMG03, ZHH09]. β_{2} [YKK09]. β_{2}m [WCF04]. c [BS06, JKL08, KK01a, OON01]. · [VS02, ZY01]. χ [OMNH08]. d [ASS^{+}02]. Δ [CFR06]. ΔΔ [AGI^{+}07]. η^{2} [KJP^{+}07]. η^{6} [FKŠ^{+}09, GM01]. f [ASS^{+}02]. g [KRM^{+}02]. γ [HL08, Ike04, TK08]. k [DLW06]. λ [CN03, KB09]. ↔ [UCT^{+}03]. m = 1 [LYZ^{+}08]. μ_{3} [KJP^{+}07]. N [BY06, CCK01, GSB09, HT05, HRBKB03, HJCP01, JD09, KK09, KKH^{+}07, KK01a, MMPK01, ZPL07, MGMM07a, MGMM07b, VVBV02]. n = 0 [FHF^{+}01, mJlZyL^{+}08, KSN01, LLXS02, TJM^{+}03, ZZZ^{+}06]. n = 1 [Han01, KS05a, UM03, WLL07a, WDZS07, WDX^{+}02, XLL^{+}02, ZXL^{+}04]. n = 1, 2 [DRAS05]. n = 14 [LMGR06]. n = 2 [QB05, ZGXX06]. n = 4 [HXD08]. n = 5 [LMO09]. n = 8 [LMGR06]. nd^{10}(n + 1)s^{0} [GPSP06]. o [OSA06]. Φ [SO09]. π [Ang09, BY06, CXZ^{+}09, DMZT08, EFQD09, GYMN07, HIM07, IINK09, LC09, LFR07, MM05, OO08, RRCA08, SSS^{+}09, ST01, WRP^{+}06, Wou00, WJ00, YTH^{+}07, ZGZX07]. π^{∗} [Ang09, XZ04, ZX04]. pK_{a} [KK08b, SMG09]. Ψ [SO09]. → [Ang09, CMD^{+}04, Gog08, HTN03, LW04a, MGLL03, PGNG03, SOOF05, SLL^{+}04b, ZZL04]. s [ATM^{+}07, GRO^{+}03]. s/d [Hol05, SZT08]. σ [CXZ^{+}09, HIM07]. sp^{2} [TTB01a]. V [Ang09]. x = 4 [KGL07].

-1 [PWFS01, ZPL07]. -2-chalcogena-trans-hydrindans [HKHN08]. -2-methyloxirane [ZPL07]. -34 [PHH^{+}08]. -acetaldehyde [Lu09, Lu09]. -adrenergic-based [YKK09]. -alkanes [VVBV02]. -amino [CLA^{+}00, LDY^{+}08]. -amyloid [MS03]. -arene [FKŠ^{+}09]. -assisted [KT08]. -ATPase [HLB09]. -based [JD09]. -benzodiazepinone [SPGS08]. -benzylideneaniline [BY06]. -block [ATM^{+}07, ASS^{+}02]. -C [KJP^{+}07]. -carbon [KLS02]. -carboranes [OSA06]. -center [MGMM07a, MGMM07b]. -channel [CXZ^{+}09]. -cis [GRO^{+}03]. -conjugated [LC09, LFR07]. -D [BAH^{+}02]. -dependent [PMM05]. -di-substitution [ZWS^{+}02]. -diketiminate [GTC06]. -diketonate [RMP01]. -dimethylcarbamate [KKH^{+}07]. -Dynamics [KB09]. -effect [RY09]. -electrocyclization [ZGZX07]. -electron [BY06]. -electrons [WJ00]. -formyl- [HJCP01]. -formyl-alanine [GSB09]. -formyl-glycine [GSB09]. -formylglycinamide-water [HRBKB03]. -glucans [CMD^{+}04]. -glutamine [WC08]. -H [LLL07]. -Hairpin [ZHH09, CJW^{+}09, IGNH03]. -Hairpins [IGNH03]. -Helical [CPML08b, Van08, PCO^{+}07b, PCO^{+}07a]. -helices [IGNH03]. -helix [JS07b]. -hydrogen [ZGZX07]. -jun [KK01a]. -lactamase [AGO^{+}02, APG05, SDM02]. -lactamases [ESM06, MK02]. -lyase [PMM05]. -maltotriose [SWBM08]. -mercaptocarboxamides [TFZRG01]. -metal [Wu06]. -methoxycarbonyl [KK09]. -methyl [CCK01]. -methylacetamide [MMPK01]. -methylimidazole [HT05]. -methyloxaziridine [ZPL07]. -Mn [AZM03]. -peptide [BSP06a, CSJ01, CLA^{+}00, ZALMG03]. -peptides [BCP04, LLW02]. -phthalocyanines [LS02]. -pleated [PGC05]. -porphyrazines [LS02]. -quartz [ZWPR^{+}04]. -R [LZZC09]. -radical [CXZ^{+}09, CXZ^{+}09]. -scaled [CN03]. -sheet [KF03, PP08a]. -Si [TK08]. -sigmatropic [LLKC06]. -SO [ZZW09]. -stacked [RRCA08]. -stacking [WRP^{+}06]. -step [DLW06]. -strand [GAS04]. -substituents [PSF^{+}08]. -TCNE [GYMN07]. -tensors [KRM^{+}02]. -terminal [KK01a]. -the [VHRR07b]. -TMPyP [AZM03]. -turns [HL08]. -type [OON01].

/As [KS05a]. /CBS [Lu09]. /CCl [ZLLS06a]. /CF [YLWL09]. /Cs [GLRL02]. /empirical [CYM02]. /free [BG00]. /Ge [LLXS02]. /GeH [LLXS02]. /Mn [BL00]. /poly [BSJ01].

0 [Bic09, CG06, Lip00, Sta00]. 0-470-03735-0 [Bic09]. 0-471-33135-X [Sta00]. 0-471-96588-X [Lip00].

1 [AJNG01, AVS09, BPC01, BWE05, KF08, NLL^{+}09, SPT^{+}03, VVS07, WC09, WHF08, YT04]. 1- [CUS00]. 1-amino-3-propenal [FDSA00]. 1-jk [SPGS08]. 1-naphthoic [CMLS05]. 1-pK [ZCS04]. 1-X-bicyclo [BPC01]. 1.0 [AGSFA^{+}05]. 10-endoperoxide [CG08]. 10-membered [ZW09]. 1174 [WWC^{+}05]. 12 [HDO^{+}02]. 12-crown-O [HDO^{+}02]. 1291 [Ano06a]. 1328 [Ano05b]. 142 [Ano06b]. 1629 [WB04a]. 1Z [Lu09].

2 [BL08, LBG08, MM02, PFR04b, Höf05, HTN03, MG00, PO03, RC04, WC09]. 2-allyl-2 [ZGZX07]. 2-amino-2-imidazoline [XKG^{+}05]. 2-amino-2-oxazoline [XKG^{+}05]. 2-amino-2-thiazoline [XKG^{+}05]. 2’-aminoarabinonucleosides [BL08]. 2-Aminopurine [SC01]. 2’-aminoribonucleosides [BL08]. 2’-azido-2’-deoxyribonucleoside [PFR04b]. 2-chloro- [JKM08]. 2’-deoxyguanosine [MM02]. 2’-deoxyribose [LBG08]. 2-difluoroethane [CUS00]. 2-dihaloethanes [WFR08]. 2-dihydro-pyrimidinyl [WJX^{+}08]. 2-dimethoxyethane [LCGA03]. 2-electron [ABF^{+}03]. 2-H [LDC^{+}07]. 2-oxazolidones [OY01]. 2-oxo-1 [WJX^{+}08]. 2-oxoimidazoles [JKM08]. 2-phenoxycarboxylic [XKKL03]. 2-substituted [OSA06, WW04]. 2-thiouracil [LMGO^{+}09]. 2.0 [GZM09]. 21C7 [GLRL02]. 21D8 [UTH^{+}03]. 22 [Bof01, Qua01]. 24 [Ano06b]. 25 [Kne05, WB04a, WWC^{+}05]. 26 [Ano05b]. 27 [Ano06a, Ano06c]. 28 [HNWF12, Van08]. 2B6 [LCC09]. 2B6-substrate [LCC09]. 2H [dSVA^{+}09]. 2H-tetrazole [dSVA^{+}09].

3 [AAP00, KAK^{+}09, LEV^{+}09, NYTH09, Rud05c]. 3-butadiene [Hir08, WR07]. 3-dihydroxypyridine [YXZ^{+}04]. 3-dimethylallene [ZPL07]. 3-fluorobutanal [NSB08]. 3-hydroxy-2-mercaptopyridine [YXZ^{+}04]. 3-trimethylsilyl-1-pyrazoline [LLKC06]. 311 [Wib04]. 31G [BRV^{+}07]. 31G* [RRP^{+}01, FKJ^{+}01, NL08]. 31G** [WD04]. 31G*/AMBER [FSFK05]. 34 [PHH^{+}08]. 3a [HKHN08]. 3D [DHW^{+}08, SGPS09, CGMPT^{+}08, CPUGD09, MH08b, VB09]. 3D-chiral [CGMPT^{+}08]. 3D-RISM [MH08b]. 3D-structure-function [CPUGD09]. 3DEX [SHBD05].

4 [Lu09]. 4-carbonate [vDSSvA04]. 4-dihydropyridine-based [HSMT04]. 4-dimethylamino-benzonitrile [ZH08]. 4-dithiacyclohexane [FD03]. 4-dithiane [FD03]. 4-Spinor [PV03]. 43C9 [CBS^{+}03]. 45-ns [SO07]. 4d [CWWS07]. 4R [BISB02]. 4R-hydroxyproline [BISB02].

5 [LDY^{+}08, PFR04b, PMM05, TAS07]. 5-di-tert-butyl-o-diiminobenzosemiquinonate [Bac09]. 5’-diphosphates [PFR04b]. 5-hexadiene [PA05]. 5-hexadiyne-3-enes [PWFS01]. 5-hexatrienaldehyde [ZGZX07]. 5-nitro-3-carboxybenzisoxazole [UTH^{+}03]. 5’-phosphate [PMM05]. 5’-phosphate-dependent [LDY^{+}08]. 5-phospho- [RGP^{+}07]. 500-MHz [CMD^{+}04]. 53A5 [OVMV04]. 53A6 [CLWL09, OVMV04].

6 [BRV^{+}07, Wib04]. 6-311 [Wib04]. 6-31G [BRV^{+}07]. 6-31G* [FSFK05, RRP^{+}01, FKJ^{+}01, NL08]. 6-31G** [WD04]. 6-bisphosphatase [MRS^{+}07]. 67-residue [MH09]. 6G [VSW^{+}03].

7-species [WG02]. 790 [Ano06c]. 7a [HKHN08].

8-nitroguanine [JM07a]. 8-oxoguanine [FPN^{+}05, JM07a, Pin03].

9-heterofluorenes [CZFH07].

= [Bac09, CPJ00, CRC^{+}08, GHLK^{+}02, GPSP06, Han01, HT05, HYA02, HKHN08, HZ09, Hua09a, JJK^{+}00, KBL08, LS08a, LZZC09, LYZ^{+}08, Mar03, Mck07a, Mck07b, OS08, RB01, STC^{+}08, WLLS04, WZZ^{+}09, WD08, WWS07, XFF06, ZJM^{+}07, ZY01, ZXY03, ZL09b].

A*0201 [WCF04]. AA [KB02, KDSV02, POJ01, PB05, XLT07]. AA/L [KOML08]. Ab-initio [Haf08, HELM09]. abasic [FPN^{+}05]. ABEEM [YZ06]. ABEEM/MM [YZ06]. abilities [OYH09]. Ability [GM01, RRZA08]. ablation [KZW^{+}05]. ablation-mass [KZW^{+}05]. ABO [WD08]. ABSINTH [VP09]. Absolute [BWE05, ZMZ09, BMRF01, DHF^{+}05]. absorption [MWL^{+}08, MSH^{+}06a, YXZ^{+}04]. abstraction [AST06, CUS00, CUSS03, GAIMVB01, mJlZsLyL07, LW04a, LLL07, TGLL07, WLLS04, XLL^{+}02]. Ac [GHLK^{+}02]. accelerate [Gon07]. Accelerated [LSG06, LDG02, FSM09, Har04]. Accelerating [FEV^{+}09, LEV^{+}09, SPF^{+}07]. Acceleration [KG02]. accelerator [ATMK03]. acceptance [KBB09]. acceptor [BL06, RM07]. accessibility [AG03, ENM^{+}04, GP06, Tot04]. accessible [BHW00, BMLV04, GB04, HHS^{+}05, LFBSK07, RP07d, TSMNG01, TRS02, ZCL09]. accessible-surface-area [ZCL09]. Accompanying [Ish02]. account [May07, SN06, Vya01]. Accuracy [FII^{+}07, GG09, PSMB05, SKK^{+}07, UBDPJ04, Bie04b, FKFG08, JS07b, KC01b, MKGA06, MHW04, RK05, SM03, SW06]. Accurate [ABWT09, BdPRMAI00, EK06, GK09, Gri04, HdS06, HD06, HMSM06, Ish04, LLZL09, MSH^{+}06a, Tot04, WFHP01, WHP02, WHF08, WX09, ALKH04, Bie04a, BLB09, CGBF05, CF06, GKK07, HdMdS05, JJK^{+}00, Rud05a, Rud05b, Rud05c, Vas02, WCC08, ZFW08, vEMK01]. accurately [IGL07, SBI08]. acetaldehyde [Lu09, YLZ08, Lu09]. Acetalization [RUPH06]. acetals [MGG06]. acetamide [CCK01]. acetate [GWM08, PGG06]. acetic [MH08a, YT03]. acetonitrile [ELK^{+}09, GJK00, NL07]. acetylcholine [GCD^{+}08, MCK05]. acetylcholinesterase [MCK05]. acetylcholinesterase-catalyzed [MCK05]. acetylene [DLD^{+}02]. AcF [GHLK^{+}02]. acid [CJK^{+}02, CMLS05, CJDK09, CCK01, DP03, DP04, DLHC06, DHW^{+}07, FZL07, HFHL06, HLC09, IT03, IKYM09, JPF^{+}00, JKM08, JCL05, KLB03, LL07, MT03, MMLC05, MSF^{+}08, MH08a, Nak07, NHH05, NLL^{+}09, Pac06, RR05, RKH03, SKGS00, SYC03, SL04, SBG09b, SWR06, SHK^{+}05, UNHYT06, VM02, XSHC06, XLC08, XLT07, YT03, YXL^{+}09, ZZY07, ZZY08, ZOJ^{+}06, vDSSvA04]. acid-3 [vDSSvA04]. acid-catalyzed [RR05]. acid-nucleotide [MSF^{+}08]. acidity [ELK^{+}09]. acids [BE06, CADW03, CLA^{+}00, FM00, HWTL03, HP04, IKYM09, KS01a, LDY^{+}08, MB00, MM05, NHH05, OMNH08, PPYS08, Van02a, XKKL03, YLL^{+}09, ZLD09]. AcO [GHLK^{+}02]. across [HZX04, SSM08, SRB06]. act [GM01]. actinide [AB00, GHLK^{+}02, NSO^{+}07, VMA03]. Activation [EL06, BGC^{+}09, BZL05, CC09, CFER04, KT08, Lu09, PV07, RRS06, Vya01]. active [AG00, BSDM04, CFR06, CFS^{+}09, FCP^{+}04a, FCP^{+}05, HBM06, HFS^{+}07, HYR06, JHPRSM^{+}05, KSK00, KZRO03, LLL03, MDA08, PMM06, RZWS07, SS05, SFR07, TDH06, XLZ08]. active-site [SFR07]. activities [HMMS09, MS04, ZWB09]. activity [AGMPRG^{+}08, Bou01, CW02, DD08, DA01, DHW^{+}08, DHW^{+}09, FTLV01, GDPP08, LC09, MRS09, Sha02, WZY04, Zer08]. acylation [MCK05, MK02]. adaptable [KF08]. Adaptation [HLM05]. adapted [FCP^{+}04b, HdS06, HD06, LWX07, PTC01]. Adaptive [BHW00, HBW00, HBW01, HW03, HLSH05, DK01, GY08, OM04, RNG03, SJJ^{+}04, WCS09, LS08c]. addition [BLO^{+}02, DGD^{+}05, LL00, Mui05, RAGLL09b, RR05, WCW08, WSC09]. addition-elimination [Mui05]. additions [AVB00]. Additive [GGK^{+}08, CCK01, LKA01]. adenine [KKMMS04, SG07a]. adenine-thymine [KKMMS04]. adenosine [MRS^{+}07, YKK09]. ADF [tVBB^{+}01]. adiabatic [SLRC01, TVL^{+}03]. ADMA [EM03b]. admissible [WG02]. adrenergic [YKK09]. adsorbate [BWI^{+}02]. adsorbed [DR09, PBZ00, XPW09]. Adsorption [ATH^{+}03, BRS00, BRS01, HSF08, ZTP^{+}08, NK06, SURG06, ST04, WLX^{+}05, ZCS04, ZSC05]. Adun [JGVF05]. advanced [LAEL01]. advances [MMRVH07]. Advancing [PP08b]. affect [AST06]. affinities [AVS09, DJ04, KS05a, KKMMS04, LLXS02, MRS^{+}07, SRB06, WSM^{+}08, ZJM^{+}07, ZXL^{+}04, dSGCG00]. affinity [ABÅ04, FO08, GCD^{+}08, KFB05, KS08, Lee09, LXW^{+}09, MML^{+}06, RTG00, SOOF05, SWV^{+}05, ZWB09]. affording [OY01]. after [TJM^{+}03]. Ag [GPSP06, LYZ^{+}08, NA06, SG07a, WCS09]. against [SSS^{+}09]. AGBNP [GL04a]. AgBr [Sha02]. agent [LHJ^{+}06]. agent-based [LHJ^{+}06]. aggregate [KHF^{+}09]. aggregates [AB08]. aggregation [IM06, OGH05]. AgN [ZX08]. agonists [GCD^{+}08, SBG^{+}09a]. Agreement [LS05b]. aided [PJB^{+}07]. AIM [SFC04, CFS03, WW03]. AIM2000 [Ano01a, BKS02]. AIM2000-Program [BKS02]. Al [OS08, WZZ^{+}09, Van08, KKJH08, QB05, WZZ^{+}09]. al. [RKH03]. Ala [PC00]. alanine [DSR^{+}07, ECA06, GAIMVB01, GSB09, GKTS04, HHP04, HMK02, JW06, LSW^{+}01, MOP^{+}07, MFR07, PFJ^{+}03, Qua07, WD04, ZW09, ZM06]. albicans [RGP^{+}07]. Alchemical [Blo04]. alcohol [FBDG06, JJH01, KBN02]. alcohol/water [FBDG06]. Alcohols [ACLD03, LCDA03, LCGA03, LCA03]. aldehydes [LLA01d]. Alder [Hir08]. algebra [AT02]. algebraic [Tor02]. Algorithm [GGB07a, GGB07b, KSU03, MO01, WM12, ASWG07, AM06a, AM06b, AGSFAL05, BP02, BED02, BA08, BMTSC01, BSH07, BHH^{+}09, CS02, CLZX09, CWV^{+}05, CF04, DDKV07, DP03, DP04, DSR^{+}07, DBGV07, DAK08, FKFG08, FCP^{+}04b, GF08, GLD08, GCD04, GM04, Gon07, GKK07, HHJ03, HH04, HM06, HLM05, Ish04, IPN06, IPN07, JCA^{+}02, KH01, KM00, KMA^{+}07, KUB07, KvGH01, KH06, LJZ^{+}07, LJS05, LM03, LMO09, MP03a, MP03b, MVLG06, MM07, Nak02, OR05, PV03, Rap06, RHL09, SKSH07, SJJ^{+}04, SWR06, SSMW09, Ste04, SBH02, TYO^{+}02, TGD05, VGO^{+}07, WRBV03, WCS09, YK00, YGZZ05, ZBS03, ZZ08, ZA07, vLBBR12]. Algorithms [LMH^{+}09, BYQS03, BdPRMAI00, CKMC04, CSRST04, CHMI05, DB06, FS98, FS00a, HWDB03, KOFF09, OTL08, OGH05, RLER04a, SE08, TP01a, VW04, WK01, YL06, vLBBR12, HBW00, HBW01]. alignment [BA08, CLZ^{+}09, CLZX09, HHG^{+}09, RI07, VGDSU08, Leh06]. alignment-free [VGDSU08]. aliphatic [MSR04, SDvG01]. alkali [JHMB^{+}09, JHMB^{+}11, WWT08, ZWY^{+}09]. alkalimetal [BSG07]. alkaline [JHMB^{+}09, JHMB^{+}11]. alkaloid [BMRF01]. alkane [HGMB04]. alkanes [BL09, VVBV02]. alkyl [BE06, CC07, EB04, LLZL09]. alkyl-cyanobiphenyl [CC07]. alkylation [EL06, VBGL^{+}00]. alkynes [WCW08]. All-atom [FM00, MB00, VGO^{+}07, GB04, IT03, MT03, PHH^{+}08, RG08, WS07, JS07a]. All-electron [EL09, ITN^{+}05, IS07]. all-purpose [JGVF05]. all-siliceous [LST08, LTV08]. allene [WMRW^{+}01, ZKZ^{+}07]. allenes [WCHW09]. allocation [SKSH07]. allosteric [LGB^{+}09]. Allostery [Sen06]. alloys [GD09, KGD06]. allyl [ZGZX07]. AlNC [MLCD01]. alpha [GKK07]. Alpha7 [GCD^{+}08]. AlPO [PHH^{+}08]. altered [DLRZ09]. alternating [YFR05]. alternation [JPCA08]. aluminophosphate [LMV07]. aluminum [TBG00]. always [Kol04]. Alzheimer [MS03]. AM1 [DC02, FRS05, JBJB00, JJB02, LMMW04, RFSS06, TCT03, VGGMM05]. AM1-BCC [JBJB00, JJB02]. AMBER [Ano06c, WWC^{+}05, CCD^{+}05, FSFK05, JS07b, JM07b, MRC03, OYH05, TdMSD^{+}08, WCC08, WWC^{+}04, WZW^{+}06, WS07]. AMBER*C [CLA^{+}00]. AMBER95 [ONHN00]. ambiguous [BS01]. Amica [GBL^{+}05]. amidase [CBS^{+}03]. amide [CCK01, GSB09, LKA01, SJW09, YSA^{+}03, IINK09]. Amide- [IINK09]. amides [CMLS05, KS06]. amine [OY03, PS03]. amines [CLFA07]. amino [CADW03, CJDK09, CLA^{+}00, DLHC06, DHW^{+}07, FDSA00, HFHL06, HLC09, IT03, IKYM09, JPF^{+}00, KS01a, KLB03, LDY^{+}08, LL07, MT03, MM05, MSF^{+}08, NLL^{+}09, RKH03, SURG06, SKGS00, VM02, XSHC06, XLC08, XLT07, XKG^{+}05, YLL^{+}09, YXL^{+}09, ZLD09, ZOJ^{+}06]. amino-cyclopentene [SURG06]. aminoarabinonucleosides [BL08]. Aminopurine [SC01]. aminoribonucleosides [BL08]. aminosubstituted [TKS^{+}01]. ammonia [HT05]. ammonolysis [UNM^{+}01]. among [IGNH03, LLL07, WS07]. amorphous [CA04, CA07a, CA07b, SHH07]. AMP [FKM^{+}06, FKM^{+}07]. amplitude [KS05b]. Amyloid [BTP09, MS03]. amyloidogenic [CP09]. amylose [NK01]. anabolic [AGMPRG^{+}08]. analgesic [CMBC08]. analogs [AS09, BWE05, FPG^{+}06, FKJ^{+}01, HFHL06, LFEdL06, MT03, VM02]. analogue [LRI^{+}02, WC08]. analogues [BISB02, WHP02, XYN^{+}06, ZSE08]. analyses [BLT03, CZ05, DRAS05, HN02, Hir08, RP09]. Analysis [CGSdST06, EKO^{+}01, HG08, KS01a, LWK08, UTM^{+}02, UTT^{+}04, hYDN^{+}08, AJNG01, ALB09, AVS09, BM07, BLF02, BL00, BAH^{+}02, BPCD07, CG03, CS01, CCP04, CSRST04, CA07b, DDKV07, DRAS04, DSR^{+}07, ECM^{+}03, FK07a, FC06, FSS00, Gly06, GS07, Gra07, GHBB04, HHWG08, HS00, HLSH05, HP05, IN08, JPF^{+}00, JMD^{+}02, JFG04, KB02, KK08a, KMH02, KSU03, KN04, KSK00, KMA^{+}07, Kni00, LRI^{+}02, LS05a, MGCA07, MS03, MWE02, MHW04, MMP^{+}07, MA05, NK06, NSU^{+}02, OML^{+}00, ON07, PFC03, PGH^{+}04, PP08b, PYEA03, PAS07, PYCD03, PYS05, PC07, PLC08, RMHK03, RS07b, SH09, SMGE08, SSHT03, SFC04, SCF^{+}09, SvDS01, TYN05, TCR^{+}02, TT01, TD06, TTB01a, TTB01b, UTH^{+}03, VGB08, VKCK09, WRBV03, YNW05, YK08, ZSE08, Zer08, ZWB09, ZHH09, ZB07, NYK^{+}09, Ruv07, VB09, RS07a, VB07]. analytic [ASWG07, DOSG06, GL04a, IK00, KBT03, LFSB03a, LFSB03b, SJW09]. Analytical [HNWF07, HNWF12, PDC^{+}08, QCK01, QCK02, RLR^{+}04, WL02, DSR^{+}07, HC08, HHS^{+}05]. analyze [AGMPRG^{+}08, Ham07, MCF07]. analyzing [DW08, LD05b]. andradite [ZWTP^{+}08]. anesthetics [TZX01b, TZX01a]. AnF [Han01]. Anglada [Bof01, Qua01]. angle [CIB05, FWH^{+}07, LI07, OFIK09, YL06]. angles [FPG^{+}06, FKZ09, HK08c, WHRG08]. angular [GY08]. anharmonic [BP07, DB07, GBDP05, LMO09]. anilido [GTC06]. anilido-imine [GTC06]. Anion [JT06, CW02, DMN05, EFQD09, GS04, wQZsLyZ02, QZZZ03, VDM06, WLZ^{+}07]. anion- [EFQD09]. anionic [ALC08, HHWG08, PGG06]. anions [CTFC08, KKMMS04, Lee09, LLXS02, Owe05, ZXL^{+}04]. anisotropic [BCIB05]. anisotropy [GKTS04]. ANN [HSMT04]. annealing [ADM^{+}06, CS02, CCP04, HPP00, KH01, LCKL05, LJKL08, MCF05, WG02]. annexin [MHJS06]. Announcement [Nor04]. Announcements [Ano05a]. annulation [GLH^{+}08]. anomeric [CPJ00, CPJ01, CKF01, LCDA03, VM07]. ant [CLZ^{+}09, CLZX09, DDKV07]. antara [LFS^{+}07]. antara-antara [LFS^{+}07]. anthracene [CG08, CDPL09, HIA03]. anthracene-9 [CG08]. antibacterial [YCW^{+}09]. antibody [CBS^{+}03, SOOF05, TH02, UTH^{+}03]. antibody-catalyzed [TH02]. anticancer [BZL05, PFR04a, SMM^{+}08]. antifungals [GDPP08]. antigenic [WCF04]. antiinflammatory [CMBC08]. antilipid [MRS09]. antimicrobials [GDPP08]. antitumor [KC01a, WM01]. apatite [RD06]. Applicability [DC02, PRS04, QTdG^{+}08]. applicable [PB05]. Application [ASDP^{+}06, CIB05, DSR^{+}07, GWS^{+}02, HKMS01, HSMT04, HM08, HMMS09, KSS08, KAK^{+}09, KBK^{+}01, Mat03, MW09, TBSM09, TKN^{+}08, BLN01, COS01, CSJ01, DLWV07, DV02, FOK^{+}04, FKU^{+}05, GLD08, GWM^{+}00, HHBH00, JW06, JGVF05, KTM02, KK01a, LL00, LFSB03a, LFSB03b, LI07, LSHR04, LCSZ09, LXL07, LLL03, LMO09, MG00, MS04, PB04, RI07, Rud05a, SG07b, Tot04, VM00, Wil01b, WR07, ZLY07, ZNLL07, dGWH01, CMJ08, CWV^{+}05, CMBC08, GRO^{+}03, GBDP05, GS04, IN08, IS03, LJZ^{+}07, LSY02, MB00, MCF05, MBC08, MTB09, ON07, PHJ^{+}08, Qua04, QCK01, QCK02, Rud05b, Rud05c, SL09, SDCG02, SFC04, TH02, VVS07, VKCK09, WRP^{+}06, WFR08, vGGB00]. Applications [HLLN06, BWZ08, BBG^{+}04, CGMPT^{+}08, CR09b, Est07, LW06, PMC^{+}08, PTC01, SPF^{+}07, Tor02, ECM^{+}03, GGB07b, HT03, RSN^{+}02, WNH03]. applied [ADM^{+}06, BL09, BT00, BS06, DMN03, FII^{+}07, GHLK^{+}02, MKGA06, OM04, TLKT00, YAÇ^{+}02, ZZ08, dVB01, BBC^{+}05]. Applying [AGMPRG^{+}08, Woo01, You11, GS02]. appreciation [vRS01]. approach [AB00, AGK03, AVS09, BG03, BME05, BS03, BGJ01a, BLF02, BMTSC01, BPCD07, CCWH02, CYM02, CGBF05, CAGR08, CMCB08, CG05, EM03b, GAIMVB01, GAI06, GJL^{+}08, GRO^{+}03, GDV03, GS02, GDPP08, GdSuM^{+}07, GdAcV^{+}07, GCD^{+}08, HMMS09, Hin00, HRR05, HT03, HLSH05, IT03, KKG^{+}09, KN04, KHY00, KHF^{+}09, KBT03, LFKL00, LCC09, LMJ02, LYS08, LD05b, LWZ09, MAF^{+}07, MR04, MFR07, Mor02, NYTH09, Oos09, PS09a, SAM06, SPGS08, SGPS09, SL04, SWZS04, VW03, Van02b, WWL^{+}09, XWC09, XKKL03, YS00, YK08, ZCS04, ZS04]. Approached [LL07, XSHC06]. approaches [BP07, Con02, CSD05, COL^{+}06, MLJ03, PSF^{+}08, PMM05, RLDI09, RSS09, SM08a, YCW^{+}09]. appropriate [Bac07]. approximate [Cul08, GB04, Hol05, KS02a, SZT08, SYC03]. approximated [PSF^{+}08]. Approximately [EA06]. approximating [MR04]. approximation [AB09, BRS00, BRS01, CLP^{+}05, CCK01, Der09, EA08, GMA04, GWS^{+}02, ION07, Kri09a, Lai07, LFSB03a, LFSB03b, LN01, MTE04, Nee03, OCB02, RP07d, SHSF05, ZFL^{+}05]. approximations [Dya02]. APS [CBC^{+}08]. APX [ZJM^{+}07]. aqua [RMP01]. aqua- [RMP01]. aqueous [BISB02, CPJ00, CPJ01, CW02, CCK01, DA01, EK06, FHRR07, HMWC03, HRR05, HDO^{+}02, IV04, IvSV06, JM07b, KEH^{+}02, KPR04, Kri08, Kri09b, LRI^{+}02, LR03b, Loe03, LMIF06, MM02, NL07, PK04, PHRR08, SH09, SMKM00, SBB02, VP09]. arabinonate [RGP^{+}07]. arabinonohydroxamate [RGP^{+}07]. arbitrary [KH06, LMV07]. architecture [TDK07]. architectures [TYO^{+}02]. area [GCD^{+}08, GB04, HHS^{+}05, Lab08, LFBSK07, RP07d, VP02, ZCL09]. areas [BHH^{+}09, TRS02]. arene [FKŠ^{+}09, PCMG09, RRZA08]. arene-containing [RRZA08]. arginine [ČJPZS08, SMGE08]. arginine-bound [ČJPZS08]. argon [BWW^{+}08]. argument [Ish04]. ARIs [PS09a]. arising [CCSJ00]. armed [KLM^{+}09]. ArOCS [ZGXX06]. Aromatic [CPML08b, PCO^{+}07b, Van08, Bor03, FVB08, HLC09, MM05, MGMM07a, ST01, SMV^{+}09, TDK07, VS08, WFHP01, XLT07, PCO^{+}07a]. Aromatic-Backbone [CPML08b, Van08, PCO^{+}07b, PCO^{+}07a]. aromatic-type [HLC09]. Aromaticity [BPCD07, FMPS08, JHMB^{+}09, JHMB^{+}11, LWW^{+}06, LTF^{+}07, MGMM07b]. Array [FJP07, ABF^{+}03]. arsenic [ALC08, KS05a, ZXL^{+}04]. ARTE [VB07]. ARTE-QSAR [VB07]. artifacts [CCSJ00]. Artificial [PS09a, RWBH09, dVB01, CLC03, Gol09, NINAT^{+}07, TCSM03]. arylamide [VIP^{+}06]. ascorbate [HBM06]. ASIC [NYTH09]. aspartic [ZZY08]. Aspects [HHBH00, MO01, BMRDB01, BRS07, Sie01, TT02]. assemblies [DFGB09]. assembly [DPRR05]. Assessing [IB04, FGR07]. Assessment [BP03, CCWH02, DGI^{+}08, KS08, LWH06, SSS^{+}09, WSM^{+}08, CKMC04, FMPS08, GT03, LLS03, SP05, GGT08, GHBB04, TFZRG01]. assignment [BB05, BMRF01, PRJ02]. assignments [PF06]. assisted [BA04b, KT08, WJX^{+}08]. assists [BM07]. associated [SWR06, TT08]. Associative [ABYM08, NL08]. asymmetric [WR07, WFR08]. Asynchronous [GLP08]. atmospheric [GCCVB00, PGNG03]. Atom [RP07a, RM00, BPC01, BR04, BWW^{+}08, CCK01, FM00, GWS^{+}02, GB04, HLLS05, IT03, JS07a, mJlZyL^{+}08, LMK01, MT03, MB00, PHH^{+}08, RG08, RS08, SSB^{+}03, SBLK01, SLL^{+}04a, TGLL07, VK06, VGO^{+}07, WLL07a, WBSR03, WS07, WLL^{+}03, YLWL09]. atom-bond [VK06]. atom-centered [SSB^{+}03]. Atomic [DVP^{+}02, FDM00, AS00, Bac07, BSÇ^{+}01, BCNs07, BSP06b, BK00, BLT03, BAÅ07, CN03, FS04, GC02, Ish03, JBJB00, JJB02, KRM^{+}02, Kau07, KS01a, KC01b, Lab08, LMV07, LST08, LTV08, Nil09, PGG06, PSMB05, RKA^{+}09, SvDS01, TBSM09, UBDPJ04, VZVG06, VC04, XLT07, YOB^{+}08, dSGCG00, DVRP^{+}03]. atomic-centered [TBSM09]. atomistic [CA04, IDMC09, KK01b, RPMP03, SPGS08, ZALMG03]. Atoms [VM07, YK08, ALTB06, AD00, ASS^{+}02, BHTCG07, BKS02, BS03, CMJ08, CDS09, EdlVR^{+}03, FS04, HSF08, mJlZsLyL07, KGN07, KS02b, LDC^{+}07, Mat03, NSO^{+}07, RP07b, RRP^{+}01, RLR^{+}04, RLER04b, SO09, SNM^{+}06, SFC04, Wil01b, WDX^{+}02, XLL^{+}02, PFB05]. Atoms-in-molecules [YK08, RLR^{+}04]. ATP [FCP^{+}04a, GS04]. ATP-binding [GS04]. ATP-dependent [FCP^{+}04a]. ATPase [HLB09]. attachment [LBG08, XWXC08]. attack [CBS^{+}03]. Attaining [Rud05a, Rud05b, Rud05c]. attending [HT05]. aug [Wib04]. aug-cc-pVDZ [Wib04]. Auger [OKE^{+}02]. augmented [JČHS07, KDG^{+}09, LFK05, MOP^{+}07]. autoantigen [KVS^{+}06]. AutoDock4 [MHL^{+}09]. AutoDockTools4 [MHL^{+}09]. Automated [CKMC04, LMO09, HR08, LR03a, MM03, VSW^{+}03, MHL^{+}09]. Automatic [CHMI05, WK01, AGI^{+}00, AGI^{+}07]. automaton [XWC09]. auxiliary [GKH05, JSHG07]. available [SCF^{+}09]. average [TRS02]. averaged [CP08, PYCD03, PYS05, PC07, PLC08, SMAdV00]. averages [Rap06]. averaging [BSÇ^{+}01]. avian [DLRZ09]. avoidance [WCFH02]. axial [BMRF01, CN05]. axis [OMNH08]. azaglycine [LKJ^{+}04]. azide [MSR04]. azides [ZX08]. azido [PFR04b]. azole [SMM^{+}08]. azole-bridged [SMM^{+}08]. azurin [PMGL03].

B [OS08, Sta00, WZZ^{+}09, WD08, ZZZ^{+}06, ALKH04, BAL^{+}01, FH01, GL04b, JS07a, KVS^{+}06, LMGR06, Maz01, Pin03, VDM06, ZWB09]. B-DNA [Maz01, Pin03]. B-domain [JS07a]. B-spline [ALKH04]. B-splines [GL04b]. B3LYP [CLP^{+}05, FSFK05, HWGB01, NL08, TCT03, WC04, WX09]. B3LYP/6 [FSFK05, NL08]. B3LYP/6-31G* [NL08]. B3LYP/6-31G*/AMBER [FSFK05]. Ba [WD08, XB08]. Bacitracin [Dra00]. back [BB05]. Backbone [CPML08b, GKK07, Van08, Adc04, Ano06c, AHGK09, CLWL09, HSWN01, KLS02, LKA01, MFB04, MLL08a, PCO^{+}07b, SP05, WZW^{+}06, YL06, PCO^{+}07a]. backward [KM07]. bacterial [Ano06b, CPM03, GS04]. bacteriochlorophylls [LKT04]. bacteriopheophytin [IN01]. bacteriorhodopsin [RG02]. Bacteroides [SDM02]. Bader [GHBB04, SKSH07]. Baker [WB05]. balance [Ano06c, WZW^{+}06]. balanced [PB05]. Balancing [CF06]. band [AJ03, JCA^{+}02, ZZW09]. Baoshan [JW12]. bare [KT08]. barrier [CRGN07, KSTC01, LSG06, MG06]. barriers [DBM03, EL06, HFHL06, PBF09]. bars [MDI04]. base [CCK01, DP04, FZL07, HWTL03, KKMMS04, MMLC05, MSF^{+}08, MHS05, NL08, OY01, PG04, PSS^{+}04, PSMB05, SKGS00, SG07a, ŠBL05, SC01, SYC03]. base-catalyzed [OY01]. base-pairing [DP04]. based [Adc04, ALTB06, ALB09, AB09, BdPRMAI00, BMRF01, BDW00, BMTSC01, CGMPT^{+}08, CKR08, CLZ^{+}09, CLZX09, CFK08, CHA^{+}07, CPUGD09, CMBC08, CRGN07, CA04, CSB^{+}03, DLW06, DMLI05, DHW^{+}00, DB02, DDVD09, DHW^{+}07, DHW^{+}09, DWC^{+}03, FCK^{+}08, FCP^{+}04a, FCP^{+}05, FM00, FZL^{+}06, FRLN09, GZL02, GRCD01, Gon07, GDPCPU07, GDPP08, Gra07, GAS04, HSMT04, HS07b, HLM05, HZ06a, HZ06b, HMOG07, IIK09, IR03, JD09, JKII08, JGVF05, KLS02, KBA^{+}04, KK08c, KBK^{+}01, Kob03, KIM^{+}09, KZW^{+}05, KVS^{+}06, LFKL00, LHJ^{+}06, Leh06, LXZ06, LJZ^{+}07, LH05, LLM09, LM03, MLL06, MSF^{+}08, MSH^{+}06a, MBH^{+}02, NLL^{+}09, NMAT01, OVMV04, PS09a, PFR04b, PA05, PAS07, PRJ02, PF06, PRDS08, Pul05, QLHL09, Rao00b, RSE07, RSER09, RLDI09, RSN^{+}02, RKA^{+}09, RUPH06, RRS09, RSS09, Ruv07, SAM06, SKSH07, SGPS09, SPL^{+}02, SBB02, SE08]. based [SZW^{+}05, TTBM09, Tot04, VSK^{+}04, VB09, VGDSU08, WL09a, WL00, WS07, WRBV03, XYN^{+}06, XL02, YWHZ03, YNW05, hYDN^{+}08, YJF06, YXL^{+}09, YKK09, ZCL09, ZLY07, ZLD09, ZWB09, dSR08]. basepair [BCP03]. bases [CCK01, Nak07, RTG00, RKH03, SL04, WRP^{+}06]. basic [Rud05a]. basicity [EK06, Lee09]. basins [CFS03, MP03a]. Basis [AHK02, BRLS08, BRLS12, JJK^{+}00, Wib04, ABF^{+}03, ALKH04, Bac07, BY06, BR04, BT00, BSOB05, BRV^{+}07, CMJ08, CRS05, Cul04, CGSdST06, DMZT08, EA08, EdlVR^{+}03, EL09, FZL07, GKH05, HdMdS05, HdS06, HD06, IO08, JSHG07, KK08a, LFK05, Lai07, LMV07, LST08, LTV08, MV06, Mas04, MLL^{+}08b, MC06, MY08b, MY08a, NSO^{+}07, OBBS05, PSC^{+}01, Pen06, PFJ^{+}03, PSMB05, RRP^{+}01, RLRE01, RLER07, SSB^{+}03, SNM^{+}06, TSSSG08, VKP^{+}08, VB03, Var09, VKCK09, WMGK07, WTKM06, Wei08, ZWPR^{+}04]. Basis-set [AHK02, MV06, Pen06, VKCK09]. BCC [JBJB00, JJB02]. be [HdMdS05, HdS06, HD06, IGL07, STSF02, WCF04, BP01, LWW^{+}06]. beam [BAL^{+}01]. bearing [NL08]. Becke [AAP00]. Becke-3 [AAP00]. Becke3 [PDS01]. Becke3-LYP [PDS01]. BeH [PRSMV08]. behave [PB02]. behavior [Ama02b, Bac05, BISB02, LB05, OO04, RP07c, SBI08]. behaviors [LML^{+}00]. being [OCB02]. benchmark [Ano01b, BSB05, DGD^{+}05]. Benchmarking [Hol05, SZT08, WS07]. Bennett [KBB09]. Benzdiynes [ASY01]. benzene [BE09, BRLS08, BRLS12, HT05, HRG07, IINK09, LWX07, Sch00, SG07b, ZTP^{+}08]. benzenes [PB05, WRP^{+}06]. benzo [GLRL02]. benzocryptand [WWT08]. benzodiazepinone [SPGS08]. benzodioxoles [MRS09]. benzoic [BE06]. benzonitrile [ZH08]. benzylideneaniline [BY06]. benzylpenicillin [DSS03]. Beowulf [BMRDB01]. Bergman [PWFS01]. Besalú [Bof01, Qua01]. Bessel [DBS08]. Beta [LHI09, BTP09]. Beta-hairpin [LHI09]. between [AD00, AZM03, BS03, CFR06, DRAS05, EFQD09, FG03, FL07, FO08, FKM^{+}06, FKM^{+}07, GWM08, HPP00, HRBKB03, HFHL06, HN02, Hir08, IINK09, JPCA08, KWK^{+}01, KWK^{+}02, LDC^{+}07, Li01, LL01, LFZS04, LLL03, LS05b, MST^{+}08, MBH^{+}02, OY01, PSF^{+}08, PMPGP05, PS03, RLRE01, SM08a, SBLK01, Sim07, SWM04, SKK^{+}07, SP05, TYN05, TK08, TDH06, UTM^{+}02, UTT^{+}04, WLX^{+}05, YT03, YQQH09, Yos02, ZZTS09]. beyond [CLP^{+}05, CCK01, Haf08, PP08b]. BH [QZL^{+}04, SAS05]. Bi [LS08a, WL09b, HZ09]. bi-transition [WL09b]. bias [OM04, SY09]. Biased [MLG04, KV00]. bicyclic [EBDPM00]. bicyclo [BE07, BPC01]. bifurcation [CPFL02]. bilayer [CEP07, HNL08, MCR08]. bilayers [JM07b, RG08]. bimetallic [WCS09]. bimolecular [ML00]. binary [Kle02, Kle03, LCSZ09]. Binding [ABÅ04, AGO^{+}02, BCP03, RGP^{+}07, ABYM08, AM06b, APG05, AVS09, BWE05, BSP06b, DLRZ09, Dra00, ECM^{+}03, FKU^{+}05, GCD^{+}08, GS04, HT05, HNWF07, HNWF12, IO08, JMD^{+}02, JZD^{+}09, KFB05, KS08, LXW^{+}09, MK02, MHJS06, MLL^{+}08b, MRS^{+}07, NyHN06, NHN06, OYH09, OFIK09, PMGL03, RSP03, RGG08, RK05, Ruv07, SOOF05, STSF02, SWV^{+}05, TGGP^{+}00, VGGMM05, WM04, WHF08, Wou00, XL02, ZGFL01, ZWB09, KEB04]. binodal [MM07]. binuclear [GS04, PLC08]. bio [KH01]. bio-molecules [KH01]. bioactive [BLB09, SD09]. bioactivity [LJZ^{+}07, SJJ^{+}04]. bioinorganic [MSH^{+}06b, SGD06]. bioisosterism [DPM09]. biological [CCK01, CMGDAC^{+}07, GdAcV^{+}07, HMMS09, LDTS07, Mac04, TH02, WCK00, YPNE09, vdVGDM00]. biologically [CSU05, LLL03, RZWS07]. biomarkers [VGDSU08]. biomembrane [WEE01]. biomimetic [FO08]. biomolecular [BHW00, BBM^{+}09, CCD^{+}05, CHB^{+}05, CvG08, FWH^{+}07, JTR05, KAK^{+}09, KYT^{+}08, LSO04, OVMV04, WB04a, WB04b, WB05, WL09a, ZFW08]. biomolecule [ABWT09]. biomolecules [ECM^{+}03, Est07, FEVM01, HMD06, KHY00, MMLC05, QSS01, YNZ^{+}08, YJF06]. bionanosystem [MO09]. biophysical [Mat03]. BiOX [HZ09, Hua09a]. biphenyl [PCMG09]. biradicals [KC01a]. bis [BLN01, CDL06, PYS05]. bis-heteropentalenes [CDL06]. bishomoaromaticity [HWGB01]. bisphosphatase [MRS^{+}07]. bispidine [ACM^{+}06]. bits [PM02]. black [MBP09]. Blind [GZM09]. block [ATM^{+}07, ASS^{+}02]. blocked [RRS09]. blockers [HSMT04]. blocks [SSB^{+}03]. blood [CRGN07, HMSM06]. Blue [CPDZH08, HRG07, CR02, McD08, SRK^{+}00]. blue-shifted [McD08]. Blue-shifting [HRG07]. BLYP [TCT03]. board [ATMK03, KAK^{+}09]. boat [RP09]. bodies [FS98, FS00a]. body [CCK01, FII^{+}07, FBDG06, Ike04, Loe03, SM03, TKH07, LR03b]. Bofill [Qua01]. Boltzmann [WB04a, WB05, ABWT09, BHW00, BBP09, BH03, BF04, BF07, GPN01, GCD^{+}08, GGT08, Höf05, HBW00, HBW01, KWHH07, LDG02, NYTH09, PZS04, SATO04, VZM^{+}08, WB04b]. Bond [CGMPT^{+}08, CRC^{+}08, JG03, MGCA07, May07, SH08, WM12, Bic09, BL06, CMLS05, CPFL02, CPDZH08, CJW^{+}09, Cul08, DR09, DGD^{+}05, DMZT08, FH01, FO08, GYMN07, GR07, GS07, HRG07, HS07a, Hir08, JPCA08, JP09, Kle03, KBLP09, LC07, LZZC09, LS08c, LS05b, MG00, OO04, Pac06, PSC^{+}01, PAS07, PYS05, PV07, Rao00a, RM07, RCJ02a, RD00, SEKS09, Sha07, Sim07, SPT^{+}03, SWZS04, SMZW05, SSMW09, ST01, SSW^{+}07, TJM^{+}03, Tru07, VK06, VBGL^{+}00, WHRG08, WJ00, XWXC08, vLBBR12]. Bond-based [CGMPT^{+}08]. bond-order [LS08c]. bonded [CPDZH08, Gon07, HT03, IO08, LB05, LDL^{+}09, LZF^{+}09, McD08, MH08a, NBTN04a, NBTN04b, NL08, PHFC04, ZH08, vEMK01, vE01]. bonding [AM07, AG00, Bac04, Bac05, Bac07, BHTCG07, BM07, BSG07, CWWS07, CQ04, CCK01, EFQD09, FLK^{+}07, FK07b, Jac09, Kau07, KJP^{+}07, KBL08, Kle02, Kle03, KGD06, LW07, LWK08, LDL^{+}09, PG01, PYCD03, PLC08, RPNJ07, RP04, RS07a, RS07b, SM08a, SG07a, SCP08, Wil01a, WD08, WWS07, XZ04, XK08, Yos02, ZX04, ZW09, ZB07]. bonds [Bac05, BUMCMRL00, BRS07, CRC^{+}08, DR07, HA04, Mit01, NHH05, OO08, PGG06, PC05, PC07, SO09, SGD06, SJW09, YT04]. Book [Bic09, Lip00, Sta00, Woo01]. borane [ZZZ^{+}06]. borate [HT05]. Born [LFSB03a, BC06, CF06, DLG00, FOL^{+}04, FC06, GZL02, ILB03, Lab08, LFSB03b, MTE04, MCM04, OCB02, Tot04, XL02, YJF06, ZGFL01, ZWZ09]. Born/volume [Lab08]. borohydride [QZL^{+}04]. Boron [JBGK08, LMGR05, LMGR06, LX07, LWLS07, SRS07, ZB07]. Boron-doped [JBGK08, LWLS07]. boronyls [LMGR05]. BOSS [JTR05]. both [HdMdS05, HdS06, HD06]. bound [ČJPZS08, WC09]. Boundary [BH03, ABWT09, Ara04, BVW04, BF04, BF07, HH04, KWHH07, QSS01, TK08, WM06]. bovine [MBC08]. bowls [LMGR06]. box [LM03, WM06]. box-counting-based [LM03]. boxes [MBP09]. Boyd [Sta00]. Br [FHF^{+}01, HZ09, Hua09a, KBL08, Mar03, RB01, STC^{+}08, WLLS04, ZJM^{+}07, ZY01, ZL09b, HYA02, LDC^{+}07, RFSS06, SLL^{+}04b, WLLS05, ZZW^{+}07]. brain [CRGN07, HMSM06]. BrCl [WLLS05]. breakage [SWR06]. breakage/closure [SWR06]. breaking [HRR05, Pac06, VHRR07b]. breast [VGDSU08]. Breit [Ish03]. bridge [CFC^{+}08, PAS07]. bridged [LHP01, SMM^{+}08]. bridges [KF03]. broadly [PB05]. broken [ATH^{+}03, BB08]. bromine [WyLG^{+}09]. bromine-substituted [WyLG^{+}09]. BSSE [GAI06, PSC^{+}01, SPDS01, SAM06]. BSSE-corrected [PSC^{+}01]. BSSE-free [SAM06]. Buckminsterfullerene [NRKH02]. buckycatcher [Won09]. Buff [KS06]. build [BMTSC01]. build-up [BMTSC01]. Building [BCIB05, BAÅ07, GKK07, HP05, MABM09, PC00, PFC03, Fau01]. built [GFS05]. bulk [BACJCT01, BGC^{+}09, EBL^{+}08, JBGK08, LZZC09, LLL03, PB04]. bulk-doping [JBGK08]. butadiene [GRO^{+}03, Hir08, WR07]. butanal [NSB08]. butanes [WW04]. butyl [Bac09].

C [Ano00, BAL^{+}01, BPC01, BSB05, Bic09, CTFC08, HK07, mJlZyL^{+}08, KYFW07, KJP^{+}07, KSN01, LYZ^{+}08, Mck07a, Mck07b, Owe05, SLL^{+}04b, VBS09, WLL01, WDXS06, WZZ^{+}09, XFF06, YHD^{+}06, ATBLS04, Ber03, CPDZH08, CCB04, CS01, CRSB03, CTFC08, DGD^{+}05, DRAS04, Ell07, FHF^{+}01, GYCZ04, GZ07, GB02, HBM06, HYA02, HK07, HA04, IN08, LDMR01, LMK01, Mit01, OO04, PRSMM03, PV07, RD00, RFSS06, SLL^{+}04b, STC^{+}08, TYN05, WLPF05, WDWS06, Wil01b, XDS06a, XWXC08, ZZL04, ZW09, ZLD09, ZZvRSC08, dRLMS00]. C-PCM [CRSB03]. C12A [BRDC02]. C12A-p8 [BRDC02]. C96 [ONHN00]. Ca [WZZ^{+}09, WD08, XWC09, HLB09, PNG08]. cabonyl [RUPH06]. CaCO [SCP08]. cage [CS01, KFD06, WLPF05, WL09b]. cages [CJS^{+}03, Wan09]. calcium [HSMT04, HLB09, LGB^{+}09, MHJS06]. calcium-induced [LGB^{+}09]. calculate [BACJCT01, CSD04, IS07, Kar01, Kne05, KBLP09, OV03, RSN^{+}02, SFRS01, WW03, YS00]. calculated [BE06, BE07, GG09, Gra07, LMV07, RSSKB03, RM00, Wib04, WM04, ZXY08, Kle03]. calculates [ATMK03]. Calculating [Chu07, CG05, DRMD03, DF04, LN01, MC06, PDC^{+}08, PMM05, RSE07, SYC03, WCK00]. Calculation [BK00, CPML08a, DJ04, KKY01, KRM^{+}02, LSW^{+}01, MT03, MO01, MRS^{+}07, TS05, VM02, VC04, WKYU01, ZWPR^{+}04, BP02, BSÇ^{+}01, BH03, ECA06, FOL^{+}04, FROD08, GKRG08, GLMV09, GAdGM08, GGT08, HTKG08, II02, ITN^{+}05, JJK^{+}00, KFNH08, LFK05, Li01, LLZL09, MGLO03, NRKH02, Nil09, OCP02, PZWG^{+}04, PRSMV08, RNG03, Rap06, RLER04a, RKA^{+}09, RRFC^{+}03, SOOF05, SKGS00, STSF02, SSM08, SHH07, TZX01b, TZX01a, TLOG00, TP01b, TRS02, TKN^{+}08, UKN04, UIHN09, WLLS05, WLL^{+}07b, WSM^{+}08, WM01, XOW^{+}00, ZWB09, ZL09b]. Calculations [YH07, Ano01b, ALKH04, Bac09, BP07, BSOB05, BMRDB01, Blo04, BMB07, BMRF01, BWI^{+}02, CLP09, CMJ08, CN03, CRS05, Chi03, CS03, CMA^{+}08, DB07, DPDG05, DSS03, DMJV05, DWC^{+}03, EKO^{+}01, EBL^{+}08, EL09, FKFG08, FL08, FMSA06, FR06, FO04, GJL^{+}08, GBDP05, GM04, GWM^{+}00, GBB07, GPSP06, HHBH00, HLLN06, HYA02, HMSM06, HFHL06, HKHN08, HT03, HWGB01, HZ09, Hua09a, Hua09b, ILKR09, IGL07, IPN06, IPN07, JCA^{+}02, JBGK08, JČHS07, KGL07, KWHH07, KRLD09, KTM02, Kob03, KBLP09, KS02b, KKS04, LLA01a, LLA01b, LLA01c, LLA01d, LLA03, LFSB03b, LDC^{+}07, LHI09, LWH06, LC06, LKW04, LZF^{+}09, LS05b, LDG02, LMO09, MMLC05, Maz08, MBL^{+}00, MA05, MLCD01, MLJ03, MS01, Mui05, NYK^{+}09, NTH09, NSU^{+}02, OBBS05, OYH05, OS06, OKH^{+}02, Oos09, OKE^{+}02, PB06, PMGL03, PVdJB00, PV03]. calculations [PBZ00, PRSMM02, PSS^{+}04, PSMB05, RSP03, RMP01, RP02, RRCA08, ROG00, RWBH09, RS05, RRS07, RJLR06, SBI08, SSB^{+}03, SCS07, SBG^{+}09a, SHM04, Sha02, SSL02, SN06, SFRS01, SMG09, SG01, TK08, TY03, UKNS01, UTM^{+}02, VKP^{+}08, VZVG06, Van02b, VE09, VIP^{+}06, WL02, WTKM06, Whe08, WC04, WHF08, WR07, WFR08, WHG^{+}07, WJ00, XLZ08, YTH01, YK08, YSA^{+}03, ZSE08, ZXYF09, ZM03, vGGB00, vdVGDM00, LFSB03a]. Calibration [OKH^{+}02, LLZL09]. calix [RRZA08]. camphor [AST06]. Can [DSB^{+}02, DDBP09, LC07, MSH^{+}06b, PB06, SBI08, STSF02, WS07, IGL07, Mck07a, WCF04]. cancer [VGDSU08]. Candida [RGP^{+}07]. candidate [SF07]. canonical [BP02, EMP07, ITN^{+}05, IS07, KM00, Kni00]. CAOs [PS03]. capabilities [GCD04]. capped [CZ05]. caps [ZCZ03]. capsid [KCL06]. captopril [AGO^{+}02, APG05]. capture [YXC^{+}07]. Car [JP09, Sch04]. carbapenem [BBSS06]. carbazole [YFR05]. carbenes [HA04]. Carbohydrate [KBN02]. carbohydrates [ACLD03, HR08, KDSV02, LR03a, LCDA03, LCGA03, LCA03, LH05, MW00, Sto05, KYT^{+}08]. Carbon [KK08c, LMGR05, BSB05, BG07, CZ05, CDPL09, DWS^{+}09, GKK07, HT05, KT08, KLS02, KK01b, LMK01, MMRVH07, PAS07, wQZsLyZ02, SRS07, Wan09, WSC09, XWL^{+}09, ZKZ^{+}07, ZWY^{+}09, ZZvRSC08]. carbon- [ZWY^{+}09]. carbon-centered [WSC09]. carbon-rich [CZ05]. carbonate [vDSSvA04]. Carbonyl [RD00, DLR^{+}08, LL00, LLA01a, LLA01b, LLA01c, LLA01d, LLA03]. carbonyls [BRV^{+}07, LMGR05, PLC08]. carboplatin [WM01]. carboranes [JRJ01, OSA06]. carboxybenzisoxazole [UTH^{+}03]. carboxylates [ČJPZS08]. CarC [BBSS06]. carcinogenic [EL06]. carcinogenicity [VS08]. Carlo [AGSFAL05, AGSFA^{+}05, BR03, BHG03, Der00, FCK^{+}08, FKFG08, GHH07, HMD06, IM06, IKYM09, KLS02, KM07, KKC05, LML^{+}00, LZA02, LRWG03, MH09, Nak02, NA06, NCO^{+}05, OM04, SKGS00, SCS07, SBJ08, SM08b, SWR06, TS05, XKG^{+}05, ZCS04]. Carlo-with-Minimization [NCO^{+}05]. carma [Gly06]. carrying [Tor02]. Cartesian [LPK07, PHR^{+}05]. CAS [PRSMM03, BMB07, JHPRSM^{+}05, PRSMM02]. CAS-SCF [JHPRSM^{+}05]. CAS-SDCI [BMB07, PRSMM02]. case [AB00, AS00, BUMCMRL00, BWW^{+}08, BS06, CFS^{+}08, ČJPZS08, MV06, MDI04, Pac06, PGG06, PC00, Pog03, RUPH06, SBI08, SN06, Van02a, WRBV03]. cases [BCF^{+}09]. CASPT [PO03]. CASPT2 [KRLD09, WLZ^{+}07]. CASSCF [Bac04, Bac07, DOSG06, WLZ^{+}07, dSVA^{+}09]. CASSCF/CASPT2 [WLZ^{+}07]. Catalysis [UTH^{+}03, DWS^{+}09, QZZZ03, TH02, VBGL^{+}00, ZDS^{+}05]. catalyst [VBGL^{+}00]. catalysts [YXC^{+}07]. catalytic [CGB^{+}09, KSK00, LS08b, MS04, NTH00, RWBH09, SPT^{+}03, WC08, ZAT07]. catalyzed [AST06, BTP09, GVATG03, GLH^{+}08, HSWW00, MCK05, OY01, PHKG07, RR05, Sie01, TH02, TGLL07, WCW08, WCHW09, ZWS^{+}09]. catastrophe [PA05]. cathepsin [ZWB09]. cation [DSB^{+}02, Don08, LB08, OO04, PV07, QZL^{+}04, SLRC01, VL00, WLZ^{+}07, WSM^{+}08, Wou00, ZL05]. cation-water [DSB^{+}02]. Cationic [JRJ01, TBG00]. cations [ALC08, GSB09, GS04, GWL07, HIA03, Hol05, NSB08, RRS06, SZT08, WWT08, ZWY^{+}09]. caused [LPK07, TT08]. cavities [BCIB05, BHH^{+}09, IME02]. cavity [RRZA08, ZFL^{+}05]. C − [CJW^{+}09]. CBr [FHF^{+}01]. CBS [Lu09]. cc [Wib04, GYMN07]. cc-pVTZ [WD04]. CCH [ZKZ^{+}07, ZKZ^{+}07]. CCl [ZLLS06a, FHF^{+}01, WDZS07]. cclib [OTL08]. CCN [JDWS06]. CCSD [BBI^{+}09, Lu09, PFJ^{+}03, PV03]. Cd [GPSP06, XB08, BMRF01, BBI^{+}09]. CD38 [UNHYT06]. CDOCKER [WRBV03]. Ce [SNM^{+}06]. Ceccarelli [Ano06b]. cell [Gon07, KVS^{+}06, KS05c, LEV^{+}09]. cell-based [Gon07]. cells [CCCJ09]. cellular [XWC09]. CeN [VP08]. centauric [PA05]. center [BRS07, GGA00, IN01, Lai07, MGMM07a, MGMM07b, NR04, OON01, SGPS09, TBG00]. centered [CCK01, SSB^{+}03, TBSM09, WSC09]. centers [GYMN07, JKL08]. central [CGMPT^{+}08, CM09]. CeO [CCCJ09]. ceramic [HZX04]. cesium [HD06]. CF [mJlZsLyL07, YLWL09, LDC^{+}07, gThDjL^{+}01, UTM^{+}02, UTT^{+}04, WLL^{+}07b, YLW^{+}08, YLWL09]. CFCl [mJlZsLyL07]. CFF91 [TTB01a]. CFMC [NCO^{+}05]. CH [CPJ00, GBDP05, HTN03, IN08, mJlZyL^{+}08, LW04a, LDT^{+}02b, MGLL03, Mui05, RD00, SLL^{+}04a, TJM^{+}03, WLL07a, WLL^{+}07b, WDX^{+}02, YLW^{+}08, ZZL04, ZZW^{+}07, ZZL^{+}08, ZZL^{+}09, HKHN08, FHF^{+}01, GD06, HLLS05, mJlZsLyL07, mJlZyL^{+}08, KZY09, KKJH08, LW04a, LWY^{+}09, MM05, Mas01b, OO04, OO08, SEKS09, SLL^{+}04a, SSS^{+}09, TJM^{+}03, UTM^{+}02, UTT^{+}04, WLL07a, WDZS07, XLL^{+}02, YLW^{+}08, ZLLS06b, ZKZ^{+}07, dOMSL01]. CH/ [MM05, OO08, SSS^{+}09]. chain [BHG03, DLW06, Der09, Din00, ENM^{+}04, GT03, HFHL06, JPF^{+}00, KG02, Kró03, LLA01a, LL01, MT03, PFC03, SMG09, TBG00, SWR06]. chains [Cri04, CA07b, DLHC06, MMLC05, MSR04, VM02, XLT07, ZM06]. chair [RP09]. chalcogena [HKHN08]. chalcogenides [JT06]. challenge [FKRE08]. Challenges [Fie02]. challenging [BS06, NGTB03]. chameleonic [PA05]. Change [KIFK07, OFIK09]. changes [HHP04, JO02, Kar06]. channel [CXZ^{+}09, FCP^{+}04a, FCP^{+}05, HSMT04, MCR08, RAGLL09b]. channels [DAK08]. chaos [LSY02]. character [ALC08, OV03]. characteristic [YGZZ05]. characteristics [DF06, KZY09, LZZC09, SLC^{+}09]. Characterization [FGR07, GGP09, GTC06, HYT05, HTSR04, DLW06, FDSA00, KHY00, LXL07, Mit01, SC01, TSMNG01, YPNE09]. characterizations [LS08a]. characterize [Wou00]. characterizing [PHJ^{+}08]. Charge [CM09, HT05, JKL08, MZ05, SWM04, ZY01, BB05, BSP06b, Chi03, DWC^{+}03, ECM^{+}03, FHRR07, GY08, GDV03, GGLR00, GY06, GHBB04, HMOG07, IC08, Jac09, JVVK09, KS01a, LFZS04, LLS03, OR05, PB04, PMB04, PP08b, PMPGP05, RLER05, RSN^{+}02, SL09, San01, SKSH07, SHSF05, SRB02, SvDS01, TCT03, WSM^{+}09, WM04, XL02, XLT07, YK08, ZBS03, ZH08, SDCG02]. charge-based [HMOG07]. charge-density [ECM^{+}03, XL02]. charge-scaling [GY06]. charge-transfer [GGLR00, LLS03, ZH08]. charged [PPYS08]. charges [BSÇ^{+}01, BCNs07, CR09a, CGBF05, DVP^{+}02, GHBB04, HS01, JBJB00, JJB02, KGL07, KC01b, LMV07, PGG06, SR09, TBSM09, TGGP^{+}00, UBDPJ04, WMS06, XLT07, YOB^{+}08, dSGCG00]. CHARMM [BBM^{+}09, HNL08, HMD06, JKII08, LLL03, MM05, MMMY07, PB04, PMB04, WHG^{+}07, WRBV03]. CHARMm-based [WRBV03]. CHARMM-GUI [JKII08]. CHBr [ZWL^{+}05]. CHCl [LDT^{+}02a, ZLLS04b, ZLLS06a]. CHCLOCHF [YLWL09]. Chebyshev [II02]. chelation [TFZRG01]. Chem [Bof01, HNWF12, Kne05, KWK^{+}00, Qua01, Van08, WHG^{+}07]. Chemical [BHTCG07, BBC^{+}05, GCB03, HLS07, Jac09, MGMM07a, PBF07, PYCD03, PYS05, WPS02, AM07, AGMPRG^{+}08, ATH^{+}03, CZFH07, CDD^{+}02, CWWS07, DF06, DBS07, DA01, DPM09, DSS03, DMN05, Dra00, DHW^{+}07, Fau01, FVB08, FR06, FLK^{+}07, FK07b, FHF^{+}01, FO04, GR07, GGB07a, GGB07b, GBB07, GS04, HWFN01, HHP04, JHZ09, KFNH08, Kau07, KBL08, Kle03, KIM^{+}09, KC01b, KGD06, LWK08, LZZC09, LHP01, LDTS07, LKA01, Mat03, MA05, MC06, NRKH02, OYH05, OKH^{+}02, PPXP01, PAS08, PFC03, PG04, PHKG07, PC05, PC07, PRS04, PV07, RNG03, RM07, Rud05b, Rud05c, RSS09, RON02, SAM06, SM08a, Sch03, Sch00, Sha07, SC01, SS05, SHH07, SFRS01, SCP08, TLOG00, Tru07, TT02, UNM^{+}01, VBS09, VBGL^{+}00, VKCK09, WS05a, WFHP01, WHP02, WWS07, WZXY07, XYN^{+}06, ZB07, ZMH^{+}09, HP05]. chemically [AVS09, Bud07, SB01, PP08b]. chemicals [CMGDAC^{+}07]. chemisorption [KKJH08]. Chemist [SH08, Bic09, Gan09]. Chemistries [Duk01, EA08]. Chemistry [Ano05b, Ano06a, Ano06b, Ano06c, GBL^{+}05, vRS98, WB04a, WWC^{+}05, Woo01, You11, tVBB^{+}01, APG05, BWM^{+}09, BT00, BMRDB01, BS06, BSJ01, CMàGL^{+}04, CFS^{+}08, CPUGD09, CMGDAC^{+}07, CMCB08, CMA^{+}08, DBM03, FJP07, FKRE08, GDPCPU07, GDPP08, GdSuM^{+}07, GdAcV^{+}07, KSB^{+}02, KBA^{+}04, KJVW08, KYL03, KC01a, LX07, MGCA07, MR09, MBP09, MMRVH07, MPF00, Nye07, OTL08, SH07, Sha07, SBB02, SGD06, TKH07, Vis02, Ano01c, Ano04b, LB99, Lip00, Sta00]. chemistry-based [SBB02]. chemists [Pra01]. Chemometric [HPL03, MRS09]. chemometrics [BLF02]. chemotaxis [FC06]. CheY [FC06]. CHF [UTT^{+}04]. CHFOCHF [YLWL09]. Chichester [Lip00]. chief [Lip00]. CHIMERA [NSU^{+}02, PGH^{+}04]. chip [Höf05]. Chiral [ZPL07, CGMPT^{+}08, Sza08, ZOJ^{+}06]. chirality [CGMPT^{+}08, PDC^{+}08]. chiraly [ST04]. chlorinated [DA01, WDZS07]. chlorine [mJlZsLyL07, mJlZyL^{+}08, WLL07a, XLL^{+}02]. chloro [JKM08]. chloroform [CCK01]. chloroform/water [CCK01]. chlorophylls [LKT04]. Choice [Duk01]. Cholesky [SKDO08]. chorismate [HHBH00]. chromium [RRS06]. chromophore [DHM^{+}03, HFS^{+}07, KHF^{+}09, XZ05]. Chun [Ano06c]. CI [Ano01b, DHM^{+}03, HFS^{+}07, HKHN08, IK00, dSVA^{+}09]. Cieplak [Ano06c]. Circular [AB08, MM00, HKHN08]. cis [DMN05, GRO^{+}03]. cis-diammineplatinum [DMN05]. cisplatin [BZL05, RP04, WM01]. CL [FKŠ^{+}09, CRC^{+}08, DMN05, GZ07, Han01, HYA02, HTN03, HZ09, Hua09a, mJlZyL^{+}08, KBL08, LF02, Mar03, RB01, STC^{+}08, WLLS04, WDZS07, XFF06, ZY01, ZL09b, BS03, HLLS05, mJlZsLyL07, RFSS06, SLL^{+}04a, WLLS05, YLW^{+}08, YLWL09, ZWL^{+}05]. class [CKR08, EBD^{+}01, LL07, VIP^{+}06, YNW05, hYDN^{+}08, Car02, PCS04]. classes [CLF^{+}09, KH06, QLHL09, XSHC06, XLC08, XWC09]. Classic [SRE08]. Classical [ATBLS04, COL^{+}06, DK01, LLM09, MA05, Nil09, RP07a, STH02, Zho06]. classification [GDPCPU07, dGWH01]. classifier [CLF^{+}09]. clathrate [EM03a]. clay [ATH^{+}03, DJT08]. cleavage [CLXC02, NLL^{+}09]. cleavages [XWXC08]. cleft [SPT^{+}03]. closed [DSB^{+}02]. closed-shell [DSB^{+}02]. closo [JRJ01]. closure [CSJD04, DLSVY00, Mak08, SWR06]. Cluster [AHGK09, Cri04, BR04, BP01, BGJ01a, BWI^{+}02, IN08, JBGK08, KKJH08, KSTC01, LMJ02, LWX07, LYS08, Mck07a, MS01, Mor02, NK06, PSF^{+}08, SSB07, VDM06, WKYU01, Whe08, WJX^{+}08]. cluster-continuum [WJX^{+}08]. cluster/adsorbate [BWI^{+}02]. clustered [FPN^{+}05]. clustering [CCWH02, FKZ09, LZ05a, LOL^{+}08, RLA01, ZS04]. clusters [BP00, BACJCT01, CGG06, CAG07, DSB^{+}02, GBBH09, HXD08, HYR06, JHMB^{+}09, JHMB^{+}11, JG03, KGL07, KDG^{+}09, KZW^{+}05, LML^{+}00, LWLS07, LJS05, Mck07a, Mck07b, NBJ04, OSHS03, OS08, PBZ00, Pul05, QB05, SCC04, SYC08, SW06, WLL01, WZZ^{+}09, WCS09, XZ04, YCS07, ZLJS03, ZXL^{+}04, ZWC^{+}09, ZXY03, ZGXX06, ZX09, ZB07, Est07]. CM1 [UBDPJ04]. CM3 [UBDPJ04]. CN [LW04a, TJM^{+}03, TYN05]. CO [GD06, PLC08, RD00, TJM^{+}03, VS02, KT08, ABYM08, DLD^{+}02, FLK^{+}07, PBZ00, SS05, PLC08]. co-ligation [KT08]. CoA [LLL^{+}08]. coagulation [PDP02]. Coarse [CA07a, EBAN07, VTT^{+}08, CP09, DR07, DJB02, HXLS09, MBC08, PSHP08, SBJ08, WWL^{+}09]. Coarse-grained [VTT^{+}08, CP09, DJB02, HXLS09, MBC08, SBJ08, WWL^{+}09]. Coarse-graining [CA07a, EBAN07]. coastline [UTH^{+}03]. cobalamins [KPZK06]. cobalt [LMIF06]. cocaine [ZDS^{+}05]. cocrystals [CWV^{+}05]. code [BDW00, FROD08, GBL^{+}05, GJK^{+}06, GY08, PZWG^{+}04]. codification [CGMPT^{+}08]. coding [LCSZ09]. coefficient [CSB^{+}03, YSJ09]. coefficients [CCK01, DA01, GS09, Gol09, LZZC09, Whe08]. coenzyme [WC08]. cofactor [Mck07a, Mck07b]. cofactors [Ano06b, CPM03]. Cohesive [VP08]. coil [OCP02]. coiled [OCP02]. coiled-coil [OCP02]. collagen [PP08a, PRKP05]. collagen-like [PRKP05]. collapse [Mei02]. Colle [IKN08]. collections [CDD^{+}02]. Collective [CCSJ00, HSWN01, LV08, SWR06]. collide [CKF01]. collinear [Van02b]. collision [VW00, VW04, TYN05]. Collisions [Pan07]. collocated [CJDK09]. collocation [CKR08]. collocation-based [CKR08]. colony [CLZ^{+}09, CLZX09]. comb [LAEL01]. comb-like [LAEL01]. Combination [MH08b, DLD^{+}02, GHH07, HSWN01, KKS04, dSR08]. combinations [ZOJ^{+}06]. combinatorial [AL01, GD09]. Combined [CYM02, GWM08, KBLP09, RG02, ZKZ^{+}07, BAL^{+}01, BCNs07, BME05, COS01, DPT03, Der09, FSFK05, KH01, KM00, LMCD09, MC06, Oos09, TCR^{+}02, VMF^{+}03, WX09, XLZ08]. Combining [MBC08, SS00, HTKG08, KN04, LLZL09, ZMZ09, KK08b, LEK07]. comblike [LZA02, ZALMG03]. CoMFA [JFG04]. Comment [CPML08b, JW12, Kne05, Qua01, Van08, Bof01, CSD05, WM12, vLBBR12]. Comments [MBP09]. Common [ZAT07]. commonly [ESP04, SCF^{+}09]. Comp [HNWF12, Kne05]. compact [Kri09a]. comparably [PB02]. Comparative [CCP04, GHLK^{+}02, LS02, SBJ08, SCF^{+}09, Sto05, VMA03, CMD^{+}04, FC06, FO08, GM01, HLMR06, JFG04, KS01b, LPK07, McD08, PHKG07, PMM05, TBG00, WC09, EBL^{+}08, JHZ09]. compared [IO08, JČHS07]. Comparing [JCL05]. Comparison [AEE^{+}03, Bou00, CV09, DB07, DLG00, GGLR00, Kró03, SL06, SDL07, TCR^{+}02, WMGK07, WMW03, YLL^{+}09, ZCL09, AE06, BL05, FOL^{+}04, FDSA00, GDPCPU07, Gra07, HRBKB03, JARM02, LC07, LOL^{+}08, LZF^{+}09, MZ05, MS00, MST^{+}08, MC06, PK04, PS03, RDM^{+}08, SBLK01, ŠBL05, SSB07, WB04a, WB04b, WB05, BRDC02, CCT^{+}03, COL^{+}06, LKT04, NCO^{+}05, ONHN00, PSF^{+}08, SPDS01, SL04, SKK^{+}07, WMW04, YKK09, ZGFL01, ZOJ^{+}06]. Comparisons [GPK05]. COMPASS [MSR04]. compatible [BSDM04, KH06]. competing [HA04]. competitive [FG03]. complementarity [EKB02b]. Complete [BT00, MLL^{+}08b, Pog03, PRT^{+}07, PRT^{+}08, MC06, Var09, WMRW^{+}01]. completeness [AHK02, MV06]. completeness-optimized [MV06]. Complex [DFGB09, AS06, Bac09, BRS00, BRS01, GC04, HDF^{+}07, HDO^{+}02, HMK02, IV04, IvSV06, Ish04, JLHF03, Kle02, KVS^{+}06, LB05, LDL^{+}09, MM03, MCF05, MY08b, MY08a, NHN06, Pac06, ZWB09]. complexation [AGI^{+}07, HT05, LMMW04, SRK^{+}00, SLRC01]. complexed [Pin03, SDM02, WCF04]. Complexes [APG05, AB00, Ano06a, ACM^{+}06, BTP09, BR04, BL06, BM00, BGC^{+}09, BZL05, CZ05, CG03, CBC^{+}08, CSB08, CBH^{+}03, DPT03, DF04, FRS05, FO08, FRLN09, FKŠ^{+}09, GTC06, GL04b, GM01, GPSP06, GPK05, Gri04, GZM09, GWL07, HLLN06, HRG07, IN08, IO08, IGL07, JMD^{+}02, JD09, JČHS07, KT08, KRM^{+}02, KJP^{+}07, LL00, LHJ^{+}06, LPP06, LH02, LMGR06, LLS03, LMMW04, LWZ09, LZF^{+}09, LS05b, Mas04, McD08, MHJS06, MSBS01, MLL^{+}08b, NyHN06, NR04, NMAT01, PGG06, QTdG^{+}08, RPNJ07, RMP01, RRFC^{+}03, SG07a, SCF^{+}09, SBH02, ST06, SVV^{+}08, TGGP^{+}00, UM03, VS02, VMA03, VL00, WB07, WWT08, Won09, ZY01, ZBS03, ZWY^{+}09, ZTS09, dVB01]. complexes* [GK09]. complexity [BT00, PK05, XSHC06]. Component [KBA^{+}04, CCT^{+}03, GPSP06, JMD^{+}02, PVdJB00, PV03, SH02, SM08b, Van02b, WG02]. Component-based [KBA^{+}04]. Components [KJVW08, Car02, LL07, TGGP^{+}00]. composition [HM06, KWHH07, LL07, PAS08, XSHC06, XLC08]. Compound [CN05, BR07, HBM06, RD00, XZ05]. compounds [ACLD03, BB08, BLO^{+}02, CYM02, DA01, EBDPM00, EDAJ04, EBD^{+}01, FJ08, FROD08, FO04, Gol09, Gor01, JLHF03, KFD06, LLA01a, LLA01b, LLA01c, LLA01d, LLA03, LD05a, LWK08, LTF^{+}07, LWH06, LCDA03, LCGA03, LCA03, LJZ^{+}07, LLM09, MDI04, NBTN04a, NBTN04b, PZWG^{+}04, PYCD03, POJ01, RUPH06, SJJ^{+}04, ST01, TTBM09, YCW^{+}09, YSA^{+}03]. Comprehensive [LF02, ZL09b, ZB07, DLR^{+}08, JPF^{+}00, ŠBL05]. compression [BG07, MBWP03]. comprising [Rud05b, Rud05c]. Comput [Bof01, Qua01, Van08]. Computation [BLL^{+}06, Car02, CDD^{+}02, FZL07, GS09, LFEdL06, TNS00, ATMK03, Gon07, NKIS02, PAS08, RK05, WZXY07, ZCZ03, vW06]. Computational [Ano01c, Ano04b, Ano05b, Ano06a, Ano06b, Ano06c, BLMS08, CZ05, CPUGD09, CMGDAC^{+}07, CMCB08, CG05, CA07b, GBL^{+}05, GDPCPU07, GdSuM^{+}07, HMK02, LKJ^{+}04, LB99, Lip00, MW00, MFR07, OY03, Pac06, SPGS08, vRS98, Sta00, TMBM02, WB04a, WWC^{+}05, WXK08, You11, ZOJ^{+}06, ZWB09, Bac04, BG00, BAL^{+}01, BTP09, BMRDB01, BZL05, CMLS05, CFS^{+}08, CBC^{+}08, CJS^{+}03, CRH^{+}07, FG03, FJP07, FD03, FKRE08, Gan09, GD09, GDPP08, GdAcV^{+}07, GGB07a, KYL03, MCF07, McD03, MH08a, OTL08, OY01, OSA06, Pan07, PMM05, PWFS01, PRS04, RGG08, SF07, SM03, STCJ08, TD08, Vis02, VZM^{+}08, WSM^{+}09, WOC^{+}03, YT03, YT04, YDWS06, ZZ08, ZGZX07, ZTP^{+}08, ZM06, Woo01]. Computationally [KM00, KFZ03]. Computations [Bou01, Dib05, CLP^{+}05, GPK05, Lu09, ME06, Pog06, SMZW05, TT08]. compute [BDW00, RKA^{+}09]. computed [PFJ^{+}03, PK05, TDH06]. computer [HFSD03, Höf05, NK01, PPXP01, PHJ^{+}08, TRS02, UIHN09, VB07, VKCK09, YNZ^{+}08, Zer08]. Computing [HHW^{+}03, WL00, BHH^{+}09, DGHR02, DP04, DK01, GLD08, HHS^{+}05, KFB05, LM03, MA09, ZP03]. CONAN [SSHT03]. concave [Won09]. concentration [GGT08]. concept [LSY02, Rao00b, Rud05a]. Conceptual [VB09]. Concerning [FG03, Bor03]. concerted [LFS^{+}07, Mck07b, NSB08, LLKC06]. condensed [CLP09, DGI^{+}08, DWC^{+}03, FM00, GLMV09, Mor02, SDvG01, ZSK07]. condensed-phase [DWC^{+}03]. condition [SK08]. conditions [BVW04, CEP07, HH04, WM06]. Condon [Ama02a, LMCD09, TP01b]. conduction [SM06]. conductor [DHW^{+}00, FZL^{+}06]. conductor-like [DHW^{+}00, FZL^{+}06]. Configuration [LSAS01, SLRC01, ASS^{+}02, BMRF01, DLD^{+}02, KBT03, LWX07, MLL06, PRSMM02, PRSMV08, SSL02, SWZS04]. configurational [HDF^{+}07, HTKG08]. configurations [MM03, MABM09]. Conformation [GS04, BY06, BR03, BLB09, DPM09, LXL07, PFJ^{+}03, YL06, YZ06]. Conformation-dependent [GS04]. Conformational [AJNG01, AZS^{+}04, BLF02, DSS03, HW09, KK09, LRI^{+}02, LGB^{+}09, OYK^{+}09, OM04, PSDM00, SH09, SMGE08, SSHT03, SR09, TTB01a, TTB01b, WHRG08, WW04, BISB02, BMTSC01, BAH^{+}02, CN03, CSJ01, CvG08, CS01, CCP04, CSRST04, CKT^{+}08, DDKV07, DSR^{+}07, ECA06, FGR07, FC06, FD03, GT03, GSB09, HYA02, IB04, IZA06, JW06, JO02, KB02, KK08a, KH01, Kle02, KK01a, KF08, LFKL00, LKJ^{+}04, LCKL05, LJKL08, MFB04, Mak08, MH08b, MA05, MGJAARC00, NKIS02, OML^{+}00, OGH05, PRT^{+}07, PRT^{+}08, SPL^{+}02, SWBM08, SHBD05, SD09, SSBE06, WCK00, YXL^{+}09, ZA07, CG03, HJCP01, JPF^{+}00]. Conformations [NHH05, CLWL09, CIB05, yCkHmY08, DBM03, ENM^{+}04, FWH^{+}07, GBB07, GB04, HDO^{+}02, HP05, LZKT04, MW00, OFB08, OKH^{+}02, PC00, PFC03, Rao00b, RP09, TLKT00]. conformers [HS00, HHP04, KAS^{+}07, PG01, RSSKB03, TT02, WD04, YXL^{+}09]. Congo [SRK^{+}00]. conical [IK00]. conjecture [Pog03]. conjugated [BG00, CZ05, DDBP09, LYS08, LC09, LFR07, WJ00]. connectivity [EDAJ04, Pog06]. conquer [AKN07, MLJ03, vdVGDM00]. consensus [GP06, JMD^{+}02, LLL^{+}08, RHL09]. Consequences [RSS09]. considerations [GRCD01, PB05]. Consistent [RP02, BWI^{+}02, ECM^{+}03, KK08a, KBN02, NUH02, SMD02, VTT^{+}08, WM04, XL02]. Constant [MCM04, DRMD03, Sch00, Vas02, WLLS05, WLL^{+}07b]. constants [Chu07, FCW06, GGB07a, GGB07b, HWFN01, JHZ09, Kle03, MMLC05, MGLL03, MDI04, PJPJdPRMI07, RRFC^{+}03, SLL^{+}04b, SFRS01, TLOG00, WDX^{+}02, WZXY07, ZXY08]. constrained [COS01, EC06, LFKL00, MM00]. Constraining [AM09, HSWW00]. constraint [BL09, FS98, FS00a, KvGH01, YXC^{+}07]. constraints [BVW04, Bud07, BRS00, BRS01, ECA06, Pen06, PJB^{+}07, VMF^{+}03]. constructed [Gri06, YCS07]. Constructing [ZBS03]. construction [HH04, RSN^{+}02, TYO^{+}02, UIHN09]. contact [ENM^{+}04]. contained [LH02, SH07]. containing [BS06, FLOD07, FPN^{+}05, JPF^{+}00, LKJ^{+}04, LWW^{+}06, LFR07, LLM09, MSH^{+}06b, PFC03, RRZA08, SL09, STC^{+}08, WL04, Wil01b]. contemporary [CFS^{+}08]. content [CLC03]. context [KMH02, KBT03, OCP02]. contexts [Sim07]. Continuous [FZL^{+}06, LF04, LFZS04, ZFL^{+}05, HHHS01, SM08b, PDC^{+}08]. Continuum [FCP^{+}04b, MGLO03, ABWT09, BCIB05, CFK08, COL01, CCT^{+}03, COL^{+}06, FKL^{+}06, FEVM01, FBLO08, GS02, GWS^{+}02, HC08, HS01, HHP04, KKS04, LRI^{+}02, LJ04, MWL^{+}08, Pom04, RSP03, Sch00, STSF02, TJE03, VP09, WJX^{+}08, YÇBM00, ZFW08]. continuum-solvation [Sch00]. contours [YGZZ05]. contracted [GS09]. contraction [CGSdST06]. contribution [BCIB05, CR09a, DBS07, KC01a, LKA01, PWHF^{+}03, PWHF^{+}04, PMPGP05, RI07]. contributions [CBC^{+}08, CR08, CSB^{+}03, COL^{+}06, GPK05, HIM07, MGLO03, RM07, SWV^{+}05]. Control [Kar06, DB06, LR06, MN02, RS05]. conventional [WMW03, WMW04]. Convergence [KGN07, LST08, GG09, LTV08, LJ04, Mas04, Rud05a, Rud05b, Rud05c]. convergent [PAS08, Zho06]. Converging [GC04, KF02b]. conversion [CC09, CFD04, PHR^{+}05, RR05]. converting [RM00]. Cool [BHG03]. cooperation [ATMK03]. cooperative [HLB09, RRCA08]. coordinate [BGC^{+}09, HDBD04, Ish02, KTA03, LN01, RWBH09]. coordinated [GWL07, Sha02, SBH02]. Coordinates [EA06, Din00, EC06, GKK07, KSU03, LPK07, NKIS02, QCK01, QCK02, TNS00]. coordinating [JRJ01]. Coordination [KZRO03, Gor01, HXD08, SFR07, TBG00]. coordinations [DSB^{+}02]. Cope [PA05]. copolymer [CHA^{+}07]. copolymers [YFR05]. copper [CR02, DBS07, FNP^{+}06, PMGL03, PBZ00, PS03, ZWC^{+}09, ACM^{+}06]. corannulene [Won09]. core [ATM^{+}07, CM09, FR06, HXD08, HYR06, HJCP01, ION07, LFK05, LK03, LK04, NTH09, ON07, Pog03, TJM^{+}03, TLKT00, TJE03, TKN^{+}08, YCS07]. core-excitation [ON07]. core-excited-state [TKN^{+}08]. Cornell [RKH03]. Coronavirus [LXZ06]. corrected [NTH09, PSC^{+}01]. Correction [Duk01, CFC^{+}08, GAI06, Gri06, HK08c, HK08d, IKN08, LBT07, MGLDS00, Mue01, QCK01, QCK02, SPDS01, STSF02, TKH03, WX09]. corrections [BF07, Gri04, KSS08, WB07]. correctly [LF04]. corrector [Kol04]. correlate [Kle03]. Correlated [GBB07, BWW^{+}08, BLT03, DBM03, GPSP06, KMA^{+}07, PFB05, TBG00, WMW03, WMW04]. correlating [NSO^{+}07, SNM^{+}06]. Correlation [LRWG03, TDH06, AAP00, AGI^{+}07, AS00, BL00, CKT^{+}08, EL09, FDM00, GKTS04, HN02, HR08, Hir08, IKN08, JJK^{+}00, JSHG07, KGN07, KK08a, KC01b, LMJ02, Mat03, PJPJdPRMI07, PMC^{+}08, SRE08, TKH07, WMW03, WL04, YH09, dSGCG00, WMW04]. correlations [DR09, HHW^{+}03]. Correspondence [RLRE01]. corrphycene [NyHN06]. COSMO [EK06, ELK^{+}09, KEH^{+}02]. COSMO-RS [EK06, ELK^{+}09, KEH^{+}02]. COSMOS90 [SO07]. costs [SSL02]. Coulomb [WMW04, BWP07, Nee03, WMW03]. counterions [JD09]. counterpoise [GAI06]. counting [HYT05, LM03]. Coupled [BSP06b, MO01, CXZ^{+}09, DOSG06, IN08, KSTC01, LMJ02, LYS08, PSF^{+}08, SSB07, WKYU01, Whe08, XWC09, SMAdV00]. coupled-cluster [IN08, KSTC01, Whe08]. coupling [CR08, DXW08, GdAcV^{+}07, KTM02, KBLP09, LB08, MDI04, PJPJdPRMI07, QTdG^{+}08, RI08, RLDI09, RRFC^{+}03, SM08a]. couplings [BPC01, NR04, TP01b]. CoV [LZ05b]. Covalent [BSG07, BMTFR08, PML03, RS07a, RS07b]. covalently [PHFC04]. coverage [SURG06]. covering [RKH03]. COX [WC09]. COX-1 [WC09]. COX-2 [WC09]. CP [ZKZ^{+}07]. CPHF [ASWG07]. Cr [KPR04, Kri08]. Crehuet [Bof01, Qua01]. criteria [Kle03]. criterion [ALTB06, GLD08, PSDM00]. Critical [GT03, BMLV04, BLN01, BAÅ07, CRC^{+}08, CKMC04, FMPS08, LFR^{+}04, MP03b]. cross [Gan09, MY08b, MY08a]. crosscorrelation [HWDB03]. crossed [BAL^{+}01]. crossing [LI07]. crossings [LSG06]. crossover [KRLD09]. crown [GLRL02, HDO^{+}02, LWW^{+}06, ZWY^{+}09]. crown-shaped [LWW^{+}06]. cryogenic [HN02]. cryptand [WWT08]. Crystal [KOFF09, Van02a, DPT03, EL09, HN02, KP05, TD08, VVBV02, vDSSvA04, vEMK01, vE01, DRMD03, FROD08, PZWG^{+}04]. crystalline [AS00, CADW03, JB04, PZWG^{+}04, Wil01a, ZLD09]. crystallographic [RON02]. crystals [BCF^{+}09, CC07, FÁ01a, GAdGM08, GBJ03, PMC^{+}08, RD06, WMS06, Wil01b]. Cs [GWL07, GLRL02]. CSA [NCO^{+}05]. CSOV [GPSP06, PMPGP05]. Cu [BTP09, GPSP06, Sha02, HSF08, NK06, TDK07, WCS09, ZTP^{+}08]. CuN [ZX08]. cuprates [MDI04]. Current [NYTH09, CDPL09, Vis02]. curvature [TRS02]. curved [ABWT09]. curves [BBI^{+}09, MM07, SSS^{+}09, ZLY07]. Customized [BDW00]. cut [BME05]. cutoff [GGT08, KLM^{+}09]. CuX [KBL08]. cyanines [BG00]. cyano [PA05]. cyanoacetylene [YDWS06]. cyanobiphenyl [CC07]. cyanoboranes [WCW08]. cyanomethylidyne [WDS06]. cycle [ZAT07]. Cyclic [KJP^{+}07, BGJ01a, CLA^{+}00, FKM^{+}06, FKM^{+}07, JBGK08, LXL07, OYK^{+}09, VVS07, WOC^{+}03]. cyclic-AMP [FKM^{+}06, FKM^{+}07]. cyclization [PWFS01]. cyclizations [SGS03]. cyclo [TDK07]. cyclo-Cu [TDK07]. cycloalkanes [SSBE06]. cyclobutane [QZZZ03]. cyclobutene [SRE08]. cyclohexane [MT03, RP09]. Cycloketones [LLA01b]. cyclononane [SSBE06]. cyclononatriene [ZSE08]. cyclooctatetraene [CPFL02]. cyclopentadienyl [ML00]. cyclopentene [SURG06]. cyclopeptidic [FL07]. cycloreversion [QZZZ03]. CYP2A6 [VB09]. cysteine [CN05, MOP^{+}07, PMM06]. cysteine-6 [PMM06]. cysteines [CFR06]. cytochrome [AST06, ATBLS04, HBM06, JKL08, LCC09, OYH05, OON01, ZAT07, BS06]. cytosine [KKMMS04, MDA08, MHS05, MH08a, SBI08, SG07a, SC01]. cytosine- [MHS05]. cytosine-5-acetic [MH08a]. CZ [CRC^{+}08].

D [IS03, PF06, SHBD05, AGO^{+}02, BAH^{+}02, CPC^{+}00, DDBP09, DMC05, FROD08, GDPCPU07, GdSuM^{+}07, GdAcV^{+}07, HP05, LW04b, LXZ06, LW06, MP03b, OYK^{+}09, RSSKB03, RGP^{+}07, SFC04, YNW05, hYDN^{+}08, ZTS09, vDSSvA04, TGLL07]. D- [AGO^{+}02]. D-arabinonate [RGP^{+}07]. D-arabinonohydroxamate [RGP^{+}07]. D-Epitope-Explorer [SHBD05]. D-erythronic [vDSSvA04]. D-galactose [RSSKB03]. D-QSAR [DMC05]. D-RNA-coupling [GdAcV^{+}07]. D/ [PF06]. D180 [NYK^{+}09]. d2 cluster [CCWH02]. damage [FPN^{+}05]. dangers [MBP09]. data [ASWG07, BRDC02, BK00, CDD^{+}02, CRGN07, FOK^{+}04, FM00, HHJ03, HSWN01, KMH02, KMA^{+}07, LEK07, MBWP03, MMP^{+}07, PFJ^{+}03, PF06, RLA01, RRS07, RRS09, RON02, SY09, SFC04, WG02]. database [DPM09, LFKL00]. databases [BR07, PPXP01]. dative [FH01]. David [Woo01, Ano05b]. day [GR07]. DD [ZLY07]. DD-curves [ZLY07]. Dead [YFS07, Adc04, GLD08, KUB07, PSDM00]. Dead-end [YFS07, Adc04, GLD08, KUB07, PSDM00]. deaminase [MDA08]. dearomatization [HT05]. debates [Nye07]. decarboxylase [HLC09, LLL^{+}08]. decarboxylation [UTH^{+}03]. decker [RPNJ07]. decomposable [VZM^{+}08]. decomposition [BM07, CBH^{+}03, FK07a, FPG^{+}06, Hir08, KZY09, KN04, LBG08, ML00, SKDO08, TBSM09, TCR^{+}02, ZZL04]. decompositions [GPSP06, PBF07]. decoys [LZ05a, SRCD03]. defect [ZMH^{+}09]. defects [JT08]. Definition [EA06, LFSB03a, LFSB03b]. Definitions [PBF07]. Definitive [dOMSL01]. deformation [GHBB04]. deformations [Din00]. deformed [RLER04b]. degenerate [NUH02]. degradation [PCMG09]. degree [CC09, RLER07]. degrees [DHF^{+}05, MZL08]. dehalogenase [NYK^{+}09]. dehydration [TT02]. dehydrogenase [SS05]. dehydrogenases [JJH01]. dehydrohalogenation [TT02]. deletion [SHH07]. delineate [MP03a]. delocalization [BY06, BI06, FVB08, FS02, Kar06, MGMM07a, MGMM07b, WMW03, WW03, WMW04, Wan09]. deltorphin [OM04, YAÇ^{+}02]. deMon2k [GJK^{+}06]. denatured [GB04]. dendrimeric [SCG04]. densities [GY08, GBJ03, HSWW00, KCK^{+}08, LMV07, RLR^{+}04, VZVG06, Van02b]. Density [BP01, FG02, Han01, JČHS07, KWK^{+}01, KWK^{+}02, MSBS01, QZZZ03, QZL^{+}04, VL00, WCW08, AB00, ABYM08, AEE^{+}03, ASY01, Bac09, BP03, BMLV04, BB08, BAÅ07, CLP^{+}05, CRC^{+}08, CFK08, CRS05, CR08, CSB08, CAG07, CPML08a, Cul04, CGSdST06, DVP^{+}02, DVRP^{+}03, DF04, ECM^{+}03, FCW06, FZL07, FDM00, FS04, GHLK^{+}02, GLRL02, Gri04, Gri06, GBBH09, GHBB04, HGMB04, HLS07, HNWF07, HNWF12, HN02, Hir08, Hol05, II02, ION07, IN08, IB04, ITN^{+}05, IS07, JNV08, Jac09, JCA^{+}02, JFG04, KGL07, KRM^{+}02, KN04, KSS08, Kle03, Kni00, KZW^{+}05, Kri09a, KS01b, LRI^{+}02, Leh06, LV08, LMB08, LMGR05, LLS03, LWH06, LKT04, LF02, LLZL09, LDL^{+}09, LZF^{+}09, MP03a, MW09, MS00, NK06, NTH09, NAT07, OKE^{+}02, PSF^{+}08, RB01, RK04, RLER04b, RDM^{+}08, RR05, RZWS07, SH07, SZT08, SPT^{+}03, SCF^{+}09]. density [SLRC01, SSB07, SW06, TBG00, TST^{+}08, TKN^{+}08, TKH03, Van02b, VMA03, VBS09, VC04, VKCK09, WRP^{+}06, WB07, WZY04, WMRW^{+}01, WL02, WCHW09, WM04, WCL05, WZXY07, WM01, XB08, XL02, XPW09, YTH01, YL09, YK08, YYW07, YLL^{+}09, ZZL04, ZH08, Zho06, ZM03, vGGB00, Haf08, LWK08, MW00, XYN^{+}06, GM01]. density-functional [HNWF07, HNWF12, LLS03, LWH06, TST^{+}08, XB08, Haf08]. density/polarization [YL09]. deoxyguanosine [MM02]. deoxyribonucleoside [PFR04b]. deoxyribose [LBG08, SA07]. Dependence [ASS^{+}02, MGLL03, BRLS08, BRLS12, BL00, KH06, NK06, SR09, TJM^{+}03, VKCK09, ZP03, ZXY08, DvG00, DPM09, MG06]. dependencies [FHF^{+}01]. dependency [OKH^{+}02]. dependent [Bac09, CFK08, FCW06, FCP^{+}04a, Gog08, GS04, HNWF07, HNWF12, HS01, ION07, LDY^{+}08, LDL^{+}09, LSW^{+}01, MML02, MW09, MY08a, NTH09, ONHN00, PSF^{+}08, TST^{+}08, TKN^{+}08, Whe08, WC08, YH07, ZH08, ZM03, vGGB00, PMM05]. depiction [ZTS09]. deposit [JG03]. deposition [UNM^{+}01]. deprotonated [Mas04]. Derivation [EBD^{+}01, JFG04, TT05, TTB01a, EBD^{+}01, HZ06a, Tor02, Tot04]. derivative [CNN07]. derivatives [BT00, Bor03, BC06, CJK^{+}02, COMR^{+}04, DMC05, DOSG06, FL08, GLRL02, IS03, PSF^{+}08, PA05, QCK01, QCK02, RP09, SPGS08, SGPS09, Sch00, STC^{+}08, TNS00]. derived [GBJ03, HSWN01, Ish02, KŠB09, KS06, KFNH08, MLJ03, SvDS01, TBSM09, WMS06]. Deriving [RPMP03]. desaturation [BBSS06]. descreening [MTE04]. describe [DDBP09, IDMC09, MSH^{+}06b, RLDI09, SBI08]. describing [CMàGL^{+}04, HK08a, HK08b]. Description [ION07, MHT01, BUMCMRL00, BME05, CLWL09, CHRL09, Gri04, HGMB04, SM08a, VMA03]. descriptions [SB08]. descriptor [CDS09, RSS09, TCSM03, XYN^{+}06, ZNLL07]. descriptors [AGMPRG^{+}08, BAÅ07, DA01, EDAJ04, HM08, HMMS09, Jac09, LXW^{+}09, MGMM07b, RUPH06, Tie09, TTBM09, Wou00]. Design [AG03, KV00, BSP06a, BMTSC01, BLMS08, CRH^{+}07, CMBC08, DB06, DHW^{+}07, DHW^{+}09, GHMP03, Ham07, HM06, HLTLP09, HLM05, JGVF05, LFBSK07, LZ05b, LFS^{+}07, MWE02, NHH05, PS09a, SPGS08, SRS07, SHM04, STCJ08, VGGMM05, VZM^{+}08, YFS07, ZZ08, ZL09a]. designed [GT03]. designing [GDV03]. Desirability [CMBC08]. Desirability-based [CMBC08]. desolvation [HMOG07, SWV^{+}05]. Detailed [PB05, WRBV03]. details [GGB07a]. detecting [BHH^{+}09]. Detection [WHH^{+}06, BAL^{+}01, CMCB08, OYH09]. determinant [GS09]. determinants [BCP03, Bou00]. Determination [BLT03, CFR06, CR08, DLD^{+}02, FSS00, Vas02, BL08, BR03, BCNs07, BdPRMAI00, CC09, Chi03, CAGR08, FAR02, GCCVB00, HP05, Mar03, MGLDS00, MM07, PC00, PFC03, PABK03, RI07, RTG00, SCF^{+}09, TBSM09, vDSSvA04]. determinations [YXL^{+}09]. determine [DDVD09, KUB07, OO08, RI08, YH06]. determined [OYH05, TDH06]. determining [BY06, DV02, LR06, PHJ^{+}08]. Deterministic [LS05a]. detonation [JWB05]. detoxification [ZWS^{+}09]. developed [CRS05, KMH02, RG08]. Development [ATMK03, BGJ01a, HHJ03, IS07, KSB^{+}02, KOML08, KVL^{+}04, LAT05, LK03, MSR04, MRC03, WWC^{+}04, WWC^{+}05, WS05b, XYN^{+}06, Yan04, BA08, COS01, CMGDAC^{+}07, KLB03, NG04, BG03, IKYM09, SM08b]. developments [FCP^{+}04b, HS07a, SMD02]. DFT [BRLS12, ASDP^{+}06, ACM^{+}06, BWP07, BPC01, BP07, BSB05, BM08, BB08, BE07, BRLS08, BBSS06, BZL05, CMJ08, CCCJ09, CHA^{+}07, CG06, CS03, CMA^{+}08, DGD^{+}05, Der09, DDBP09, ESP04, EKO^{+}01, EBL^{+}08, FO08, FO04, FKŠ^{+}09, GCCVB00, GKH05, GPSP06, GKTS04, HLLN06, HT05, HSWW00, HK07, HZ09, Hua09a, JPF^{+}00, Kle02, Kle03, KTM02, Kri09a, KPZK06, LMV07, LYK^{+}04, LDC^{+}07, LWLS07, LS08b, LWZ09, LS05b, MML^{+}06, MOP^{+}07, MGG06, MBWP03, PFJ^{+}03, PMPGP05, PMM06, RM00, SBI08, SWBM08, ŠBL05, SN06, SCG04, SSBE06, SRB06, Tie09, VS02, VB09, WMGK07, WLX^{+}05, WWT08, WL09b, XLL^{+}02, XKG^{+}05, YS00, YK08, ZSE08, ZKZ^{+}07, ZWS^{+}02, ZWY^{+}09]. DFT-D [DDBP09]. DFT/MRCI [KTM02]. DFTB [ECM^{+}03]. di- [CU01, GBB07]. di-arsenic [KS05a]. diabetes [PS09a]. Diagonalization [LSAS01, BdPRMAI00, PU09]. diagram [Hir08]. diamide [HHP04]. diaminoguanidine [BI06]. diaminosilylenes [TKS^{+}01]. diammineplatinum [DMN05]. diamond [EKO^{+}01, JBGK08, ZMH^{+}09]. dianionic [OSA06]. diatomic [ALKH04, FCW06, TLOG00, WWS07]. diatomics [Cul08]. Diatropicity [CdML06]. diazonium [EL06, EL07]. diborane [wQZsLyZ02]. dibromomethane [LXSF08]. dicarboxylic [NHH05]. dication [Bac09]. dichlorides [LHP01]. dichloromethane [RRZA08]. dichroic [MM00]. dichroism [AB08, HKHN08]. didehydropyridine [KC01a]. didehydropyridinium [KC01a]. Dielectric [HS01, DRMD03, GS03, HMWC03, HLLN06, LZZC09, MML02, Vas02, ZFW08]. Diels [Hir08]. DIESEL [ME06]. differ [SRK^{+}00]. difference [ALC08, Bie04a, Bie04b, BF04, PMPGP05, PZS04, Rud05a, Rud05b, Rud05c, VZM^{+}08]. differences [CV09, GG09, OV03, YZ04]. different [ABÅ04, BL05, CEP07, CMGDAC^{+}07, DSB^{+}02, MCF07, MN02, VC04, WM01]. differential [DD08]. differently [HSF08]. diffraction [HHJ03, dGWH01]. Diffuse [GS07]. Diffusion [VW04, BZP09, Bie04a, CCCJ09, Rud05a, VW00, PK04]. difluoroethane [CUS00]. dihaloethanes [WFR08]. dihedral [FKZ09, HK08c, OFIK09, YL06]. dihydro [WJX^{+}08]. dihydrodiol [PCMG09]. dihydrofolate [GGLR00]. dihydrogen [Mck07b]. Dihydrophospholophosphole [CDL06]. dihydropyridine [HSMT04]. dihydroxypyridine [YXZ^{+}04]. diiminobenzosemiquinonate [Bac09]. diiron [BB08]. diketiminate [GTC06]. diketonate [RMP01]. dilute [HRR05, Kri09b, XZ04]. dilution [DA01]. dimension [TSMNG01]. dimensional [BP01, Bie04a, CVR08, DHW^{+}08, LAR^{+}03, LR06, MP03a, MVLG06, RSS09, SHBD05, Wan09]. dimensionality [CDGS09]. dimensions [AHK02]. dimer [CWY09, GYCZ04, Kró03, LZJ03, MMPK01, McD03, NK01, RRCA08, SBI08, SG07b, YTH01, ZGXX06]. dimerization [HK07, JJK^{+}00, WXK08]. dimers [BBI^{+}09, FKRE08, GYMN07, LMGO^{+}09, NL08, OKE^{+}02, Owe05, RB01, VC04, WXK08, ZOJ^{+}06]. dimethoxyethane [LCGA03]. dimethyl [GGGLL05, GWM^{+}00, WLL^{+}03, WJX^{+}08]. dimethyl-2-iodobenzoylphosphonate [GWM^{+}00]. dimethylacetylene [MTB09]. dimethylallene [ZPL07]. dimethylamino [ZH08]. dimethylcarbamate [KKH^{+}07]. dimethylhydrazone [Lu09]. dinitrogen [Ano06a, ST06]. dinuclear [SMM^{+}08, SDM02]. diodes [LFR07]. diol [Kle02, Kle03]. diol-water [Kle02]. diols [Kle02, Kle03]. dioncophylline [BMRF01]. dioxide [KT08, KZW^{+}05, KK01b, gThDjL^{+}01, ZLLS04a]. dioxide-based [KZW^{+}05]. dioxygen [BLO^{+}02, SSW^{+}07]. dipalmitoyl [CEP07]. dipeptide [BISB02, ECA06, HLMR06, JW06, KK09, LRI^{+}02, LL07, PFJ^{+}03, Qua07, WD04, YXL^{+}09]. dipeptides [LSW^{+}01, TTB01b, YXL^{+}09]. diphosphates [PFR04b]. dipolar [RI08, San01]. dipole [DVP^{+}02, EDW07, HN02, HK08a, HK08d, HK08b, KFZ03, MLA00]. dipole-quadrupole [HK08a, HK08b]. dipoles [DVRP^{+}03]. Dirac [HDBD04, PVdJB00, TW03, Vis02]. Direct [CBS^{+}03, CAG07, JJH01, LW04a, TY03, WLLS04, WDX^{+}02, YLZ08, GLMV09, HP05, mJlZsLyL07, PC00, PFC03, SLL^{+}04a, SHH07, SVT09, UKN04, WLL07a, WyLG^{+}09, WLL^{+}03, XLL^{+}02, ZWL^{+}05]. direct-particle-deletion [SHH07]. directly [SFRS01]. disaccharide [FKJ^{+}01]. disaccharides [SRB02]. DISCO [ZBS03]. DISCOtech [JFG04]. discover [LHJ^{+}06]. discovery [HS07b, KV00]. discrete [DXW08, MGLO03, QLHL09, YL06, ZBS03]. discretization [Bie04a, Bie04b, RP07b]. discriminant [ZHH09]. discriminants [FTLV01]. discriminating [yCkHmY08]. Discrimination [LDTS07, ZPL07]. discriminative [WHH^{+}06]. discussion [CDGS09]. disilenes [TKS^{+}01]. dismutase [PMM06]. dismutases [RJLR06]. Dispersion [COL^{+}06, RDM^{+}08, CLZX09, GYMN07, Gri06, JČHS07, KSS08, Lab08, Whe08]. dispersive [BCF^{+}09]. dissimilarity [hYDN^{+}08]. dissipative [YCXY03]. dissociation [CJW^{+}09, KWK^{+}01, KWK^{+}02, LS05b, TJM^{+}03, WZZ^{+}09]. dissociative [ABYM08, KKJH08]. dissolution [SBG09b]. distal [IGNH03]. Distance [MML02, RSS09, Ano05b, BL00, Cri04, IZA06, KvGH01, KH06, LI07, LHI09, PYEA03, Sha05]. distance-dependence [BL00]. distance-limited [Ano05b, Sha05]. distances [Var09]. Distorted [KS01b]. distortion [LMO09]. distributed [ASWG07, DGHR02, FOK^{+}04, IS07, KMA^{+}07, SKK^{+}07, TYO^{+}02, ZP03]. distribution [ACM^{+}06, BBP09, CFS03, CV09, CMGDAC^{+}07, JVVK09, KS01a, KS02a, LBT07, MZ05, MLG04, PP08b, SK09]. distributions [AEE^{+}03, BSP06b, Chi03, HLS07, LV08, MFB04, RLER05]. disulfide [DWS^{+}09, KF03, wQZsLyZ02]. dithiacyclohexane [FD03]. dithiane [FD03]. divalent [GS04]. diverse [AGMPRG^{+}08, AVS09]. diversity [ZHH09]. divide [AKN07, vdVGDM00, MLJ03]. divide-and-conquer [AKN07, MLJ03]. dizincocene [GXK09]. DL POLY [KSY^{+}00]. DMPC [HNL08]. DMS [RAGLL09a, RAGLL09b]. DMS-OH [RAGLL09a]. DMSO [RAGLL09a, RAGLL09b]. DNA [AB08, AZM03, BCP03, DLW06, DLWV07, EL06, EL07, FPN^{+}05, FKM^{+}06, FKM^{+}07, JMD^{+}02, JCL05, LW04b, LD05b, LXZ06, MB00, Maz01, PG04, Pin01, Pin03, PSHP08, PSS^{+}04, PSMB05, RTG00, SG07a, SHD^{+}08, WRP^{+}06, WWL^{+}09, hYDN^{+}08, YS00, ZLY07]. DNA-base [PG04]. do [SRK^{+}00, YJF06]. Dock [BS08, CWV^{+}05]. docked [NMAT01, ZWB09]. Docking [BTLP03, RGZM09, WS02b, AGI^{+}00, AGI^{+}07, AB09, BS05, BS08, CKMC04, CBC^{+}08, CLH^{+}07, CGBF05, CWV^{+}05, CR09b, GZM09, HR08, HLM05, HW09, KG02, KCL06, LR03a, LCKL05, MKT04, MM03, MCR08, MHL^{+}09, RK05, Ruv07, SBG^{+}09a, TH02, TFN04, TJE03, TP01a, Tot04, VVS07, WRBV03, YK00, Yan04, YKK09]. dodecamer [JCL05]. Does [RY09, RS07a, RS07b, WCK00]. DOIT [SFRS01]. domain [IGNH03, JS07a, OO08, PAT^{+}09, PYCD03, PYS05, PC07, PLC08]. domain-averaged [PYCD03, PYS05]. domains [GCDL^{+}05, PC05, PC07, SCS07, WCF04]. dominant [LMB08]. DommiMOE [DFWH05]. Donald [Sta00]. donation [HT05]. donor [RM07, SEKS09]. dopamine [FPG^{+}06]. dopants [CM09]. doped [JBGK08, LWLS07, SCP08, WZZ^{+}09, XWL^{+}09]. Doping [SM06, JBGK08]. dot [CLZ^{+}09]. Double [LB05, AZM03, CMJ08, DLRZ09, LMGO^{+}09, Won09, YS00]. double-stranded [AZM03]. doubles [IN08, WKYU01, dSVA^{+}09]. doubly [CHRL09, LDL^{+}09]. doubly-linked [CHRL09]. Douglas [YH09]. Douglas-Kroll-Hess [YH09]. downward [LMO09]. DP [CDGS09]. DQ2/DQ7 [KVS^{+}06]. DQ7 [KVS^{+}06]. DQ8 [KVS^{+}06]. drag [YSJ09]. Dramatic [AM06a, KT08]. dressed [MW09]. driven [MH09, PV03, SVT09, WPS02]. drives [LFS^{+}07]. driving [AM07]. Drude [LLM09]. drug [CMCB08, DHW^{+}09, HS07b, LLW^{+}09, MCR08, PPXP01, PFR04a, SPGS08]. drug-induced [CMCB08]. drugs [BLB09, KEH^{+}02, KC01a, SMM^{+}08, VGGMM05, WM01]. DsbA [CFR06]. Dual [WyLG^{+}09, WLL^{+}03, ZWL^{+}05, mJlZsLyL07, TST^{+}08, ZZL04]. Dual-level [WyLG^{+}09, WLL^{+}03, ZWL^{+}05, TST^{+}08, ZZL04]. Duan [Ano06c]. due [Car02, JM07a]. duplexes [BL08, NL08]. duration [CCSJ00]. during [IZA06]. dyad [CHRL09]. dye [BG00, KS05c]. dye-sensitized [KS05c]. Dynamic [SDCG02, XLZ08, CC07, CVR08, CEP07, FEV^{+}09, JW06, mJlZyL^{+}08, LW04a, LDG02, LEV^{+}09, QCK01, QCK02, SDL^{+}09, SCC04, SYC08, WM06, XLC08, YCS07, YCXY03]. dynamical [CKW09, EM03a, Kri09b, LDTS07, MS03, LPK07]. dynamically [CvG08]. Dynamics [BBG^{+}04, DJB02, KB09, KIM^{+}09, SSBE06, Yos02, ALB09, ATMK03, AM06b, BL09, BB05, BWE05, BRDC02, BSJ01, BG07, CLP09, CLC09, CADW03, CW02, CIB05, CCSJ00, CF06, CPC^{+}00, CCP04, CMD^{+}04, DLG00, DSS03, DBGV07, ES00, EMP07, ESM06, FSM09, FHRR07, FG02, FBDG06, FEVM01, FKZ09, FNP^{+}06, FPN^{+}05, GL04a, GL04b, Gly06, GS02, GS03, Gon07, GSDT09, GJK00, HB09, HGMB04, HSWN01, HN02, HTSR04, HW03, HTN03, HLB09, Ike04, ITS05, ITS06, IC08, JS07a, JP09, JCL05, mJlZsLyL07, JHZ09, KMH02, KŠB09, KM00, KM07, KSY^{+}00, KAK^{+}09, KZRO03, Kol04, KvGH01, KPR04, Kri08, Kri09b, Kró03, LLM08, LSG06, LPK07, LGB^{+}09, LWY^{+}09, LR03b, Loe03, LMIF06, LM03, LPB03, MB00, MFB04, MN02, MM03, MABM09, MBC08, MCR08, MDA08, MOP^{+}07]. dynamics [Maz01, MVL^{+}05, MO09, MH08b, MCM04, MST^{+}08, MS01, NK01, NBJ04, NYK^{+}09, OO04, OO06, OR05, ON07, PMGL03, PRKP05, PMB04, PBW^{+}05, Pin01, Pin03, PZS04, PPYS08, PHH^{+}08, PHRR08, PB02, PNG08, QNF09, RRZA08, RMHK03, RG08, SO07, SH09, SBJ08, SBG^{+}09a, SLL^{+}04a, SM03, STH02, Ste04, SDM02, TYN05, TLKT00, TFN04, TK08, Tru07, UTH^{+}03, VSK^{+}04, VCM01, WCF04, WLLS04, WLL07a, WyLG^{+}09, WWL^{+}09, WEE01, WD08, WDX^{+}02, WLL^{+}03, XLL^{+}02, XLZ08, YXC^{+}07, YLZ08, YGLvG06, YTH^{+}07, ZZL04, ZWL^{+}05, ZWS^{+}09, ZWZ09, ZSK07, SO07]. dynamics-quantum [ZSK07]. dynamics/order [MO09]. dynamo [FAB^{+}00]. Dyson [SVT09].

E-state [SPGS08]. E1 [YT04]. E2 [RY09]. EADock [GZM09]. early [CMCB08]. earth [JHMB^{+}09, JHMB^{+}11, SO07]. earths [LZZC09]. ECEPP [Sen06]. economic [FZL07]. edge [XWL^{+}09]. edge-doped [XWL^{+}09]. Edited [Sta00]. Editor [JW12, WM12, vLBBR12, Lip00]. editor-in-chief [Lip00]. Editorial [Bro05]. Editors [BFS07, FA01b, FBS09]. educing [BS01]. Effect [CXZ^{+}09, CN05, CEP07, KGL07, Mue01, WMW04, BB08, CPJ00, CPJ01, CGB^{+}09, CSB08, CKT^{+}08, DMJV05, GT03, HK08a, HK08b, KT08, KMM07, KCL00, Kri08, Lee09, LL01, LCDA03, Mas04, MZL08, PCS04, RY09, RR05, SOOF05, SPDS01, SCG04, SDL07, VM07, WM06, WDX^{+}02, XWL^{+}09, ZY01, ZZS^{+}07, ZWPR^{+}04, CPDZH08, HFS^{+}07, JD09, WMW03, WSC09, vE01]. Effective [OCB02, SBLK01, VBGL^{+}00, BCF^{+}09, CR09a, DPT03, HMWC03, HSWW00, LFK05, MML02, NGTB03, NG04, RPMP03, SG07b, Vas02]. effectively [SMGE08]. Effects [DXW08, KKH^{+}07, RLP08, XWXC08, AD00, ASS^{+}02, BA03, BA04a, BA04b, BPC01, BE07, BDW00, BBI^{+}09, CC07, CKF01, CDPL09, Don08, Dra00, ECA06, FGR07, FHF^{+}01, GWM08, GVATG03, GM04, GGLR00, GKTS04, HRR05, IC08, JJK^{+}00, JWB05, KSK00, LS08b, LR03b, Loe03, LFR^{+}04, MDA08, PBF09, PSF^{+}08, PWFS01, RRCA08, SF07, SL09, SMAdV00, SURG06, SM06, SN06, SRB06, VDM06, WMGK07, WD04, Wib04, XYN^{+}06, YXZ^{+}04, YH09, ZX04, ZWS^{+}02, CM09]. efficacy [KSM05]. Efficiency [IO08, MKGA06, RLER04a, YAÇ^{+}02, CN03, FSM09, GF08, KK08a, LJ04, LJS05, LKW04, PSMB05, SM08b, SM03, SE08, vLBBR12]. Efficient [AT02, BP00, BB05, CSJ01, DMZT08, EA08, FL08, FKZ09, GHH07, GB04, HMWC03, HTKG08, KMA^{+}07, KCL00, Nil09, OD09, Oos09, SATO04, SSM08, TP01b, WM12, YZ04, Ami00, BL09, BP02, BdPRMAI00, CGG06, CIB05, CY09, CY13, DBS08, FG02, GJL^{+}08, JBJB00, JJB02, KM00, LSO04, LCKL05, PRSMM02, RKA^{+}09, SAM06, SSMW09, TS05, Tot04, WW03, YXL^{+}09, vLBBR12, FS98, FS00a]. efficiently [IGL07, LR06]. eigenvalue [SSL02]. eight [CWV^{+}05]. elastic [AJ03, BED02]. Electric [LTV08, SF07, WMS06, ÁCD^{+}03, BSOB05, CM09, Kar06, LST08, Mar03, OBBS05]. electrical [KCL00]. electride [LWW^{+}06]. electrochemical [Bie04a, Bie04b, Rud05c]. electrocyclization [ZGZX07]. electrolyte [CCCJ09, YSJ09, ZCS04]. Electron [AS00, BK08, BWW^{+}08, BLN01, CFS03, FS02, Li01, OON01, PC05, PC07, TKH07, TD06, ZJM^{+}07, AEE^{+}03, ABF^{+}03, BG03, Bac07, BG00, BY06, BMLV04, BI06, BRS07, BL00, BAÅ07, CRC^{+}08, CXZ^{+}09, CR09a, CHRL09, CKT^{+}08, DVP^{+}02, DVRP^{+}03, EL09, FLOD07, FR06, FS04, FZL^{+}06, FSS00, GKRG08, GGA00, GR07, GBJ03, HLS07, Hir08, HSWW00, IKN08, ITN^{+}05, IS07, IN01, Jac09, KS05a, KK08c, Kle03, KFD06, Kri09a, KKMMS04, Lai07, LMV07, Lee09, Leh06, LV08, LL01, LLXS02, LH02, LB05, LBG08, MLL06, MGCA07, MGMM07a, MS00, MA09, OKE^{+}02, PFB05, PA05, PAS07, QZZZ03, RS07a, RS07b, RTG00, SFC04, VC04, VKCK09, WMRW^{+}01, Wan09, WL00, XWXC08, Yas08, YCXY03, YH09, ZXL^{+}04, PC05]. electron-correlation [YH09]. electron-pair [FS04]. electron-repulsion [Kri09a]. electron-sharing [BRS07]. electron-transfer [QZZZ03]. electron/four [GYMN07]. electronegativities [dSGCG00]. electronegativity [ALC08, BCNs07, JVVK09, VK06]. Electronic [CWWS07, DHM^{+}03, FLK^{+}07, GGGLL05, Hua09a, IME02, KHY00, KMM07, KGD06, LPP06, LTF^{+}07, OS08, QB05, RPNJ07, SCP08, Wu06, WWS07, XZ05, AJ03, AEE^{+}03, Ama02a, AZS^{+}04, AGSFA^{+}05, ASS^{+}02, Bac09, BBG^{+}04, CMàGL^{+}04, CZFH07, CN05, CNN07, CAG07, CRSB03, CSV^{+}07, CTFC08, DD00, FL08, GJL^{+}08, GBL^{+}05, GM04, HMMS09, HZ09, Hua09b, Kar01, KRM^{+}02, KJP^{+}07, KIFK07, KWK^{+}00, Kri09a, KPZK06, LWK08, LWLS07, LWZ09, LFR07, LB08, LDL^{+}09, LMRVFH^{+}09, MKGA06, MM02, NYH02, PP08b, PMC^{+}08, QCK01, QCK02, RRCA08, RS05, SSB^{+}03, ŠBL05, TD08, TT01, TD06, TDK07, WMRW^{+}01, WLX^{+}05, WL00, YXZ^{+}04, YFR05, YS00, ZZY07, ZZS^{+}07, ZZY08, ZXYF09, ZX08, ZL05, ZL07, ZL09b, ZM03, SMKM00]. electrons [HIM07, Pog03, WJ00]. electroosmotic [YSJ09]. electrophilicity [RUPH06]. electrophoresis [WWL^{+}09]. Electrostatic [CCT^{+}03, GYMN07, PK05, PML03, RLER04b, SG07b, ABWT09, BCNs07, CPUGD09, CHMI05, DWNB01, FOL^{+}04, GY06, GPK05, GBJ03, KFZ03, KLH^{+}04, KCK^{+}08, MMPK01, PMB04, PP08b, RLP08, SMAdV00, SHSF05, VGDSU08, VC04, WCK00, YH06]. electrostatics [HS01, MLJ03, RSP03, STSF02, VVBV02, WMS06, ZFW08]. electrotopological [SPGS08]. element [Ara04, BHW00, BK08, BH03, BF04, HBW00, HBW01, Li01, SRB06]. elements [ABWT09, ATM^{+}07, ASS^{+}02, Ell07, JGH00, RP07b, RRS09, VB03, WL04, vW06]. elevated [TK08]. ELF [SFC04, CFS03, FSS00, PC05, PC07]. ELI [BWW^{+}08]. ELIA [BWW^{+}08]. elimination [Adc04, CFD03, GLD08, GS08, KUB07, Mui05, PSDM00, YFS07]. Elongation [KLM^{+}09, MKGA06]. Elucidating [DBS07]. elucidation [GZ07, GLH^{+}08]. embedded [CEP07, GGLR00]. embedding [Agr03, JNV08, KS02b]. emission [MLCD01, RGG08]. emitting [LFR07]. Empirical [CBC^{+}08, LS08c, Mac04, SP05, CYM02, FM00, Gri04, GGK^{+}08, HRBKB03, JČHS07, KK08b, LR03a, LLM09, MB00, RKH03, ZNLL07, VBGL^{+}00]. employing [MHT01, THHN01]. empty [CZA03]. enantiomerization [Qua07]. encapsulated [WL09b]. enclose [ZBS03]. Encyclopedia [vRS98, Lip00]. end [Adc04, GLD08, KUB07, PSDM00, YFS07]. endohedral [KSN01]. endoperoxide [BLO^{+}02, CG08]. enediyne [KC01a]. Energetic [DRAS04, DRAS05, JW12, RP09, BCP03, ECM^{+}03, JD09, KCK^{+}08, PBF07, SLHW09]. Energetics [KRLD09, AHGK09, DBGV07, Hua09b, ILKR09, LD05a, MFB04, Mas01b, MOP^{+}07, SDCG02, WSC09, ZXY03, ZX09, DLG00]. Energies [CRSB03, BP02, BWE05, BLL^{+}06, BE07, BDW00, CHA^{+}07, CCK01, CPML08a, CG05, DB07, DMZT08, FOL^{+}04, FJP07, FKU^{+}05, GS04, GKTS04, HT05, HYA02, Hol05, IO08, JJK^{+}00, JCL05, JZD^{+}09, JG03, JSHG07, JČHS07, KWK^{+}01, KWK^{+}02, KSTC01, KC01b, LN01, LSW^{+}01, LFEdL06, MT03, Mas04, MLL^{+}08b, NHH05, PK05, PSF^{+}08, RSSKB03, RSE07, RSN^{+}02, RM00, SZT08, STSF02, SYC03, SA07, SLRC01, TKS^{+}01, UBDPJ04, Var09, VC04, WCK00, WLX^{+}05, WS05b, WW04, XLT07, ZMZ09, ZW09, ZM03, vGGB00]. Energy [CBH^{+}03, HFHL06, IN08, KLS02, MSF^{+}08, NK06, SSB07, WM12, dSR08, AMR04, AJ03, ABÅ04, AE06, AM06b, ABBC01a, ABBC01b, AGSFAL05, BM07, BCIB05, Ber03, BL05, Blo04, Bof01, BRLS08, BRLS12, BACJCT01, BF07, BLB09, CC09, CN03, CCB04, CY09, CY13, CJW^{+}09, Chi03, yCkHmY08, CV09, CMGDAC^{+}07, CA04, COL^{+}06, DLD^{+}02, DLRZ09, DMJV05, DK01, EGSG00, FSM09, FK07a, FKJ^{+}01, FZL^{+}06, GZL02, GMA04, GLMV09, GAdGM08, GG09, GC04, GS02, GS03, GPSP06, GB02, GWS^{+}02, HKMS01, HP01, HR08, HMOG07, IGNH03, ILKR09, IGL07, IPN06, IPN07, Jac09, JMD^{+}02, KGN07, KN04, KKC05, KUB07, KB09, Kob03, KC01b, Kri09a, Lab08, LR03a, LMK01, LF04, LFZS04, LJ04, LBG08, LKW04, Lu09, MG06, MCF05, MAF^{+}07, Maz08, MH09, MGJAARC00, MGLO03, MRS^{+}07, Nak02, NKIS02, NA06]. energy [OD09, OFB08, ONHN00, OKH^{+}02, OV03, Oos09, PSC^{+}01, PMGL03, PK04, PAT^{+}09, PMPGP05, Qua01, RP07a, Rao00b, Rap06, RSE07, RRCA08, RWBH09, RHL09, SOOF05, SPDS01, SKGS00, SPL^{+}02, Sch03, SMGE08, Sen06, SRCD03, SSM08, SY09, SG07b, SSMW09, SMD02, SJW09, SSBE06, TJE03, TGGP^{+}00, TCR^{+}02, UTH^{+}03, VE09, VM02, Vya01, WL02, WD04, Whe08, WHF08, XZZ04, YXC^{+}07, YZ04, YHD^{+}06, ZCZ03, ZZ08, ZGXX06, vEMK01, vLBBR12, Hir08]. Energy-based [KLS02, MSF^{+}08]. energy-consistent [SMD02]. energy-transfer [MAF^{+}07]. energy/one [Oos09]. energy/one-step [Oos09]. enes [PWFS01]. engine [MVL^{+}05]. Enhanced [KG02, DAK08, NYTH09]. enhancement [AB08]. enhancements [AM06a]. enkephalin [ZCL09]. enones [SLRC01]. enough [VGGMM05]. Ensemble [Blo04, BSÇ^{+}01, EMP07, OO06, SM08b, SM03]. ensembles [GLD08, Ike04]. ensure [FKFG08]. enterovirus [KCL06]. Enthalpies [EB04, WC04, BE06, LS05b, RM00, TTBM09, VGGMM05]. enthalpy [OVMV04]. entire [ZAT07]. entropic [CBC^{+}08, FGR07]. entropy [DHF^{+}05, HDF^{+}07, HTKG08, KKH^{+}07, LM03, RK05, Ruv07, STSF02, WG02]. enumeration [AL01]. Enveloping [BHH^{+}09, CV09]. environment [DFWH05, DPM09, GT03, HFS^{+}07, PMM05, GBL^{+}05]. environmental [CMGDAC^{+}07, FGR07, MDA08, TP01a, VW03]. environmentally [EDAJ04]. environments [MPF00, ZFW08]. enzymatic [PCMG09]. enzyme [CFER04, Fie02, GGLR00, GS04, MDA08, Pin03, TDH06, VB09, ZL09a]. enzymes [BS06, CPUGD09]. enzymic [CG05, TCR^{+}02]. Epimerization [BBSS06]. Epitope [SHBD05]. epitopes [KVS^{+}06, SHBD05]. epoxide [Owe05]. epoxides [OY01]. epoxy [OY03]. epoxy-amine [OY03]. EPR [SN06]. equalization [BCNs07, JVVK09, VK06]. equation [AMR04, ABWT09, BHW00, BH03, BF04, BF07, BRS00, BRS01, FS00b, Höf05, HBW00, HBW01, SATO04, TW03, Vas02, Vis02, Zho06]. equations [Bie04a, CF04, DOSG06, Har04, KvGH01, LPK07, LMJ02, QNF09, RI07, Rud05a, Rud05b, Rud05c]. equilibrated [CA07a]. Equilibration [SDCG02]. Equilibration-Morse [SDCG02]. equilibria [FGR07]. equilibrial [Kli01]. Equilibrium [KSTC01, BBP09, ECA06, KBLP09, LS08b, MMLC05, PAS08, ST04, WMGK07]. EQUIPATH [Kli01]. Equivalent [ZZY07, WBSR03, ZZY08]. equivalents [RCJ02a, RM00]. ERE [MCF07]. Erratum [ABBC01a, Ano05b, Ano06a, Ano06b, Ano06c, BA04a, BRLS12, CY13, FS00a, HNWF12, HBW01, HK08a, JHMB^{+}11, KWK^{+}02, LFSB03a, LR03b, NBTN04a, PCO^{+}07a, PWHF^{+}04, RS07a, TZX01b, WB04a, WB05, WMW04, WWC^{+}05]. error [IO08, KMA^{+}07, Kob03, Mas04, MDI04, RS05, Rud05a, Rud05b, Rud05c, TBSM09, VKP^{+}08]. error-ranked [TBSM09]. errors [CS03]. erythronic [vDSSvA04]. ESFF [SYY^{+}03]. Essay [BHTCG07, FK07b, GR07, Kut07, MGCA07, Nye07, Sha07, Sim07, Tru07]. established [SB01]. ester [TH02]. esters [POJ01]. estimate [KC01b, YZ04]. estimated [ZMZ09]. estimates [GC04, HT05, MDI04, SY09]. estimating [HDF^{+}07]. Estimation [DHF^{+}05, ZW09, CV09, DDVD09, HLTLP09, KC01b, PYEA03, Lab08]. estimators [GZL02]. estrogen [FKU^{+}05, KBK^{+}01]. ethane [DGD^{+}05]. ethanes [WyLG^{+}09]. ethene [Ang09]. ether [GLRL02, WD04, WLL^{+}03, YLW^{+}08]. ethers [ACLD03, LCDA03, LCGA03, LCA03, ZWY^{+}09]. ethyl [KKH^{+}07]. Ethylene [TBG00, BSJ01, Hir08, NTH00, SBB02]. Euclidean [RRS09].

EUDOC [PPXP01]. Euler [SG01]. evaluate [GGB07a, GGB07b, LF04, OSHS03, TSSGS07]. evaluated [ABBC01a, ABBC01b, Bof01, Qua01]. evaluating [FO08, Con02]. Evaluation [BMLV04, DR07, KSM05, NMAT01, OYH09, VKP^{+}08, YSJ09, Ano05b, AGSFAL05, CGG06, CAG07, DSR^{+}07, DBS08, ESP04, FMAMVK06, FKZ09, GGA00, HMWC03, JSHG07, KJVW08, KH06, LMV07, LYS08, MSH^{+}06a, Mor02, PRS04, Sha05, VP02, WL02, Yan04, Yas08, CBC^{+}08, GKTS04, OGH05, ZSK07]. evaluations [SF05]. Evans [SRK^{+}00]. EVEBAT [CZA03]. Even [CVVB04, CC07, VVBV02]. Even-tempered [CVVB04]. evidence [BLO^{+}02, IO08, SFR07]. Evolution [SPL^{+}02, Der09, Mck07a]. Evolutionary [DPRR05, CS02, CKR08, yCkHmY08, DB06, KH05, VGO^{+}07, YK00, Yan04]. Ewald [Ami00, BYQS03, KM00, KSY^{+}00]. Exact [GC02, GHMP03, DLSVY00, TRS02]. examination [CZA03, LJKL08, QNF09, WL04, ABÅ04]. example [AS00, JHPRSM^{+}05, MGLL03]. examples [CMA^{+}08, HBW00, HBW01, SDL^{+}09]. exchange [AAP00, AKN07, BWI^{+}02, EL09, FSM09, FGR07, FDM00, GLP08, LMIF06, RRFC^{+}03, SM08a, SM08b, WL04, Wei08, NCO^{+}05]. exchange-correlation [AAP00, EL09, FDM00, WL04]. Excitation [HKHN08, BMB07, Che01, CG08, Hol05, LWZ09, LFEdL06, ON07, PSF^{+}08, SZT08, SA07, SLRC01, XZ04, ZM03, vGGB00]. excitations [DHM^{+}03, ION07, MA09, TJM^{+}03, XZ05, ZX04]. Excited [CHA^{+}07, HFS^{+}07, Ang09, FCW06, FDSA00, HNWF07, HNWF12, IR03, LWX07, LDL^{+}09, MW09, NBTN04a, NBTN04b, NTH09, PO03, PSS^{+}04, SBI08, SMKM00, TY03, TKN^{+}08, WLZ^{+}07, ZH08]. excited-state [LDL^{+}09, NTH09, PSS^{+}04]. exclusively [RI08]. exercise [FLK^{+}07]. exist [RY09]. existence [WPH^{+}07]. expanding [Bac07, Bie04a, IZA06]. expansion [AHGK09, GS09, GKH05, HTKG08, II02, Ish02, LZZC09, SvDS01, WBSR03, ZFL^{+}05]. expansions [Bou01, DWNB01, GC02, JSHG07, RLER04b, SG01]. experiment [BE09, GBJ03, LS05b, Mat03]. experimental [BE06, JARM02, LEK07]. experimentally [KBN02, TDH06]. experiments [CVR08, HP05, OD09, PC00, PFC03, SL04]. Explicit [EC06, PPYS08, RI07, AL01, DMJV05, FC06, HM02, JZD^{+}09, KIFK07, KIM^{+}09, LSO04, PK05, RKA^{+}09, WB04a, WB04b, WB05, ZGFL01]. explicit/implicit [LSO04]. exploiting [JSHG07]. exploration [CSJ01, HLB09, LXW^{+}09, LMO09]. exploratory [PGH^{+}04]. explore [ILKR09]. Explorer [SHBD05]. Exploring [BL05, HPP00, HXLS09, KF08, Sch03, Tie09, SPL^{+}02]. exponent [WTKM06]. exponential [Rud05a, Rud05b, Rud05c]. exponentially [Bie04a]. exponents [MY08b]. exposure [MML02]. expression [dGWH01]. expressions [TNS00]. Extended [LMH^{+}09, TVL^{+}03, Bie04b, Cul04, DXW08, KUB07, QNF09, SS00, ST01].

Extending [GCD04, MFB04]. extensible [SYY^{+}03, GBL^{+}05]. Extension [CR09b, FBLO08, GY08, TBGRJ04]. Extensive [JW12, LB08, SLHW09, YXL^{+}09, ZL05, SMG09]. external [CM09, EC06]. extra [LW07]. extra-valence [LW07]. Extracting [HM02]. extraction [OD09]. extrapolated [KSTC01, Var09]. extrapolation [MO09, MC06, PSC^{+}01, PFJ^{+}03]. extremal [ZZ08]. extremely [GFS05]. Eyring [Nye07].

F [CRC^{+}08, FO04, Gog08, HYA02, HZ09, Hua09a, IV04, KS05a, KBL08, Mar03, RB01, STC^{+}08, UTT^{+}04, WLLS04, WLL07a, XLL^{+}02, ZY01, ZLLS06b, ZL09b, HK07, KS05a, RFSS06, SOOF05, Sha02, WDWS06, YWHZ03]. facility [SWZS04]. factor [LMCD09, WL00, XSHC06]. factorization [EC06]. factors [AST06, SBH02, TP01b]. FACTS [HC08]. family [CFS^{+}09, DMC05, NAT07, WTKM06]. FapydG [SHD^{+}08]. farnesyl [SFR07]. Fast [JBJB00, JJB02, JSHG07, NG04, RS08, SYC03, SFC04, Ami00, ATMK03, Ano05b, CS02, CRG01, CHMI05, CZA03, GY08, GKK07, HH04, HLM05, Ish04, KM00, KLM^{+}09, KvGH01, KH06, PZS04, RK05, San01, SCC04, Sha05, TRS02, VP02, WCC08, HC08, LZ05a, VLH^{+}05]. faster [SF05, AM06a]. faujasite [TLOG00]. faujasite-type [TLOG00]. FB [DHW^{+}09]. FB-QSAR [DHW^{+}09]. FBP28WW [PAT^{+}09]. FDS [TJE03]. Fe [BTP09, HLLN06, HYR06, KRLD09, AGK03, KT08, DF04, Mck07a, Mck07b, NyHN06, NHN06, PLC08]. feasibility [MWE02]. feed [SJJ^{+}04]. feed-forward [SJJ^{+}04]. feedforward [LJZ^{+}07]. Felix [Ano06a, Ano06a]. FeMo [Mck07a, Mck07b]. Fermi [Kri09a, PYCD03, PYS05, PC07, PLC08]. ferrocene [Kan07, MBP09, ZZS^{+}07]. ferromagnet [TD08]. ferromagnetism [SK08]. FeS [Mck07a]. Feynman [RLER07]. Field [MO01, AS06, ACLD03, Ano06b, Ano06c, ATBLS04, BWI^{+}02, CLP09, CLWL09, CPM03, CM09, CGB03, CLA^{+}00, CR02, CSU05, DvG00, DPT03, DRMD03, DFWH05, DMLI05, DGI^{+}08, DHW^{+}08, DWC^{+}03, EBD^{+}01, FHRR07, FBDG06, FAR02, FM00, GZL02, GMA04, GRO^{+}03, GGK^{+}08, HP01, HGMB04, HXLS09, HIM07, HNL08, HMOG07, IDMC09, IT03, IKYM09, JS07b, JCL05, JM07b, JFG04, KB02, KSB^{+}02, KS06, Kar06, KFNH08, KTA03, KOFF09, KLB03, KYT^{+}08, KOML08, KDSV02, KVL^{+}04, KBN02, LL00, LST08, LTV08, LFZS04, LAT05, LH05, LLM09, MT03, MB00, MM05, MP03b, MBC08, MMMY07, MSR04, MRC03, MHJS06, NUH02, NCO^{+}05, OYH05, OMNH08, ONHN00, OKH^{+}02, OVMV04, OBT09, PB04, PMB04, PS09b, PWHF^{+}03, PWHF^{+}04, Pom04, PHH^{+}08, POJ01, PB05, RSN^{+}02, RKH03, SF07, SO09, SDL^{+}09, SDvG01, SAS05]. field [SDCG02, SSS^{+}09, SYY^{+}03, SHK^{+}05, SP05, SK05, TAS07, TTB01a, VSW^{+}03, VCM01, VTT^{+}08, WK01, WWC^{+}04, WWC^{+}05, WZW^{+}06, WMS06, Wil01a, Wil01b, XLT07, YCXY03, ZWC^{+}09, vDSSvA04]. field-based [DMLI05]. field-derived [WMS06]. field-induced [CGB03]. fields [ABÅ04, Car02, EBD^{+}01, HRBKB03, LLM08, Mac04, MFB04, OSHS03, PK04, PB02, RP07a, RLER04b, RG08, SL09]. files [FJP07]. fill [RRZA08]. find [HQ02, WS07]. Finding [BS01, Qua07, GF08, Rao00b]. fine [VSK^{+}04]. fine-grained [VSK^{+}04]. fingerprints [LHJ^{+}06]. Finite [Ell07, MO01, AB09, ALKH04, BHW00, BP01, Bie04a, Bie04b, BF04, DRMD03, Der09, GM04, HBW00, HBW01, KGD06, PZS04, RP07b, Rud05a, Rud05b, Rud05c, VZM^{+}08]. finite-chain [Der09]. finite-difference [Bie04a, Bie04b, Rud05a, Rud05b, Rud05c, VZM^{+}08]. finite-temperature [KGD06]. fire [LDC^{+}07]. First [CS01, HZX04, Hua09b, TK08, WZZ^{+}09, WD08, ZDS^{+}05, ZXYF09, ZHMW09, AD00, BP03, CJK^{+}02, EBL^{+}08, FO08, GJL^{+}08, GD09, JPCA08, KK08c, LWZ09, LK03, Mck07b, MLJ03, Rud05b, VP08, WLX^{+}05, XWL^{+}09, KSB^{+}02]. first-order [Rud05b]. First-principle [ZDS^{+}05, GJL^{+}08]. First-principles [CS01, HZX04, Hua09b, TK08, WZZ^{+}09, WD08, ZXYF09, EBL^{+}08, GD09, WLX^{+}05]. first-row [AD00, BP03, LK03]. fit [BCNs07, SY09]. fitted [YOB^{+}08]. Fitting [KC01b, MCF05, Wei08]. five [SBH02, Van02a]. five-coordinated [SBH02]. fixed [HM06]. fixed-composition [HM06]. Flex [GCD04]. flexibility [BL08, BCP04, KG02, KTA03, MHL^{+}09, OV03]. Flexible [COS01, NGTB03, YK00, AGI^{+}00, AGI^{+}07, AJ03, AHGK09, BZP09, BTLP03, BS08, CCL06, CKMC04, CLH^{+}07, DDKV07, GCD04, HW09, JNV08, KOFF09, MH08b, SSBE06, TFN04, TP01a, Tot04, VLH^{+}05, vEMK01, vE01, TJE03]. flexible-backbone [AHGK09]. flexible-ligand [HW09]. flexible-protein [HW09]. Flooding [LSG06]. Fluctuating [OR05, KMH02, PB04, PMB04, Yos02]. Fluctuation [MHW04, PC05, SBLK01]. fluctuations [AZS^{+}04, WMGK07]. fluid [BCIB05, CLC09]. fluorene [CHA^{+}07, YFR05]. fluorene-pyridine [CHA^{+}07]. fluorene/carbazole [YFR05]. fluorescence [CHA^{+}07, MAF^{+}07]. fluorescent [DHM^{+}03, NAT07, VSW^{+}03, XZ05]. fluoride [BSG07, IV04]. fluorides [KS05a]. fluorinated [CUSS03]. fluorobenzene [ZTP^{+}08]. fluorobutanal [NSB08]. fluorocarbons [JARM02]. fluoroglycine [HS00]. fluoromethylene [ZLLS04a]. flux [DAK08, RKA^{+}09, Rud05a, Rud05b, Rud05c]. fly [KMA^{+}07]. FMO [FOK^{+}04, FKL^{+}06, KIM^{+}09]. FO [Gog08]. focal [KK08a]. Fock [RRS07, TW03, WMW04, AKN07, Bou00, Cul04, DD00, GAdGM08, HDBD04, MS00, MBWP03, PFJ^{+}03, PVdJB00, TYO^{+}02, UIHN09, WMW03, Wei08, YH07, vDSSvA04]. Fock/Kohn [RRS07]. Focus [Mat03]. focusing [KBK^{+}01]. fold [DB06, ZM06]. folded [CP08, GB04]. Folding [HEP^{+}02, ADM^{+}06, CCC03, DvG00, HG08, IM06, JS07a, JIK09, KH05, MLG04, MH09, Mei02, MWE02, RSER09, RLP08, VW00, VW04, VGO^{+}07, ZP03, dSR08]. folds [BS01, ZS04]. following [DLD^{+}02, LMO09]. For-Gly-NH [PC00]. For-L-Ala-NH [PC00]. Force [CLP09, JCL05, OMNH08, OBT09, SO09, SL09, ZWC^{+}09, AM07, AS06, ACLD03, ABÅ04, Ano06b, Ano06c, ATBLS04, CLWL09, Car02, CPM03, CLA^{+}00, CR02, CSU05, DvG00, DGI^{+}08, DWC^{+}03, EBD^{+}01, FBDG06, FAR02, FM00, GRO^{+}03, GGK^{+}08, HP01, HGMB04, HXLS09, HRBKB03, HFSD03, HNL08, HMOG07, IDMC09, IT03, IKYM09, JS07b, JM07b, KB02, KSB^{+}02, KS06, KFNH08, KTA03, KOFF09, KLB03, KYT^{+}08, Kle03, KOML08, KDSV02, KVL^{+}04, KBN02, LLM08, LL00, LMGO^{+}09, LHI09, LAT05, LH05, LLM09, MT03, MB00, MM05, Mac04, MFB04, MMLC05, MBC08, MMMY07, MSR04, MRC03, MHJS06, NCO^{+}05, NMAT01, OYH05, OSHS03, ONHN00, OKH^{+}02, OVMV04, PB04, PMB04, PK04, PS09b, PHH^{+}08, POJ01, PB02, PB05, RP07a, RNG03, RI07, RG08, RKH03, SDL^{+}09, SDvG01, SAS05, SDCG02, SF05, SSS^{+}09]. force [SYY^{+}03, SHK^{+}05, SP05, SMM^{+}08, SK05, TAS07, TTB01a, VSW^{+}03, VCM01, VTT^{+}08, WK01, WWC^{+}04, WWC^{+}05, WZW^{+}06, Wil01a, Wil01b, XLT07]. Force-field [CLP09, OBT09, SO09, HGMB04, IDMC09, KLB03, MBC08, NCO^{+}05, OKH^{+}02, OVMV04, SP05, VCM01]. forced [CAG07, LPB03]. forcefield [Adc04]. forces [BCF^{+}09, BH03, HNWF07, HNWF12, JS07a, LPB03, PK05, RLP08, WB04a, WB04b, WB05]. Foreword [DF08, Fre00, FJ02, FH06, FS07, Gad03]. form [AT02, Bac07, BRS01, CR02, LC07]. formaldehyde [WCL05]. formalism [AS00, FLGW00, YCXY03]. formalisms [CF06]. formamide [IINK09, Pac06]. formamidine [WJX^{+}08]. format [TDK07]. Formate [ČJPZS08, NK06]. Formate-Lyase [ČJPZS08]. formates [CUSS03]. Formation [JM07a, RAGLL09a, RAGLL09b, BE06, BMTFR08, CS03, EB04, HIA03, JWB05, Kle02, LLA01c, LYZ^{+}08, Nee03, RCJ02a, RM00, TT08, TTBM09, WC04, WX09, ZZW09, dOMSL01, JKM08]. formed [LLW02, LSW^{+}01]. formic [Pac06]. forming [PP08a]. forms [SPT07]. formulas [Ish02, Tor02]. formulation [BF07, Cul08, PK05]. formyl [GSB09, HJCP01, PFC03]. formylglycinamide [HRBKB03]. forward [KM07, SJJ^{+}04]. Four [SH02, FJ08, GPSP06, Lai07, PVdJB00, PV03, SBH02]. four- [Lai07, SBH02]. four-centers [GYMN07]. Four-component [SH02, GPSP06, PVdJB00, PV03]. four-index [PVdJB00]. four-membered [FJ08]. Fourier [BWP07, CGG06, HLM05, TYN05]. fourth [Bie04a, Rud05a, Rud05b, Rud05c]. fourth-order [Bie04a, Rud05a, Rud05b, Rud05c]. FPT [BPC01]. fractal [Pan07, TT08, XOW^{+}00]. fractional [MGLO03, SM08b]. fragilis [SDM02]. Fragment [DHW^{+}09, CFK08, DPM09, FOK^{+}04, FKL^{+}06, FII^{+}07, FKU^{+}05, FKM^{+}06, FKM^{+}07, IIK09, KIFK07, MLG04, MLL08a, NYK^{+}09, NGTB03, NG04, OO08, SG07b, ZMZ09, KIM^{+}09]. Fragment-based [DHW^{+}09]. fragmental [CSB^{+}03]. fragmentation [Gor01]. fragments [AM09, DWNB01, DPRR05, KS01a, LV08, NG04, PBF07]. Framework [JGVF05, CR08, EA08, FS04, TAS07, Tie09]. Framework-based [JGVF05]. Franck [Ama02a, LMCD09, TP01b]. Free [DLRZ09, GS03, JMD^{+}02, MH09, PMGL03, YXC^{+}07, AM06b, BG00, BWE05, BLL^{+}06, BCIB05, Blo04, CN03, CM09, CY09, CY13, Chi03, CV09, CCK01, CMGDAC^{+}07, CG05, COL^{+}06, DMJV05, FSM09, GZL02, GMA04, GLMV09, GG09, GC04, GS02, Gra07, GWS^{+}02, HKMS01, HR08, HMOG07, ILKR09, IGL07, JZD^{+}09, KDG^{+}09, KAS^{+}07, KKC05, KUB07, KB09, Kob03, KK01a, Lab08, LR03a, LF04, LSW^{+}01, LKW04, MG06, MT03, MGLO03, MRS^{+}07, OD09, ONHN00, OKH^{+}02, OV03, OVMV04, Oos09, PK04, PAT^{+}09, RSE07, RWBH09, SOOF05, SAM06, SKGS00, STSF02, ŠBL05, SSM08, SY09, UBDPJ04, UTH^{+}03, VLH^{+}05, VE09, VGDSU08, VM02, WHF08, XLT07, YZ04, ZMZ09]. free-base [ŠBL05]. Free-energy [JMD^{+}02, AM06b, CY09, GMA04, ONHN00, RWBH09, SKGS00]. Free-energy-driven [MH09]. freedom [DHF^{+}05, MZL08]. freeze [BME05]. freeze-and-cut [BME05]. frequencies [BRV^{+}07, Han01, Kle03, KBN02, LMB08, PZWG^{+}04, WM04, ZWPR^{+}04]. frequency [DF06, DR09, MY08a, VSW^{+}03, YH07]. frequency-dependent [MY08a, YH07]. friction [JS07a]. frozen [AEE^{+}03, GWS^{+}02, JNV08]. frozen-density [JNV08]. fructose [MRS^{+}07]. fuel [CCCJ09]. Fujitsu [KSY^{+}00]. Fukui [DVP^{+}02, FS04, TSSGS07]. fulfillment [RLER07]. Full [PRSMV08, GD09, IR03, KGD06, RS08, ZCZ03]. full-atom [RS08]. fullerene [CHRL09, CTFC08, GYCZ04, GXK09, Kan07]. fullerene-dizincocene [GXK09]. fullerenes [GZ07, GM01]. Fully [GWM^{+}00, XZZ04, WTKM06]. function [Bac04, BS05, BdPRMAI00, CFS^{+}09, Che01, yCkHmY08, CPUGD09, Con02, DMZT08, DP03, FSS00, GCB03, GS09, GdAcV^{+}07, GPN01, HMWC03, HZ06a, HZ06b, ILB03, ILKR09, Ish04, KK08c, Kni00, KFD06, LR03a, LHI09, LBT07, MP03a, MML02, MY08b, Nak07, NKIS02, PP08b, PA05, PAS07, SFC04, SJW09, TLKT00, TW03, TJE03, TT05, TSSGS07, VVS07, YLL^{+}09]. functional [AAP00, AB00, ABYM08, ASY01, Bac09, BP01, BE09, CLP^{+}05, CFK08, CRS05, CR08, CSB08, CPML08a, Cul04, DVRP^{+}03, ECM^{+}03, EL09, FCW06, FZL07, FG02, GHLK^{+}02, GM01, GLRL02, Gri04, Gri06, GBBH09, Haf08, Han01, HNWF07, HNWF12, Hol05, ION07, IB04, ITN^{+}05, IS07, JFG04, JČHS07, KGL07, KSS08, KWK^{+}01, KWK^{+}02, KZW^{+}05, Kri09a, LRI^{+}02, LWK08, LMB08, LMGR05, LLS03, LWH06, LKT04, LF02, LLZL09, LDL^{+}09, LZF^{+}09, MW09, MSBS01, MW00, NTH09, NAT07, OKE^{+}02, PSF^{+}08, PU09, PDS01, QZZZ03, QZL^{+}04, RB01, RK04, RDM^{+}08, RR05, RZWS07, SH07, SZT08, SPT^{+}03, SPT07, SLRC01, SSB07, SW06, TBG00, TST^{+}08, TKN^{+}08, TKH03, Van02b, VMA03, VL00, VBS09, WRP^{+}06, WB07, WZY04, WMRW^{+}01, WL02, WCW08, WCHW09, WM04, WSC09, WCL05, WZXY07, WM01, XYN^{+}06, XB08, XWC09, XL02]. functional [XPW09, YYW07, YLL^{+}09, ZZL04, ZH08, Zho06, ZM03, vGGB00]. functional/continuum [LRI^{+}02]. functionals [BP03, DF04, Han01, ION07, JPCA08, KRM^{+}02, KS01b, PJPJdPRMI07, SCF^{+}09, WL04, YTH01]. functions [AE06, Bac07, Bou00, CGB03, CGSdST06, DVP^{+}02, GFS05, GLD08, GS07, GBJ03, IT03, MLL06, MY08a, NUH02, OFB08, PFB05, RHL09, Ruv07, SS00, TS05, TD06, WG02, YH06, ZM03]. Fundamental [LMB08]. fungal [LPP06]. funnels [HEP^{+}02]. furfural [COMR^{+}04]. Further [GPK05]. fusion [CRGN07]. fuzzy [ALTB06, EKB02a, RLA01]. FVII [PDP02].

G [AGI^{+}07, Kró03, XWC09]. G** [Wib04]. G-protein [XWC09]. G. [CSD05, Sim07]. G2 [RY09, ZKZ^{+}07]. GA [HSMT04, HSMT04, LLL^{+}08]. GA-MLR [HSMT04, LLL^{+}08]. gadolinium [AB00]. gain [HP05]. gains [NYTH09]. galabiose [RSSKB03]. galactose [RSSKB03]. GAMESS [UKNS01, UKN04]. gamma [Ish04]. gap [KUB07]. gap-free [KUB07]. Gas [BAL^{+}01, POJ01, CPJ00, CPJ01, DR09, EGSG00, JJK^{+}00, JHZ09, KSB^{+}02, KT08, KKH^{+}07, LRI^{+}02, Lee09, LZA02, LB05, LXSF08, MFB04, Mas01a, Mas01b, MM02, Pan07, PV07, wQZsLyZ02, RRS06, ROG00, SMGE08, STSF02, SMKM00, SK05, TDH06, UCT^{+}03, UNM^{+}01, WD04, XKKL03, XKG^{+}05, YQQH09]. Gas-phase [BAL^{+}01, POJ01, JHZ09, KKH^{+}07, Lee09, MFB04, wQZsLyZ02, TDH06, UCT^{+}03, XKKL03, YQQH09]. Gaseous [WDWS06, PG01]. gases [SRB06]. gauge [Ish03]. gauge-including [Ish03]. Gauss [DBS08]. Gaussian [TdMSD^{+}08, CMJ08, Duk01, EDW07, GC02, HdMdS05, HdS06, HD06, IO08, Leh06, Lu09, MV06, MY08a, RC04, TW03, WTKM06, YJF06]. Gaussian-2 [RC04]. Gaussian-4 [Lu09]. Gaussian-type [Leh06, MY08a, TW03]. GAUSSIAN94 [Kli01]. GB [GC04, WHF08, GWS^{+}02, YJF06]. GB/SA [GWS^{+}02]. GBR [FPG^{+}06]. GDDI [FOK^{+}04]. Ge [LLXS02, WDXS06, CJS^{+}03, LLXS02]. GeD [WDX^{+}02]. GeH [LLXS02]. gemcitabine [PFR04a]. geminal [TT05]. geminals [TT01]. general [AM07, BRS01, DGI^{+}08, EBD^{+}01, JM07b, KSU03, KBT03, NUH02, RG08, SAM06, TZX01b, TZX01a, WWC^{+}04, WWC^{+}05, XYN^{+}06, BRS07, EA08, FLGW00]. generalizable [KYT^{+}08]. Generalized [ADM^{+}06, ILB03, AB08, BC06, Bud07, CPJ00, CF06, Cul08, DLG00, FOL^{+}04, FL08, FC06, GZL02, ION07, Lab08, LFSB03a, LFSB03b, MTE04, MCM04, OCB02, SHH07, Tot04, XL02, YH07, YJF06, ZGFL01, dGWH01, BPCD07, FOK^{+}04]. generate [BWI^{+}02, CA07a, BAH^{+}02]. generated [Kri09a, LAR^{+}03]. generating [AMR04, CA04]. Generation [RLER07, BSOB05, BAH^{+}02, Ell07, EKB02a, JBJB00, JJB02, KSB^{+}02, LS08b, MM03, PVdJB00, PABK03, YJF06]. generator [Fau01, HDBD04, LN01, VW00]. generic [yCkHmY08, Yan04]. Genetic [LSY02, MM07, YL06, BMTSC01, CKMC04, HHJ03, HWDB03, HMSM06, KOFF09, LJS05, SPT07, SBH02, TP01a, WK01]. Geometric [CSRST04, ZZTS09, Est07, LDL^{+}09, RSN^{+}02, ZXY08]. geometrical [GRCD01, Kle03, MLL^{+}08b, PJB^{+}07, SCF^{+}09, GCD04]. geometries [BB08, Han01, IZA06, JČHS07, KKY01, KJP^{+}07, WB07, Wib04]. Geometry [Bud07, LHP01, RK04, VMF^{+}03, BP00, Bie04a, BM00, Cri04, GPSP06, HHBH00, IZA06, KKG^{+}09, KHF^{+}09, Kle03, LJ04, MBP09, MW00, PO03, Pul05, RON02, SCP08, WPS02, ZZS^{+}07]. GEPOL [PTC01]. Germanium [LLXS02]. GFP [HFS^{+}07, NINAT^{+}07]. GGA [DDBP09, Gri06, RLDI09]. GGA-type [Gri06]. GIAO [FO04]. Gibbs [EGSG00, HR08, IGL07, Lu09, SM08b]. Gilbert [GR07]. Gillespie [RMP01]. give [JJK^{+}00]. glasses [NA06]. Global [CZB07, FTLV01, JHZ09, CS02, CMBC08, LS05a, RUPH06, SE08, TSSGS07, TSSSG08, WS02a, WG02, UKN04]. Globally [PAS08, SPT07]. Glu441 [PCS04]. glucans [CMD^{+}04]. glutamate [FTLV01]. glutamic [ZZY07]. glutamine [WC08]. Gly [PC00, VKP^{+}08]. GLYCAM06 [KYT^{+}08, SDL^{+}09]. glycinamide [LB05]. glycine [BA03, BA04a, BA04b, GAIMVB01, GSB09, GKTS04, KAS^{+}07, LB05, LSW^{+}01, MOP^{+}07, PG01, ROG00, ZW09]. glycol [Pin01, RR05]. glycol-lesioned [Pin01]. glycosidase [BMTFR08]. glycosidase-inhibitor [BMTFR08]. glycosidase-substrate [BMTFR08]. glycosidic [SO09, SDL^{+}09]. glycyl [KOML08]. going [CCK01]. gold [BR04, CZ05]. gold-capped [CZ05]. GolP [IDMC09]. good [VGGMM05]. GPCR [XWC09]. GPCR-CA [XWC09]. GPU [NYTH09]. Gradient [SE07, DLD^{+}02, DSR^{+}07, FRLN09, GMA04, ION07, Ish02, IPN07, LST08, TNS00, WL02]. gradient-based [FRLN09]. Gradients [WM12, BWP07, HHS^{+}05, IK00, KBT03, LJ04, SSMW09, vLBBR12]. grain [PSHP08]. grained [CP09, DR07, DJB02, HXLS09, MBC08, SBJ08, VSK^{+}04, VTT^{+}08, WWL^{+}09]. graining [CA07a, EBAN07]. grand [EMP07]. GRAPE [Höf05]. graph [CLZX09, MGMM07a, Pog03]. graphene [KK08c]. Graphical [LD05b, DPDG05, JKII08, KMH02, KBT03, LW04b, LXZ06, Pra01, YNW05, hYDN^{+}08]. graphically [GS09]. graphics [FEV^{+}09, SPF^{+}07, Yas08]. graphite [BCF^{+}09, EKO^{+}01]. Gravitational [WS02a]. greedy [TGD05]. green [DHM^{+}03, XZ05, KK08c, KFD06, ZM03]. grey [XLC08]. grid [ALB09, CG06, Pom04, RSN^{+}02, RKA^{+}09, SKSH07, STH02, WL00, WRBV03, YK08]. grid-based [ALB09, RSN^{+}02, RKA^{+}09, SKSH07, WL00, WRBV03]. GridMAT [ALB09]. GridMAT-MD [ALB09]. grids [Bie04a, SFC04, THHN01]. GROMACS [KVF^{+}07, LSG06, VLH^{+}05]. GROMOS [CLWL09, CHB^{+}05, LH05, OVMV04, SHK^{+}05]. GROMOS05 [CHB^{+}05]. GROMOS96 [SDvG01]. groove [BCP03]. grossular [ZWTP^{+}08]. Ground [HM01, PO03, PSS^{+}04, BBI^{+}09, CWY09, FCW06, FDSA00, IR03, Kri09a, LMK01, ZOJ^{+}06]. Ground- [PSS^{+}04]. Ground-state [HM01, Kri09a, LMK01, ZOJ^{+}06]. group [ATBLS04, CQ04, DVRP^{+}03, Ell07, EB04, JWB05, JGH00, KBT03, LW07, MBM^{+}00, MA05, RCJ02a, RZWS07, RKH03, SGPS09, SN00, TD06, dSGCG00]. groups [BE09, EB04, FJ08, Van02a, WSC09]. growing [Qua07]. grown [WHH^{+}06]. Growth [TDK07, HMK02]. Grubbs [YXC^{+}07]. Grubbs- [YXC^{+}07]. GS [MH09]. GS- [MH09]. GTO [CGB03, RLRE01]. guanine [EL07, GWL07, HHWG08, JM07a, KKMMS04, MSBS01, MHS05, SMKM00]. guanine- [MHS05]. guanine-cytosine [KKMMS04]. guess [Qua07]. guest [LMMW04, Oos09]. GUI [JKII08, SD09]. Guide [SH08, Woo01, You11, Bic09]. guideline [MWE02].

H [AGI^{+}07, BAL^{+}01, BPC01, BL00, CPJ00, CS01, DRAS04, GPSP06, HYA02, IN08, IS03, LDMR01, LMK01, LLXS02, LW04a, LYZ^{+}08, LMO09, Mas01a, Mas01b, MGLL03, SLL^{+}04b, TYN05, UCT^{+}03, WDXS06, XDS06a, ZZL04, ZZZ^{+}06, dRLMS00, CPDZH08, CJW^{+}09, CGB03, DLD^{+}02, Don08, EdlVR^{+}03, Gog08, HK07, ITS06, LC07, LDC^{+}07, LLXS02, LB05, LN01, LLL07, LS05b, LMO09, MR02, McD08, MY08a, NL08, OO04, PGRRNG03, PRSMM03, PV07, RFSS06, RWBH09, SOOF05, SEKS09, SLL^{+}04a, WDS06, WTKM06, Wei08, Wil01b, WDX^{+}02, YTY07, ZY01]. H-bonded [LB05, McD08, NL08]. H-NMR [AGI^{+}07]. H5N1 [DLRZ09]. Hairpin [ZHH09, CJW^{+}09, IGNH03, LHI09]. Hairpins [IGNH03, Der00]. Half [FMAMVK06, PS03, PMM06]. Half-numerical [FMAMVK06]. half-reaction [PS03, PMM06]. halide [RC04, CW02]. halides [AB00, LYK^{+}04, LSY02, ZJM^{+}07]. Hall [SPGS08]. halo [TT02]. halo-hydroxyformaldoxime [TT02]. haloacid [NYK^{+}09]. haloalkane [CS03]. halogen [BS03, FHF^{+}01, GGP09, LZF^{+}09]. halogen-bonded [LZF^{+}09]. halogenated [STC^{+}08, TZX01b, TZX01a]. halogens [TBGRJ04]. halothane [TZX01b, TZX01a]. Hamiltonian [FGR07, FBLO08, MR02, SAM06, ZWPR^{+}04]. Hamiltonians [CV09]. hand [DFGB09]. handle [GCD04, GM04]. Hansen [BBG^{+}04]. Haptic [MR09]. hard [TGGP^{+}00, ZHMW09]. hardness [PRS04, TSSGS07, TSSSG08]. hardware [ATMK03]. harmonic [CLP^{+}05, Ish02, TFN04]. Harris [Cul04]. Hartree [WMW04, AKN07, Bou00, Cul04, DD00, GAdGM08, MS00, MBWP03, PFJ^{+}03, RRS07, WMW03, Wei08, YH07]. HAsXH [LS08a]. having [WJ00]. haystack [BS01]. HBCC [BAL^{+}01]. HBOP [OYH09]. HBr [SLL^{+}04b]. HBSITE [OYH09]. HCCX [Mar03]. HCl [BL06, WDS06]. HCO [JPF^{+}00, dRLMS00]. HCO-L-SER-NH [JPF^{+}00]. HDMR [LRWG03, LAR^{+}03, LSHR04]. head [HSWN01]. heart [TKH07]. heartland [Sha07]. heat [dOMSL01]. heats [CS03, JWB05, RCJ02a, WX09, LLA01c]. heavier [ZJM^{+}07]. heavy [BPC01, SL09, WL04, ZX08]. heavy-metal [ZX08]. HeC [Var09]. Helical [CPML08b, Van08, Der00, KF02a, LC09, PCO^{+}07b, PCO^{+}07a, ZALMG03]. helicenes [VKP^{+}08]. helices [IGNH03]. Helix [BRDC02, JS07b, LI07, PP08a, YS00]. Hellmann [RLER07]. hemagglutinin [DLRZ09]. heme [ATBLS04, MBM^{+}00, OYH05, RGZM09, RZWS07]. hemicarcerand [LMMW04]. hemoglobin [MML^{+}06, SO07, Sen06]. Henry [Sch00, TLOG00]. hept [STC^{+}08]. hept-C [STC^{+}08]. heptafluoropropane [LDC^{+}07]. heptagon [STC^{+}08]. heptagon-containing [STC^{+}08]. heptapeptide [OM04, YAÇ^{+}02]. herbicidal [XYN^{+}06]. hERG [MCR08]. Hess [YH09]. Hessian [KK01a, NKIS02]. Hessian-free [KK01a]. Hessians [ASWG07, Chu07]. heteroaromatic [LLM09]. heterobimetallic [RD00]. heterochiral [ZOJ^{+}06]. heterocycles [FSS00, MGMM07b]. heterofluorenes [CZFH07]. heterogeneity [HS01, ZSC05]. heterogeneous [FCK^{+}08, ZCS04]. heterohelicenes [LC09]. heterolevel [EA08]. heteropentalenes [CDL06]. heteropolymers [SBJ08]. Heuristic [DMC05, DLHC06, CAGR08, IZA06]. Heusler [GD09, KGD06]. hexadiene [PA05]. hexadiyne [PWFS01]. hexagonal [BK08, LTF^{+}07]. Hexahelicene [LC09]. hexamer [NK01]. hexatrienaldehyde [ZGZX07]. hexopyranose [GGK^{+}08, LH05]. hexopyranose-based [LH05]. HF [BRLS12, BRLS08, FKJ^{+}01, GKTS04, PMPGP05, WW03]. HF/6 [FKJ^{+}01]. HF/6-31G* [FKJ^{+}01]. HF/DFT [BRLS12, BRLS08]. HF/MP2 [GKTS04]. HFCO [JHPRSM^{+}05]. Hg [GPSP06, BBI^{+}09, WTKM06]. HH [CMàGL^{+}04]. HI [KKJH08]. Hiberty [Bic09]. hidden [FWH^{+}07, HLT^{+}05, RP07a]. hierarchic [RRS07]. Hierarchical [LMH^{+}09, CWV^{+}05, DJB02, FOK^{+}04, LCC09, UIHN09]. High [BB08, GAdGM08, LAR^{+}03, AZM03, BACJCT01, CCWH02, CN05, DPT03, GL04a, GY08, HGMB04, JBJB00, JJB02, KWK^{+}00, KVF^{+}07, LR06, Mck07a, Mck07b, MTB09, RP07c, RLER07, RSS09, SSS^{+}09, WMRW^{+}01, WS05b, XK08, UTM^{+}02]. High-dimensional [LAR^{+}03, LR06, RSS09]. high-latency [KVF^{+}07]. high-level [WS05b]. high-performance [CCWH02, KWK^{+}00]. High-precision [GAdGM08]. high-quality [JBJB00, JJB02, SSS^{+}09]. high-rank [RP07c]. high-resolution [GL04a, WMRW^{+}01]. High-spin [BB08, DPT03, Mck07a, Mck07b]. high-valent [AZM03, CN05]. higher [BdPRMAI00, LMGR05]. Highly [ZFW08, BWW^{+}08, CKMC04, CLH^{+}07, PPYS08]. Hildebrand [BBG^{+}04]. hindering [HFSD03]. HINT [CPUGD09]. Hirshfeld [DVP^{+}02, DVRP^{+}03, GHBB04]. histogram [Kob03]. histogram-based [Kob03]. HIV [AJNG01, AVS09, BWE05, CLXC02, DLG00, KF08, NLL^{+}09, SPT^{+}03, SVV^{+}08, VVS07, WHF08]. HIV-1 [AJNG01, AVS09, BWE05, KF08, NLL^{+}09, SPT^{+}03, VVS07, WHF08]. HLA [KVS^{+}06, WCF04]. HLA-A*0201 [WCF04]. HLA-DQ2 [KVS^{+}06]. HLA-DQ2/DQ7 [KVS^{+}06]. HLA-DQ8 [KVS^{+}06]. HMLP [DMC05, DLHC06]. Hoboken [Bic09]. hOGG1 [Pin03]. hole [Li01, SZW^{+}05]. hole-particle [SZW^{+}05]. holes [PYCD03, PYS05, PC07, PLC08]. homoalanine [MM00]. homochiral [ZOJ^{+}06]. homoconjugation [MMLC05]. homogeneous [FCK^{+}08, Pog03]. homologous [CC07]. homology [KCL06, KVS^{+}06, OO08, SGS03, YKK09]. homopolypeptides [JS07b]. homotops [TDK07]. Hongxing [Ano06c]. HOO [BL06]. Hoover [QNF09]. hormone [HMK02]. host [CS01, LMMW04, Oos09]. host-guest [Oos09]. Hou [JW12]. HOX [WLLS04]. HP [VGDSU08]. HP-lattice [VGDSU08]. HSAB [PRS04]. Hua [JW12]. Hückel [Kut07]. human [FKU^{+}05, HN02, HMK02, LCC09, PDP02, VGDSU08]. HX [RB01]. Hybrid [BF04, HTN03, WRP^{+}06, ZSK07, AAP00, BG00, BBSS06, DMN03, DDBP09, FAR02, FAB^{+}00, FMSA06, GRO^{+}03, GLD08, HHBH00, Han01, ION07, JPCA08, JIK09, KRM^{+}02, KN04, LSO04, LWZ09, LS05b, MBM^{+}00, MSH^{+}06a, PDS01, RDM^{+}08, Sza08, THHN01, TFN04, WWL^{+}09, XLZ08, ZZL04]. Hybridized [SJJ^{+}04]. hybrids [GXK09, Kan07]. hydrate [IME02]. hydrated [ITS05, XZ04, YSJ09]. hydrates [EM03a]. Hydration [BZL05, CFC^{+}08, HN02, BLL^{+}06, CMD^{+}04, GZL02, HB09, HKMS01, HM02, Lab08, LSW^{+}01, MS03, NTH00, OVMV04, PK04, Pin01, RSP03, UBDPJ04, XLT07]. hydration-parametrized [RSP03]. hydrazines [BLN01]. hydrazone [Lu09]. Hydride [GVATG03, JJH01, LLXS02]. hydrides [KS01b, SRB06, dSGCG00]. hydridotris [HT05]. hydrindans [HKHN08]. hydrocarbon [CS01, KFD06, LC06, Wan09, WEE01]. hydrocarbons [Bor03, BS03, EB04, FVB08, LS08c, MGMM07a, SDvG01, VS08, WFHP01, WJ00, ZKZ^{+}07]. Hydrodynamic [BZP09]. Hydrogen [AG00, Kle03, RP04, XZ04, ZX04, AD00, AST06, BM07, BUMCMRL00, BL06, CUS00, CUSS03, CPDZH08, CVVB04, CCK01, CDPL09, DR07, EFQD09, GAIMVB01, HdMdS05, HdS06, HRG07, HIA03, HT03, HA04, IO08, JP09, mJlZsLyL07, Kle02, LC07, LDC^{+}07, LW04a, LDL^{+}09, Mck07a, MH08a, NHH05, OO08, PG01, Pac06, PGG06, Pog06, Rao00a, RM07, SPT^{+}03, SJW09, TGLL07, WLLS04, WZZ^{+}09, Wil01a, WLL^{+}03, XLL^{+}02, YT04, ZSC05, ZW09, ZH08, ZGZX07, ZX09, vEMK01, vE01, Yos02]. hydrogen-abstraction [WLLS04]. hydrogen-bond [RM07, SPT^{+}03]. hydrogen-bonded [CPDZH08, LDL^{+}09, MH08a, ZH08, vEMK01, vE01]. Hydrogen-bonding [AG00, ZW09, Yos02]. hydrogenase [TDH06]. hydrolases [OBT09]. hydrolyses [DWS^{+}09, LYK^{+}04]. Hydrolysis [WOC^{+}03, DLR^{+}08, MBL^{+}00, RP04, TH02, WJX^{+}08]. hydroperoxy [BL06]. hydrophilicity [DLHC06]. Hydrophobic [MBH^{+}02, CJDK09, HJCP01, SDL07]. hydroxide [CBS^{+}03]. hydroxo [AZM03]. hydroxy [YXZ^{+}04]. hydroxyacetone [WXX03]. hydroxyaromatic [BLO^{+}02]. hydroxyformaldoxime [TT02]. hydroxyl [CUS00]. hydroxylase [HLC09]. hydroxymatairesinol [SH09]. hydroxyproline [BISB02, PRKP05]. hyper [Mar03, vGGB00]. hyperconjugation [CPDZH08]. hyperconjugative [BPC01]. Hyperpolarizabilities [MO01, CJK^{+}02, LWZ09, Tor02]. hyperpolarizability [JPCA08, XWL^{+}09]. hypersurface [SSBE06]. hypersurfaces [PSC^{+}01]. hypochlorous [JKM08]. hypothetical [LD05a]. hypoxanthine [KKMMS04]. hypoxanthine-cytosine [KKMMS04].

I/O [SSL02]. ICFF [KTA03]. icosahedral [Ell07, LML^{+}00, OSA06]. IDE [Gan09]. IDEA [DBGV07]. ideal [Pan07, STSF02]. ideal-gas [STSF02]. identical [CSD05]. Identification [CP09, FWH^{+}07, KS05b, PPXP01, SLC^{+}09]. identify [LHJ^{+}06, ZS04]. Identifying [CCCJ09, DB02, CLS^{+}09, HLT^{+}05]. identity [DSR^{+}07, Nee03, RC04]. idiosyncratic [CMCB08]. II [ACM^{+}06, DPT03, DF04, FNP^{+}06, FKŠ^{+}09, GPK05, LPP06, NK06, Sha02, TGGP^{+}00, WM01, BHW00, BA04b, Ber03, Car02, CSB08, EBD^{+}01, EKB02b, GGB07b, HZ06b, Ish03, JJB02, Kle03, LLA01a, LFZS04, LCDA03, LMIF06, MB00, OBBS05, PMB04, PV03, Rud05b, TJM^{+}03, TCR^{+}02, WHP02, WNH03, WFR08, vEMK01]. III [BB08, DF04, FRS05, KPR04, VHRR07a, YIN03, CCB04, CG05, KEB04, LLA01b, LCGA03, ZFL^{+}05, vE01]. Illustration [KS02a]. image [IC08, XWC09]. IMiCMO [MS01]. imidazole [JKM08, PGG06]. imidazoline [XKG^{+}05]. imide [CXZ^{+}09]. imine [Bor03, GTC06]. imipenem [SDM02]. immune [WCS09]. immunoglobulin [Kró03]. IMOMO [VM00]. IMPACT [BBC^{+}05]. Implementation [AKN07, CKW09, DRMD03, KBT03, LI07, PZS04, RNG03, SVT09, YTH^{+}07, BMRDB01, BLMS08, DFWH05, DBS08, FROD08, GY06, JNV08, NBJ04, PZWG^{+}04, SAM06, SM08b, TT01, VK06, WCC08, YCXY03, LLL03, VW03, YOB^{+}08]. implementations [FL08]. implemented [HP01, MP03b]. implementing [OR05]. Implications [Ano06a, JS07b, KS05c, CBH^{+}03, NHH05, ST06]. implicit [BBHD04, BLL^{+}06, GL04a, JS07a, JZD^{+}09, KTA03, Kró03, Lab08, LSO04, MCM04, PZS04, PPYS08, SBLK01, SL06, WL09a, ZGFL01]. Importance [CGBF05, ENM^{+}04, ZM06, HLC09, JW06, OCB02, PMPGP05, PMM05, TS05]. important [CSU05, EDAJ04, Tor02]. improper [TNS00]. improve [FSM09, XLT07]. Improved [CN03, CLA^{+}00, Gri03, HQ02, KK08a, KK08b, LK04, RCJ02a, SKSH07, TGD05, Wil01a, Wil01b, CMBC08, DSR^{+}07, KDSV02, MP03b, Maz08, MFR07, PABK03, PRS04, SDvG01, STSF02, SHK^{+}05, VZM^{+}08, LK03, RCJ02b]. Improvement [SM08b, UKN04, Nee03]. improves [CLWL09, RK05]. Improving [BUMCMRL00, Bie04b, GF08, LJ04, LKW04, GRO^{+}03, GP06, SMG09]. in-core [FR06]. inactivation [PFR04a]. incidence [YWHZ03]. InCl [ZL05]. Including [IC08, AKN07, DP03, DP04, Gri04, Ish03, LB08, SL09, Wil01a]. Inclusion [HK08a, HK08b, PWHF^{+}03, PWHF^{+}04]. incomplete [FWH^{+}07, Ish04]. incompressible [ZHMW09]. incorporate [KTA03]. Incorporating [CLS^{+}09, DMJV05, HLT^{+}05, SLC^{+}09, HS01, LL07, RD06]. incorporation [SM06]. Increasing [ZWZ09, BT00, LJS05, YAÇ^{+}02]. independent [FVB08, OTL08, Van02a]. index [COS01, JLHF03, MBH^{+}02, PVdJB00, Pog03, YWHZ03, YWH04, YYW07]. indexing [HWDB03]. indicators [BWW^{+}08, HIM07]. indices [BLT03, CGMPT^{+}08, FZL07, FMPS08, FVB08, FS04, GDPCPU07, MGMM07a, MGMM07b, May07, Rao00a, SPGS08, TSMNG01, TSSSG08, WMW03, WW03, WMW04]. indium [ZL05]. individual [ZM06]. INDO [PBZ00, TY03]. INDO/SCI [TY03]. indole [LL01]. induced [CGB03, CMCB08, EDW07, HIM07, HHP04, KIFK07, LGB^{+}09, LBG08, MLA00, PSF^{+}08, RSN^{+}02, ST01, ZALMG03]. Induction [HK08c, HK08d, ZOJ^{+}06]. Inductive [BE07]. inelastic [BACJCT01]. inexact [Har04]. inexpensive [KFZ03]. infinite [DA01]. Influence [GSB09, JS07a, JT08, LZA02, BGC^{+}09, SBH02, SLRC01, DB06, EL09]. Information [Ham07, GCB03, HTKG08, HP05, NL08]. information-bearing [NL08]. infrared [CVR08, Kle03, LDL^{+}09, MGLDS00, TDH06, Zer08]. Inherent [BYQS03]. inhibition [PFR04b, WC08]. Inhibitor [VVS07, BMTFR08, CWV^{+}05, FPG^{+}06, MBC08, SVV^{+}08]. inhibitors [AJNG01, AGO^{+}02, APG05, AVS09, LLL^{+}08, LZ05b, PB06, RGP^{+}07, VB09, VVS07, WZY04, WHF08, ZWB09]. inhibitory [DD08]. inhomogeneous [MZL08]. initial [MM03, MABM09, Qua07, UNM^{+}01]. initialization [FKFG08]. initiated [RAGLL09b]. initiation [GGGLL05]. initio [AJ03, Ama02a, Ama02b, Ano06b, ASY01, BG03, BG00, BL08, BSB05, BS01, BL06, BLO^{+}02, BSJ01, BZL05, CPJ00, CPM03, CUS00, CU01, CUSS03, CLC09, CYM02, CJW^{+}09, CHRL09, DGD^{+}05, DWS^{+}09, DPM09, DB06, Dra00, EM03b, FG02, FAR02, FO08, FNP^{+}06, FKU^{+}05, FKM^{+}06, FKM^{+}07, FKŠ^{+}09, GKRG08, GD06, GBB07, GGLR00, GKTS04, GPK05, GBBH09, Haf08, HSF08, HYA02, HELM09, HSMT04, HMSM06, HFHL06, HS07a, HTSR04, HRR05, HSWW00, HJCP01, ITS05, ITS06, JCA^{+}02, JFG04, JČHS07, KSB^{+}02, KP05, KFNH08, KKG^{+}09, Kle02, KF02a, KIFK07, KWK^{+}00, KPR04, Kri08, Kri09b, LMK01, LYK^{+}04, LDC^{+}07, LMCD09, LMB08, LZJ03, LWLS07, LKT04, LF02, LXSF08, LLL07, LZF^{+}09, LMO09, Mar03, Mas01a, Mas01b, MDA08, MGLDS00, MLL^{+}08b, MM02, MS00, MA09, MG00, MLCD01, MW00]. initio [MS01, Mor02, ML00, Mui05, NYK^{+}09, OO04, OON01, OS08, ON07, OO08, PP08a, PGNG03, PGRRNG03, PHRR08, RSSKB03, RSE07, RRCA08, ROG00, RGG08, SS00, SG07a, SAS05, ŠBL05, Sha02, SLL^{+}04a, SSS^{+}09, SRE08, SMKM00, SMZW05, SK05, SSBE06, TAS07, TYN05, TZX01b, TZX01a, TGGP^{+}00, TFZRG01, UCT^{+}03, UM03, UTM^{+}02, VSK^{+}04, VS02, VIP^{+}06, WMGK07, WLLS04, WLL07a, WS05b, WOC^{+}03, WDX^{+}02, WXX03, WCL05, XLL^{+}02, YXC^{+}07, ZSE08, ZZL04, ZZZ^{+}06, ZWZ09, ZGXX06, ZX08, ZWTP^{+}08, ZL09b, vDSSvA04, vEMK01, vE01, CSV^{+}07, MDI04, SH07, TBG00]. inner [Pog03]. inner-core [Pog03]. inorganic [CMA^{+}08, SYY^{+}03]. inserted [BL08]. Insertion [ZZvRSC08, RD00, TBG00]. insight [MDA08]. Insights [BTP09, PSHP08, SBG^{+}09a, MCK05, PAS07, SGS03]. instabilities [DD00, vW06]. instability [LPK07]. instead [Lab08]. insulator [RDM^{+}08]. insulin [ITN^{+}05, KVS^{+}06, ZM06]. insulin-B [KVS^{+}06]. Integral [JCA^{+}02, BRS00, BRS01, CC09, CJDK09, CF04, DBS08, GWM08, JP09, KJVW08, Lab08, MBWP03, SVT09, UKN04, Vya01, Yas08, PVdJB00]. integral-driven [SVT09]. integrals [ABF^{+}03, FL08, FMAMVK06, FR06, GGA00, Ish03, Lai07, RLER04a]. integrands [GC03]. integrate [MP03a]. Integrated [BBC^{+}05, HT03, LOL^{+}08, MS01, VKCK09, VM00]. integration [BK00, Blo04, CLF^{+}09, ESP04, HW03, JCA^{+}02, KBA^{+}04, LRWG03, Pom04, RP07b, SY09]. integrations [PFB05]. integrator [CF06, KM00, QNF09]. intensities [WKYU01]. Intensity [Tor02]. Intensity-carrying [Tor02]. Inter [HRBKB03, MAF^{+}07, RP02, SWV^{+}05]. Inter- [HRBKB03, RP02, SWV^{+}05]. inter-phycocyanin [MAF^{+}07]. interacting [RGG08, YCXY03]. Interaction [BE09, GBBH09, LSAS01, MHS05, ROG00, AKN07, ABÅ04, AZM03, BRLS08, BRLS12, CLC09, CF04, CFC^{+}08, CPML08a, DLD^{+}02, DMZT08, FK07a, FL07, GDV03, GWM^{+}00, GS04, GKTS04, HZ06a, IDMC09, Ish03, KN04, KBT03, KS05c, LDMR01, LWX07, LPB03, MLL06, MN02, Mas04, NTH09, Oos09, PPXP01, PMPGP05, PRSMM02, PRSMV08, RRCA08, SKGS00, SSL02, SYC03, SWZS04, SLRC01, TKH03, UTM^{+}02, VC04, WS05b, ZCZ03, ZWP08, EB04, JČHS07, LFZS04].

Interactions [CPML08b, FKŠ^{+}09, Van08, ALC08, AG00, ASDP^{+}06, ATMK03, Ano05b, AS00, BPC01, BM07, BUMCMRL00, BSJ01, CMàGL^{+}04, CEP07, DDBP09, EFQD09, FÁ01a, FO08, FNP^{+}06, FKM^{+}06, FKM^{+}07, GGP09, Gon07, HLC09, HA04, HZ06a, HZ06b, IINK09, IB04, JWB05, KF02b, KH06, LHJ^{+}06, LCC09, LZJ03, LS08c, MM05, MMLC05, MCF07, NK06, Nil09, PG01, PCO^{+}07b, PCO^{+}07a, PK05, PNG08, RZWS07, SOOF05, San01, Sha05, SSS^{+}09, SWV^{+}05, SG01, SL06, SDL07, SMV^{+}09, UTT^{+}04, VW03, WRP^{+}06, WPH^{+}07, Won09, YT03, YTH^{+}07, ZTS09]. interactive [DFGB09]. interactively [SB01]. Interatomic [RD06, AMR04, SS00, SPT07]. interconnected [SB08]. Interconversion [OO04]. interconversions [FD03]. Interesting [Kri09a]. interface [CW02, DPDG05, FOK^{+}04, GKRG08, HHBH00, HZX04, JKII08, KKG^{+}09, LLL03, LPB03, PHJ^{+}08, SWM04, TdMSD^{+}08, ZCS04, DBGV07]. interfaced [FKL^{+}06]. interfaces [BSH07, ZZTS09]. interfacial [CW02, MWL^{+}08, PHJ^{+}08]. Interfacing [WHG^{+}07]. interior [SYC08]. interlayer [ALC08]. interlayers [DJT08]. intermediates [BLO^{+}02, BMTFR08, IGNH03, MMMY07, OBT09, WSM^{+}09]. Intermolecular [PSC^{+}01, AS00, CMàGL^{+}04, CLC09, FÁ01a, FKM^{+}06, FKM^{+}07, GGP09, GS04, IGL07, KS05c, LZJ03, Mas04, PMPGP05, RRCA08, SPDS01, SJW09, UTM^{+}02, UTT^{+}04, Wil01a, Wil01b, ZDS^{+}05]. Internal [EA06, BHH^{+}09, CFD03, CFD04, COMR^{+}04, DHF^{+}05, Din00, HFSD03, KSU03, KTA03, LPK07, MGLL03, NKIS02, SWR06, TNS00, WR07, WFR08]. internal-rotation [DHF^{+}05]. interparticle [PK05]. Interplay [EFQD09, SP05]. interpolated [YK08]. interpolation [BB05, IS03]. interpretation [CPJ00, HLS07, VM07]. intersection [SSHT03]. intersections [IK00]. interval [LS05a]. interwall [ZZvRSC08]. Intra [FÁ01a, FKM^{+}06, FKM^{+}07, MAF^{+}07]. Intra- [FÁ01a, FKM^{+}06, FKM^{+}07]. intra-phycocyanin [MAF^{+}07]. Intramolecular [GKTS04, HA04, PG01, TFZRG01, AGK03, BA03, BA04a, FDSA00, HRBKB03, HK08d, Li01, NHH05, RP02, SWV^{+}05, VKP^{+}08, VIP^{+}06, ZDS^{+}05, ZW09, ZH08, Kle02]. Intraprotein [MLJ03]. intraresidue [IB04]. intrinsic [JS07b, JT08, YGZZ05]. intrinsically [NAT07]. Intruder [CWY09, WCFH02]. intuitive [PP08b]. invariant [Est07, ZLY07]. Inverse [BR03, MLL08a, Nil09]. inversion [KSTC01, RC04, ZSE08]. investigate [DWNB01]. investigated [HN02, Kle03, YH09]. Investigation [LZZC09, YTH01, AST06, BL00, CW02, CHA^{+}07, CG08, FG03, GS04, Hir08, JJH01, KYFW07, LH02, LXSF08, NSB08, PV07, QZZZ03, QZL^{+}04, RM07, RC04, RY09, SL04, TGGP^{+}00, TFZRG01, UCT^{+}03, WL09b, WLZ^{+}07, ZXY08, ZKZ^{+}07, ZHMW09, ZGZX07, GBJ03, JBGK08]. Investigations [JP09, WG02]. involvement [BLO^{+}02]. involving [LL01, MM05, ZGZX07]. iodides [CM09]. iodine [GWM^{+}00]. iodobenzoylphosphonate [GWM^{+}00]. Ion [DAK08, BM08, Dra00, EL07, FHRR07, FL07, GWM08, Gor01, IvSV06, KPR04, Kri08, Kri09b, MSBS01, PPYS08, PHRR08, RC04, VHRR07a, VHRR07b, ZZW09, dOMSL01]. ion-pair [RC04]. ionic [Ang09, BM08, BSG07, BGJ01a, CFC^{+}08, DMJV05, GGT08, HTN03, LR03b, Loe03]. ionizable [OS06]. ionization [GSB09, KFD06, RTG00, SVT09]. ionospheric [LSHR04]. ions [CXZ^{+}09, DMJV05, EL06, FG03, HTSR04, HLB09, JRJ01, KT08, KZRO03, LSWB00, LMIF06, MHS05, RMP01, SL09, ZSC05, ZZZ^{+}06]. IPR [GZ07]. IR [NRKH02, ZWTP^{+}08]. iron [DPT03, GK09, HLLN06, LWH06, MSH^{+}06b, OYH05, RJLR06, SW06, TGLL07, TDH06, CN05, LPP06]. iron-containing [MSH^{+}06b]. irregular [ZBS03]. irrelevance [VVBV02]. ISBN [Bic09, Lip00, Sta00]. ISBN-10 [Bic09]. Iso [GWL07, Rap06]. iso-energy [Rap06]. Iso-guanine [GWL07]. isobaric [SM03, Ste04]. isoconversional [CC09, Vya01]. isocyanates [OY01]. isocyanide [HT05]. isocyanurates [OY01]. isoelectronic [Che01]. isolated [RZWS07]. isomer [HLLN06]. isomerase [AGK03, GVATG03, RGP^{+}07]. isomeric [GCCVB00, XFF06]. isomerism [ACM^{+}06]. isomerization [Lu09, ML00, Qua07]. isomerization/decomposition [ML00]. Isomers [Ber03, CCB04, GB02, CDL06, CTFC08, HYA02, KYFW07, KZY09, LS08a, LMO09, Sau04, ZZZ^{+}06, ZZvRSC08]. isomorphic [CRGN07]. isoprene [Dib05]. isoprene-OH-O [Dib05]. isothermal [SM03, Ste04]. isothermal-isobaric [SM03]. isothianaphthene [CFD04]. isotope [GWM08, WDX^{+}02, GM04]. isotopologues [LMB08]. isotropic [BCIB05]. isozymes [WC09]. issues [Mac04, PHFC04]. Iterative [PU09, Rao00b, CC09, HZ06a, HZ06b]. itinerant [SM08a]. IV [AZM03, CN05, CDL06, DMN05, FZL^{+}06, LLA01c, LLA03, LCA03].

J [BPC01, Bof01, HNWF12, Kne05, Qua01, Van08]. Jacobi [IR03]. Jacobians [SWR06]. Jahn [Kri08, VDM06]. Jason [WB04a]. Jersey [Bic09]. Jicun [JW12]. jk [SPGS08]. JNK3 [KK01a]. John [Ano04a, Bic09, Lip00]. Joint [RP07d, DF06, PA05]. Jones [CYM02, FSFK05, Pul05, SCC04, SYC08, YCS07]. Journal [Ano05b, Ano06a, Ano06b, Ano06c, WB04a, WWC^{+}05, Ano01c, Ano04b]. journey [PSCD^{+}09]. jun [KK01a]. junction [DWNB01]. Junmei [WWC^{+}05].

K151 [NYK^{+}09]. Kenny [Sta00]. Kepert [RMP01]. kernel [BRS00, BRS01, DDVD09, TSSGS07, TSSSG08]. kernel-based [DDVD09]. ketene [MGG06]. ketones [LLA01a]. key [HEP^{+}02]. Kick [AM09]. Kier [SPGS08]. Kier-Hall [SPGS08]. kinase [FCP^{+}04a, FCP^{+}05, GdSuM^{+}07, HLT^{+}05, HW09, KK01a, PB06, SWV^{+}05]. kinase-channel-phosphatase [FCP^{+}05]. kinase-specific [HLT^{+}05]. kinases [SWM04]. kind [LX07]. kinematic [CSJD04]. kinematics [CCC03, LFKL00, MLL08a]. kinematics-based [LFKL00]. Kinetic [mJlZsLyL07, Bie04b, BZL05, GWM08, Jac09, Kri09a, WDX^{+}02]. Kinetics [CUS00, ST04, Gog08, HLSH05, MG06, NSU^{+}02, RLP08, UNM^{+}01, VW00, VW04]. kinking [BCP03]. Kirchhoff [YOB^{+}08]. Kirkwood [KS06].

Kirkwood-Buff [KS06]. Kneller [CSD05]. knowledge [Adc04, HZ06a, HZ06b, LWW^{+}06, NMAT01, dSR08]. knowledge-based [Adc04, HZ06a, HZ06b, NMAT01, dSR08]. Kohn [RRS07, Bou00, SH02]. Kollman [JVVK09]. Kr [CMJ08, CGB03]. Kroll [YH09]. Krylov [Har04]. krypton [CVVB04].

L [Bac09, HT05, JPF^{+}00, PC00, AGO^{+}02, HT05, HJCP01, KOML08, NYK^{+}09, OYK^{+}09]. L-2-haloacid [NYK^{+}09]. L-captopril [AGO^{+}02]. L-peptides [OYK^{+}09]. L-phenylalaninamide [HJCP01]. L-valinamide [HJCP01]. La/SSB [KVS^{+}06]. label [VCM01]. labeling [SN00]. lactamase [AGO^{+}02, APG05, SDM02]. lactamases [ESM06, MK02]. LADH [DMC05]. LaN [VP08]. Lanczos [MO01]. Landau [GHH07]. landscape [IGNH03, PAT^{+}09, SPL^{+}02]. landscapes [OKH^{+}02, SSB07]. Langmuir [BRS00]. lanthanide [AB00, FRS05, RMP01, SNM^{+}06, VMA03]. lanthanides [RD06]. Lanthanum [AB00]. Large [WCF04, ARL01, AB08, AS00, BG03, BP01, BdPRMAI00, BME05, CJK^{+}02, CDD^{+}02, CG06, DMN03, DJB02, Ell07, FZL07, HB09, HMSM06, IME02, IS07, JO02, JW00, KS05b, KKG^{+}09, KK01a, KH06, LMJ02, MKGA06, MH09, MHW04, MH08b, MPF00, ME06, NRKH02, PFJ^{+}03, RRS07, SYC08, SSL02, TYO^{+}02, VSK^{+}04, WWL^{+}09, YCS07, vGGB00, WS07]. large-amplitude [KS05b]. Large-scale [WCF04, DMN03, JO02, KK01a, MH09, MHW04, MPF00, ME06, RRS07, SSL02, TYO^{+}02]. larger [VKP^{+}08]. lariat [ZWY^{+}09]. laser [Sha02, KZW^{+}05]. latency [KVF^{+}07]. latter [LPK07]. Lattice [OGH05, SG01, HP01, KWK^{+}01, KWK^{+}02, KF02b, LJKL08, MH08a, SCC04, SYC08, TK08, VGDSU08, YCS07, vE01]. law [Sch00]. LCAO [EBL^{+}08, EL09]. LDA [RLDI09]. lead [RS07a, RS07b]. leads [PPXP01]. learning [YCW^{+}09]. least [CSD05, Gol09, LLZL09]. least-square [LLZL09]. LEDO [GKH05]. legacy [Sha07]. Lei [Ano06c]. length [CRC^{+}08, DR09, JPCA08]. length-frequency [DR09]. lengths [PSC^{+}01]. Lennard [CYM02, FSFK05, Pul05, SCC04, SYC08, YCS07]. Lennard-Jones [CYM02, FSFK05, Pul05, SCC04, SYC08, YCS07]. lesion [Pin01, SHD^{+}08]. lesioned [Pin01, Pin03]. Letter [BFS07]. Letters [JW12, WM12, vLBBR12]. level [BUMCMRL00, BLT03, BL00, DPM09, JMD^{+}02, mJlZsLyL07, KK08c, PFJ^{+}03, RC04, TBG00, TST^{+}08, UTM^{+}02, WyLG^{+}09, WS05b, WLL^{+}03, ZZL04, ZWL^{+}05]. levels [BACJCT01, Cul04, DJB02, PFJ^{+}03, WW03]. Lewis [BHTCG07, GR07, Sha07, Sim07]. LF [PWHF^{+}03, PWHF^{+}04]. Li [CRC^{+}08, GBDP05, JW12, HDO^{+}02, LWK08, LWW^{+}06, LAT05, WWT08]. libraries [AL01, KV00, LZ05b, ZMZ09]. library [CRH^{+}07, FAB^{+}00, KSM05, OTL08, SH07]. LiF [EL09, UM03]. lifetime [CHA^{+}07]. Ligand [MKT04, AM06b, BSP06b, BGC^{+}09, BS08, BMTSC01, CGB^{+}09, CLH^{+}07, CN05, DFWH05, FO08, GZM09, HZ06a, HZ06b, HW09, JZD^{+}09, KS08, LXW^{+}09, Mue01, NR04, NMAT01, OFIK09, PWHF^{+}03, PWHF^{+}04, RK05, Ruv07, SOOF05, STSF02, TFN04, TJE03, VGGMM05, XZZ04, YK00, Yan04, ZGFL01, ZWS^{+}02, BDW00, HLC09]. ligand-charge [BSP06b]. ligand-protein [VGGMM05]. ligands [BS05, CKMC04, FO08, FKU^{+}05, GTC06, GM01, GGLR00, JFG04, RGG08, SWM04, TP01a, TGGP^{+}00, WS02b]. ligation [KT08]. light [Kró03, LFR07]. light-emitting [LFR07]. lignin [PS09b]. LiH [McD03]. like [BCIB05, DHW^{+}00, DB02, FZL^{+}06, JD09, Kut07, LAEL01, PRKP05, WL09b]. limit [MV06, MLL^{+}08b, PSC^{+}01, SAS05, Var09]. limitations [BYQS03, LFEdL06, PRDS08, MFB04]. limited [Ano05b, Sha05]. Limits [OV03]. line [RHL09]. Linear [Con02, DLWV07, KDG^{+}09, LMJ02, OS06, OFIK09, SKDO08, vdVGDM00, AT02, ABÅ04, BH03, BPCD07, CC09, CGMPT^{+}08, GCDL^{+}05, Gol09, GGLR00, Har04, KLM^{+}09, MW09, McD08, Oos09, PK05, RI07, RS05, SSB^{+}03, TCR^{+}02, vGGB00]. Linear-scaling [OS06, SKDO08, GGLR00, TCR^{+}02]. linearized [ABWT09]. link [GdAcV^{+}07, KS02b]. linkages [SDL^{+}09]. linked [CMD^{+}04, CHRL09, FS98, FS00a]. LiPF [BSJ01]. lipid [HNL08, RGG08, SSM08, WC09]. Lipkowitz [Sta00]. lipophilicity [DMC05, DLHC06, DLHC06]. lipoxygenase [TGLL07]. lipoxygenase-1 [TGLL07]. LiPt [LWK08]. liquid [BM08, CC07, EGSG00, GDV03, GJK00, HPL03, MN02, MM07, NL07, PHJ^{+}08, PB04, POJ01, YGLvG06]. liquid-state [POJ01]. liquids [CF04]. list [PABK03]. lists [KUB07]. lithium [HXD08, LWLS07, RC04, SLRC01, YSA^{+}03]. liver [CMCB08]. LMO [BY06]. Local [Din00, LYS08, AGSFAL05, CPML08a, ION07, IS03, JHPRSM^{+}05, KMA^{+}07, LMJ02, MA09, PMC^{+}08, RUPH06, SL09, SEKS09, SB08, TT05, TSSSG08, VKCK09]. localizability [BK08, BWW^{+}08]. Localization [Che01, ALTB06, FS02, FSS00, GBJ03, PP08b, PA05, PAS07, PC05, PC07, SFC04, ST01, WMW03, WW03, WMW04, SHBD05]. Localized [ABF^{+}03, AB09, Bac04, Bac05, Bac07, BME05, FMSA06, GFS05, ITN^{+}05, TT01]. locally [TYO^{+}02]. locate [ABBC01a, ABBC01b, Bof01, GMA04, MP03b, Qua01]. locating [WSM^{+}09]. Log [Tot04]. London [Lab08]. Long [RP07c, CCSJ00, CPC^{+}00, CEP07, CSRST04, Gri06, KSS08, LYS08, MN02, MBC08, RLP08, San01, VVBV02]. long-duration [CCSJ00]. long-range [CEP07, Gri06, KSS08, MN02, RLP08, San01]. long-time [CPC^{+}00]. lookup [Nil09]. loop [CSJD04, KK01a, Mak08, OFB08, PRT^{+}07, PRT^{+}08, TLKT00]. loops [CSRST04]. LoProp [SKK^{+}07]. Low [DPT03, MG06, AG00, BS05, GS03, KUB07, KK01a, KK01b, LAR^{+}03, LBG08, LB08, PFJ^{+}03, PRSMV08, Rao00b, Sha02, WS02b, ZL05, ZL07, ZL09b, dSVA^{+}09, BS08]. Low- [DPT03]. low-energy [Rao00b]. low-level [PFJ^{+}03]. low-lying [LB08, ZL05, ZL07, ZL09b, dSVA^{+}09]. low-mode [KK01a]. low-resolution [BS05, WS02b, BS08]. lowest [FDSA00, OSA06, XZ04, ZX04]. lowest-lying [FDSA00]. lp [LAR^{+}03]. lp-RS-HDMR [LAR^{+}03]. LR [ZWB09, NSO^{+}07]. LR-MMPBSA [ZWB09]. LSCF [FAR02]. Lu [SNM^{+}06]. lumiflavin [CNN07]. LUMMOX [MS04]. lutetium [AB00]. lyase [PMM05, ČJPZS08]. lying [FDSA00, LB08, PRSMV08, ZL05, ZL07, ZL09b, dSVA^{+}09]. LYP [PDS01]. Lysine [DJT08]. lysozyme [HN02].

M [Bof01, GPSP06, JJK^{+}00, LYZ^{+}08, OS08, Qua01, WWC^{+}05, JJK^{+}00, LMGR06, OS08]. machine [CLS^{+}09, HL08, LJZ^{+}07, YCW^{+}09]. machines [CLXC02, QLHL09, YMT04]. macro [Wou00]. macrocycle [RRZA08]. macrocycles [FLOD07]. Macrocyclic [SCG04, KB02]. macromolecular [Ara04, Con02, EA06, FM00, JO02, KS01a, KHY00, RP07d, ZBS03]. macromolecule [BVW04, HMWC03, YH06]. macromolecules [BTLP03, BR03, BH03, EM03b, HHS^{+}05, Mac04, PPXP01]. macropolyhedral [ZZZ^{+}06]. Macrotricyclic [CW02]. Magic [KZW^{+}05, HXD08, KKJH08]. magnesium [ZZW09]. magnetic [BACJCT01, CDL06, CDPL09, DXW08, HWFN01, HIM07, KCL00, KGD06, MV06, MDI04, PJPJdPRMI07, QTdG^{+}08, RLDI09, TDK07, WZXY07, ZPL07, ZXY08]. magnetism [Hua09b]. magnetizabilities [YH07, YH09]. magnetizability [CDL06]. magnets [FKRE08, SM08a]. MAGPACK [BACJCT01]. main [Din00, JGH00, LW07]. main-chain [Din00]. maingroup [SRB06]. maingroup-element [SRB06]. Maintaining [LFBSK07]. maleimide [RP09]. malonyl [LLL^{+}08]. malonyl-CoA [LLL^{+}08]. maltose [MW00]. maltotriose [SWBM08]. Manager [FCK^{+}08]. manganese [AZM03, CWY09, GK09, LMIF06, LS05b, RJLR06]. Many [Loe03, BM08, TKH07, YCXY03, LR03b]. Many-body [Loe03, TKH07, LR03b]. many-electron [YCXY03]. map [MLL08a, SKGS00, HLTLP09]. mapping [FKZ09, RLA01]. maps [PRT^{+}07, PRT^{+}08, TTB01b, WD04]. Marchi [Ano06b]. Marcus [BLN01]. Markov [BHG03, CPUGD09, CMGDAC^{+}07, DLW06, FWH^{+}07, HLT^{+}05, SK09]. Markovian [YCXY03]. mass [GM04, Gor01, KZW^{+}05, LHJ^{+}06, ZWZ09]. masses [CN03]. Massimo [Ano06b]. Massive [TP01a, RLA01]. massively [DGHR02]. master [FR06, FS00b]. master-slave [FR06]. matching [SMM^{+}08, VSW^{+}03]. Material [JW12, SLHW09]. materials [BCF^{+}09, Haf08, LLXS02, LMRVFH^{+}09, Tie09, XK08, YPNE09]. mathematical [DDVD09]. Matrices [LSAS01, AT02, BdPRMAI00, CZA03]. matrix [CGSdST06, Ell07, GHH07, IS07, JCA^{+}02, Li01, NKIS02, Nee03, RS05, RRS07, RRS09, SK09, SSB^{+}03, TYO^{+}02, UIHN09, YWHZ03, vDSSvA04, vW06]. Matteo [Ano06b]. matter [ASDP^{+}06]. maxima [MSH^{+}06a]. Maximal [GCDL^{+}05]. maximization [BWI^{+}02]. maximizing [AM07]. Maximum [MWE02, SCS07, HXD08]. MBO [CPC^{+}00]. MC [HMD06, MLG04]. MCCE2 [SMG09]. MCDP [LAEL01]. McLafferty [NSB08]. MCM [NCO^{+}05]. MCPRO [JTR05]. MCSCF [IR03]. MD [KIM^{+}09, MDA08, ALB09, BMRF01, CADW03, HRR05, Höf05, PHRR08, WRBV03]. MD-based [BMRF01]. MD-GRAPE-2 [Höf05]. MDGRAPE [KAK^{+}09, NYTH09]. MDGRAPE-3 [KAK^{+}09, NYTH09]. MDSimAid [CHMI05]. mean [GMA04, HFSD03, LHI09, MMLC05, NMAT01, RNG03, YCXY03]. meanfield [KRM^{+}02]. meaningful [AE06, Bud07]. means [Bac05, Bie04b, BLF02, Kar06, KBL08, SMAdV00, WKYU01, dOMSL01]. measure [XSHC06, ZHH09, PDC^{+}08]. measurement [YZ04]. measurements [KBLP09]. measures [BDW00, DW08, Ham07, Leh06, PYEA03, PCA^{+}08, PDC^{+}08]. mechanical [AVB00, BISB02, CLP09, CGBF05, CCK01, COL^{+}06, DWC^{+}03, ECA06, ESM06, EBD^{+}01, FHRR07, FÁ01a, FAB^{+}00, FKU^{+}05, GAIMVB01, GGLR00, JJH01, KKS04, MP03a, MBL^{+}00, Sau04, TCR^{+}02, VHRR07a, VHRR07b, XZZ04, XLZ08, YPNE09, ZCZ03, ZAT07]. mechanical/molecular [CGBF05, FÁ01a, TCR^{+}02, VHRR07a, VHRR07b, XLZ08, ZAT07]. mechanics [AS06, AS09, AD00, AM06b, AGO^{+}02, APG05, BDW00, CLFA07, CR02, CSU05, DPT03, DFWH05, DWC^{+}03, EC06, FEVM01, GCD^{+}08, GS04, GKTS04, GPK05, HWTL03, JČHS07, KLB03, KZRO03, LL00, LLA01a, LLA01b, LLA01c, LLA01d, LLA03, LSWB00, MLA00, MFB04, MPF00, OSHS03, PRKP05, PS09b, PWHF^{+}03, PWHF^{+}04, RMP01, RSE07, RP02, ROG00, RM00, RGP^{+}07, SS00, SHD^{+}08, TGGP^{+}00, TFZRG01, TT05, VSW^{+}03, XOW^{+}00, YSA^{+}03, ZSK07]. mechanics-based [BDW00, RSE07]. mechanics/molecular [MPF00]. Mechanism [CJK^{+}02, LWY^{+}09, PFR04a, Rao00a, AM07, AGK03, BTP09, BLO^{+}02, BS03, CGB^{+}09, CBS^{+}03, DBS07, HP04, HLB09, mJlZsLyL07, JDWS06, JJH01, LMGO^{+}09, LPP06, LL01, LDT^{+}02a, LDT^{+}02b, LYZ^{+}08, LFS^{+}07, MCK05, Mui05, PGNG03, PFR04b, PS03, PMM06, wQZsLyZ02, QZZZ03, RJLR06, SMKM00, gThDjL^{+}01, WDWS06, WCW08, WCHW09, WXX03, WJX^{+}08, YQQH09, ZLLS04b, ZLLS04a, ZLLS05, ZLLS06b, ZKZ^{+}07, ZGZX07, GVATG03]. mechanism-based [PFR04b]. mechanisms [AGI^{+}00, AGI^{+}07, BS06, CCCJ09, CG05, ILKR09, KZY09, KKJH08, LLKC06, MK02, NSB08, RC04, Sie01, TMBM02]. Mechanistic [BMTFR08, SGS03, TT02, Ano06a, ST06, WDS06, XDS06a, ZLLS06a]. media [HLLN06, MM02, SMKM00]. mediated [HIA03]. medium [FZL^{+}06, HXLS09, LF04, LFZS04, SHH07, ZFL^{+}05]. medium-resolution [HXLS09]. medium-sized [SHH07]. melatonin [CKT^{+}08]. mellitus [PS09a]. Melting [LML^{+}00, KT02]. membered [FJ08, ZW09]. membrane [ALB09, CJDK09, DAK08, FCP^{+}04a, GAS04, ILKR09, JM07b, LPB03, MHJS06]. membranes [Ike04, SSM08, WC09]. memoriam [Ano00]. memories [WHRG08]. memory [TYO^{+}02]. mercaptocarboxamides [TFZRG01]. mercaptocarboxylate [APG05]. mercaptopyridine [YXZ^{+}04]. mercury [FNP^{+}06]. Merging [PJPJdPRMI07]. Merz [JVVK09]. Merz-Kollman-Singh [JVVK09]. mesh [BYQS03, KM00, KSY^{+}00]. mesoscale [RPMP03, ZBS03]. Mesoscopic [YPNE09]. Met-enkephalin [ZCL09]. meta [DDBP09, ION07, ZTP^{+}08, Gan09]. meta-di-fluorobenzene [ZTP^{+}08]. meta-generalized [ION07]. metabolites [PCMG09]. metabolizing [VB09]. metabotropic [FTLV01]. metadynamics [BBP09]. Metal [SGD06, ABYM08, Ano06a, Bac05, BTP09, BS06, BRV^{+}07, BM00, BWI^{+}02, CM09, CXZ^{+}09, CWWS07, DSB^{+}02, Dra00, FO08, FKRE08, GM01, GS04, HZX04, Hol05, HSWW00, IC08, JHMB^{+}09, JHMB^{+}11, JKL08, KT08, KRM^{+}02, KEM08, LW07, LML^{+}00, LWK08, LMGR06, LS02, LWZ09, MSBS01, NR04, PYCD03, PLC08, RRFC^{+}03, SL09, SZT08, ST06, TAS07, VHRR07b, WB07, WWT08, Wu06, YTH01, ZSC05, ZWY^{+}09, ZX08]. metal-catalyzed [HSWW00]. metal-free [CM09]. metal-organic [TAS07]. metal-porphyrins [LS02]. metal-rich [LWK08]. metallic [ALC08, KWK^{+}01, KWK^{+}02, SK08, WLX^{+}05, WL09b, dVB01]. metallo [AGO^{+}02, APG05]. metallo- [AGO^{+}02, APG05]. metalloenzyme [BSDM04]. metalloenzymes [Sie01]. metallofullerene [CTFC08, KJP^{+}07]. metallofullerenes [KSN01]. metalloproteins [CR09b, SN06]. metals [BP03, BGJ01b, CM09, LD05a, LK03, LK04, WLX^{+}05]. metathesis [PHKG07]. Methane [RRS06, CLC09, EL06, EL07, EM03a, HLLN06, HYR06, IME02, PV07, TMBM02, YQQH09]. methanediamines [CPJ01]. methanol [CCK01, YGLvG06, ZH08, ZWP08]. methionine [BTP09]. Method [KFB05, MO01, YGZZ05, AKN07, ABÅ04, AE06, Ami00, ATMK03, AB09, Ano05b, Ara04, BL09, BWP07, BP01, BVW04, BCNs07, Bie04b, BH03, BF04, BM00, BGC^{+}09, BHG03, BHH^{+}09, CC09, CCL06, CSJ01, CG03, yCkHmY08, CRG01, CPC^{+}00, CAG07, CA04, DPT03, DRMD03, DPRR05, DMLI05, DDVD09, ECM^{+}03, FOK^{+}04, FKL^{+}06, FII^{+}07, FÁ01a, FAR02, FS00b, FSFK05, FRLN09, FKU^{+}05, GMA04, GS09, GY06, GB04, HDBD04, Har04, HFS^{+}07, HHS^{+}05, HNWF07, HNWF12, HQ02, HDF^{+}07, HSWW00, IIK09, IR03, IK00, JO02, JVVK09, KKY01, KDG^{+}09, Kar01, KTA03, KN04, KM00, KLS02, KKC05, KSY^{+}00, KK08c, Kol04, KIM^{+}09, KBB09, KLM^{+}09, KC01b, LSO04, LRWG03, LR06, LWX07, LZZC09, LXW^{+}09, LCSZ09, LZ05b, MLL06, MKGA06, MKT04, MBM^{+}00, MCF05, MSH^{+}06a, MH08b, MG00, MS01, MY08b]. method [MPF00, MRS^{+}07, NKIS02, NA06, NGTB03, NG04, Nye07, OM04, PHJ^{+}08, PABK03, PMC^{+}08, PZS04, Qua07, RNG03, Rao00b, RSE07, RLR^{+}04, RK05, SK09, SMAdV00, Sch04, SCC04, SYC08, Sha05, SF05, SG07b, SWZS04, SVT09, THHN01, TFN04, TVL^{+}03, TY03, TKN^{+}08, VK06, VZM^{+}08, VM00, Vya01, WS05a, WCC08, WW03, WLZ^{+}07, WL00, WBSR03, WM04, WX09, XL02, Yan04, YZ06, YCS07, YAÇ^{+}02, YH06, YH07, YH09, YXL^{+}09, ZWS^{+}02, ZL09a, vdVGDM00, JBJB00, KTM02, SRCD03]. method-based [KIM^{+}09]. methodological [FDSA00, MFR07]. Methodology [KS02a, SPL^{+}02]. Methods [LSAS01, SB01, AJ03, AGMPRG^{+}08, BB05, BL05, CMàGL^{+}04, CG05, DLD^{+}02, DB07, DMN03, DBM03, DDBP09, FOL^{+}04, FO08, GHLK^{+}02, GHH07, GY08, GD09, GPN01, GGLR00, GHBB04, HTKG08, IB04, JFG04, KK08b, KBL08, Kle02, Kle03, KB09, LMV07, LMB08, LLS03, LWH06, LHP01, LK03, LKW04, LZF^{+}09, MN02, MZ05, MBWP03, MHT01, MC06, OFB08, PDC^{+}08, PSMB05, Qua04, RSSKB03, RCJ02b, RK04, SAM06, Sch03, ŠBL05, SYC03, STH02, SE08, SDL07, SSB07, TBG00, TCR^{+}02, TBGRJ04, VMA03, VC04, VMF^{+}03, WMGK07, WMW03, WMW04, WSM^{+}08, WCL05, WM01, YLL^{+}09, ZM03, vDSSvA04]. methoxycarbonyl [KK09]. Methyl [CADW03, CCK01, DBM03, DMN05, HT05, RC04, WLX^{+}05, WDZS07, WSM^{+}08, WLL^{+}03, WC08, ZLLS05]. methylacetamide [MMPK01]. methylacetylene [ZKZ^{+}07]. methylamine [LMB08]. methylation [EL07, HM08, SLC^{+}09]. methylene [LFS^{+}07]. methylenimine [dOMSL01]. methylenimmonium [dOMSL01]. methylimidazole [HT05]. methyloxaziridine [ZPL07]. methyloxirane [ZPL07]. methyltransferase [WC08]. MF [DHW^{+}08]. MF-3D-QSAR [DHW^{+}08]. Mg [WZZ^{+}09, AS06, LST08, ŠBL05]. Mg-porphin [ŠBL05]. mGluR1 [FTLV01]. MgO [SBG09b]. MHz [CMD^{+}04]. MICCs [YGZZ05]. micelles [KS02a]. Michael [Ano00]. microcanonical [Rap06]. Microiterations [VMF^{+}03]. microscopic [PK04]. microsecond [MST^{+}08]. microsolvation [UM03]. microwave [ZGXX06]. middle [RPNJ07]. migration [HLB09, ZGZX07]. MILCH [BL09]. mimic [WMS06]. minima [CA04, DLD^{+}02, KG02, XOW^{+}00]. Minimal [BWZ08, BVW04, Maz01]. minimization [EGSG00, OM04, Sen06, WG02, vEMK01, ABF^{+}03, NCO^{+}05]. minimizations [dSR08]. minimized [GLD08]. Minimum [AM06b, AJ03, CZB07, CY09, CY13, SBLK01]. mining [CDD^{+}02, KG02]. minor [BCP03]. misfolded [CP08, WS07]. mispairs [SG07a]. mixed [FÁ01a, HELM09, MPF00, MC06]. mixed-valence [HELM09]. mixing [AMR04]. mixture [WG02]. mixtures [FBDG06, NL07, YGLvG06]. MLR [HSMT04, LLL^{+}08]. MM [CGBF05, MPF00, AST06, CR09b, CG05, FAR02, FMSA06, FSFK05, GWM08, GC04, GWM^{+}00, HHBH00, HBM06, HNR08, HRR05, HTN03, IV04, IvSV06, ITS05, ITS06, KHF^{+}09, KPR04, Kri08, Kri09b, LLL03, MBM^{+}00, MK02, MSH^{+}06a, MG00, MLJ03, NGTB03, PB06, RG02, SBG^{+}09a, SN06, SMM^{+}08, SVV^{+}08, THHN01, TdMSD^{+}08, VMF^{+}03, WCC08, WHF08, WHG^{+}07, WC08, YZ06, ZWZ09]. MM-PB [GC04, WHF08]. MM-PBSA [PB06]. MM2 [KKY01]. MM3 [AD00, LSWB00, YSA^{+}03, FH01, Sto05, SBH02, TAS07]. MM3/MM4 [AD00]. MM4 [AD00, CLFA07, LLA01d, LLA03, ACLD03]. MMPBSA [ZWB09]. MMVB [GRO^{+}03]. Mn [AZM03, BL00, GD06, BL00, PYS05]. MNDO [DC02, RCJ02b, TBGRJ04]. MNDOC [IK00]. Mo [LZZC09, ML00, DRAS04, DRAS05, LDMR01]. MO/statistical [ML00]. mobility [HIM07]. MoCalc [DPDG05]. Modal [ES00]. mode [EBAN07, HNR08, KSU03, KJP^{+}07, KK01a, SFRS01, TVL^{+}03]. mode-specific [SFRS01]. Model [CPML08b, Duk01, EL07, Van08, AB08, ABBC01a, ABBC01b, ATM^{+}07, APG05, BG00, BBHD04, BLL^{+}06, BT00, BCIB05, BB08, BLN01, Bof01, BLMS08, CFK08, CF04, CFC^{+}08, CP08, CP09, CRSB03, COMR^{+}04, CMGDAC^{+}07, CA07b, COL01, COL^{+}06, DLWV07, DHW^{+}00, DAK08, EA08, EDW07, EKO^{+}01, FKL^{+}06, FCK^{+}08, FEVM01, FC06, FBLO08, FZL^{+}06, FNP^{+}06, GZL02, GL04a, GK09, GJK00, GKTS04, GGT08, GWS^{+}02, HB09, HPL03, HS01, HK08c, HK08d, HHP04, ILB03, JBJB00, JJB02, JPF^{+}00, JBGK08, KSB^{+}02, KFZ03, Lab08, LRI^{+}02, LAR^{+}03, LFZS04, LJ04, LSHR04, LR06, LK03, LK04, LKA01, MWL^{+}08, MMPK01, MCF05, MBC08, Nak07, NL07, OKE^{+}02, PCO^{+}07b, PCO^{+}07a, PMB04, PFC03, Pom04, Qua01, RMP01, Sen06, TJE03, TCT03, TGGP^{+}00, Tot04, TT02, VW00, VW04, VGDSU08, VP09, WCK00, WKYU01]. model [WS05b, WEE01, WOC^{+}03, WJX^{+}08, XL02, XZ05, XLT07, YL09, YPNE09, YJF06, vDSSvA04, FCP^{+}04b]. modeled [PB05, vDSSvA04]. Modeling [ECM^{+}03, FRS05, MCR08, Mck07a, Mck07b, MTB09, PSCD^{+}09, SEKS09, Sie01, BA03, BA04a, BZP09, BM08, BSH07, CLP09, ČJPZS08, DJT08, DMN05, EDAJ04, ENM^{+}04, GL04a, Gor01, HBM06, HMSM06, HMMS09, HRBKB03, HP04, Hin00, JM07b, JTR05, KCL06, KS01a, KJP^{+}07, KVS^{+}06, KPZK06, LEK07, LLL^{+}08, LFR^{+}04, MMLC05, MBM^{+}00, MPF00, OFB08, SPGS08, SGS03, SYY^{+}03, SS05, SPF^{+}07, Sto05, TTBM09, VBGL^{+}00, XLC08, YKK09, ZBS03, ZMH^{+}09, BBC^{+}05]. modelling [PSHP08]. Models [JB04, AS09, AHGK09, ACM^{+}06, CCK01, CPUGD09, CRGN07, CA04, CA07a, CCT^{+}03, DMN03, DLG00, DR07, DDVD09, EC06, FWH^{+}07, FK07b, GSB09, GDV03, GS02, GDPCPU07, GDPP08, GS08, HdMdS05, HdS06, HD06, HP01, HLT^{+}05, HG08, HJCP01, JPF^{+}00, KS02a, Kró03, LJKL08, Leh06, LS08b, LDTS07, MTE04, MA05, MC06, OGH05, OYH05, PFJ^{+}03, PA05, QSS01, RD06, RSP03, RSER09, RS08, RR05, SBLK01, Sch00, SSS^{+}09, SRB02, SKK^{+}07, SB01, SL06, VBS09, WB04a, WB04b, WB05, WZXY07, YÇBM00, YGLvG06, YJF06, YKK09, ZLJS03, ZCL09, ZGFL01, ZLD09, ZWP08, TDH06]. Modern [PB02, FLK^{+}07, Pra01]. modes [Gra07, LSY02, MGLL03, OR05, Tor02]. Modification [HNL08, Vya01, YWH04, CM09, KFZ03]. Modified [LC06, NTH09, RC04, AVS09, CLA^{+}00, KKY01, NA06, VVS07, WCS09]. Modifying [XLT07]. Modular [EA06]. module [HMD06]. moduli [LZZC09]. MOF [TAS07]. MOF-5 [TAS07]. Moffitt [Kar01]. moiety [LBG08]. Moldyn [RMHK03]. Molecular [AS09, BBG^{+}04, BG07, BDW00, CLC09, CLFA07, CCK01, DJT08, EMP07, FEVM01, FPN^{+}05, GJK00, HLB09, Ish03, JTR05, KŠB09, KAS^{+}07, KLB03, KIM^{+}09, LLA01a, LLA01b, LLA01c, LLA01d, LLA03, LSWB00, MLA00, Maz01, MO09, MS00, MST^{+}08, NBJ04, Pin01, RMP01, RRZA08, SHD^{+}08, SDM02, VSW^{+}03, WEE01, YSA^{+}03, YGLvG06, YKK09, ZCZ03, ZWS^{+}09, AM09, ARL01, AS06, AG00, ALB09, AD00, AGMPRG^{+}08, ATMK03, AM06b, AB09, AGO^{+}02, APG05, AS00, BG03, BP00, BR07, BA08, BB05, BWE05, BRDC02, BWZ08, BVW04, BT00, BSOB05, BME05, BSJ01, BPCD07, CMJ08, CC07, CDS09, CCL06, COS01, CW02, CIB05, CDD^{+}02, CFK08, CCSJ00, CGBF05, CF06, CF04, CPC^{+}00, CR02, CCP04, CEP07, CMD^{+}04, CSU05, CBH^{+}03, DvG00, DB07, DPT03, DFWH05, DFGB09, DPDG05]. molecular [DSS03, DMC05, DLHC06, DK01, DWC^{+}03, ESM06, EKB02a, EKB02b, FSM09, FHRR07, FG02, FOK^{+}04, FKL^{+}06, FII^{+}07, FÁ01a, FBDG06, FLOD07, FAB^{+}00, FKZ09, FEV^{+}09, FKRE08, FNP^{+}06, FRLN09, FKU^{+}05, FKM^{+}06, FKM^{+}07, GL04a, GLP08, GKRG08, GCCVB00, GLMV09, GFS05, GLD08, GRCD01, Gly06, GS02, GS03, Gon07, GSDT09, GCD^{+}08, GS04, GKTS04, GPK05, Ham07, HB09, HHWG08, HYT05, HGMB04, HH04, HELM09, HM08, HLS07, HSWN01, HN02, HP04, Hin00, HIA03, HTKG08, HK08d, HW03, HTN03, IKN08, ITS05, IC08, Ish04, IIK09, IKYM09, JS07a, JMD^{+}02, JP09, JGVF05, JFG04, JHPRSM^{+}05, KMH02, KFZ03, KKG^{+}09, KM00, KCL06, KEB04, KKC05, KSY^{+}00, KAK^{+}09, KCK^{+}08, KBL08, Kni00, KZRO03, Kol04, KIFK07, KvGH01, KH06, KC01b, Kri09a, KPR04, Kri08, Kri09b, Kró03]. molecular [KKS04, LFK05, LL00, LSG06, LFSB03a, LFSB03b, LCKL05, LPK07, LJ04, LTF^{+}07, LS05a, LKT04, LM03, LPB03, LEV^{+}09, MB00, MFB04, MN02, MBP09, MM03, MABM09, Mas01a, Mas01b, MDA08, MVL^{+}05, MH08b, MG00, MCM04, MS01, MPF00, NK01, NYK^{+}09, OBBS05, OS06, OO04, OO06, OSHS03, OR05, OCP02, ON07, OO08, PMGL03, PRKP05, PMB04, PS09b, PBW^{+}05, Pin03, PWHF^{+}03, PWHF^{+}04, PDS01, Pog06, PTC01, PRT^{+}07, PRT^{+}08, PZS04, PPYS08, PHH^{+}08, PB02, PNG08, QTdG^{+}08, QNF09, RP07a, Rao00b, RP02, RLRE01, RLER04a, RLDI09, RSN^{+}02, RMHK03, ROG00, RM00, RGP^{+}07, RHL09, SF07, SO07, SDL^{+}09, SH09, SBJ08, SEKS09, SBG^{+}09a, SYY^{+}03, STH02, Ste04, SPF^{+}07, SVT09, TJM^{+}03, TYN05, TLKT00, TT08, TFN04, TGGP^{+}00, TFZRG01, TCR^{+}02, TT01, TT05, TRS02, TBGRJ04, UTH^{+}03]. molecular [UIHN09, VSK^{+}04, VW03, VCM01, VP02, VHRR07a, VHRR07b, VC04, VM00, WCF04, WZZ^{+}09, WWL^{+}09, WTKM06, WMS06, WNH03, WD08, WSC09, WRBV03, WM01, XLZ08, XOW^{+}00, YGZZ05, YXC^{+}07, YL09, YH06, YH07, YZ04, YTH^{+}07, ZMZ09, ZAT07, ZWZ09, ZSK07, vdVGDM00, BRV^{+}07, CLP09, HWTL03, LPP06, LLXS02, LGB^{+}09, MCR08]. molecular-dynamics [WD08]. molecular-loop [PRT^{+}07, PRT^{+}08]. molecule [ALKH04, BAL^{+}01, CLP09, FM00, KK08c, MZ05, NK06, Oos09, Pin03, RGG08, SM08b, SFC04, SSW^{+}07, SVV^{+}08, WLPF05, WDS06, XDS06b, ZLLS06a]. molecule-based [KK08c]. Molecules [LLXS02, BG03, BHTCG07, BG00, BZP09, BT00, BKS02, CLP^{+}05, CG03, Che01, CRSB03, Cul08, DDKV07, DHF^{+}05, DR09, DRAS04, Ell07, EKO^{+}01, FCW06, FÁ01a, FS02, FS04, GCDL^{+}05, GCD04, GBBH09, HIM07, HDF^{+}07, IS07, IC08, JVVK09, JG03, KGN07, KH01, KP05, KKG^{+}09, KLH^{+}04, KCK^{+}08, KOFF09, Kol04, KvGH01, Leh06, LZA02, LYS08, MWL^{+}08, Mat03, MSH^{+}06b, MSR04, MA09, MH08b, MBH^{+}02, NRKH02, PO03, Pan07, PHJ^{+}08, PDS01, PPYS08, PSS^{+}04, PSMB05, RLR^{+}04, RLER04a, RLER04b, RRZA08, RZWS07, SD09, SHH07, TLOG00, TD06, Tor02, VKP^{+}08, Van02a, VM07, WCK00, WL04, Wil01b, WR07, WFR08, Wou00, YT03, YGZZ05, YK08, Yos02, ZZY07, vEMK01, vE01, vGGB00, PFB05]. MOLFDIR [PVdJB00]. Møller [CPML08a, DSR^{+}07, FII^{+}07, Gri03, IN08, JSHG07, Var09, WCFH02, YH09]. MOLMAP [HMMS09]. MOLPRO [VW03]. molybdenum [BGC^{+}09, SS05]. moment [HK08a, HK08d, HK08b]. moments [DVP^{+}02, GdSuM^{+}07, MLA00, Mar03, PP08b, RP07c, VC04, YOB^{+}08]. momentum [GY08, WMRW^{+}01]. mono [CU01, GWL07, MGMM07b, HT05]. mono- [CU01]. monoanion [Bac09]. monochloride [ZL05]. monocopper [GTC06]. monoderivatives [EB04]. monofluoride [LB08]. monohalides [CWWS07]. monohydrates [GGP09]. monomer [YXC^{+}07, ZM06]. mononitrides [Wu06]. mononuclear [CZ05, GK09]. monooxygenase [HLLN06, HYR06, TMBM02]. monophosphate [MRS^{+}07]. monophosphide [ZKZ^{+}07]. monopole [ZFL^{+}05]. monoprotonated [HA04]. monosaccharides [GGK^{+}08]. monosilicide [HXD08]. Monosilicon [YDWS06]. Monosilicon-substituted [YDWS06]. monosubstituted [COMR^{+}04, Lee09]. monoxide [GGP09, HT05, YQQH09]. Monte [NCO^{+}05, SCS07, AGSFAL05, AGSFA^{+}05, BR03, BHG03, Der00, FCK^{+}08, FKFG08, GHH07, HMD06, IM06, IKYM09, KLS02, KM07, KKC05, LML^{+}00, LZA02, LRWG03, MH09, Nak02, NA06, OM04, SKGS00, SBJ08, SM08b, SWR06, TS05, XKG^{+}05, ZCS04]. Monte-Carlo [KLS02]. montmorillonite [DJT08]. MoO [LZZC09]. MOPED [SRCD03]. MORPHY [MP03b]. Morse [SDCG02]. Mössbauer [HLLN06]. most [KAS^{+}07]. motif [HHW^{+}03, LLL07]. Motifs [HWTL03, WHH^{+}06]. motion [BRDC02, CCSJ00, LPK07]. motions [HSWN01, KS05b, LV08]. Mott [RDM^{+}08]. MOVB [MG00]. move [SM08b]. moving [CvG08]. MP2 [WD04, BP02, EA08, FJP07, GCCVB00, GKTS04, IPN06, IPN07, JPF^{+}00, ME06, PFJ^{+}03, PMC^{+}08, SAM06, WD04]. MP2/cc [WD04]. MP2/cc-pVTZ//MP2/6-31G** [WD04]. MpProp [SKK^{+}07]. MPSim [CWV^{+}05]. MRCI [KTM02, SZW^{+}05]. MSINDO [BGJ01b, JGH00, JW00, NBJ04, SBG09b]. MST [COL01, CSB^{+}03, FBLO08, MBH^{+}02]. MST-based [MBH^{+}02]. Mulliken [GHBB04]. multi [ABWT09, SL09]. multi-region [ABWT09]. multi/heavy [SL09]. multiatom [SSB^{+}03]. Multibaric [OO06]. multibody [CPC^{+}00]. Multicanonical [HHHS01, SKGS00, YÇBM00, KH01, YAÇ^{+}02]. multicenter [DBS08, MS01]. multicentered [DWNB01, HT03, WBSR03]. multicomponent [ST04]. multiconfiguration [NUH02]. multiconfigurational [GD06, PJPJdPRMI07]. multicore [KHF^{+}09]. Multicut [LSHR04]. Multicut-HDMR [LSHR04]. Multidimensional [AL01, ARL01, Chu07, HP05, PC00, PFC03, RNG03]. Multidimensionality [FVB08]. multiensemble [HKMS01]. multiexponential [GC03]. multifarious [Sim07]. multifield [BRDC02]. multigrid [BB05]. Multiisotopic [Gor01]. multilayer [LJZ^{+}07, SJJ^{+}04]. multilayered [MR04]. multilevel [BHW00, HBW00, HBW01, JNV08]. multiobjective [CMBC08]. Multiple [CLF^{+}09, CLZX09, DHW^{+}08, JW06, SK09, STH02, BYQS03, CV09, Gol09, KM00, KH06, LHJ^{+}06, MST^{+}08, PAT^{+}09, STCJ08, XOW^{+}00, PYCD03]. multiplications [SSB^{+}03]. multiply [HT03]. multipoint [WS05b]. multipolar [DWNB01]. multipole [Ami00, ATMK03, BH03, CRG01, GY08, KM00, KLM^{+}09, Mar03, RP07c, SF07, SG01, SvDS01, TFZRG01, VC04, WL09a, WBSR03, YOB^{+}08, ZFL^{+}05]. multipole-based [WL09a]. multipoles [KS01a, SKK^{+}07]. multipopulation [HHJ03]. Multireference [WNH03, CWY09, DLD^{+}02, HELM09, KBT03, MLL06, ME06, QTdG^{+}08, UKNS01, UKN04, WCFH02, dSVA^{+}09]. multireference-MP2 [ME06]. Multiscale [San01, OFB08]. multiscaling [VTT^{+}08]. Multispecies [GDPP08]. Multistate [JHPRSM^{+}05, FSM09, MST^{+}08, YFS07]. multithermal [OO06]. mutagenesis [MFR07]. mutant [DLRZ09]. mutants [GDPCPU07, MRS^{+}07]. mutase [HHBH00]. mutations [HFS^{+}07]. mutual [HTKG08]. mutual-information [HTKG08]. MXO [HT05]. myoglobin [AZS^{+}04]. myosin [HSWN01].

N [GR07, JD09, KYFW07, KSN01, LS08a, LWW^{+}06, Lu09, Mck07a, Mck07b, OS08, Sim07, TK08, WDXS06, WD08, WJX^{+}08, XWXC08, YHD^{+}06, ZY01, ZXY03, RMHK03, XZ04, ZX04, CPC^{+}00, DRAS05, FJ08, FH01, HDO^{+}02, KKH^{+}07, KBL08, Lu09, LKA01, McD08, Mck07b, PFC03, RFSS06, SN00, SRE08, WLL01, Wil01b, WC08, WJX^{+}08, ZW09, ZX09]. N’- [WJX^{+}08]. N-dimethyl [WJX^{+}08]. N-dimethylhydrazone [Lu09]. N-formyl-serinamide [PFC03]. NaCl [PK04]. NAMD [PBW^{+}05]. Nano [Est07]. nanoalloy [LJS05]. nanocomposite [DJT08]. nanomaterials [GJL^{+}08]. nanomedicine [PSCD^{+}09]. nanoneedles [PSCD^{+}09]. nanoparticles [CGG06, KEM08, ZWC^{+}09]. nanotube [KK08c, XWL^{+}09]. nanotubes [BG07, ZZvRSC08]. nanowire [KK08c]. naphthalene [CDPL09, HRG07, WL09b]. naphthalene-like [WL09b]. naphthoic [CMLS05]. naphthylisoquinoline [BMRF01]. native [BS01, yCkHmY08, DB02, MMMY07, WS07, ZCL09, ZS04]. native-like [DB02]. Natural [FLGW00, Bac07, GS07, LZ05b, PU09, RD06]. naturally [CJW^{+}09]. Nature [CQ04, SK08, Ang09, KSK00, PYS05]. Nb [WD08]. nbo [Kar01, BPC01]. NCH [KRLD09]. NCN [LD05a]. NDDO [CSB^{+}03, FÁ01a]. NDDO-based [CSB^{+}03]. near [BVW04, PABK03, YL09, ZS04]. near-minimal-volume [BVW04]. near-native [ZS04]. near-neighbor [PABK03]. near-solute [YL09]. Nearest [HDF^{+}07, HTKG08]. Nearest-neighbor [HDF^{+}07, HTKG08]. NEB [GF08]. needle [BS01]. Negative [BLO^{+}02]. Neglect [Lai07]. neglected [WCF04]. neighbor [HDF^{+}07, HTKG08, PABK03]. neighbors [RP07d]. nematogenic [CLP09]. neopentyl [YTY07]. Nernst [DAK08]. net [BED02]. Network [KYL03, AG03, CLC03, Gol09, GDPP08, GAS04, HMSM06, KEB04, LSY02, MVLG06, NINAT^{+}07, SJJ^{+}04, SPT^{+}03, UIHN09, WX09, LMH^{+}09]. network-based [GDPP08, GAS04]. networks [BMRDB01, BSH07, FCK^{+}08, KVF^{+}07, LJZ^{+}07, PS09a, RLA01, TCSM03, VGDSU08, dVB01]. networks-based [PS09a]. Neural [GAS04, AG03, CLC03, Gol09, HMSM06, KEB04, LJZ^{+}07, NINAT^{+}07, PS09a, RLA01, SJJ^{+}04, TCSM03, WX09, dVB01]. neuronal [SBG^{+}09a].

Neutral [DWS^{+}09, ASS^{+}02, Bac09, CYM02, DLR^{+}08, EBDPM00, FCP^{+}05, MT03, OSA06, PGG06, ROG00, VM02, Wan09]. neutrals [LLXS02]. neutron [BACJCT01, RMHK03]. nevirapine [AJNG01]. new-generation [YJF06]. Newly [CRS05]. News [Ano04b, BACJCT01, DvL01, Gly06, JVVK09]. Newton [Har04, Qua07]. NF [FJ08]. NH [DMN05, LF02, dOMSL01, DRAS04, ITS05, JPF^{+}00, KT08, KSTC01, LDMR01, MR02, Mck07b, PC00, SEKS09, dOMSL01]. NHC [ZZL^{+}08]. NHNH [LWY^{+}09]. Ni [Bac09, KGL07, PMM06]. NiAt [ZL07]. niches [TP01a]. nickel [Bac09, GK09, LMIF06, YQQH09]. nicotine [VB09]. nicotinic [GCD^{+}08, SBG^{+}09a]. NiH [ZL07]. NiO [SBG09b]. niobium [Tie09]. NiSOD [PMM06]. nitrate [CGB^{+}09]. nitrenium [FG03]. nitric [JDWS06, LPP06]. nitride [UNM^{+}01]. nitrides [LX07]. nitrido [Bac05]. nitriles [POJ01]. nitrite [DBS07]. nitro [MA05, POJ01, UTH^{+}03]. nitroethane [GWM08]. nitrogen [BGC^{+}09, LLM09, gThDjL^{+}01, WC08, XWL^{+}09, ZLLS04a, ZLLS05, ZWY^{+}09, ZMH^{+}09]. nitrogen-containing [LLM09]. nitrogen-pivot [ZWY^{+}09]. nitrogen-vacancy [ZMH^{+}09]. nitrogenase [Ano06a, Mck07b, ST06]. nitroguanine [JM07a]. nitromethane [LZJ03]. nitrones [MGG06]. nitroxyl [LPP06, VCM01]. NLOPredict [MMP^{+}07]. NMR [AGI^{+}07, BPC01, BRDC02, CADW03, CMD^{+}04, CMA^{+}08, FO04, HP05, Kle03, KBLP09, MC06, PC00, PFC03, PF06, RI07, RSER09, ZPL07]. NO [LDT^{+}02b, PGNG03, PGRRNG03, ZLLS04b, ZLLS06b, IS03, FJ08, GBBH09, LDT^{+}02a, MR02, PGNG03, RAGLL09a, XDS06a, ZLLS06a]. Nobelium [HdMdS05]. noble [SRB06]. nodal [HYT05]. nodes [Kau07]. NOE [AGI^{+}07, PF06]. non [GZ07, Gon07, SVT09]. non-bonded [Gon07]. non-Dyson [SVT09]. non-IPR [GZ07]. Nonadditive [Don08, PMB04, ZWP08]. nonbonded [ASDP^{+}06, ATMK03, DK01, GWM^{+}00, KH06, PABK03, SF05]. nonbonding [IB04, ZTS09]. noncentrosymmetric [GBJ03]. Noncollinear [Van02b]. Noncovalent [Won09, JČHS07, SP05, SMV^{+}09, TH02]. noncovalently [PHFC04]. nonelectrostatic [KF02b]. Nonempirical [KSK00]. nonequilibrium [FZL^{+}06, GG09, KK08c, LF04, LFZS04, OD09, YZ04, ZFL^{+}05]. nonhybrid [DF04]. nonisomorphic [CRGN07]. Nonlinear [RLA01, BF04, BF07, Har04, HLSH05, MMP^{+}07, WCL05]. nonlinearity [LPK07]. nonmetallic [ALC08]. nonnative [yCkHmY08]. nonnucleoside [AJNG01]. nonorthogonal [SMZW05]. nonparametric [HDF^{+}07]. nonplanar [Din00]. nonplanarity [RKH03]. nonpolar [GZL02]. nonrelativistic [WL02]. Nonspecific [LPB03, RGG08]. nonstable [GDPCPU07]. Nonuniform [SHSF05, Bie04a]. norbornadiene [WXK08]. norm [RRS09]. normal [EBAN07, KSU03, OR05]. nose [BBG^{+}04, QNF09]. Note [Ano04b, Ano04a, FBS09]. Notes [CDGS09, CDS09]. Novel [ILKR09, JLHF03, NL08, TRS02, WL09b, YWHZ03, CKR08, CMBC08, GZL02, GDV03, LXW^{+}09, RI08, TYO^{+}02, YJF06, ZNLL07, ZL09a]. novo [LEK07, VGO^{+}07]. Np [Han01, GZL02, Ike04, GHLK^{+}02]. NPAT [Ike04]. NPT [Ike04]. NQR [MH08a]. NR [CPJ00]. ns [CMD^{+}04, SO07]. NSAIDs [CMBC08]. Nt [ZNLL07]. nuclear [CR09a, CDL06, GM04, HWFN01, IKN08, QCK01, QCK02, WZXY07, ZPL07, ZXY08]. nuclearity [BACJCT01]. nucleation [CKW09]. nuclei [CDPL09]. nucleic [CCK01, DP03, DP04, FZL07, FM00, HWTL03, JCL05, MB00, Nak07, OMNH08, PPYS08, RKH03, SYC03, SL04, SWR06, SHK^{+}05]. nucleobases [FKŠ^{+}09, SBI08]. nucleophilic [BSB05, SSB07]. nucleoside [Wil01b]. nucleosides [SA07]. nucleosome [VTT^{+}08]. nucleotide [Mak08, MSF^{+}08]. nucleotides [XWXC08]. nucleus [FVB08, HdMdS05, HdS06, HD06, IKN08]. NUCS [SHSF05]. nudged [AJ03]. number [CDS09, HXD08, KZRO03, KZW^{+}05, KH06, TGGP^{+}00, WWL^{+}09]. numbered [GYCZ04]. numbers [GdAcV^{+}07]. Numerical [DLW06, LXL07, MO01, QNF09, TT08, WL04, vW06, DB07, ESP04, FMAMVK06, IO08, WL00, YK08, WG02]. Numerov [Bie04a, Bie04b]. Numerov-type [Bie04a].

O [BL00, GCCVB00, GPSP06, HYR06, ITS06, mJlZyL^{+}08, KGL07, LZZC09, LMO09, Mas01a, Mas01b, NA06, Owe05, PGNG03, PGRRNG03, UCT^{+}03, XWXC08, YHD^{+}06, ZJM^{+}07, ZY01, ZZL04, ZZL^{+}08, ZXY03, Bac09, CCCJ09, DRAS05, Dib05, HM08, HDO^{+}02, IS03, LC07, LW04a, LS05b, MGLL03, NyHN06, NHN06, RAGLL09a, RFSS06, SSL02, SRE08, Wil01b, YTY07, ZX09, KZY09]. O-methylation [HM08]. O3LYP [BP03]. OB [NA06]. object [CRH^{+}07, FL08, MVL^{+}05]. object-oriented [CRH^{+}07, MVL^{+}05]. objective [WG02]. objectives [STCJ08]. objects [RSN^{+}02]. observables [MG06]. observations [FWH^{+}07]. observed [VBS09]. obtain [BVW04]. obtained [HFSD03, VC04, WMW03, WMW04, WHH^{+}06]. obtaining [Bac04, YGZZ05, SK09]. occupied [HHWG08]. occurring [CJW^{+}09]. OCF [UTM^{+}02, UTT^{+}04]. OCHF [YLW^{+}08]. OCl [HLLS05]. OCLO [WLZ^{+}07]. OCO [VM07]. OCS [VS02, ZGXX06]. octahedral [OSA06]. octan [BE07]. octan-1-yloxy [BE07]. octanol [COL01, CSB^{+}03, Gol09, Tot04]. octanol/water [CSB^{+}03, Tot04]. octet [GR07]. Odd [CC07, GYCZ04]. odd-numbered [GYCZ04]. off [HP01, LJKL08, XLT07]. off-lattice [HP01, LJKL08]. off-plane [XLT07]. OH [Dib05, Gog08, HTN03, IvSV06, LW04a, Mas01b, WLL07a, CU01, CUSS03, GAIMVB01, GCCVB00, GGGLL05, HTN03, KZY09, Kle03, KBLP09, LC07, LWY^{+}09, Mas01a, MGLL03, Mui05, RAGLL09a, RAGLL09b, SEKS09, UCT^{+}03, WLLS04, WLL^{+}07b, WyLG^{+}09, YLW^{+}08, YLWL09, ZZL^{+}08]. OH-initiated [RAGLL09b]. OH-rotamer [KBLP09]. OH-stretch [Kle03]. OH/Cl [YLW^{+}08]. OHO [Wil01a]. OHS [JP09]. olefin [PHKG07, YXC^{+}07]. olefins [AVB00]. oligomeric [EL07]. oligomers [BSG07, CSJ01, Der09, LFR07, SBB02, WCL05, ZOJ^{+}06]. oligopeptides [MGJAARC00]. Oligovalent [KS02b]. OLYP [BP03]. on-the-fly [KMA^{+}07]. On-the-path [CY09, CY13]. One [CR09a, BG03, Bac07, Bie04a, GKRG08, KFD06, Kri09a, Lai07, LB05, ZWS^{+}02]. one- [Lai07]. one-dimensional [Bie04a]. One-electron [CR09a, BG03, Bac07, GKRG08, KFD06, Kri09a, LB05]. one-step [Oos09]. ONIOM [BGC^{+}09, MDA08, MC06, VMF^{+}03, XKKL03]. ONIOM-molecular [MDA08]. ONO [FJ08]. onto [NK06]. OOPSE [MVL^{+}05]. open [CSV^{+}07, FS02, PRSMM03, LLA01a]. Open-chain [LLA01a]. open-shell [FS02]. open-source [CSV^{+}07]. opening [SRE08]. OPEP [ÁCD^{+}03]. operating [DFWH05]. operation [PCA^{+}08, SYC08]. operators [KRM^{+}02, Qua04]. OPFMM [CRG01]. OPLS [KB02, KOML08, KDSV02, MT03, POJ01, PB05, XLT07]. OPLS-AA [KOML08, POJ01, PB05, XLT07]. OPLS-AA/L [KOML08]. OPO [KZY09]. Oppenheimer [ZWZ09]. opposite [JSHG07]. opsin [RG02]. optical [Bou01, CZFH07, CTFC08, Hua09a, KŠB09, LC09, LFR07, MA09, SN06, TDK07, WCL05, YFR05, Zer08, ZX08]. optics [MMP^{+}07]. Optimal [GFS05, ÁCD^{+}03, BSP06b, Blo04, CRG01, DDVD09, SPT07, TTBM09]. optimal-parameter [CRG01]. Optimization [Ano06c, GL04b, GKH05, IK00, WCS09, WM12, AJ03, AM06a, BP00, BdPRMAI00, BM00, Bud07, BLMS08, CS02, CZB07, COS01, CYM02, CLH^{+}07, CY09, CY13, CHMI05, CMBC08, DMN03, DV02, FM00, FRLN09, GHMP03, HHBH00, HLTLP09, KKG^{+}09, KBA^{+}04, KHF^{+}09, KK01a, LJKL08, LJ04, LS05a, LJS05, MKT04, MM03, MM00, MW00, MGJAARC00, Pen06, PU09, Pul05, RK04, SCC04, SYC08, SWM04, SSMW09, SE08, SBH02, STCJ08, VMF^{+}03, WS05a, WS02a, WPS02, XZZ04, YL06, YCS07, ZBS03, ZZ08, vLBBR12, WZW^{+}06]. optimizations [IR03, PO03, RON02]. Optimized [KM07, VB03, VK06, BSDM04, FKFG08, LJZ^{+}07, MV06, MY08b, MY08a, WTKM06, WNH03]. optimizer [KG02]. optimizing [QSS01, SRCD03]. Orbit [Duk01, CR08, CR09a, DXW08, KRM^{+}02, KTM02, LB08, LXSF08]. Orbital [KIM^{+}09, Pen06, WM12, ALTB06, AB09, AS00, CFK08, FOK^{+}04, FKL^{+}06, FII^{+}07, FKU^{+}05, FKM^{+}06, FKM^{+}07, GCCVB00, GS07, HHWG08, Hir08, IKN08, IIK09, KDG^{+}09, KBL08, KIFK07, LKT04, Mas01a, Mas01b, MG00, MY08b, NYK^{+}09, OS06, OO08, PU09, SRE08, SSMW09, TBGRJ04, UIHN09, VM00, WPH^{+}07, vLBBR12, vdVGDM00]. orbital-based [CFK08]. orbital-correlation [SRE08]. Orbital-orthogonality [Pen06]. orbital-valence [MG00]. orbitals [ABF^{+}03, Bac07, BME05, CVVB04, EdlVR^{+}03, FMSA06, GFS05, GC02, HYT05, ITN^{+}05, Ish03, JHPRSM^{+}05, Kau07, Kni00, LW07, MS01]. order [Bie04a, DSR^{+}07, FS04, Gri03, IN08, JSHG07, KGN07, LAR^{+}03, LS08c, May07, MO09, QTdG^{+}08, Rud05a, Rud05b, Rud05c, YH09]. ordering [SM08a]. organic [ATH^{+}03, BLL^{+}06, BT00, CCK01, DA01, EDAJ04, EBD^{+}01, Gol09, HELM09, HP04, JLHF03, JVVK09, JTR05, KLH^{+}04, LH02, LJZ^{+}07, LMRVFH^{+}09, PO03, PB04, SJJ^{+}04, SYY^{+}03, TAS07, Van02a, WCK00, YGZZ05]. organizing [BA08, ZA07]. organocatalytic [WSM^{+}09]. organocopper [YIN03]. organocuprate [YIN03]. organometallic [Gor01, SYY^{+}03, TD08, TTBM09]. orientation [BL00, MWL^{+}08]. oriented [CRH^{+}07, FL08, MVL^{+}05, RMHK03]. Origin [JS07b, GYMN07, KMM07]. Orthogonal [Bac07]. orthogonality [Pen06]. orthogonalized [Lai07]. orthonormality [ABF^{+}03]. oscillation [CAG07]. oscillations [DF06, SMKM00]. oscillator [AB08, LLM09]. other [PS03, VMF^{+}03]. outer [BF07, GAS04]. outline [SE08]. overall [LZ05a]. overcome [Vis02]. overdetermined [RI07]. overlap [LKW04, SGPS09]. overview [Sch03, Mac04]. oxazolidones [OY01]. oxazoline [XKG^{+}05]. oxidase [BS06, JKL08, WZY04]. oxidases [PS03]. oxidation [BTP09, DRAS05, GCCVB00, GGGLL05, LB05, RAGLL09b, SS05, XPW09]. Oxidative [DGD^{+}05, LL00, PMM06]. oxide [BSJ01, CFS^{+}08, CCCJ09, JDWS06, JT08, LPP06, PV07, RRS06, SBB02, ZCS04]. oxide/electrolyte [ZCS04]. oxides [ZSC05, ZLLS05]. oxidized [CNN07, CR02]. oxidoreductases [CFS^{+}09]. oxidosqualene [SGS03]. oxo [CN05, WJX^{+}08]. oxo-porphyrins [CN05]. oxocarbenium [LSWB00]. oxoguanine [FPN^{+}05, JM07a, Pin03]. oxohydrocarbons [Wil01a]. oxoimidazoles [JKM08]. oxonols [BG00]. oxygen [GTC06, GWM^{+}00, MML^{+}06, SO09, WSC09, XPW09]. oxygen-adsorbed [XPW09]. oxyl [AZM03]. oxyl- [AZM03]. oxynitrides [WD08]. ozone [YLZ08].

P [BAL^{+}01, Gog08, KZY09, LS08a, Lip00, OS08, QB05, WZZ^{+}09, ZY01, MK02, CCP04, Mit01, RPNJ07, RFSS06, Tot04, KZY09]. P450 [AST06, HBM06, LCC09, ZAT07]. P450-catalyzed [AST06]. P450nor [LPP06]. P450s [OYH05]. p8 [BRDC02]. P Anhar [GBDP05]. package [AGSFA^{+}05, BACJCT01, CSV^{+}07, GSDT09, IM06, KSY^{+}00, Kli01, KWK^{+}00, MABM09, OTL08, PVdJB00, RMHK03]. package-independent [OTL08]. packed [AT02]. Packing [MM03, CM09, CA07b]. Packmol [MABM09]. PAH [Don08]. Pair [FK07a, FS04, GR07, HZX04, KSTC01, LBT07, MLL06, MGCA07, Nil09, PC05, RC04, SC01, SYC03, Sim07, PC07]. pairing [DP04, HWTL03, PC05, PC07]. pairlist [HH04]. pairlist-construction [HH04]. pairs [BM08, CJDK09, FZL07, KKMMS04, PABK03, ZZW09]. pairwise [Ano05b, CLZ^{+}09, MTE04, Sha05, VP02, VZM^{+}08]. Palermo [Van08]. palladium [WCW08, WCHW09]. palladium-catalyzed [WCW08, WCHW09]. pancreatic [MBC08]. paper [JW12, WM12]. para [ASDP^{+}06, KC01a, ZX09]. para-didehydropyridine [KC01a]. para-didehydropyridinium [KC01a]. para-hydrogen [ZX09]. paradox [CDGS09]. Parallel [BWP07, DOSG06, MBWP03, TGGP^{+}00, UIHN09, ASWG07, Ano05b, AGSFA^{+}05, BP02, BWM^{+}09, CRG01, GBDP05, GS04, GKTS04, GPK05, HHJ03, HHHS01, IS07, IPN06, IPN07, KKC05, KOFF09, KVF^{+}07, MVL^{+}05, MGJAARC00, NKIS02, NG04, Sha05, SPT07, TYO^{+}02, TFZRG01].

Parallelization [GJK^{+}06, PVdJB00, PV03, SZW^{+}05, UKNS01, CCWH02, FOK^{+}04, FCK^{+}08, UKN04, VSK^{+}04, vGGB00]. parallelized [TP01a, VK06]. parallelizing [SO07]. parameter [BLMS08, CRG01, CHMI05, HXLS09, MO09, OVMV04, SHK^{+}05, FM00]. Parameterization [KB02, PNG08, SMM^{+}08, TCT03, BGJ01b, FH01, JKL08, JGH00, LSWB00, MTE04, PB04, RKH03, TGLL07, VSW^{+}03, WK01, JJB02, JVVK09]. parameterized [GB04]. parameters [AAP00, AMR04, Ano06c, ATBLS04, BBG^{+}04, BSDM04, BZL05, CYM02, DB06, DDVD09, FAR02, FSFK05, FRS05, HPL03, KFNH08, KOML08, KVL^{+}04, KC01b, MMMY07, MSR04, MRC03, MLL^{+}08b, MC06, OYH05, OMNH08, OBT09, PRKP05, RRCA08, SO09, SEKS09, SRCD03, SHD^{+}08, SCF^{+}09, TT05, TTB01a, VCM01, VIP^{+}06, WZW^{+}06, ZSK07]. Parametrization [PDS01, COL01, SBH02, WS05b]. parametrized [RSP03, TAS07]. paraoxon [ZWS^{+}09]. parent [MDI04, YLW^{+}08]. Pareto [STCJ08]. Paris [HP04]. Parr [Kri09a]. Parrinello [JP09, Sch04]. part [AGI^{+}07, CDS09, ESP04, GDPP08, LLA01d, vDSSvA04, AGI^{+}00, Rud05c, vEMK01, vE01]. Partial [Ike04, BSÇ^{+}01, Gol09, KC01b]. partially [SVT09]. particle [Ano05b, BYQS03, CZB07, GY08, KM00, KSY^{+}00, LJZ^{+}07, Sha05, SJJ^{+}04, SHH07, SZW^{+}05]. particles [BCIB05, WWL^{+}09]. particularly [BS06]. partition [CCK01, CSB^{+}03, DP03, GLD08, Gol09, RM07, TS05]. partitioning [ÁCD^{+}03, DVP^{+}02, DVRP^{+}03, HMSM06, RP07a, VC04, WNH03]. path [ABBC01a, ABBC01b, Blo04, Bof01, CY09, GF08, GWM08, JP09, Kli01, Qua01, UCT^{+}03, VGB08, WLPF05, WHG^{+}07, CY13]. paths [FG03]. pathway [LGB^{+}09, WLL01]. pathways [AJ03, JW06, LZKT04, MAF^{+}07, Qua04, RAGLL09a, RAGLL09b]. Pattern [DGHR02, EKB02a, EKB02b, KEB04, AGMPRG^{+}08, HWDB03, dGWH01, EKB02a]. patterns [CGG06, Gor01]. Pauli [Ish03]. PB [GC04, WHF08]. PBCAID [QSS01]. PBSA [PB06]. PbTiO [ZXYF09]. PC [Ano01b, BMRDB01, HSMT04, OSHS03]. PC-GA-ANN [HSMT04]. PCM [FKL^{+}06, CRSB03]. PCs [HS07b]. Pd [DGD^{+}05, GBBH09]. PDDG [RCJ02b, TBGRJ04]. PDDG/MNDO [RCJ02b, TBGRJ04]. PDDG/PM3 [RCJ02b, TBGRJ04]. PDE7 [DD08]. pea [PS03]. penicillin [MK02]. penicillin-binding [MK02]. penicillins [DSS03]. pentacarbonylmanganese [PYS05]. pentanes [BPC01]. PEPCAT [OML^{+}00]. Peptide [Adc04, DHW^{+}07, HJCP01, JPF^{+}00, ONHN00, PFJ^{+}03, BTP09, BWE05, BSP06a, CLWL09, CSJ01, CJW^{+}09, CLA^{+}00, CP09, DvG00, DWNB01, GSB09, IGNH03, LHJ^{+}06, LL01, MS03, MHT01, MST^{+}08, OGH05, OKH^{+}02, PHFC04, SDL07, Tot04, WCF04, Wil01b, YZ06, YÇBM00, ZALMG03, ZCZ03, WHP02, KVS^{+}06]. peptide/HLA [KVS^{+}06]. peptide/HLA-DQ8 [KVS^{+}06]. Peptides [CPML08b, Van08, Ano06c, BBHD04, BCP04, BAH^{+}02, CP08, DJ04, EA08, HHP04, IKYM09, LKJ^{+}04, LLW02, LXL07, MM00, MC06, OML^{+}00, OYK^{+}09, OO08, OM04, PCO^{+}07b, PRKP05, PFC03, SJW09, WZW^{+}06, YAÇ^{+}02, ZW09, ZLD09, ZOJ^{+}06, PCO^{+}07a]. peptidomimetics [BAH^{+}02]. percolation [Mei02]. Perfect [Wan09, OCB02]. perfluoro [FO04]. perfluorosulfonate [YSJ09]. perform [ME06, WCK00, WHG^{+}07]. Performance [BM00, Cul04, CA04, DMN03, FOL^{+}04, JM07b, KPZK06, LLS03, RLDI09, VBS09, ZM03, AM06a, BL05, BRV^{+}07, BLMS08, CCWH02, DF04, DB06, DGI^{+}08, FMPS08, KWK^{+}00, KEM08, KS05c, LWH06, MA09, NYTH09, SF07, SCF^{+}09, Sto05, SBH02, UKN04, WMGK07, WL04, WSM^{+}09, WM01, BP07]. Performances [CLP^{+}05]. Periodic [PMC^{+}08, Ami00, BVW04, DRMD03, FROD08, HH04, Kau07, KSS08, KAK^{+}09, QSS01, SRB06, WM06, ZLD09]. peripheral [BGC^{+}09]. periplasmic [CGB^{+}09]. permeability [CRGN07]. permittivity [GPN01, PZS04]. permutation [SN00]. perovskite [WD08]. perovskite-type [WD08]. peroxa [BLO^{+}02]. peroxidase [HBM06]. peroxidative [MRS09]. peroxides [LLZL09]. peroxy [Dib05]. peroxynitrite [JM07a]. personal [May07]. perspective [KRLD09, LMGO^{+}09, PBF09]. perspectives [Fie02]. perturbation [CWY09, CPML08a, CG05, DRMD03, DSR^{+}07, FII^{+}07, Gri03, IN08, LKW04, MRS^{+}07, NUH02, Oos09, PMGL03, Pog06, QTdG^{+}08, RSE07, SWZS04, UTH^{+}03, UKNS01, UKN04, Var09, WCFH02, WNH03, YH09]. perturbations [OV03]. perturbed [DOSG06, ZZW09]. pesticides [KEH^{+}02]. PH [RD00, DR09, WDS06, MCM04]. pharmaceutical [KV00]. pharmacophore [BA08, JFG04, LFKL00, HHG^{+}09]. pharmacophore-constrained [LFKL00]. phase [BAL^{+}01, CPJ00, CPJ01, DR09, DGI^{+}08, DWC^{+}03, FBDG06, FM00, GLMV09, JJK^{+}00, JHZ09, KSB^{+}02, KT08, KFNH08, KKH^{+}07, LRI^{+}02, Lee09, LB05, LLL03, MFB04, Mas01a, Mas01b, MM02, Mor02, POJ01, PV07, wQZsLyZ02, QNF09, RRS06, ROG00, SMGE08, SDvG01, SMKM00, TK08, TDH06, UCT^{+}03, UNM^{+}01, WD04, XKKL03, XKG^{+}05, YQQH09, ZALMG03, ZSK07]. phase-space [QNF09]. phases [ALC08, CLP09, LXSF08, SK05, XB08]. PhAST [HHG^{+}09]. Phe [VKP^{+}08]. Phe-Gly-Phe [VKP^{+}08]. Phen [ZWS^{+}02]. phenol [LL01]. phenols [HM08]. phenomena [KK08c, RSS09]. phenoxycarboxylic [XKKL03]. phenyl [WZY04]. phenylalaninamide [HJCP01]. phenylalanine [SMV^{+}09]. phenylene [ASDP^{+}06]. Philippe [Bic09]. phillipsites [LST08]. phonon [EL09]. phosphatase [AG00, FCP^{+}04a, FCP^{+}05]. phosphate [LDY^{+}08, MBL^{+}00, PMM05, PHRR08]. phosphates [WOC^{+}03]. phosphatidylcholine [CEP07]. phosphine [HT05, LL00, MGLDS00]. phosphinine [FLOD07]. phosphinine-containing [FLOD07]. phosphininium [LTF^{+}07]. phospho [RGP^{+}07]. phosphodiesterase [XLZ08]. phosphodiesterase-5 [XLZ08]. phosphohistidine [KVL^{+}04]. phosphoimidazole [KVL^{+}04]. phosphole [LFR07]. phosphole-containing [LFR07]. phospholipid [MCR08, RG08]. phosphomannose [RGP^{+}07]. phosphonic [CJK^{+}02]. phosphorus [LYK^{+}04, LTF^{+}07, Mit01]. phosphoryl [ZJM^{+}07]. phosphorylation [HLT^{+}05]. phosphotriesterase [KZRO03, ZWS^{+}09]. phosphotyrosyl [OO08]. photoabsorption [CHRL09]. photoadsorption [ZMH^{+}09]. photocatalysts [HZ09]. photochemical [Ama02b]. Photochemistry [GD06, SRE08]. photodetachment [LMCD09]. photodissociation [JHPRSM^{+}05, LXSF08, WXX03]. photoelectron [VDM06]. photoemission [RDM^{+}08]. photoexcited [SRE08]. photographic [Sha02]. Photoionization [MY08b, MY08a]. photoisomerization [GRO^{+}03]. photosynthesis [Ano06b, CPM03]. photosynthetic [IN01, OON01]. photovoltaic [LMRVFH^{+}09]. phthalocyanine [CM09]. phthalocyanines [LS02]. phycobilisomes [MAF^{+}07]. phycocyanin [MAF^{+}07]. phylogeny [LXZ06, ZLY07]. physical [BRS07, DHW^{+}07, OFB08, OS08, SRCD03]. physically [AE06]. physico [AGMPRG^{+}08, Mat03, SB01]. physico-chemical [AGMPRG^{+}08, Mat03]. physico-chemically [SB01]. physicochemical [CP08, CP09, FTLV01, KLH^{+}04, KEM08]. physics [DB02, SPL^{+}02, WS07]. physics-based [DB02, SPL^{+}02, WS07]. physio [CDD^{+}02]. physio-chemical [CDD^{+}02]. piano [FKŠ^{+}09]. piano-stool [FKŠ^{+}09]. picture [VBGL^{+}00]. Piero [Ano06b]. pinacol [YTY07]. Piotr [Ano06c]. pivot [ZWY^{+}09]. pK [KKS04, ZCS04]. pKa [CFR06, OS06]. planar [CSB08, MMRVH07, SRS07, SBG09b, Wan09]. Planck [DAK08]. plane [PSS^{+}04, PSMB05, RLDI09, VSK^{+}04, XLT07]. plane-wave [PSS^{+}04, VSK^{+}04]. plane-wave-based [RLDI09]. planewave [YK08]. plaster [HP04]. Plastocyanin [SN06]. platform [Gan09]. platinum [CSB08, SMM^{+}08, WM01]. plausible [CBS^{+}03, SB01]. play [YJF06]. Playstation [LEV^{+}09]. pleated [PGC05]. Plesset [CPML08a, DSR^{+}07, FII^{+}07, Gri03, IN08, JSHG07, Var09, WCFH02, YH09]. plot [KMH02]. plots [CLZ^{+}09, SDL^{+}09, SRE08]. PLP [PMM05]. plus [AGMPRG^{+}08, CG05, IKN08]. PM3 [BM00, BSDM04, DC02, GM01, MSH^{+}06b, RCJ02b, TGLL07, TCT03, TBGRJ04]. PM3-compatible [BSDM04]. PM3/d [TGLL07]. PM5 [LKT04]. PMF [Mue01]. pocket [BS08, MDA08, OYH09]. pocket-specific [BS08]. Point [Est07, BCNs07, Bie04a, CRC^{+}08, DWC^{+}03, GDV03, GGLR00, KGL07, KK08a, KFZ03, MGCA07, SRB02, TBSM09, WMS06, ZMH^{+}09]. point-charge [DWC^{+}03, GGLR00, SRB02]. points [BMLV04, BAÅ07, DLD^{+}02, GMA04, HQ02, MP03b]. Poisson [WB04a, WB05, ABWT09, BHW00, BH03, BF04, BF07, DLG00, DAK08, FOL^{+}04, GPN01, GCD^{+}08, GGT08, Höf05, HBW00, HBW01, KWHH07, LDG02, NYTH09, PZS04, SATO04, Vas02, VZM^{+}08, WB04b, ZGFL01]. Polanyi [Nye07]. polar [BAÅ07, CYM02, CPML08a, EB04, FÁ01a, HLLN06, HSF08, JPF^{+}00, PFC03, ZXYF09]. polar-neutral [CYM02]. polarisabilities [ZPL07]. polarizabilities [FROD08, LFK05, LYS08, MLA00, MY08a, SKK^{+}07, Tor02, Whe08]. polarizability [BP01, HK08a, HK08b, Mar03, Mor02, QCK01, QCK02, vGGB00].

Polarizable [CFK08, LLM09, Nak07, Ano06c, AGO^{+}02, APG05, BCIB05, COL^{+}06, DGI^{+}08, FKL^{+}06, GWM^{+}00, GS04, GKTS04, GPK05, HHP04, JZD^{+}09, KSB^{+}02, Kol04, LJ04, MMPK01, MBC08, OR05, PWHF^{+}03, PWHF^{+}04, Pom04, RGP^{+}07, TFZRG01, WZW^{+}06, YGLvG06, FCP^{+}04b]. polarization [CGB03, CBH^{+}03, EDW07, GGLR00, GKTS04, HK08a, HK08b, KFZ03, MR04, Maz08, RP02, SL09, WL09a, YL09]. Polarized [EdlVR^{+}03, BSOB05, OBBS05]. poly [ASDP^{+}06, BSJ01, CHA^{+}07, CFD04, MGMM07b, Qua07, SBB02, ZALMG03]. poly-isothianaphthene [CFD04]. poly-para-phenylene [ASDP^{+}06]. polyacenes [BPCD07]. polyacetylene [PM02]. polyacrylates [LZA02]. polyalcohols [KBLP09]. polyatomic [GGB07a, GGB07b, RLER04a]. polyatomics [TP01b]. polyazidocubanes [JWB05]. polycoordinated [TGGP^{+}00]. polycyclic [Bor03, CA07b, FVB08, MGMM07a, VS08]. polyenes [MW09]. Polyethylene [BCF^{+}09]. polyketides [KB02]. Polymer [Mei02, BBG^{+}04, CZA03, DJT08, MM07, RRZA08, YSJ09]. polymeric [Fau01, JCA^{+}02]. polymerization [BG07, YXC^{+}07]. polymers [CFD04, CA04, CA07a, DC02, Der09, Din00, DDBP09, HM01, LAEL01, OKE^{+}02, SHH07, VIP^{+}06, YYW07]. polymorphism [VVBV02]. polynomial [HDBD04]. polynuclear [HYR06, RRFC^{+}03]. polyoxoanions [LFR^{+}04]. polypeptide [Cri04]. polypeptides [CPML08a, IB04, KF02a, KF03, Nak02, VP09]. polyphosphate [MRC03]. polythiophene [CA07b]. POPC [JM07b]. Pople [Ano04a, EA08]. population [BLT03, BPCD07, Pul05]. population-based [Pul05]. populations [KBN02]. porphin [ŠBL05, ŠBL05]. porphycene [NyHN06]. porphyrazines [LS02]. porphyrin [AZM03, CHRL09, LPP06, NyHN06, NHN06]. porphyrin-fullerene [CHRL09]. porphyrins [CN05, LS02, LWH06]. portable [SH07]. positron [RGG08]. Possibilities [PRDS08]. possibility [LMGR05, LBG08, TT05]. Possible [HIA03, OCP02, WLL01]. post [WW03]. post-HF [WW03]. posteriori [SPDS01]. potassium [MCR08, MHS05]. potential [AMR04, AE06, ABBC01a, ABBC01b, BCNs07, BL05, Bof01, BBI^{+}09, DMLI05, DMC05, DLHC06, DK01, FSFK05, FKRE08, HPP00, HRBKB03, HPL03, HFSD03, IS03, IT03, JZD^{+}09, LFK05, LMK01, LS08c, MMLC05, MCF05, Nak07, NG04, NMAT01, PSC^{+}01, Qua01, RD06, RNG03, RHL09, RTG00, SPDS01, SS00, Sch03, SMGE08, SSS^{+}09, SHH07, SG07b, SBB02, SJW09, TBSM09, TLKT00, WCC08, WL09a, WCK00, WS07, YH06, YHD^{+}06, ZCS04, ZZY07, ZZY08, ZGXX06]. potential-derived [TBSM09]. potentials [ATM^{+}07, CLC09, CPUGD09, CKW09, DB02, FAB^{+}00, FNP^{+}06, GK09, GBJ03, HZX04, HHHS01, HZ06a, IKYM09, KLH^{+}04, KCK^{+}08, KK01b, LI07, LHI09, LK03, LK04, LLW^{+}09, MCF05, MWE02, OR05, PML03, RPMP03, RLER04b, SMAdV00, SPT07, VGDSU08, dSR08]. Powder [HWDB03, HHJ03, dGWH01]. powerful [PSDM00]. pp [Bic09, Lip00, Sta00]. PQS [BWM^{+}09]. PR [AVS09, VVS07]. Practical [BMRDB01, PHR^{+}05, Woo01, You11, Blo04, Sch03, SHSF05, SWZS04, WW03]. precalculated [ZMZ09]. preceding [CSD05]. Precise [Ami00, Ara04]. precision [CN03, GAdGM08]. precursors [CFD03, CFD04, DJT08]. predict [HL08, HZ06a, HZ06b, LL07, PB06, PJPJdPRMI07, XSHC06, XLC08, YMT04]. Predicted [PDP02, IGL07, JARM02, KCL06, WS02b, ZCL09, ZGXX06]. Predicting [DR09, Der00, LKA01, ZLJS03, AG03, CLXC02, CRGN07, IO08, KS02a, XWC09]. Prediction [AVS09, CLC03, CKR08, CJDK09, DA01, ELK^{+}09, FCW06, Gol09, JIK09, KLH^{+}04, KCK^{+}08, KEH^{+}02, KF03, KKS04, LCC09, NINAT^{+}07, OFB08, Sch00, YCW^{+}09, YYW07, ABÅ04, BED02, CLF^{+}09, CLA^{+}00, DB06, EK06, GP06, GAS04, HEP^{+}02, HMSM06, HG08, KZY09, KP05, KFNH08, KEB04, KK08b, KS08, KOFF09, KF02a, LEK07, LXW^{+}09, LHP01, LLZL09, LLW^{+}09, MSF^{+}08, MS04, NCO^{+}05, NLL^{+}09, PJB^{+}07, QLHL09, RGG08, TKS^{+}01, TLKT00, Tot04, VGDSU08, WFHP01, WHP02, WHF08, WX09, ZHH09, AGI^{+}07, GCD^{+}08, KVS^{+}06, ZCS04]. predictions [BS01, BLB09, CP08, Ruv07, Van02a, ZLD09, vEMK01, vE01]. predictor [Kol04]. preface [FA01b]. preferences [GSB09, KK09, LKJ^{+}04]. preferred [DV02]. preliminary [KMH02, PMC^{+}08]. Preprocessing [SHM04]. prerequisite [WHF08]. presence [LZA02, RAGLL09a]. present [GR07]. Presentation [Rud05a]. preserving [QNF09]. pressure [Car02, MTB09]. pressures [TK08]. primary [HB09, JIK09, KBN02]. primitive [MV06]. principle [GJL^{+}08, PRS04, ZDS^{+}05]. principles [CS01, EBL^{+}08, GD09, HZX04, Hua09b, KK08c, MLJ03, TK08, VP08, WLX^{+}05, WZZ^{+}09, WD08, ZXYF09, ZHMW09]. prion [IIK09]. priori [SPDS01]. prismatic [WL09b]. probabilistic [PJB^{+}07]. probabilities [DP04]. probability [CFS03, DLW06, GCDL^{+}05, Kni00, SK09, SCS07]. probe [CVR08, DMLI05, TH02, VSW^{+}03]. Probing [PAT^{+}09, WMGK07]. problem [ABBC01a, ABBC01b, Ano06a, Bof01, CCL06, HLTLP09, Qua01, ST06, TKH07, XOW^{+}00]. problem-size [HLTLP09]. Problems [You11, ABWT09, Mat03, Vis02, Woo01]. Procacci [Ano06b]. procedure [AM09, BR03, CA07a, DLSVY00, GP06, KBT03, RS08, SSL02, SMM^{+}08, YÇBM00, Zho06]. procedures [GT03, HSMT04]. process [BZL05, LGB^{+}09, ML00, Pac06]. processes [Che01, GG09, KEM08, LDTS07]. processing [AGI^{+}00, AGI^{+}07, FEV^{+}09]. processor [LEV^{+}09, Yas08]. processors [SPF^{+}07]. Producing [KBN02]. product [SFR07, YLW^{+}08, YLWL09]. production [YQQH09]. products [KYFW07, LZ05b]. PROFASI [IM06]. profile [Ber03, CCB04, CCP04, GB02, ONHN00, Zho06]. profiles [AHK02, CMBC08, OD09, YXC^{+}07]. program [AJ03, BBM^{+}09, BAH^{+}02, DRMD03, GRCD01, Gly06, GM04, IS07, Kli01, KWK^{+}00, MP03b, ME06, PPXP01, PRJ02, QSS01, RMHK03, SFRS01, SMZW05, TRS02, UIHN09, VB07, VKCK09, Zer08, BBC^{+}05, BKS02]. programmable [Gan09]. programming [SPT07]. programs [CCD^{+}05, KS08, MBP09, SH07]. projection [FS00b, GKH05, GY06, Qua04, TKN^{+}08]. projector [MOP^{+}07]. projector-augmented [MOP^{+}07]. prokaryotes [WHH^{+}06]. prolapse [HdMdS05, HdS06, HD06, TW03]. proline [BISB02, KK09]. promising [JRJ01]. promolecular [Leh06]. promolecule [MS00]. promoted [SBG09b]. promotion [KMM07]. propagator [SVT09]. propanal [RR05]. propanone [RR05]. propargyl [LMK01]. propellanes [PAS07]. propenal [FDSA00]. propene [BS03]. properties [AB00, AEE^{+}03, ÁCD^{+}03, Ara04, AZS^{+}04, BG03, BZP09, BT00, BSOB05, BACJCT01, CMJ08, CDGS09, CDS09, CPDZH08, CLC09, CVR08, CZFH07, CDD^{+}02, CHA^{+}07, CRSB03, CTFC08, CMA^{+}08, DD08, DXW08, DWNB01, DVRP^{+}03, DD00, DPM09, DSS03, DHW^{+}07, EM03a, EM03b, Fau01, FTLV01, GKRG08, Hua09a, HJCP01, JPF^{+}00, JWB05, JT08, KHY00, KLH^{+}04, KJP^{+}07, KCL00, Kri09a, Kri09b, KGD06, KPZK06, KK01b, LTF^{+}07, LWLS07, LC06, LFR07, LLZL09, LMRVFH^{+}09, MV06, MM02, MA09, NA06, NINAT^{+}07, NAT07, OBBS05, OS08, PMB04, PK04, PBF07, PTC01, PSS^{+}04, POJ01, RKA^{+}09, SBJ08, SRK^{+}00, TZX01b, TZX01a, Tor02, TDK07, UM03, VB09, VKCK09, VP08, WLX^{+}05, WM06, WCL05, YFR05, ZY01, ZXYF09, ZWP08, ZX08, ZSK07, ZMH^{+}09]. properties-based [VB09]. property [BAÅ07, JLHF03, NLL^{+}09, PSCD^{+}09]. propylene [QZL^{+}04, RR05]. propynyl [Lee09]. prosthetic [ATBLS04]. protease [BWE05, CLXC02, DLG00, LZ05b, NLL^{+}09, SPT^{+}03, SVV^{+}08, WHF08]. protease-inhibitor [SVV^{+}08]. Protein [LEK07, NCO^{+}05, PJB^{+}07, ADM^{+}06, AG00, AHGK09, BED02, BRDC02, BMLV04, BS01, BSP06b, BS05, BSH07, BLMS08, CCC03, CLXC02, CLC03, CLWL09, CLS^{+}09, CIB05, CLH^{+}07, CKR08, CLF^{+}09, CJDK09, yCkHmY08, CRH^{+}07, CPUGD09, CSRST04, DHM^{+}03, DPRR05, DB06, DB02, EBAN07, FOL^{+}04, FC06, FKM^{+}06, FKM^{+}07, GLD08, GHH07, GL04b, GC04, GDPCPU07, GdSuM^{+}07, GHMP03, GKK07, GZM09, GB04, HEP^{+}02, HFS^{+}07, HP01, HS01, HM06, HLTLP09, HLM05, HLT^{+}05, HZ06a, HZ06b, HG08, HW09, HP05, ILKR09, IM06, IIK09, IT03, JS07a, JMD^{+}02, JIK09, KFB05, KFNH08, KLS02, KCL06, KHF^{+}09, KK01a, KIFK07, KH05, KF08, LFBSK07, LHJ^{+}06, LJKL08, LV08, LZKT04, LXW^{+}09, LL07, LW06, MFB04, MKT04, MLG04, MH09, Mei02, MWE02, MLL08a, MHT01, MPF00, NMAT01, OFB08, OFIK09, PB06, PHR^{+}05, PC00, PFC03]. protein [PSHP08, PMM05, PB02, PF06, PNG08, QLHL09, RI08, RSER09, RLP08, RS08, RK05, Ruv07, SHM04, SLC^{+}09, SWM04, SWV^{+}05, SN06, SR09, SMG09, STCJ08, SL06, TLKT00, TGD05, VW00, VW04, VGO^{+}07, VGDSU08, VGGMM05, VZM^{+}08, WS05a, WS07, XZZ04, XSHC06, XLC08, XWC09, Yan04, YL06, YFS07, YPNE09, Yos02, ZP03, ZGFL01, ZS04, ZZ08, ZTS09, ZM06, dSR08, HLC09, PMB04, ZZTS09]. protein-DNA [PSHP08]. protein-environment [HFS^{+}07]. protein-ligand [LXW^{+}09, RK05]. protein-tyrosine [AG00]. Protein/solvent [PMB04]. proteinogenic [IKYM09]. Proteins [LMH^{+}09, AG03, Ano06c, BBHD04, BCP03, BHH^{+}09, CR02, DWNB01, DMN03, DR07, DV02, DJ04, DJB02, DWC^{+}03, ES00, ENM^{+}04, FNP^{+}06, GAS04, HB09, HHHS01, HM02, HS01, HHW^{+}03, HL08, HJCP01, Ike04, IDMC09, IN01, KSB^{+}02, KT02, KKS04, LR03a, LHJ^{+}06, LKA01, MK02, MSH^{+}06a, MZL08, NAT07, OS06, OSHS03, OM04, PB04, PMB04, PRJ02, RGZM09, RON02, SL09, SPL^{+}02, SHBD05, SHSF05, SMV^{+}09, VBS09, WZW^{+}06, WM06, WS05b, WHH^{+}06, XZ05]. Protocol [AGI^{+}00]. Proton [SRB06, AGK03, BA03, BA04a, CXZ^{+}09, FDSA00, FO08, GWM08, HFHL06, LLM08, LMGO^{+}09, LB05, MA05, PGG06, PCS04, SM06, WFHP01, WHP02, XKG^{+}05, ZCS04, dSGCG00]. proton-coupled [CXZ^{+}09]. protonated [CPDZH08, ZDS^{+}05]. protonation [Bac05, CG05, DHM^{+}03, HP05, KYFW07, WHF08, XZ05]. protoporphyrinogen [WZY04]. prototype [Ang09, CS01, ASDP^{+}06]. prototypes [SSS^{+}09]. proximity [Agr03]. pruning [TCSM03]. pseudo [LL07, VDM06, XSHC06, XLC08]. pseudo-Jahn [VDM06]. pseudofolding [VGDSU08]. pseudoknots [DP03, DP04]. Pseudomonas [NYK^{+}09]. Pseudopericyclic [LFS^{+}07]. pseudopotential [FMAMVK06, LK03, VW03, vW06]. pseudopotentials [PSS^{+}04, PSMB05, SMD02]. PSI3 [CSV^{+}07]. psoralen [NBTN04a, NBTN04b]. Pt [DMN05, LWK08, LF02, RD00]. PtCl [LF02]. PtF [LF02]. PtH [LF02]. Pu [Han01]. Publisher [Ano04a, Ano04b]. pump [CVR08]. pump-probe [CVR08]. PUPIL [TdMSD^{+}08]. Pure [WG02, Rud05a, SDCG02, SCP08]. purpose [DGI^{+}08, JGVF05, KAK^{+}09]. Putting [MDI04]. pVDZ [Wib04]. pVTZ//MP2/6 [WD04]. PW [EBL^{+}08]. PyFrag [VGB08]. pyrazine [LWX07]. pyrazole [DMC05]. pyrazoline [LLKC06]. pyrazolyl [HT05]. pyrene [HIA03]. pyridine [CHA^{+}07, HT05]. pyridines [WRP^{+}06]. pyridoxal [LDY^{+}08, PMM05]. pyrimidine [LWX07, XWXC08]. pyrimidinyl [WJX^{+}08]. pyrolysis [KKH^{+}07, XKKL03]. pyrope [ZWTP^{+}08]. Pyruvate [ČJPZS08]. pyVib [Zer08].

Q [BS08, KWK^{+}00, WHG^{+}07]. Q-Chem [WHG^{+}07, KWK^{+}00]. Q-Dock [BS08]. QCISD [ZKZ^{+}07]. QCT [DPM09]. QM [CGBF05, MPF00, AGK03, AST06, AB09, CR09b, CG05, FAR02, FMSA06, FSFK05, GWM08, GWM^{+}00, HHBH00, HBM06, HNR08, HRR05, HT03, HTN03, IV04, IvSV06, ITS05, ITS06, KHF^{+}09, KBLP09, KPR04, Kri08, Kri09b, LLL03, MBM^{+}00, MK02, MSH^{+}06a, MG00, MLJ03, NGTB03, RG02, SURG06, SBG^{+}09a, SN06, SMM^{+}08, SVV^{+}08, THHN01, TdMSD^{+}08, VMF^{+}03, WCC08, WHG^{+}07, WC08, ZWZ09]. QM/FE [AGK03]. QM/MM [CGBF05, MPF00, AST06, CR09b, CG05, FAR02, FMSA06, FSFK05, GWM08, GWM^{+}00, HHBH00, HBM06, HNR08, HRR05, HTN03, IV04, IvSV06, ITS05, ITS06, KHF^{+}09, KPR04, Kri08, Kri09b, MBM^{+}00, MSH^{+}06a, MG00, MLJ03, NGTB03, RG02, SBG^{+}09a, SN06, SMM^{+}08, SVV^{+}08, THHN01, TdMSD^{+}08, VMF^{+}03, WCC08, WHG^{+}07, WC08, ZWZ09]. QM/QM [AB09, HT03, SURG06]. QMCF [PHRR08]. Qmd [KMH02]. Qmd-plot [KMH02]. QMPFF3 [DGI^{+}08]. QMQSAR [DMLI05]. QSAR [DHW^{+}08, DHW^{+}09, SGPS09, CGMPT^{+}08, CMBC08, CRGN07, DMLI05, DMC05, GDPP08, HSMT04, HMMS09, LLL^{+}08, LJZ^{+}07, LSY02, MRS09, PS09a, SJJ^{+}04, TCSM03, VB07, VB09, VGDSU08, XYN^{+}06, ZNLL07]. QSAR-analysis [VB07]. QSAR/QSPR [TCSM03]. QSPR [CDGS09, CDS09, CDGS09, GS08, HM08, TTBM09, TCSM03, ZNLL07]. QSPR/QSAR [ZNLL07]. QTAIM [MGMM07b, RKA^{+}09]. quadratic [ABBC01a, ABBC01b, Bof01, HG08, Qua01, ZHH09]. quadrature [CG06, DBS08, GC03]. quadrilaterals [GKK07]. quadrupolar [CMA^{+}08]. quadrupole [HLLN06, HK08a, HK08b]. quality [BG03, CMJ08, EM03b, FKZ09, JBJB00, JJB02, SSS^{+}09, TSSGS07]. quantifying [GT03]. Quantitative [Mit01, WZY04, YNZ^{+}08, BAÅ07, CDGS09, CDS09, DHW^{+}08, DHW^{+}09, Gra07]. quantization [GLMV09]. Quantum [AVB00, BWM^{+}09, BISB02, BS06, DMN05, ECA06, ESM06, EDAJ04, FHRR07, LBT07, MBL^{+}00, MA05, NRKH02, PM02, RM07, RON02, SC01, SS05, TLOG00, VHRR07a, ZMH^{+}09, AGMPRG^{+}08, AGO^{+}02, APG05, ATH^{+}03, AGSFAL05, AGSFA^{+}05, BSJ01, BPCD07, CLP09, CDGS09, CZFH07, Con02, CKW09, COL^{+}06, DBS07, DBM03, DA01, DWC^{+}03, EBD^{+}01, FCK^{+}08, FÁ01a, FAB^{+}00, FKFG08, FR06, FKU^{+}05, GAIMVB01, GVATG03, Gog08, GBB07, GGLR00, GS04, HM08, HHP04, JJH01, JČHS07, KSB^{+}02, KFNH08, KJVW08, KHY00, KZRO03, KLM^{+}09, LX07, LHP01, MFB04, MP03a, MGCA07, MKT04, MR09, MBP09, Mat03, MC06, MPF00, OYH05, OKH^{+}02, PG04, PHKG07, PDS01, PV07, RP07b, RSE07, RGP^{+}07, SF07, SH07, SS00, Sau04, Sch00, SFRS01, SBB02, TCR^{+}02, TT02, VHRR07b, Vis02, VKCK09, WS05a]. quantum [WOC^{+}03, XYN^{+}06, XZZ04, XLZ08, ZCZ03, ZAT07, ZSK07, SB08, CGBF05, DSS03, KBL08, PFB05, SCS07]. quantum-chemical [DA01, SFRS01, VKCK09, XYN^{+}06]. Quantum-connectivity [EDAJ04]. QUantum-regions [SB08]. quartet [MSBS01]. quartet/metal [MSBS01]. quartets [MSBS01]. quartic [SAS05]. quartz [ZWPR^{+}04]. Quasi [AGI^{+}07, NUH02, AGI^{+}00, ITN^{+}05, VMA03, YH07]. quasi-canonical [ITN^{+}05]. Quasi-degenerate [NUH02]. quasi-flexible [AGI^{+}00]. quasi-relativistic [VMA03, YH07]. Quasirelativistic [HWFN01]. quaternary [CW02, SO07]. quaternions [CSD04, CSD05, Kne05]. Quick [LMV07]. QUILD [SB08]. quinolines [KS05c]. quinoprotein [JJH01]. quintet [GWL07].

R [Bof01, CPJ00, LZZC09, Lip00, Qua01, ZY01, LZZC09, ZPL07]. rack [OCP02]. Radial [GC03, ESP04, Kau07, Kni00]. Radical [XDS06a, AVB00, BL06, CUS00, CU01, CUSS03, CXZ^{+}09, GSB09, HIA03, JDWS06, KOML08, KKMMS04, LC07, LMK01, NSB08, OO04, gThDjL^{+}01, WDWS06, WDS06, WDZS07, WyLG^{+}09, WLZ^{+}07, WLL^{+}03, XDS06b, YLWL09, ZLLS04a, ZLLS05, ZLLS06a, ZZL^{+}09, CXZ^{+}09, QZZZ03]. radical-molecule [ZLLS06a]. radicals [BE07, Dib05, Lee09, WLLS04, WDZS07, WSC09, YLW^{+}08, YLWL09, ZM03]. radii [OCB02, PML03]. radon [HD06]. Raf [GC04]. Ramachandran [SDL^{+}09, GSB09, HHP04, PFJ^{+}03]. Ramachandran-type [SDL^{+}09]. Raman [Bou01, LC09, NRKH02, OBBS05, Zer08]. RAMSES [BMRDB01]. random [CY09, CY13, CA04, HXLS09, JS07a]. randomized [LFKL00]. range [CEP07, GPK05, Gri06, HGMB04, IZA06, JPCA08, KSS08, MN02, RP07c, RLP08, San01]. range-separated [JPCA08]. rank [RP07c]. ranked [TBSM09]. ranking [KSM05]. Rapid [GGA00, RSN^{+}02, BH03, Gra07, KMH02, KC01b, WS05a, PABK03]. Rapidly [KF02b, Zho06]. rare [LZZC09]. Ras [GC04]. rate [Chu07, GGB07a, GGB07b, JHZ09, MGLL03, NSU^{+}02, SLL^{+}04b, SFRS01, UCT^{+}03, WLLS05, WLL^{+}07b, WDX^{+}02, ZP03]. rates [HG08, JIK09]. Ratio [LR06, KBB09]. Rational [Chi03, BSP06a, Ham07, VGGMM05]. rationalized [Bac05]. rattle [FS98, FS00a]. raw [RON02]. ray [HSWN01, HN02, WKYU01]. Rb [GLRL02, GWL07, HRR05]. Re [LJKL08]. Re-examination [LJKL08]. Reaction [CU01, JKM08, JDWS06, KKJH08, MGG06, Qua04, WCHW09, ABBC01a, ABBC01b, BAL^{+}01, Bie04a, Bof01, BS03, CUSS03, CG05, DRAS04, DRAS05, FG03, GZL02, GF08, GWM08, Gog08, HLLS05, Hir08, HLSH05, HTN03, IN01, JJH01, JHZ09, LMGO^{+}09, LMK01, LL01, LFZS04, LW04a, LDT^{+}02a, LDT^{+}02b, LWY^{+}09, LLL07, MGLL03, MG00, MS04, Mui05, NSU^{+}02, OON01, PGNG03, PGRRNG03, Pom04, PS03, PMM06, wQZsLyZ02, QZL^{+}04, Qua01, Qua07, RSN^{+}02, RD00, RWBH09, RJLR06, RR05, SLL^{+}04a, SLL^{+}04b, SRE08, TYN05, gThDjL^{+}01, TGLL07, TMBM02, UCT^{+}03, VGB08, WLL01, WDWS06, WLL^{+}07b, WCW08, WHG^{+}07, XDS06a, YLZ08, YQQH09, ZZL04, ZLLS04a, ZWL^{+}05, ZLLS05, ZLLS06a, ZLLS06b, ZZW^{+}07, ZZL^{+}08, ZZL^{+}09, dRLMS00]. reaction-diffusion [Bie04a]. reactions [AM07, BS03, CUS00, CFD03, Fie02, GAIMVB01, GMA04, GLH^{+}08, GGB07a, GGB07b, HFHL06, HSWW00, JM07a, mJlZsLyL07, mJlZyL^{+}08, JHZ09, KYFW07, KIM^{+}09, LL00, LDC^{+}07, MBL^{+}00, NTH00, OY01, OY03, RNG03, Rao00a, RC04, RY09, Rud05b, Rud05c, Sch03, Sie01, SSB07, TT08, TCR^{+}02, UNM^{+}01, VBGL^{+}00, WLLS04, WLLS05, WDS06, WLL07a, WyLG^{+}09, WDX^{+}02, WLL^{+}03, XLL^{+}02, XDS06b, YT04, YLW^{+}08, YLWL09, ZLLS04b, ZKZ^{+}07]. Reactive [LLM08, Hir08, MMMY07]. reactivities [HTSR04, YIN03]. reactivity [Ano06a, BM08, Bor03, BL00, CN05, FZL07, GTC06, MTB09, ST06, Tie09, TSSSG08]. reagent [DHW^{+}07]. Real [Woo01, You11, PBF07, PBF09, Sch04, THHN01]. Real-World [You11]. rearrangement [NSB08, PA05, ZGZX07]. rearrangements [LLKC06, YTY07]. Reassessment [DBM03]. Reassociation [DWNB01]. recently [RG08]. RECEP [KC01b]. receptor [DLRZ09, FKU^{+}05, FKM^{+}06, FKM^{+}07, GCD^{+}08, HMK02, KBK^{+}01, MHL^{+}09, SBG^{+}09a, TFN04, TJE03, WS02b, XWC09]. receptors [CW02, FTLV01, NHH05, YKK09]. ReCO [HT05]. Recognition [UNHYT06, AGI^{+}00, AGI^{+}07, AGMPRG^{+}08, BR07, CW02, DGHR02, EKB02a, EKB02b, GdSuM^{+}07, KEB04, MSF^{+}08, PSHP08].

Reconstructing [BBP09]. reconstruction [Adc04, GKK07, KLS02, RS08, TGD05, WG02]. recoverin [LGB^{+}09]. red [McD08, SRK^{+}00]. red- [McD08]. redesign [GLD08]. redistribution [ZY01]. RedMD [GSDT09]. redox [GK09]. Reduced [BR04, BSOB05, OBBS05, ABBC01a, ABBC01b, BMLV04, Bof01, CNN07, CP08, DLD^{+}02, Ell07, GSDT09, HP01, Qua01, RS08, WEE01]. Reduced-size [BSOB05, OBBS05]. Reducing [PRSMM02, SSL02, SY09]. reductase [CFER04, CGB^{+}09, CBC^{+}08, DBS07, GGLR00, HLLN06, LPP06, PCS04, PFR04a, PFR04b, TMBM02]. reduction [CCCJ09, DBS07, DMN05, HLTLP09, LRWG03, Mck07b]. reductive [PS03]. reevaluation [Kle03]. Reference [ZZ08, CF04, CFC^{+}08, LZ05a, NUH02, OV03]. Refinement [HB09, Ruv07, BHW00, MM05]. refinements [GPK05]. Refining [CLWL09, SB01]. refractive [YYW07]. regarding [KZY09]. region [ABWT09, Ama02a, HHBH00, WEE01]. Regional [TKH03, NTH09]. regions [HYT05, SB08]. regioselectivity [AVB00]. Registering [GBL^{+}05]. regression [DLWV07, Gol09, GS08, LCC09, SY09]. Rehybridization [AM07]. related [ALC08, ACLD03, CFD04, KC01a, LXW^{+}09, LWH06, LCDA03, LCGA03, LCA03, ML00]. Relation [SM08a, DVRP^{+}03]. relationship [DHW^{+}08, DHW^{+}09, JPCA08, KWK^{+}01, WLX^{+}05, KWK^{+}02]. relationships [BAÅ07, CDGS09, CDS09, CPUGD09, JLHF03, PSCD^{+}09, WZY04]. Relative [SWV^{+}05, BLB09, CG05, MML^{+}06, MRS^{+}07, RSE07, ZOJ^{+}06]. Relativistic [FHF^{+}01, NYH02, NSO^{+}07, SNM^{+}06, SMD02, WTKM06, YH09, ASS^{+}02, BBI^{+}09, Dya02, GHLK^{+}02, GPSP06, HdMdS05, HdS06, HD06, LF02, SH02, Van02b, VMA03, WL02, YH07]. relax [GFS05]. relaxation [BRDC02, HS01]. relaxed [AEE^{+}03, CA07a]. relevance [Ano06a, MGCA07, ST06]. reliability [IB04, LKW04]. reliable [BE06, WHF08]. remove [LZ05a]. removing [PCS04]. reordering [TVL^{+}03]. reorganization [FZL^{+}06, KMM07]. repair [Pin03]. reparameterization [RFSS06]. repeat [NK01, NL08]. Repeated [KH01]. Repeated-annealing [KH01]. replica [FSM09, FGR07, GLP08, NCO^{+}05]. replica-exchange [FGR07, NCO^{+}05]. Reply [Bof01, CPML08b, WM12, CSD05]. Representation [GPK05, BB08, CKR08, CF06, JIK09, LW04b, LXZ06, LW06, RLR^{+}04, WEE01, hYDN^{+}08]. representations [BMLV04, LAR^{+}03, LR06, RS08, SN00, YNW05]. representative [YLL^{+}09, YXL^{+}09]. reproduce [VBS09, WS05b]. reproducing [MFB04]. repulsion [COL^{+}06, Kri09a]. repulsions [HGMB04, PBF09]. repulsive [BDW00, CFC^{+}08]. requirements [AM06b]. research [JLHF03, PGH^{+}04]. residual [RI08]. residue [MH09, NBTN04b, PMM06, NBTN04a]. residues [CFS^{+}09, DHW^{+}07, HJCP01, JPF^{+}00, OS06, UNHYT06, XLT07]. resolution [BS05, BS08, CDGS09, DSR^{+}07, GL04a, HXLS09, Nee03, WMRW^{+}01, WS02b]. resolution-of-the-identity [DSR^{+}07]. resonance [BM07, FLGW00, MAF^{+}07, WZXY07]. resonances [LMB08, PF06]. RESP [WCK00]. respect [QCK01, QCK02]. response [HG08, OFIK09, vGGB00]. Restrained [SRB02, WCK00]. restraint [LI07, LHI09]. restraints [BS08, HWTL03]. restricted [BdPRMAI00]. restrictions [KŠB09]. results [CSD05, LKT04, PFJ^{+}03]. Retardation [HP04]. retention [RC04]. reticulum [HLB09]. retinal [BL05, LFEdL06, MSH^{+}06a]. Retrieval [CVR08]. reuptake [FPG^{+}06]. reveal [DLRZ09]. revealed [HW09]. reveals [Pin01]. reverse [AJNG01, ML00]. Reversible [DvG00, Kol04, NHN06]. Review [Bic09, CvG08, Lip00, Sta00, Woo01]. Reviews [LB99, Sta00]. Revised [ATM^{+}07, SBB02]. revisited [ASY01, CVVB04, PCS04]. Revisiting [GPSP06, JPCA08, LN01]. Rg [ZXY03, ZXY03]. RGF [HQ02]. rhenium [SBH02]. RHF [EA08, JPF^{+}00]. rhodamine [VSW^{+}03]. rhodium [GLH^{+}08, LL00]. rhodium-catalyzed [GLH^{+}08]. Rhodopseudomonas [IN01, OON01]. rhodopsin [CEP07, YKK09]. rhodopsin- [YKK09]. ribonuclease [KSK00, WOC^{+}03]. Ribonucleotide [CFER04, HLLN06, PCS04, PFR04a, PFR04b, TMBM02]. ribose [SA07]. ribosomal [SB01]. ribozymes [MMMY07]. rich [CZ05, LWK08]. Rigid [SM03, DPRR05, Din00, ECA06, FS98, FS00a, Ike04, Leh06, LV08, KP05]. Rigid-body [SM03, Ike04]. Ring [ZSE08, BE09, CDPL09, DC02, DLSVY00, FJ08, RPNJ07, SRE08, ZW09]. ring-structured [DC02]. RISM [MH08b]. rival [DDVD09]. RM1 [FBLO08, RFSS06]. RMSD [Kne05, CSD04]. Rn [Wei08]. RNA [AM06b, DW08, GdAcV^{+}07, LhWX07, LCSZ09, LOL^{+}08, MB00, Mak08, RTG00, SB01, YNW05, ZZTS09]. RNA-ligand [AM06b]. RNACluster [LOL^{+}08]. RNase [RWBH09]. ro [LN01]. ro-vibrational [LN01]. robust [GS08, HEP^{+}02, YK00]. ROCR [CPJ00]. rod [BCIB05]. rod-like [BCIB05]. Role [BCF^{+}09, CPJ01, CFS^{+}09, CPFL02, Ruv07, SVV^{+}08, ZSC05, BY06, CDS09, CFER04, Kau07, TFZRG01, VBGL^{+}00, YT04, YTY07, YJF06, PMM06]. Roles [ALC08, IN01, NYK^{+}09]. roll [FS98, FS00a]. room [TD08]. Roothaan [TW03]. roots [BdPRMAI00, Nil09]. Rotamer [HLTLP09, GHMP03, KBLP09, LFBSK07, SMG09]. Rotamers [LMH^{+}09, SHM04]. rotation [CMLS05, COMR^{+}04, DHF^{+}05, DBM03, HFSD03, HK08c, LHI09, LZ05a, MGLL03, OMNH08, PBF09]. Rotational [CSD05, BVW04, KBN02, TS05]. rotations [IR03]. rotors [WR07, WFR08]. rough [Pan07]. rough/fractal [Pan07]. roughness [PHJ^{+}08]. routes [GGGLL05]. routine [Kli01]. routines [AT02]. row [AD00, BP03, BGJ01b, JGH00, LK03, LK04, RRP^{+}01, YTH01]. Royal [LWW^{+}06]. RS [LAR^{+}03, EK06, ELK^{+}09, KEH^{+}02, LRWG03]. RS-HDMR [LRWG03]. RT [TYO^{+}02]. Ru [ZWS^{+}02]. rule [GR07]. rules [AMR04]. runs [EL07]. ruthenium [FKŠ^{+}09, PHKG07]. Rydberg [PRSMV08, ZM03, dSVA^{+}09].

S [BSB05, Bic09, Gog08, HKHN08, JJK^{+}00, KYFW07, MGLL03, Mck07a, Mck07b, WWS07, ZJM^{+}07, ZY01, XZ04, ZALMG03, DLD^{+}02, HTN03, MVLG06, MG00, RC04, RFSS06, SN00, WDS06, YT04]. S-network [MVLG06]. s-tetrazine [XZ04]. SA [GC04, GWS^{+}02, WHF08]. SAAP [IT03, IT03, IKYM09]. SAC [DHM^{+}03, HFS^{+}07, HKHN08]. SAC/SAC [DHM^{+}03]. saddle [DLD^{+}02, GMA04, HQ02]. SAFE p [AVS09, VVS07]. Sakurai [TKN^{+}08]. Sakurai-Sugiura [TKN^{+}08]. salts [JHMB^{+}09, KWK^{+}01, KWK^{+}02, JHMB^{+}11]. Salvetti [IKN08]. Salvetti-type [IKN08]. SAM [WC08]. SAM-dependent [WC08]. Sammon [FKZ09]. sampled [IZA06]. Sampling [LZKT04, BHG03, CN03, CIB05, CY09, CY13, CV09, CvG08, CEP07, DDVD09, FKZ09, GT03, HKMS01, IS03, JW06, KH01, KM07, LKW04, Mak08, MH08b, MST^{+}08, Nak02, NA06, RNG03, Rap06, SD09, SMG09, TS05, YL06, ZA07]. sandwich [JD09, RPNJ07]. sandwich-like [JD09]. sandwiched [MHS05]. sarcoplasmic [HLB09]. SARS [LZ05b]. SARS-CoV [LZ05b]. SASMIC [EA06]. Sason [Bic09]. Sb [LS08a, XB08, XK08]. Scalable [PBW^{+}05, VSK^{+}04, Ano05b, KKC05, Sha05, VGO^{+}07]. scalar [GPSP06, KBLP09, MP03b]. scale [DMN03, JO02, KK01a, MH09, MHW04, MPF00, ME06, Nak02, NA06, RRS07, SSL02, TYO^{+}02, WCF04, WS07]. scale-transformed [Nak02, NA06]. scaled [CN03, JSHG07]. scaling [AL01, ARL01, Con02, FR06, GGLR00, GY06, KLM^{+}09, LMJ02, OS06, RS05, SSB^{+}03, SHSF05, SKDO08, SP05, TCR^{+}02, ZWZ09, vGGB00, vdVGDM00]. scanning [HMK02, MFR07, SMGE08, ZM06]. scattering [BACJCT01, Est07, HSWN01, RMHK03, WKYU01]. SCC [ECM^{+}03]. SCC-DFTB [ECM^{+}03]. SCF [JHPRSM^{+}05, PFJ^{+}03, PVdJB00, SAM06, VZVG06]. SCH [ZZW^{+}07]. scheme [Bac04, FOK^{+}04, IS03, JCA^{+}02, JVVK09, LMV07, Maz08, MSH^{+}06b, RKA^{+}09, SN00, SHH07, WS05b]. schemes [Bac04, Bac05, Bac07, PRS04, SPDS01]. Schleyer [Lip00]. SCI [TY03]. SCMP [FÁ01a]. SCMP-NDDO [FÁ01a]. ScO [LMCD09]. Scope [LFEdL06]. scoring [BS05, GLD08, HZ06a, HZ06b, Mue01, NMAT01, OFB08, Ruv07, SBG^{+}09a, VVS07]. SCPF [Maz08]. screen [MHW04]. screened [RDM^{+}08, VVS07]. screening [DHW^{+}00, FZL^{+}06, KV00, KSM05, LFKL00, LZ05b, PRDS08, SHSF05, YOB^{+}08]. SCRF [CCT^{+}03]. SCUD [LZ05a]. ScX [WWS07]. SD [WLLS05]. SDCI [BMB07, PRSMM02, PRSMM03]. Se [HKHN08, JJK^{+}00, WWS07]. search [AM09, BR07, BMTSC01, CSJ01, CA04, GLD08, HHG^{+}09, HXLS09, HM06, IZA06, KK01a, LFKL00, MGJAARC00, NL08, OGH05, OM04, Pul05, RHL09, Sau04, SE07, SE08, WK01, ZZ08]. searches [CZB07, YXL^{+}09]. Searching [SPT07, STC^{+}08, CvG08, Nak02, OYH09, SCC04, SYC08, YCS07]. Second [BC06, FS04, MO01, AGK03, DSR^{+}07, DOSG06, FO08, IN08, JSHG07, LK04, QTdG^{+}08, QCK01, QCK02, Rud05a, Rud05b, Rud05c, YTH01, YH09]. second- [LK04, Rud05a, Rud05b, Rud05c, YTH01]. Second-order [FS04, DSR^{+}07, IN08, JSHG07, QTdG^{+}08, Rud05c, YH09]. secondary [CLC03, CLA^{+}00, DW08, DP03, GdAcV^{+}07, IGNH03, LhWX07, LCSZ09, LW06, LOL^{+}08, LLL07, MHT01, WPH^{+}07, YNW05]. Section [Ano01c, Ano04b]. sections [MY08b, MY08a]. seedling [PS03]. segment [YS00]. segmented [CGSdST06]. segments [BTLP03, GAS04, KF02a, YMT04]. segregation [Sza08]. Selected [BMB07, LSAS01, Mat03, PRSMM02, ZOJ^{+}06]. Selecting [HXLS09]. selection [HW09]. Selective [TCSM03, XPW09, AM06b, MHL^{+}09]. selectivity [BSP06b, GLRL02, OO08, ZZvRSC08]. Self [ZA07, BA08, BWI^{+}02, ECM^{+}03, NUH02, NTH09, NL08, SH07, TKH03, VTT^{+}08, WM04, XL02]. self-associative [NL08]. self-consistent [BWI^{+}02, NUH02, VTT^{+}08, WM04, XL02]. self-contained [SH07]. self-interaction [NTH09, TKH03]. Self-organizing [ZA07, BA08]. semi [BBHD04, UKN04, ZNLL07]. semi- [UKN04]. semi-empirical [ZNLL07]. semi-implicit [BBHD04]. Semiautomatic [PRJ02]. semibullvalenes* [HWGB01]. semiclassical [KM07]. semicore [HZ09]. semidirect [MBWP03]. Semiempirical [Gri06, LKT04, TT01, BUMCMRL00, BM00, Der09, DMLI05, FBLO08, FO08, GGLR00, HMOG07, IK00, JFG04, KSS08, KBT03, LMMW04, MAF^{+}07, MSH^{+}06b, MHT01, Nye07, RSSKB03, RCJ02a, RCJ02b, ŠBL05, TCR^{+}02, TBGRJ04, vdVGDM00]. semiempirical-DFT [Der09]. Semiglobal [DV02]. semimicroscopic [KK08b]. semirigid [CLP^{+}05]. sensitivity [HLSH05]. sensitization [Sha02]. sensitized [KS05c]. sensors [BBG^{+}04]. Separable [EA06]. separated [JPCA08, MLL06, WSM^{+}09]. separating [CN03]. separation [FBDG06]. sequence [AM06b, CCWH02, CKR08, CLZX09, Dya02, JIK09, LXZ06, LSW^{+}01, PRJ02, WHH^{+}06, ZLY07]. sequence-dependent [LSW^{+}01]. sequence-specific [PRJ02]. sequences [CP09, DLW06, DLWV07, Der00, JIK09, LW04b, LD05b, LW06, MCF07, PP08a, hYDN^{+}08]. sequential [TT05, ZGZX07]. SER [JPF^{+}00]. serially [KMA^{+}07]. series [CC07, KMH02, PDS01]. serinamide [PFC03]. serine [OBT09]. serotonin [HLC09]. serve [Mck07a]. set [ABF^{+}03, ALKH04, AHK02, BR04, BT00, BSOB05, BRV^{+}07, BRLS08, BRLS12, CMJ08, Che01, DMZT08, EL09, EKB02a, FZL07, FMPS08, GGT08, HdMdS05, HMSM06, IO08, JJK^{+}00, LFK05, MV06, Mas04, MLL^{+}08b, MC06, PSC^{+}01, PRKP05, Pen06, PFJ^{+}03, PSMB05, RRP^{+}01, SSB^{+}03, SHK^{+}05, TW03, VKP^{+}08, Var09, VKCK09, WMGK07, Wib04, WG02, ZWPR^{+}04]. set-up [GGT08]. sets [BY06, BSOB05, CRS05, Cul04, EA08, EdlVR^{+}03, GKH05, HdS06, HD06, IO08, KK08a, LST08, LTV08, MV06, NSO^{+}07, OBBS05, OVMV04, RLA01, RLRE01, RLER07, SNM^{+}06, VB03, WTKM06, Wei08]. setting [HP04]. setup [ZAT07]. several [KS05b, XLT07]. sevoflurane [TZX01a, TZX01b]. SG [CG06]. SG-0 [CG06]. SGB [GZL02]. SGB/NP [GZL02]. SH [Mas01b, MGLL03, Mui05, SSS^{+}09, WLLS05]. SH/ [SSS^{+}09]. SH3 [IGNH03]. Shaik [Bic09]. shake [KFD06, BL09, FS98, FS00a, KvGH01]. shake-up [KFD06]. Sham [Bou00, RRS07, SH02]. Shannon [LM03]. shape [BR07, PRDS08, WM06]. shape-based [PRDS08]. shaped [LWW^{+}06]. Shapelets [PRDS08]. shapes [BR07, KS02a]. shaping [HJCP01]. shared [Sim07]. sharing [BRS07, RS07a, RS07b]. SHARPEN [LMH^{+}09]. Shaw [Ano05b]. sheet [KF03, PP08a]. sheets [LLW02, PGC05]. shell [DSB^{+}02, FS02, FO08, HB09, PRSMM03]. shielding [CDL06, CDPL09, HWFN01, MC06, PFC03, WZXY07, ZPL07, ZXY08, ZLD09]. shift [Dra00, HP05, LFZS04, MA05, RG02, WPS02, XZ04, ZFL^{+}05]. shifted [McD08]. shifting [CPFL02, HRG07]. shifts [CPDZH08, FVB08, FHF^{+}01, FO04, HWFN01, HLLN06, Kle03, KKS04, LFS^{+}07, LKA01, VBS09, WFHP01, WHP02, WZXY07]. SHOP [YXC^{+}07]. SHOP-type [YXC^{+}07]. Short [TYN05, GPK05, HGMB04, IKYM09]. short-range [GPK05, HGMB04]. Short-time [TYN05]. shorter [MST^{+}08]. Si [BSB05, TK08, WZZ^{+}09, YHD^{+}06, CJS^{+}03, SURG06, WL09b]. sialic [UNHYT06]. SIBFA [PWHF^{+}03, PWHF^{+}04, ROG00]. SIBFA-LF [PWHF^{+}03, PWHF^{+}04]. side [DLHC06, ENM^{+}04, GT03, HFHL06, JPF^{+}00, KG02, LL01, MT03, MMLC05, PFC03, SMG09, VM02, XLT07, ZM06]. side-chain [ENM^{+}04, GT03, JPF^{+}00, KG02, MT03, PFC03]. sieve [PHH^{+}08]. sieves [LMV07]. SiF [LAT05]. sigma [JFG04, KMM07]. sigmatropic [LLKC06, LFS^{+}07]. SiH [ZZL^{+}09]. silastannation [WCHW09]. SiLi [XFF06, HXD08]. silica [SDCG02]. siliceous [LST08, LTV08]. silico [LLW^{+}09, MHW04, PHR^{+}05]. silicon [BSB05, HXD08, KZW^{+}05, KS01b, LB08, NBJ04, ZLJS03]. silsesquioxanes [JW00]. silyl [MGG06]. silylenoid [XFF06]. SiMe [XFF06]. similar [BR07]. similarities [HPP00]. Similarity [Leh06, LhWX07, ARL01, BPCD07, COS01, Con02, HM08, MBH^{+}02, PDS01, RSS09, YNW05, hYDN^{+}08, ZZTS09, dGWH01]. similarity-based [RSS09]. similarity/dissimilarity [hYDN^{+}08]. Simple [MO01, Ste04, ACLD03, Bac04, BLMS08, GRO^{+}03, GDV03, Gon07, ILB03, IT03, KS02a, LLW^{+}09, MCF05, SF05, BG00]. simplex [DV02, MCF05]. simplex-annealing [MCF05]. simplified [OYH05, WOC^{+}03]. simulate [LAEL01]. simulated [ADM^{+}06, AB08, CCP04, HPP00, RLP08, WM06, WG02]. Simulating [Fie02]. Simulation [FBDG06, WWL^{+}09, BBHD04, BVW04, BG07, BBM^{+}09, CCD^{+}05, CV09, CHB^{+}05, Dra00, EA06, EMP07, FHRR07, FEV^{+}09, FPN^{+}05, GS02, GJK00, HN02, HLB09, ITS05, IM06, IKYM09, JO02, JGVF05, KSY^{+}00, KB09, KEM08, KPR04, Kri09b, LMCD09, LEV^{+}09, MLG04, MMMY07, MVL^{+}05, MLCD01, MST^{+}08, ON07, OBT09, Pin01, PHH^{+}08, SO07, SL06, SDL07, TYN05, VHRR07b, WEE01, XKG^{+}05, YAÇ^{+}02, YTH^{+}07, ZALMG03, ZWTP^{+}08, ZL09b, ZSK07, OBBS05]. Simulations [FCP^{+}05, MZL08, ATMK03, Ano06c, BWE05, BRDC02, Bie04b, BSJ01, CLP09, CLWL09, CLC09, CCSJ00, CF06, CPC^{+}00, CEP07, CMD^{+}04, CBH^{+}03, DHF^{+}05, DLRZ09, DFGB09, Der00, DSS03, DWC^{+}03, ESM06, FGR07, FG02, FCP^{+}04a, FAB^{+}00, FC06, FKZ09, GL04a, GLP08, GWM08, GHH07, GS03, Gon07, Haf08, HB09, HGMB04, HHHS01, HH04, HM02, HPL03, Hin00, HFSD03, HTKG08, HTSR04, HTN03, HMD06, IC08, JNV08, JCL05, JZD^{+}09, KMH02, KFZ03, KM00, KKC05, KAK^{+}09, KvGH01, KH06, Kri08, Kró03, KBN02, LML^{+}00, LSO04, LGB^{+}09, LM03, LPB03, MB00, MFB04, MN02, MABM09, MBC08, MO09, MG00, NK01, NL07, Nil09, OO06, OR05, PRKP05, PHJ^{+}08, PB04, PMB04, PK04, PB02, PNG08, RPMP03, RSER09, RMHK03, SK09, SDL^{+}09, Sch04, SBG^{+}09a, SWR06, SR09, SDM02, VCM01]. simulations [VHRR07a, VP09, WL09a, WCF04, WZW^{+}06, XLZ08, YNZ^{+}08, YGLvG06, ZCS04, ZSC05, ZGFL01, ZWS^{+}09, ZWZ09, ZSK07]. simulator [JGVF05, KIM^{+}09, MS04, SO07]. simultaneous [DDVD09]. Singh [JVVK09]. Single [OV03, BG07, CV09, HSF08, IT03, IKYM09, LFZS04, WTKM06, XWL^{+}09, ZZvRSC08, Mak08]. single-family [WTKM06]. Single-nucleotide [Mak08]. single-sphere [LFZS04]. Single-step [OV03]. single-walled [XWL^{+}09, ZZvRSC08]. singles [IN08, WKYU01, dSVA^{+}09]. singlet [BLO^{+}02, CZ05, CG08, FG03, LS08a, OSA06]. singlet-dioxygen [BLO^{+}02]. singly [HHWG08]. Singular [FPG^{+}06, TBSM09]. Sir [Ano04a]. Site [CJW^{+}09, LLL07, AG00, CFR06, CFS^{+}09, CF04, CFC^{+}08, FPN^{+}05, GJK00, GS04, HFS^{+}07, HYR06, KSK00, KEB04, KZRO03, MDA08, NL07, NLL^{+}09, PMM06, SS05, SPT^{+}03, SFR07, TDH06, XLZ08]. sites [APG05, BSP06b, BSDM04, CLXC02, CLS^{+}09, DV02, FPN^{+}05, FSS00, GDV03, HM02, HN02, HLT^{+}05, MHJS06, PPXP01, SEKS09, SLC^{+}09, Tie09, Wou00]. six [GJK00, NL07]. six-site [GJK00, NL07]. sixth [CGB^{+}09]. sizable [CAG07]. size [BSOB05, EL09, HLTLP09, KS02a, KH06, NK06, OBBS05, OV03, YAÇ^{+}02]. sized [SHH07]. Slater [CVVB04, EdlVR^{+}03, GC02, KDG^{+}09, RLER04a, VB03]. Slater-type [CVVB04, EdlVR^{+}03, GC02, KDG^{+}09, VB03]. slave [FR06]. small [CN03, Che01, CG06, FM00, IME02, IO08, JARM02, KvGH01, Leh06, LZA02, Oos09, PO03, PBZ00, PDS01, RRS09, RZWS07, SHH07, TYO^{+}02, Van02a, WS02b, ZP03, ZOJ^{+}06, ZX09]. small- [SHH07]. smallest [SRS07]. SMART [TTBM09]. SMART-based [TTBM09]. smooth [GPN01, KSY^{+}00, PZS04]. smooth-particle [KSY^{+}00]. smooth-permittivity [PZS04]. smoothed [LV08]. smoothing [HPP00, ILB03, WS02a]. snapshot [YNZ^{+}08]. SnCl [RD00]. sodium [FL07, MHS05, YSJ09]. SODOCK [CLH^{+}07]. Soft [yCkHmY08, ASDP^{+}06, TLKT00, TJE03, TGGP^{+}00]. soft-core [TLKT00, TJE03]. Software [Ano04b, BACJCT01, DvL01, Gly06, JVVK09, CHB^{+}05, GBDP05, KBA^{+}04, MMP^{+}07, NSU^{+}02, BLMS08]. solar [KS05c]. solid [CFS^{+}08, CCCJ09, CMA^{+}08, EGSG00, Ish02, KCK^{+}08, SK05]. solid-state [CMA^{+}08]. Solids [vDSSvA04, JB04]. solubilities [SHH07]. solubility [BBG^{+}04, EDAJ04, KEH^{+}02, LLW^{+}09]. solute [BRLS08, BRLS12, FCP^{+}05, LFZS04, MR04, YL09, ZSK07]. solutes [BLL^{+}06, HMSM06]. solution [ABWT09, BHW00, BP07, BISB02, BH03, CPJ00, CCK01, CRSB03, DA01, EK06, ELK^{+}09, FHRR07, FG02, GMA04, HHJ03, HMWC03, HSWN01, HRR05, HBW00, HBW01, HDO^{+}02, KPR04, Kri08, Kri09b, KBN02, LRI^{+}02, LXSF08, LMIF06, MB00, MH08b, PDP02, PTC01, PHRR08, RNG03, RRZA08, SH09, SATO04, SBB02, TDH06, Vas02, VBS09, YH06]. solution-phase [TDH06]. solutions [Blo04, CPJ01, Loe03, PK04, VP09, XZ04, XZ05, ZWP08, LR03b]. solvated [HTSR04, HRR05, KHY00, QSS01, RSP03, BSÇ^{+}01]. Solvation [COL01, HHP04, WB04a, WB04b, WB05, WD04, BCIB05, CRSB03, CCT^{+}03, COL^{+}06, DV02, DHW^{+}00, FOL^{+}04, FBLO08, FZL^{+}06, GS02, GS03, GPN01, GWS^{+}02, HC08, HLMR06, IV04, IvSV06, KIM^{+}09, LF04, LFZS04, LS08b, MGLO03, OVMV04, PZS04, PPYS08, RSE07, RP04, Sch00, SDL07, VM02, VP09, XL02, ZFL^{+}05, ZFL^{+}05]. solvation-effect [SDL07]. solvatochromatic [XZ04]. solve [KvGH01, XOW^{+}00, Zho06]. Solvent [BA03, BA04b, SMAdV00, ZP03, ZGFL01, AG03, BHW00, BBHD04, BMLV04, BRLS08, BRLS12, ENM^{+}04, FEVM01, FC06, GZL02, GL04a, GP06, GB04, HHS^{+}05, HN02, JS07a, JZD^{+}09, KIFK07, Kró03, KKS04, Lab08, LRI^{+}02, LFBSK07, LSO04, LL01, LFR^{+}04, MBC08, MM07, MCM04, MS01, PMB04, RRZA08, RP07d, SBLK01, STSF02, SHSF05, SL06, TJE03, TSMNG01, Tot04, VBGL^{+}00, WB04a, WB04b, WB05, WWL^{+}09, YTY07, YXZ^{+}04, YL09, ZX04, ZCL09, BA04a, FZL^{+}06]. solvent-accessible [BHW00, BMLV04, HHS^{+}05, TSMNG01]. solvents [GS03, IT03, THHN01]. Solving [FS00b, Höf05, BF04, CCL06, CF04, LMJ02, SATO04]. Some [VE09, FMPS08, JARM02, KCL06, McD08, Rao00a, Sch03, WL04, YLL^{+}09, CMA^{+}08]. sometimes [BE06]. Song [JW12]. Sons [Bic09, Lip00]. source [CSV^{+}07, GCB03]. soybean [TGLL07]. sp [NYK^{+}09]. space [Bie04a, BMTSC01, CSJ01, CvG08, CKT^{+}08, CZA03, GT03, HXLS09, JO02, KF08, LCKL05, LJKL08, Nak02, NA06, OFIK09, PHR^{+}05, PBF07, PBF09, PRSMM02, QNF09, Sch04, THHN01, Van02a, YL06]. spaces [JHPRSM^{+}05, PRSMM02, PRT^{+}07, PRT^{+}08, RSS09]. spacing [ZZvRSC08]. spanning [SN00]. Sparkle [FRS05]. Sparkle/AM1 [FRS05]. Sparse [SSB^{+}03, AGSFAL05, LEK07, RS05, RRS07]. sparsity [JSHG07]. spatial [Bie04a, Bie04b, RP07b]. special [KAK^{+}09]. specialized [Höf05]. species [CFC^{+}08, DR09, GHLK^{+}02, HBM06, KZY09, WG02, YIN03, LMGR05]. Specific [FAR02, LR03a, BS08, HLT^{+}05, Pin01, PRJ02, SFRS01, TGLL07, TST^{+}08, UIHN09, WCF04]. specificities [PB06]. specificity [CJW^{+}09, DLRZ09, LLL07]. specified [Fau01]. spectra [Bac09, BACJCT01, CNN07, CG08, Gor01, HKHN08, JARM02, KŠB09, KFD06, LDL^{+}09, MLCD01, NRKH02, OBBS05, OKE^{+}02, ŠBL05, SN06, TDH06, WM01, YXZ^{+}04, ZGXX06, ZWTP^{+}08, dGWH01]. Spectral [II02, CVR08, GdSuM^{+}07, LFZS04, NINAT^{+}07, NAT07, SMKM00, WG02, ZSK07]. spectrometric [KZW^{+}05]. spectroscopic [Ano06a, FCW06, KCL00, ST06]. spectroscopy [ACM^{+}06, RDM^{+}08, VDM06, WMRW^{+}01, ZPL07]. spectrum [EL09, LMCD09, MWL^{+}08, MGLDS00, PRSMM03]. Speeding [KVF^{+}07]. speedup [BYQS03]. sphere [HdMdS05, HdS06, HD06, LFZS04, SFR07]. spherical [BCIB05, ZFL^{+}05]. spheriphane [CS01]. SPICKER [ZS04]. Spin [Duk01, HYR06, KTM02, LXSF08, Van02b, ACM^{+}06, BB08, BACJCT01, CR08, CR09a, DXW08, DPT03, DF04, JSHG07, KRM^{+}02, KRLD09, KK08c, LB08, Mck07a, Mck07b, VCM01]. spin-crossover [KRLD09]. spin-label [VCM01]. Spin-orbit [KTM02, LXSF08, CR08, CR09a, DXW08, KRM^{+}02, LB08]. Spinor [PV03]. spins [JD09]. Spiro [HELM09]. spiroquinazolinones [DD08]. spline [ALKH04]. splines [GL04b]. split [EA08]. split-valence [EA08]. splitting [PSDM00]. splittings [HLLN06, SFRS01]. Spontaneous [Sza08]. square [CSB08, LLZL09, Nil09]. square-planar [CSB08]. squares [CSD05, Gol09]. Sr [WD08, SCP08, XB08]. Sr-doped [SCP08]. Src [OO08]. SrFeO [Hua09b]. SrZrO [SM06]. SSB [KVS^{+}06]. st2nmr [PRJ02]. stabilities [ACM^{+}06, CTFC08, GYCZ04, STC^{+}08, WDXS06]. Stability [JD09, Owe05, PHFC04, WSC09, CJS^{+}03, CF06, DB07, HXD08, JS07b, JBGK08, OCP02, PGC05, QB05, XFF06, ZXYF09, ZOJ^{+}06, ZM06]. Stabilization [EBDPM00, HYA02]. stabilized [HSF08]. stabilizing [GZ07]. Stable [HDO^{+}02, GDPCPU07, KYFW07, KZY09, KAS^{+}07, Kol04, LMO09, PP08a, PZS04, STC^{+}08]. stable/nonstable [GDPCPU07]. stacked [RRCA08, SBI08]. stacking [CM09, DDBP09, HWTL03, KKY01, WRP^{+}06]. standard [ASDP^{+}06, CG06, FBDG06, KOFF09, LFSB03a, LFSB03b, SSS^{+}09, SL04]. standing [KDG^{+}09]. staphylococcal [JS07a]. Starting [VZVG06, BWI^{+}02]. state [Ang09, BBI^{+}09, CWY09, CFS^{+}08, CHA^{+}07, Chu07, CAG07, CMA^{+}08, HM01, HNWF07, HNWF12, Hir08, HP05, IME02, JHZ09, KT02, Kri09a, LMK01, LZ05a, LDL^{+}09, NTH09, PO03, PSS^{+}04, POJ01, Qua07, SPGS08, Sen06, SRE08, TH02, TST^{+}08, TY03, TKN^{+}08, WCFH02, WHF08, ZH08, ZOJ^{+}06]. state-correlation [SRE08]. state-specific [TST^{+}08]. states [Ang09, ABBC01b, Bof01, Bou01, CWY09, CNN07, DHM^{+}03, DF04, EL07, FCW06, FDSA00, HFS^{+}07, HYR06, HZ09, IR03, KUB07, LS08a, LWX07, LB08, MW09, MLCD01, NBTN04a, NBTN04b, OSA06, OV03, PRSMV08, Qua01, SBI08, SMKM00, VW00, WLZ^{+}07, XZ05, ZL05, ZL07, ZL09b, ZM03, dSVA^{+}09, ABBC01a]. static [FROD08, LDG02, Mar03, XWL^{+}09]. stationary [SK09]. Statistical [HFSD03, PYEA03, DW08, EC06, Kob03, ML00, RK05, SBJ08]. statistical-thermodynamic [RK05]. step [BYQS03, BCP03, DLW06, KM00, KH06, MK02, Mck07b, OV03, Oos09, ZWZ09]. step/particle [BYQS03]. stepwise [LLKC06, LFS^{+}07, NSB08]. Stereodynamics [CMLS05]. Stereoelectronic [DD08, PBF09]. stereoisomers [PCMG09]. stereoselective [AGI^{+}00, AGI^{+}07]. Stereospecific [PF06]. Steric [PBF09, BDW00, XYN^{+}06]. steroids [AGMPRG^{+}08]. sterols [CSU05]. STFs [DBS08]. stiff [ECA06]. stilbene [CJK^{+}02]. STO [CGB03, RLRE01, RLER05]. Stochastic [Agr03, Fau01, KEM08, Sau04, AM09, CSJ01, DHF^{+}05, GdSuM^{+}07, MKT04, ZZ08]. Stoner [SK08]. stool [FKŠ^{+}09]. storage [FR06]. stored [AT02, MBWP03]. stored-integral [MBWP03]. Strain [ST01]. strained [ST01]. strand [GAS04, JS07b]. stranded [AZM03]. Strategies [DBS08, LJS05, YL09, EKB02a, EKB02b, KEB04, Vis02]. strategy [BME05, CZA03, LLL^{+}08, MCF07, RI08, SMGE08, Wan09, WS02a]. streamlining [VGB08]. strength [DMJV05, FO08, SEKS09]. strengths [RM07]. stretch [Kle03, SDCG02]. stretching [CPDZH08]. strictly [FMSA06, TT01, TT05]. Strike [Ano06c, WZW^{+}06]. string [Qua04, Qua07]. strong [LC07, PGG06]. strongest [VHRR07b]. strongly [ONHN00]. strontium [RD06]. Structural [CZFH07, EM03a, Kri09b, LWLS07, LFR07, MS03, BCP03, CKR08, CLF^{+}09, ECM^{+}03, GZ07, HYA02, HHP04, KZY09, Kar06, KPZK06, LL07, LJS05, NAT07, OFIK09, PK04, QLHL09, SLC^{+}09, SVV^{+}08, SRB06, XSHC06, XLC08, ZLJS03, ZWP08, CA07b]. structurally [AGMPRG^{+}08]. Structure [BMTSC01, CDL06, HHWG08, HRR05, HS07b, ITS05, KCL06, KPR04, MN02, PGC05, PLC08, PHRR08, RG08, SG07a, AJ03, AGSFA^{+}05, BED02, BS01, BAH^{+}02, BAÅ07, CLC03, CZB07, CDGS09, CDS09, CMàGL^{+}04, CJS^{+}03, CN05, CLA^{+}00, CPUGD09, CSV^{+}07, DP03, DHW^{+}08, DHW^{+}09, Fau01, FL07, FLOD07, FCP^{+}04a, FL08, FLK^{+}07, GBL^{+}05, GTC06, GGGLL05, GdAcV^{+}07, HHJ03, HEP^{+}02, HN02, HP05, ILKR09, JCA^{+}02, JLHF03, KP05, KFNH08, KHY00, KOFF09, KBK^{+}01, KBL08, KIFK07, KWK^{+}00, KGD06, LEK07, LJKL08, Lee09, LPP06, LZA02, LZ05a, LCSZ09, LW06, LOL^{+}08, LYZ^{+}08, LDL^{+}09, LR03b, Loe03, LLL07, MOP^{+}07, MM00, MLL08a, MM02, NK01, NYH02, NCO^{+}05, OS08, PDP02, PSCD^{+}09, PJB^{+}07, PMM05, PRJ02, PF06, QB05, RPNJ07, RI07, RS05, RRS07, SSB^{+}03, SJJ^{+}04, SB01, SCP08, TD08, TT01]. structure [TGD05, Van02a, VHRR07b, WZY04, WMRW^{+}01, WD08, WS07, XZZ04, YXL^{+}09, ZZY07, ZZS^{+}07, ZZY08, ZLD09, ZX08, vDSSvA04, vEMK01, vE01, RRCA08]. Structure-based [BMTSC01, HS07b, KBK^{+}01]. Structure-breaking [HRR05, VHRR07b]. structure-properties [CDGS09, CDS09]. structure-property [JLHF03, PSCD^{+}09]. structured [DC02]. Structures [AB00, CTFC08, HXD08, KS05a, PCMG09, RSSKB03, Ama02a, BK08, BSP06a, BS05, BRV^{+}07, CUSS03, CWWS07, CRSB03, DW08, DF04, DSB^{+}02, DB02, DJB02, EBAN07, FOL^{+}04, GYCZ04, GJL^{+}08, GCCVB00, HM01, HZ09, Hua09a, Hua09b, IGNH03, Kar01, KAS^{+}07, KKMMS04, LD05a, LWK08, LV08, LLXS02, LhWX07, LS05a, MKGA06, Mas01b, MHT01, PYCD03, PYS05, RI08, RSER09, RM00, RRS09, SO07, SHBD05, SWR06, TZX01b, TZX01a, TT08, TD06, VVBV02, WS05a, WB04a, WB04b, WB05, WDXS06, WS02b, WS07, Wu06, WWS07, WM01, XFF06, XLZ08, YIN03, YNW05, ZCL09, ZXY03, ZGXX06, ACM^{+}06, FZL07, STC^{+}08, UM03, ZXL^{+}04, ZX09]. studied [AGO^{+}02, CFC^{+}08, DPT03, HFS^{+}07, KBL08, LML^{+}00, RGP^{+}07, RJLR06, Sen06, SRE08]. Studies [JW12, ZWS^{+}02, AB00, ADM^{+}06, Ano06a, BY06, BPC01, BI06, BSJ01, BMTFR08, CMBC08, DBS07, EBAN07, FJ08, GYCZ04, GC04, GJK00, Han01, HSWW00, mJlZyL^{+}08, KCL06, KWK^{+}01, KWK^{+}02, Kle02, KZW^{+}05, LYK^{+}04, LWLS07, LDY^{+}08, LJZ^{+}07, LF02, LWY^{+}09, LSY02, MCR08, MK02, MOP^{+}07, MW00, RPNJ07, RZWS07, SGPS09, SS00, SWBM08, SJJ^{+}04, SLL^{+}04a, SLHW09, SFR07, ST06, TJM^{+}03, TMBM02, TCSM03, VS02, VMA03, VL00, VS08, WLL01, WLLS04, WLL07a, WyLG^{+}09, WLL^{+}03, WXX03, WCL05, WZXY07, WXK08, XLL^{+}02, XKKL03, YIN03, YFR05, YLWL09, YLL^{+}09, YKK09, ZDS^{+}05, ZWL^{+}05, ZZZ^{+}06, ZAT07, ZXY03, ZL05]. studio [Gan09]. Studt [Ano06a]. Study [LSAS01, ŠBL05, YZ06, ZCS04, ABYM08, ASDP^{+}06, Ama02a, Ama02b, ATH^{+}03, AVB00, AZM03, BAL^{+}01, BTP09, BSB05, BISB02, BL06, BLO^{+}02, BRLS08, BRLS12, Bor03, BBSS06, BGC^{+}09, BBI^{+}09, BZL05, CMLS05, CC07, CFS^{+}08, CUS00, CU01, CUSS03, CJS^{+}03, CLFA07, CZFH07, CCCJ09, CJW^{+}09, Che01, CN05, CNN07, CSB08, CFD03, CFD04, CPUGD09, ČJPZS08, CMD^{+}04, CG05, CSB^{+}03, DHM^{+}03, DGD^{+}05, DLR^{+}08, DWS^{+}09, Der09, DMC05, EA08, EBL^{+}08, FÁ01a, FL07, FC06, FD03, FO08, FKU^{+}05, FKM^{+}06, FKM^{+}07, FKŠ^{+}09, GXK09, GHLK^{+}02, GD09, Gog08, GM01, GLRL02, GGGLL05, GD06, GKTS04, GBBH09, HWFN01, HLLS05, HSF08, HELM09, HSMT04, HM08, HPL03, HK07, HJCP01, HHP04, HLMR06, IB04, IV04, IvSV06, IIK09, mJlZsLyL07, JHZ09, JW00, JFG04, Kan07, KWHH07, KFD06, KSN01, KIFK07]. study [KS01b, KKMMS04, KS05c, LD05a, LMK01, LKJ^{+}04, LPK07, LWK08, LMCD09, Lee09, LZA02, LL01, LZJ03, LW04a, LX07, LWX07, LLL^{+}08, LS02, LLS03, LMMW04, LS08b, LLKC06, LLW02, LKT04, LDT^{+}02a, LDT^{+}02b, LB08, LYZ^{+}08, LDL^{+}09, LMRVFH^{+}09, LLL07, MWL^{+}08, MGMM07a, MML^{+}06, Mas01a, Mas01b, MDA08, McD03, McD08, MH09, MSBS01, MGG06, MH08a, MM02, ML00, MDI04, NTH00, NBTN04a, NBTN04b, NAT07, OO04, OON01, OY01, OY03, OSA06, OS08, OO08, Pac06, PP08a, Pan07, PGNG03, PGRRNG03, PC00, PFR04b, PG04, PHKG07, Pin03, PMPGP05, PMC^{+}08, PRSMM03, PA05, PAS07, PMM05, PMM06, PWFS01, PHRR08, PB05, wQZsLyZ02, QTdG^{+}08, RG02, RAGLL09a, RAGLL09b, RB01, RRS06, RP04, RD00, RGG08, RUPH06, RR05, SF07, SOOF05, SURG06, SBJ08, SG07a, SLL^{+}04b, SMKM00, SBG09b, SN06, SPT^{+}03, SCG04]. study [SE08, SSW^{+}07, SK05, SLRC01, SVV^{+}08, SMV^{+}09, SSBE06, TBG00, gThDjL^{+}01, TD08, TT02, UNHYT06, UNM^{+}01, UM03, VHRR07b, WZY04, WC09, WMRW^{+}01, WD04, WLX^{+}05, WLLS05, WDS06, WDXS06, WDZS07, WLL^{+}07b, WCW08, WWT08, WZZ^{+}09, WM04, WOC^{+}03, WD08, Won09, WDX^{+}02, WRBV03, WC08, WJX^{+}08, XYN^{+}06, XB08, XDS06a, XFF06, XWXC08, XKG^{+}05, XPW09, YT03, YT04, YTY07, YXZ^{+}04, YDWS06, YXC^{+}07, YLZ08, YLW^{+}08, YQQH09, YHD^{+}06, ZP03, ZPL07, ZZL04, ZLLS04b, ZLLS04a, ZLLS05, ZLLS06a, ZLLS06b, ZZW^{+}07, ZZL^{+}08, ZZL^{+}09, ZH08, ZWY^{+}09, ZOJ^{+}06, ZNLL07, ZX08, ZTP^{+}08, ZL07, dSVA^{+}09, dRLMS00, NBJ04]. studying [AGI^{+}00, AGI^{+}07, dVB01]. styrene [Ama02a, Ama02b, XPW09]. subspace [FS00b, Har04]. substance [CCP04]. substances [ATH^{+}03]. substantially [RK05]. Substituent [JWB05, Lee09, PWFS01, BPC01, HMMS09]. substituents [PSF^{+}08]. substituted [AVB00, BE06, BE07, HM08, HWGB01, LST08, MMMY07, MRS09, OSA06, PB05, RP09, TTB01b, WyLG^{+}09, WW04, YDWS06, WFHP01]. Substitution [ZZS^{+}07, BSB05, JT06, LFBSK07, SOOF05, SSB07, ZWS^{+}02]. substitutional [FSS00]. substitutions [CM09]. substrate [BMTFR08, LCC09]. subtype [FTLV01]. subunit [OON01]. subunits [MML^{+}06, PHFC04]. successful [CLA^{+}00]. successive [IR03]. successively [YK08]. suffice [LLW^{+}09]. Sugiura [TKN^{+}08]. suitability [FMSA06]. suitable [GL04a]. sulfate [ZZW09]. sulfation [CLS^{+}09]. sulfide [DLR^{+}08, GGGLL05]. sulfonylurea [XYN^{+}06]. sulfoxide [CFD04, EBDPM00]. sulfur [BT00, CGB^{+}09, SO09, SW06, TTB01a, WS05b, FNP^{+}06]. sulfuryl [CQ04]. sum [Bou01, SG01, UKN04]. sum-over-states [Bou01]. summation [Ami00, San01, SG01]. sums [KF02b]. super [RP07a]. superatoms [LWW^{+}06]. supercell [EL09]. supercells [MOP^{+}07]. supercomputers [VSK^{+}04]. superconducting [MDI04]. supercritical [HTN03]. superexchange [WL00]. superimposition [ZA07]. Superlinear [FR06]. supermolecule [BA03, BA04a, vDSSvA04]. superoxide [PMM06, RJLR06]. superposition [COS01, CSD05, GRCD01, GCD04, IO08, Mas04, VKP^{+}08, VZVG06]. supershort [XWL^{+}09]. Support [CLXC02, LJZ^{+}07, YMT04, CLS^{+}09, HL08, LCC09, QLHL09, Mui05]. supported [GTC06, KEM08, SD09]. suppressants [LDC^{+}07]. supramolecular [AM07, CMàGL^{+}04, Won09]. surface [ABBC01a, ABBC01b, BMLV04, BL05, Bof01, BHH^{+}09, CF06, DLD^{+}02, DR09, EBL^{+}08, GZL02, GCD^{+}08, GB04, HHS^{+}05, HG08, JBGK08, KLH^{+}04, KCK^{+}08, Lab08, LMK01, LFBSK07, LJ04, MG06, NK06, Pan07, PHJ^{+}08, PFJ^{+}03, PTC01, Qua01, RSN^{+}02, RP07d, SURG06, Sau04, SMGE08, SL06, SDL07, Sza08, TSMNG01, TRS02, VP02, YHD^{+}06, YJF06, ZCS04, ZSC05, ZCL09, ZGFL01, ZBS03, ZXYF09, ZGXX06, EKB02b]. surface-adsorbed [DR09]. surface-doping [JBGK08]. surface-generalized [YJF06]. surfaces [ATH^{+}03, BHW00, BWZ08, DLD^{+}02, EKB02a, EKB02b, ENM^{+}04, FKJ^{+}01, GSB09, HSF08, HHP04, IC08, IDMC09, JB04, KEB04, MCF05, MS00, PBZ00, Rap06, RKA^{+}09, SPDS01, Sch03, Sha02, SBG09b, ST04, YJF06, ZBS03]. surrogate [Mck07a]. surrounding [KGL07, Yos02]. survey [HS07a]. SuSi [CA04]. SVD [CSD05, WG02]. SVM^{T M} [YMT04]. swarm [CZB07, LJZ^{+}07, SJJ^{+}04, CLH^{+}07]. switch [SF07]. switching [GG09]. symmetric [AT02]. symmetrically [Lai07]. Symmetry [PDC^{+}08, PCA^{+}08, BB08, CAGR08, FCP^{+}04b, LWX07, SZW^{+}05, WLPF05, Ell07, PV03, PTC01]. symmetry-adapted [FCP^{+}04b]. Symmetry-driven [PV03]. Symmetry-generation [Ell07]. Synergistic [GS08]. syngas [YQQH09]. synthase [BBSS06]. synthesis [HLC09, PHR^{+}05, WLL01]. synthetic [NHH05, WG02]. system [BL00, HELM09, HRBKB03, IS03, KYL03, LHJ^{+}06, LCGA03, LDTS07, MM03, PGH^{+}04, PRSMM03, Rud05a, YOB^{+}08, ZAT07]. Systematic [AST06, CS03, KWHH07, Kob03, LSAS01, MV06, PK04, PG04, RS05, WM04, WZXY07, ZXY08, Dya02, PWFS01, PV07, SYY^{+}03, WK01, EA06, LMH^{+}09]. systems [AS00, BHW00, BP01, BME05, BGJ01a, BWI^{+}02, CN03, CG06, CvG08, CCK01, CMGDAC^{+}07, DXW08, DRMD03, Don08, DK01, EGSG00, Ell07, FZL07, Fau01, GLMV09, HT03, JCA^{+}02, JTR05, JG03, KSS08, KKC05, KAK^{+}09, KBL08, Kle02, Kle03, Kri09a, LMJ02, LC09, LLL03, LDG02, MMLC05, MKGA06, MTB09, MM07, MS01, Oos09, RLDI09, RSN^{+}02, Rud05b, Rud05c, SRS07, SS00, SYY^{+}03, SWV^{+}05, ST01, TH02, TT08, WWL^{+}09, WNH03, YCXY03, YZ04, vdVGDM00].

T [BBI^{+}09, Lu09, PFJ^{+}03, ZKZ^{+}07, DLD^{+}02, Ike04, KVS^{+}06]. T-cell [KVS^{+}06]. table [Kau07, Nil09, SRB06]. tableaux [SN00]. tables [ARL01]. tabu [MGJAARC00, SE07, SE08]. tailoring [BG03, KKG^{+}09]. taking [SN06]. TaN [ZHMW09]. tandem [UNHYT06]. tantalum [Tie09]. target [FM00]. tautomeric [LS08b]. tautomerism [YXZ^{+}04]. tautomerization [BA03, BA04a, BA04b]. tautomers [HHWG08, PG04]. TCNE [GYMN07, TD08]. tCONCOORD [SD09]. tCONCOORD-GUI [SD09]. TD [CHA^{+}07, SBI08]. TD-DFT [CHA^{+}07, SBI08]. TDDFT [SL04]. TDHF [QCK01, QCK02]. Te [HKHN08, WWS07, HWFN01]. technique [COS01, GKH05, KLM^{+}09, SATO04, TS05]. Techniques [Woo01, You11, AM06b, DC02, FSM09, KH05, PAT^{+}09, PDS01, VE09, WSM^{+}09, vGGB00]. Teller [Kri08, VDM06]. temperature [FGR07, JS07a, KT02, KGD06, MN02, TD08, XK08]. temperatures [KK01b, TK08, WHH^{+}06]. tempered [BBP09, CVVB04]. tempering [SPT07]. templated [ST04]. tensor [BZP09, BAÅ07, RI07]. tensors [CDL06, KRM^{+}02, ZLD09]. term [JČHS07, SP05]. terminal [KK01a]. Terms [Duk01, BMLV04, HP01, LAR^{+}03, RP07a, YZ06]. ternary [Don08, MM07]. territory [Sha07]. tert [Bac09]. tertiary [CMLS05, PRJ02, PF06, SO07]. tessellation [LJ04, PTC01]. tessellationless [Pom04]. test [BCF^{+}09, BUMCMRL00, BLN01, BE06, CF04, FMPS08, KTM02, SBI08, SM03]. Testing [CMàGL^{+}04, BG03, PZS04, WWC^{+}04, WWC^{+}05]. tests [KSB^{+}02, NGTB03]. tetraammonium [CW02]. tetraazanaphthalenes [CdML06]. tetrachloride [DMN05]. tetracoordinate [MMRVH07, SRS07, Wan09]. tetracyanoethylene [LH02]. tetracyanoethylene-contained [LH02]. tetracycline [AS06, AS09]. tetrads [MHS05]. tetrahedral [LSY02, OBT09]. tetrahydroimidazo [SPGS08]. tetrahydroimidazo- [SPGS08]. tetrahydroxouranylate [IvSV06]. tetramer [RRCA08]. Tetraoxide [JW12, SLHW09]. tetrapeptide [DSR^{+}07]. tetrapeptides [GKTS04]. Tetrazine [JW12, SLHW09, XZ04]. Tetrazino [JW12, SLHW09]. Tetrazino-Tetrazine-Tetraoxide [JW12, SLHW09]. tetrazole [dSVA^{+}09]. TGSA [GRCD01, GCD04]. TGSA-Flex [GCD04]. Th [NSO^{+}07]. their [Bac04, Bac05, Bac07, BWZ08, BHH^{+}09, DVRP^{+}03, FL08, GCD^{+}08, JHMB^{+}09, JHMB^{+}11, Lee09, Owe05, PCMG09, SWM04, SRK^{+}00, WWT08, YLWL09, ZXL^{+}04, ZWY^{+}09]. them [YNW05]. theorem [Kar01, RLER07]. theoretic [SWZS04]. Theoretical [Ano06a, AZM03, BY06, BI06, Bor03, BS03, CNN07, CFD03, CFD04, CG08, COMR^{+}04, FJ08, FL07, FDSA00, GYCZ04, GLH^{+}08, HLLS05, HS00, IGL07, IIK09, mJlZyL^{+}08, JW12, KYFW07, KZY09, KSN01, KS05c, LS08a, LH02, LWX07, LDY^{+}08, LLW02, LDT^{+}02a, LDT^{+}02b, LYZ^{+}08, LMRVFH^{+}09, MMLC05, MCK05, MBM^{+}00, NBTN04b, OCP02, OKE^{+}02, PFR04b, wQZsLyZ02, RTG00, RZWS07, SLL^{+}04b, SFR07, ST06, TKS^{+}01, TJM^{+}03, gThDjL^{+}01, VS08, WLL01, WLLS05, WDS06, WDXS06, WDZS07, WLL^{+}07b, WCL05, WJX^{+}08, XFF06, XKKL03, YTY07, YIN03, YFR05, YLW^{+}08, YQQH09, YLWL09, YHD^{+}06, ZLLS04b, ZLLS04a, ZLLS05, ZLLS06a, ZLLS06b, ZZW^{+}07, ZZL^{+}08, ZZL^{+}09, ZZW09, ZXY03, ZL07, BGC^{+}09, CN05, DLR^{+}08, GXK09, Ham07, HRBKB03, HLMR06, Kan07, KKMMS04, LC07, LD05a, Lee09, LL01, LLKC06, LB08, MM02, MDI04, NTH00, NSB08, Nye07, OON01, PGRRNG03, PC00, PAS07]. theoretical [RAGLL09a, RAGLL09b, RRS06, RP04, RJLR06, SLHW09, SMV^{+}09, UNHYT06, WLZ^{+}07, WSM^{+}08, Wou00, XWXC08, YXZ^{+}04, ZL05, dRLMS00, Li01, NBTN04a, RD00, UNM^{+}01, ZPL07, ZLD09]. theoretically [WS02b]. theories [JHZ09]. Theory [BBC^{+}05, SH08, WM12, ALTB06, ASDP^{+}06, ASY01, BC06, CWY09, CFK08, CR08, Chu07, CKW09, CPML08a, Cul08, CGSdST06, DPM09, DSR^{+}07, Ell07, EKB02a, FCW06, FZL07, FG02, FII^{+}07, FLGW00, FS04, FLK^{+}07, FZL^{+}06, GM01, Gri03, Gri04, Haf08, HSMT04, HS07a, Hol05, ION07, IKN08, IN08, JČHS07, KSS08, KWK^{+}01, KWK^{+}02, KK08c, KZW^{+}05, Kut07, LMB08, LF04, LFZS04, LMGR05, LF02, LLZL09, LDL^{+}09, Lu09, MGMM07a, Mat03, MW09, MA09, MH08b, ML00, NYH02, NUH02, NTH09, OFIK09, OKE^{+}02, PFJ^{+}03, PSF^{+}08, PU09, PA05, QTdG^{+}08, RB01, RDM^{+}08, RZWS07, SH07, SH02, SZT08, SSMW09, SSB07, SW06, TST^{+}08, Tru07, TKN^{+}08, TKH03, WRP^{+}06, WB07, WZY04, WMRW^{+}01, WW03, WCHW09, WL00, WCFH02, WNH03, WCL05, XYN^{+}06, XB08, XL02, XPW09, YCXY03, YH09, YYW07, YLL^{+}09, ZZL04, ZH08, Zho06]. theory [ZFL^{+}05, dOMSL01, vGGB00, vLBBR12, Blo04, BE09, CGMPT^{+}08, GGB07a, PFB05, PMC^{+}08, SG07b, ZSC05, Bic09]. theory-based [XL02]. Theozyme [UTH^{+}03]. there [KT02]. thermal [LZZC09, LLKC06, Lu09, SFRS01, WXK08]. Thermally [ZALMG03]. Thermochemical [BT00]. thermochemistry [LLXS02, ZXL^{+}04]. Thermodynamic [NA06, WR07, WFR08, ZWP08, Blo04, BZL05, KS05b, KK01b, LC06, LLZL09, RK05, SY09]. thermodynamical [KZY09]. Thermodynamics [UNM^{+}01, HFSD03, JMD^{+}02, MH09, NSU^{+}02, PCMG09]. thermoelectric [XK08]. thermostats [MZL08]. these [LL00]. thiamin [LS08b]. thiazoline [XKG^{+}05]. thio [MMMY07, TTB01b]. thio-substituted [MMMY07, TTB01b]. thioacetalization [RUPH06]. thioamide [LKJ^{+}04]. thioamide-containing [LKJ^{+}04]. thioether [SFR07]. thiolate [DMN05, SGD06]. thiomandelate [APG05]. thiopeptides [TTB01a, TTB01b]. thiophene [KTM02, PSF^{+}08, RRCA08]. thiophosphoryl [ZJM^{+}07]. thioredoxin [CFR06, CFS^{+}09]. thiouracil [LMGO^{+}09]. third [BGJ01b, Gri03, JGH00, KGN07, LK04, RRP^{+}01, YTH01]. third-order [Gri03]. third-row [BGJ01b, JGH00, LK04, RRP^{+}01, YTH01]. Thomas [Kri09a]. thorium [AB00]. threading [BS08]. three [BY06, BP01, Bie04a, BB08, BGC^{+}09, CV09, DHW^{+}08, FII^{+}07, FLOD07, GDV03, HK08c, KBLP09, Lai07, MP03a, MVLG06, SHBD05, Wan09]. three-body [FII^{+}07]. three-bond [KBLP09]. three-center [Lai07]. three-coordinate [BGC^{+}09]. three-dimensional [BP01, MP03a, MVLG06, SHBD05, Wan09]. three-point [Bie04a]. threshold [Mei02]. thymidine [LBG08]. thymine [KKMMS04, MHS05, NBTN04a, NBTN04b, Pin01]. thymine- [MHS05]. TIBO [AJNG01, SGPS09]. TiCl [UNM^{+}01]. tight [ECM^{+}03, HNWF07, HNWF12, WM04, XL02]. tight-binding [HNWF07, HNWF12, WM04, XL02]. tilt [LHI09]. TIM [AGK03, LD05a]. Time [Bac09, CP08, Gog08, Kol04, LDL^{+}09, Whe08, ZH08, BYQS03, CFK08, CPC^{+}00, DF06, FCW06, HNWF07, HNWF12, HS01, ION07, KMH02, KM00, KH06, MW09, NTH09, PK05, PSF^{+}08, TYN05, TST^{+}08, TKN^{+}08, YH07, ZWZ09, ZM03, vGGB00]. Time-averaged [CP08]. Time-dependent [Bac09, Gog08, LDL^{+}09, Whe08, ZH08, CFK08, FCW06, HNWF07, HNWF12, HS01, ION07, MW09, NTH09, PSF^{+}08, TST^{+}08, TKN^{+}08, YH07, ZM03, vGGB00]. time-frequency [DF06]. Time-reversible [Kol04]. time-step [KM00]. times [DDVD09]. timescale [MST^{+}08]. TiN [JD09]. TINKER^{T M} [Sto05]. TiO [FHRR07]. TIP4P [HPL03, THHN01]. tires [LMGR06]. titanium [UNM^{+}01]. Tl [VHRR07b, VHRR07a]. TlN [ZX08]. TlX [ZL09b]. tm [BM00, GM01]. TMPyP [AZM03]. toluene [GCCVB00]. tomography [RGG08]. tool [ALB09, ÁCD^{+}03, CCL06, CA07b, Gra07, HHG^{+}09, LAEL01, LOL^{+}08, NSU^{+}02, OML^{+}00, Pra01]. toolkit [Hin00]. tools [MRS09, Nye07]. topo [GRCD01, GCD04]. topo-geometrical [GRCD01, GCD04]. topological [CAGR08, DRAS04, DRAS05, FSS00, GDPCPU07, HM08, JLHF03, MP03a, RP07c, RP07b, SFC04, YWHZ03, ZSE08, ZNLL07]. Topology [RSER09, FCP^{+}05, GdAcV^{+}07, Kle03, KF03, KBL08]. Topology-based [RSER09]. Tork [CG03]. torsion [Ano06c, CIB05, FWH^{+}07, FKZ09, MGLDS00, OMNH08, PHR^{+}05, TNS00, WZW^{+}06]. Torsional [PSF^{+}08, DHF^{+}05, FPG^{+}06]. torsions [SP05]. total [RP07a]. toxicity [CMCB08]. Tp [HT05]. TpMXO [HT05]. trace [KLS02]. tracing [Kli01]. tracking [HNR08]. training [AG03, LJZ^{+}07, LSY02, SJJ^{+}04]. trajectories [Ham07, MST^{+}08]. trajectory [Qua07]. trans [BZL05, CSB08, HKHN08, KMM07]. trans- [BZL05]. transamination [LDY^{+}08]. transcriptase [AJNG01]. transfer [AGK03, BA03, BLN01, BL00, CXZ^{+}09, CHRL09, FDSA00, FZL^{+}06, GWM08, GVATG03, GGLR00, HFHL06, IN01, JJH01, LLM08, LMGO^{+}09, Li01, LL01, LH02, LB05, LLS03, MT03, MAF^{+}07, Mck07b, OON01, PGG06, PMPGP05, QZZZ03, Rao00a, SL09, TBG00, WL00, WC08, YS00, ZY01, ZH08, BA04a]. Transferability [CSB^{+}03, TT05, FDM00, KS01a, OSHS03, RSP03, TFZRG01]. Transferable [WBSR03, HXLS09]. transferase [SFR07]. transferred [GFS05]. transfers [XKG^{+}05]. Transform [BWP07, HLM05, ON07, QLHL09, TYN05]. transform-based [HLM05]. transformation [PVdJB00]. transformational [CN03]. transformations [WSM^{+}09]. transformed [Nak02, NA06, vDSSvA04]. Transition [FKRE08, LMGR06, TH02, ABYM08, ABBC01a, ABBC01b, Ano06a, Bac05, BP03, BS06, Bof01, BRV^{+}07, BM00, BGJ01b, CWWS07, Chu07, DLW06, Dib05, EL07, GHH07, GM01, Hol05, JHZ09, KRM^{+}02, LW07, LD05a, LH02, LGB^{+}09, LWZ09, LK03, LK04, NR04, PYCD03, Qua01, Qua07, RRFC^{+}03, SK09, SZT08, ST06, TKS^{+}01, WB07, WL09b, YTH01, ZALMG03]. transition-metal [Ano06a, ST06]. transitions [CZ05, FC06, JW06, OYK^{+}09, SMKM00]. Translation [RLER05]. transmembrane [GAS04, YMT04]. transport [Ara04, CM09, FCP^{+}04a, FCP^{+}05, KK08c]. treating [MA09]. treatment [BCF^{+}09, BZL05, CLA^{+}00, CBH^{+}03, HC08, HHBH00, IB04, JB04, KS05b, KCL00, LS08c, MFB04, MR02, MGLL03, RI07, RP02, XL02]. treatments [CEP07, DWNB01]. tree [GY08]. treecode [DK01]. Trends [SRB06]. triangulation [BHH^{+}09]. triazines [ZX04]. triazolinones [WZY04]. tribenzo [GLRL02]. trichloroacetaldehyde [CU01]. tricoordinated [LTF^{+}07]. trigonal [JHMB^{+}09, JHMB^{+}11]. trimer [LZJ03, RRCA08]. trimers [ABYM08, VS02]. trimethylamine [CPDZH08]. trimethylsilyl [LLKC06]. triosephosphate [AGK03]. tripeptide [VKP^{+}08]. triphosphate [GS04]. triple [PP08a, RPNJ07]. triple-decker [RPNJ07]. triplet [CZ05, CG08, FDSA00, LS08a, OSA06]. tripodal [HA04]. trisaccharide [GBB07]. tRNA [GGT08]. truncated [KK08a]. Truncation [RRS09, MN02]. Trypan [SRK^{+}00]. Trypsin [JZD^{+}09, CWV^{+}05, MBC08]. Trypsin-ligand [JZD^{+}09]. tryptophan [HLC09, Li01, LL01, MM05]. Tsi [XFF06]. TTTO [JW12, SLHW09]. tubular [FL07]. Tuczek [Ano06a]. tumor [WCF04]. tumor-specific [WCF04]. Tuning [JHMB^{+}09, JHMB^{+}11]. tunneling [Chu07, MKT04, RWBH09, SFRS01]. TURBOMOLE [LLL03]. turns [HL08]. Two [PFB05, Yas08, AMR04, AHK02, BRS07, BE09, CVR08, CCK01, DHF^{+}05, FBDG06, FR06, GGP09, GYMN07, GGA00, HK08c, JJK^{+}00, KAS^{+}07, KT02, Lai07, LDC^{+}07, LWW^{+}06, Sen06, Sto05, Van02a, Van02b, YFR05, dSR08]. two- [Lai07]. two-body [FBDG06]. two-center [BRS07, GGA00]. two-component [Van02b]. two-dimensional [CVR08]. Two-electron [PFB05, Yas08, FR06, GYMN07, GGA00, Lai07]. two-electron/four-centers [GYMN07]. two-state [KT02, Sen06]. type [Bie04a, CXZ^{+}09, CJDK09, CVVB04, EdlVR^{+}03, GC02, Gri06, HLC09, IKN08, KDG^{+}09, Leh06, MY08a, OON01, SDL^{+}09, TW03, TLOG00, VB03, WD08, YXC^{+}07]. types [BY06, GGP09, KS05b, MLL^{+}08b]. Typical [SMV^{+}09, MLL^{+}08b]. tyrosine [AG00, CLS^{+}09, LRI^{+}02, Li01, LL01, OO08].

U [Han01, CCCJ09, GHLK^{+}02, RLDI09]. ubiquinone [IN01]. ubiquitin [KIFK07]. UCSF [PGH^{+}04]. UF [Han01]. UK [Lip00]. ulcerogenic [CMBC08]. ultra [ZHMW09]. ultra-incompressible [ZHMW09]. Ultrafast [BR07]. ultrasoft [PSS^{+}04, PSMB05]. umbrella [RNG03]. unbiased [Pul05, SYC08]. uncertainty [SY09]. unconstrained [DMN03]. uncorrected [PSC^{+}01]. understand [DSB^{+}02]. Understanding [CAGR08, CDPL09, BRS07, ZZW09, CFER04, HP04]. Unicorns [FK07b]. Unified [GDPP08, CMGDAC^{+}07]. uniform [HdMdS05, HdS06, HD06, Rap06]. Unimolecular [ML00, FS00b, KZY09, ZZL04]. unique [KT02]. unit [VM07, Yas08]. unitary [KBT03]. units [CXZ^{+}09, FEV^{+}09, HP05, NK01, NL08, PC00, PFC03]. Universal [DHW^{+}00, HdMdS05]. Unorthodox [KBB09]. unphysical [OV03]. UNRES [HXLS09, NCO^{+}05]. unrestricted [YH07]. unsaturated [BS03, KFD06, MTB09, Wan09, ZKZ^{+}07]. Unusual [XK08]. UO [IV04, IvSV06, RDM^{+}08]. Update [BKS02]. updated [Chu07]. Updates [Ano04b, BACJCT01, DvL01, Gly06, JVVK09]. upon [OFIK09]. uracil [LMGO^{+}09, MSBS01, MHS05]. uracil-base [MHS05]. uranyl [IV04]. Urea [SK05, AS00, VVS07]. ureases [ESM06]. Use [BWI^{+}02, DW08, Wou00, ALB09, FC06, JNV08, Kli01, MRC03, OCP02, PRKP05, PRS04, RCJ02a, RSN^{+}02, Ruv07, SH07, SVT09, VGGMM05, YTH01, YZ04]. used [DvG00, ESP04, HdMdS05, HdS06, HD06]. user [DPDG05, JKII08]. uses [KBB09]. Using [CSD04, FSM09, HL08, Kne05, LL07, MO01, OSHS03, QLHL09, SWR06, XSHC06, XLC08, XOW^{+}00, Adc04, AJ03, ABWT09, AM06b, AS00, BWP07, BMLV04, BVW04, BME05, BGC^{+}09, Bud07, CLWL09, CN03, CSJ01, CLC09, CKR08, CLA^{+}00, Chu07, CP08, CP09, CCP04, CPML08a, CGSdST06, DLD^{+}02, DWNB01, DR09, DVP^{+}02, DB06, DB02, DMJV05, EKO^{+}01, EKB02a, EM03b, FCW06, FMPS08, FBDG06, FAB^{+}00, FEVM01, FR06, FSFK05, GMA04, GL04b, GdSuM^{+}07, GPSP06, GGLR00, Haf08, HWDB03, Han01, HSMT04, HMSM06, HMMS09, HG08, HHP04, II02, IS07, IS03, IT03, IK00, JBGK08, JIK09, JVVK09, JTR05, JFG04, JSHG07, KRM^{+}02, KKG^{+}09, KM00, KLH^{+}04, KK08c, KOFF09, Kle02, Kle03, KBT03, KKS04, Lab08, LCKL05, LCC09, LZ05a, LLL^{+}08, LZZC09, LS05a, LZ05b, LSY02, LKW04, MWL^{+}08, MT03, MMLC05, MKT04]. using [MV06, MBP09, MOP^{+}07, MTE04, MRS09, MRS^{+}07, NCO^{+}05, NINAT^{+}07, OFB08, OKE^{+}02, PMB04, PS09a, PAT^{+}09, PP08b, PDS01, PZS04, PSS^{+}04, RI07, RI08, RMP01, RG08, RON02, SDL^{+}09, SPGS08, SSB^{+}03, Sch00, SRCD03, SBG^{+}09a, SY09, SPT07, SMV^{+}09, TP01a, Tie09, TCSM03, UBDPJ04, VSW^{+}03, Van02a, WLZ^{+}07, WL00, WEE01, WG02, WOC^{+}03, WCS09, XLT07, YK00, YYW07, ZCS04, ZBS03, ZWP08, ZHH09, vdVGDM00, PRSMM02]. utility [KMH02]. utilization [GS08, DMLI05]. utilizing [NYTH09, Wan09].

v [Lip00, ZZW09, GBJ03, Kri09b, PFC03, TD08]. vacancy [ZMH^{+}09]. vacuum [BISB02]. Valence [LW07, SH08, Tru07, WM12, Bic09, BLT03, Cul08, EA08, HELM09, HS07a, May07, MG00, PRSMV08, SWZS04, SMZW05, SSMW09, SSW^{+}07, VBGL^{+}00, WMRW^{+}01, WJ00, dSVA^{+}09, vLBBR12]. valent [AZM03, CN05, GWL07]. Validation [BAÅ07, VCM01, AGI^{+}00, DGD^{+}05, JJB02, JCL05, JVVK09, MSR04, MTE04, SRB06, APG05, HZ06b, NGTB03]. valinamide [HJCP01]. value [FPG^{+}06, TBSM09]. values [OS06, PMPGP05]. vanadium [PV07, Tie09]. vapor [PHJ^{+}08, UNM^{+}01]. variable [CFS^{+}09, GS08, WHRG08]. variables [SWR06]. variance [Blo04, LRWG03]. variate [LR06]. Variation [AAP00, NAT07, PGG06, Rao00a, Vya01]. Variational [MR02, AB09, Chu07, GY06, HdMdS05, HdS06, HD06, RS07a, RS07b]. Variations [TGGP^{+}00]. variety [ŠBL05]. various [BP07, HMMS09, IT03, Kró03, KS01b, MLL^{+}08b, PP08a, PFJ^{+}03, PMM05, RR05, WHH^{+}06, ZCL09]. varying [CC09]. VASP [Haf08]. VBSCF [vLBBR12]. VCH [Sta00]. VDD [GHBB04, GHBB04]. vector [CLXC02, CLS^{+}09, HL08, LCC09, LJZ^{+}07, QLHL09, YMT04]. vectorizing [SO07]. vectors [BWI^{+}02]. versatile [KF08, TdMSD^{+}08]. version [HDBD04]. versions [Sto05]. versus [ABYM08, ALC08, BSG07, BB08, JS07b, JBGK08, LST08, LTV08, PSMB05, Van02b]. Vertical [PRSMM03, SA07, CG08, LWX07, LFEdL06, SLRC01, TKS^{+}01]. VI [HP05]. via [BA03, BA04a, Bou01, BRS00, BRS01, CZB07, CAG07, DFGB09, Hua09a, JP09, KSB^{+}02, KRLD09, LMO09, SMM^{+}08, SG01, ZPL07]. Viabilty [KK01b]. viable [LMGR05]. Vibalizer [Gra07]. vibrating [Yos02]. vibration [CCL06, LSY02, ZWPR^{+}04]. Vibrational [BP07, CLP^{+}05, LC09, NR04, WB07, BRV^{+}07, DB07, GBDP05, Gra07, Han01, HNR08, JARM02, LMB08, LN01, MR02, NRKH02, NAT07, PZWG^{+}04, Tor02, WM04, WM01]. vibrations [CPDZH08, DR09, KCL00, vE01]. vibronic [TP01b]. view [CSJD04, Jac09, JMD^{+}02, MGCA07]. VIII [EBD^{+}01]. vinyl [YYW07]. vinylphosphine [MGLDS00]. viridis} [IN01, OON01]. virtual [GFS05, KSM05, LZ05b, PRDS08, YOB^{+}08]. virus [AJNG01, DLRZ09, KCL06]. viscosity [ZP03]. Visualization [MMP^{+}07, RP07b, ARL01, KYL03, PGH^{+}04]. Visually [SD09]. VMD [Pra01]. VMFCI [CCL06]. VO [PV07]. voltammetry [KJP^{+}07]. Volume [Sta00, BVW04, Lab08, LFSB03a, LFSB03b, Mue01, QNF09]. volume-preserving [QNF09]. volumes [BHH^{+}09, Rao00b, SBLK01]. Voronoi [GHBB04, MVLG06, SBLK01]. VP1 [KCL06]. VPP700 [KSY^{+}00]. Vpu [KF08]. vs [CXZ^{+}09, LLKC06, MA05, SCG04, Wib04].

W [UM03, WWC^{+}05, MH09]. W2 [dOMSL01]. Waals [AD00, CPUGD09, GdSuM^{+}07, Gri04, KLH^{+}04, LS08c, VS02]. walk [CY09, CY13]. walking [BHG03]. wall [BG07]. walled [XWL^{+}09, ZZvRSC08]. Wang [Ano06c, GHH07, JW12]. warping [JO02]. Water [LMIF06, Mor02, NK01, BLL^{+}06, BUMCMRL00, BRLS08, BRLS12, BSH07, CCK01, CSB^{+}03, DLR^{+}08, DWS^{+}09, DSB^{+}02, ES00, FG03, FBDG06, FKŠ^{+}09, GWM08, GDV03, Gol09, HM02, HRBKB03, HPL03, HFHL06, HN02, HTN03, HLMR06, ITS06, IDMC09, KFNH08, Kle02, Kle03, LPB03, MT03, MN02, MZ05, MG00, OCP02, PHJ^{+}08, PPYS08, RR05, SO07, SSM08, SJW09, SVV^{+}08, THHN01, Tot04, UM03, VHRR07a, VHRR07b, VL00, WL09a, WD04, WDWS06, XKG^{+}05, YT03, YSJ09, Yos02, YGLvG06, ZX04, ZCZ03, ZZY07, ZZY08, ZWP08, BA04b, WJX^{+}08]. water-addition [RR05]. Water-assisted [BA04b, WJX^{+}08]. water-phase [KFNH08]. WATGEN [BSH07]. Wave [GBJ03, Bac04, Bac07, Bou00, GFS05, MLL06, PFB05, PSS^{+}04, PSMB05, RLDI09, TT05, VSK^{+}04, YH06, YLL^{+}09]. wavefunction [IS07]. wavefunctions [BWW^{+}08, KTM02, PJPJdPRMI07]. Wavelet [CF04, ON07, QLHL09]. waves [MOP^{+}07]. way [GZ07, HSWW00, PRSMM02, VBGL^{+}00]. weak [QTdG^{+}08]. weakly [CPML08a, JRJ01]. web [KKG^{+}09, GKRG08, Gra07, JKII08]. Web-based [Gra07, JKII08]. web-interface [KKG^{+}09]. WebMTA [KKG^{+}09]. WebProp [GKRG08]. Wei [Ano06c]. weight [AG00]. weighted [FS00b, HWDB03]. weights [Bac04, Bac05, Bac07, Kar01]. Weinhold [GHBB04]. well [BBP09, WCK00]. well-tempered [BBP09]. where [HYA02]. which [SBH02]. Wiener [YWH04]. Wiley [Bic09, Lip00, Sta00]. Wiley-VCH [Sta00]. Will [LLW^{+}09]. Windock [HS07b]. Windows [HS07b]. Windows-based [HS07b]. wise [Nil09]. within [Der09, FDM00, GS09, KC01b, MY08b, NYTH09, NAT07, SHBD05, SPT^{+}03]. without [AL01, HdMdS05, HdS06, HD06, HZ09, Nil09, PK05, Qua07, RKA^{+}09, TW03]. WNO [HT05]. worker [FCK^{+}08]. worker-based [FCK^{+}08]. Working [Nye07]. World [Woo01, You11, FK07b]. worlds [Sim07]. Wu [Ano06c].

X [BAL^{+}01, BPC01, CRC^{+}08, Dib05, HZ09, Hua09a, KBL08, LS08a, Mck07a, Mck07b, OS08, RB01, STC^{+}08, WLLS04, WZZ^{+}09, WWS07, ZJM^{+}07, ZY01, ZL09b, HYA02, HSWN01, HN02, Lip00, Mar03, Sta00, WKYU01]. X-ray [HSWN01, HN02, WKYU01]. xanthate [CFD03]. XCu [KBL08]. xenobiotics [PCMG09]. xenon [HdS06]. XES [EKO^{+}01]. Xiang [Ano06c]. Xinli [JW12]. XMVB [SMZW05]. XPS [EKO^{+}01]. XX [CRC^{+}08]. XXIII [JPF^{+}00]. XXXI [HJCP01]. XXXIII [PFJ^{+}03]. XY [ZY01]. xylose [GVATG03].

yield [CSD05]. YL [NYK^{+}09]. ylide [LS08b]. ylides [Mit01]. Ylidic [XDS06b]. yloxy [BE07]. Yong [Ano06c]. young [SN00, Woo01]. ytterbium [FRS05].

Z [CRC^{+}08, PWFS01]. zeolites [LTV08, SDCG02, TLOG00]. zeolitic [Tie09]. Zerner [Ano00]. zero [RKA^{+}09]. zero-flux [RKA^{+}09]. zeta [CMJ08]. Zhang [Ano06c]. Zhi [Ano06c]. Zhi-Xiang [Ano06c]. Zinc [CFS^{+}08, BSDM04, ECM^{+}03, JT06, JT08, KZRO03, RGP^{+}07, SFR07, SCF^{+}09, SDM02]. zinc- [SDM02]. Zintl [XB08, ALC08]. zirconocene [LHP01]. Zn [BTP09, GPSP06, BBI^{+}09, ESM06, FO08, GPK05, ROG00, TGGP^{+}00, TFZRG01]. Zn-biomimetic [FO08]. ZnO [HSF08]. ZnX [WWS07]. Zori [AGSFA^{+}05]. ZPE [MGLDS00]. ZPVA [QCK01, QCK02]. zwitterionic [ROG00]. zwitterions [KLB03]. zymogen [PDP02]. ZZ [CRC^{+}08].

References

Abu-Awwad:2000:VPB

[AAP00] Fakhr Abu-Awwad and Peter Politzer. Variation of parameters in Becke-3 hybrid exchange-correlation functional. Journal of Computational Chemistry, 21(3):227–238, February 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Adamo:2000:SPL

[AB00] Carlo Adamo and Vincenzo Barone. Structures and properties of lanthanide and actinide complexes by a new density functional approach: Lanthanum, gadolinium, lutetium, and thorium halides as case studies. Journal of Computational Chemistry, 21(13):1153–1166, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Andrushchenko:2008:CDE

[AB08] Valery Andrushchenko and Petr Bouř. Circular dichroism enhancement in large DNA aggregates simulated by a generalized oscillator model. Journal of Computational Chemistry, 29(16):2693–2703, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anisimov:2009:QQD

[AB09] Victor M. Anisimov and Vladislav L. Bugaenko. QM/QM docking method based on the variational finite localized molecular orbital approximation. Journal of Computational Chemistry, 30(5):784–798, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Almlof:2004:BAP

[ABÅ04] Martin Almlöf, BjØrn O. Brandsdal, and Johan Åqvist. Binding affinity prediction with different force fields: Examination of the linear interaction energy method. Journal of Computational Chemistry, 25(10):1242–1254, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anglada:2001:EQR

[ABBC01a] Josep Maria Anglada, Emili Besalú, Josep Maria Bofill, and Ramon Crehuet. Erratum: On the quadratic reaction path evaluated in a reduced potential energy surface model and the problem to locate transition states. Journal of Computational Chemistry, 22(7):803, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [ABBC01b].

Anglada:2001:QRP

[ABBC01b] Josep Maria Anglada, Emili Besalú, Josep Maria Bofill, and Ramon Crehuet. On the quadratic reaction path evaluated in a reduced potential energy surface model and the problem to locate transition states. Journal of Computational Chemistry, 22(4):387–406, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See comment [Qua01], reply [ABBC01b], and erratum [ABBC01a].

Anikin:2003:LBO

[ABF^{+}03] N. A. Anikin, V. L. Bugaenko, M. V. Frash, L. Gorb, and J. Leszczynski. Localized basis orbitals: Minimization of 2electron integrals array and orthonormality of basis set. Journal of Computational Chemistry, 24(9):1132–1141, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Altman:2009:ASM

[ABWT09] Michael D. Altman, Jaydeep P. Bardhan, Jacob K. White, and Bruce Tidor. Accurate solution of multi-region continuum biomolecule electrostatic problems using the linearized Poisson–Boltzmann equation with curved boundary elements. Journal of Computational Chemistry, 30(1):132–153, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Addicoat:2008:AVD

[ABYM08] Matthew A. Addicoat, Mark A. Buntine, Brian Yates, and Gregory F. Metha. Associative versus dissociative binding of CO to 4d transition metal trimers: a density functional study. Journal of Computational Chemistry, 29(9):1497–1506, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Angyan:2003:OTO

[ÁCD^{+}03] János G. Ángyán, Christophe Chipot, François Dehez, Christof Hättig, Georg Jansen, and Claude Millot. OPEP: a tool for the optimal partitioning of electric properties. Journal of Computational Chemistry, 24(8):997–1008, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Allinger:2003:AEC

[ACLD03] Norman L. Allinger, Kuo-Hsiang Chen, Jenn-Huei Lii, and Kathleen A. Durkin. Alcohols, ethers, carbohydrates, and related compounds. I. The MM4 force field for simple compounds. Journal of Computational Chemistry, 24(12):1447– 1472, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Atanasov:2006:DMC

[ACM^{+}06] Mihail Atanasov, Peter Comba, Bodo Martin, Vera Müller, Gopalan Rajaraman, Heidi Rohwer, and Steffen Wunderlich. DFT models for copper(II) bispidine complexes: Structures, stabilities, isomerism, spin distribution, and spectroscopy. Journal of Computational Chemistry, 27(12):1263– 1277, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Allinger:2000:VWE

[AD00] Norman L. Allinger and Kathleen A. Durkin. Van der Waals effects between hydrogen and first-row atoms in molecular mechanics (MM3/MM4). Journal of Computational Chemistry, 21(14):1229–1242, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Adcock:2004:PBR

[Adc04] Stewart A. Adcock. Peptide backbone reconstruction using dead-end elimination and a knowledge-based forcefield. Journal of Computational Chemistry, 25(1):16–27, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Agostini:2006:GSA

[ADM^{+}06] Flavia P. Agostini, Diogo De O. Soares-Pinto, Marcelo A. Moret, Carla Osthoff, and Pedro G. Pascutti. Generalized simulated annealing applied to protein folding studies. Journal of Computational Chemistry, 27(11):1142–1155, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alonso:2006:PMM

[AE06] José Luis Alonso and Pablo Echenique. A physically meaningful method for the comparison of potential energy functions. Journal of Computational Chemistry, 27(2):238–252, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alkorta:2003:CED

[AEE^{+}03] Ibon Alkorta, Jose Elguero, Enrique Espinosa, Ignasi Mata, and Elies Molins. Comparison of electron density properties in frozen and relaxed electronic distributions. Journal of Computational Chemistry, 24(4):416–421, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alhambra:2000:HBI

[AG00] Cristobal Alhambra and Jiali Gao. Hydrogen-bonding interactions in the active site of a low molecular weight proteintyrosine phosphatase. Journal of Computational Chemistry, 21(13):1192–1203, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ahmad:2003:DTN

[AG03] Shandar Ahmad and M. Michael Gromiha. Design and training of a neural network for predicting the solvent accessibility of proteins. Journal of Computational Chemistry, 24(11): 1313–1320, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alcaro:2000:QFA

[AGI^{+}00] S. Alcaro, F. Gasparrini, O. Incani, S. Mecucci, D. Misiti, M. Pierini, and C. Villani. A “quasi-flexible” automatic docking processing for studying stereoselective recognition mechanisms. Part I. Protocol validation. Journal of Computational Chemistry, 21(7):515–530, May 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alcaro:2007:QFA

[AGI^{+}07] S. Alcaro, F. Gasparrini, O. Incani, L. Caglioti, M. Pierini, and C. Villani. “quasi flexible” automatic docking processing for studying stereoselective recognition mechanisms, part 2: Prediction of ΔΔG of complexation and ^{1}H-NMR NOE correlation. Journal of Computational Chemistry, 28(6):1119–1128, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alagona:2003:IMS

[AGK03] Giuliano Alagona, Caterina Ghio, and Peter A. Kollman. The intramolecular mechanism for the second proton transfer in triosephosphate isomerase (TIM): a QM/FE approach. Journal of Computational Chemistry, 24(1):46–56, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alvarez-Ginarte:2008:APR

[AGMPRG^{+}08] Yoanna María Alvarez-Ginarte, Yovani Marrero-Ponce, José Alberto Ruiz-García, Luis Alberto Montero-Cabrera, Jose Manuel García De La Vega, Pedro Noheda Marin, Rachel Crespo-Otero, Francisco Torrens Zaragoza, and Ramón García-Domenech. Applying pattern recognition methods plus quantum and physico-chemical molecular descriptors to analyze the anabolic activity of structurally diverse steroids. Journal of Computational Chemistry, 29(3):317–333, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Antony:2002:BDC

[AGO^{+}02] Jens Antony, Nohad Gresh, Lars Olsen, Lars Hemmingsen, Christopher J. Schofield, and Rogert Bauer. Binding of D- and L-captopril inhibitors to metallo-β-lactamase studied by polarizable molecular mechanics and quantum mechanics. Journal of Computational Chemistry, 23(13):1281–1296, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Agrafiotis:2003:SPE

[Agr03] Dimitris K. Agrafiotis. Stochastic proximity embedding. Journal of Computational Chemistry, 24(10):1215–1221, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aspuru-Guzik:2005:ZPQ

[AGSFA^{+}05] Alán Aspuru-Guzik, Romelia Salomón-Ferrer, Brian Austin, Raul Perusquía-Flores, Mary A. Griffin, Ricardo A. Oliva, David Skinner, Dominik Domin, and William A. Lester Jr. Zori 1.0: a parallel quantum Monte Carlo electronic structure package. Journal of Computational Chemistry, 26(8):856–862, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aspuru-Guzik:2005:SAE

[AGSFAL05] Alán Aspuru-Guzik, Romelia Salomón-Ferrer, Brian Austin, and William A. Lester Jr. A sparse algorithm for the evaluation of the local energy in quantum Monte Carlo. Journal of Computational Chemistry, 26(7):708–715, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Apgar:2009:CEM

[AHGK09] James R. Apgar, Seungsoo Hahn, Gevorg Grigoryan, and Amy E. Keating. Cluster expansion models for flexiblebackbone protein energetics. Journal of Computational Chemistry, 30(15):2402–2413, November 30, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Auer:2002:BSC

[AHK02] Alexander A. Auer, Trygve Helgaker, and Wim Klopper. Basis-set completeness profiles in two dimensions. Journal of Computational Chemistry, 23(3):420–425, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alfonso:2003:FNE

[AJ03] Dominic R. Alfonso and Kenneth D. Jordan. A flexible nudged elastic band program for optimization of minimum energy pathways using ab initio electronic structure methods. Journal of Computational Chemistry, 24(8):990–996, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Abrahao-Junior:2001:CAH

[AJNG01] Odonírio Abrahão-Júnior, Paulo G. B. D. Nascimento, and Sérgio E. Galembeck. Conformational analysis of the HIV-1 virus reverse transcriptase nonnucleoside inhibitors: TIBO and nevirapine. Journal of Computational Chemistry, 22(15): 1817–1829, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Akama:2007:IDC

[AKN07] Tomoko Akama, Masato Kobayashi, and Hiromi Nakai. Implementation of divide-and-conquer method including Hartree– Fock exchange interaction. Journal of Computational Chemistry, 28(12):2003–2012, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Agrafiotis:2001:MSC

[AL01] Dimitris K. Agrafiotis and Victor S. Lobanov. Multidimensional scaling of combinatorial libraries without explicit enumeration. Journal of Computational Chemistry, 22(14):1712– 1722, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Allen:2009:GMG

[ALB09] William J. Allen, Justin A. Lemkul, and David R. Bevan. GridMAT-MD: a grid-based membrane analysis tool for use with molecular dynamics. Journal of Computational Chemistry, 30(12):1952–1958, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alemany:2008:RCE

[ALC08] Pere Alemany, Miquel Llunell, and Enric Canadell. Roles of cations, electronegativity difference, and anionic interlayer interactions in the metallic versus nonmetallic character of Zintl phases related to arsenic. Journal of Computational Chemistry, 29(13):2144–2153, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Artemyev:2004:FBS

[ALKH04] A. N. Artemyev, E. V. Ludeña, V. V. Karasiev, and A. J. Hernández. A finite B-spline basis set for accurate diatomic molecule calculations. Journal of Computational Chemistry, 25(3):368–374, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alcoba:2006:OLC

[ALTB06] Diego R. Alcoba, Luis Lain, Alicia Torre, and Roberto C. Bochicchio. An orbital localization criterion based on the theory of “fuzzy” atoms. Journal of Computational Chemistry, 27(5):596–608, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Allen:2006:DPE

[AM06a] Benjamin D. Allen and Stephen L. Mayo. Dramatic performance enhancements for the FASTER optimization algorithm. Journal of Computational Chemistry, 27(10):1071–1075, July 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anderson:2006:MSR

[AM06b] Peter C. Anderson and Sandro Mecozzi. Minimum sequence requirements for selective RNA-ligand binding: a molecular mechanics algorithm using molecular dynamics and freeenergy techniques. Journal of Computational Chemistry, 27 (14):1631–1640, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alabugin:2007:RGM

[AM07] Igor V. Alabugin and Mariappan Manoharan. Rehybridization as a general mechanism for maximizing chemical and supramolecular bonding and a driving force for chemical reactions. Journal of Computational Chemistry, 28(1):373–390, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Addicoat:2009:KCS

[AM09] Matthew A. Addicoat and Gregory F. Metha. Kick: Constraining a stochastic search procedure with molecular fragments. Journal of Computational Chemistry, 30(1):57–64, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Amatatsu:2002:ISE

[Ama02a] Yoshiaki Amatatsu. Ab initio study on the electronic structures of styrene in the Franck–Condon region. Journal of Computational Chemistry, 23(9):928–937, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Amatatsu:2002:ISP

[Ama02b] Yoshiaki Amatatsu. Ab initio study on the photochemical behavior of styrene. Journal of Computational Chemistry, 23 (10):950–956, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Amisaki:2000:PEE

[Ami00] Takashi Amisaki. Precise and efficient Ewald summation for periodic fast multipole method. Journal of Computational Chemistry, 21(12):1075–1087, September 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Al-Matar:2004:GEM

[AMR04] Ali Khalaf Al-Matar and David A. Rockstraw. A generating equation for mixing rules and two new mixing rules for interatomic potential energy parameters. Journal of Computational Chemistry, 25(5):660–668, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Angeli:2009:NIE

[Ang09] Celestino Angeli. On the nature of the π → π^{∗} ionic excited states: the V state of ethene as a prototype. Journal of Computational Chemistry, 30(8):1319–1333, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2000:MMC

[Ano00] Anonymous. In memoriam: Michael C. Zerner. Journal of Computational Chemistry, 21(8):591, June 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2001:A

[Ano01a] Anonymous. AIM2000. Journal of Computational Chemistry, 22(5):545–559, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2001:CBC

[Ano01b] Anonymous. CI benchmark calculations on PC. Journal of Computational Chemistry, 22(5):560–563, April 15, 2001. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2001:NSJ

[Ano01c] Anonymous. New section of the Journal of Computational Chemistry. Journal of Computational Chemistry, 22(1):1, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2004:PNSa

[Ano04a] Anonymous. Publisher’s note: Sir John A. Pople, 1925–2004. Journal of Computational Chemistry, 25(9):fmv–viii, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2004:PNSb

[Ano04b] Anonymous. Publisher’s note: “Software News and Updates” section of the Journal of Computational Chemistry. Journal of Computational Chemistry, 25(14):i, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2005:A

[Ano05a] Anonymous. Announcements. Journal of Computational Chemistry, 26(5):v, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anonymous:2005:EDS

[Ano05b] Anonymous. Erratum: David E. Shaw. “A fast, scalable method for the parallel evaluation of distance-limited pairwise particle interactions”, Journal of Computational Chemistry (2005) 26 (13) 1318–1328. Journal of Computational Chemistry, 26(16):1803, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [Sha05].

Anonymous:2006:EFS

[Ano06a] Anonymous. Erratum: Felix Studt, Felix Tuczek, “Theoretical, spectroscopic, and mechanistic studies on transition-metal dinitrogen complexes: Implications to reactivity and relevance to the nitrogenase problem,” Journal of Computational Chemistry (2006) 27(12) 1278–1291. Journal of Computational Chemistry, 27(13):1621, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [ST06].

Anonymous:2006:EMC

[Ano06b] Anonymous. Erratum: Matteo Ceccarelli, Piero Procacci, Massimo Marchi “An ab initio force field for the cofactors of bacterial photosynthesis,” Journal of Computational Chemistry (2003) 24(2) 129–142. Journal of Computational Chemistry, 27(13):1620, October 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [CPM03, Van08, CPML08b].

Anonymous:2006:EZX

[Ano06c] Anonymous. Erratum: Zhi-Xiang Wang, Wei Zhang, Chun Wu, Hongxing Lei, Piotr Cieplak, Yong Duan, “Strike a balance: Optimization of backbone torsion parameters of AM-BER polarizable force field for simulations of proteins and peptides”, Journal of Computational Chemistry (2006) 27(6) 781–790. Journal of Computational Chemistry, 27(8):994, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [WZW^{+}06].

Antony:2005:CTC

[APG05] Jens Antony, Jean-Philip Piquemal, and Nohad Gresh. Complexes of thiomandelate and captopril mercaptocarboxylate inhibitors to metallo-β-lactamase by polarizable molecular mechanics. Validation on model binding sites by quantum chemistry. Journal of Computational Chemistry, 26(11):1131– 1147, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aragon:2004:PBE

[Ara04] Sergio Aragon. A precise boundary element method for macromolecular transport properties. Journal of Computational Chemistry, 25(9):1191–1205, July 15, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Agrafiotis:2001:MSV

[ARL01] Dimitris K. Agrafiotis, Dmitrii N. Rassokhin, and Victor S. Lobanov. Multidimensional scaling and visualization of large molecular similarity tables. Journal of Computational Chemistry, 22(5):488–500, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ayala:2000:ECL

[AS00] Philippe Y. Ayala and Gustavo E. Scuseria. Electron correlation in large molecular systems using the atomic orbital formalism. The case of intermolecular interactions in crystalline urea as an example. Journal of Computational Chemistry, 21 (16):1524–1531, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aleksandrov:2006:TMC

[AS06] Alexey Aleksandrov and Thomas Simonson. The tetracycline: Mg^{2+} complex: a molecular mechanics force field. Journal of Computational Chemistry, 27(13):1517–1533, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aleksandrov:2009:MMM

[AS09] Alexey Aleksandrov and Thomas Simonson. Molecular mechanics models for tetracycline analogs. Journal of Computational Chemistry, 30(2):243–255, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alves-Santos:2006:ASD

[ASDP^{+}06] Marcelo Alves-Santos, Liliana Y. A. Dávila, Helena M. Petrilli, Rodrigo B. Capaz, and Marília J. Caldas. Application of standard DFT theory for nonbonded interactions in soft matter: Prototype study of poly-para-phenylene. Journal of Computational Chemistry, 27(2):217–227, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Autschbach:2002:DRE

[ASS^{+}02] J. Autschbach, S. Siekierski, M. Seth, P. Schwerdtfeger, and W. H. E. Schwarz. Dependence of relativistic effects on electronic configuration in the neutral atoms of d- and f-block elements. Journal of Computational Chemistry, 23(8):804– 813, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Altun:2006:SQM

[AST06] Ahmet Altun, Sason Shaik, and Walter Thiel. Systematic QM/MM investigation of factors that affect the cytochrome P450-catalyzed hydrogen abstraction of camphor. Journal of Computational Chemistry, 27(12):1324–1337, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Alexeev:2007:PDD

[ASWG07] Yuri Alexeev, Michael W. Schmidt, Theresa L. Windus, and Mark S. Gordon. A parallel distributed data CPHF algorithm for analytic Hessians. Journal of Computational Chemistry, 28(10):1685–1694, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Arulmozhiraja:2001:BRI

[ASY01] Sundaram Arulmozhiraja, Tadatake Sato, and Akira Yabe. Benzdiynes revisited: ab initio and density functional theory. Journal of Computational Chemistry, 22(9):923–930, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ahlrichs:2002:ELA

[AT02] Reinhart Ahlrichs and Kakha Tsereteli. Efficient linear algebra routines for symmetric matrices stored in packed form. Journal of Computational Chemistry, 23(2):306–309, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Autenrieth:2004:CFF

[ATBLS04] Felix Autenrieth, Emad Tajkhorshid, Jerome Baudry, and Zaida Luthey-Schulten. Classical force field parameters for the heme prosthetic group of cytochrome c. Journal of Computational Chemistry, 25(13):1613–1622, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aquino:2003:AOS

[ATH^{+}03] Adélia J. A. Aquino, Daniel Tunega, Georg Haberhauer, Martin H. Gerzabek, and Hans Lischka. Adsorption of organic substances on broken clay surfaces: a quantum chemical study. Journal of Computational Chemistry, 24(15):1853– 1863, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Anjima:2007:RMC

[ATM^{+}07] Hidenori Anjima, Shinya Tsukamoto, Hirotoshi Mori, Masaki Mine, Mariusz Klobukowski, and Eisaku Miyoshi. Revised model core potentials of s-block elements. Journal of Computational Chemistry, 28(15):2424–2430, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Amisaki:2003:DHA

[ATMK03] Takashi Amisaki, Shinjiro Toyoda, Hiroh Miyagawa, and Kunihiro Kitamura. Development of hardware accelerator for molecular dynamics simulations: a computation board that calculates nonbonded interactions in cooperation with fast multipole method. Journal of Computational Chemistry, 24 (5):582–592, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Arnaud:2000:QMS

[AVB00] Roger Arnaud, Valentina Vetere, and Vincenzo Barone. Quantum mechanical study of regioselectivity of radical additions to substituted olefins. Journal of Computational Chemistry, 21 (8):675–691, June 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Arenas:2009:PAB

[AVS09] Miguel Arenas, M. Carmen Villaverde, and Fredy Sussman. Prediction and analysis of binding affinities for chemically diverse HIV-1 PR inhibitors by the modified SAFE p approach. Journal of Computational Chemistry, 30(8):1229–1240, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Arnaud:2003:TSI

[AZM03] Philippe Arnaud, Krystyna Zakrzewska, and Bernard Meunier. Theoretical study of the interaction between a highvalent manganese porphyrin oxyl-(hydroxo)-Mn(IV)-TMPyP and double-stranded DNA. Journal of Computational Chemistry, 24(7):797–805, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aschi:2004:CFE

[AZS^{+}04] Massimiliano Aschi, Costantino Zazza, Riccardo Spezia, Cecilia Bossa, Alfredo Di Nola, Maurizio Paci, and Andrea Amadei. Conformational fluctuations and electronic properties in myoglobin. Journal of Computational Chemistry, 25 (7):974–984, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Balata:2003:SEG

[BA03] Bülent Balata and Viktorya Aviyente. Solvent effects on glycine. I. A supermolecule modeling of tautomerization via intramolecular proton transfer. Journal of Computational Chemistry, 24(14):1789–1802, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [BA04a].

Balata:2004:ESE

[BA04a] Bülent Balata and Viktorya Aviyente. Erratum: Solvent effects on glycine. I. A supermolecule modeling of tautomerization via intramolecular proton transfer. Journal of Computational Chemistry, 25(1):151, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [BA03].

Balta:2004:SEG

[BA04b] Bülent Balta and Viktorya Aviyente. Solvent effects on glycine II. Water-assisted tautomerization. Journal of Computational Chemistry, 25(5):690–703, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bandyopadhyay:2008:SOA

[BA08] Deepak Bandyopadhyay and Dimitris K. Agrafiotis. A selforganizing algorithm for molecular alignment and pharmacophore development. Journal of Computational Chemistry, 29(6):965–982, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Buttingsrud:2007:VCP

[BAÅ07] Bård Buttingsrud, Bjørn K. Alsberg, and Per-Olof Åstrand. Validation of critical points in the electron density as descriptors by building quantitative structure–property relationships for the atomic polar tensor. Journal of Computational Chemistry, 28(13):2130–2139, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bachler:2004:SCS

[Bac04] Vinzenz Bachler. A simple computational scheme for obtaining localized bonding schemes and their weights from a CASSCF wave function. Journal of Computational Chemistry, 25(3):343–367, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bachler:2005:BTM

[Bac05] Vinzenz Bachler. The behavior of transition metal nitrido bonds towards protonation rationalized by means of localized bonding schemes and their weights. Journal of Computational Chemistry, 26(6):532–551, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bachler:2007:ONA

[Bac07] Vinzenz Bachler. Orthogonal natural atomic orbitals form an appropriate one-electron basis for expanding CASSCF wave functions into localized bonding schemes and their weights. Journal of Computational Chemistry, 28(12):2013– 2019, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bachler:2009:TDD

[Bac09] Vinzenz Bachler. Time-dependent density functional calculations on the electronic spectra of the neutral nickel complex [Ni(L^{ISQ})_{2}] (L^{ISQ} = 3,5-di-tert-butyl-o-diiminobenzosemiquinonate(1^{−})) and its monoanion and dication. Journal of Computational Chemistry, 30(13):2087–2098, October 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Borras-Almenar:2001:SNU

[BACJCT01] J. J. Borrás-Almenar, J. M. Clemente-Juan, E. Coronado, and B. S. Tsukerblat. Software news and updates: MAGPACK: a package to calculate the energy levels, bulk magnetic properties, and inelastic neutron scattering spectra of high nuclearity spin clusters. Journal of Computational Chemistry, 22(9): 985–991, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bultinck:2002:GPD

[BAH^{+}02] Patrick Bultinck, Stijn Augustynen, Hans W. Hilbers, Ed E. Moret, and Jan P. Tollenaere. Generate: a program for 3-D structure generation and conformational analysis of peptides and peptidomimetics. Journal of Computational Chemistry, 23(7):746–754, May 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Balucani:2001:GPD

[BAL^{+}01] N. Balucani, O. Asvany, Y. T. Lee, R. I. Kaiser, N. Galland, M. T. Rayez, and Y. Hannachi. Gas-phase detection of the HBCC (X^{1}Σ) molecule: a combined crossed beam and computational study of the B(^{2}P) + C_{2}H_{2}(^{1}Σ_{g}^{+}) reaction. Journal of Computational Chemistry, 22(13):1359–1365, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Banerjee:2005:ECA

[BB05] Sanjay Banerjee and John A. Board Jr. Efficient charge assignment and back interpolation in multigrid methods for molecular dynamics. Journal of Computational Chemistry, 26(9): 957–967, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Binning:2008:HSV

[BB08] R. C. Binning Jr. and Daniel E. Bacelo. High-spin versus broken symmetry — effect of DFT spin density representation on the geometries of three diiron (III) model compounds. Journal of Computational Chemistry, 29(5):716–723, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Banks:2005:IMP

[BBC^{+}05] Jay L. Banks, Hege S. Beard, Yixiang Cao, Art E. Cho, Wolfgang Damm, Ramy Farid, Anthony K. Felts, Thomas A. Halgren, Daniel T. Mainz, Jon R. Maple, Robert Murphy, Dean M. Philipp, Matthew P. Repasky, Linda Y. Zhang, Bruce J. Berne, Richard A. Friesner, Emilio Gallicchio, and Ronald M. Levy. Integrated Modeling Program, Applied Chemical Theory (IMPACT). Journal of Computational Chemistry, 26(16):1752–1780, December 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Belmares:2004:HHS

[BBG^{+}04] M. Belmares, M. Blanco, W. A. Goddard III, R. B. Ross, G. Caldwell, S.-H. Chou, J. Pham, P. M. Olofson, and Cristina Thomas. Hildebrand and Hansen solubility parameters from molecular dynamics with applications to electronic nose polymer sensors. Journal of Computational Chemistry, 25(15): 1814–1826, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Basdevant:2004:SIS

[BBHD04] Nathalie Basdevant, Daniel Borgis, and Tap Ha-Duong. A semi-implicit solvent model for the simulation of peptides and proteins. Journal of Computational Chemistry, 25(8):1015– 1029, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bucinsky:2009:REG

[BBI^{+}09] Lukáš Bučinský, Stanislav Biskupič, Michal Ilčin, Vladimír Lukeš, and Viliam Laurinc. On relativistic effects in ground state potential curves of Zn_{2}, Cd_{2}, and Hg_{2} dimers. A CCSD(T) study. Journal of Computational Chemistry, 30(1): 65–74, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brooks:2009:CBS

[BBM^{+}09] B. R. Brooks, C. L. Brooks III, A. D. Mackerell Jr., L. Nilsson, R. J. Petrella, B. Roux, Y. Won, G. Archontis, C. Bartels, S. Boresch, A. Caflisch, L. Caves, Q. Cui, A. R. Dinner, M. Feig, S. Fischer, J. Gao, M. Hodoscek, W. Im, K. Kuczera, T. Lazaridis, J. Ma, V. Ovchinnikov, E. Paci, R. W. Pastor, C. B. Post, J. Z. Pu, M. Schaefer, B. Tidor, R. M. Venable, H. L. Woodcock, X. Wu, W. Yang, D. M. York, and M. Karplus. CHARMM: the biomolecular simulation program. Journal of Computational Chemistry, 30(10):1545–1614, July 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bonomi:2009:REB

[BBP09] M. Bonomi, A. Barducci, and M. Parrinello. Reconstructing the equilibrium Boltzmann distribution from well-tempered metadynamics. Journal of Computational Chemistry, 30(11):

1615–1621, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Borowski:2006:EDC

[BBSS06] Tomasz Borowski, Ewa Broclawik, Christopher J. Schofield, and Per E. M. Siegbahn. Epimerization and desaturation by carbapenem synthase (CarC). A hybrid DFT study. Journal of Computational Chemistry, 27(6):740–748, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brown:2006:SDG

[BC06] Russell A. Brown and David A. Case. Second derivatives in generalized Born theory. Journal of Computational Chemistry, 27(14):1662–1675, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barone:2009:RET

[BCF^{+}09] Vincenzo Barone, Maurizio Casarin, Daniel Forrer, Michele Pavone, Mauro Sambi, and Andrea Vittadini. Role and effective treatment of dispersive forces in materials: Polyethylene and graphite crystals as test cases. Journal of Computational Chemistry, 30(6):934–939, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Benzi:2005:BCF

[BCIB05] Caterina Benzi, Maurizio Cossi, Roberto Improta, and Vincenzo Barone. Building cavities in a fluid of spherical or rodlike particles: a contribution to the solvation free energy in isotropic and anisotropic polarizable continuum model. Journal of Computational Chemistry, 26(11):1096–1105, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Berente:2007:CEE

[BCNs07] Imre Berente, Eszter Czinki, and Gábor Náray-szabó. A combined electronegativity equalization and electrostatic potential fit method for the determination of atomic point charges. Journal of Computational Chemistry, 28(12):1936– 1942, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bosch:2003:BPM

[BCP03] David Bosch, Mercedes Campillo, and Leonardo Pardo. Binding of proteins to the minor groove of DNA: What are the structural and energetic determinants for kinking a basepair step? Journal of Computational Chemistry, 24(6):682–691, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Beke:2004:FP

[BCP04] Tamás Beke, Imre G. Csizmadia, and András Perczel. On the flexibility of β-peptides. Journal of Computational Chemistry, 25(2):285–307, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bofill:2000:AED

[BdPRMAI00] Josep Maria Bofill, Ibério de Pinho Ribeiro Moreira, Josep Maria Anglada, and Francesc Illas. Accurate and efficient determination of higher roots in diagonalization of large matrices based in function restricted optimization algorithms. Journal of Computational Chemistry, 21(15):1375– 1386, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bubel:2000:MMB

[BDW00] Robert J. Bubel, Warthen Douglass, and David P. White. Molecular mechanics-based measures of steric effects: Customized code to compute Ligand repulsive energies. Journal of Computational Chemistry, 21(3):239–246, February 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bohm:2006:CEF

[BE06] Stanislav Böhm and Otto Exner. Are calculated enthalpies of formation sometimes more reliable than experimental? A test on alkyl substituted benzoic acids. Journal of Computational Chemistry, 27(5):571–577, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bohm:2007:IER

[BE07] Stanislav Böhm and Otto Exner. Inductive effects in radicals calculated from DFT energies; substituted bicyclo[2.2.2]octan-1-yloxy radicals. Journal of Computational Chemistry, 28(11):

1783–1789, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bohm:2009:ITF

[BE09] Stanislav Böhm and Otto Exner. Interaction of two functional groups through the benzene ring: Theory and experiment. Journal of Computational Chemistry, 30(7):1069–1074, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ball:2002:ENA

[BED02] Keith D. Ball, Burak Erman, and Ken A. Dill. The elastic net algorithm and protein structure prediction. Journal of Computational Chemistry, 23(1):77–83, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Beran:2003:ICE

[Ber03] Kyle A. Beran. Isomers of C_{20}: an energy profile II. Journal of Computational Chemistry, 24(11):1287–1290, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Boschitsch:2004:HBE

[BF04] Alexander H. Boschitsch and Marcia O. Fenley. Hybrid boundary element and finite difference method for solving the nonlinear Poisson–Boltzmann equation. Journal of Computational Chemistry, 25(7):935–955, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Boschitsch:2007:NOB

[BF07] Alexander H. Boschitsch and Marcia O. Fenley. A new outer boundary formulation and energy corrections for the nonlinear Poisson–Boltzmann equation. Journal of Computational Chemistry, 28(5):909–921, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brooks:2007:LE

[BFS07] Charles L. Brooks III, Gernot Frenking, and Shigeyoshi Sakaki. Letter from the Editors. Journal of Computational Chemistry, 28(11):1775, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2000:HIS

[BG00] Thomas A. Baker and Gregory I. Gellene. A hybrid ab initio /free electron computational model for conjugated dye molecules: Simple cyanines and oxonols. Journal of Computational Chemistry, 21(11):943–953, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Babu:2003:IQO

[BG03] K. Babu and Shridhar R. Gadre. Ab initio quality oneelectron properties of large molecules: Development and testing of molecular tailoring approach. Journal of Computational Chemistry, 24(4):484–495, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Braga:2007:MDS

[BG07] S. F. Braga and D. S. Galvão. Molecular dynamics simulation of single wall carbon nanotubes polymerization under compression. Journal of Computational Chemistry, 28(10):1724– 1734, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brookes:2009:IPL

[BGC^{+}09] Nigel J. Brookes, David C. Graham, Gemma Christian, Robert Stranger, and Brian F. Yates. The influence of peripheral ligand bulk on nitrogen activation by three-coordinate molybdenum complexes — a theoretical study using the ONIOM method. Journal of Computational Chemistry, 30 (13):2146–2156, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bredow:2001:DCC

[BGJ01a] Thomas Bredow, Gerald Geudtner, and Karl Jug. Development of the cyclic cluster approach for ionic systems. Journal of Computational Chemistry, 22(1):89–101, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bredow:2001:MPT

[BGJ01b] Thomas Bredow, Gerald Geudtner, and Karl Jug. MSINDO parameterization for third-row transition metals. Journal of Computational Chemistry, 22(8):861–887, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bordner:2003:BES

[BH03] A. J. Bordner and G. A. Huber. Boundary element solution of the linear Poisson–Boltzmann equation and a multipole method for the rapid calculation of forces on macromolecules in solution. Journal of Computational Chemistry, 24(3):353–367, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brown:2003:CWN

[BHG03] Scott Brown and Teresa Head-Gordon. Cool walking: a new Markov chain Monte Carlo sampling method. Journal of Computational Chemistry, 24(1):68–76, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Busa:2009:ETM

[BHH^{+}09] Ján Buša, Shura Hayryan, Chin-Kun Hu, Jaroslav Skřivánek, and Ming-Chya Wu. Enveloping triangulation method for detecting internal cavities in proteins and algorithm for computing their surface areas and volumes. Journal of Computational Chemistry, 30(3):346–357, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bader:2007:ECB

[BHTCG07] R. F. W. Bader, J. Hernández-Trujillo, and F. Cortés-Guzmán. Essay: Chemical bonding: from Lewis to atoms in molecules. Journal of Computational Chemistry, 28(1):4– 14, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2000:AMF

[BHW00] N. Baker, M. Holst, and F. Wang. Adaptive multilevel finite element solution of the Poisson–Boltzmann equation II. Refinement at solvent-accessible surfaces in biomolecular systems. Journal of Computational Chemistry, 21(15):1343–1352, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bharatam:2006:TSE

[BI06] Prasad V. Bharatam and Pansy Iqbal. Theoretical studies on electron delocalization in diaminoguanidine. Journal of Computational Chemistry, 27(3):334–343, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bickelhaupt:2009:BRB

[Bic09] F. Matthias Bickelhaupt. Book review: A chemist’s guide to valence bond theory, by Sason S. Shaik and Philippe C. Hiberty, John Wiley & Sons, Inc., Hoboken, New Jersey, 2008, 311 pp. ISBN-10 0-470-03735-0. Journal of Computational Chemistry, 30(16):2784, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bieniasz:2004:FOA

[Bie04a] Lesław K. Bieniasz. A fourth-order accurate, Numerovtype, three-point finite-difference discretization of electrochemical reaction-diffusion equations on nonuniform (exponentially expanding) spatial grids in one-dimensional space geometry. Journal of Computational Chemistry, 25(12):1515– 1521, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bieniasz:2004:IAS

[Bie04b] Lesław K. Bieniasz. Improving the accuracy of the spatial discretization in finite-difference electrochemical kinetic simulations, by means of the extended Numerov method. Journal of Computational Chemistry, 25(8):1075–1083, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Benzi:2002:QMS

[BISB02] Caterina Benzi, Roberto Improta, Giovanni Scalmani, and Vincenzo Barone. Quantum mechanical study of the conformational behavior of proline and 4R-hydroxyproline dipeptide analogues in vacuum and in aqueous solution. Journal of Computational Chemistry, 23(3):341–350, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Biegler-Konig:2000:CAI

[BK00] Friedrich Biegler-König. Calculation of atomic integration data. Journal of Computational Chemistry, 21(12):1040–1048, September 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baranov:2008:ELH

[BK08] A. I. Baranov and M. Kohout. Electron localizability for hexagonal element structures. Journal of Computational Chemistry, 29(13):2161–2171, October 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Biegler-Konig:2002:UAP

[BKS02] Friedrich Biegler-König and Jens Schönbohm. Update of the AIM2000-Program for atoms in molecules. Journal of Computational Chemistry, 23(15):1489–1494, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bu:2000:ODD

[BL00] Yuxiang Bu and Chengbu Liu. The orientation and distancedependence analysis of the electron transfer reactivity: an electron correlation level investigation of Mn^{2+}(H_{2}O)_{2}/ Mn^{3+}(H_{2}O)_{2} system. Journal of Computational Chemistry, 21(15):1387–1404, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Blomgren:2005:EPE

[BL05] Fredrik Blomgren and Sven Larsson. Exploring the potential energy surface of retinal, a comparison of the performance of different methods. Journal of Computational Chemistry, 26 (7):738–742, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bil:2006:HRH

[BL06] Andrzej Bil and Zdzisław Latajka. The hydroperoxy radical as a hydrogen bond acceptor. HOO HCl complexes — ab initio study. Journal of Computational Chemistry, 27(3):287–295, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barbe:2008:IDF

[BL08] Sophie Barbe and Marc Le Bret. Ab initio determination of the flexibility of 2’-aminoribonucleosides and 2’- aminoarabinonucleosides inserted in duplexes. Journal of Computational Chemistry, 29(9):1353–1363, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bailey:2009:MSE

[BL09] A. G. Bailey and C. P. Lowe. MILCH SHAKE: an efficient method for constraint dynamics applied to alkanes. Journal of Computational Chemistry, 30(15):2485–2493, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Butler:2009:TAR

[BLB09] Keith T. Butler, F. Javier Luque, and Xavier Barril. Toward accurate relative energy predictions of the bioactive conformation of drugs. Journal of Computational Chemistry, 30(4): 601–610, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bruni:2002:CAN

[BLF02] Aline Thaís Bruni, Vitor B. P. Leite, and Márcia M. C. Ferreira. Conformational analysis: a new approach by means of chemometrics. Journal of Computational Chemistry, 23(2): 222–236, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Basilevsky:2006:CHF

[BLL^{+}06] Mikhail V. Basilevsky, Igor V. Leontyev, Sofia V. Luschekina, Olga A. Kondakova, and Vladimir B. Sulimov. Computation of hydration free energies of organic solutes with an implicit water model. Journal of Computational Chemistry, 27(5):552– 570, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Busch:2008:CPD

[BLMS08] Marcel Schmidt Am Busch, Anne Lopes, David Mignon, and Thomas Simonson. Computational protein design: Software implementation, parameter optimization, and performance of a simple model. Journal of Computational Chemistry, 29(7): 1092–1102, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Blomgren:2001:ETB

[BLN01] Fredrik Blomgren, Sven Larsson, and Stephen F. Nelsen. Electron transfer in bis(hydrazines), a critical test for application of the Marcus model. Journal of Computational Chemistry, 22 (6):655–664, April 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bobrowski:2002:ISM

[BLO^{+}02] Maciej Bobrowski, Adam Liwo, Stanisław Ołdziej, Danuta Jeziorek, and Tadeusz Ossowski. Ab initio study of the mechanism of singlet-dioxygen addition to hydroxyaromatic compounds: Negative evidence for the involvement of peroxa and endoperoxide intermediates. Journal of Computational Chemistry, 23(11):1076–1089, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Blondel:2004:EVF

[Blo04] Arnaud Blondel. Ensemble variance in free energy calculations by thermodynamic integration: Theory, optimal “Alchemical” path, and practical solutions. Journal of Computational Chemistry, 25(7):985–993, May 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bochicchio:2003:DAV

[BLT03] R. Bochicchio, L. Lain, and A. Torre. Determination of atomic valence indices from population analyses at correlated level. Journal of Computational Chemistry, 24(15):1902–1909, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bosque:2000:PSP

[BM00] Ramón Bosque and Feliu Maseras. Performance of the semiempirical PM3 (tm) method in the geometry optimization of transition metal complexes. Journal of Computational Chemistry, 21(7):562–571, May 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Beck:2007:HRA

[BM07] John Frederick Beck and Yirong Mo. How resonance assists hydrogen bonding interactions: an energy decomposition analysis. Journal of Computational Chemistry, 28(1):455–466, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bessac:2008:DMR

[BM08] Fabienne Bessac and Feliu Maseras. DFT modeling of reactivity in an ionic liquid: How many ion pairs? Journal of Computational Chemistry, 29(6):892–899, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bories:2007:SEC

[BMB07] Benoît Bories, Daniel Maynau, and Marie-Laure Bonnet. Selected excitation for CAS-SDCI calculations. Journal of Computational Chemistry, 28(3):632–643, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Borini:2005:CFC

[BME05] Stefano Borini, Daniel Maynau, and Stefano Evangelisti. A combined freeze-and-cut strategy for the description of large molecular systems using a localized orbitals approach. Journal of Computational Chemistry, 26(10):1042–1051, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Becue:2004:EPS

[BMLV04] A. Becue, N. Meurice, L. Leherte, and D. P. Vercauteren. Evaluation of the protein solvent-accessible surface using reduced representations in terms of critical points of the electron density. Journal of Computational Chemistry, 25(9):1117–1126, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Beuve-Mery:2001:PAC

[BMRDB01] Daniel Beuve-Mery, Mehdi Rahman, Philippe Ducarme, and Robert Brasseur. Practical aspects of computational chemistry calculations through PC networks: the RAMSES– Beowulf implementation. Journal of Computational Chemistry, 22(2):172–177, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bringmann:2001:MBC

[BMRF01] Gerhard Bringmann, Jörg Mühlbacher, Charlotte Repges, and Jörg Fleischhauer. MD-based CD calculations for the assignment of the absolute axial configuration of the naphthylisoquinoline alkaloid dioncophylline A. Journal of Computational Chemistry, 22(12):1273–1278, September 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bras:2008:MSF

[BMTFR08] Natércia F. Brás, Sara A. Moura-Tamames, Pedro A. Fernandes, and Maria J. Ramos. Mechanistic studies on the formation of glycosidase-substrate and glycosidase-inhibitor covalent intermediates. Journal of Computational Chemistry, 29 (15):2565–2574, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Budin:2001:SBL

[BMTSC01] Nicolas Budin, Nicolas Majeux, Catherine Tenette-Souaille, and Amedeo Caflisch. Structure-based ligand design by a build-up approach and genetic algorithm search in conformational space. Journal of Computational Chemistry, 22(16): 1956–1970, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bofill:2001:RCQ

[Bof01] Josep Maria Bofill. Reply on the comment “On the quadratic reaction path evaluated in a reduced potential energy surface model and the problem to locate transition states” [by J. M. Anglada, E. Besalú, J. M. Bofill, and R. Crehuet, J Comput Chem 2001, 22, 4, 387–406]. Journal of Computational Chemistry, 22(5):541–544, April 15, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [ABBC01b, Qua01].

Borosky:2003:TSC

[Bor03] Gabriela L. Borosky. Theoretical study concerning the reactivity of imine derivatives of polycyclic aromatic hydrocarbons. Journal of Computational Chemistry, 24(5):601–608, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bour:2000:CHF

[Bou00] Petr Bouř. Comparison of Hartree–Fock and Kohn–Sham determinants as wave functions. Journal of Computational Chemistry, 21(1):8–16, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bour:2001:CRO

[Bou01] Petr Bouř. Computations of the Raman optical activity via the sum-over-states expansions. Journal of Computational Chemistry, 22(4):426–435, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2000:EGO

[BP00] Jon Baker and Peter Pulay. Efficient geometry optimization of molecular clusters. Journal of Computational Chemistry, 21 (1):69–76, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Begue:2001:DFF

[BP01] Didier Bégué and Claude Pouchan. Density functional finite cluster method for polarizability of large Be_{N} threedimensional systems. Journal of Computational Chemistry, 22(2):230–240, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2002:EPA

[BP02] Jon Baker and Peter Pulay. An efficient parallel algorithm for the calculation of canonical MP2 energies. Journal of Computational Chemistry, 23(12):1150–1156, September 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2003:AOO

[BP03] Jon Baker and Peter Pulay. Assessment of the OLYP and O3LYP density functionals for first-row transition metals. Journal of Computational Chemistry, 24(10):1184–1191, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Begue:2007:VAC

[BP07] Didier Begue and Claude Pouchan. Vibrational anharmonic calculations in solution: Performance of various DFT approaches. Journal of Computational Chemistry, 28(8):1456– 1462, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barone:2001:NJC

[BPC01] Verónica Barone, Juan E. Peralta, and Rubén H. Contreras. NMR ^{3}J(C_{1},H_{3}) couplings in 1-X-bicyclo[1.1.1]pentanes. FPT–DFT and NBO studies of hyperconjugative interactions and heavy atom substituent effects. Journal of Computational Chemistry, 22(14):1615–1621, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bultinck:2007:ALP

[BPCD07] Patrick Bultinck, Robert Ponec, and Ramon Carbó-Dorca. Aromaticity in linear polyacenes: Generalized population analysis and molecular quantum similarity approach. Journal of Computational Chemistry, 28(1):152–160, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bathe:2003:IMC

[BR03] Mark Bathe and Gregory C. Rutledge. Inverse Monte Carlo procedure for conformation determination of macromolecules. Journal of Computational Chemistry, 24(7):876–890, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Basch:2004:RBS

[BR04] Harold Basch and Mark A. Ratner. Reduced basis set for the gold atom in cluster complexes. Journal of Computational Chemistry, 25(7):899–906, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ballester:2007:USR

[BR07] Pedro J. Ballester and W. Graham Richards. Ultrafast shape recognition to search compound databases for similar molecular shapes. Journal of Computational Chemistry, 28(10):1711– 1723, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barthe:2002:HMP

[BRDC02] Philippe Barthe, Christian Roumestand, Hélène Déméné, and Laurent Chiche. Helix motion in protein C12A-p8^{MT CP} 1: Comparison of molecular dynamics simulations and multifield NMR relaxation data. Journal of Computational Chemistry, 23(16):1577–1586, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bondesson:2008:BSD

[BRLS08] Laban Bondesson, Elias Rudberg, Yi Luo, and Paweł Sałek. Basis set dependence of solute–solvent interaction energy of benzene in water: a HF/DFT study. Journal of Computational Chemistry, 29(11):1725–1732, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [BRLS12].

Bondesson:2012:EBS

[BRLS12] Laban Bondesson, Elias Rudberg, Yi Luo, and Paweł Sałek. Erratum: “Basis set dependence of solute–solvent interaction energy of benzene in water: a HF/DFT study”. Journal of Computational Chemistry, 33(3):354, January 30, 2012. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [BRLS08].

Brooks:2005:E

[Bro05] Charles L. Brooks III. Editorial. Journal of Computational Chemistry, 26(16):1667, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bushenkov:2000:AIE

[BRS00] Vladimir A. Bushenkov, J. P. Prates Ramalho, and Georgi V. Smirnov. Adsorption integral equation via complex approximation with constraints: the Langmuir kernel. Journal of Computational Chemistry, 21(3):191–200, February 2000. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bushenkov:2001:AIE

[BRS01] Vladimir A. Bushenkov, J. P. Prates Ramalho, and Georgi V. Smirnov. Adsorption integral equation via complex approximation with constraints: kernel of general form. Journal of Computational Chemistry, 22(10):1058–1066, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bitter:2007:TPU

[BRS07] T. Bitter, K. Ruedenberg, and W. H. E. Schwarz. Toward a physical understanding of electron-sharing two-center bonds. I. General aspects. Journal of Computational Chemistry, 28 (1):411–422, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bolshakov:2007:PNB

[BRV^{+}07] Vladimir I. Bolshakov, Vladimir V. Rossikhin, Eugene O. Voronkov, Sergiy I. Okovytyy, and Jerzy Leszczynski. The performance of the new 6-31G^{##} basis set: Molecular structures and vibrational frequencies of transition metal carbonyls. Journal of Computational Chemistry, 28(4):778–782, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Betancourt:2001:FNH

[BS01] Marcos R. Betancourt and Jeffrey Skolnick. Finding the needle in a haystack: educing native folds from ambiguous ab initio protein structure predictions. Journal of Computational Chemistry, 22(3):339–353, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brana:2003:TAM

[BS03] Pedro Braña and José A. Sordo. Theoretical approach to the mechanism of reactions between halogen atoms and unsaturated hydrocarbons: the Cl + propene reaction. Journal of Computational Chemistry, 24(16):2044–2062, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bindewald:2005:SFD

[BS05] Eckart Bindewald and Jeffrey Skolnick. A scoring function for docking ligands to low-resolution protein structures. Journal of Computational Chemistry, 26(4):374–383, March 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Blomberg:2006:QCA

[BS06] Margareta R. A. Blomberg and Per E. M. Siegbahn. Quantum chemistry applied to the mechanisms of transition metal containing enzymes — Cytochrome c oxidase, a particularly challenging case. Journal of Computational Chemistry, 27(12): 1373–1384, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brylinski:2008:QDL

[BS08] Michal Brylinski and Jeffrey Skolnick. Q-Dock: Lowresolution flexible ligand docking with pocket-specific threading restraints. Journal of Computational Chemistry, 29(10): 1574–1588, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bento:2005:IDB

[BSB05] A. Patrícia Bento, Miquel Solà, and F. Matthias Bickelhaupt. Ab initio and DFT benchmark study for nucleophilic substitution at carbon (S_{N}2@C) and silicon (S_{N}2@Si). Journal of Computational Chemistry, 26(14):1497–1504, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Basma:2001:SEA

[BSÇ^{+}01] M. Basma, S. Sundara, D. Çalgan, T. Vernali, and R. J. Woods. Solvated ensemble averaging in the calculation of partial atomic charges. Journal of Computational Chemistry, 22 (11):1125–1137, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brothers:2004:PCZ

[BSDM04] Edward N. Brothers, Dimas Suarez, David W. Deerfield II, and Kenneth M. Merz Jr. PM3-compatible zinc parameters optimized for metalloenzyme active sites. Journal of Computational Chemistry, 25(14):1677–1692, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bickelhaupt:2007:CVI

[BSG07] F. M. Bickelhaupt, M. Solà, and C. Fonseca Guerra. Covalent versus ionic bonding in alkalimetal fluoride oligomers. Journal of Computational Chemistry, 28(1):238–250, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bui:2007:WAM

[BSH07] Huynh-Hoa Bui, Alexandra J. Schiewe, and Ian S. Haworth. WATGEN: an algorithm for modeling water networks at protein–protein interfaces. Journal of Computational Chemistry, 28(14):2241–2251, November 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Borodin:2001:IQC

[BSJ01] Oleg Borodin, Grant D. Smith, and Richard L. Jaffe. Ab initio quantum chemistry and molecular dynamics simulations studies of LiPF_{6} /poly(ethylene oxide) interactions. Journal of Computational Chemistry, 22(6):641–654, April 30, 2001.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Benkova:2005:RSP

[BSOB05] Zuzana Benkova, Andrzej J. Sadlej, Roma E. Oakes, and Steven E. J. Bell. Reduced-size polarized basis sets for calculations of molecular electric properties. I. The basis set generation. Journal of Computational Chemistry, 26(2):145–153, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Beke:2006:TRD

[BSP06a] Tamás Beke, Csaba Somlai, and András Perczel. Toward a rational design of β-peptide structures. Journal of Computational Chemistry, 27(1):20–38, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bhat:2006:CAC

[BSP06b] Sathesh Bhat, Traian Sulea, and Enrico O. Purisima. Coupled atomic charge selectivity for optimal ligand-charge distributions at protein binding sites. Journal of Computational Chemistry, 27(16):1899–1907, December 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Benassi:2000:CBS

[BT00] Rois Benassi and Fernando Taddei. Complete basis set model chemistry applied to molecules of increasing molecular complexity: Thermochemical properties of organic sulfur derivatives. Journal of Computational Chemistry, 21(15):1405–1418, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bastard:2003:DMF

[BTLP03] Karine Bastard, Aurélien Thureau, Richard Lavery, and Chantal Prévost. Docking macromolecules with flexible segments. Journal of Computational Chemistry, 24(15):1910– 1920, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barman:2009:IMM

[BTP09] Arghya Barman, Woody Taves, and Rajeev Prabhakar. Insights into the mechanism of methionine oxidation catalyzed by metal (Cu^{2+}, Zn^{2+}, and Fe^{3+}) — amyloid beta (Aβ) peptide complexes: a computational study. Journal of Computational Chemistry, 30(9):1405–1413, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Budzelaar:2007:GOU

[Bud07] Peter H. M. Budzelaar. Geometry optimization using generalized, chemically meaningful constraints. Journal of Computational Chemistry, 28(13):2226–2236, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bernal-Uruchurtu:2000:IDH

[BUMCMRL00] M. I. Bernal-Uruchurtu, M. T. C. Martins-Costa, C. Millot, and M. F. Ruiz-López. Improving description of hydrogen bonds at the semiempirical level: water–water interactions as test case. Journal of Computational Chemistry, 21(7):572– 581, May 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bekker:2004:MON

[BVW04] Henk Bekker, Jur P. Van Den Berg, and Tsjerk A. Wassenaar. A method to obtain a near-minimal-volume molecular simulation of a macromolecule, using periodic boundary conditions and rotational constraints. Journal of Computational Chemistry, 25(8):1037–1046, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bartels:2005:AFE

[BWE05] Christian Bartels, Armin Widmer, and Claus Ehrhardt. Absolute free energies of binding of peptide analogs to the HIV-1 protease from molecular dynamics simulations. Journal of Computational Chemistry, 26(12):1294–1305, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Buenker:2002:UEM

[BWI^{+}02] R. J. Buenker, J. L. Whitten, E. I. Izgorodina, H.-P. Liebermann, and D. B. Kokh. Use of exchange maximization to generate starting vectors for self-consistent field calculations on metal cluster/adsorbate systems. Journal of Computational Chemistry, 23(10):943–949, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2009:QCP

[BWM^{+}09] Jon Baker, Krzysztof Wolinski, Massimo Malagoli, Don Kinghorn, Pawel Wolinski, Gábor Magyarfalvi, Svein Saebo, Tomasz Janowski, and Peter Pulay. Quantum chemistry in parallel with PQS. Journal of Computational Chemistry, 30 (2):317–335, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Baker:2007:PDG

[BWP07] Jon Baker, Krzysztof Wolinski, and Peter Pulay. Parallel DFT gradients using the Fourier Transform Coulomb method. Journal of Computational Chemistry, 28(16):2581–2588, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bezugly:2008:ELI

[BWW^{+}08] Viktor Bezugly, Pawel Wielgus, Frank R. Wagner, Miroslav Kohout, and Yuri Grin. Electron localizability indicators ELI and ELIA: the case of highly correlated wavefunctions for the argon atom. Journal of Computational Chemistry, 29(8):1198– 1207, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bates:2008:MMS

[BWZ08] P. W. Bates, G. W. Wei, and Shan Zhao. Minimal molecular surfaces and their applications. Journal of Computational Chemistry, 29(3):380–391, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Bao:2006:TSR

[BY06] Peng Bao and Zhong-Heng Yu. Theoretical studies on the role of π-electron delocalization in determining the conformation of N-benzylideneaniline with three types of LMO basis sets. Journal of Computational Chemistry, 27(7):809–824, May 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barash:2003:ISL

[BYQS03] Danny Barash, Linjing Yang, Xiaoliang Qian, and Tamar Schlick. Inherent speedup limitations in multiple time step/ particle mesh Ewald algorithms. Journal of Computational Chemistry, 24(1):77–88, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Burda:2005:HPA

[BZL05] Jaroslav V. Burda, Michal Zeizinger, and Jerzy Leszczynski. Hydration process as an activation of trans- and cisplatin complexes in anticancer treatment. DFT and ab initio computational study of thermodynamic and kinetic parameters. Journal of Computational Chemistry, 26(9):907–914, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Barone:2009:HMD

[BZP09] Vincenzo Barone, Mirco Zerbetto, and Antonino Polimeno. Hydrodynamic modeling of diffusion tensor properties of flexible molecules. Journal of Computational Chemistry, 30(1): 2–13, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curco:2004:PSM

[CA04] David Curcó and Carlos Alemán. Performance of SuSi: a method for generating atomistic models of amorphous polymers based on a random search of energy minima. Journal of Computational Chemistry, 25(6):790–798, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curco:2007:CGP

[CA07a] David Curcó and Carlos Alemán. Coarse-graining: a procedure to generate equilibrated and relaxed models of amorphous polymers. Journal of Computational Chemistry, 28(12): 1929–1935, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curco:2007:CTM

[CA07b] David Curcó and Carlos Alemán. Computational tool to model the packing of polycyclic chains: Structural analysis of amorphous polythiophene. Journal of Computational Chemistry, 28(10):1743–1749, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chatfield:2003:MDC

[CADW03] David C. Chatfield, Alberto Augsten, Cassian D’Cunha, and Sergio E. Wong. Methyl dynamics in crystalline amino acids: MD and NMR. Journal of Computational Chemistry, 24(9): 1052–1058, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Conte:2007:DEF

[CAG07] R. Conte, G. P. Arrighini, and C. Guidotti. Direct evaluation via forced oscillation method of the electronic state density of sizable clusters. Journal of Computational Chemistry, 28(2): 584–593, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Contreras:2008:UTS

[CAGR08] M. L. Contreras, J. Alvarez, D. Guajardo, and R. Rozas. Understanding topological symmetry: a heuristic approach to its determination. Journal of Computational Chemistry, 29(4): 588–600, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carpenter:2002:CPC

[Car02] John E. Carpenter. Computation of pressure components due to Class II force fields. Journal of Computational Chemistry, 23(6):667–672, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chang:2008:EEC

[CBC^{+}08] Max W. Chang, Richard K. Belew, Kate S. Carroll, Arthur J. Olson, and David S. Goodsell. Empirical entropic contributions in computational docking: Evaluation in APS reductase complexes. Journal of Computational Chemistry, 29(11): 1753–1761, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curutchet:2003:EDM

[CBH^{+}03] Carles Curutchet, Josep María Bofill, Begoña Hernández, Modesto Orozco, and F. Javier Luque. Energy decomposition in molecular complexes: Implications for the treatment of polarization in molecular simulations. Journal of Computational Chemistry, 24(10):1263–1275, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chong:2003:DHA

[CBS^{+}03] Lillian T. Chong, Pradipta Bandyopadhyay, Thomas S. Scanlan, Irwin D. Kuntz, and Peter A. Kollman. Direct hydroxide attack is a plausible mechanism for amidase antibody 43C9. Journal of Computational Chemistry, 24(12):1371– 1377, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Capar:2007:OEE

[CC07] Mine Ilk Capar and Emine Cebe. Odd–even effects in the homologous series of alkyl-cyanobiphenyl liquid crystals: a molecular dynamic study. Journal of Computational Chemistry, 28(13):2140–2146, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cai:2009:NIL

[CC09] Junmeng Cai and Siyu Chen. A new iterative linear integral isoconversional method for the determination of the activation energy varying with the conversion degree. Journal of Computational Chemistry, 30(13):1986–1991, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chavez:2004:ICE

[CCB04] Jennifer I. Chavez, Maira M. Carrillo, and Kyle A. Beran. Isomers of C_{20}: an energy profile III. Journal of Computational Chemistry, 25(3):322–327, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cahill:2003:KPF

[CCC03] Sean Cahill, Michael Cahill, and Kevin Cahill. On the kinematics of protein folding. Journal of Computational Chemistry, 24(11):1364–1370, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:IDR

[CCCJ09] Hsin-Tsung Chen, Jee-Gong Chang, Hui-Lung Chen, and Shin-Pon Ju. Identifying the O_{2} diffusion and reduction mechanisms on CeO_{2} electrolyte in solid oxide fuel cells: a DFT + U study. Journal of Computational Chemistry, 30(15):2433– 2442, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Case:2005:ABS

[CCD^{+}05] David A. Case, Thomas E. Cheatham III, Tom Darden, Holger Gohlke, Ray Luo, Kenneth M. Merz Jr., Alexey Onufriev, Carlos Simmerling, Bing Wang, and Robert J. Woods. The AM-BER biomolecular simulation programs. Journal of Computational Chemistry, 26(16):1668–1688, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cieplak:2001:MMM

[CCK01] Piotr Cieplak, James Caldwell, and Peter Kollman. Molecular mechanical models for organic and biological systems going beyond the atom centered two body additive approximation: aqueous solution free energies of methanol and N-methyl acetamide, nucleic acid base, and amide hydrogen bonding and chloroform/water partition coefficients of the nucleic acid bases. Journal of Computational Chemistry, 22(10):1048– 1057, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cassam-Chenai:2006:VMF

[CCL06] P. Cassam-Chenaï and J. Liévin. The VMFCI method: a flexible tool for solving the molecular vibration problem. Journal of Computational Chemistry, 27(5):627–640, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Corcho:2004:CAC

[CCP04] Francesc J. Corcho, Josep Canto, and Juan J. Perez. Comparative analysis of the conformational profile of substance P using simulated annealing and molecular dynamics. Journal of Computational Chemistry, 25(16):1937–1952, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chiu:2000:CMA

[CCSJ00] See-Wing Chiu, Michael Clark, Shankar Subramaniam, and Eric Jakobsson. Collective motion artifacts arising in longduration molecular dynamics simulations. Journal of Computational Chemistry, 21(2):121–131, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curutchet:2003:ECS

[CCT^{+}03] Carles Curutchet, Christopher J. Cramer, Donald G. Truhlar, Manuel F. Ruiz-López, Daniel Rinaldi, Modesto Orozco, and F. Javier Luque. Electrostatic component of solvation: Comparison of SCRF continuum models. Journal of Computational Chemistry, 24(3):284–297, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carpenter:2002:APA

[CCWH02] John E. Carpenter, Alan Christoffels, Yael Weinbach, and Winston A. Hide. Assessment of the parallelization approach of d2 cluster for high-performance sequence clustering. Journal of Computational Chemistry, 23(7):755–757, May 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (elec-

tronic). URL http://www.sanbi.ac.za/CODES.

Cheng:2002:CPC

[CDD^{+}02] Ailan Cheng, David J. Diller, Steven L. Dixon, William J. Egan, George Lauri, and Kenneth M. Merz Jr. Computation of the physio-chemical properties and data mining of large molecular collections. Journal of Computational Chemistry, 23(1):172–183, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carbo-Dorca:2009:NQSa

[CDGS09] Ramon Carbó-Dorca, Ana Gallegos, and Ángel J. Sánchez. Notes on quantitative structure-properties relationships (QSPR) (1): a discussion on a QSPR dimensionality paradox (QSPR DP) and its quantum resolution. Journal of Computational Chemistry, 30(7):1146–1159, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cuesta:2006:SMN

[CDL06] Inmaculada García Cuesta, Alfredo M. J. Sánchez De Merás, and Paolo Lazzeretti. Structure, magnetizability, and nuclear magnetic shielding tensors of bis-heteropentalenes. IV. Dihydrophospholophosphole isomers. Journal of Computational Chemistry, 27(3):344–351, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cuesta:2006:DT

[CdML06] I. García Cuesta, A. Sánchez de Merás, and P. Lazzeretti. Diatropicity of tetraazanaphthalenes. Journal of Computational Chemistry, 27(16):1980–1989, December 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cuesta:2009:URC

[CDPL09] Inmaculada Garíca Cuesta, Alfredo Sánchez De Merás, Stefano Pelloni, and Paolo Lazzeretti. Understanding the ring current effects on magnetic shielding of hydrogen and carbon nuclei in naphthalene and anthracene. Journal of Computational Chemistry, 30(4):551–564, March 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carbo-Dorca:2009:NQSb

[CDS09] Ramon Carbó-Dorca and Ana Gallegos Saliner. Notes on quantitative structure-properties relationships (QSPR) part 2: the role of the number of atoms as a molecular descriptor. Journal of Computational Chemistry, 30(13):2099–2104, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cordomi:2007:EDT

[CEP07] Arnau Cordomí, Olle Edholm, and Juan J. Perez. Effect of different treatments of long-range interactions and sampling conditions in molecular dynamic simulations of rhodopsin embedded in a dipalmitoyl phosphatidylcholine bilayer. Journal of Computational Chemistry, 28(6):1017–1030, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chuev:2004:WAS

[CF04] Gennady N. Chuev and Maxim V. Fedorov. Wavelet algorithm for solving integral equations of molecular liquids. A test for the reference interaction site model. Journal of Computational Chemistry, 25(11):1369–1377, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chocholousova:2006:BAR

[CF06] Jana Chocholoušová and Michael Feig. Balancing an accurate representation of the molecular surface in generalized Born formalisms with integrator stability in molecular dynamics simulations. Journal of Computational Chemistry, 27(6):719–729, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chuev:2008:HIS

[CFC^{+}08] Gennady N. Chuev, Maxim V. Fedorov, Sandro Chiodo, Nino Russo, and Emilia Sicilia. Hydration of ionic species studied by the reference interaction site model with a repulsive bridge correction. Journal of Computational Chemistry, 29(14):2406– 2415, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Claes:2003:TSI

[CFD03] L. Claes, J.-P. François, and M. S. Deleuze. Theoretical study of the internal elimination reactions of xanthate precursors. Journal of Computational Chemistry, 24(16):2023–2031, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Claes:2004:TSI

[CFD04] Luc Claes, Jean-Pierre François, and Michael S. Deleuze. Theoretical study of the internal conversion of sulfoxide precursors of poly-isothianaphthene and related polymers. Journal of Computational Chemistry, 25(1):40–50, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cerqueira:2004:RAE

[CFER04] Nuno M. F. S. A. Cerqueira, Pedro Alexandrino Fernandes, Leif A. Eriksson, and Maria João Ramos. Ribonucleotide activation by enzyme ribonucleotide reductase: Understanding the role of the enzyme. Journal of Computational Chemistry, 25(16):2031–2037, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chiba:2008:PCM

[CFK08] Mahito Chiba, Dmitri G. Fedorov, and Kazuo Kitaura. Polarizable continuum model with the fragment molecular orbitalbased time-dependent density functional theory. Journal of Computational Chemistry, 29(16):2667–2676, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carvalho:2006:DBA

[CFR06] Alexandra T. P. Carvalho, P. A. Fernandes, and Maria J. Ramos. Determination of the ΔpKa between the active site cysteines of thioredoxin and DsbA. Journal of Computational Chemistry, 27(8):966–975, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chamorro:2003:EPD

[CFS03] E. Chamorro, P. Fuentealba, and A. Savin. Electron probability distribution in AIM and ELF basins. Journal of Computational Chemistry, 24(4):496–504, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Catlow:2008:ZOC

[CFS^{+}08] C. Richard A. Catlow, Samuel A. French, Alexey A. Sokol, Abdullah A. Al-Sunaidi, and Scott M. Woodley. Zinc oxide: a case study in contemporary computational solid state chemistry. Journal of Computational Chemistry, 29(13):2234–2249, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carvalho:2009:RVA

[CFS^{+}09] Alexandra T. P. Carvalho, Pedro A. Fernandes, Marcel Swart, Joost N. P. Van Stralen, F. Matthias Bickelhaupt, and Maria J. Ramos. Role of the variable active site residues in the function of thioredoxin family oxidoreductases. Journal of Computational Chemistry, 30(5):710–724, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chang:2003:TCA

[CG03] Chia-En Chang and Michael K. Gilson. Tork: Conformational analysis method for molecules and complexes. Journal of Computational Chemistry, 24(16):1987–1998, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cummins:2005:CMS

[CG05] Peter L. Cummins and Jill E. Gready. Computational methods for the study of enzymic reaction mechanisms III: a perturbation plus QM/MM approach for calculating relative free energies of protonation. Journal of Computational Chemistry, 26(6):561–568, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chien:2006:SSS

[CG06] Siu-Hung Chien and Peter M. W. Gill. SG-0: a small standard grid for DFT quadrature on large systems. Journal of Computational Chemistry, 27(6):730–739, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Corral:2008:TIA

[CG08] Inés Corral and Leticia González. Theoretical investigation of anthracene-9,10-endoperoxide vertical singlet and triplet excitation spectra. Journal of Computational Chemistry, 29(12): 1982–1991, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chong:2003:SGF

[CGB03] Delano P. Chong, Myrta Grüning, and Evert Jan Baerends. STO and GTO field-induced polarization functions for H to Kr. Journal of Computational Chemistry, 24(13):1582–1591, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cerqueira:2009:ESS

[CGB^{+}09] N. M. F. S. A. Cerqueira, P. J. Gonzalez, C. D. Brondino, M. J. Romão, C. C. Romão, I. Moura, and J. J. G. Moura. The effect of the sixth sulfur ligand in the catalytic mechanism of periplasmic nitrate reductase. Journal of Computational Chemistry, 30(15):2466–2484, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cho:2005:IAC

[CGBF05] Art E. Cho, Victor Guallar, Bruce J. Berne, and Richard Friesner. Importance of accurate charges in molecular docking: Quantum mechanical/molecular mechanical (QM/MM) approach. Journal of Computational Chemistry, 26(9):915–931, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cervellino:2006:EEF

[CGG06] Antonio Cervellino, Cinzia Giannini, and Antonietta Guagliardi. On the efficient evaluation of Fourier patterns for nanoparticles and clusters. Journal of Computational Chemistry, 27(9):

995–1008, July 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Castillo-Garit:2008:BBC

[CGMPT^{+}08] Juan A. Castillo-Garit, Yovani Marrero-Ponce, Francisco Torrens, Ramón García-Domenech, and Vicente Romero-Zaldivar. Bond-based 3D-chiral linear indices: Theory and QSAR applications to central chirality codification. Journal of Computational Chemistry, 29(15):2500–2512, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Custodio:2006:ASC

[CGSdST06] Rogério Custodio, André Severo Pereira Gomes, Fabrício Ronil Sensato, and Júlio Murilo dos Santos Trevas. Analysis of the segmented contraction of basis functions using density matrix theory. Journal of Computational Chemistry, 27(15): 1822–1829, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chidthong:2007:ESP

[CHA^{+}07] Rungtiwa Chidthong, Supa Hannongbua, Adelia J. A. Aquino, Peter Wolschann, and Hans Lischka. Excited state properties, fluorescence energies, and lifetime of a poly(fluorene-pyridine) copolymer, based on TD-DFT investigation. Journal of Computational Chemistry, 28(10):1735–1742, July 30, 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Christen:2005:GSB

[CHB^{+}05] Markus Christen, Philippe H. Hünenberger, Dirk Bakowies, Riccardo Baron, Roland Bürgi, Daan P. Geerke, Tim N. Heinz, Mika A. Kastenholz, Vincent Kräutler, Chris Oostenbrink, Christine Peter, Daniel Trzesniak, and Wilfred F. van Gunsteren. The GROMOS software for biomolecular simulation: GROMOS05. Journal of Computational Chemistry, 26(16): 1719–1751, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chesnut:2001:LFS

[Che01] D. B. Chesnut. Localization function study of excitation processes in a set of small isoelectronic molecules. Journal of Computational Chemistry, 22(14):1702–1711, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chipot:2003:RDC

[Chi03] Christophe Chipot. Rational determination of charge distributions for free energy calculations. Journal of Computational Chemistry, 24(4):409–415, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Crocker:2005:MAP

[CHMI05] Michael S. Crocker, Scott S. Hampton, Thierry Matthey, and Jesús A. Izaguirre. MDSimAid: Automatic parameter optimization in fast electrostatic algorithms. Journal of Computational Chemistry, 26(10):1021–1031, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cramariuc:2009:IDP

[CHRL09] Oana Cramariuc, Terttu I. Hukka, Tapio T. Rantala, and Helge Lemmetyinen. Ab initio description of photoabsorption and electron transfer in a doubly-linked porphyrin-fullerene dyad. Journal of Computational Chemistry, 30(8):1194–1201, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chuang:2007:CRC

[Chu07] Yao-Yuan Chuang. Calculating rate constants with updated Hessians using variational transition state theory with multidimensional tunneling. Journal of Computational Chemistry, 28(11):1809–1816, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2005:ATA

[CIB05] Jianhan Chen, Wonpil Im, and Charles L. Brooks III. Application of torsion angle molecular dynamics for efficient sampling of protein conformations. Journal of Computational Chemistry, 26(15):1565–1578, November 30, 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:PIM

[CJDK09] Ke Chen, Yingfu Jiang, Li Du, and Lukasz Kurgan. Prediction of integral membrane protein type by collocated hydrophobic amino acid pairs. Journal of Computational Chemistry, 30(1):

163–172, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Calaminici:2002:MLF

[CJK^{+}02] Patrizia Calaminici, Karl Jug, Andreas M. Köster, Cécile Arbez-Gindre, and Constantinos G. Screttas. Mechanism for large first hyperpolarizabilities of phosphonic acid stilbene derivatives. Journal of Computational Chemistry, 23(2):291– 297, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Condic-Jurkic:2008:MAB

[ČJPZS08] Karmen Čondić-Jurkić, V. Tamara Perchyonok, Hendrik Zipse, and David M. Smith. On the modeling of argininebound carboxylates: a case study with Pyruvate Formate-Lyase. Journal of Computational Chemistry, 29(14):2425– 2433, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2003:SSS

[CJS^{+}03] Zhongfang Chen, Haijun Jiao, Gotthard Seifert, Anselm H. C. Horn, Dengke Yu, Tim Clark, Walter Thiel, and Paul von Ragué Schleyer. The structure and stability of Si_{60} and Ge_{60} cages: a computational study. Journal of Computational Chemistry, 24(8):948–953, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cheng:2009:SSC

[CJW^{+}09] Wan-Chun Cheng, Soonmin Jang, Chen-Chang Wu, Ren-Jie Lin, Hsiu-Feng Lu, and Feng-Yin Li. Site specificity of the αC − H bond dissociation energy for a naturally occurring β-hairpin peptide — an ab initio study. Journal of Computational Chemistry, 30(3):407–414, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cramer:2001:WAE

[CKF01] Christopher J. Cramer, Anne-Marie Kelterer, and Alfred D. French. When anomeric effects collide. Journal of Computational Chemistry, 22(11):1194–1204, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cecchini:2004:ADH

[CKMC04] Marco Cecchini, Peter Kolb, Nicolas Majeux, and Amedeo Caflisch. Automated docking of highly flexible ligands by genetic algorithms: a critical assessment. Journal of Computational Chemistry, 25(3):412–422, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2008:PPS

[CKR08] Ke Chen, Lukasz A. Kurgan, and Jishou Ruan. Prediction of protein structural class using novel evolutionary collocationbased sequence representation. Journal of Computational Chemistry, 29(10):1596–1604, July 30, 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Csontos:2008:EEC

[CKT^{+}08] József Csontos, Péter Kálmán, Gyula Tasi, Miklós Kálmán, Richard F. Murphy, and Sándor Lovas. The effect of electron correlation on the conformational space of melatonin. Journal of Computational Chemistry, 29(9):1466–1471, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Crosby:2009:IDN

[CKW09] Lonnie D. Crosby, Shawn M. Kathmann, and Theresa L. Windus. Implementation of dynamical nucleation theory with quantum potentials. Journal of Computational Chemistry, 30 (5):743–749, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Christianson:2000:ITC

[CLA^{+}00] Laurie A. Christianson, Melissa J. Lucero, Daniel H. Appella, Daniel A. Klein, and Samuel H. Gellman. Improved treatment of cyclic β-amino acids and successful prediction of β-peptide secondary structure using a modified force field: AMBER*C. Journal of Computational Chemistry, 21(9):763–773, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cai:2003:PPS

[CLC03] Yu-Dong Cai, Xiao-Jun Liu, and Kuo-Chen Chou. Prediction of protein secondary structure content by artificial neural network. Journal of Computational Chemistry, 24(6):727–731, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chao:2009:MDS

[CLC09] Shih-Wei Chao, Arvin Huang-Te Li, and Sheng D. Chao. Molecular dynamics simulations of fluid methane properties using ab initio intermolecular interaction potentials. Journal of Computational Chemistry, 30(12):1839–1849, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:MCI

[CLF^{+}09] Lei Chen, Lin Lu, Kairui Feng, Wenjin Li, Jie Song, Lulu Zheng, Youlang Yuan, Zhenbin Zeng, Kaiyan Feng, Wencong Lu, and Yudong Cai. Multiple classifier integration for the prediction of protein structural classes. Journal of Computational Chemistry, 30(14):2248–2254, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2007:MMM

[CLFA07] Kuo-Hsiang Chen, Jenn-Huei Lii, Yi Fan, and Norman L. Allinger. Molecular mechanics (MM4) study of amines. Journal of Computational Chemistry, 28(15):2391–2412, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2007:SSO

[CLH^{+}07] Hung-Ming Chen, Bo-Fu Liu, Hui-Ling Huang, Shiow-Fen Hwang, and Shinn-Ying Ho. SODOCK: Swarm optimization for highly flexible protein–ligand docking. Journal of Computational Chemistry, 28(2):612–623, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carbonniere:2005:VCB

[CLP^{+}05] Philippe Carbonniere, Tecla Lucca, Claude Pouchan, Nadia Rega, and Vincenzo Barone. Vibrational computations beyond the harmonic approximation: Performances of the B3LYP density functional for semirigid molecules. Journal of Computational Chemistry, 26(4):384–388, March 2005. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cacelli:2009:FFM

[CLP09] Ivo Cacelli, Carlo Federico Lami, and Giacomo Prampolini. Force-field modeling through quantum mechanical calculations: Molecular dynamics simulations of a nematogenic molecule in its condensed phases. Journal of Computational Chemistry, 30(3):366–378, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chang:2009:ISV

[CLS^{+}09] Wen-Chi Chang, Tzong-Yi Lee, Dray-Ming Shien, Justin Bo-Kai Hsu, Jorng-Tzong Horng, Po-Chiang Hsu, Ting-Yuan Wang, Hsien-Da Huang, and Rong-Long Pan. Incorporating support vector machine for identifying protein tyrosine sulfation sites. Journal of Computational Chemistry, 30(15): 2526–2537, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cao:2009:RDP

[CLWL09] Zanxia Cao, Zhixiong Lin, Jun Wang, and Haiyan Liu. Refining the description of peptide backbone conformations improves protein simulations using the GROMOS 53A6 force field. Journal of Computational Chemistry, 30(4):645–660, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cai:2002:SVM

[CLXC02] Yu-Dong Cai, Xiao-Jun Liu, Xue-Biao Xu, and Kuo-Chen Chou. Support vector machines for predicting HIV protease cleavage sites in protein. Journal of Computational Chemistry, 23(2):267–274, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:ACP

[CLZ^{+}09] Weiyang Chen, Bo Liao, Wen Zhu, Hao Liu, and Qingguang Zeng. An ant colony pairwise alignment based on the dot plots. Journal of Computational Chemistry, 30(1):93–97, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:MSA

[CLZX09] Weiyang Chen, Bo Liao, Wen Zhu, and Xuyu Xiang. Multiple sequence alignment algorithm based on a dispersion graph and ant colony algorithm. Journal of Computational Chemistry, 30(13):2031–2038, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:CTS

[CM09] Shuang Chen and Jing Ma. Charge transport in stacking metal and metal-free phthalocyanine iodides. Effects of packing, dopants, external electric field, central metals, core modification, and substitutions. Journal of Computational Chemistry, 30(13):1959–1972, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cuny:2008:DCQ

[CMA^{+}08] Jerome Cuny, Sabri Messaoudi, Veronique Alonzo, Eric Furet, Jean-François Halet, Eric Le Fur, Sharon E. Ashbrook, Chris J. Pickard, Regis Gautier, and Laurent Le Polles. DFT calculations of quadrupolar solid-state NMR properties: Some examples in solid-state inorganic chemistry. Journal of Computational Chemistry, 29(13):2279–2287, October 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Casadesus:2004:TES

[CMàGL^{+}04] Ricard Casadesús, Miquel Moreno, àngels González-Lafont, José M. Lluch, and Matthew P. Repasky. Testing electronic structure methods for describing intermolecular H⋯H interactions in supramolecular chemistry. Journal of Computational Chemistry, 25(1):99–105, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cruz-Monteagudo:2008:DBM

[CMBC08] Maykel Cruz-Monteagudo, Fernanda Borges, and M. Natália D. S. Cordeiro. Desirability-based multiobjective optimization for global QSAR studies: Application to the design of novel NSAIDs with improved analgesic, antiinflammatory, and ulcerogenic profiles. Journal of Computational Chemistry, 29 (14):2445–2459, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cruz-Monteagudo:2008:CCA

[CMCB08] Maykel Cruz-Monteagudo, M. Natália D. S. Cordeiro, and Fernanda Borges. Computational chemistry approach for the early detection of drug-induced idiosyncratic liver toxicity.

Journal of Computational Chemistry, 29(4):533–549, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Corzana:2004:HSL

[CMD^{+}04] Francisco Corzana, Mohammed S. Motawia, Catherine Hervé Du Penhoat, Serge Perez, Sarah M. Tschampel, Robert J. Woods, and Søren B. Engelsen. A hydration study of (1→4) and (1→6) linked α-glucans by comparative 10 ns molecular dynamics simulations and 500-MHz NMR. Journal of Computational Chemistry, 25(4):573–586, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cruz-Monteagudo:2007:CCD

[CMGDAC^{+}07] Maykel Cruz-Monteagudo, Humberto González-Díaz, Guillermín Agüero-Chapín, Lourdes Santana, Fernanda Borges, Elena Rosa Domínguez, Gianni Podda, and Eugenio Uriarte. Computational chemistry development of a unified free energy Markov model for the distribution of 1300 chemicals to 38 different environmental or biological systems. Journal of Computational Chemistry, 28(11):1909–1923, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Camiletti:2008:GBS

[CMJ08] G. G. Camiletti, S. F. Machado, and F. E. Jorge. Gaussian basis set of double zeta quality for atoms K through Kr: Application in DFT calculations of molecular properties. Journal of Computational Chemistry, 29(14):2434–2444, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Campomanes:2005:SBR

[CMLS05] Pablo Campomanes, M. Isabel Menéndez, Ramón López, and Tomás L. Sordo. Stereodynamics of bond rotation in tertiary 1-naphthoic acid amides: a computational study. Journal of Computational Chemistry, 26(4):365–373, March 2005. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carlsson:2003:IPE

[CN03] Peter Carlsson and Lennart Nilsson. Improved precision and efficiency of free energy calculations for small systems using λ-scaled atomic masses and separating conformational and transformational sampling. Journal of Computational Chemistry, 24(12):1383–1389, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Choe:2005:EAC

[CN05] Yoong-Kee Choe and Shigeru Nagase. Effect of the axial cysteine ligand on the electronic structure and reactivity of high-valent iron(IV) oxo-porphyrins (Compound I): a theoretical study. Journal of Computational Chemistry, 26(15): 1600–1611, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Choe:2007:TSE

[CNN07] Yoong-Kee Choe, Shigeru Nagase, and Kichisuke Nishimoto. Theoretical study of the electronic spectra of oxidized and reduced states of lumiflavin and its derivative. Journal of Computational Chemistry, 28(4):727–739, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curutchet:2001:SOP

[COL01] Carles Curutchet, Modesto Orozco, and F. Javier Luque. Solvation in octanol: parametrization of the continuum MST model. Journal of Computational Chemistry, 22(11):1180– 1193, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curutchet:2006:DRC

[COL^{+}06] Carles Curutchet, Modesto Orozco, F. Javier Luque, Benedetta Mennucci, and Jacopo Tomasi. Dispersion and repulsion contributions to the solvation free energy: Comparison of quantum mechanical and classical approaches in the polarizable continuum model. Journal of Computational Chemistry, 27 (15):1769–1780, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Crespo-Otero:2004:TMI

[COMR^{+}04] Rachel Crespo-Otero, Luis A. Montero, Giselle Rosquete, J. Alexander Padrón-García, and Raúl H. González-Jonte. Theoretical model of internal rotation in monosubstituted derivatives of furfural. Journal of Computational Chemistry, 25(3):429–438, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Constans:2002:LSA

[Con02] Pere Constans. Linear scaling approaches to quantum macromolecular similarity: Evaluating the similarity function. Journal of Computational Chemistry, 23(14):1305–1313, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chae:2001:FMS

[COS01] Chong Hak Chae, Dong Gweon Oh, and Whanchul Shin. Flexible molecular superposition: development of a combined similarity index and application of the constrained optimization technique. Journal of Computational Chemistry, 22(8):888– 900, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Clarke:2008:TAP

[CP08] Oliver J. Clarke and Martin J. Parker. Time-averaged predictions of folded and misfolded peptides using a reduced physicochemical model. Journal of Computational Chemistry, 29(7): 1177–1185, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Clarke:2009:IAP

[CP09] Oliver J. Clarke and Martin J. Parker. Identification of amyloidogenic peptide sequences using a coarse-grained physicochemical model. Journal of Computational Chemistry, 30(4): 621–630, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chun:2000:MDM

[CPC^{+}00] Hon M. Chun, Carlos E. Padilla, Donovan N. Chin, Masakatsu Watanabe, Valeri I. Karlov, Howard E. Alper, Keto Soosaar, Kim B. Blair, Oren M. Becker, Leo S. D. Caves, Robert Nagle, David N. Haney, and Barry L. Farmer. MBO(N)D: a multibody method for long-time molecular dynamics simulations. Journal of Computational Chemistry, 21(3):159–184, February 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chandra:2008:BSC

[CPDZH08] Asit K. Chandra, Salma Parveen, Subojit Das, and Thérèse Zeegers-Huyskens. Blue shifts of the C H stretching vibrations in hydrogen-bonded and protonated trimethylamine. Effect of hyperconjugation on bond properties. Journal of Computational Chemistry, 29(9):1490–1496, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Castano:2002:RBB

[CPFL02] Obis Castaño, Raúl Palmeiro, Luis Manuel Frutos, and José Luisandrés. Role of bifurcation in the bond shifting of cyclooctatetraene. Journal of Computational Chemistry, 23(7):732– 736, May 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carballeira:2000:IIG

[CPJ00] Luis Carballeira and Ignacio Pérez-Juste. An ab initio interpretation in gas phase and aqueous solution of the generalized anomeric effect in R − O − CR_{2} − NR_{2} (R = H, CH_{3}). Journal of Computational Chemistry, 21(6):462–477, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Carballeira:2001:RAE

[CPJ01] Luis Carballeira and Ignacio Pérez-Juste. Role of the anomeric effect in methanediamines in the gas phase and aqueous solutions. Journal of Computational Chemistry, 22(2):135–150, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ceccarelli:2003:IFF

[CPM03] Matteo Ceccarelli, Piero Procacci, and Massimo Marchi. An ab initio force field for the cofactors of bacterial photosynthesis. Journal of Computational Chemistry, 24(2):129–142, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [Ano06b] and comments [Van08, CPML08b].

Csontos:2008:CWP

[CPML08a] József Csontos, Nicholas Y. Palermo, Richard F. Murphy, and Sándor Lovas. Calculation of weakly polar interaction energies in polypeptides using density functional and local Møller– Plesset perturbation theory. Journal of Computational Chemistry, 29(8):1344–1352, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Csontos:2008:RCA

[CPML08b] József Csontos, Nicholas Y. Palermo, Richard F. Murphy, and Sándor Lovas. Reply to “Comment on Aromatic-Backbone Interactions in Model α-Helical Peptides”. Journal of Computational Chemistry, 29(1):4–7, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [Van08].

Concu:2009:CCS

[CPUGD09] Riccardo Concu, Gianni Podda, Eugenio Uriarte, and Humberto González-Díaz. Computational chemistry study of 3Dstructure-function relationships for enzymes based on Markov models for protein electrostatic, HINT, and van der Waals potentials. Journal of Computational Chemistry, 30(9):1510– 1520, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chesnut:2004:NBS

[CQ04] D. B. Chesnut and L. D. Quin. Nature of bonding in the sulfuryl group. Journal of Computational Chemistry, 25(5): 734–738, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Comba:2002:NMM

[CR02] Peter Comba and Rainer Remenyi. A new molecular mechanics force field for the oxidized form of blue copper proteins. Journal of Computational Chemistry, 23(7):697–705, May 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chiodo:2008:DSO

[CR08] Sandro Chiodo and Nino Russo. Determination of spin-orbit coupling contributions in the framework of density functional theory. Journal of Computational Chemistry, 29(6):912–920, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chiodo:2009:OES

[CR09a] Sandro G. Chiodo and Nino Russo. One-electron spin-orbit contribution by effective nuclear charges. Journal of Computational Chemistry, 30(5):832–839, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cho:2009:EQM

[CR09b] Art E. Cho and David Rinaldo. Extension of QM/MM docking and its applications to metalloproteins. Journal of Computational Chemistry, 30(16):2609–2616, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Castillo:2008:BLE

[CRC^{+}08] Norberto Castillo, Katherine N. Robertson, S. C. Choi, Russell J. Boyd, and Osvald Knop. Bond length and the electron density at the bond critical point: X − X, Z − Z, and C − Z bonds (X = Li–F, Z = Na–Cl). Journal of Computational Chemistry, 29(3):367–379, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Choi:2001:NPO

[CRG01] Cheol Ho Choi, Klaus Ruedenberg, and Mark S. Gordon. New parallel optimal-parameter fast multipole method (OPFMM). Journal of Computational Chemistry, 22(13):1484–1501, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cuadrado:2007:QMB

[CRGN07] Manuel Urbano Cuadrado, Irene Luque Ruiz, and Miguel Ángel Gómez-Nieto. QSAR models based on isomorphic and nonisomorphic data fusion for predicting the blood brain barrier permeability. Journal of Computational Chemistry, 28(7): 1252–1260, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chowdry:2007:OOL

[CRH^{+}07] Arnab B. Chowdry, Kimberly A. Reynolds, Melinda S. Hanes, Mark Voorhies, Navin Pokala, and Tracy M. Handel. An object-oriented library for computational protein design. Journal of Computational Chemistry, 28(14):2378–2388, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Crippen:2004:CDG

[Cri04] Gordon M. Crippen. Cluster distance geometry of polypeptide chains. Journal of Computational Chemistry, 25(10):1305– 1312, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chiodo:2005:NDB

[CRS05] S. Chiodo, N. Russo, and E. Sicilia. Newly developed basis sets for density functional calculations. Journal of Computational Chemistry, 26(2):175–184, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cossi:2003:ESE

[CRSB03] Maurizio Cossi, Nadia Rega, Giovanni Scalmani, and Vincenzo Barone. Energies, structures, and electronic properties of molecules in solution with the C-PCM solvation model. Journal of Computational Chemistry, 24(6):669–681, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cioslowski:2001:FPC

[CS01] Jerzy Cioslowski and Agnieszka Szarecka. First-principles conformational analysis of the C_{36}H_{36} spheriphane, a prototype hydrocarbon host cage. Journal of Computational Chemistry, 22(12):1279–1286, September 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cai:2002:FAE

[CS02] Wensheng Cai and Xueguang Shao. A fast annealing evolutionary algorithm for global optimization. Journal of Computational Chemistry, 23(4):427–435, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cloud:2003:SED

[CS03] C. F. Cloud III and M. Schwartz. Systematic errors in DFT calculations of haloalkane heats of formation. Journal of Computational Chemistry, 24(5):640–646, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curutchet:2003:TFC

[CSB^{+}03] C. Curutchet, A. Salichs, X. Barril, Modesto Orozco, and F. Javier Luque. Transferability of fragmental contributions to the octanol/water partition coefficient: an NDDO-based MST study. Journal of Computational Chemistry, 24(1):32– 45, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chval:2008:TES

[CSB08] Zdenek Chval, Miroslav Sip, and Jaroslav V. Burda. The trans effect in square-planar platinum(II) complexes — a density functional study. Journal of Computational Chemistry, 29(14):2370–2381, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Coutsias:2004:UQC

[CSD04] Evangelos A. Coutsias, Chaok Seok, and Ken A. Dill. Using quaternions to calculate RMSD. Journal of Computational Chemistry, 25(15):1849–1857, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See comment [Kne05] and reply [CSD05].

Coutsias:2005:RSL

[CSD05] Evangelos A. Coutsias, Chaok Seok, and Ken A. Dill. Rotational superposition and least squares: the SVD and quaternions approaches yield identical results. Reply to the preceding comment by G. Kneller. Journal of Computational Chemistry, 26(15):1663–1665, November 30, 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [CSD04, Kne05].

Chandrasekhar:2001:EEC

[CSJ01] Jayaraman Chandrasekhar, Martin Saunders, and William L. Jorgensen. Efficient exploration of conformational space using the stochastic search method: application to β-peptide oligomers. Journal of Computational Chemistry, 22(14):1646– 1654, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Coutsias:2004:KVL

[CSJD04] Evangelos A. Coutsias, Chaok Seok, Matthew P. Jacobson, and Ken A. Dill. A kinematic view of loop closure. Journal of Computational Chemistry, 25(4):510–528, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cortes:2004:GAC

[CSRST04] J. Cortés, T. Siméon, M. Remaud-Siméon, and V. Tran. Geometric algorithms for the conformational analysis of long protein loops. Journal of Computational Chemistry, 25(7):956– 967, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cournia:2005:MMF

[CSU05] Zoe Cournia, Jeremy C. Smith, and G. Matthias Ullmann. A molecular mechanics force field for biologically important sterols. Journal of Computational Chemistry, 26(13):1383– 1399, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Crawford:2007:POS

[CSV^{+}07] T. Daniel Crawford, C. David Sherrill, Edward F. Valeev, Justin T. Fermann, Rollin A. King, Matthew L. Leininger, Shawn T. Brown, Curtis L. Janssen, Edward T. Seidl, Joseph P. Kenny, and Wesley D. Allen. PSI3: an open-source Ab Initio electronic structure package. Journal of Computational Chemistry, 28(9):1610–1616, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cui:2008:SSE

[CTFC08] Yan-Hong Cui, Wei Quan Tian, Ji-Kang Feng, and De-Li Chen. Structures, stabilities, electronic, and optical properties of C_{64} fullerene isomers, anions (C and C_{64}^{4−}), metallofullerene Sc_{2}@C_{64}, and Sc_{2}C_{2}@C_{64}. Journal of Computational Chemistry, 29(16):2623–2630, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chandra:2001:ROR

[CU01] Asit K. Chandra and Tadafumi Uchimaru. Reaction of OH radical with mono-, di-, and trichloroacetaldehyde: an ab initio study. Journal of Computational Chemistry, 22(13):1509– 1521, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cullen:2004:PHF

[Cul04] John Cullen. Performance of the Harris functional for extended basis sets at the Hartree–Fock and density functional levels. Journal of Computational Chemistry, 25(5):637–648, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cullen:2008:ADM

[Cul08] John Cullen. An approximate diatomics in molecules formulation of generalized valence bond theory. Journal of Computational Chemistry, 29(4):497–504, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chandra:2000:KHA

[CUS00] Asit K. Chandra, Tadafumi Uchimaru, and Masaaki Sugie. Kinetics of the hydrogen abstraction reactions of 1,1- and 1,2difluoroethane with hydroxyl radical: an ab initio study. Journal of Computational Chemistry, 21(14):1305–1318, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chandra:2003:ISS

[CUSS03] Asit K. Chandra, Tadafumi Uchimaru, Masaaki Sugie, and Akira Sekiya. Ab initio study on the structures of fluorinated formates and hydrogen abstraction reaction with OH radical. Journal of Computational Chemistry, 24(3):396–407, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Christ:2009:CTE

[CV09] Clara D. Christ and Wilfred F. Van Gunsteren. Comparison of three enveloping distribution sampling Hamiltonians for the estimation of multiple free energy differences from a single simulation. Journal of Computational Chemistry, 30(11): 1664–1679, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Christen:2008:SSD

[CvG08] Markus Christen and Wilfred F. van Gunsteren. On searching in, sampling of, and dynamically moving through conformational space of biomolecular systems: a review. Journal of Computational Chemistry, 29(2):157–166, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chelli:2008:RSD

[CVR08] Riccardo Chelli, Victor V. Volkov, and Roberto Righini. Retrieval of spectral and dynamic properties from twodimensional infrared pump-probe experiments. Journal of Computational Chemistry, 29(9):1507–1516, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chong:2004:ETS

[CVVB04] Delano P. Chong, Erik Van Lenthe, Stan Van Gisbergen, and Evert Jan Baerends. Even-tempered Slater-type orbitals revisited: From hydrogen to krypton. Journal of Computational Chemistry, 25(8):1030–1036, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chaumont:2002:MQT

[CW02] Alain Chaumont and Georges Wipff. Macrotricyclic quaternary tetraammonium receptors: Halide anion recognition and interfacial activity at an aqueous interface. A molecular dynamics investigation. Journal of Computational Chemistry, 23(16):1532–1543, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cho:2005:MDH

[CWV^{+}05] Art E. Cho, John A. Wendel, Nagarajan Vaidehi, Peter M. Kekenes-Huskey, Wely B. Floriano, Prabal K. Maiti, and William A. Goddard III. The MPSim–Dock hierarchical docking algorithm: Application to the eight trypsin inhibitor cocrystals. Journal of Computational Chemistry, 26(1):48–71, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cheng:2007:ESC

[CWWS07] L. Cheng, M. Y. Wang, Z. J. Wu, and Z. M. Su. Electronic structures and chemical bonding in 4d transition metal monohalides. Journal of Computational Chemistry, 28(13):2190– 2202, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Camacho:2009:ISM

[CWY09] Cristopher Camacho, Henryk A. Witek, and Shigeyoshi Yamamoto. Intruder states in multireference perturbation theory: the ground state of manganese dimer. Journal of Computational Chemistry, 30(3):468–478, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:EMI

[CXZ^{+}09] Xiaohua Chen, Dianxiang Xing, Liang Zhang, Robert I. Cukier, and Yuxiang Bu. Effect of metal ions on radical type and proton-coupled electron transfer channel: σ-Radical vs πradical and σ-channel vs π-channel in the imide units. Journal of Computational Chemistry, 30(16):2694–2705, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2009:PRW

[CY09] Mengen Chen and Wei Yang. On-the-path random walk sampling for efficient optimization of minimum free-energy path. Journal of Computational Chemistry, 30(11):1649–1653, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [CY13].

Chen:2013:EPR

[CY13] M. E. Chen and W. Yang. Erratum: On-the-path random walk sampling for efficient optimization of minimum free energy path. Journal of Computational Chemistry, 34(27):2400, October 15, 2013. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [CY09].

Chen:2002:CIS

[CYM02] I. Jen Chen, Daxu Yin, and Alexander D. MacKerell Jr. Combined ab initio /empirical approach for optimization of Lennard-Jones parameters for polar-neutral compounds. Journal of Computational Chemistry, 23(2):199–213, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Cao:2005:CAS

[CZ05] Zexing Cao and Qianer Zhang. Computational analyses of singlet–singlet and singlet–triplet transitions in mononuclear gold-capped carbon-rich conjugated complexes. Journal of Computational Chemistry, 26(12):1214–1221, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Curco:2003:EFS

[CZA03] David Curcó, David Zanuy, and Carlos Alemán. EVEBAT: a fast strategy for the examination of the empty space in polymer matrices. Journal of Computational Chemistry, 24(10): 1208–1214, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Call:2007:GMS

[CZB07] Seth T. Call, Dmitry Yu. Zubarev, and Alexander I. Boldyrev. Global minimum structure searches via particle swarm optimization. Journal of Computational Chemistry, 28(7):1177– 1186, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chen:2007:SEO

[CZFH07] Run-Feng Chen, Chao Zheng, Qu-Li Fan, and Wei Huang. Structural, electronic, and optical properties of 9-heterofluorenes: a quantum chemical study. Journal of Computational Chemistry, 28(13):2091–2101, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Delgado:2001:PID

[DA01] Eduardo J. Delgado and Joel B. Alderete. Prediction of infinite dilution activity coefficients of chlorinated organic compounds in aqueous solution from quantum-chemical descriptors. Journal of Computational Chemistry, 22(15):1851–1856, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dyrka:2008:IFT

[DAK08] Witold Dyrka, Andy T. Augousti, and Małgorzata Kotulska. Ion flux through membrane channels — an enhanced algorithm for the Poisson–Nernst–Planck model. Journal of Computational Chemistry, 29(12):1876–1888, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dominy:2002:INL

[DB02] Brian N. Dominy and Charles L. Brooks III. Identifying native-like protein structures using physics-based potentials. Journal of Computational Chemistry, 23(1):147–160, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Djurdjevic:2006:IPF

[DB06] Dusan P. Djurdjevic and Mark J. Biggs. Ab initio protein fold prediction using evolutionary algorithms: Influence of design and control parameters on performance. Journal of Computational Chemistry, 27(11):1177–1195, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Danecek:2007:CNS

[DB07] Petr Danĕček and Petr Bouř. Comparison of the numerical stability of methods for anharmonic calculations of vibrational molecular energies. Journal of Computational Chemistry, 28 (10):1617–1624, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Duca:2007:IID

[DBGV07] D. Duca, G. Barone, S. Giuffrida, and Zs. Varga. IDEA: Interface dynamics and energetics algorithm. Journal of Computational Chemistry, 28(15):2483–2499, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DelRio:2003:RMR

[DBM03] Alberto Del Rio, Abdou Boucekkine, and Jean Meinnel. Reassessment of methyl rotation barriers and conformations by correlated quantum chemistry methods. Journal of Computational Chemistry, 24(16):2093–2100, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DeMarothy:2007:EMR

[DBS07] S. A. De Marothy, M. R. A. Blomberg, and P. E. M. Siegbahn. Elucidating the mechanism for the reduction of nitrite by copper nitrite reductase — a contribution from quantum chemical studies. Journal of Computational Chemistry, 28(2): 528–539, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Duret:2008:SEI

[DBS08] Stefan Duret, Ahmed Bouferguene, and Hassan Safouhi. Strategies for an efficient implementation of the Gauss–Bessel quadrature for the evaluation of multicenter integral over STFs. Journal of Computational Chemistry, 29(6):934–944, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Davila:2002:AMT

[DC02] L. Y. A. Dávila and M. J. Caldas. Applicability of MNDO techniques AM1 and PM3 to ring-structured polymers. Journal of Computational Chemistry, 23(12):1135–1142, September 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dehareng:2000:HFI

[DD00] Dominique Dehareng and Georges Dive. Hartree–Fock instabilities and electronic properties. Journal of Computational Chemistry, 21(6):483–504, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Daga:2008:SPS

[DD08] Pankaj R. Daga and Robert J. Doerksen. Stereoelectronic properties of spiroquinazolinones in differential PDE7 inhibitory activity. Journal of Computational Chemistry, 29 (12):1945–1954, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dkhissi:2009:CHM

[DDBP09] Ahmed Dkhissi, Jean Marie Ducéré, Ralf Blossey, and Claude Pouchan. Can the hybrid meta GGA and DFT-D methods describe the stacking interactions in conjugated polymers? Journal of Computational Chemistry, 30(8):1179–1184, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Daeyaert:2007:AAC

[DDKV07] Frits Daeyaert, Marc De Jonge, Luc Koymans, and Maarten Vinkers. An ant algorithm for the conformational analysis of flexible molecules. Journal of Computational Chemistry, 28 (5):890–898, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Donckels:2009:KBM

[DDVD09] Brecht M. R. Donckels, Dirk J. W. De Pauw, Peter A. Vanrolleghem, and Bernard De Baets. A kernel-based method to determine optimal sampling times for the simultaneous estimation of the parameters of rival mathematical models. Journal of Computational Chemistry, 30(13):2064–2077, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Derreumaux:2000:PHH

[Der00] Philippe Derreumaux. Predicting helical hairpins from sequences by Monte Carlo simulations. Journal of Computational Chemistry, 21(7):582–589, May 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Derosa:2009:CSD

[Der09] Pedro A. Derosa. A combined semiempirical-DFT study of oligomers within the finite-chain approximation, evolution from oligomers to polymers. Journal of Computational Chemistry, 30(8):1220–1228, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Deeth:2004:PND

[DF04] Robert J. Deeth and Natalie Fey. The performance of nonhybrid density functionals for calculating the structures and spin states of Fe(II) and Fe(III) complexes. Journal of Computational Chemistry, 25(15):1840–1848, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Darowicki:2006:JTF

[DF06] K. Darowicki and W. Felisiak. On the joint time-frequency characteristics of chemical oscillations. Journal of Computational Chemistry, 27(8):961–965, June 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dronskowski:2008:F

[DF08] Richard Dronskowski and Gernot Frenking. Foreword. Journal of Computational Chemistry, 29(13):2039–2043, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Delalande:2009:CMA

[DFGB09] Olivier Delalande, Nicolas Férey, Gilles Grasseau, and Marc Baaden. Complex molecular assemblies at hand via interactive simulations. Journal of Computational Chemistry, 30(15):

2375–2387, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Deeth:2005:DIL

[DFWH05] Robert J. Deeth, Natalie Fey, and Benjamin Williams-Hubbard. DommiMOE: an implementation of ligand field molecular mechanics in the molecular operating environment. Journal of Computational Chemistry, 26(2):123–130, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DeJong:2005:OAE

[DGD^{+}05] G. Theodoor De Jong, Daan P. Geerke, Axel Diefenbach, Miquel Solà, and F. Matthias Bickelhaupt. Oxidative addition of the ethane C C bond to Pd. An ab initio benchmark and DFT validation study. Journal of Computational Chemistry, 26(10):1006–1020, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Davies:2002:PRM

[DGHR02] E. Keith Davies, Meir Glick, Karl N. Harrison, and W. Graham Richards. Pattern recognition and massively distributed computing. Journal of Computational Chemistry, 23(16): 1544–1550, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Donchev:2008:APG

[DGI^{+}08] Alexander G. Donchev, Nikolay G. Galkin, Alexey A. Illarionov, Oleg V. Khoruzhii, Michael A. Olevanov, Vladimir D. Ozrin, Leonid B. Pereyaslavets, and Vladimir I. Tarasov. Assessment of performance of the general purpose polarizable force field QMPFF3 in condensed phase. Journal of Computational Chemistry, 29(8):1242–1249, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

deGelder:2001:GES

[dGWH01] R. de Gelder, R. Wehrens, and J. A. Hageman. A generalized expression for the similarity of spectra: application to powder diffraction pattern classification. Journal of Computational Chemistry, 22(3):273–289, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Darian:2005:EAI

[DHF^{+}05] Eva Darian, Vladimir Hnizdo, Adam Fedorowicz, Harshinder Singh, and Eugene Demchuk. Estimation of the absolute internal-rotation entropy of molecules with two torsional degrees of freedom from stochastic simulations. Journal of Computational Chemistry, 26(7):651–660, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Das:2003:EEG

[DHM^{+}03] Abhijit K. Das, Jun-Ya Hasegawa, Tomoo Miyahara, Masahiro Ehara, and Hiroshi Nakatsuji. Electronic excitations of the green fluorescent protein chromophore in its protonation states: SAC/SAC–CI study. Journal of Computational Chemistry, 24(12):1421–1431, September 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dolney:2000:USM

[DHW^{+}00] Derek M. Dolney, Gregory D. Hawkins, Paul Winget, Daniel A. Liotard, Christopher J. Cramer, and Donald G. Truhlar. Universal solvation model based on conductor-like screening model. Journal of Computational Chemistry, 21(5): 340–366, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Du:2007:PRD

[DHW^{+}07] Qi-Shi Du, Ri-Bo Huang, Yu-Tuo Wei, Cheng-Hua Wang, and Kuo-Chen Chou. Peptide reagent design based on physical and chemical properties of amino acid residues. Journal of Computational Chemistry, 28(12):2043–2050, September 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Du:2008:MFT

[DHW^{+}08] Qi-Shi Du, Ri-Bo Huang, Yu-Tuo Wei, Li-Qin Du, and Kuo-Chen Chou. Multiple field three dimensional quantitative structure–activity relationship (MF-3D-QSAR). Journal of Computational Chemistry, 29(2):211–219, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Du:2009:FBQ

[DHW^{+}09] Qi-Shi Du, Ri-Bo Huang, Yu-Tuo Wei, Zong-Wen Pang, Li-Qin Du, and Kuo-Chen Chou. Fragment-based quantitative structure–activity relationship (FB-QSAR) for fragmentbased drug design. Journal of Computational Chemistry, 30 (2):295–304, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dibble:2005:CXT

[Dib05] Theodore S. Dibble. Computations on the Ã–X^{−}; transition of isoprene-OH-O_{2} peroxy radicals. Journal of Computational Chemistry, 26(8):836–845, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dinner:2000:LDP

[Din00] Aaron R. Dinner. Local deformations of polymers with nonplanar rigid main-chain internal coordinates. Journal of Computational Chemistry, 21(13):1132–1144, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Donnini:2004:CAP

[DJ04] Serena Donnini and André H. Juffer. Calculation of affinities of peptides for proteins. Journal of Computational Chemistry, 25(3):393–411, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Doruker:2002:DLP

[DJB02] Pemra Doruker, Robert L. Jernigan, and Ivet Bahar. Dynamics of large proteins through hierarchical levels of coarsegrained structures. Journal of Computational Chemistry, 23 (1):119–127, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Davis:2008:MMP

[DJT08] Alicia M. Davis, Gary Joanis, and Lorena Tribe. Molecular modeling of polymer–clay nanocomposite precursors: Lysine in montmorillonite interlayers. Journal of Computational Chemistry, 29(6):983–987, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Duan:2001:ATC

[DK01] Zhong-Hui Duan and Robert Krasny. An adaptive treecode for computing nonbonded potential energy in classical molecular systems. Journal of Computational Chemistry, 22(2):184–195, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dallos:2002:DEM

[DLD^{+}02] Michal Dallos, Hans Lischka, Elizete Ventura Do Monte, Michael Hirsch, and Wolfgang Quapp. Determination of energy minima and saddle points using multireference configuration interaction methods in combination with reduced gradient following: the S_{0} surface of H_{2}CO and the T_{1} and T_{2} surfaces of acetylene. Journal of Computational Chemistry, 23 (5):576–583, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

David:2000:CGB

[DLG00] Laurent David, Ray Luo, and Michael K. Gilson. Comparison of generalized Born and Poisson models: Energetics and dynamics of HIV protease. Journal of Computational Chemistry, 21(4):295–309, March 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Du:2006:HML

[DLHC06] Qi-Shi Du, Da-Peng Li, Wen-Zhang He, and Kuo-Chen Chou. Heuristic molecular lipophilicity potential (HMLP): Lipophilicity and hydrophilicity of amino acid side chains. Journal of Computational Chemistry, 27(6):685–692, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Deng:2008:CTS

[DLR^{+}08] Chao Deng, Qiang-Gen Li, Yi Ren, Ning-Bew Wong, San-Yan Chu, and Hua-Jie Zhu. A comprehensive theoretical study on the hydrolysis of carbonyl sulfide in the neutral water. Journal of Computational Chemistry, 29(3):466–480, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Das:2009:FES

[DLRZ09] Payel Das, Jingyuan Li, Ajay K. Royyuru, and Ruhong Zhou. Free energy simulations reveal a double mutant avian H5N1 virus hemagglutinin with altered receptor binding specificity. Journal of Computational Chemistry, 30(11):1654–1663, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DiBenedetto:2000:NPE

[DLSVY00] M. D. Di Benedetto, P. Lucibello, A. L. Sangiovanni-Vincentelli, and K. Yamaguchi. A new procedure for exact ring closure. Journal of Computational Chemistry, 21(10): 870–881, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dai:2006:NCD

[DLW06] Qi Dai, Xiao-Qing Liu, and Tian-Ming Wang. Numerical characterization of DNA sequences based on the k-step Markov chain transition probability. Journal of Computational Chemistry, 27(15):1830–1842, November 30, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dai:2007:LRM

[DLWV07] Qi Dai, Xiao-Qing Liu, Tian-Ming Wang, and Damir Vukicevic. Linear regression model of DNA sequences and its application. Journal of Computational Chemistry, 28(8):1434– 1445, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Du:2005:HML

[DMC05] Qishi Du, P. G. Mezey, and Kuo-Chen Chou. Heuristic molecular lipophilicity potential (HMLP): a 2D-QSAR study to LADH of molecular family pyrazole and derivatives. Journal of Computational Chemistry, 26(5):461–470, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Donnini:2005:IEI

[DMJV05] Serena Donnini, Alan E. Mark, André H. Juffer, and Alessandra Villa. Incorporating the effect of ionic strength in free energy calculations using explicit ions. Journal of Computational Chemistry, 26(2):115–122, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dixon:2005:QUS

[DMLI05] Steve Dixon, Kenneth M. Merz Jr., Giorgio Lauri, and James C. Ianni. QMQSAR: Utilization of a semiempirical probe potential in a field-based QSAR method. Journal of Computational Chemistry, 26(1):23–34, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Das:2003:PHM

[DMN03] B. Das, H. Meirovitch, and I. M. Navon. Performance of hybrid methods for large-scale unconstrained optimization as applied to models of proteins. Journal of Computational Chemistry, 24(10):1222–1231, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dobrogorskaia-Mereau:2005:QCM

[DMN05] Ia. I. Dobrogorskaia-Méreau and A. V. Nemukhin. Quantum chemical modeling of the reduction of cis-diammineplatinum(IV) tetrachloride [Pt(NH_{3})_{2}Cl_{4}] by methyl thiolate anion. Journal of Computational Chemistry, 26(9):865–870, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ding:2008:EBF

[DMZT08] Yun Ding, Ye Mei, John Z. H. Zhang, and Fu-Ming Tao. Efficient bond function basis set for π-π interaction energies. Journal of Computational Chemistry, 29(2):275–279, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

deOliveira:2001:DHF

[dOMSL01] Glênisson de Oliveira, Jan M. L. Martin, Indira K. C. Silwal, and Joel F. Liebman. Definitive heat of formation of methylenimine, CH_{2}NH, and of methylenimmonium ion, CH_{2}NH_{2}^{+}, by means of W2 theory. Journal of Computational Chemistry, 22(13):1297–1305, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Donchev:2008:NET

[Don08] Alexander G. Donchev. Nonadditive effects in ternary H_{2} – cation–PAH systems. Journal of Computational Chemistry, 29(11):1733–1739, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dudley:2006:PCP

[DOSG06] Timothy J. Dudley, Ryan M. Olson, Michael W. Schmidt, and Mark S. Gordon. Parallel coupled perturbed CASSCF equations and analytic CASSCF second derivatives. Journal of Computational Chemistry, 27(3):352–362, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dirks:2003:PFA

[DP03] Robert M. Dirks and Niles A. Pierce. A partition function algorithm for nucleic acid secondary structure including pseudoknots. Journal of Computational Chemistry, 24(13):1664– 1677, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dirks:2004:ACN

[DP04] Robert M. Dirks and Niles A. Pierce. An algorithm for computing nucleic acid base-pairing probabilities including pseudoknots. Journal of Computational Chemistry, 25(10):1295– 1304, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Depizzol:2005:MNG

[DPDG05] Daniela Bertolini Depizzol, Marcia Helena Moreira Paiva, Thiago Oliveira Dos Santos, and Anderson Coser Gaudio. Mo-Calc: a new graphical user interface for molecular calculations. Journal of Computational Chemistry, 26(2):142–144, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Devereux:2009:TIF

[DPM09] Michael Devereux, Paul L. A. Popelier, and Iain M. McLay. Toward an ab initio fragment database for bioisosterism: Dependence of QCT properties on level of theory, conformation, and chemical environment. Journal of Computational Chemistry, 30(8):1300–1318, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DeSancho:2005:EMA

[DPRR05] David De Sancho, Lidia Prieto, Ana M. Rubio, and Antonio Rey. Evolutionary method for the assembly of rigid protein fragments. Journal of Computational Chemistry, 26(2):131– 141, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Darkhovskii:2003:LHS

[DPT03] M. B. Darkhovskii, I. V. Pletnev, and A. L. Tchougréeff. Lowand high-spin iron (II) complexes studied by effective crystal field method combined with molecular mechanics. Journal of Computational Chemistry, 24(14):1703–1719, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DeSancho:2007:ECG

[DR07] David De Sancho and Antonio Rey. Evaluation of coarse grained models for hydrogen bonds in proteins. Journal of Computational Chemistry, 28(7):1187–1199, May 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Das:2009:PPV

[DR09] Ujjal Das and Krishnan Raghavachari. Predicting PH vibrations of gas phase molecules and surface-adsorbed species using bond length-frequency correlations. Journal of Computational Chemistry, 30(12):1872–1881, September 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Drablos:2000:ISC

[Dra00] Finn Drabløs. Ab initio simulation of chemical shift effects from metal ion binding in Bacitracin A. Journal of Computational Chemistry, 21(1):1–7, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DelCarmenMichelini:2004:ETA

[DRAS04] Maria Del Carmen Michelini, Nino Russo, Mohammad Esmaïl Alikhani, and Bernard Silvi. Energetic and topological analysis of the reaction of Mo and Mo_{2} with NH_{3}, C_{2}H_{2}, and C_{2}H_{4} molecules. Journal of Computational Chemistry, 25(13): 1647–1655, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DelCarmenMichelini:2005:ETA

[DRAS05] Maria Del Carmen Michelini, Nino Russo, Mohammad Esmaïl Alikhani, and Bernard Silvi. Energetic and topological analyses of the oxidation reaction between Mo_{n} (n = 1, 2) and N_{2}O. Journal of Computational Chemistry, 26(12):1284–1293, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

delRio:2000:TSR

[dRLMS00] E. del Río, R. López, M. I. Menéndez, and T. L. Sordo. A theoretical study of the reaction of HCO^{+} with C_{2}H_{2}. Journal of Computational Chemistry, 21(1):35–42, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Darrigan:2003:IFF

[DRMD03] Clovis Darrigan, Michel Rérat, Giuseppe Mallia, and Roberto Dovesi. Implementation of the finite field perturbation method in the CRYSTAL program for calculating the dielectric constant of periodic systems. Journal of Computational Chemistry, 24(11):1305–1312, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Derepas:2002:CWU

[DSB^{+}02] A.-L. Derepas, J.-M. Soudan, V. Brenner, J.-P. Dognon, and Ph. Millié. Can we understand the different coordinations and structures of closed-shell metal cation-water clusters? Journal of Computational Chemistry, 23(10):1013–1030, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

deSouza:2000:CPA

[dSGCG00] Hermes F. de Souza, Paulo H. Guadagnini, Rogério Custodio, and John D. Goddard. The correlation of proton affinities with atomic charges and electronegativities for the group 14 to 17 hydrides. Journal of Computational Chemistry, 21(13): 1119–1131, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Distasio:2007:IAA

[DSR^{+}07] Robert A. Distasio, Jr., Ryan P. Steele, Young Min Rhee, Yihan Shao, and Martin Head-Gordon. An improved algorithm for analytical gradient evaluation in resolution-of-the-identity second-order Møller–Plesset perturbation theory: Application to alanine tetrapeptide conformational analysis. Journal of Computational Chemistry, 28(5):839–856, April 15, 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

deSancho:2008:EMC

[dSR08] David de Sancho and Antonio Rey. Energy minimizations with a combination of two knowledge-based potentials for protein folding. Journal of Computational Chemistry, 29(10):1684– 1692, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Diaz:2003:CPP

[DSS03] Natalia Díaz, Dimas Suárez, and Tomás L. Sordo. Conformational properties of penicillins: Quantum chemical calculations and molecular dynamics simulations of benzylpenicillin. Journal of Computational Chemistry, 24(15):1864–1873, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

deSouza:2009:CMC

[dSVA^{+}09] Miguel A. F. de Souza, Elizete Ventura, Regiane C. M. U. Araújo, Mozart N. Ramos, and Silmar A. do Monte. CASSCF and multireference CI with singles and doubles study of lowlying valence and Rydberg states of 2H-tetrazole. Journal of Computational Chemistry, 30(7):1075–1081, May 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Duke:2001:CSO

[Duk01] Brian J. Duke. Choice of spin–orbit correction terms in Gaussian model chemistries. Journal of Computational Chemistry, 22(13):1552–1556, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dennis:2002:SSO

[DV02] S. Dennis and S. Vajda. Semiglobal simplex optimization and its application to determining the preferred solvation sites of proteins. Journal of Computational Chemistry, 23(3):319–334, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

deViterbo:2001:ANN

[dVB01] V. D. de Viterbo and J. C. Belchior. Artificial neural networks applied for studying metallic complexes. Journal of Computational Chemistry, 22(14):1691–1701, November 15, 2001. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Damm:2000:RPF

[DvG00] Wolfgang Damm and Wilfred F. van Gunsteren. Reversible peptide folding: Dependence on molecular force field used. Journal of Computational Chemistry, 21(9):774–787, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dijkstra:2001:SNU

[DvL01] Fokke Dijkstra and Joop H. van Lenthe. Software news and updates. Journal of Computational Chemistry, 22(6):665–672, April 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DeProft:2002:ACD

[DVP^{+}02] F. De Proft, C. Van Alsenoy, A. Peeters, W. Langenaeker, and P. Geerlings. Atomic charges, dipole moments, and Fukui functions using the Hirshfeld partitioning of the electron density. Journal of Computational Chemistry, 23(12):1198–1209, September 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

DeProft:2003:HPE

[DVRP^{+}03] F. De Proft, R. Vivas-Reyes, A. Peeters, C. Van Alsenoy, and P. Geerlings. Hirshfeld partitioning of the electron density: Atomic dipoles and their relation with functional group properties. Journal of Computational Chemistry, 24(4):463–470, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dai:2008:USM

[DW08] Qi Dai and Tian-Ming Wang. Use of statistical measures for analyzing RNA secondary structures. Journal of Computational Chemistry, 29(8):1292–1305, June 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Duan:2003:PCF

[DWC^{+}03] Yong Duan, Chun Wu, Shibasish Chowdhury, Mathew C. Lee, Guoming Xiong, Wei Zhang, Rong Yang, Piotr Cieplak, Ray Luo, Taisung Lee, James Caldwell, Junmei Wang, and Peter Kollman. A point-charge force field for molecular mechanics simulations of proteins based on condensed-phase quantum mechanical calculations. Journal of Computational Chemistry, 24(16):1999–2012, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dardenne:2001:RFU

[DWNB01] L. E. Dardenne, A. S. Werneck, M. Oliveira Neto, and P. M. Bisch. Reassociation of fragments using multicentered multipolar expansions: peptide junction treatments to investigate electrostatic properties of proteins. Journal of Computational Chemistry, 22(7):689–701, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Deng:2009:NHC

[DWS^{+}09] Chao Deng, Xiao-Peng Wu, Xiao-Ming Sun, Yi Ren, and Ying-Hong Sheng. Neutral hydrolyses of carbon disulfide: an ab initio study of water catalysis. Journal of Computational Chemistry, 30(2):285–294, January 30, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dai:2008:ESO

[DXW08] Dadi Dai, Hongjun Xiang, and Myung-Hwan Whangbo. Effects of spin-orbit coupling on magnetic properties of discrete and extended magnetic systems. Journal of Computational Chemistry, 29(13):2187–2209, October 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dyall:2002:SSR

[Dya02] Kenneth G. Dyall. A systematic sequence of relativistic approximations. Journal of Computational Chemistry, 23(8):

786–793, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Echenique:2006:DSA

[EA06] Pablo Echenique and J. L. Alonso. Definition of Systematic, Approximately Separable, and Modular Internal Coordinates (SASMIC) for macromolecular simulation. Journal of Computational Chemistry, 27(10):1076–1087, July 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Echenique:2008:EMC

[EA08] Pablo Echenique and José Luis Alonso. Efficient model chemistries for peptides. I. General framework and a study of the heterolevel approximation in RHF and MP2 with Pople split-valence basis sets. Journal of Computational Chemistry, 29(9):1408–1422, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Exner:2004:EFM

[EB04] Otto Exner and Stanislav Böhm. Enthalpies of formation of monoderivatives of hydrocarbons: Interaction of polar groups with an alkyl group. Journal of Computational Chemistry, 25 (16):1979–1986, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Eom:2007:CGP

[EBAN07] Kilho Eom, Seung-Chul Baek, Jung-Hee Ahn, and Sungsoo Na. Coarse-graining of protein structures for the normal mode studies. Journal of Computational Chemistry, 28(8):1400– 1410, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ewig:2001:DCI

[EBD^{+}01] Carl S. Ewig, Rajiv Berry, Uri Dinur, Jörg-Rüdiger Hill, Ming-Jing Hwang, Haiying Li, Chris Liang, Jon Maple, Zhengwei Peng, Thomas P. Stockfisch, Thomas S. Thacher, Lisa Yan, Xiangshan Ni, and Arnold T. Hagler. Derivation of class II force fields. VIII. Derivation of a general quantum mechanical force field for organic compounds. Journal of Computational Chemistry, 22(15):1782–1800, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

El-Bergmi:2000:SNB

[EBDPM00] Redouan El-Bergmi, J. A. Dobado, Dolores Portal, and José Molina Molina. Stabilization in neutral bicyclic sulfoxide compounds. Journal of Computational Chemistry, 21(4): 322–327, March 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Evarestov:2008:FPD

[EBL^{+}08] R. A. Evarestov, A. V. Bandura, M. V. Losev, E. A. Kotomin, Yu. F. Zhukovskii, and D. Bocharov. A first-principles DFT study of UN bulk and (001) surface: Comparative LCAO and PW calculations. Journal of Computational Chemistry, 29(13):2079–2087, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Echenique:2006:EFE

[EC06] Pablo Echenique and Iván Calvo. Explicit factorization of external coordinates in constrained statistical mechanics models. Journal of Computational Chemistry, 27(14):1748–1755, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Echenique:2006:QMC

[ECA06] Pablo Echenique, Iván Calvo, and J. L. Alonso. Quantum mechanical calculation of the effects of stiff and rigid constraints in the conformational equilibrium of the alanine dipeptide. Journal of Computational Chemistry, 27(14):1733– 1747, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Elstner:2003:MZB

[ECM^{+}03] Marcus Elstner, Qiang Cui, Petra Munih, Efthimios Kaxiras, Thomas Frauenheim, and Martin Karplus. Modeling zinc in biomolecules with the self consistent charge-density functional tight binding (SCC-DFTB) method: Applications to structural and energetic analysis. Journal of Computational Chemistry, 24(5):565–581, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Estrada:2004:QCD

[EDAJ04] Ernesto Estrada, Eduardo J. Delgado, Joel B. Alderete, and Gonzalo A. Jaña. Quantum-connectivity descriptors in modeling solubility of environmentally important organic compounds. Journal of Computational Chemistry, 25(14):1787– 1796, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ema:2003:PBS

[EdlVR^{+}03] I. Ema, J. M. García de la Vega, G. Ramírez, R. López, J. Fernández Rico, H. Meissner, and J. Paldus. Polarized basis sets of Slater-type orbitals: H to Ne atoms. Journal of Computational Chemistry, 24(7):859–868, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Elking:2007:GID

[EDW07] Dennis Elking, Tom Darden, and Robert J. Woods. Gaussian induced dipole polarization model. Journal of Computational Chemistry, 28(7):1261–1274, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Escudero:2009:IBA

[EFQD09] Daniel Escudero, Antonio Frontera, David Quiñonero, and Pere M. Deyà. Interplay between anion-π and hydrogen bonding interactions. Journal of Computational Chemistry, 30(1): 75–82, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ebel:2000:GEM

[EGSG00] Denton S. Ebel, Mark S. Ghiorso, Richard O. Sack, and Lawrence Grossman. Gibbs energy minimization in gas + liquid + solid systems. Journal of Computational Chemistry, 21 (4):247–256, March 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Eckert:2006:APB

[EK06] Frank Eckert and Andreas Klamt. Accurate prediction of basicity in aqueous solution with COSMO-RS. Journal of Computational Chemistry, 27(1):11–19, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Exner:2002:PRSa

[EKB02a] Thomas E. Exner, Matthias Keil, and Jürgen Brickmann. Pattern recognition strategies for molecular surfaces. I. Pattern generation using fuzzy set theory. Journal of Computational Chemistry, 23(12):1176–1187, September 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Exner:2002:PRSb

[EKB02b] Thomas E. Exner, Matthias Keil, and Jürgen Brickmann. Pattern recognition strategies for molecular surfaces. II. Surface complementarity. Journal of Computational Chemistry, 23 (12):1188–1197, September 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Endo:2001:AXX

[EKO^{+}01] K. Endo, S. Koizumi, T. Otsuka, M. Suhara, T. Morohasi, E. Z. Kurmaev, and D. P. Chong. Analysis of XPS and XES of diamond and graphite by DFT calculations using model molecules. Journal of Computational Chemistry, 22(1):102– 108, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ekanayake:2006:ABD

[EL06] Kaushalya S. Ekanayake and Pierre R. Lebreton. Activation barriers for DNA alkylation by carcinogenic methane diazonium ions. Journal of Computational Chemistry, 27(3):277– 286, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ekanayake:2007:MTS

[EL07] Kaushalya S. Ekanayake and Pierre R. Lebreton. Model transition states for methane diazonium ion methylation of guanine runs in oligomeric DNA. Journal of Computational Chemistry, 28(14):2352–2365, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Evarestov:2009:AEL

[EL09] R. A. Evarestov and M. V. Losev. All-electron LCAO calculations of the LiF crystal phonon spectrum: Influence of the basis set, the exchange-correlation functional, and the supercell size. Journal of Computational Chemistry, 30(16):2645–2655, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Eckert:2009:PAA

[ELK^{+}09] Frank Eckert, Ivo Leito, Ivari Kaljurand, Agnes Kütt, Andreas Klamt, and Michael Diedenhofen. Prediction of acidity in acetonitrile solution with COSMO-RS. Journal of Computational Chemistry, 30(5):799–810, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ellzey:2007:FGT

[Ell07] M. L. Ellzey Jr. Finite group theory for large systems. 3. Symmetry-generation of reduced matrix elements for icosahedral C_{20} and C_{60} molecules. Journal of Computational Chemistry, 28(4):811–817, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

English:2003:SDP

[EM03a] Niall J. English and J. M. D. Macelroy. Structural and dynamical properties of methane clathrate hydrates. Journal of Computational Chemistry, 24(13):1569–1581, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Exner:2003:IQP

[EM03b] Thomas E. Exner and Paul G. Mezey. Ab initio quality properties for macromolecules using the ADMA approach. Journal of Computational Chemistry, 24(16):1980–1986, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Eslami:2007:MDS

[EMP07] Hossein Eslami and Florian Müller-Plathe. Molecular dynamics simulation in the grand canonical ensemble. Journal of Computational Chemistry, 28(10):1763–1773, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Eyal:2004:ISA

[ENM^{+}04] Eran Eyal, Rafael Najmanovich, Brendan J. Mcconkey, Marvin Edelman, and Vladimir Sobolev. Importance of solvent accessibility and contact surfaces in modeling side-chain conformations in proteins. Journal of Computational Chemistry, 25(5):712–724, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Elezgaray:2000:MDP

[ES00] J. Elezgaray and Y. H. Sanejouand. Modal dynamics of proteins in water. Journal of Computational Chemistry, 21(14):

1274–1282, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Estiu:2006:QMM

[ESM06] Guillermina Estiu, Dimas Suárez, and Kenneth M. Merz. Quantum mechanical and molecular dynamics simulations of ureases and Zn β-lactamases. Journal of Computational Chemistry, 27(12):1240–1262, September 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

El-Sherbiny:2004:ERP

[ESP04] Aisha El-Sherbiny and Raymond A. Poirier. An evaluation of the radial part of numerical integration commonly used in DFT. Journal of Computational Chemistry, 25(11):1378– 1384, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Estrada:2007:PSN

[Est07] Ernesto Estrada. Point scattering: a new geometric invariant with applications from (Nano)clusters to biomolecules. Journal of Computational Chemistry, 28(4):767–777, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ferenczy:2001:III

[FÁ01a] György G. Ferenczy and János G. Ángyán. Intra- and intermolecular interactions in crystals of polar molecules. A study by the mixed quantum mechanical/molecular mechanical SCMP-NDDO method. Journal of Computational Chemistry, 22(14):1679–1690, November 15, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2001:EP

[FA01b] Gernot Frenking and Norman L. Allinger. Editors’ preface. Journal of Computational Chemistry, 22(13):vii–viii, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Field:2000:DLM

[FAB^{+}00] Martin J. Field, Marc Albe, Céline Bret, Flavien Proust-De Martin, and Aline Thomas. The dynamo library for molecular simulations using hybrid quantum mechanical and molecular mechanical potentials. Journal of Computational Chemistry, 21(12):1088–1100, September 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ferre:2002:SFF

[FAR02] Nicolas Ferré, Xavier Assfeld, and Jean-Louis Rivail. Specific force field parameters determination for the hybrid ab initio QM/MM LSCF method. Journal of Computational Chemistry, 23(6):610–624, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Faulon:2001:SGC

[Fau01] Jean-Loup Faulon. Stochastic generator of chemical structure. 4. Building polymeric systems with specified properties. Journal of Computational Chemistry, 22(6):580–590, April 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ferrari:2006:SPS

[FBDG06] E. S. Ferrari, R. C. Burton, R. J. Davey, and A. Gavezzotti. Simulation of phase separation in alcohol/water mixtures using two-body force field and standard molecular dynamics. Journal of Computational Chemistry, 27(11):1211–1219, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Forti:2008:EMC

[FBLO08] Flavio Forti, Xavier Barril, F. Javier Luque, and Modesto Orozco. Extension of the MST continuum solvation model to the RM1 semiempirical Hamiltonian. Journal of Computational Chemistry, 29(4):578–587, March 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2009:EN

[FBS09] Gernot Frenking, Charles L. Brooks III, and Shigeyoshi Sakiki. Editors’ note. Journal of Computational Chemistry, 30(1):1, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Formaneck:2006:UGB

[FC06] Mark S. Formaneck and Qiang Cui. The use of a generalized Born model for the analysis of protein conformational transitions: a comparative study with explicit solvent simulations for chemotaxis Y protein (CheY). Journal of Computational Chemistry, 27(16):1923–1943, December 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Feldmann:2008:MWB

[FCK^{+}08] Michael T. Feldmann, Julian C. Cummings, David R. Kent IV, Richard P. Muller, and William A. Goddard III. Manager– worker-based model for the parallelization of quantum Monte Carlo on heterogeneous and homogeneous networks. Journal of Computational Chemistry, 29(1):8–16, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fiaty:2004:ADA

[FCP^{+}04a] K. Fiaty, C. Charcosset, B. Perrin, R. Couturier, and B. Maïsterrena. ATP-dependent active transport simulations based on a phosphatase–channel–kinase membrane structure. Journal of Computational Chemistry, 25(10):1264–1276, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frediani:2004:NDS

[FCP^{+}04b] Luca Frediani, Roberto Cammi, Christian S. Pomelli, Jacopo Tomasi, and Kenneth Ruud. New developments in the symmetry-adapted algorithm of the Polarizable Continuum Model. Journal of Computational Chemistry, 25(3):375–385, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fiaty:2005:SAT

[FCP^{+}05] K. Fiaty, C. Charcosset, B. Perrin, R. Couturier, and B. Maïsterrena. Simulations of the active transport of a neutral solute based on a kinase-channel-phosphatase topology. Journal of Computational Chemistry, 26(3):201–213, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Falzon:2006:PSC

[FCW06] Chantal T. Falzon, Delano P. Chong, and Feng Wang. Prediction of spectroscopic constants for diatomic molecules in the ground and excited states using time-dependent density functional theory. Journal of Computational Chemistry, 27(2):

163–173, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Freeman:2003:CSC

[FD03] Fillmore Freeman and Elika Derek. A computational study of conformational interconversions in 1,4-dithiacyclohexane (1,4dithiane). Journal of Computational Chemistry, 24(8):909– 919, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fradera:2000:ATW

[FDM00] Xavier Fradera, Miquel Duran, and Jordi Mestres. Atomic transferability within the exchange-correlation density. Journal of Computational Chemistry, 21(15):1361–1374, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fores:2000:TCI

[FDSA00] Marta Forés, Miquel Duran, Miquel Solà, and Ludwik Adamowicz. Theoretical characterization of intramolecular proton transfer in the ground and the lowest-lying triplet excited states of 1-amino-3-propenal: a methodological comparison. Journal of Computational Chemistry, 21(4):257–269, March 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Friedrichs:2009:AMD

[FEV^{+}09] Mark S. Friedrichs, Peter Eastman, Vishal Vaidyanathan, Mike Houston, Scott Legrand, Adam L. Beberg, Daniel L. Ensign, Christopher M. Bruns, and Vijay S. Pande. Accelerating molecular dynamic simulation on graphics processing units. Journal of Computational Chemistry, 30(6):864–872, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fogolari:2001:MMD

[FEVM01] F. Fogolari, G. Esposito, P. Viglino, and H. Molinari. Molecular mechanics and dynamics of biomolecules using a solvent continuum model. Journal of Computational Chemistry, 22 (15):1830–1842, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fattebert:2002:DFT

[FG02] Jean-Luc Fattebert and François Gygi. Density functional theory for efficient ab initio molecular dynamics simulations in solution. Journal of Computational Chemistry, 23(6):662–666, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Facchini:2003:CRB

[FG03] Paola Facchini and Felice Grandinetti. Concerning the reaction between singlet nitrenium ions and water: a computational investigation on competitive reaction paths. Journal of Computational Chemistry, 24(5):547–564, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Faraldo-Gomez:2007:CCE

[FGR07] José D. Faraldo-Gómez and Benoît Roux. Characterization of conformational equilibria through Hamiltonian and temperature replica-exchange simulations: Assessing entropic and environmental effects. Journal of Computational Chemistry, 28(10):1634–1647, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fisher:2001:MPB

[FH01] Luke Fisher and Thomas Holme. MM3 parameterization for the B N dative bond. Journal of Computational Chemistry, 22(9):913–922, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2006:F

[FH06] Gernot Frenking and Max C. Holthausen. Foreword. Journal of Computational Chemistry, 27(12):1221–1222, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fukawa:2001:REH

[FHF^{+}01] Shuhei Fukawa, Masahiko Hada, Ryoichi Fukuda, Shinji Tanaka, and Hiroshi Nakatsuji. Relativistic effects and the halogen dependencies in the ^{13}C chemical shifts of CH_{4−n}I_{n}, CH_{4−n}Br_{n}, CCl_{4−n}I_{n}, and CBr_{4−n}I_{n} (n = 0–4). Journal of Computational Chemistry, 22(5):528–536, April 15, 2001.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fatmi:2007:QMC

[FHRR07] M. Qaiser Fatmi, Thomas S. Hofer, Bernhard R. Randolf, and Bernd M. Rode. Quantum mechanical charge field molecular dynamics simulation of the TiO^{2+} ion in aqueous solution. Journal of Computational Chemistry, 28(10):1704–1710, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Field:2002:SER

[Fie02] Martin J. Field. Simulating enzyme reactions: Challenges and perspectives. Journal of Computational Chemistry, 23(1):48– 58, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fedorov:2007:ATB

[FII^{+}07] Dmitri G. Fedorov, Kazuya Ishimura, Toyokazu Ishida, Kazuo Kitaura, Peter Pulay, and Shigeru Nagase. Accuracy of the three-body fragment molecular orbital method applied to Møller–Plesset perturbation theory. Journal of Computational Chemistry, 28(9):1476–1484, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2002:F

[FJ02] Gernot Frenking and William L. Jorgensen. Foreword. Journal of Computational Chemistry, 23(16):vi–vii, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fan:2008:TSF

[FJ08] Xiao-Wei Fan and Xue-Hai Ju. Theoretical studies on fourmembered ring compounds with NF_{2}, ONO_{2}, N_{3}, and NO_{2} groups. Journal of Computational Chemistry, 29(4):505–513, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ford:2007:AFC

[FJP07] Alan R. Ford, Tomasz Janowski, and Peter Pulay. Array files for computational chemistry: MP2 energies. Journal of Computational Chemistry, 28(7):1215–1220, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fedorov:2007:PIE

[FK07a] Dmitri G. Fedorov and Kazuo Kitaura. Pair interaction energy decomposition analysis. Journal of Computational Chemistry, 28(1):222–237, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2007:EUW

[FK07b] Gernot Frenking and Andreas Krapp. Essay: Unicorns in the world of chemical bonding models. Journal of Computational Chemistry, 28(1):15–24, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fisher:2008:OIA

[FKFG08] Daniel R. Fisher, David R. Kent IV, Michael T. Feldmann, and William A. Goddard III. An optimized initialization algorithm to ensure accuracy in quantum Monte Carlo calculations. Journal of Computational Chemistry, 29(14):2335– 2343, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

French:2001:HES

[FKJ^{+}01] Alfred D. French, Anne-Marie Kelterer, Glenn P. Johnson, Michael K. Dowd, and Christopher J. Cramer. HF/6-31G* energy surfaces for disaccharide analogs. Journal of Computational Chemistry, 22(1):65–78, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fedorov:2006:PCM

[FKL^{+}06] Dmitri G. Fedorov, Kazuo Kitaura, Hui Li, Jan H. Jensen, and Mark S. Gordon. The polarizable continuum model (PCM) interfaced with the fragment molecular orbital method (FMO). Journal of Computational Chemistry, 27(8):976–985, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fukuzawa:2006:III

[FKM^{+}06] Kaori Fukuzawa, Yuto Komeiji, Yuji Mochizuki, Akifumi Kato, Tatsuya Nakano, and Shigenori Tanaka. Intra- and intermolecular interactions between cyclic-AMP receptor protein and DNA: ab initio fragment molecular orbital study. Journal of Computational Chemistry, 27(8):948–960, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fukuzawa:2007:III

[FKM^{+}07] Kaori Fukuzawa, Yuto Komeiji, Yuji Mochizuki, Akifumi Kato, Tatsuya Nakano, and Shigenori Tanaka. Intra- and intermolecular interactions between cyclic-AMP receptor protein and DNA: ab initio fragment molecular orbital study. Journal of Computational Chemistry, 28(13):2237–2239, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fritsch:2008:TMD

[FKRE08] Daniel Fritsch, Klaus Koepernik, Manuel Richter, and Helmut Eschrig. Transition metal dimers as potential molecular magnets: a challenge to computational chemistry. Journal of Computational Chemistry, 29(13):2210–2219, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Futera:2009:IPS

[FKŠ^{+}09] Zdenĕk Futera, Julia Klenko, Judit E. Šponer, Jiří Šponer, and Jaroslav V. Burda. Interactions of the “piano-stool” [ruthenium(II) (η^{6}-arene)(en)CL]^{+} complexes with water and nucleobases; ab initio and DFT study. Journal of Computational Chemistry, 30(12):1758–1770, September 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fukuzawa:2005:IQM

[FKU^{+}05] Kaori Fukuzawa, Kazuo Kitaura, Masami Uebayasi, Kotoko Nakata, Tsuguchika Kaminuma, and Tatsuya Nakano. Ab initio quantum mechanical study of the binding energies of human estrogen receptor α with its ligands: an application of fragment molecular orbital method. Journal of Computational Chemistry, 26(1):1–10, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frickenhaus:2009:EES

[FKZ09] Stephan Frickenhaus, Srinivasaraghavan Kannan, and Martin Zacharias. Efficient evaluation of sampling quality of molecular dynamics simulations by clustering of dihedral torsion angles and Sammon mapping. Journal of Computational Chemistry, 30(3):479–492, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ferrante:2007:TSI

[FL07] Francesco Ferrante and Gianfranco La Manna. Theoretical study of the interaction between sodium ion and a cyclopeptidic tubular structure. Journal of Computational Chemistry, 28(13):2085–2090, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Flocke:2008:EEI

[FL08] N. Flocke and V. Lotrich. Efficient electronic integrals and their generalized derivatives for object oriented implementations of electronic structure calculations. Journal of Computational Chemistry, 29(16):2722–2736, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Feldgus:2000:NRT

[FLGW00] Steven Feldgus, Clark R. Landis, Eric D. Glendening, and Frank Weinhold. Natural resonance theory. I. General formalism. Journal of Computational Chemistry, 21(5):411–413, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2007:ESC

[FLK^{+}07] Gernot Frenking, Christoph Loschen, Andreas Krapp, Stefan Fau, and Steven H. Strauss. Electronic structure of CO — an exercise in modern chemical bonding theory. Journal of Computational Chemistry, 28(1):117–126, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ferro:2007:MES

[FLOD07] V. R. Ferro, R. López, S. Omar, and J. M. García De La Vega. On the molecular electron structure of three phosphininecontaining macrocycles. Journal of Computational Chemistry, 28(5):958–966, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Foloppe:2000:AAE

[FM00] Nicolas Foloppe and Alexander D. MacKerell, Jr. All-atom empirical force field for nucleic acids: I. Parameter optimization based on small molecule and condensed phase macromolecular target data. Journal of Computational Chemistry, 21(2):86–104, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Flores-Moreno:2006:HNE

[FMAMVK06] Roberto Flores-Moreno, Rodrigo J. Alvarez-Mendez, Alberto Vela, and Andreas M. Köster. Half-numerical evaluation of pseudopotential integrals. Journal of Computational Chemistry, 27(9):1009–1019, July 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Feixas:2008:PSA

[FMPS08] Ferran Feixas, Eduard Matito, Jordi Poater, and Miquel Solà. On the performance of some aromaticity indices: a critical assessment using a test set. Journal of Computational Chemistry, 29(10):1543–1554, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fornili:2006:SSL

[FMSA06] Arianna Fornili, Yohann Moreau, Maurizio Sironi, and Xavier Assfeld. On the suitability of strictly localized orbitals for hybrid QM/MM calculations. Journal of Computational Chemistry, 27(4):515–523, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fuchs:2006:NMP

[FNP^{+}06] Jean-François Fuchs, Hristo Nedev, David Poger, Michel Ferrand, Valérie Brenner, Jean-Pierre Dognon, and Serge Crouzy. New model potentials for sulfur–copper(I) and sulfur– mercury(II) interactions in proteins: From ab initio to molecular dynamics. Journal of Computational Chemistry, 27(7): 837–856, May 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fukaya:2004:DGC

[FO04] Haruhiko Fukaya and Taizo Ono. DFT–GIAO calculations of 19F NMR chemical shifts for perfluoro compounds. Journal of Computational Chemistry, 25(1):51–60, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frison:2008:CSS

[FO08] Gilles Frison and Gilles Ohanessian. A comparative study of semiempirical, ab initio, and DFT methods in evaluating metal–ligand bond strength, proton affinity, and interactions between first and second shell ligands in Zn-biomimetic complexes. Journal of Computational Chemistry, 29(3):416–433, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fedorov:2004:NHP

[FOK^{+}04] Dmitri G. Fedorov, Ryan M. Olson, Kazuo Kitaura, Mark S. Gordon, and Shiro Koseki. A new hierarchical parallelization scheme: Generalized distributed data interface (GDDI), and an application to the fragment molecular orbital method (FMO). Journal of Computational Chemistry, 25(6):872–880, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Feig:2004:PCG

[FOL^{+}04] Michael Feig, Alexey Onufriev, Michael S. Lee, Wonpil Im, David A. Case, and Charles L. Brooks III. Performance comparison of generalized Born and Poisson methods in the calculation of electrostatic solvation energies for protein structures. Journal of Computational Chemistry, 25(2):265–284, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fiorentino:2006:SVD

[FPG^{+}06] Anna Fiorentino, Deepangi Pandit, Kathleen M. Gilbert, Milind Misra, Rose Dios, and Carol A. Venanzi. Singular value decomposition of torsional angles of analogs of the dopamine reuptake inhibitor GBR 12909. Journal of Computational Chemistry, 27(5):609–620, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fujimoto:2005:MDS

[FPN^{+}05] Hirofumi Fujimoto, Miroslav Pinak, Toshiyuki Nemoto, Peter O’Neill, Etsuo Kume, Kimiaki Saito, and Hideaki Maekawa. Molecular dynamics simulation of clustered DNA damage sites containing 8-oxoguanine and abasic site. Journal of Computational Chemistry, 26(8):788–798, June 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fossgaard:2006:SSM

[FR06] Eirik Fossgård and Kenneth Ruud. Superlinear scaling in master-slave quantum chemical calculations using in-core storage of two-electron integrals. Journal of Computational Chemistry, 27(3):326–333, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2000:F

[Fre00] Gernot Frenking. Foreword. Journal of Computational Chemistry, 21(16):v, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fuhrmann:2009:NMG

[FRLN09] Jan Fuhrmann, Alexander Rurainski, Hans-Peter Lenhof, and Dirk Neumann. A new method for the gradient-based optimization of molecular complexes. Journal of Computational Chemistry, 30(9):1371–1378, July 15, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ferrero:2008:CSP

[FROD08] Mauro Ferrero, Michel Rérat, Roberto Orlando, and Roberto Dovesi. The calculation of static polarizabilities of 1–3D periodic compounds. the implementation in the CRYSTAL code. Journal of Computational Chemistry, 29(9):1450–1459, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Freire:2005:MLC

[FRS05] Ricardo O. Freire, Gerd B. Rocha, and Alfredo M. Simas. Modeling lanthanide complexes: Sparkle/AM1 parameters for ytterbium (III). Journal of Computational Chemistry, 26(14): 1524–1528, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Forester:1998:SRR

[FS98] Timothy R. Forester and William Smith. SHAKE, rattle, and roll: Efficient constraint algorithms for linked rigid bodies. Journal of Computational Chemistry, 19(1):102–111, January 15, 1998. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [FS00a].

Forester:2000:ESR

[FS00a] Timothy R. Forester and William Smith. Erratum: SHAKE, rattle, and roll: Efficient constraint algorithms for linked rigid bodies. Journal of Computational Chemistry, 21(2):157, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [FS98].

Frankcombe:2000:SUM

[FS00b] Terry J. Frankcombe and Sean C. Smith. Solving the unimolecular master equation with a weighted subspace projection method. Journal of Computational Chemistry, 21(8):592– 606, June 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fradera:2002:ELD

[FS02] Xavier Fradera and Miquel Solà. Electron localization and delocalization in open-shell molecules. Journal of Computational Chemistry, 23(14):1347–1356, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fradera:2004:SOA

[FS04] Xavier Fradera and Miquel Solà. Second-order atomic Fukui indices from the electron-pair density in the framework of the atoms in molecules theory. Journal of Computational Chemistry, 25(3):439–446, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Frenking:2007:F

[FS07] Gernot Frenking and Sason Shaik. Foreword. Journal of Computational Chemistry, 28(1):1–3, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Freindorf:2005:LJP

[FSFK05] Marek Freindorf, Yihan Shao, Thomas R. Furlani, and Jing Kong. Lennard-Jones parameters for the combined QM/MM method using the B3LYP/6-31G*/AMBER potential. Journal of Computational Chemistry, 26(12):1270–1278, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fajer:2009:UMF

[FSM09] Mikolai Fajer, Robert V. Swift, and J. Andrew McCammon. Using multistate free energy techniques to improve the efficiency of replica exchange accelerated molecular dynamics. Journal of Computational Chemistry, 30(11):1719–1725, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fuster:2000:DSS

[FSS00] Franck Fuster, Alain Sevin, and Bernard Silvi. Determination of substitutional sites in heterocycles from the topological analysis of the electron localization function (ELF). Journal of Computational Chemistry, 21(7):509–514, May 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Filizola:2001:GPP

[FTLV01] Marta Filizola, Silvina M. Tasso, Gilda H. Loew, and Hugo O. Villar. Global physicochemical properties as activity discriminants for the mGluR1 subtype of metabotropic glutamate receptors. Journal of Computational Chemistry, 22(16):2018– 2027, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fias:2008:MDI

[FVB08] Stijn Fias, Sofie Van Damme, and Patrick Bultinck. Multidimensionality of delocalization indices and nucleus independent chemical shifts in polycyclic aromatic hydrocarbons. Journal of Computational Chemistry, 29(3):358–366, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fischer:2007:IBC

[FWH^{+}07] Alexander Fischer, Sonja Waldhausen, Illia Horenko, Eike Meerbach, and Christof Schütte. Identification of biomolecular conformations from incomplete torsion angle observations by hidden Markov models. Journal of Computational Chemistry, 28(15):2453–2464, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fu:2006:CMT

[FZL^{+}06] Ke-Xiang Fu, Quan Zhu, Xiang-Yuan Li, Zhen Gong, Jian-Yi Ma, and Rong-Xing He. Continuous medium theory for nonequilibrium solvation: IV. Solvent reorganization energy of electron transfer based on conductor-like screening model. Journal of Computational Chemistry, 27(3):368–374, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fan:2007:CLS

[FZL07] W. J. Fan, R. Q. Zhang, and Shubin Liu. Computation of large systems with an economic basis set: Structures and reactivity indices of nucleic acid base pairs from density functional theory. Journal of Computational Chemistry, 28(5):967–974, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gadre:2003:F

[Gad03] Shridhar R. Gadre. Foreword. Journal of Computational Chemistry, 24(4):vii, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gillan:2008:HPC

[GAdGM08] M. J. Gillan, D. Alfè, S. de Gironcoli, and F. R. Manby. Highprecision calculation of Hartree–Fock energy of crystals. Journal of Computational Chemistry, 29(13):2098–2106, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Galano:2006:NAC

[GAI06] Annia Galano and J. Raúl Alvarez-Idaboy. A new approach to counterpoise correction to BSSE. Journal of Computational Chemistry, 27(11):1203–1210, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Galano:2001:OHA

[GAIMVB01] Annia Galano, J. Raúl Alvarez-Idaboy, Luis A. Montero, and Annik Vivier-Bunge. OH hydrogen abstraction reactions from alanine and glycine: a quantum mechanical approach. Journal of Computational Chemistry, 22(11):1138–1153, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ganesh:2009:MSC

[Gan09] V. Ganesh. MeTA studio: a cross platform, programmable IDE for computational chemist. Journal of Computational Chemistry, 30(4):661–672, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gromiha:2004:NNB

[GAS04] M. Michael Gromiha, Shandar Ahmad, and Makiko Suwa. Neural network-based prediction of transmembrane β-strand segments in outer membrane proteins. Journal of Computational Chemistry, 25(5):762–767, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Greene:2002:ICE

[GB02] Katherine R. Greene and Kyle A. Beran. Isomers of C_{20}: an energy profile. Journal of Computational Chemistry, 23(9): 938–942, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Guvench:2004:EAA

[GB04] Olgun Guvench and Charles L. Brooks III. Efficient approximate all-atom solvent accessible surface area method parameterized for folded and denatured protein conformations. Journal of Computational Chemistry, 25(8):1005–1014, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gould:2007:CIQ

[GBB07] Ian R. Gould, Hoda Abdel-Aal Bettley, and Richard A. Bryce. Correlated ab initio quantum chemical calculations of di- and trisaccharide conformations. Journal of Computational Chemistry, 28(12):1965–1973, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grybos:2009:INM

[GBBH09] Robert Grybos, Lubomir Benco, Tomas Bučko, and Jürgen Hafner. Interaction of NO molecules with Pd clusters: Ab initio density–functional study. Journal of Computational Chemistry, 30(12):1910–1922, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gohaud:2005:NPS

[GBDP05] Neil Gohaud, Didier Begue, Clovis Darrigan, and Claude Pouchan. New parallel software (P Anhar) for anharmonic vibrational calculations: Application to (CH_{3}Li)_{2}. Journal of Computational Chemistry, 26(7):743–754, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grimwood:2003:WFD

[GBJ03] Daniel J. Grimwood, Ian Bytheway, and Dylan Jayatilaka. Wave functions derived from experiment. V. Investigation of electron densities, electrostatic potentials, and electron localization functions for noncentrosymmetric crystals. Journal of Computational Chemistry, 24(4):470–483, March 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gdanitz:2005:RAE

[GBL^{+}05] Robert J. Gdanitz, Gary D. Black, Carina S. Lansing, Bruce J. Palmer, and Karen L. Schuchardt. Registering the Amica electronic structure code in the Extensible Computational Chemistry Environment. Journal of Computational Chemistry, 26 (3):214–225, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gomes:2002:EGE

[GC02] André Severo Pereira Gomes and Rogério Custodio. Exact Gaussian expansions of Slater-type atomic orbitals. Journal of Computational Chemistry, 23(10):1007–1012, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gill:2003:RQM

[GC03] Peter M. W. Gill and Siu-Hung Chien. Radial quadrature for multiexponential integrands. Journal of Computational Chemistry, 24(6):732–740, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gohlke:2004:CFE

[GC04] Holger Gohlke and David A. Case. Converging free energy estimates: MM-PB(GB)SA studies on the protein–protein complex Ras–Raf. Journal of Computational Chemistry, 25(2): 238–250, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gatti:2003:CIS

[GCB03] Carlo Gatti, Fausto Cargnoni, and Luca Bertini. Chemical information from the source function. Journal of Computational Chemistry, 24(4):422–436, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Garcia-Cruz:2000:DMM

[GCCVB00] Isidoro García-Cruz, Miguel Castro, and Annik Vivier-Bunge. DFT and MP2 molecular orbital determination of OH– toluene–O_{2} isomeric structures in the atmospheric oxidation of toluene. Journal of Computational Chemistry, 21(9):716–730, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Girones:2004:TFE

[GCD04] Xavier Gironés and Ramon Carbó-Dorca. TGSA-Flex: Extending the capabilities of the Topo-Geometrical superposition algorithm to handle flexible molecules. Journal of Computational Chemistry, 25(2):153–159, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grazioso:2008:ANA

[GCD^{+}08] Giovanni Grazioso, Andrea Cavalli, Marco De Amici, Maurizio Recanatini, and Carlo De Micheli. Alpha7 nicotinic acetylcholine receptor agonists: Prediction of their binding affinity through a molecular mechanics Poisson–Boltzmann surface area approach. Journal of Computational Chemistry, 29(15): 2593–2602, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gallegos:2005:MPD

[GCDL^{+}05] A. Gallegos, R. Carbó-Dorca, F. Lodier, E. Cancès, and A. Savin. Maximal probability domains in linear molecules. Journal of Computational Chemistry, 26(5):455–460, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez:2006:PCM

[GD06] Leticia González and Chantal Daniel. Photochemistry of CH_{3}Mn(CO)_{5}: a multiconfigurational ab initio study. Journal of Computational Chemistry, 27(15):1781–1786, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gillessen:2009:CSF

[GD09] Michael Gilleßen and Richard Dronskowski. A combinatorial study of full Heusler alloys by first-principles computational methods. Journal of Computational Chemistry, 30(8):1290– 1299, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez-diaz:2007:RCN

[GdAcV^{+}07] Humberto González-díaz, Guillermín Agüero-chapin, Javier Varona, Reinaldo Molina, Giovanna Delogu, Lourdes Santana, Eugenio Uriarte, and Gianni Podda. 2D-RNA-coupling numbers: a new computational chemistry approach to link secondary structure topology with biological function. Journal of Computational Chemistry, 28(6):1049–1056, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez-Diaz:2007:CCC

[GDPCPU07] Humberto González-Díaz, Yunierkis Pérez-Castillo, Gianni Podda, and Eugenio Uriarte. Computational chemistry comparison of stable/nonstable protein mutants classification models based on 3D and topological indices. Journal of Computational Chemistry, 28(12):1990–1995, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez-Diaz:2008:UQN

[GDPP08] Humberto González-Díaz and Francisco J. Prado-Prado. Unified QSAR and network-based computational chemistry approach to antimicrobials, part 1: Multispecies activity models for antifungals. Journal of Computational Chemistry, 29(4): 656–667, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez-diaz:2007:CCA

[GdSuM^{+}07] Humberto González-díaz, Liane Saíz-urra, Reinaldo Molina, Yenny González-díaz, and Angeles Sánchez-gonzález. Computational chemistry approach to protein kinase recognition using 3D stochastic van der Waals spectral moments. Journal of Computational Chemistry, 28(6):1042–1048, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Glattli:2003:NAD

[GDV03] Alice Glättli, Xavier Daura, and Wilfred F. Van Gunsteren. A novel approach for designing simple point charge models for liquid water with three interaction sites. Journal of Computational Chemistry, 24(9):1087–1096, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Galvan:2008:IEN

[GF08] Ignacio Fdez. Galván and Martin J. Field. Improving the efficiency of the NEB reaction path finding algorithm. Journal of Computational Chemistry, 29(1):139–143, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Genoni:2005:OVO

[GFS05] Alessandro Genoni, Arianna Fornili, and Maurizio Sironi. Optimal virtual orbitals to relax wave functions built up with transferred extremely localized molecular orbitals. Journal of Computational Chemistry, 26(8):827–835, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goette:2009:ACF

[GG09] Maik Goette and Helmut Grubmüller. Accuracy and convergence of free energy differences calculated from nonequilibrium switching processes. Journal of Computational Chemistry, 30 (3):447–456, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gill:2000:RET

[GGA00] Peter M. W. Gill, Andrew T. B. Gilbert, and Terry R. Adams. Rapid evaluation of two-center two-electron integrals. Journal of Computational Chemistry, 21(16):1505–1510, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez:2007:AERa

[GGB07a] Javier González, Xavier Giménez, and Josep Maria Bofill. Algorithm to evaluate rate constants for polyatomic chemical reactions. I. Theory and computational details. Journal of Computational Chemistry, 28(13):2102–2110, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez:2007:AERb

[GGB07b] Javier González, Xavier Giménez, and Josep Maria Bofill. Algorithm to evaluate rate constants for polyatomic chemical reactions. II. Applications. Journal of Computational Chemistry, 28(13):2111–2121, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonzalez-Garcia:2005:ESS

[GGGLL05] Núria González-García, Àngels González-Lafont, and José M. Lluch. Electronic structure study of the initiation routes of the dimethyl sulfide oxidation by OH. Journal of Computational Chemistry, 26(6):569–583, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Guvench:2008:AEF

[GGK^{+}08] Olgun Guvench, Shannon N. Greene, Ganesh Kamath, John W. Brady, Richard M. Venable, Richard W. Pastor, and Alexander D. Mackerell Jr. Additive empirical force field for hexopyranose monosaccharides. Journal of Computational Chemistry, 29(15):2543–2564, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Greatbanks:2000:CEP

[GGLR00] Stephen P. Greatbanks, Jill E. Gready, Ajay C. Limaye, and Alistair P. Rendell. Comparison of enzyme polarization of ligands and charge-transfer effects for dihydrofolate reductase using point-charge embedded ab initio quantum mechanical and linear-scaling semiempirical quantum mechanical methods. Journal of Computational Chemistry, 21(9):788–811, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Galvez:2009:CTT

[GGP09] Oscar Gálvez, Pedro C. Gómez, and Luis F. Pacios. Characterization of two types of intermolecular interactions on halogen monoxide monohydrates. Journal of Computational Chemistry, 30(15):2538–2549, November 30, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gruziel:2008:PBM

[GGT08] Magdalena Gruziel, Pawel Grochowski, and Joanna Trylska. The Poisson–Boltzmann model for tRNA: Assessment of the calculation set-up and ionic concentration cutoff. Journal of Computational Chemistry, 29(12):1970–1981, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Guerra:2004:VDD

[GHBB04] Célia Fonseca Guerra, Jan-Willem Handgraaf, Evert Jan Baerends, and F. Matthias Bickelhaupt. Voronoi deformation density (VDD) charges: Assessment of the Mulliken, Bader, Hirshfeld, Weinhold, and VDD methods for charge analysis. Journal of Computational Chemistry, 25(2):189–210, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ghulghazaryan:2007:ECW

[GHH07] Ruben G. Ghulghazaryan, Shura Hayryan, and Chin-Kun Hu. Efficient combination of Wang–Landau and transition matrix Monte Carlo methods for protein simulations. Journal of Computational Chemistry, 28(3):715–726, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Garcia-Hernandez:2002:CSR

[GHLK^{+}02] Maite García-Hernández, Christa Lauterbach, Sven Krüger, Alexei Matveev, and Notker Rösch. Comparative study of relativistic density functional methods applied to actinide species AcO_{2}^{2+} and AcF_{6} for Ac = U, Np. Journal of Computational Chemistry, 23(8):834–846, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gordon:2003:ERO

[GHMP03] D. Benjamin Gordon, Geoffrey K. Hom, Stephen L. Mayo, and Niles A. Pierce. Exact rotamer optimization for protein design. Journal of Computational Chemistry, 24(2):232–243, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grabuleda:2000:MDS

[GJK00] Xavier Grabuleda, Carlos Jaime, and Peter A. Kollman. Molecular dynamics simulation studies of liquid acetonitrile: New six-site model. Journal of Computational Chemistry, 21 (10):901–908, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Geudtner:2006:PDC

[GJK^{+}06] Gerald Geudtner, Florian Janetzko, Andreas M. Köster, Alberto Vela, and Patrizia Calaminici. Parallelization of the deMon2k code. Journal of Computational Chemistry, 27(4): 483–490, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gao:2008:EFP

[GJL^{+}08] Bin Gao, Jun Jiang, Kai Liu, Ziyu Wu, Wei Lu, and Yi Luo. An efficient first-principle approach for electronic structures calculations of nanomaterials. Journal of Computational Chemistry, 29(3):434–444, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Galstyan:2009:ARP

[GK09] Artur Galstyan and Ernst-Walter Knapp. Accurate redox potentials of mononuclear iron, manganese, and nickel model complexes*. Journal of Computational Chemistry, 30(2):203– 211, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gotz:2005:OAB

[GKH05] Andreas W. Götz, Christian Kollmar, and Bernd A. Hess. Optimization of auxiliary basis sets for the LEDO expansion and a projection technique for LEDO–DFT. Journal of Computational Chemistry, 26(12):1242–1253, September 2005. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gront:2007:BBQ

[GKK07] Dominik Gront, Sebastian Kmiecik, and Andrzej Kolinski. Backbone building from quadrilaterals: a fast and accurate algorithm for protein backbone reconstruction from alpha carbon coordinates. Journal of Computational Chemistry, 28(9): 1593–1597, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ganesh:2008:WWI

[GKRG08] V. Ganesh, Ritwik Kavathekar, Anuja Rahalkar, and Shridhar R. Gadre. WebProp: Web interface for ab initio calculation of molecular one-electron properties. Journal of Computational Chemistry, 29(3):488–495, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gresh:2004:IIE

[GKTS04] Nohad Gresh, Sherif A. Kafafi, Jean-François Truchon, and Dennis R. Salahub. Intramolecular interaction energies in model alanine and glycine tetrapeptides. Evaluation of anisotropy, polarization, and correlation effects. A parallel ab initio HF/MP2, DFT, and polarizable molecular mechanics study. Journal of Computational Chemistry, 25(6):823–834, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gallicchio:2004:AAI

[GL04a] Emilio Gallicchio and Ronald M. Levy. AGBNP: an analytic implicit solvent model suitable for molecular dynamics simulations and high-resolution modeling. Journal of Computational Chemistry, 25(4):479–499, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gillilan:2004:ODP

[GL04b] Richard E. Gillilan and Ryan H. Lilien. Optimization and dynamics of protein–protein complexes using B-splines. Journal of Computational Chemistry, 25(13):1630–1646, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Georgiev:2008:MDE

[GLD08] Ivelin Georgiev, Ryan H. Lilien, and Bruce R. Donald. The minimized dead-end elimination criterion and its application to protein redesign in a hybrid scoring and search algorithm for computing partition functions over molecular ensembles. Journal of Computational Chemistry, 29(10):1527–1542, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Geng:2008:TER

[GLH^{+}08] Cai-Yun Geng, Ji-Lai Li, Xu-Ri Huang, Hui-Ling Liu, Zhuo Li, and Chia-Chung Sun. Theoretical elucidation of the rhodiumcatalyzed [4 + 2] annulation reactions. Journal of Computational Chemistry, 29(5):686–693, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Geerke:2009:DCF

[GLMV09] Daan P. Geerke, Sandra Luber, Koni H. Marti, and Wilfred F. Van Gunsteren. On the direct calculation of the free energy of quantization for molecular systems in the condensed phase. Journal of Computational Chemistry, 30(4):514–523, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gallicchio:2008:ARE

[GLP08] Emilio Gallicchio, Ronald M. Levy, and Manish Parashar. Asynchronous replica exchange for molecular simulations. Journal of Computational Chemistry, 29(5):788–794, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Golebiowski:2002:RSC

[GLRL02] Jérôme Golebiowski, Véronique Lamare, and Manuel F. Ruiz-López. Rb^{+} /cs^{+} selectivity of benzo and tribenzo derivatives of the 21C7 crown ether. A density functional study. Journal of Computational Chemistry, 23(7):724–731, May 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Glykos:2006:SNU

[Gly06] Nicholas M. Glykos. Software news and updates carma: a molecular dynamics analysis program. Journal of Computational Chemistry, 27(14):1765–1768, November 15, 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goh:2001:AFA

[GM01] Sor-Koon Goh and Dennis S. Marynick. Ability of fullerenes to act as η^{6} ligands in transition metal complexes. A comparative PM3(tm)–density functional theory study. Journal of Computational Chemistry, 22(16):1881–1886, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goncalves:2004:NAH

[GM04] Cristina P. Gonçalves and José R. Mohallem. A new algorithm to handle finite nuclear mass effects in electronic calculations: the ISOTOPE program. Journal of Computational Chemistry, 25(14):1736–1739, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Galvan:2004:NML

[GMA04] I. Fdez. Galván, M. E. Martín, and M. A. Aguilar. A new method to locate saddle points for reactions in solution by using the free-energy gradient method and the mean field approximation. Journal of Computational Chemistry, 25(10): 1227–1233, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gogtas:2008:TDQ

[Gog08] Fahrettin Gogtas. Time-dependent quantum study of the kinetics of the H(^{2}S) + FO(^{2}Π) → OH(^{2}Π) + F(^{2}P) reaction. Journal of Computational Chemistry, 29(12):1889–1894, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Golmohammadi:2009:POW

[Gol09] Hassan Golmohammadi. Prediction of octanol–water partition coefficients of organic compounds by multiple linear regression, partial least squares, and artificial neural network. Journal of Computational Chemistry, 30(15):2455–2465, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gonnet:2007:SAA

[Gon07] Pedro Gonnet. A simple algorithm to accelerate the computation of non-bonded interactions in cell-based molecular dynamics simulations. Journal of Computational Chemistry, 28(2):570–573, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goraczko:2001:MMF

[Gor01] Andrzej J. Gora̧czko. Multiisotopic modeling of fragmentation ion patterns in mass spectra of organometallic and coordination compounds. Journal of Computational Chemistry, 22(3): 354–365, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gianese:2006:CPI

[GP06] Giulio Gianese and Stefano Pascarella. A consensus procedure improving solvent accessibility prediction. Journal of Computational Chemistry, 27(5):621–626, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gresh:2005:RZI

[GPK05] Nohad Gresh, Jean-Philip Piquemal, and Morris Krauss. Representation of Zn(II) complexes in polarizable molecular mechanics. Further refinements of the electrostatic and shortrange contributions. Comparisons with parallel ab initio computations. Journal of Computational Chemistry, 26(11):1113– 1130, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grant:2001:SPF

[GPN01] J. Andrew Grant, Barry T. Pickup, and Anthony Nicholls. A smooth permittivity function for Poisson–Boltzmann solvation methods. Journal of Computational Chemistry, 22(6):608– 640, April 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gourlaouen:2006:RGM

[GPSP06] Christophe Gourlaouen, Jean-Philip Piquemal, Trond Saue, and Olivier Parisel. Revisiting the geometry of nd^{10}(n + 1)s^{0} [M(H_{2}O)]^{p+} complexes using four-component relativistic DFT calculations and scalar relativistic correlated CSOV energy decompositions (M^{p+} = Cu^{+}, Zn^{2+}, Ag^{+}, Cd^{2+}, Au^{+}, Hg^{2+}). Journal of Computational Chemistry, 27(2):142–156, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gillespie:2007:EGL

[GR07] R. J. Gillespie and E. A. Robinson. Essay: Gilbert N. Lewis and the chemical bond: the electron pair and the octet rule from 1916 to the present day. Journal of Computational Chemistry, 28(1):87–97, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grafton:2007:VFW

[Gra07] Anthony K. Grafton. Vibalizer: a free, Web-based tool for rapid, quantitative comparison and analysis of calculated vibrational modes. Journal of Computational Chemistry, 28(7): 1290–1305, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Girones:2001:TMS

[GRCD01] Xavier Gironés, David Robert, and Ramon Carbó-Dorca. TGSA: a molecular superposition program based on topogeometrical considerations. Journal of Computational Chemistry, 22(2):255–263, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grimme:2003:ITO

[Gri03] Stefan Grimme. Improved third-order Møller–Plesset perturbation theory. Journal of Computational Chemistry, 24(13): 1529–1537, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grimme:2004:ADV

[Gri04] Stefan Grimme. Accurate description of van der Waals complexes by density functional theory including empirical corrections. Journal of Computational Chemistry, 25(12):1463–1473, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grimme:2006:SGT

[Gri06] Stefan Grimme. Semiempirical GGA-type density functional constructed with a long-range dispersion correction. Journal of Computational Chemistry, 27(15):1787–1799, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Garavelli:2003:SAI

[GRO^{+}03] Marco Garavelli, Fabrizio Ruggeri, François Ogliaro, Michael J. Bearpark, Fernando Bernardi, Massimo Olivucci, and Michael A. Robb. A simple approach for improving the hybrid MMVB force field: Application to the photoisomerization of s-cis butadiene. Journal of Computational Chemistry, 24(11): 1357–1363, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goncalves:2002:NAF

[GS02] Paulo F. B. Gonçalves and Hubert Stassen. New approach to free energy of solvation applying continuum models to molecular dynamics simulation. Journal of Computational Chemistry, 23(7):706–714, May 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goncalves:2003:FES

[GS03] Paulo F. B. Gonçalves and Hubert Stassen. Free energy of solvation from molecular dynamics simulations for low dielectric solvents. Journal of Computational Chemistry, 24(14):1758– 1765, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gresh:2004:CDI

[GS04] Nohad Gresh and Gen-Bin Shi. Conformation-dependent intermolecular interaction energies of the triphosphate anion with divalent metal cations. Application to the ATP-binding site of a binuclear bacterial enzyme. A parallel quantum chemical and polarizable molecular mechanics investigation. Journal of Computational Chemistry, 25(2):160–168, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Goodman:2007:DFN

[GS07] Lionel Goodman and Ronald R. Sauers. Diffuse functions in natural bond orbital analysis. Journal of Computational Chemistry, 28(1):269–275, January 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grohmann:2008:TRQ

[GS08] Rainer Grohmann and Torsten Schindler. Toward robust QSPR models: Synergistic utilization of robust regression and variable elimination. Journal of Computational Chemistry, 29 (6):847–860, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gidofalvi:2009:CDE

[GS09] Gergely Gidofalvi and Ron Shepard. Computation of determinant expansion coefficients within the graphically contracted function method. Journal of Computational Chemistry, 30 (15):2414–2419, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gil:2009:IIC

[GSB09] Adrià Gil, Mariona Sodupe, and Juan Bertran. Influence of ionization on the conformational preferences of peptide models. Ramachandran surfaces of N-formyl-glycine amide and N-formyl-alanine amide radical cations. Journal of Computational Chemistry, 30(12):1771–1784, September 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gorecki:2009:RRM

[GSDT09] Adam Górecki, Marcin Szypowski, Maciej Długosz, and Joanna Trylska. RedMD — reduced molecular dynamics package. Journal of Computational Chemistry, 30(14):2364–2373, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gautier:2003:CAS

[GT03] R. Gautier and P. Tufféry. Critical assessment of side-chain conformational space sampling procedures designed for quantifying the effect of side-chain environment. Journal of Computational Chemistry, 24(15):1950–1961, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gherman:2006:CSR

[GTC06] Benjamin F. Gherman, William B. Tolman, and Christopher J. Cramer. Characterization of the structure and reactivity of monocopper–oxygen complexes supported by βdiketiminate and anilido-imine ligands. Journal of Computational Chemistry, 27(16):1950–1961, December 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tao:2001:TSR

[gThDjL^{+}01] Yu guo Tao, Yi hong Ding, Jian jun Liu, Ze sheng Li, Xu ri Huang, and Chia-Chung Sun. Theoretical study on reaction mechanism of the CF radical with nitrogen dioxide. Journal of Computational Chemistry, 22(16):1907–1919, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Garcia-Viloca:2003:HTC

[GVATG03] Mireia Garcia-Viloca, Cristóbal Alhambra, Donald G. Truhlar, and Jiali Gao. Hydride transfer catalyzed by xylose isomerase: Mechanism and quantum effects. Journal of Computational Chemistry, 24(2):177–190, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gu:2007:IGQ

[GWL07] Jiande Gu, Jing Wang, and Jerzy Leszczynski. Iso-guanine quintet complexes coordinated by mono valent cations (Na^{+}, K^{+}, Rb^{+}, and Cs^{+}). Journal of Computational Chemistry, 28(11):1790–1795, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gooding:2000:FPQ

[GWM^{+}00] Stuart R. Gooding, Peter J. Winn, Richard I. Maurer, György G. Ferenczy, John R. Miller, Jayne E. Harris, D. Vaughan Griffiths, and Christopher A. Reynolds. Fully polarizable QM/MM calculations: an application to the nonbonded iodine–oxygen interaction in dimethyl-2-iodobenzoylphosphonate. Journal of Computational Chemistry, 21(6):478–482, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gao:2008:CQM

[GWM08] Jiali Gao, Kin-Yiu Wong, and Dan T. Major. Combined QM/ MM and path integral simulations of kinetic isotope effects in the proton transfer reaction between nitroethane and acetate ion in water. Journal of Computational Chemistry, 29(4):514– 522, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Guvench:2002:AFA

[GWS^{+}02] Olgun Guvench, Jörg Weiser, Peter Shenkin, István Kolossváry, and W. Clark Still. Application of the frozen atom approximation to the GB/SA continuum model for solvation free energy. Journal of Computational Chemistry, 23(2):214–221, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gao:2009:TSF

[GXK09] Guohua Gao, Xiang Xu, and Hong Seok Kang. A theoretical study on fullerene-dizincocene hybrids. Journal of Computational Chemistry, 30(6):978–982, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gregersen:2006:CSI

[GY06] Brent A. Gregersen and Darrin M. York. A charge-scaling implementation of the variational electrostatic projection method. Journal of Computational Chemistry, 27(1):103–115, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Giese:2008:EAT

[GY08] Timothy J. Giese and Darrin M. York. Extension of adaptive tree code and fast multipole methods to high angular momentum particle charge densities. Journal of Computational Chemistry, 29(12):1895–1904, September 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gao:2004:TSS

[GYCZ04] Xingfa Gao, Hui Yuan, Zhenling Chen, and Yuliang Zhao. Theoretical studies of structures and stabilities of a new oddnumbered fullerene dimer: C_{141}. Journal of Computational Chemistry, 25(16):2023–2030, December 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Garcia-Yoldi:2007:OTE

[GYMN07] Iñigo García-Yoldi, Fernando Mota, and Juan J. Novoa. The origin of the two-electron/four-centers C − C bond in π-TCNE_{2}^{2−} dimers: Electrostatic or dispersion? Journal of Computational Chemistry, 28(1):326–334, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gao:2007:WSN

[GZ07] Xingfa Gao and Yuliang Zhao. The way of stabilizing non-IPR fullerenes and structural elucidation of C_{54}Cl_{8}. Journal of Computational Chemistry, 28(4):795–801, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Gallicchio:2002:SNH

[GZL02] Emilio Gallicchio, Linda Yu Zhang, and Ronald M. Levy. The SGB/NP hydration free energy model based on the surface generalized Born solvent reaction field and novel nonpolar hydration free energy estimators. Journal of Computational Chemistry, 23(5):517–529, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Grosdidier:2009:BDP

[GZM09] Aurélien Grosdidier, Vincent Zoete, and Olivier Michielin. Blind docking of 260 protein–ligand complexes with EADock 2.0. Journal of Computational Chemistry, 30(13):2021–2030, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Howard:2004:ICH

[HA04] Siân T. Howard and Colin D. Abernethy. Intramolecular C H… C_{carbene} hydrogen bonds and competing interactions in monoprotonated tripodal carbenes. Journal of Computational Chemistry, 25(5):649–659, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hafner:2008:ISM

[Haf08] Jürgen Hafner. Ab-initio simulations of materials using VASP: Density-functional theory and beyond. Journal of Computational Chemistry, 29(13):2044–2078, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hamacher:2007:ITM

[Ham07] K. Hamacher. Information theoretical measures to analyze trajectories in rational molecular design. Journal of Computational Chemistry, 28(16):2576–2580, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Han:2001:DFS

[Han01] Young-Kyu Han. Density functional studies of AnF_{6} (An = U, Np, and Pu) and UF_{6−n}Cl_{n} (n = 1–6) using hybrid functionals: geometries and vibrational frequencies. Journal of Computational Chemistry, 22(16):2010–2017, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Harrison:2004:KSA

[Har04] Robert J. Harrison. Krylov subspace accelerated inexact Newton method for linear and nonlinear equations. Journal of Computational Chemistry, 25(3):328–334, February 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hamaneh:2009:RPH

[HB09] Mehdi Bagheri Hamaneh and Matthias Buck. Refinement of the primary hydration shell model for molecular dynamics simulations of large proteins. Journal of Computational Chemistry, 30(16):2635–2644, December 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Harvey:2006:QMM

[HBM06] Jeremy N. Harvey, Christine M. Bathelt, and Adrian J. Mulholland. QM/MM modeling of compound I active species in cytochrome P450, cytochrome C peroxidase, and ascorbate peroxidase. Journal of Computational Chemistry, 27(12): 1352–1362, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Holst:2000:AMF

[HBW00] M. Holst, N. Baker, and F. Wang. Adaptive multilevel finite element solution of the Poisson–Boltzmann equation I. Algorithms and examples. Journal of Computational Chemistry, 21(15):1319–1342, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [HBW01].

Holst:2001:EAM

[HBW01] M. Holst, N. Baker, and F. Wang. Erratum: Adaptive multilevel finite element solution of the Poisson–Boltzmann equation I. Algorithms and examples. Journal of Computational Chemistry, 22(4):475, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [HBW00].

Haberthur:2008:FFA

[HC08] Urs Haberthür and Amedeo Caflisch. FACTS: Fast analytical continuum treatment of solvation. Journal of Computational Chemistry, 29(5):701–715, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haiduke:2006:ARAb

[HD06] Roberto L. A. Haiduke and Albérico B. F. Da Silva. Accurate relativistic adapted Gaussian basis sets for cesium through radon without variational prolapse and to be used with both uniform sphere and Gaussian nucleus models. Journal of Computational Chemistry, 27(16):1970–1979, December 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haiduke:2004:PVG

[HDBD04] Roberto L. A. Haiduke, Luiz G. M. De Macedo, Rugles C. Barbosa, and Albérico B. F. Da Silva. A polynomial version of the generator coordinate Dirac–Fock method. Journal of Computational Chemistry, 25(15):1904–1909, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hnizdo:2007:NNN

[HDF^{+}07] Vladimir Hnizdo, Eva Darian, Adam Fedorowicz, Eugene Demchuk, Shengqiao Li, and Harshinder Singh. Nearestneighbor nonparametric method for estimating the configurational entropy of complex molecules. Journal of Computational Chemistry, 28(3):655–668, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haiduke:2005:ARU

[HdMdS05] Roberto L. A. Haiduke, Luiz G. M. de Macedo, and Albérico B. F. da Silva. An accurate relativistic universal Gaussian basis set for hydrogen through Nobelium without variational prolapse and to be used with both uniform sphere and Gaussian nucleus models. Journal of Computational Chemistry, 26 (9):932–940, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hori:2002:SCC

[HDO^{+}02] Kenzi Hori, Nobumitsu Dou, Katsuhiko Okano, Ai Ohgami, and Hiroshi Tsukube. Stable conformations of 12-crown-O_{3}N and its Li^{+} complex in aqueous solution. Journal of Computational Chemistry, 23(13):1226–1235, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haiduke:2006:ARAa

[HdS06] Roberto L. A. Haiduke and Albérico B. F. da Silva. Accurate relativistic adapted Gaussian basis sets for hydrogen through xenon without variational prolapse and to be used with both uniform sphere and Gaussian nucleus models. Journal of Computational Chemistry, 27(1):61–71, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Helal:2009:IMS

[HELM09] Wissam Helal, Stefano Evangelisti, Thierry Leininger, and Daniel Maynau. Ab-initio multireference study of an organic mixed-valence Spiro molecular system. Journal of Computational Chemistry, 30(1):83–92, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hardin:2002:FFK

[HEP^{+}02] Corey Hardin, Michael P. Eastwood, Michael Prentiss, Z. Luthey-Schulten, and Peter G. Wolynes. Folding funnels: the key to robust protein structure prediction. Journal of Computational Chemistry, 23(1):138–146, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Herzog:2006:EBP

[HFHL06] Elena Herzog, Tomaso Frigato, Volkhard Helms, and C. Roy D. Lancaster. Energy barriers of proton transfer reactions between amino acid side chain analogs and water from ab initio calculations. Journal of Computational Chemistry, 27(13):1534–1547, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hasegawa:2007:ESG

[HFS^{+}07] Jun-Ya Hasegawa, Kazuhiro Fujimoto, Ben Swerts, Tomoo Miyahara, and Hiroshi Nakatsuji. Excited states of GFP chromophore and active site studied by the SAC–CI method: Effect of protein-environment and mutations. Journal of Computational Chemistry, 28(15):2443–2452, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hnizdo:2003:STI

[HFSD03] Vladimir Hnizdo, Adam Fedorowicz, Harshinder Singh, and Eugene Demchuk. Statistical thermodynamics of internal rotation in a hindering potential of mean force obtained from computer simulations. Journal of Computational Chemistry, 24(10):1172–1183, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2008:APP

[HG08] Liang-Tsung Huang and M. Michael Gromiha. Analysis and prediction of protein folding rates using quadratic response surface models. Journal of Computational Chemistry, 29(10): 1675–1683, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hayes:2004:FFD

[HGMB04] Joseph M. Hayes, James C. Greer, and David A. Morton-Blake. A force-field description of short-range repulsions for high density alkane molecular dynamics simulations. Journal of Computational Chemistry, 25(16):1953–1966, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Heinz:2004:FPC

[HH04] Tim N. Heinz and Philippe H. Hünenberger. A fast pairlistconstruction algorithm for molecular simulations under periodic boundary conditions. Journal of Computational Chemistry, 25(12):1474–1486, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hall:2000:AHQ

[HHBH00] Richard J. Hall, Sally A. Hindle, Neil A. Burton, and Ian H. Hillier. Aspects of hybrid QM/MM calculations: the treatment of the QM/MM interface region and geometry optimization with an application to chorismate mutase. Journal of Computational Chemistry, 21(16):1433–1441, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hahnke:2009:PPA

[HHG^{+}09] Volker Hähnke, Bettina Hofmann, Tomislav Grgat, Ewgenij Proschak, Dieter Steinhilber, and Gisbert Schneider. PhAST: Pharmacophore alignment search tool. Journal of Computational Chemistry, 30(5):761–771, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hayryan:2001:MPS

[HHHS01] Shura Hayryan, Chin-Kun Hu, Shun-Yun Hu, and Rung-Ji Shang. Multicanonical parallel simulations of proteins with continuous potentials. Journal of Computational Chemistry, 22(12):1287–1296, September 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Habershon:2003:DMP

[HHJ03] Scott Habershon, Kenneth D. M. Harris, and Roy L. Johnston. Development of a multipopulation parallel genetic algorithm for structure solution from powder diffraction data. Journal of Computational Chemistry, 24(14):1766–1774, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hudaky:2004:SMI

[HHP04] Ilona Hudáky, Péter Hudáky, and András Perczel. Solvation model induced structural changes in peptides. A quantum chemical study on Ramachandran surfaces and conformers of alanine diamide using the polarizable continuum model. Journal of Computational Chemistry, 25(12):1522–1531, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hayryan:2005:NAM

[HHS^{+}05] Shura Hayryan, Chin-Kun Hu, Jaroslav Skřivánek, Edik Hayryane, and Imrich Pokorný. A new analytical method for computing solvent-accessible surface area of macromolecules and its gradients. Journal of Computational Chemistry, 26 (4):334–343, March 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Horng:2003:CMC

[HHW^{+}03] Jorng-Tzong Horng, Hsien-Da Huang, Shih-Hsien Wang, Ming-You Chen, Shir-Ly Huang, and Jenn-Kang Hwang. Computing motif correlations in proteins. Journal of Computational Chemistry, 24(16):2032–2043, December 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haranczyk:2008:SSO

[HHWG08] Maciej Haranczyk, John Holliday, Peter Willett, and Maciej Gutowski. Structure and singly occupied molecular orbital analysis of anionic tautomers of guanine. Journal of Computational Chemistry, 29(8):1277–1291, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hirama:2003:PMH

[HIA03] Mutsumi Hirama, Toshimasa Ishida, and Jun-Ichi Aihara. Possible molecular hydrogen formation mediated by the radical cations of anthracene and pyrene. Journal of Computational Chemistry, 24(12):1378–1382, September 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Heine:2007:CIM

[HIM07] Thomas Heine, Rafael Islas, and Gabriel Merino. σ and π contributions to the induced magnetic field: Indicators for the mobility of electrons in molecules. Journal of Computational Chemistry, 28(1):302–309, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hinsen:2000:MMT

[Hin00] Konrad Hinsen. The molecular modeling toolkit: a new approach to molecular simulations. Journal of Computational Chemistry, 21(2):79–85, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hirao:2008:RBO

[Hir08] Hajime Hirao. A reactive bond orbital investigation of the Diels–Alder reaction between 1,3-butadiene and ethylene: Energy decomposition, state correlation diagram, and electron density analyses. Journal of Computational Chemistry, 29(9): 1399–1407, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hudaky:2001:PMX

[HJCP01] Péter Hudáky, Imre Jákli, Attila G. Császár, and András Perczel. Peptide models XXXI. Conformational properties of hydrophobic residues shaping the core of proteins. An ab initio study of N-formyl- L-valinamide and N-formyl- Lphenylalaninamide. Journal of Computational Chemistry, 22 (7):732–751, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hou:2007:DSD

[HK07] Jin Qiang Hou and Hong Seok Kang. A DFT study on the dimerization of C_{62}, H_{2}C_{62}, and F_{2}C_{62}. Journal of Computational Chemistry, 28(8):1417–1426, June 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Holt:2008:EIQ

[HK08a] Asbjørn Holt and Gunnar Karlström. Erratum: Inclusion of the quadrupole moment when describing polarization. The effect of the dipole-quadrupole polarizability. Journal of Computational Chemistry, 29(14):2485–2486, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [HK08b].

Holt:2008:IQM

[HK08b] Asbjørn Holt and Gunnar Karlström. Inclusion of the quadrupole moment when describing polarization. The effect of the dipole-quadrupole polarizability. Journal of Computational Chemistry, 29(12):2033–2038, September 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [HK08a].

Holt:2008:ICM

[HK08c] Asbjørn Holt and Gunnar Karlstrøm. Induction correction model for rotation of two or three dihedral angles. Journal of Computational Chemistry, 29(12):1905–1911, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Holt:2008:IIC

[HK08d] Asbjørn Holt and Gunnar Karlström. An intramolecular induction correction model of the molecular dipole moment. Journal of Computational Chemistry, 29(7):1084–1091, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Honda:2008:ECD

[HKHN08] Yasushi Honda, Atsushi Kurihara, Masahiko Hada, and Hiroshi Nakatsuji. Excitation and circular dichroism spectra of (-)-(3a S, 7a S)-2-chalcogena-trans-hydrindans (Ch = S, Se, Te): SAC and SAC–CI calculations. Journal of Computational Chemistry, 29(4):612–621, March 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Han:2001:AMS

[HKMS01] Kyu-Kwang Han, Ki Hae Kim, Byung Jin Mhin, and Hyeon S. Son. Application of the multiensemble sampling to the hydration free energy. Journal of Computational Chemistry, 22(10): 1004–1009, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hu:2008:USV

[HL08] Xiuzhen Hu and Qianzhong Li. Using support vector machine to predict β- and γ-turns in proteins. Journal of Computational Chemistry, 29(12):1867–1875, September 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2009:MDS

[HLB09] Yongqi Huang, Huifang Li, and Yuxiang Bu. Molecular dynamics simulation exploration of cooperative migration mechanism of calcium ions in sarcoplasmic reticulum Ca^{2+}-ATPase. Journal of Computational Chemistry, 30(13):2136–2145, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hofto:2009:IAT

[HLC09] Laura R. Hofto, Caroline E. Lee, and Mauricio Cafiero. The importance of aromatic-type interactions in serotonin synthesis: Protein–ligand interactions in tryptophan hydroxylase and aromatic amino acid decarboxylase. Journal of Computational Chemistry, 30(7):1111–1115, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Han:2006:DCM

[HLLN06] Wen-Ge Han, Tiqing Liu, Timothy Lovell, and Louis Noodleman. DFT calculations of ^{57}Fe Mössbauer isomer shifts and quadrupole splittings for iron complexes in polar dielectric media: Applications to methane monooxygenase and ribonucleotide reductase. Journal of Computational Chemistry, 27 (12):1292–1306, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

He:2005:TSR

[HLLS05] Hong-Qing He, Jing-Yao Liu, Ze-Sheng Li, and Chia-Chung Sun. Theoretical study for the reaction of CH_{3}OCl with Cl atom. Journal of Computational Chemistry, 26(6):642–650, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2005:AFF

[HLM05] Po-Ssu Huang, John J. Love, and Stephen L. Mayo. Adaptation of a fast Fourier transform-based docking algorithm for protein design. Journal of Computational Chemistry, 26(12): 1222–1232, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hugosson:2006:CTS

[HLMR06] Håkan W. Hugosson, Alessandro Laio, Patrick Maurer, and Ursula Rothlisberger. A comparative theoretical study of dipeptide solvation in water. Journal of Computational Chemistry, 27(5):672–684, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Henn:2007:CIM

[HLS07] Julian Henn, Dirk Leusser, and Dietmar Stalke. Chemical interpretation of molecular electron density distributions. Journal of Computational Chemistry, 28(14):2317–2324, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Horenko:2005:AAN

[HLSH05] Illia Horenko, Sönke Lorenz, Christof Schütte, and Wilhelm Huisinga. Adaptive approach for nonlinear sensitivity analysis of reaction kinetics. Journal of Computational Chemistry, 26 (9):941–948, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2005:IHM

[HLT^{+}05] Hsien-Da Huang, Tzong-Yi Lee, Shih-Wei Tzeng, Li-Cheng Wu, Jorng-Tzong Horng, Ann-Ping Tsou, and Kuan-Tsae Huang. Incorporating hidden Markov models for identifying protein kinase-specific phosphorylation sites. Journal of Computational Chemistry, 26(10):1032–1041, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hong:2009:ROP

[HLTLP09] Eun-Jong Hong, Shaun M. Lippow, Bruce Tidor, and Tomás Lozano-Pérez. Rotamer optimization for protein design through MAP estimation and problem-size reduction. Journal of Computational Chemistry, 30(12):1923–1945, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hauck:2001:GSS

[HM01] J. Hauck and K. Mika. Ground-state structures of polymers. Journal of Computational Chemistry, 22(16):1944–1955, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Henchman:2002:EHS

[HM02] Richard H. Henchman and J. Andrew McCammon. Extracting hydration sites around proteins from explicit water simulations. Journal of Computational Chemistry, 23(9):861–869, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hom:2006:SAF

[HM06] Geoffrey K. Hom and Stephen L. Mayo. A search algorithm for fixed-composition protein design. Journal of Computational Chemistry, 27(3):375–378, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hemmateenejad:2008:AQT

[HM08] Bahram Hemmateenejad and Afshan Mohajeri. Application of quantum topological molecular similarity descriptors in QSPR study of the O-methylation of substituted phenols. Journal of Computational Chemistry, 29(2):266–274, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hu:2006:MCS

[HMD06] Jie Hu, Ao Ma, and Aaron R. Dinner. Monte Carlo simulations of biomolecules: the MC module in CHARMM. Journal of Computational Chemistry, 27(2):203–216, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huo:2002:CAS

[HMK02] Shuanghong Huo, Irina Massova, and Peter A. Kollman. Computational alanine scanning of the 1:1 human growth hormone–receptor complex. Journal of Computational Chemistry, 23(1):15–27, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hemmateenejad:2009:AMA

[HMMS09] Bahram Hemmateenejad, Ahmad R. Mehdipour, Ramin Miri, and Mojtaba Shamsipur. Application of MOLMAP approach for QSAR modeling of various biological activities using substituent electronic descriptors. Journal of Computational Chemistry, 30(13):2001–2009, October 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huey:2007:SFE

[HMOG07] Ruth Huey, Garrett M. Morris, Arthur J. Olson, and David S. Goodsell. A semiempirical free energy force field with chargebased desolvation. Journal of Computational Chemistry, 28 (6):1145–1152, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hemmateenejad:2006:APB

[HMSM06] Bahram Hemmateenejad, Ramin Miri, Mohammad A. Safarpour, and Ahmad R. Mehdipour. Accurate prediction of the blood–brain partitioning of a large set of solutes using ab initio calculations and genetic neural network modeling. Journal of Computational Chemistry, 27(11):1125–1135, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Haberthur:2003:EEE

[HMWC03] Urs Haberthür, Nicolas Majeux, Philipp Werner, and Amedeo Caflisch. Efficient evaluation of the effective dielectric function of a macromolecule in aqueous solution. Journal of Computational Chemistry, 24(15):1936–1949, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Higo:2002:HSH

[HN02] Junichi Higo and Masayoshi Nakasako. Hydration structure of human lysozyme investigated by molecular dynamics simulation and cryogenic X-ray crystal structure analyses: On the correlation between crystal water sites, solvent density, and solvent dipole. Journal of Computational Chemistry, 23(14): 1323–1336, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hogberg:2008:MCF

[HNL08] Carl-Johan Högberg, Alexei M. Nikitin, and Alexander P. Lyubartsev. Modification of the CHARMM force field for DMPC lipid bilayer. Journal of Computational Chemistry, 29(14):2359–2369, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Herrmann:2008:QMV

[HNR08] Carmen Herrmann, Johannes Neugebauer, and Markus Reiher. QM/MM vibrational mode tracking. Journal of Computational Chemistry, 29(14):2460–2470, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Heringer:2007:AES

[HNWF07] D. Heringer, T. A. Niehaus, M. Wanko, and Th. Frauenheim. Analytical excited state forces for the time-dependent densityfunctional tight-binding method. Journal of Computational Chemistry, 28(16):2589–2601, December 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [HNWF12].

Heringer:2012:EAE

[HNWF12] D. Heringer, T. A. Niehaus, M. Wanko, and Th. Frauenheim. Erratum: “Analytical excited state forces for the time-dependent density-functional tight-binding method” [J. Comp. Chem. 28, 2589]. Journal of Computational Chemistry, 33(5):593, February 15, 2012. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [HNWF07].

Hofinger:2005:SPB

[Höf05] Siegfried Höfinger. Solving the Poisson–Boltzmann equation with the specialized computer chip MD-GRAPE-2. Journal of Computational Chemistry, 26(11):1148–1154, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Holthausen:2005:BAD

[Hol05] Max C. Holthausen. Benchmarking approximate density functional theory. I. s/d excitation energies in 3d transition metal cations. Journal of Computational Chemistry, 26(14):1505– 1518, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hassinen:2001:NET

[HP01] Tommi Hassinen and Mikael Peräkylä. New energy terms for reduced protein models implemented in an off-lattice force field. Journal of Computational Chemistry, 22(12):1229–1242, September 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hill:2004:RSP

[HP04] Jörg-Rüdiger Hill and Johann Plank. Retardation of setting of plaster of Paris by organic acids: Understanding the mechanism through molecular modeling. Journal of Computational Chemistry, 25(12):1438–1448, September 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hudaky:2005:TDD

[HP05] Péter Hudáky and András Perczel. Toward direct determination of conformations of protein building units from multidimensional NMR experiments VI. Chemical shift analysis of his to gain 3D structure and protonation state information. Journal of Computational Chemistry, 26(13):1307–1317, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hernandes:2003:CSL

[HPL03] Marcelo Zaldini Hernandes, João Bosco P. Da Silva, and Ricardo L. Longo. Chemometric study of liquid water simulations. I. The parameters of the TIP4P model potential. Journal of Computational Chemistry, 24(8):973–981, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hart:2000:ESB

[HPP00] Reece K. Hart, Rohit V. Pappu, and Jay W. Ponder. Exploring the similarities between potential smoothing and simulated annealing. Journal of Computational Chemistry, 21 (7):531–552, May 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hirsch:2002:IRM

[HQ02] Michael Hirsch and Wolfgang Quapp. Improved RGF method to find saddle points. Journal of Computational Chemistry, 23(9):887–894, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hill:2008:GFE

[HR08] Anthony D. Hill and Peter J. Reilly. A Gibbs free energy correlation for automated docking of carbohydrates. Journal of Computational Chemistry, 29(7):1131–1141, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hermida-Ramon:2003:IIP

[HRBKB03] Jose Manuel Hermida-Ramón, Steve Brdarski, Gunnar Karlström, and Ulf Berg. Inter- and intramolecular potential for the N-formylglycinamide-water system. A comparison between theoretical modeling and empirical force fields. Journal of Computational Chemistry, 24(2):161–176, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hermida-Ramon:2007:BSH

[HRG07] Jose M. Hermida-Ramón and Ana M. Graña. Blueshifting hydrogen bond in the benzene–benzene and benzene– naphthalene complexes. Journal of Computational Chemistry, 28(2):540–546, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hofer:2005:SBE

[HRR05] Thomas S. Hofer, Bernhard R. Randolf, and Bernd M. Rode. Structure-breaking effects of solvated Rb(I) in dilute aqueous solution — an ab initio QM/MM MD approach. Journal of Computational Chemistry, 26(9):949–956, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Headley:2000:TAF

[HS00] Allan D. Headley and Stephen D. Starnes. Theoretical analysis of fluoroglycine conformers. Journal of Computational Chemistry, 21(6):426–431, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hofinger:2001:DRP

[HS01] Siegfried Höfinger and Thomas Simonson. Dielectric relaxation in proteins: a continuum electrostatics model incorporating dielectric heterogeneity of the protein and timedependent charges. Journal of Computational Chemistry, 22 (3):290–305, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hiberty:2007:SRD

[HS07a] Philippe C. Hiberty and Sason Shaik. A survey of recent developments in ab initio valence bond theory. Journal of Computational Chemistry, 28(1):137–151, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hu:2007:WSB

[HS07b] Zengjian Hu and William Southerland. Windock: Structurebased drug discovery on Windows-based PCs. Journal of Computational Chemistry, 28(14):2347–2351, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hegemann:2008:ASC

[HSF08] Ilka Hegemann, Andreas Schwaebe, and Karin Fink. Adsorption of single Cu atoms at differently stabilized polar ZnO surfaces: an ab initio study. Journal of Computational Chemistry, 29(13):2302–2310, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hemmateenejad:2004:AIT

[HSMT04] Bahram Hemmateenejad, Mohammad A. Safarpour, Ramin Miri, and Fariba Taghavi. Application of ab initio theory to QSAR study of 1,4-dihydropyridine-based calcium channel blockers using GA-MLR and PC-GA-ANN procedures. Journal of Computational Chemistry, 25(12):1495–1503, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Higo:2001:CMM

[HSWN01] Junichi Higo, Yasunobu Sugimoto, Katsuzo Wakabayashi, and Haruki Nakamura. Collective motions of myosin head derived from backbone molecular dynamics and combination with Xray solution scattering data. Journal of Computational Chemistry, 22(16):1983–1994, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hong:2000:CED

[HSWW00] Gongyi Hong, Marek Strajbl, Tomasz A. Wesolowski, and Arieh Warshel. Constraining the electron densities in DFT method as an effective way for ab initio studies of metalcatalyzed reactions. Journal of Computational Chemistry, 21 (16):1554–1561, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hopkins:2003:MAI

[HT03] Brian W. Hopkins and Gregory S. Tschumper. A multicentered approach to integrated QM/QM calculations. Applications to multiply hydrogen bonded systems. Journal of Computational Chemistry, 24(13):1563–1568, October 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Harman:2005:CDD

[HT05] W. Dean Harman and Carl Trindle. Charge donation to and dearomatization of benzene attending complexation: DFT estimates of binding energies of TpMXO(L) with benzene, for Tp = hydridotris(pyrazolyl) borate, MXO = MoNO, ReCO, and WNO, and L = ammonia, N-methylimidazole, pyridine, phosphine, methyl isocyanide, and carbon monoxide. Journal of Computational Chemistry, 26(2):194–200, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hnizdo:2008:ECC

[HTKG08] Vladimir Hnizdo, Jun Tan, Benjamin J. Killian, and Michael K. Gilson. Efficient calculation of configurational entropy from molecular simulations by combining the mutualinformation expansion and nearest-neighbor methods. Journal of Computational Chemistry, 29(10):1605–1614, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hori:2003:HQM

[HTN03] Takumi Hori, Hideaki Takahashi, and Tomoshige Nitta. Hybrid QM/MM molecular dynamics simulations for an ionic S_{N}2 reaction in the supercritical water: OH^{−} + CH_{3}Cl → CH_{3}OH + Cl^{−}. Journal of Computational Chemistry, 24(2): 209–221, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hofer:2004:CDR

[HTSR04] Thomas S. Hofer, Hung T. Tran, Christian F. Schwenk, and Bernd M. Rode. Characterization of dynamics and reactivities of solvated ions by ab initio simulations. Journal of Computational Chemistry, 25(2):211–217, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2009:ESO

[Hua09a] Wen Lai Huang. Electronic structures and optical properties of BiOX (X = F, Cl, Br, I) via DFT calculations. Journal of Computational Chemistry, 30(12):1882–1891, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2009:FPC

[Hua09b] Wen Lai Huang. First-principles calculations on the energetics, electronic structures and magnetism of SrFeO_{2}. Journal of Computational Chemistry, 30(16):2684–2693, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Horenko:2003:AIM

[HW03] Illia Horenko and Martin Weiser. Adaptive integration of molecular dynamics. Journal of Computational Chemistry, 24(15):1921–1929, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2009:CSP

[HW09] Zunnan Huang and Chung F. Wong. Conformational selection of protein kinase A revealed by flexible-ligand flexible-protein docking. Journal of Computational Chemistry, 30(4):631–644, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hageman:2003:PPI

[HWDB03] J. A. Hageman, R. Wehrens, R. De Gelder, and L. M. C. Buydens. Powder pattern indexing using the weighted crosscorrelation and genetic algorithms. Journal of Computational Chemistry, 24(9):1043–1051, July 15, 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hada:2001:QSN

[HWFN01] Masahiko Hada, Jian Wan, Ryoichi Fukuda, and Hiroshi Nakatsuji. Quasirelativistic study of ^{125}Te nuclear magnetic shielding constants and chemical shifts. Journal of Computational Chemistry, 22(13):1502–1508, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hrovat:2001:BCB

[HWGB01] David A. Hrovat, Richard Vaughan Williams, Alan C. Goren, and Weston Thatcher Borden. B3LYP calculations on bishomoaromaticity in substituted semibullvalenes*. Journal of Computational Chemistry, 22(13):1565–1573, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Harvey:2003:MNA

[HWTL03] Stephen C. Harvey, Chunlin Wang, Stephane Teletchea, and Richard Lavery. Motifs in nucleic acids: Molecular mechanics restraints for base pairing and base stacking. Journal of Computational Chemistry, 24(1):1–9, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

He:2008:SSL

[HXD08] Ning He, Hong-Bin Xie, and Yi-Hong Ding. Structures and stability of lithium monosilicide clusters SiLi_{n} (n = 4–16): What is the maximum number, magic number, and core number for lithium coordination to silicon? Journal of Computational Chemistry, 29(11):1850–1858, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

He:2009:EPS

[HXLS09] Yi He, Yi Xiao, Adam Liwo, and Harold A. Scheraga. Exploring the parameter space of the coarse-grained UNRES force field by random search: Selecting a transferable mediumresolution force field. Journal of Computational Chemistry, 30(13):2127–2135, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Helal:2002:ICS

[HYA02] Mustafa R. Helal, Yaser A. Yousef, and Akef T. Afaneh. Ab initio calculations of the stabilization energies of the conformational and the structural isomers of C_{3}H_{7}X where X = F, Cl, and Br. Journal of Computational Chemistry, 23(10):966–976, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yao:2008:ASD

[hYDN^{+}08] Yu hua Yao, Qi Dai, Xu-Ying Nan, Ping-An He, Zuo-Ming Nie, Song-Ping Zhou, and Yao-Zhou Zhang. Analysis of similarity/dissimilarity of DNA sequences based on a class of 2D graphical representation. Journal of Computational Chemistry, 29(10):1632–1639, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Herrmann:2006:SSP

[HYR06] Carmen Herrmann, Lian Yu, and Markus Reiher. Spin states in polynuclear clusters: The [Fe_{2}O_{2}] core of the methane monooxygenase active site. Journal of Computational Chemistry, 27(12):1223–1239, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hatano:2005:CMO

[HYT05] Yasuyo Hatano, Shigeyoshi Yamamoto, and Hiroshi Tatewaki. Characterization of molecular orbitals by counting nodal regions. Journal of Computational Chemistry, 26(4):325–333, March 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2006:IKBa

[HZ06a] Sheng-You Huang and Xiaoqin Zou. An iterative knowledgebased scoring function to predict protein–ligand interactions: I. Derivation of interaction potentials. Journal of Computational Chemistry, 27(15):1866–1875, November 30, 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2006:IKBb

[HZ06b] Sheng-You Huang and Xiaoqin Zou. An iterative knowledgebased scoring function to predict protein–ligand interactions: II. Validation of the scoring function. Journal of Computational Chemistry, 27(15):1876–1882, November 30, 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Huang:2009:DCE

[HZ09] Wen Lai Huang and Qingshan Zhu. DFT calculations on the electronic structures of BiOX (X = F, Cl, Br, I) photocatalysts with and without semicore Bi 5d states. Journal of Computational Chemistry, 30(2):183–190, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Han:2004:FPP

[HZX04] Xiaoping Han, Yue Zhang, and Huibin Xu. First-principles pair potentials across the metal–ceramic interface. Journal of Computational Chemistry, 25(7):968–973, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Improta:2004:ARD

[IB04] Roberto Improta and Vincenzo Barone. Assessing the reliability of density functional methods in the conformational study of polypeptides: the treatment of intraresidue nonbonding interactions. Journal of Computational Chemistry, 25(11): 1333–1341, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Iori:2008:IIC

[IC08] F. Iori and S. Corni. Including image charge effects in the molecular dynamics simulations of molecules on metal surfaces. Journal of Computational Chemistry, 29(10):1656–1666, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Iori:2009:GAF

[IDMC09] F. Iori, R. Di Felice, E. Molinari, and S. Corni. GolP: an atomistic force-field to describe the interaction of proteins with Au(111) surfaces in water. Journal of Computational Chemistry, 30(9):1465–1476, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Isayev:2007:TCC

[IGL07] Olexandr Isayev, Leonid Gorb, and Jerzy Leszczynski. Theoretical calculations: Can Gibbs free energy for intermolecular complexes be predicted efficiently and accurately? Journal of Computational Chemistry, 28(9):1598–1609, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ikeda:2003:HHI

[IGNH03] Kazuyoshi Ikeda, Oxana V. Galzitskaya, Haruki Nakamura, and Junichi Higo. β-hairpins, α-helices, and the intermediates among the secondary structures in the energy landscape of a peptide from a distal β-hairpin of SH3 domain. Journal of Computational Chemistry, 24(3):310–318, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ikegami:2002:SDC

[II02] Tsutomu Ikegami and Suehiro Iwata. Spectral density calculation by using the Chebyshev expansion. Journal of Computational Chemistry, 23(2):310–318, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishikawa:2009:TSP

[IIK09] Takeshi Ishikawa, Takakazu Ishikura, and Kazuo Kuwata. Theoretical study of the prion protein based on the fragment molecular orbital method. Journal of Computational Chemistry, 30(16):2594–2601, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Imai:2009:AIB

[IINK09] Yumi N. Imai, Yoshihisa Inoue, Isao Nakanishi, and Kazuo Kitaura. Amide-π interactions between formamide and benzene. Journal of Computational Chemistry, 30(14):2267–2276, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Izzo:2000:OCI

[IK00] Roberto Izzo and Martin Klessinger. Optimization of conical intersections using the semiempirical MNDOC–CI method with analytic gradients. Journal of Computational Chemistry, 21(1):52–62, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ikeguchi:2004:PRB

[Ike04] Mitsunori Ikeguchi. Partial rigid-body dynamics in NPT, NPAT and NPγT ensembles for proteins and membranes. Journal of Computational Chemistry, 25(4):529–541, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Imamura:2008:CST

[IKN08] Yutaka Imamura, Hiroyoshi Kiryu, and Hiromi Nakai. Colle– Salvetti-type correction for electron–nucleus correlation in the nuclear orbital plus molecular orbital theory. Journal of Computational Chemistry, 29(5):735–740, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Iwaoka:2009:SFF

[IKYM09] Michio Iwaoka, Naoki Kimura, Daisuke Yosida, and Toshiya Minezaki. The SAAP force field: Development of the single amino acid potentials for 20 proteinogenic amino acids and Monte Carlo molecular simulation for short peptides. Journal of Computational Chemistry, 30(13):2039–2055, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Im:2003:GBM

[ILB03] Wonpil Im, Michael S. Lee, and Charles L. Brooks III. Generalized Born model with a simple smoothing function. Journal of Computational Chemistry, 24(14):1691–1702, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Im:2009:NFE

[ILKR09] Wonpil Im, Jinhyuk Lee, Taehoon Kim, and Huan Rui. Novel free energy calculations to explore mechanisms and energetics of membrane protein structure and function. Journal of Computational Chemistry, 30(11):1622–1633, August 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Irback:2006:PMC

[IM06] Anders Irbäck and Sandipan Mohanty. PROFASI: a Monte Carlo simulation package for protein folding and aggregation. Journal of Computational Chemistry, 27(13):1548–1555, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ida:2002:ESS

[IME02] T. Ida, M. Mizuno, and K. Endo. Electronic state of small and large cavities for methane hydrate. Journal of Computational Chemistry, 23(11):1071–1075, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ito:2001:RPE

[IN01] Hiroyuki Ito and Hiroshi Nakatsuji. Roles of proteins in the electron transfer in the photosynthetic reaction center of Rhodopseudomonas viridis: bacteriopheophytin to ubiquinone. Journal of Computational Chemistry, 22(3):265– 272, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Imamura:2008:EDA

[IN08] Yutaka Imamura and Hiromi Nakai. Energy density analysis for second-order Møller–Plesset perturbation theory and coupled-cluster theory with singles and doubles: Application to C_{2}H_{4}CH_{4} complexes. Journal of Computational Chemistry, 29(10):1555–1563, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Inada:2008:ENB

[IO08] Yasuji Inada and Hideo Orita. Efficiency of numerical basis sets for predicting the binding energies of hydrogen bonded complexes: Evidence of small basis set superposition error compared to Gaussian basis sets. Journal of Computational Chemistry, 29(2):225–232, January 30, 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Imamura:2007:DCE

[ION07] Yutaka Imamura, Takao Otsuka, and Hiromi Nakai. Description of core excitations by time-dependent density functional theory with local density approximation, generalized gradient approximation, meta-generalized gradient approximation, and hybrid functionals. Journal of Computational Chemistry, 28 (12):2067–2074, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishimura:2006:NPA

[IPN06] Kazuya Ishimura, Peter Pulay, and Shigeru Nagase. A new parallel algorithm of MP2 energy calculations. Journal of Computational Chemistry, 27(4):407–413, March 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishimura:2007:NPA

[IPN07] Kazuya Ishimura, Peter Pulay, and Shigeru Nagase. New parallel algorithm for MP2 energy gradient calculations. Journal of Computational Chemistry, 28(12):2034–2042, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ivanic:2003:MMG

[IR03] Joseph Ivanic and Klaus Ruedenberg. A MCSCF method for ground and excited states based on full optimizations of successive Jacobi rotations. Journal of Computational Chemistry, 24(10):1250–1262, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishida:2003:LIS

[IS03] Toshimasa Ishida and George C. Schatz. A local interpolation scheme using no derivatives in potential sampling: Application to O(^{1}D) + H_{2} system. Journal of Computational Chemistry, 24(9):1077–1086, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Inaba:2007:DPD

[IS07] Toru Inaba and Fumitoshi Sato. Development of parallel density functional program using distributed matrix to calculate all-electron canonical wavefunction of large molecules. Journal of Computational Chemistry, 28(5):984–995, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishida:2002:ACE

[Ish02] Kazuhiro Ishida. Accompanying coordinate expansion formulas derived with the solid harmonic gradient. Journal of Computational Chemistry, 23(3):378–393, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishida:2003:MIG

[Ish03] Kazuhiro Ishida. Molecular integrals over the gauge-including atomic orbitals. II. The Breit–Pauli interaction. Journal of Computational Chemistry, 24(15):1874–1890, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ishida:2004:AFA

[Ish04] Kazuhiro Ishida. Accurate and fast algorithm of the molecular incomplete gamma function with a complex argument. Journal of Computational Chemistry, 25(5):739–748, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Iwaoka:2003:SFF

[IT03] Michio Iwaoka and Shuji Tomoda. The SAAP force field. A simple approach to a new all-atom protein force field by using single amino acid potential (SAAP) functions in various solvents. Journal of Computational Chemistry, 24(10):1192– 1200, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Inaba:2005:AED

[ITN^{+}05] Toru Inaba, Saisei Tahara, Nobutaka Nisikawa, Hiroshi Kashiwagi, and Fumitoshi Sato. All-electron density functional calculation on insulin with quasi-canonical localized orbitals. Journal of Computational Chemistry, 26(10):987–993, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Intharathep:2005:SDH

[ITS05] Pathumwadee Intharathep, Anan Tongraar, and Kritsana Sagarik. Structure and dynamics of hydrated NH: an ab initio QM/MM molecular dynamics simulation. Journal of Computational Chemistry, 26(13):1329–1338, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Intharathep:2006:IQM

[ITS06] Pathumwadee Intharathep, Anan Tongraar, and Kritsana Sagarik. Ab initio QM/MM dynamics of H_{3}O^{+} in water. Journal of Computational Chemistry, 27(14):1723–1732, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Infante:2004:QMS

[IV04] Ivan Infante and Lucas Visscher. QM/MM study of aqueous solvation of the uranyl fluoride [UO_{2}F] complex. Journal of Computational Chemistry, 25(3):386–392, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Infante:2006:QMS

[IvSV06] Ivan Infante, Bas van Stralen, and Lucas Visscher. A QM/MM study on the aqueous solvation of the tetrahydroxouranylate [UO_{2} (OH)_{4}]^{2−} complex ion. Journal of Computational Chemistry, 27(11):1156–1162, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Izrailev:2006:DGH

[IZA06] Sergei Izrailev, Fangqiang Zhu, and Dimitris K. Agrafiotis. A distance geometry heuristic for expanding the range of geometries sampled during conformational search. Journal of Computational Chemistry, 27(16):1962–1969, December 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jacobsen:2009:CBV

[Jac09] Heiko Jacobsen. Chemical bonding in view of electron charge density and kinetic energy density descriptors. Journal of Computational Chemistry, 30(7):1093–1102, May 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jiang:2002:CPE

[JARM02] He Jiang, Dominic Appadoo, Evan Robertson, and Don Mc-Naughton. A comparison of predicted and experimental vibrational spectra in some small fluorocarbons. Journal of Computational Chemistry, 23(13):1220–1225, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jug:2004:MTC

[JB04] Karl Jug and Thomas Bredow. Models for the treatment of crystalline solids and surfaces. Journal of Computational Chemistry, 25(13):1551–1567, October 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Janetzko:2008:BDD

[JBGK08] Florian Janetzko, Thomas Bredow, Gerald Geudtner, and Andreas M. Köster. Boron-doped diamond: Investigation of the stability of surface-doping versus bulk-doping using cyclic cluster model calculations. Journal of Computational Chemistry, 29(13):2295–2301, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jakalian:2000:FEG

[JBJB00] Araz Jakalian, Bruce L. Bush, David B. Jack, and Christopher I. Bayly. Fast, efficient generation of high-quality atomic charges. AM1-BCC model: I. Method. Journal of Computational Chemistry, 21(2):132–146, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jacquemin:2002:IAD

[JCA^{+}02] Denis Jacquemin, Benoît Champagne, Jean-Marie André, Erik Deumens, and Yngve öhrn. Integral algorithm and density matrix integration scheme for ab initio band structure calculations on polymeric systems. Journal of Computational Chemistry, 23(15):1430–1444, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jurecka:2007:DFT

[JČHS07] Petr Jurečka, Jiří Černý, Pavel Hobza, and Dennis R. Salahub. Density functional theory augmented with an empirical dispersion term. Interaction energies and geometries of 80 noncovalent complexes compared with ab initio quantum mechanics calculations. Journal of Computational Chemistry, 28(2):555– 569, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jha:2005:FFV

[JCL05] Shantenu Jha, Peter V. Coveney, and Charles A. Laughton. Force field validation for nucleic acid simulations: Comparing energies and dynamics of a DNA dodecamer. Journal of Computational Chemistry, 26(15):1617–1627, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jin:2009:SBS

[JD09] Lin Jin and Yi-Hong Ding. Stability of N-based sandwich-like energetic complexes [N_{4}TiN_{4}]^{2−}: Effect of spins and counterions. Journal of Computational Chemistry, 30(8):1279–1289, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jin:2006:RMC

[JDWS06] Lin Jin, Yi-Hong Ding, Jian Wang, and Chia-Chung Sun. Reaction mechanism of the CCN radical with nitric oxide. Journal of Computational Chemistry, 27(7):883–893, May 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jung:2004:CMF

[JFG04] Dawoon Jung, Jie Floyd, and Tamara M. Gund. A comparative molecular field analysis (CoMFA) study using semiempirical, density functional, ab initio methods and pharmacophore derivation using DISCOtech on sigma 1 ligands. Journal of Computational Chemistry, 25(11):1385–1399, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jug:2003:BEM

[JG03] Karl Jug and Gerald Geudtner. Bond energies for molecules, clusters, and deposit systems. Journal of Computational Chemistry, 24(16):2013–2022, December 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jug:2000:MPT

[JGH00] Karl Jug, Gerald Geudtner, and Thorsten Homann. MSINDO parameterization for third-row main group elements. Journal of Computational Chemistry, 21(11):974–987, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Johnston:2005:FBD

[JGVF05] Michael A. Johnston, Ignacio Fdez. Galván, and Jordi Villà-Freixa. Framework-based design of a new all-purpose molecular simulation application: the Adun simulator. Journal of Computational Chemistry, 26(15):1647–1659, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jimenez-Halla:2009:TAT

[JHMB^{+}09] J. Oscar C. Jiménez-Halla, Eduard Matito, Lluís Blancafort, Juvencio Robles, and Miquel Solà. Tuning aromaticity in trigonal alkaline earth metal clusters and their alkali metal salts. Journal of Computational Chemistry, 30(16):2764–2776, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [JHMB^{+}11].

Jimenez-Halla:2011:ETA

[JHMB^{+}11] J. Oscar C. Jiménez-Halla, Eduard Matito, Lluís Blancafort, Juvencio Robles, and Miquel Solà. Erratum: Tuning aromaticity in trigonal alkaline earth metal clusters and their alkali metal salts. Journal of Computational Chemistry, 32(2):372– 373, January 30, 2011. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [JHMB^{+}09].

Junquera-Hernandez:2005:MAS

[JHPRSM^{+}05] J. M. Junquera-Hernández, J. Pitarch-Ruiz, J. Sánchez-Marín, S. Evangelisti, and D. Maynau. Multistate active spaces from local CAS-SCF molecular orbitals: the photodissociation of HFCO as an example. Journal of Computational Chemistry, 26(12):1254–1262, September 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ju:2009:GDT

[JHZ09] Li-Ping Ju, Ke-Li Han, and John Z. H. Zhang. Global dynamics and transition state theories: Comparative study of reaction rate constants for gas-phase chemical reactions. Journal of Computational Chemistry, 30(2):305–316, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jiang:2009:PPF

[JIK09] Yingfu Jiang, Paul Iglinski, and Lukasz Kurgan. Prediction of protein folding rates from primary sequences using hybrid sequence representation. Journal of Computational Chemistry, 30(5):772–783, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jakalian:2002:FEG

[JJB02] Araz Jakalian, David B. Jack, and Christopher I. Bayly. Fast, efficient generation of high-quality atomic charges. AM1-BCC model: II. Parameterization and validation. Journal of Computational Chemistry, 23(16):1623–1641, December 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jongejan:2001:DHT

[JJH01] A. Jongejan, J. A. Jongejan, and W. R. Hagen. Direct hydride transfer in the reaction mechanism of quinoprotein alcohol dehydrogenases: a quantum mechanical investigation. Journal of Computational Chemistry, 22(15):1732–1749, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jenkins:2000:BSC

[JJK^{+}00] H. Donald B. Jenkins, Luminita C. Jitariu, Ingo Krossing, Jack Passmore, and Reijo Suontamo. Basis set and correlation effects in the calculation of accurate gas phase dimerization energies of two M to give M (M = S, Se). Journal of Computational Chemistry, 21(3):218–226, February 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jo:2008:CGW

[JKII08] Sunhwan Jo, Taehoon Kim, Vidyashankara G. Iyer, and Wonpil Im. CHARMM-GUI: a Web-based graphical user interface for CHARMM. Journal of Computational Chemistry, 29(11): 1859–1865, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Johansson:2008:CPM

[JKL08] Mikael P. Johansson, Ville R. I. Kaila, and Liisa Laakkonen. Charge parameterization of the metal centers in cytochrome c oxidase. Journal of Computational Chemistry, 29(5):753–767, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jena:2008:RHA

[JKM08] N. R. Jena, P. S. Kushwaha, and P. C. Mishra. Reaction of hypochlorous acid with imidazole: Formation of 2-chloro- and 2-oxoimidazoles. Journal of Computational Chemistry, 29(1):

98–107, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jiang:2003:NTI

[JLHF03] Yu-Ren Jiang, Jing-Ya Liu, Yue-Hua Hu, and Toyohisa Fujita. Novel topological index for research on structure-property relationships of complex organic compounds. Journal of Computational Chemistry, 24(7):842–849, May 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jena:2007:FNO

[JM07a] N. R. Jena and P. C. Mishra. Formation of 8-nitroguanine and 8-oxoguanine due to reactions of peroxynitrite with guanine. Journal of Computational Chemistry, 28(8):1321–1335, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jojart:2007:PGA

[JM07b] Balázs Jójárt and Tamás A. Martinek. Performance of the general AMBER force field in modeling aqueous POPC membrane bilayers. Journal of Computational Chemistry, 28(12): 2051–2058, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jayaram:2002:FEC

[JMD^{+}02] B. Jayaram, K. McConnell, S. B. Dixit, A. Das, and D. L. Beveridge. Free-energy component analysis of 40 protein–DNA complexes: a consensus view on the thermodynamics of binding at the molecular level. Journal of Computational Chemistry, 23(1):1–14, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jacob:2008:FIF

[JNV08] Christoph R. Jacob, Johannes Neugebauer, and Lucas Visscher. A flexible implementation of frozen-density embedding for use in multilevel simulations. Journal of Computational Chemistry, 29(6):1011–1018, April 30, 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jaqaman:2002:NSW

[JO02] Khuloud Jaqaman and Peter J. Ortoleva. New space warping method for the simulation of large-scale macromolecular conformational changes. Journal of Computational Chemistry, 23 (4):484–491, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jezierska:2009:IOH

[JP09] Aneta Jezierska and Jarosław J. Panek. Investigations of an OH⋯S hydrogen bond via Car–Parrinello and path integral molecular dynamics. Journal of Computational Chemistry, 30 (8):1241–1250, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jacquemin:2008:RRB

[JPCA08] Denis Jacquemin, Eric A. Perpéte, Ilaria Ciofini, and Carlo Adamo. Revisiting the relationship between the bond length alternation and the first hyperpolarizability with rangeseparated hybrid functionals. Journal of Computational Chemistry, 29(6):921–925, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jakli:2000:PMX

[JPF^{+}00] Imre Jákli, András Perczel, Ödön Farkas, Attila G. Császár, Carlos Sosa, and Imre G. Csizmadia. Peptide models XXIII. Conformational model for polar side-chain containing amino acid residues: a comprehensive analysis of RHF, DFT, and MP2 properties of HCO-L-SER-NH_{2}. Journal of Computational Chemistry, 21(8):626–655, June 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jemmis:2001:CCC

[JRJ01] Eluvathingal D. Jemmis, M. Ramalingam, and Elambalassery G. Jayasree. Cationic closo carboranes — promising weakly coordinating ions. Journal of Computational Chemistry, 22(13):1542–1551, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jagielska:2007:ITF

[JS07a] Anna Jagielska and Harold A. Scheraga. Influence of temperature, friction, and random forces on folding of the B-domain of staphylococcal protein A: All-atom molecular dynamics in implicit solvent. Journal of Computational Chemistry, 28(6): 1068–1082, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jagielska:2007:OIH

[JS07b] Anna Jagielska and Jeffrey Skolnick. Origin of intrinsic 3_{10}- helix versus strand stability in homopolypeptides and its implications for the accuracy of the AMBER force field. Journal of Computational Chemistry, 28(10):1648–1657, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jung:2007:FES

[JSHG07] Yousung Jung, Yihan Shao, and Martin Head-Gordon. Fast evaluation of scaled opposite spin second-order Møller–Plesset correlation energies using auxiliary basis expansions and exploiting sparsity. Journal of Computational Chemistry, 28(12): 1953–1964, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jug:2006:ASZ

[JT06] Karl Jug and Viatcheslav A. Tikhomirov. Anion substitution in zinc chalcogenides. Journal of Computational Chemistry, 27(10):1088–1092, July 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jug:2008:IID

[JT08] Karl Jug and Viatcheslav A. Tikhomirov. Influence of intrinsic defects on the properties of zinc oxide. Journal of Computational Chemistry, 29(13):2250–2254, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jorgensen:2005:MMO

[JTR05] William L. Jorgensen and Julian Tirado-Rives. Molecular modeling of organic and biomolecular systems using BOSS and MCPRO. Journal of Computational Chemistry, 26(16): 1689–1700, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jirouskova:2009:SNU

[JVVK09] Zuzana Jiroušková, Radka Svobodová Vařeková, Jakub Vanĕk, and Jaroslav Koča. Software news and updates electronegativity equalization method: Parameterization and validation for organic molecules using the Merz-Kollman-Singh charge distribution scheme. Journal of Computational Chemistry, 30(7):1174–1178, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jug:2000:MSL

[JW00] Karl Jug and Daniel Wichmann. MSINDO study of large silsesquioxanes. Journal of Computational Chemistry, 21(16): 1549–1553, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jang:2006:MPC

[JW06] Hyunbum Jang and Thomas B. Woolf. Multiple pathways in conformational transitions of the alanine dipeptide: an application of dynamic importance sampling. Journal of Computational Chemistry, 27(11):1136–1141, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jorgensen:2012:LEC

[JW12] Kameron R. Jorgensen and Angela K. Wilson. Letters to the editor: Comment on the paper “Extensive Theoretical Studies of a New Energetic Material: Tetrazino-Tetrazine-Tetraoxide (TTTO)” by Xinli Song, Jicun Li, Hua Hou, and Baoshan Wang. Journal of Computational Chemistry, 33(24):1967– 1968, September 15, 2012. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [SLHW09].

Ju:2005:SEH

[JWB05] Xue-Hai Ju, Xin Wang, and Feng-Li Bei. Substituent effects on heats of formation, group interactions, and detonation properties of polyazidocubanes. Journal of Computational Chemistry, 26(12):1263–1269, September 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jiao:2009:TLB

[JZD^{+}09] Dian Jiao, Jiajing Zhang, Robert E. Duke, Guohui Li, Michael J. Schnieders, and Pengyu Ren. Trypsin-ligand binding free energies from explicit and implicit solvent simulations with polarizable potential. Journal of Computational Chemistry, 30(11):1701–1711, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kikugawa:2009:AMS

[KAK^{+}09] Gota Kikugawa, Rossen Apostolov, Narutoshi Kamiya, Makoto Taiji, Ryutaro Himeno, Haruki Nakamura, and Yasushige Yonezawa. Application of MDGRAPE-3, a special purpose board for molecular dynamics simulations, to periodic biomolecular systems. Journal of Computational Chemistry, 30(1):110–118, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kang:2007:TSF

[Kan07] Hong Seok Kang. A theoretical study of fullerene–ferrocene hybrids. Journal of Computational Chemistry, 28(2):594–600, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Karafiloglou:2001:MCW

[Kar01] Padeleimon Karafiloglou. A method to calculate the weights of nbo electronic structures from Moffitt’s theorem. Journal of Computational Chemistry, 22(3):306–315, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Karafiloglou:2006:CDS

[Kar06] Padeleimon Karafiloglou. Control of delocalization and structural changes by means of an electric field. Journal of Computational Chemistry, 27(15):1883–1891, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kasalova:2007:MST

[KAS^{+}07] Veronika Kasalová, Wesley D. Allen, Henry F. Schaefer III, Eszter Czinki, and Attila G. Császár. Molecular structures of the two most stable conformers of free glycine. Journal of Computational Chemistry, 28(8):1373–1383, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kaupp:2007:RRN

[Kau07] Martin Kaupp. The role of radial nodes of atomic orbitals for chemical bonding and the periodic table. Journal of Computational Chemistry, 28(1):320–325, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kahn:2002:POA

[KB02] Kalju Kahn and Thomas C. Bruice. Parameterization of OPLS–AA force field for the conformational analysis of macrocyclic polyketides. Journal of Computational Chemistry, 23 (10):977–996, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Knight:2009:DFE

[KB09] Jennifer L. Knight and Charles L. Brooks III. λ-dynamics free energy simulation methods. Journal of Computational Chemistry, 30(11):1692–1700, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kenny:2004:CBI

[KBA^{+}04] Joseph P. Kenny, Steven J. Benson, Yuri Alexeev, Jason Sarich, Curtis L. Janssen, Lois Curfman Mcinnes, Manojkumar Krishnan, Jarek Nieplocha, Elizabeth Jurrus, Carl Fahlstrom, and Theresa L. Windus. Component-based integration of chemistry and optimization software. Journal of Computational Chemistry, 25(14):1717–1725, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Konig:2009:UUB

[KBB09] Gerhard König, Stefan Bruckner, and Stefan Boresch. Unorthodox uses of Bennett’s acceptance ratio method. Journal of Computational Chemistry, 30(11):1712–1718, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kirchhoff:2001:ASB

[KBK^{+}01] Paul D. Kirchhoff, Rob Brown, Scott Kahn, Marvin Waldman, and C. M. Venkatachalam. Application of structurebased focusing to the estrogen receptor. Journal of Computational Chemistry, 22(10):993–1003, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kisowska:2008:SCB

[KBL08] Karolina Kisowska, Slawomir Berski, and Zdzislaw Latajka. The structure and chemical bonding in the N_{2}CuX and N_{2}⋯XCu (X = F, Cl, Br) systems studied by means of the molecular orbital and Quantum Chemical Topology methods. Journal of Computational Chemistry, 29(16):2677–2692, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kover:2009:CNT

[KBLP09] Katalin E. Kövér, Tamás Beke, András Lipták, and András Perczel. Combined NMR three-bond scalar coupling measurements and QM calculations to calculate OH-rotamer equilibrium of polyalcohols. Journal of Computational Chemistry, 30 (4):540–550, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kuttel:2002:CSS

[KBN02] Michelle Kuttel, J. W. Brady, and Kevin J. Naidoo. Carbohydrate solution simulations: Producing a force field with experimentally consistent primary alcohol rotational frequencies and populations. Journal of Computational Chemistry, 23(13):1236–1243, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Koslowski:2003:IGM

[KBT03] Axel Koslowski, Michael E. Beck, and Walter Thiel. Implementation of a general multireference configuration interaction procedure with analytic gradients in a semiempirical context using the graphical unitary group approach. Journal of Computational Chemistry, 24(6):714–726, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kraka:2001:PDP

[KC01a] Elfi Kraka and Dieter Cremer. The para-didehydropyridine, para-didehydropyridinium, and related biradicals — a contribution to the chemistry of enediyne antitumor drugs. Journal of Computational Chemistry, 22(2):216–229, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kristyan:2001:FAC

[KC01b] Sándor Kristyán and Gábor I. Csonka. Fitting atomic correlation parameters for RECEP (rapid estimation of correlation energy from partial charges) method to estimate molecular correlation energies within chemical accuracy. Journal of Computational Chemistry, 22(2):241–254, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2008:PDS

[KCK^{+}08] Chan Kyung Kim, Soo Gyeong Cho, Chang Kon Kim, Hyung-Yeon Park, Hui Zhang, and Hai Whang Lee. Prediction of densities for solid energetic molecules with molecular surface electrostatic potentials. Journal of Computational Chemistry, 29(11):1818–1824, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kirtman:2000:ETE

[KCL00] Bernard Kirtman, Benoit Champagne, and Josep M. Luis. Efficient treatment of the effect of vibrations on electrical, magnetic, and spectroscopic properties. Journal of Computational Chemistry, 21(16):1572–1588, December 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ke:2006:SVC

[KCL06] Yi-Yu Ke, Yun-Chu Chen, and Thy-Hou Lin. Structure of the virus capsid protein VP1 of enterovirus 71 predicted by some homology modeling and molecular docking studies. Journal of Computational Chemistry, 27(13):1556–1570, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kang:2009:LAS

[KDG^{+}09] K. S. Kang, J. W. Davenport, J. Glimm, D. E. Keyes, and M. McGuigan. Linear augmented Slater-type orbital method for free standing clusters. Journal of Computational Chemistry, 30(8):1185–1193, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kony:2002:IOA

[KDSV02] D. Kony, W. Damm, S. Stoll, and W. F. Van Gunsteren. An improved OPLS–AA force field for carbohydrates. Journal of Computational Chemistry, 23(15):1416–1429, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Keil:2004:PRS

[KEB04] Matthias Keil, Thomas E. Exner, and Jürgen Brickmann. Pattern recognition strategies for molecular surfaces: III. Binding site prediction with a neural network. Journal of Computational Chemistry, 25(6):779–789, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klamt:2002:PAS

[KEH^{+}02] Andreas Klamt, Frank Eckert, Martin Hornig, Michael E. Beck, and Thorsten Bürger. Prediction of aqueous solubility of drugs and pesticides with COSMO-RS. Journal of Computational Chemistry, 23(2):275–281, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kovalyov:2008:SSP

[KEM08] Evgenii V. Kovalyov, Vladimir I. Elokhin, and Aleksandr V. Myshlyavtsev. Stochastic simulation of physicochemical processes performance over supported metal nanoparticles. Journal of Computational Chemistry, 29(1):79–86, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klepeis:2002:IPH

[KF02a] J. L. Klepeis and C. A. Floudas. Ab initio prediction of helical segments in polypeptides. Journal of Computational Chemistry, 23(2):245–266, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ko:2002:RCL

[KF02b] Gwon Hee Ko and William H. Fink. Rapidly converging lattice sums for nonelectrostatic interactions. Journal of Computational Chemistry, 23(4):477–483, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klepeis:2003:PST

[KF03] J. L. Klepeis and C. A. Floudas. Prediction of β-sheet topology and disulfide bridges in polypeptides. Journal of Computational Chemistry, 24(2):191–208, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kruger:2008:ECS

[KF08] Jens Krüger and Wolfgang B. Fischer. Exploring the conformational space of Vpu from HIV-1: a versatile adaptable protein. Journal of Computational Chemistry, 29(14):2416– 2424, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Karney:2005:MCP

[KFB05] Charles F. F. Karney, Jason E. Ferrara, and Stephan Brunner. Method for computing protein binding affinity. Journal of Computational Chemistry, 26(3):243–251, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Knippenberg:2006:GFS

[KFD06] S. Knippenberg, J.-P. François, and M. S. Deleuze. Green’s function study of the one-electron and shake-up ionization spectra of unsaturated hydrocarbon cage compounds. Journal of Computational Chemistry, 27(14):1703–1722, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Katagiri:2008:IPS

[KFNH08] Daisuke Katagiri, Hideyoshi Fuji, Saburo Neya, and Tyuji Hoshino. Ab initio protein structure prediction with force field parameters derived from water-phase quantum chemical calculation. Journal of Computational Chemistry, 29(12):1930– 1944, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kaminski:2003:CIM

[KFZ03] George A. Kaminski, Richard A. Friesner, and Ruhong Zhou. A computationally inexpensive modification of the point dipole electrostatic polarization model for molecular simulations. Journal of Computational Chemistry, 24(3):267–276, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kairys:2002:EDM

[KG02] Visvaldas Kairys and Michael K. Gilson. Enhanced docking with the mining minima optimizer: Acceleration and sidechain flexibility. Journal of Computational Chemistry, 23(16): 1656–1670, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kurtulus:2006:ESC

[KGD06] Yasemin Kurtulus, Michael Gilleßen, and Richard Dronskowski. Electronic structure, chemical bonding, and finitetemperature magnetic properties of full Heusler alloys. Journal of Computational Chemistry, 27(1):90–102, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kadossov:2007:ESP

[KGL07] Evgueni B. Kadossov, Karen J. Gaskell, and Marjorie A. Langell. Effect of surrounding point charges on the density functional calculations of Ni_{x}O_{x} clusters (x = 4–12). Journal of Computational Chemistry, 28(7):1240–1251, May 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kahn:2007:CTO

[KGN07] Kalju Kahn, Alex A. Granovsky, and Jozef Noga. Convergence of third order correlation energy in atoms and molecules. Journal of Computational Chemistry, 28(2):547–554, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kamiya:2001:RAS

[KH01] Narutoshi Kamiya and Junichi Higo. Repeated-annealing sampling combined with multicanonical algorithm for conformational sampling of bio-molecules. Journal of Computational Chemistry, 22(10):1098–1106, July 30, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Koskowski:2005:TPF

[KH05] Florian Koskowski and Bernd Hartke. Towards protein folding with evolutionary techniques. Journal of Computational Chemistry, 26(11):1169–1179, August 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Krautler:2006:MTS

[KH06] Vincent Kräutler and Philippe H. Hünenberger. A multiple time step algorithm compatible with a large number of distance classes and an arbitrary distance dependence of the time step size for the fast evaluation of nonbonded interactions in molecular simulations. Journal of Computational Chemistry, 27(11):1163–1176, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kiyota:2009:MQM

[KHF^{+}09] Yasuomi Kiyota, Jun-Ya Hasegawa, Kazuhiro Fujimoto, Ben Swerts, and Hiroshi Nakatsuji. A multicore QM/MM approach for the geometry optimization of chromophore aggregate in protein. Journal of Computational Chemistry, 30(8):1351– 1359, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Khandogin:2000:ESP

[KHY00] Jana Khandogin, Anguang Hu, and Darrin M. York. Electronic structure properties of solvated biomolecules: a quantum approach for macromolecular characterization. Journal of Computational Chemistry, 21(16):1562–1571, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Komeiji:2007:CPE

[KIFK07] Yuto Komeiji, Toyokazu Ishida, Dmitri G. Fedorov, and Kazuo Kitaura. Change in a protein’s electronic structure induced by an explicit solvent: an ab initio fragment molecular orbital study of ubiquitin. Journal of Computational Chemistry, 28 (10):1750–1762, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Komeiji:2009:FMO

[KIM^{+}09] Yuto Komeiji, Takeshi Ishikawa, Yuji Mochizuki, Hiroshi Yamataka, and Tatsuya Nakano. Fragment Molecular Orbital method-based Molecular Dynamics (FMO–MD) as a simulator for chemical reactions in explicit solvation. Journal of Computational Chemistry, 30(1):40–50, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2007:CVM

[KJP^{+}07] Kyoung Hoon Kim, Jaehoon Jung, Bo Keun Park, Young-Kyu Han, and Joon T. Park. Cyclic voltammetry modeling, geometries, and electronic properties for metallofullerene complexes with μ_{3}–η^{2}: η^{2}: η^{2}-C_{60} bonding mode. Journal of Computational Chemistry, 28(6):1100–1106, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kenny:2008:CIE

[KJVW08] Joseph P. Kenny, Curtis L. Janssen, Edward F. Valeev, and Theresa L. Windus. Components for integral evaluation in quantum chemistry. Journal of Computational Chemistry, 29 (4):562–577, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kolossvary:2001:HFL

[KK01a] István Kolossváry and György M. Keserü. Hessian-free lowmode conformational search for large-scale protein loop optimization: application to c-jun N-terminal kinase JNK3. Journal of Computational Chemistry, 22(1):21–30, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kuznetsova:2001:VAP

[KK01b] Tatyana Kuznetsova and Bjørn Kvamme. Viabilty of atomistic potentials for thermodynamic properties of carbon dioxide at low temperatures. Journal of Computational Chemistry, 22 (15):1772–1781, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kahn:2008:IEF

[KK08a] Kalju Kahn and Iiris Kahn. Improved efficiency of focal point conformational analysis with truncated correlation consistent basis sets. Journal of Computational Chemistry, 29(6):900– 911, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kieseritzky:2008:IPC

[KK08b] Gernot Kieseritzky and E. W. Knapp. Improved pK_{a} prediction: Combining empirical and semimicroscopic methods. Journal of Computational Chemistry, 29(15):2575–2581, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2008:CNG

[KK08c] Woo Youn Kim and Kwang S. Kim. Carbon nanotube, graphene, nanowire, and molecule-based electron and spin transport phenomena using the nonequilibrium Green’s function method at the level of first principles theory. Journal of Computational Chemistry, 29(7):1073–1083, May 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kang:2009:CPM

[KK09] Young Kee Kang and Nam Sook Kang. Conformational preferences of N-methoxycarbonyl proline dipeptide. Journal of Computational Chemistry, 30(7):1116–1127, May 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Khan:2005:SPM

[KKC05] Malek O. Khan, Gareth Kennedy, and Derek Y. C. Chan. A scalable parallel Monte Carlo method for free energy simulations of molecular systems. Journal of Computational Chemistry, 26(1):72–77, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kavathekar:2009:WWI

[KKG^{+}09] Ritwik Kavathekar, Subodh Khire, V. Ganesh, Anuja P. Rahalkar, and Shridhar R. Gadre. WebMTA: a web-interface for ab initio geometry optimization of large molecules using molecular tailoring approach. Journal of Computational Chemistry, 30(7):1167–1173, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2007:EEG

[KKH^{+}07] Chang K. Kim, Dong J. Kim, Y. Hsieh, Hai W. Lee, Bon-Su Lee, and Chan K. Kim. Effects of entropy on the gas-phase pyrolysis of ethyl N, N-dimethylcarbamate. Journal of Computational Chemistry, 28(3):625–631, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2008:RMD

[KKJH08] Jong Chan Kim, Kyoung Hoon Kim, Jaehoon Jung, and Young-Kyu Han. Reaction mechanisms of dissociative chemisorption of HI, I_{2}, and CH_{3}I on a magic cluster Al. Journal of Computational Chemistry, 29(10):1626–1631, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kumar:2004:TSS

[KKMMS04] Anil Kumar, Michaela Knapp-Mohammady, P. C. Mishra, and Sándor Suhai. A theoretical study of structures and electron affinities of radical anions of guanine-cytosine, adeninethymine, and hypoxanthine-cytosine base pairs. Journal of Computational Chemistry, 25(8):1047–1059, June 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kuhn:2004:PPS

[KKS04] Bernd Kuhn, Peter A. Kollman, and Martin Stahl. Prediction of pK_{a} shifts in proteins using a combination of molecular mechanical and continuum solvent calculations. Journal of Computational Chemistry, 25(15):1865–1872, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kamishima:2001:CSG

[KKY01] Mayumi Kamishima, Masaaki Kojima, and Yuzo Yoshikawa. Calculation of stacking geometries by a modified MM2 method. Journal of Computational Chemistry, 22(8):835–845, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kirschner:2003:MMF

[KLB03] K. N. Kirschner, A. H. Lewin, and J. P. Bowen. Molecular mechanics force-field development for amino acid zwitterions. Journal of Computational Chemistry, 24(1):111–128, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klein:2002:ICS

[Kle02] Roger A. Klein. Ab initio conformational studies on diols and binary diol-water systems using DFT methods. Intramolecular hydrogen bonding and 1:1 complex formation with water. Journal of Computational Chemistry, 23(6):585–599, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klein:2003:HBD

[Kle03] Roger A. Klein. Hydrogen bonding in diols and binary diol– water systems investigated using DFT methods. II. Calculated infrared OH-stretch frequencies, force constants, and NMR chemical shifts correlate with hydrogen bond geometry and electron density topology. A reevaluation of geometrical criteria for hydrogen bonding. Journal of Computational Chemistry, 24(9):1120–1131, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2004:PPP

[KLH^{+}04] Chan Kyung Kim, Kyung A. Lee, Kwan Hoon Hyun, Heung Jin Park, In Young Kwack, Chang Kon Kim, Hai Whang Lee, and Bon-Su Lee. Prediction of physicochemical properties of organic molecules using van der Waals surface electrostatic potentials. Journal of Computational Chemistry, 25 (16):2073–2079, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kliesch:2001:EEP

[Kli01] Wolfgang Kliesch. EQUIPATH — an equilibrial path tracing routine for the use with the program package GAUS-SIAN94. Journal of Computational Chemistry, 22(15):1801– 1816, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Korchowiec:2009:ECT

[KLM^{+}09] Jacek Korchowiec, Jakub Lewandowski, Marcin Makowski, Feng Long Gu, and Yuriko Aoki. Elongation cutoff technique armed with quantum fast multipole method for linear scaling. Journal of Computational Chemistry, 30(15):2515–2525, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kazmierkiewicz:2002:EBR

[KLS02] Rajmund Kaźmierkiewicz, Adam Liwo, and Harold A. Scheraga. Energy-based reconstruction of a protein backbone from its α-carbon trace by a Monte-Carlo method. Journal of Computational Chemistry, 23(7):715–723, May 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kawata:2000:CEC

[KM00] Masaaki Kawata and Masuhiro Mikami. Computationally efficient canonical molecular dynamics simulations by using a multiple time-step integrator algorithm combined with the particle mesh Ewald method and with the fast multipole method. Journal of Computational Chemistry, 21(3):201–217, February 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kegerreis:2007:OMC

[KM07] Jeb Kegerreis and Nancy Makri. Optimized Monte Carlo sampling in forward–backward semiclassical dynamics. Journal of Computational Chemistry, 28(4):818–824, March 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kent:2007:EAF

[KMA^{+}07] David R. Kent IV, Richard P. Muller, Amos G. Anderson, William A. Goddard III, and Michael T. Feldmann. Efficient algorithm for “on-the-fly” error analysis of local or distributed serially correlated data. Journal of Computational Chemistry, 28(14):2309–2316, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kalat:2002:QPG

[KMH02] Sam Kalat, Geoff Mann, and Jan Hermans. Qmd-plot: a graphical utility for rapid preliminary analysis of time series of fluctuating data, developed in the context of molecular dynamics simulations. Journal of Computational Chemistry, 23 (1):184–188, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Khoroshun:2007:ERO

[KMM07] Dmitry V. Khoroshun, Djamaladdin G. Musaev, and Keiji Morokuma. Electronic reorganization: Origin of sigma trans promotion effect. Journal of Computational Chemistry, 28(1): 423–441, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kawamura:2004:HAC

[KN04] Yoshiumi Kawamura and Hiromi Nakai. A hybrid approach combining energy density analysis with the interaction energy decomposition method. Journal of Computational Chemistry, 25(15):1882–1887, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kneller:2005:CUQ

[Kne05] G. R. Kneller. Comment on “Using quaternions to calculate RMSD” [J. Comp. Chem. 25, 1849 (2004)]. Journal of Computational Chemistry, 26(15):1660–1662, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [CSD04] and reply [CSD05].

Knight:2000:RPD

[Kni00] Eugene T. Knight. A radial probability density function for analysis of canonical molecular orbitals. Journal of Computational Chemistry, 21(4):310–321, March 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kobrak:2003:SSE

[Kob03] Mark N. Kobrak. Systematic and statistical error in histogram-based free energy calculations. Journal of Computational Chemistry, 24(12):1437–1446, September 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2009:CSP

[KOFF09] Seonah Kim, Anita M. Orendt, Marta B. Ferraro, and Julio C. Facelli. Crystal structure prediction of flexible molecules using parallel genetic algorithms with a standard force field. Journal of Computational Chemistry, 30(13):1973–1985, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kolafa:2004:TRA

[Kol04] Jiří Kolafa. Time-reversible always stable predictor–corrector method for molecular dynamics of polarizable molecules. Journal of Computational Chemistry, 25(3):335–342, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Komaromi:2008:DGR

[KOML08] István Komáromi, Michael C. Owen, Richard F. Murphy, and Sándor Lovas. Development of glycyl radical parameters for the OPLS-AA/L force field. Journal of Computational Chemistry, 29(12):1999–2009, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Karamertzanis:2005:ICS

[KP05] Panagiotis G. Karamertzanis and Constantinos C. Pantelides. Ab initio crystal structure prediction — I. Rigid molecules. Journal of Computational Chemistry, 26(3):304–324, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kritayakornupong:2004:SDC

[KPR04] Chinapong Kritayakornupong, Kristof Plankensteiner, and Bernd M. Rode. Structure and dynamics of the Cr(III) ion in aqueous solution: ab initio QM/MM molecular dynamics simulation. Journal of Computational Chemistry, 25(13): 1576–1583, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kuta:2006:PDM

[KPZK06] Jadwiga Kuta, Seguei Patchkovskii, Marek Z. Zgierski, and Pawel M. Kozlowski. Performance of DFT in modeling electronic and structural properties of cobalamins. Journal of Computational Chemistry, 27(12):1429–1437, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kritayakornupong:2008:JTE

[Kri08] Chinapong Kritayakornupong. The Jahn–Teller effect of the Cr^{2+} ion in aqueous solution: ab initio QM/MM molecular dynamics simulations. Journal of Computational Chemistry, 29(1):115–121, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kristyan:2009:IPT

[Kri09a] Sandor Kristyan. Interesting properties of Thomas–Fermi kinetic and Parr electron–electron-repulsion DFT energy functional generated compact one-electron density approximation for ground-state electronic energy of molecular systems. Journal of Computational Chemistry, 30(9):1445–1453, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kritayakornupong:2009:SDP

[Kri09b] Chinapong Kritayakornupong. Structural and dynamical properties of the V^{3+} ion in dilute aqueous solution: an ab initio QM/MM molecular dynamics simulation. Journal of Computational Chemistry, 30(16):2777–2783, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kepenekian:2009:EFN

[KRLD09] Mikaël Kepenekian, Vincent Robert, Boris Le Guennic, and Coen De Graaf. Energetics of [Fe(NCH)_{6}]^{2+} via CASPT2 calculations: a spin-crossover perspective. Journal of Computational Chemistry, 30(14):2327–2333, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kaupp:2002:CET

[KRM^{+}02] Martin Kaupp, Roman Reviakine, Olga L. Malkina, Alexei Arbuznikov, Bernd Schimmelpfennig, and Vladimir G. Malkin. Calculation of electronic g-tensors for transition metal complexes using hybrid density functionals and atomic meanfield spin-orbit operators. Journal of Computational Chemistry, 23 (8):794–803, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Krol:2003:CVI

[Kró03] Marcin Król. Comparison of various implicit solvent models in molecular dynamics simulations of immunoglobulin G light chain dimer. Journal of Computational Chemistry, 24(5):531– 546, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kedzierski:2001:ATA

[KS01a] Pawel Kȩdzierski and W. Andrzej Sokalski. Analysis of the transferability of atomic multipoles for amino acids in modeling macromolecular charge distribution from fragments. Journal of Computational Chemistry, 22(10):1082–1097, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kruger:2001:DSH

[KS01b] Thomas Krüger and Alexander F. Sax. Distorted silicon hydrides— a comparative study with various density functionals. Journal of Computational Chemistry, 22(2):151–161, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kinoshita:2002:MPA

[KS02a] Masahiro Kinoshita and Yoshiki Sugai. Methodology of predicting approximate shapes and size distribution of micelles: Illustration for simple models. Journal of Computational Chemistry, 23(15):1445–1455, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kruger:2002:OLA

[KS02b] Thomas Krüger and Alexander F. Sax. Oligovalent link atoms in embedding calculations. Journal of Computational Chemistry, 23(3):371–377, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kasalova:2005:SEA

[KS05a] Veronika Kasalová and Henry F. Schaefer III. Structures and electron affinities of the di-arsenic fluorides As_{2}F_{n} /As_{2}F (n = 1–8). Journal of Computational Chemistry, 26(5):411–435, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Katzer:2005:ITT

[KS05b] Gernot Katzer and Alexander F. Sax. Identification and thermodynamic treatment of several types of large-amplitude motions. Journal of Computational Chemistry, 26(14):1438–1451, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kusama:2005:TSQ

[KS05c] Hitoshi Kusama and Hideki Sugihara. Theoretical study of quinolines–I_{2} intermolecular interaction and implications on dye-sensitized solar cell performance. Journal of Computational Chemistry, 26(13):1372–1382, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kang:2006:KBD

[KS06] Myungshim Kang and Paul E. Smith. A Kirkwood-Buff derived force field for amides. Journal of Computational Chemistry, 27(13):1477–1485, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2008:APL

[KS08] Ryangguk Kim and Jeffrey Skolnick. Assessment of programs for ligand binding affinity prediction. Journal of Computational Chemistry, 29(8):1316–1331, June 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kaminski:2002:DPF

[KSB^{+}02] George A. Kaminski, Harry A. Stern, B. J. Berne, Richard A. Friesner, Yixiang X. Cao, Robert B. Murphy, Ruhong Zhou, and Thomas A. Halgren. Development of a polarizable force field for proteins via ab initio quantum chemistry: First generation model and gas phase tests. Journal of Computational Chemistry, 23(16):1515–1531, December 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kaminsky:2009:MDR

[KŠB09] Jakub Kaminský, Jiří Šebek, and Petr Bouř. Molecular dynamics with restrictions derived from optical spectra. Journal of Computational Chemistry, 30(6):983–991, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kedzierski:2000:NAN

[KSK00] Paweł Kȩdzierski, W. Andrzej Sokalski, and Morris Krauss. Nonempirical analysis of nature of catalytic effects in ribonuclease A active site. Journal of Computational Chemistry, 21 (6):432–445, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kontoyianni:2005:ELR

[KSM05] Maria Kontoyianni, Glenn S. Sokol, and Laura M. Mcclellan. Evaluation of library ranking efficacy in virtual screening. Journal of Computational Chemistry, 26(1):11–22, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kobayashi:2001:TSE

[KSN01] Kaoru Kobayashi, Yuko Sano, and Shigeru Nagase. Theoretical study of endohedral metallofullerenes: Sc_{3−n}La_{n}N@C_{80} (n = 0–3). Journal of Computational Chemistry, 22(13):1353– 1358, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kerber:2008:ASL

[KSS08] Torsten Kerber, Marek Sierka, and Joachim Sauer. Application of semiempirical long-range dispersion corrections to periodic systems in density functional theory. Journal of Computational Chemistry, 29(13):2088–2097, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klopper:2001:EIB

[KSTC01] Wim Klopper, Claire C. M. Samson, György Tarczay, and Attila G. Császár. Equilibrium inversion barrier of NH_{3} from extrapolated coupled-cluster pair energies. Journal of Computational Chemistry, 22(13):1306–1314, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kamiya:2003:ANM

[KSU03] Kenshu Kamiya, Yoko Sugawara, and Hideaki Umeyama. Algorithm for normal mode analysis with general internal coordinates. Journal of Computational Chemistry, 24(7):826–841, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kholmurodov:2000:SPM

[KSY^{+}00] Kholmirzo Kholmurodov, William Smith, Kenji Yasuoka, Tom Darden, and Toshikazu Ebisuzaki. A smooth-particle mesh Ewald method for DL POLY molecular dynamics simulation package on the Fujitsu VPP700. Journal of Computational Chemistry, 21(13):1187–1191, October 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Klimov:2002:TUM

[KT02] D. K. Klimov and D. Thirumalai. Is there a unique melting temperature for two-state proteins? Journal of Computational Chemistry, 23(1):161–165, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Karipidis:2008:DEN

[KT08] Paraskevas A. Karipidis and Constantinos A. Tsipis. The dramatic effect of NH_{3} co-ligation on the FE^{+}-assisted activation of carbon dioxide in the gas phase: From bare metal ions to complexes. Journal of Computational Chemistry, 29(14): 2382–2396, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Katritch:2003:INM

[KTA03] Vsevolod Katritch, Maxim Totrov, and Ruben Abagyan. ICFF: a new method to incorporate implicit flexibility into an internal coordinate force field. Journal of Computational Chemistry, 24(2):254–265, January 30, 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kleinschmidt:2002:SOC

[KTM02] Martin Kleinschmidt, Jörg Tatchen, and Christel M. Marian. Spin-orbit coupling of DFT/MRCI wavefunctions: Method, test calculations, and application to thiophene. Journal of Computational Chemistry, 23(8):824–833, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kloppmann:2007:EDE

[KUB07] Edda Kloppmann, G. Matthias Ullmann, and Torsten Becker. An extended dead-end elimination algorithm to determine gap-free lists of low energy states. Journal of Computational Chemistry, 28(14):2325–2335, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kutzelnigg:2007:EWL

[Kut07] Werner Kutzelnigg. Essay: What I like about Hückel theory. Journal of Computational Chemistry, 28(1):25–34, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Koehler:2000:DSL

[KV00] Ryan T. Koehler and Hugo O. Villar. Design of screening libraries biased for pharmaceutical discovery. Journal of Computational Chemistry, 21(13):1145–1152, October 2000. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kutzner:2007:SPG

[KVF^{+}07] Carsten Kutzner, David Van Der Spoel, Martin Fechner, Erik Lindahl, Udo W. Schmitt, Bert L. De Groot, and Helmut Grubmüller. Speeding up parallel GROMACS on high-latency networks. Journal of Computational Chemistry, 28(12):2075– 2084, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Krautler:2001:FSA

[KvGH01] Vincent Kräutler, Wilfred F. van Gunsteren, and Philippe H. Hünenberger. A fast SHAKE algorithm to solve distance constraint equations for small molecules in molecular dynamics simulations. Journal of Computational Chemistry, 22(5):501– 508, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kosinsky:2004:DFF

[KVL^{+}04] Yuri A. Kosinsky, Pavel E. Volynsky, Philippe Lagant, Gerard Vergoten, Ei-Ichiro Suzuki, Alexander S. Arseniev, and Roman G. Efremov. Development of the force field parameters for phosphoimidazole and phosphohistidine. Journal of Computational Chemistry, 25(11):1313–1321, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kosmopoulou:2006:CEL

[KVS^{+}06] Aggeliki Kosmopoulou, Metaxia Vlassi, Athanassios Stavrakoudis, Constantinos Sakarellos, and Maria Sakarellos-Daitsiotis. T-cell epitopes of the La/SSB autoantigen: Prediction based on the homology modeling of HLA-DQ2/DQ7 with the insulin-B peptide/HLA-DQ8 complex. Journal of Computational Chemistry, 27(9):1033–1044, July 15, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kar:2007:SSB

[KWHH07] Parimal Kar, Yanjie Wei, Ulrich H. E. Hansmann, and Siegfried Höfinger. Systematic study of the boundary composition in Poisson Boltzmann calculations. Journal of Computational Chemistry, 28(16):2538–2544, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kong:2000:QCH

[KWK^{+}00] Jing Kong, Christopher A. White, Anna I. Krylov, David Sherrill, Ross D. Adamson, Thomas R. Furlani, Michael S. Lee, Aaron M. Lee, Steven R. Gwaltney, Terry R. Adams, Christian Ochsenfeld, Andrew T. B. Gilbert, Gary S. Kedziora, Vitaly A. Rassolov, David R. Maurice, Nikhil Nair, Yihan Shao, Nicholas A. Besley, Paul E. Maslen, Jeremy P. Dombroski, Holger Daschel, Weimin Zhang, Prakashan P. Korambath, Jon Baker, Edward F. C. Byrd, Troy Van Voorhis, Manabu Oumi, So Hirata, Chao-Ping Hsu, Naoto Ishikawa, Jan Florian, Arieh Warshel, Benny G. Johnson, Peter M. W. Gill, Martin Head-Gordon, and John A. Pople. Q-chem 2.0: a high-performance ab initio electronic structure program package. Journal of Computational Chemistry, 21(16):1532–1548, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kim:2001:DFT

[KWK^{+}01] Chang Kon Kim, Jongok Won, Hoon Sik Kim, Yong Soo Kang, Hong Guang Li, and Chan Kyung Kim. Density functional theory studies on the dissociation energies of metallic salts: relationship between lattice and dissociation energies. Journal of Computational Chemistry, 22(8):827–834, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [KWK^{+}02].

Kim:2002:EDF

[KWK^{+}02] Chang Kon Kim, Hongok Won, Hoon Sik Kim, Yong Soo Kang, Hong Guang Li, and Chan Kyung Kim. Erratum: Density functional theory studies on the dissociation energies of metallic salts: Relationship between lattice and dissociation energies. Journal of Computational Chemistry, 23(5):584, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [KWK^{+}01].

Kan:2007:TIP

[KYFW07] Wei Kan, Hai-Tao Yu, Hong-Gang Fu, and Yi-Qun Wu. Theoretical investigation on the protonation reactions and products of the stable [N,C,C,S] isomers. Journal of Computational Chemistry, 28(15):2472–2482, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kozhin:2003:NVS

[KYL03] Mikhail Kozhin, Ilya Yanov, and Jerzy Leszczynski. Network visualization system for computational chemistry. Journal of Computational Chemistry, 24(13):1678–1687, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kirschner:2008:GGB

[KYT^{+}08] Karl N. Kirschner, Austin B. Yongye, Sarah M. Tschampel, Jorge González-Outeiriño, Charlisa R. Daniels, B. Lachele Foley, and Robert J. Woods. GLYCAM06: a generalizable biomolecular force field. Carbohydrates. Journal of Computational Chemistry, 29(4):622–655, March 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Koca:2003:CNZ

[KZRO03] Jaroslav Koca, Chang-Guo Zhan, Robert C. Rittenhouse, and Rick L. Ornstein. Coordination number of zinc ions in the phosphotriesterase active site by molecular dynamics and quantum mechanics. Journal of Computational Chemistry, 24 (3):368–378, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kong:2005:MNS

[KZW^{+}05] Qingyu Kong, Li Zhao, Wenning Wang, Chen Wang, Can Xu, Wingham Zhang, Lei Liu, Kangnian Fan, Yufen Li, and Jun Zhuang. Magic number silicon dioxide-based clusters: Laser ablation-mass spectrometric and density functional theory studies. Journal of Computational Chemistry, 26(6):584– 598, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kan:2009:TPR

[KZY09] Wei Kan, Hua Zhong, and Hai-Tao Yu. Theoretical prediction regarding structural and thermodynamical characteristics of stable CH_{3}PO_{2} isomers and unimolecular decomposition mechanisms of species CH_{3}P( = O)_{2}, CH_{3}O − P = O, and CH_{2} = P( = O)OH. Journal of Computational Chemistry, 30(14):2334–2350, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Labute:2008:GBV

[Lab08] Paul Labute. The generalized Born/volume integral implicit solvent model: Estimation of the free energy of hydration using London dispersion instead of atomic surface area. Journal of Computational Chemistry, 29(10):1693–1698, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leon:2001:MAT

[LAEL01] Salvador León, Carlos Alemán, Francesc Escalé, and Manuel Laso. MCDP: an advanced tool to simulate comb-like polymers. Journal of Computational Chemistry, 22(2):162–171, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Laikov:2007:NFA

[Lai07] Dimitri N. Laikov. Neglect of four- and approximation of one-, two-, and three-center two-electron integrals in a symmetrically orthogonalized basis. Journal of Computational Chemistry, 28(3):698–702, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2003:HDM

[LAR^{+}03] Genyuan Li, Maxim Artamonov, Herschel Rabitz, Sheng wei Wang, Panos G. Georgopoulos, and Metin Demiralp. Highdimensional model representations generated from low order terms — lp-RS-HDMR. Journal of Computational Chemistry, 24(5):647–656, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liivat:2005:DFF

[LAT05] Anti Liivat, Alvo Aabloo, and John O. Thomas. Development of a force field for Li_{2}SiF_{6}. Journal of Computational Chemistry, 26(7):716–724, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lipkowitz:1999:RCC

[LB99] Kenny B. Lipkowitz and Donald B. Boyd, editors. Reviews in Computational Chemistry. Wiley-VCH, New York, NY, USA, 1999. ISBN 0-471-33135-X. LCCN ????

Li:2005:DPT

[LB05] Ping Li and Yuxiang Bu. Double proton transfer and oneelectron oxidation behavior in double H-bonded glycinamide– glycine complex in the gas phase. Journal of Computational Chemistry, 26(6):552–560, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2008:ETS

[LB08] Kun Liu and Wensheng Bian. Extensive theoretical study on the low-lying electronic states of silicon monofluoride cation including spin-orbit coupling. Journal of Computational Chemistry, 29(2):256–265, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liang:2008:PDD

[LBG08] Guoming Liang, Xiaoguang Bao, and Jiande Gu. The possibility of the decomposition of 2’-deoxyribose moiety of thymidine induced by the low energy electron attachment. Journal of Computational Chemistry, 29(16):2648–2655, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Levashov:2007:QCP

[LBT07] V. A. Levashov, S. J. L. Billinge, and M. F. Thorpe. Quantum correction to the pair distribution function. Journal of Computational Chemistry, 28(11):1865–1882, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2006:MCH

[LC06] Min Hsien Liu and Cheng Chen. Modified calculations of hydrocarbon thermodynamic properties. Journal of Computational Chemistry, 27(5):537–544, April 15, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lai:2007:COR

[LC07] Chin-Hung Lai and Pi-Tai Chou. Can an OH radical form a strong hydrogen bond? A theoretical comparison with H_{2}O. Journal of Computational Chemistry, 28(8):1357–1363, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liegeois:2009:VRO

[LC09] Vincent Liégeois and Benoît Champagne. Vibrational Raman optical activity of π-conjugated helical systems: Hexahelicene and heterohelicenes. Journal of Computational Chemistry, 30 (8):1261–1278, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lii:2003:AECc

[LCA03] Jenn-Huei Lii, Kuo-Hsiang Chen, and Norman L. Allinger. Alcohols, ethers, carbohydrates, and related compounds. IV. carbohydrates. Journal of Computational Chemistry, 24(12): 1504–1513, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leong:2009:PHC

[LCC09] Max K. Leong, Yen-Ming Chen, and Tzu-Hsien Chen. Prediction of human cytochrome P450 2B6-substrate interactions using hierarchical support vector regression approach. Journal of Computational Chemistry, 30(12):1899–1909, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lii:2003:AECa

[LCDA03] Jenn-Huei Lii, Kuo-Hsiang Chen, Kathleen A. Durkin, and Norman L. Allinger. Alcohols, ethers, carbohydrates, and related compounds. II. The anomeric effect. Journal of Computational Chemistry, 24(12):1473–1489, September 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lii:2003:AECb

[LCGA03] Jenn-Huei Lii, Kuo-Hsiang Chen, T. Bruce Grindley, and Norman L. Allinger. Alcohols, ethers, carbohydrates, and related compounds. III. The 1,2-dimethoxyethane system. Journal of Computational Chemistry, 24(12):1490–1503, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2005:EMD

[LCKL05] Kyoungrim Lee, Cezary Czaplewski, Seung-Yeon Kim, and Jooyoung Lee. An efficient molecular docking using conformational space annealing. Journal of Computational Chemistry, 26(1):78–87, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2009:BCM

[LCSZ09] Bo Liao, Weiyang Chen, Xingming Sun, and Wen Zhu. A binary coding method of RNA secondary structure and its application. Journal of Computational Chemistry, 30(14):2205– 2212, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Launay:2005:TSS

[LD05a] Maxence Launay and Richard Dronskowski. A theoretical study on the structures and energetics of hypothetical TiM(NCN)_{3} compounds of the 3d transition metals. Journal of Computational Chemistry, 26(11):1180–1188, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2005:GAA

[LD05b] Bo Liao and Kequan Ding. Graphical approach to analyzing DNA sequences. Journal of Computational Chemistry, 26(14): 1519–1523, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2007:DIC

[LDC^{+}07] Edmond P. F. Lee, John M. Dyke, Wan-Ki Chow, Foo-Tim Chau, and Daniel K. W. Mok. DFT and ab initio calculations on two reactions between hydrogen atoms and the fire suppressants 2-H heptafluoropropane and CF_{3}Br. Journal of Computational Chemistry, 28(9):1582–1592, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Luo:2002:APB

[LDG02] Ray Luo, Laurent David, and Michael K. Gilson. Accelerated Poisson–Boltzmann calculations for static and dynamic systems. Journal of Computational Chemistry, 23(13):1244–1253, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2009:TDD

[LDL^{+}09] Yufang Liu, Junxia Ding, Ruiqiong Liu, Deheng Shi, and Jinfeng Sun. Time-dependent density functional theory study on the electronic excited-state geometric structure, infrared spectra, and hydrogen bonding of a doubly hydrogen-bonded complex. Journal of Computational Chemistry, 30(16):2723– 2727, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lacaze-Dufaure:2001:IMM

[LDMR01] Corinne Lacaze-Dufaure, Tzonka Mineva, and Nino Russo. On the interaction of Mo and Mo_{2} with NH_{3}, C_{2}H_{4}, and C_{3}H_{6}. Journal of Computational Chemistry, 22(13):1557–1564, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2002:TSMa

[LDT^{+}02a] Jian-Jun Liu, Yi-Hong Ding, Yu-Guo Tao, Ji-Kang Feng, and Chia-Chung Sun. Theoretical study on the mechanism of the 1CHCl + NO reaction. Journal of Computational Chemistry, 23(6):625–649, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2002:TSMb

[LDT^{+}02b] Jian-Jun Liu, Yi-Hong Ding, Yu-Guo Tao, Ji-Kang Feng, and Chia-Chung Sun. Theoretical study on the mechanism of the 3CH_{2} + NO_{2} reaction. Journal of Computational Chemistry, 23(11):1031–1044, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lorenz:2007:DDS

[LDTS07] Sönke Lorenz, Elmar Diederichs, Regina Telgmann, and Christof Schütte. Discrimination of dynamical system models for biological and chemical processes. Journal of Computational Chemistry, 28(8):1384–1399, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2008:TSP

[LDY^{+}08] Rong-Zhen Liao, Wan-Jian Ding, Jian-Guo Yu, Wei-Hai Fang, and Ruo-Zhuang Liu. Theoretical studies on pyridoxal 5’- phosphate-dependent transamination of α-amino acids. Journal of Computational Chemistry, 29(12):1919–1929, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2009:SEE

[Lee09] Gab-Yong Lee. Substituent effect on electron affinity, gasphase basicity, and structure of monosubstituted propynyl radicals and their anions: a theoretical study. Journal of Computational Chemistry, 30(14):2181–2186, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leherte:2006:SMB

[Leh06] Laurence Leherte. Similarity measures based on Gaussiantype promolecular electron density models: Alignment of small rigid molecules. Journal of Computational Chemistry, 27(15):1800–1816, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Latek:2007:PSP

[LEK07] Dorota Latek, Dariusz Ekonomiuk, and Andrzej Kolinski. Protein structure prediction: Combining de novo modeling with sparse experimental data. Journal of Computational Chemistry, 28(10):1668–1676, July 30, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Luttmann:2009:AMD

[LEV^{+}09] Edgar Luttmann, Daniel L. Ensign, Vishal Vaidyanathan, Mike Houston, Noam Rimon, Jeppe Øland, Guha Jayachandran, Mark Friedrichs, and Vijay S. Pande. Accelerating molecular dynamic simulation on the Cell processor and Playstation 3. Journal of Computational Chemistry, 30(2): 268–274, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2002:CRI

[LF02] Wenjian Liu and Robert Franke. Comprehensive relativistic ab initio and density functional theory studies on PtH, PtF, PtCl, and Pt(NH_{3})_{2}Cl_{2}. Journal of Computational Chemistry, 23(5):564–575, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2004:CMTa

[LF04] Xiang-Yuan Li and Ke-Xiang Fu. Continuous medium theory for nonequilibrium solvation: I. How to correctly evaluate solvation free energy of nonequilibrium. Journal of Computational Chemistry, 25(4):500–509, March 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leaver-Fay:2007:MSA

[LFBSK07] Andrew Leaver-Fay, Glenn L. Butterfoss, Jack Snoeyink, and Brian Kuhlman. Maintaining solvent accessible surface area under rotamer substitution for protein design. Journal of Computational Chemistry, 28(8):1336–1341, June 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lopez:2006:CVE

[LFEdL06] Carlos Silva López, Olalla Nieto Faza, Susana López Estévez, and Angel R. de Lera. Computation of vertical excitation energies of retinal and analogs: Scope and limitations. Journal of Computational Chemistry, 27(1):116–123, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Labello:2005:AEC

[LFK05] Nicholas P. Labello, Antonio M. Ferreira, and Henry A. Kurtz. An augmented effective core potential basis set for the calculation of molecular polarizabilities. Journal of Computational Chemistry, 26(14):1464–1471, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

LaValle:2000:RKB

[LFKL00] Steven M. LaValle, Paul W. Finn, Lydia E. Kavraki, and Jean-Claude Latombe. A randomized kinematics-based approach to pharmacophore-constrained conformational search and database screening. Journal of Computational Chemistry, 21(9):731–747, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lopez:2004:SEC

[LFR^{+}04] Xavier López, Jorge A. Fernández, Susanna Romo, Jean François Paul, Leonid Kazansky, and Josep M. Poblet. Are the solvent effects critical in the modeling of polyoxoanions? Journal of Computational Chemistry, 25(12):1542–1549, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2007:SEO

[LFR07] Yan-Ling Liu, Ji-Kang Feng, and Ai-Min Ren. Structural, electronic, and optical properties of phosphole-containing πconjugated oligomers for light-emitting diodes. Journal of Computational Chemistry, 28(15):2500–2509, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lopez:2007:PDD

[LFS^{+}07] Carlos Silva López, Olalla Nieto Faza, José A. Souto, Rosana Álvarez, and Ángel R. De Lera. Pseudopericyclic design drives antara-antara [1,5] methylene sigmatropic shifts from a stepwise to a concerted mechanism. Journal of Computational Chemistry, 28(8):1411–1416, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2003:ENA

[LFSB03a] Michael S. Lee, Michael Feig, Freddie R. Salsbury Jr., and Charles L. Brooks III. Erratum: New analytic approximation to the standard molecular volume definition and its application to generalized Born calculations. Journal of Computational Chemistry, 24(14):1821, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [LFSB03b].

Lee:2003:NAA

[LFSB03b] Michael S. Lee, Michael Feig, Freddie R. Salsbury Jr., and Charles L. Brooks III. New analytic approximation to the standard molecular volume definition and its application to generalized Born calculations. Journal of Computational Chemistry, 24(11):1348–1356, August 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [LFSB03a].

Li:2004:CMTb

[LFZS04] Xiang-Yuan Li, Ke-Xiang Fu, Quan Zhu, and Min-Hua Shan. Continuous medium theory for nonequilibrium solvation: II. Interaction energy between solute charge and reaction field and single-sphere model for spectral shift. Journal of Computational Chemistry, 25(6):835–842, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2009:CTP

[LGB^{+}09] Ji-Lai Li, Cai-Yun Geng, Yuxiang Bu, Xu-Ri Huang, and Chia-Chung Sun. Conformational transition pathway in the allosteric process of calcium-induced recoverin: Molecular dynamics simulations. Journal of Computational Chemistry, 30 (7):1135–1145, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2002:TIE

[LH02] Xiang-Yuan Li and Chun-Xiu Hu. Theoretical investigation of electron transfer transition in tetracyanoethylene-contained organic complexes. Journal of Computational Chemistry, 23 (9):874–886, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lins:2005:NGF

[LH05] Roberto D. Lins and Philippe H. Hünenberger. A new GRO-MOS force field for hexopyranose-based carbohydrates. Journal of Computational Chemistry, 26(13):1400–1412, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2009:BHR

[LHI09] Jinhyuk Lee, Sihyun Ham, and Wonpil Im. Beta-hairpin restraint potentials for calculations of potentials of mean force as a function of beta-hairpin tilt, rotation, and distance. Journal of Computational Chemistry, 30(8):1334–1343, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2006:ABS

[LHJ^{+}06] Tzong-Yi Lee, Jorng-Tzong Horng, Hsueh-Fen Juan, Hsien-Da Huang, Li-Cheng Wu, Meng-Fong Tsai, and Hsuan-Cheng Huang. An agent-based system to discover protein–protein interactions, identify protein complexes and proteins with multiple peptide mass fingerprints. Journal of Computational Chemistry, 27(9):1020–1032, July 15, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Linnolahti:2001:GPB

[LHP01] Mikko Linnolahti, Pipsa Hirva, and Tapani A. Pakkanen. Geometry prediction of bridged zirconocene dichlorides by quantum chemical methods. Journal of Computational Chemistry, 22(1):51–64, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2007:SRS

[LhWX07] Chun Li, Ai hua Wang, and Lili Xing. Similarity of RNA secondary structures. Journal of Computational Chemistry, 28(2):508–512, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2001:ETB

[Li01] Xiang-Yuan Li. Electron transfer between tryptophan and tyrosine: Theoretical calculation of electron transfer matrix element for intramolecular hole transfer. Journal of Computational Chemistry, 22(6):565–579, April 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2007:IAH

[LI07] Jinhyuk Lee and Wonpil Im. Implementation and application of helix–helix distance and crossing angle restraint potentials. Journal of Computational Chemistry, 28(3):669–680, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lipkowitz:2000:BRB

[Lip00] Kenny B. Lipkowitz. Book review: Encyclopedia of Computational Chemistry, P. v. R. Schleyer, editor-in-chief, John Wiley & Sons, Chichester, UK, 1998, 3,500 pp. $3,150. ISBN 0-471-96588-X. Journal of Computational Chemistry, 21(1):77–78, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2004:IEC

[LJ04] Hui Li and Jan H. Jensen. Improving the efficiency and convergence of geometry optimization with the polarizable continuum model: New energy gradients and molecular surface tessellation. Journal of Computational Chemistry, 25(12):1449– 1462, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2008:RES

[LJKL08] Jinwoo Lee, Keehyoung Joo, Seung-Yeon Kim, and Jooyoung Lee. Re-examination of structure optimization of off-lattice protein AB models by conformational space annealing. Journal of Computational Chemistry, 29(14):2479–2484, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lloyd:2005:SIE

[LJS05] Lesley D. Lloyd, Roy L. Johnston, and Said Salhi. Strategies for increasing the efficiency of a genetic algorithm for the structural optimization of nanoalloy clusters. Journal of Computational Chemistry, 26(10):1069–1078, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lin:2007:SVM

[LJZ^{+}07] Wei-Qi Lin, Jian-Hui Jiang, Yan-Ping Zhou, Hai-Long Wu, Guo-Li Shen, and Ru-Qin Yu. Support vector machine based training of multilayer feedforward neural networks as optimized by particle swarm algorithm: Application in QSAR studies of bioactivity of organic compounds. Journal of Computational Chemistry, 28(2):519–527, January 30, 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lovallo:2003:DNP

[LK03] Christopher C. Lovallo and Mariusz Klobukowski. Development of new pseudopotential methods: Improved model core potentials for the first-row transition metals. Journal of Computational Chemistry, 24(9):1009–1015, July 15, 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lovallo:2004:IMC

[LK04] Christopher C. Lovallo and Mariusz Klobukowski. Improved model core potentials for the second- and third-row transition metals. Journal of Computational Chemistry, 25(9):1206– 1213, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Luman:2001:PAC

[LKA01] Nathaniel R. Luman, Michael P. King, and Joseph D. Augspurger. Predicting ^{15}N amide chemical shifts in proteins. I. An additive model for the backbone contribution. Journal of Computational Chemistry, 22(3):366–372, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2004:CSC

[LKJ^{+}04] Ho-Jin Lee, Jong Hyun Kim, Hee Jung Jung, Kun-Young Kim, Eun-Jung Kim, Young-Sang Choi, and Chang-Ju Yoon. Computational study of conformational preferences of thioamidecontaining azaglycine peptides. Journal of Computational Chemistry, 25(2):169–178, January 30, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Linnanto:2004:SPM

[LKT04] Juha Linnanto and Jouko Korppi-Tommola. Semiempirical PM5 molecular orbital study on chlorophylls and bacteriochlorophylls: Comparison of semiempirical, ab initio, and density functional results. Journal of Computational Chemistry, 25(1):123–138, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lu:2004:IER

[LKW04] Nandou Lu, David A. Kofke, and Thomas B. Woolf. Improving the efficiency and reliability of free energy perturbation calculations using overlap sampling methods. Journal of Computational Chemistry, 25(1):28–40, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lamprecht:2000:MMF

[LL00] Delanie Lamprecht and Gert J. Lamprecht. A molecular mechanics force field for rhodium(I) carbonyl phosphine complexes and its application on the oxidative addition reactions of these complexes. Journal of Computational Chemistry, 21 (8):692–703, June 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2001:TSS

[LL01] Xiang-Yuan Li and Ji-Feng Liu. A theoretical study of solvent effect and reaction mechanism of electron transfer reaction between indole side chain and phenol side chain of peptide involving tryptophan and tyrosine. Journal of Computational Chemistry, 22(10):1067–1081, July 30, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lin:2007:UPA

[LL07] Hao Lin and Qian-Zhong Li. Using pseudo amino acid composition to predict protein structural class: Approached by incorporating 400 dipeptide components. Journal of Computational Chemistry, 28(9):1463–1466, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Langley:2001:MMCa

[LLA01a] Charles H. Langley, Jenn-Huei Lii, and Norman L. Allinger. Molecular mechanics calculations on carbonyl compounds. II. Open-chain ketones. Journal of Computational Chemistry, 22(13):1426–1450, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Langley:2001:MMCb

[LLA01b] Charles H. Langley, Jenn-Huei Lii, and Norman L. Allinger. Molecular mechanics calculations on carbonyl compounds. III. Cycloketones. Journal of Computational Chemistry, 22(13): 1451–1475, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Langley:2001:MMCc

[LLA01c] Charles H. Langley, Jenn-Huei Lii, and Norman L. Allinger. Molecular mechanics calculations on carbonyl compounds. IV. Heats of formation. Journal of Computational Chemistry, 22(13):1476–1483, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Langley:2001:MMM

[LLA01d] Charles H. Langley, Jenn-Huei Lii, and Norman L. Allinger. Molecular mechanics (MM4) calculations on carbonyl compounds part I: aldehydes. Journal of Computational Chemistry, 22(13):1396–1425, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Langley:2003:MMM

[LLA03] Charles H. Langley, Jenn-Huei Lii, and Norman Allinger. Molecular mechanics (MM4) calculations on carbonyl compounds. I–IV. Journal of Computational Chemistry, 24(10): 1283–1286, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lim:2006:TST

[LLKC06] Chultack Lim, Hee Soon Lee, Young-Woo Kwak, and Cheol Ho Choi. A theoretical study of thermal [1,3]-sigmatropic rearrangements of 3-trimethylsilyl-1-pyrazoline: Concerted vs. stepwise mechanisms. Journal of Computational Chemistry, 27(2):228–237, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Loferer:2003:QMI

[LLL03] Markus J. Loferer, Hannes H. Loeffler, and Klaus R. Liedl. A QM–MM interface between CHARMM and TURBOMOLE:

Implementation and application to systems in bulk phase and biologically active systems. Journal of Computational Chemistry, 24(10):1240–1249, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lu:2007:SSH

[LLL07] Hsiu-Feng Lu, Feng-Yin Li, and S. H. Lin. Site specificity of α-H abstraction reaction among secondary structure motif — an ab initio study. Journal of Computational Chemistry, 28 (4):783–794, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2008:QSM

[LLL^{+}08] Jiazhong Li, Beilei Lei, Huanxiang Liu, Shuyan Li, Xiaojun Yao, Mancang Liu, and Paola Gramatica. QSAR study of malonyl-CoA decarboxylase inhibitors using GA-MLR and a new strategy of consensus modeling. Journal of Computational Chemistry, 29(16):2636–2647, December 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lammers:2008:RFF

[LLM08] Sven Lammers, Stephan Lutz, and Markus Meuwly. Reactive force fields for proton transfer dynamics. Journal of Computational Chemistry, 29(7):1048–1063, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lopes:2009:PEF

[LLM09] Pedro E. M. Lopes, Guillaume Lamoureux, and Alexander D. Mackerell Jr. Polarizable empirical force field for nitrogencontaining heteroaromatic compounds based on the classical Drude oscillator. Journal of Computational Chemistry, 30(12): 1821–1838, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2003:PAD

[LLS03] Meng-Sheng Liao, Yun Lu, and Steve Scheiner. Performance assessment of density-functional methods for study of chargetransfer complexes. Journal of Computational Chemistry, 24 (5):623–631, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lin:2002:TSS

[LLW02] Jin-Qiu Lin, Shi-Wei Luo, and Yun-Dong Wu. Theoretical study of sheets formed by β-peptides. Journal of Computational Chemistry, 23(16):1551–1558, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Luder:2009:SPD

[LLW^{+}09] Kai Lüder, Lennart Lindfors, Jan Westergren, Sture Nordholm, Rasmus Persson, and Mikaela Pedersen. In silico prediction of drug solubility: 4. Will simple potentials suffice? Journal of Computational Chemistry, 30(12):1859–1871, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2002:MMG

[LLXS02] Qian-Shu Li, Rui-Hua Lü, Yaoming Xie, and Henry F. Schaefer III. Molecules for materials: Germanium hydride neutrals and anions. Molecular structures, electron affinities, and thermochemistry of GeH_{n} /GeH (n = 0–4) and Ge_{2}H_{n} /Ge_{2}H (n = 0–6). Journal of Computational Chemistry, 23(16):1642– 1655, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2009:APT

[LLZL09] Cun-Xi Liu, Ze-Rong Li, Chong-Wen Zhou, and Xiang-Yuan Li. Accurate prediction of thermodynamic properties of alkyl peroxides by combining density functional theory calculation with least-square calibration. Journal of Computational Chemistry, 30(7):1007–1015, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lorenzo:2003:BCB

[LM03] Luis Lorenzo and Ricardo A. Mosquera. A box-counting-based algorithm for computing Shannon entropy in molecular dynamics simulations. Journal of Computational Chemistry, 24 (6):707–713, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Levi:2008:FVF

[LMB08] Chen Levi, Jan M. L. Martin, and Ilana Bar. Fundamental vibrational frequencies and dominant resonances in methylamine isotopologues by ab initio and density functional theory methods. Journal of Computational Chemistry, 29(8):1268– 1276, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2009:CIF

[LMCD09] Edmond P. F. Lee, Daniel K. W. Mok, Foo-Tim Chau, and John M. Dyke. A combined ab initio and Franck–Condon factor simulation study on the photodetachment spectrum of ScO_{2}^{−}. Journal of Computational Chemistry, 30(3):337–345, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lamsabhi:2009:MDP

[LMGO^{+}09] Al Mokhtar Lamsabhi, Otilia Mó, Soledad Gutiérrez-Oliva, Patricia Pérez, Alejandro Toro-Labbé, and Manuel Yáñez. The mechanism of double proton transfer in dimers of uracil and 2-thiouracil — the reaction force perspective. Journal of Computational Chemistry, 30(3):389–398, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2005:CBS

[LMGR05] Si-Dian Li, Chang-Qing Miao, Jin-Chang Guo, and Guang-Ming Ren. Carbon boronyls: Species with higher viable possibility than boron carbonyls at the density functional theory. Journal of Computational Chemistry, 26(8):799–802, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2006:TMB

[LMGR06] Si-Dian Li, Chang-Qing Miao, Jin-Chang Guo, and Guang-Ming Ren. Transition metal–boron complexes B_{n}M: From bowls (n = 8–14) to tires (n = 14). Journal of Computational Chemistry, 27(15):1858–1865, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Loksha:2009:SSH

[LMH^{+}09] Ilya V. Loksha, James R. Maiolo III, Cheng W. Hong, Albert Ng, and Christopher D. Snow. SHARPEN — Systematic Hierarchical Algorithms for Rotamers and Proteins on an Extended Network. Journal of Computational Chemistry, 30(6): 999–1005, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Loffler:2006:WED

[LMIF06] Hannes H. Loffler, Ahmed M. Mohammed, Yasuhiro Inada, and Shigenobu Funahashi. Water exchange dynamics of manganese(II), cobalt(II), and nickel(II) ions in aqueous solution. Journal of Computational Chemistry, 27(16):1944–1949, December 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2002:LSL

[LMJ02] Shuhua Li, Jing Ma, and Yuansheng Jiang. Linear scaling local correlation approach for solving the coupled cluster equations of large systems. Journal of Computational Chemistry, 23(2): 237–244, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Le:2001:ISC

[LMK01] Trung Ngoc Le, Alexander M. Mebel, and Ralf I. Kaiser. Ab initio study of C_{4}H_{3} potential energy surface and reaction of ground-state carbon atom with propargyl radical. Journal of Computational Chemistry, 22(13):1522–1535, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2000:MBI

[LML^{+}00] Young Joo Lee, Jae Yeol Maeng, Eok-Kyun Lee, Bongsoo Kim, Sehun Kim, and Kyu-Kwang Han. Melting behaviors of icosahedral metal clusters studied by Monte Carlo simulations. Journal of Computational Chemistry, 21(5):380–387, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liddell:2004:ASS

[LMMW04] Michael J. Liddell, Davor Margetic, Anthony S. Mitchell, and Ronald N. Warrener. An AM1 semiempirical study of host– guest complexation in hemicarcerand complexes. Journal of Computational Chemistry, 25(4):542–557, March 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Luo:2009:AES

[LMO09] Yi Luo, Satoshi Maeda, and Koichi Ohno. Automated exploration of stable isomers of H^{+}(H_{2}O)_{n} (n = 5–7) via ab initio calculations: an application of the anharmonic downward distortion following algorithm. Journal of Computational Chemistry, 30(6):952–961, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lopez-Martinez:2009:TSE

[LMRVFH^{+}09] Erika Ivonne López-Martínez, Luz María Rodríguez-Valdez, Norma Flores-Holguín, Alfredo Márquez-Lucero, and Daniel Glossman-Mitnik. Theoretical study of electronic properties of organic photovoltaic materials. Journal of Computational Chemistry, 30(7):1027–1037, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Larin:2007:QSE

[LMV07] A. V. Larin, W. J. Mortier, and D. P. Vercauteren. Quick scheme for evaluation of atomic charges in arbitrary aluminophosphate sieves on the basis of electron densities calculated with DFT methods. Journal of Computational Chemistry, 28(10):1695–1703, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2001:RGC

[LN01] Chao-Ping Liu and J. J. Soares Neto. Revisiting the generator coordinate approximation for calculating the ro-vibrational energies of H. Journal of Computational Chemistry, 22(16): 2028–2039, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Loeffler:2003:MBE

[Loe03] Hannes H. Loeffler. Many-body effects on structure and dynamics of aqueous ionic solutions. Journal of Computational Chemistry, 24(10):1232–1239, July 30, 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [?].

Liu:2008:RIT

[LOL^{+}08] Qi Liu, V. Olman, Huiqing Liu, Xiuzi Ye, Shilun Qiu, and Ying Xu. RNACluster: an integrated tool for RNA secondary structure comparison and clustering. Journal of Computational Chemistry, 29(9):1517–1526, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lorenzo:2003:NIF

[LPB03] Alicia C. Lorenzo, Pedro G. Pascutti, and Paulo M. Bisch. Nonspecific interaction forces at water–membrane interface by forced molecular dynamics simulations. Journal of Computational Chemistry, 24(3):328–339, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2007:CSM

[LPK07] Sang-Ho Lee, Kim Palmo, and Samuel Krimm. A comparative study of molecular dynamics in Cartesian and in internal coordinates: Dynamical instability in the latter caused by nonlinearity of the equations of motion. Journal of Computational Chemistry, 28(6):1107–1118, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lehnert:2006:ESI

[LPP06] Nicolai Lehnert, V. K. K. Praneeth, and Florian Paulat. Electronic structure of iron(II)–porphyrin nitroxyl complexes: Molecular mechanism of fungal nitric oxide reductase (P450nor). Journal of Computational Chemistry, 27(12):1338– 1351, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Laederach:2003:SEF

[LR03a] Alain Laederach and Peter J. Reilly. Specific empirical free energy function for automated docking of carbohydrates to proteins. Journal of Computational Chemistry, 24(14):1748– 1757, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Loeffler:2003:EMB

[LR03b] Hannes H. Loeffler and Bernd M. Rode. Erratum: Manybody effects on structure and dynamics of aqueous ionic solutions. Journal of Computational Chemistry, 24(13):1688, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [Loe03].

Li:2006:RCV

[LR06] Genyuan Li and Herschel Rabitz. Ratio control variate method for efficiently determining high-dimensional model representations. Journal of Computational Chemistry, 27(10):1112–1118, July 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Langella:2002:CAT

[LRI^{+}02] Emma Langella, Nadia Rega, Roberto Improta, Orlando Crescenzi, and Vincenzo Barone. Conformational analysis of the tyrosine dipeptide analogue in the gas phase and in aqueous solution by a density functional/continuum solvent model. Journal of Computational Chemistry, 23(6):650–661, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2003:CMV

[LRWG03] Genyuan Li, Herschel Rabitz, Sheng-Wei Wang, and Panos G. Georgopoulos. Correlation method for variance reduction of Monte Carlo integration in RS-HDMR. Journal of Computational Chemistry, 24(3):277–283, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2002:CSM

[LS02] Meng-Sheng Liao and Steve Scheiner. Comparative study of metal-porphyrins, -porphyrazines, and -phthalocyanines. Journal of Computational Chemistry, 23(15):1391–1403, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lin:2005:DGO

[LS05a] Youdong Lin and Mark A. Stadtherr. Deterministic global optimization of molecular structures using interval analysis. Journal of Computational Chemistry, 26(13):1413–1420, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lundberg:2005:ABE

[LS05b] Marcus Lundberg and Per E. M. Siegbahn. Agreement between experiment and hybrid DFT calculations for O H bond dissociation enthalpies in manganese complexes. Journal of Computational Chemistry, 26(7):661–667, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lai:2008:TCH

[LS08a] Chin-Hung Lai and Ming-Der Su. Theoretical characterizations of HAsXH (X = N, P, As, Sb, and Bi) isomers in the singlet and triplet states. Journal of Computational Chemistry, 29(15):2487–2499, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lie:2008:DSS

[LS08b] Mette Alstrup Lie and Birgit Schiøtt. A DFT study of solvation effects on the tautomeric equilibrium and catalytic ylide generation of thiamin models. Journal of Computational Chemistry, 29(7):1037–1047, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2008:EBO

[LS08c] Aibing Liu and Steven J. Stuart. Empirical bond-order potential for hydrocarbons: Adaptive treatment of van der Waals interactions. Journal of Computational Chemistry, 29(4):601– 611, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leininger:2001:SSS

[LSAS01] Matthew L. Leininger, C. David Sherrill, Wesley D. Allen, and Henry F. Schaefer III. Systematic study of selected diagonalization methods for configuration interaction matrices. Journal of Computational Chemistry, 22(13):1574–1589, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lange:2006:FGA

[LSG06] Oliver F. Lange, Lars V. Schäfer, and Helmut Grubmüller. Flooding in GROMACS: Accelerated barrier crossings in molecular dynamics. Journal of Computational Chemistry, 27(14):1693–1702, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2004:MHA

[LSHR04] Genyuan Li, Jacqueline Schoendorf, Tak-San Ho, and Herschel Rabitz. Multicut-HDMR with an application to an ionospheric model. Journal of Computational Chemistry, 25(9): 1149–1156, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2004:EHE

[LSO04] Michael S. Lee, Freddie R. Salsbury Jr., and Mark A. Olson. An efficient hybrid explicit/implicit solvent method for biomolecular simulations. Journal of Computational Chemistry, 25(16):1967–1978, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Larin:2008:CEF

[LST08] A. V. Larin, I. K. Sakodynskaya, and D. N. Trubnikov. Convergence of electric field and electric field gradient versus atomic basis sets in all-siliceous and Mg substituted phillipsites. Journal of Computational Chemistry, 29(14):2344–2358, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Loeffler:2001:CSD

[LSW^{+}01] Hannes H. Loeffler, Christoph A. Sotriffer, Rudolf H. Winger, Klaus R. Liedl, and Bernd M. Rode. Calculation of sequencedependent free energies of hydration of dipeptides formed by alanine and glycine. Journal of Computational Chemistry, 22 (8):846–860, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liang:2000:MMM

[LSWB00] Guyan Liang, Jennifer B. Sorensen, David Whitmire, and J. Phillip Bowen. Molecular mechanics (MM3) parameterization for oxocarbenium ions. Journal of Computational Chemistry, 21(5):329–339, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lu:2002:GTN

[LSY02] Qingzhang Lu, Guoli Shen, and Ruqin Yu. Genetic training of network using chaos concept: Application to QSAR studies of vibration modes of tetrahedral halides. Journal of Computational Chemistry, 23(14):1357–1365, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2007:EPT

[LTF^{+}07] Wei-Qi Li, Wei-Quan Tian, Ji-Kang Feng, Zi-Zhong Liu, Ai-Min Ren, Chia-Chung Sun, and Yuriko Aoki. Electronic properties of tricoordinated phosphorus in hexagonal phosphininium compounds and molecular aromaticity. Journal of Computational Chemistry, 28(9):1467–1475, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Larin:2008:EFC

[LTV08] A. V. Larin, D. N. Trubnikov, and D. P. Vercauteren. Electric field convergence versus atomic basis sets in all-siliceous zeolites. Journal of Computational Chemistry, 29(1):130–138, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lu:2009:GEA

[Lu09] Shih-I Lu. Gibbs energy of activation for thermal isomerization of (1Z)-acetaldehyde hydrazone and (1Z)-acetaldehyde N,N-dimethylhydrazone by Gaussian-4 theory and CCSD(T)/ CBS computations. Journal of Computational Chemistry, 30 (14):2176–2180, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leherte:2008:CMR

[LV08] Laurence Leherte and Daniel P. Vercauteren. Collective motions of rigid fragments in protein structures from smoothed electron density distributions. Journal of Computational Chemistry, 29(9):1472–1489, July 15, 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2004:DDS

[LW04a] Qian-Shu Li and Chao Yang Wang. Direct dynamic study on the hydrogen abstraction reaction CH_{3}CN + OH → CH_{2}CN + H_{2}O. Journal of Computational Chemistry, 25(2):251–257, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2004:NGR

[LW04b] Bo Liao and Tian-Ming Wang. New 2D graphical representation of DNA sequences. Journal of Computational Chemistry, 25(11):1364–1368, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2006:RPS

[LW06] Liwei Liu and Tianming Wang. 2D representation of protein secondary structure sequences and its applications. Journal of Computational Chemistry, 27(11):1119–1124, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Landis:2007:VEV

[LW07] C. R. Landis and F. Weinhold. Valence and extra-valence orbitals in main group and transition metal bonding. Journal of Computational Chemistry, 28(1):198–203, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2006:APD

[LWH06] Meng-Sheng Liao, John D. Watts, and Ming-Ju Huang. Assessment of the performance of density-functional methods for calculations on iron porphyrins and related compounds. Journal of Computational Chemistry, 27(13):1577–1592, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2008:AES

[LWK08] Changhoon Lee, Myung-Hwan Whangbo, and Jürgen Köhler. Analysis of electronic structures and chemical bonding of metal-rich compounds. I. Density functional study of Pt metal, LiPt_{2}, LiPt, and Li_{2}Pt. Journal of Computational Chemistry, 29(13):2154–2160, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2007:SEP

[LWLS07] Ying Li, Di Wu, Zhi-Ru Li, and Chia-Chung Sun. Structural and electronic properties of boron-doped lithium clusters: ab initio and DFT studies. Journal of Computational Chemistry, 28(10):1677–1684, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2006:RCS

[LWW^{+}06] Zhi-Ru Li, Fang-Fang Wang, Di Wu, Ying Li, Wei Chen, Xiao-Ying Sun, Feng Long Gu, and Yuriko Aoki. Royal crownshaped electride Li_{3}–N_{3}–Be containing two superatoms: New knowledge on aromaticity. Journal of Computational Chemistry, 27(8):986–993, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2007:TSV

[LWX07] Yongjian Li, Jian Wan, and Xin Xu. Theoretical study of the vertical excited states of benzene, pyrimidine, and pyrazine by the symmetry adapted cluster–configuration interaction method. Journal of Computational Chemistry, 28(10):1658– 1667, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2009:COR

[LWY^{+}09] Hong-Xia Liu, Ying Wang, Lei Yang, Jing-Yao Liu, Hong Gao, Ze-Sheng Li, and Chia-Chung Sun. CH_{3}NHNH_{2} + OH reaction: Mechanism and dynamics studies. Journal of Computational Chemistry, 30(14):2194–2204, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lin:2009:NHD

[LWZ09] Jian Lin, Kechen Wu, and Mingxin Zhang. A new hybrid DFT approach to electronic excitation and first hyperpolarizabilities of transition metal complexes. Journal of Computational Chemistry, 30(13):2056–2063, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2007:QCS

[LX07] Qian Shu Li and Yu Xu. A quantum chemistry study: a new kind of boron nitrides. Journal of Computational Chemistry, 28(8):1446–1455, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2007:NCC

[LXL07] Xiaoqing Liu, Zhilong Xiu, and Xiaohui Li. Numerical characterization of the conformation of cyclic peptides and its application. Journal of Computational Chemistry, 28(16):2545– 2551, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2008:SOI

[LXSF08] Yajun Liu, Hongyan Xiao, Mengtao Sun, and Weihai Fang. Spin-orbit ab initio investigation of the photodissociation of dibromomethane in the gas and solution phases. Journal of Computational Chemistry, 29(15):2513–2519, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2009:NMP

[LXW^{+}09] Shuyan Li, Lili Xi, Chengqi Wang, Jiazhong Li, Beilei Lei, Huanxiang Liu, and Xiaojun Yao. A novel method for proteinligand binding affinity prediction and the related descriptors exploration. Journal of Computational Chemistry, 30(6):900– 909, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liao:2006:CPB

[LXZ06] Bo Liao, Xuyu Xiang, and Wen Zhu. Coronavirus phylogeny based on 2D graphical representation of DNA sequence. Journal of Computational Chemistry, 27(11):1196–1202, August 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lee:2004:IDS

[LYK^{+}04] Oh-Seuk Lee, Kiyull Yang, Keum Duck Kang, In Sun Koo, Chan-Kyung Kim, and Ikchoon Lee. Ab initio and DFT studies on hydrolyses of phosphorus halides. Journal of Computational Chemistry, 25(14):1740–1748, November 15, 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2008:LAC

[LYS08] Qingxu Li, Yuanping Yi, and Zhigang Shuai. Local approach to coupled cluster evaluation of polarizabilities for long conjugated molecules. Journal of Computational Chemistry, 29(10): 1650–1655, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2008:TSS

[LYZ^{+}08] Xiao-Jing Liu, Chuan-Lu Yang, Xiang Zhang, Ke-Li Han, and Zi-Chao Tang. Theoretical study on the structure and formation mechanism of [C_{6}H_{5}M_{m}]^{−} (M = Ag, Au; m = 1–3). Journal of Computational Chemistry, 29(10):1667–1674, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2005:SFS

[LZ05a] Hongzhi Li and Yaoqi Zhou. SCUD: Fast structure clustering of decoys using reference state to remove overall rotation. Journal of Computational Chemistry, 26(11):1189–1192, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Liu:2005:SCP

[LZ05b] Bing Liu and Jiaju Zhou. SARS-CoV protease inhibitors design using virtual screening method from natural products libraries. Journal of Computational Chemistry, 26(5):484–490, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Leon:2002:IPS

[LZA02] Salvador León, David Zanuy, and Carlos Alemán. Influence of the presence of small gas molecules in the structure of comblike polyacrylates: a Monte Carlo study. Journal of Computational Chemistry, 23(7):685–696, May 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lu:2009:ICH

[LZF^{+}09] Yun-Xiang Lu, Jian-Wei Zou, Ji-Cai Fan, Wen-Na Zhao, Yong-Jun Jiang, and Qing-Sen Yu. Ab initio calculations on halogen-bonded complexes and comparison with density functional methods. Journal of Computational Chemistry, 30(5): 725–732, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2003:ISI

[LZJ03] Jinshan Li, Feng Zhao, and Fuqian Jing. An ab initio study of intermolecular interactions of nitromethane dimer and nitromethane trimer. Journal of Computational Chemistry, 24 (3):345–352, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Lei:2004:SPC

[LZKT04] Ming Lei, Maria I. Zavodszky, Leslie A. Kuhn, and M. F. Thorpe. Sampling protein conformations and pathways. Journal of Computational Chemistry, 25(9):1133–1148, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Li:2009:ICB

[LZZC09] Huaiyong Li, Siyuan Zhang, Shihong Zhou, and Xueqiang Cao. Investigation of chemical bond characteristics, thermal expansion coefficients and bulk moduli of α-R_{2}MoO_{6} and R_{2}Mo_{2}O_{7} (R = rare earths) by using a dielectric chemical bond method. Journal of Computational Chemistry, 30(12):

1892–1898, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mobli:2005:QVC

[MA05] Mehdi Mobli and Raymond J. Abraham. Quantum vs. classical models of the nitro group for proton chemical shift calculations and conformational analysis. Journal of Computational Chemistry, 26(4):389–398, March 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Miura:2009:ITT

[MA09] Masanori Miura and Yuriko Aoki. Ab initio theory for treating local electron excitations in molecules and its performance for computing optical properties. Journal of Computational Chemistry, 30(14):2213–2230, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Martinez:2009:PPB

[MABM09] L. Martínez, R. Andrade, E. G. Birgin, and J. M. Martínez. Packmol: a package for building initial configurations for molecular dynamics simulations. Journal of Computational Chemistry, 30(13):2157–2164, October 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mackerell:2004:EFF

[Mac04] Alexander D. Mackerell Jr. Empirical force fields for biological macromolecules: Overview and issues. Journal of Computational Chemistry, 25(13):1584–1604, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Matamala:2007:SAI

[MAF^{+}07] Adelio R. Matamala, Daniel E. Almonacid, Maximiliano F. Figueroa, José Martínez-oyanedel, and Marta C. Bunster. A semiempirical approach to the intra-phycocyanin and interphycocyanin fluorescence resonance energy-transfer pathways in phycobilisomes. Journal of Computational Chemistry, 28 (7):1200–1207, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mak:2008:RCS

[Mak08] C. H. Mak. RNA conformational sampling. I. Single-nucleotide loop closure. Journal of Computational Chemistry, 29(6):926– 933, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Maroulis:2003:IDE

[Mar03] George Maroulis. Ab initio determination of the electric multipole moments and static (hyper)polarizability of HCCX, X = F, Cl, Br, and I. Journal of Computational Chemistry, 24 (4):443–452, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Masamura:2001:IMOa

[Mas01a] Masao Masamura. Ab initio molecular orbital study of OH^{−}(H_{2}O)_{n} in the gas phase. Journal of Computational Chemistry, 22(1):31–37, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Masamura:2001:IMOb

[Mas01b] Masao Masamura. Ab initio molecular orbital study on the structures and energetics of CH_{3}OH(H_{2}O)_{n} and CH_{3}SH(H_{2}O)_{n} in the gas phase. Journal of Computational Chemistry, 22(1):125–131, January 15, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Masamura:2004:EBS

[Mas04] Masao Masamura. The effect of basis set superposition error on the convergence of intermolecular interaction energies for deprotonated complexes. Journal of Computational Chemistry, 25(14):1771–1778, November 15, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Matta:2003:AQT

[Mat03] Chérif F. Matta. Application of the quantum theory of atoms in molecules to selected physico-chemical and biophysical problems: Focus on correlation with experiment. Journal of Computational Chemistry, 24(4):453–463, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mayer:2007:BOV

[May07] I. Mayer. Bond order and valence indices: a personal account. Journal of Computational Chemistry, 28(1):204–221, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mazur:2001:MDM

[Maz01] Alexey K. Mazur. Molecular dynamics of minimal B-DNA. Journal of Computational Chemistry, 22(4):457–467, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mazur:2008:ISS

[Maz08] Grzegorz Mazur. An improved SCPF scheme for polarization energy calculations. Journal of Computational Chemistry, 29 (6):988–993, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

MacKerell:2000:AAE

[MB00] Alexander D. MacKerell Jr. and Nilesh K. Banavali. All-atom empirical force field for nucleic acids: II. Application to molecular dynamics simulations of DNA and RNA in solution. Journal of Computational Chemistry, 21(2):105–120, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Masella:2008:CPF

[MBC08] Michel Masella, Daniel Borgis, and Philippe Cuniasse. Combining a polarizable force-field and a coarse-grained polarizable solvent model: Application to long dynamics simulations of bovine pancreatic trypsin inhibitor. Journal of Computational Chemistry, 29(11):1707–1724, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Munoz:2002:HSB

[MBH^{+}02] J. Muñoz, X. Barril, B. Hernández, Modesto Orozco, and F. Javier Luque. Hydrophobic similarity between molecules: a MST-based hydrophobic similarity index. Journal of Computational Chemistry, 23(5):554–563, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mercero:2000:QMC

[MBL^{+}00] Jose M. Mercero, Paul Barrett, Cheuk W. Lam, Joseph E. Fowler, Jesus M. Ugalde, and Lee G. Pedersen. Quantum mechanical calculations on phosphate hydrolysis reactions. Journal of Computational Chemistry, 21(1):43–51, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marechal:2000:TMH

[MBM^{+}00] Jean-Didier Maréchal, Guada Barea, Feliu Maseras, Agustí Lledós, Liliane Mouawad, and David Pérahia. Theoretical modeling of the heme group with a hybrid QM/MM method. Journal of Computational Chemistry, 21(4):282–294, March 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Martin:2009:CMG

[MBP09] Jason Martin, Jon Baker, and Peter Pulay. Comments on the molecular geometry of ferrocene: the dangers of using quantum chemistry programs as black boxes. Journal of Computational Chemistry, 30(6):881–883, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mitin:2003:PSI

[MBWP03] Alexander V. Mitin, Jon Baker, Krzysztof Wolinski, and Peter Pulay. Parallel stored-integral and semidirect Hartree–Fock and DFT methods with data compression. Journal of Computational Chemistry, 24(2):154–160, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Moon:2006:CQC

[MC06] Seongho Moon and David A. Case. A comparison of quantum chemical models for calculating NMR shielding parameters in peptides: Mixed basis set and ONIOM methods combined with a complete basis set extrapolation. Journal of Computational Chemistry, 27(7):825–836, May 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

McDowell:2003:CSL

[McD03] Sean A. C. McDowell. A computational study of the LiH dimer. Journal of Computational Chemistry, 24(10):1201– 1207, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

McDowell:2008:CSS

[McD08] Sean A. C. McDowell. A comparative study of some redand blue-shifted linear H-bonded complexes of N_{2}. Journal of Computational Chemistry, 29(2):298–305, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marun:2005:FCP

[MCF05] Raúl A. Bustos Marún, Eduardo A. Coronado, and Juan C. Ferrero. Fitting complex potential energy surfaces to simple model potentials: Application of the simplex-annealing method. Journal of Computational Chemistry, 26(6):523–531, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marabotti:2007:NCS

[MCF07] Anna Marabotti, Giovanni Colonna, and Angelo Facchiano. New computational strategy to analyze the interactions of ERα and ERβ with different ERE sequences. Journal of Computational Chemistry, 28(6):1031–1041, April 30, 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Manojkumar:2005:TIM

[MCK05] T. K. Manojkumar, Chunzhi Cui, and Kwang S. Kim. Theoretical insights into the mechanism of acetylcholinesterasecatalyzed acylation of acetylcholine. Journal of Computational Chemistry, 26(6):606–611, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mckee:2007:MHE

[Mck07a] Michael L. Mckee. Modeling hydrogen evolution from the Fe_{4}S_{4} and Fe_{8}S_{9}X (X = N, C) clusters. Can a FeS high-spin cluster serve as a surrogate for the FeMo cofactor? Journal of Computational Chemistry, 28(11):1796–1808, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mckee:2007:MNF

[Mck07b] Michael L. Mckee. Modeling the nitrogenase FeMo cofactor with high-spin Fe_{8}S_{9}X^{+} (X = N, C) clusters. Is the first step for N_{2} reduction to NH_{3} a concerted dihydrogen transfer? Journal of Computational Chemistry, 28(8):1342–1356, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mongan:2004:CPM

[MCM04] John Mongan, David A. Case, and J. Andrew McCammon. Constant pH molecular dynamics in generalized Born implicit solvent. Journal of Computational Chemistry, 25(16):2038– 2048, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Masetti:2008:MHP

[MCR08] Matteo Masetti, Andrea Cavalli, and Maurizio Recanatini. Modeling the hERG potassium channel in a phospholipid bilayer: Molecular dynamics and drug docking studies. Journal of Computational Chemistry, 29(5):795–808, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Matsubara:2008:IEE

[MDA08] Toshiaki Matsubara, Michel Dupuis, and Misako Aida. An insight into the environmental effects of the pocket of the active site of the enzyme. Ab initio ONIOM-molecular dynamics (MD) study on cytosine deaminase. Journal of Computational Chemistry, 29(3):458–465, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Munoz:2004:PEB

[MDI04] D. Muñoz, C. De Graaf, and F. Illas. Putting error bars on the Ab Initio theoretical estimates of the magnetic coupling constants: the parent compounds of superconducting cuprates as a case study. Journal of Computational Chemistry, 25(10): 1234–1241, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Musch:2006:DMN

[ME06] Patrick Musch and Bernd Engels. DIESEL–MP2: a new program to perform large-scale multireference-MP2 computations. Journal of Computational Chemistry, 27(10):1055–1062, July 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meirovitch:2002:PCP

[Mei02] Hagai Meirovitch. Polymer collapse, protein folding, and the percolation threshold. Journal of Computational Chemistry, 23(1):166–171, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mackerell:2004:ETB

[MFB04] Alexander D. Mackerell Jr., Michael Feig, and Charles L. Brooks III. Extending the treatment of backbone energetics in protein force fields: Limitations of gas-phase quantum mechanics in reproducing protein conformational distributions in molecular dynamics simulations. Journal of Computational Chemistry, 25(11):1400–1415, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Moreira:2007:CAS

[MFR07] Irina S. Moreira, Pedro A. Fernandes, and Maria J. Ramos. Computational alanine scanning mutagenesis — an improved methodological approach. Journal of Computational Chemistry, 28(3):644–654, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mo:2000:IQM

[MG00] Yirong Mo and Jiali Gao. Ab initio QM/MM simulations with a molecular orbital-valence bond (MOVB) method: application to an S_{N}2 reaction in water. Journal of Computational Chemistry, 21(16):1458–1469, December 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ma:2006:LBK

[MG06] Hairong Ma and Martin Gruebele. Low barrier kinetics: Dependence on observables and free energy surface. Journal of Computational Chemistry, 27(2):125–134, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Malrieu:2007:EBE

[MGCA07] Jean-Paul Malrieu, Nathalie Guihéry, Carmen Jiménez Calzado, and Celestino Angeli. Essay: Bond electron pair: Its relevance and analysis from the quantum chemistry point of view. Journal of Computational Chemistry, 28(1):35–50, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Milet:2006:RNS

[MGG06] Anne Milet, Yves Gimbert, and Andrew E. Greene. Reaction of nitrones with silyl ketene acetals: a DFT study. Journal of Computational Chemistry, 27(2):157–162, January 30, 2006.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morales:2000:PTS

[MGJAARC00] L. B. Morales, R. Garduño-Juárez, J. M. Aguilar-Alvarado, and F. J. Riveros-Castro. A parallel tabu search for conformational energy optimization of oligopeptides. Journal of Computational Chemistry, 21(2):147–156, January 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Melendez:2000:IDI

[MGLDS00] F. J. Meléndez, B. Gallego-Luxan, Jean Demaison, and Yves G. Smeyers. Ab initio determination of the infrared phosphine torsion spectrum in vinylphosphine with ZPE correction. Journal of Computational Chemistry, 21(13):1167–1175, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Masgrau:2003:DRC

[MGLL03] Laura Masgrau, Àngels González-Lafont, and José M. Lluch. Dependence of the rate constants on the treatment of internal rotation modes: the reaction OH + CH_{3}SH → CH_{3}S + H_{2}O as an example. Journal of Computational Chemistry, 24(6): 701–706, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morreale:2003:CDC

[MGLO03] Antonio Morreale, Josep Lluis Gelpí, F. Javier Luque, and Modesto Orozco. Continuum and discrete calculation of fractional contributions to solvation free energy. Journal of Computational Chemistry, 24(13):1610–1623, October 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mandado:2007:CGT

[MGMM07a] Marcos Mandado, María J. González-Moa, and Ricardo A. Mosquera. Chemical graph theory and n-center electron delocalization indices: a study on polycyclic aromatic hydrocarbons. Journal of Computational Chemistry, 28(10):1625–1633, July 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mandado:2007:QCD

[MGMM07b] Marcos Mandado, María J. González-Moa, and Ricardo A. Mosquera. QTAIM n-center delocalization indices as descriptors of aromaticity in mono and poly heterocycles. Journal of Computational Chemistry, 28(1):127–136, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mirzaei:2008:CNS

[MH08a] Mahmoud Mirzaei and Nasser L. Hadipour. A computational NQR study on the hydrogen-bonded lattice of cytosine-5-acetic acid. Journal of Computational Chemistry, 29(5):832– 838, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Miyata:2008:CMD

[MH08b] Tatsuhiko Miyata and Fumio Hirata. Combination of molecular dynamics method and 3D-RISM theory for conformational sampling of large flexible molecules in solution. Journal of Computational Chemistry, 29(6):871–882, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meinke:2009:FED

[MH09] Jan H. Meinke and Ulrich H. E. Hansmann. Free-energy-driven folding and thermodynamics of the 67-residue protein GS-α_{3}W — a large-scale Monte Carlo study. Journal of Computational Chemistry, 30(11):1642–1648, August 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Merzel:2006:NFF

[MHJS06] Franci Merzel, Milan Hodošček, Dušanka Janežič, and Alain Sanson. New force field for calcium binding sites in annexin– membrane complexes. Journal of Computational Chemistry, 27(4):446–452, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morris:2009:AAA

[MHL^{+}09] Garrett M. Morris, Ruth Huey, William Lindstrom, Michel F. Sanner, Richard K. Belew, David S. Goodsell, and Arthur J. Olson. AutoDock4 and AutoDockTools4: Automated docking with selective receptor flexibility. Journal of Computational Chemistry, 30(16):2785–2791, December 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meyer:2005:ISP

[MHS05] Michael Meyer, Alexandre Hocquet, and Jürgen Sühnel. Interaction of sodium and potassium ions with sandwiched cytosine-, guanine-, thymine-, and uracil-base tetrads. Journal of Computational Chemistry, 26(4):352–364, March 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mohle:2001:DPP

[MHT01] Kerstin Möhle, Hans-Jörg Hofmann, and Walter Thiel. Description of peptide and protein secondary structures employing semiempirical methods. Journal of Computational Chemistry, 22(5):509–520, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Merlitz:2004:FAA

[MHW04] H. Merlitz, T. Herges, and W. Wenzel. Fluctuation analysis and accuracy of a large-scale in silico screen. Journal of Computational Chemistry, 25(13):1568–1575, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mitrasinovic:2001:QCP

[Mit01] Petar M. Mitrasinovic. Quantitative characterization of the P C bonds in ylides of phosphorus. Journal of Computational Chemistry, 22(13):1387–1395, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ji:2007:KMH

[mJlZsLyL07] Yue meng Ji, Xiao lei Zhao, Ze sheng Li, and Jing yao Liu. Kinetic mechanism of the hydrogen abstraction reactions of the chlorine atoms with CH_{3}CF_{2}Cl and CH_{3}CFCl_{2}: a dual level direct dynamics study. Journal of Computational Chemistry, 28(5):975–983, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ji:2008:TDS

[mJlZyL^{+}08] Yue meng Ji, Xiao lei Zhao, Jing yao Liu, Ying Wang, and Ze sheng Li. Theoretical dynamic studies on the reactions of CH_{3}C(O)CH_{3−n}Cl_{n} (n = 0–3) with the chlorine atom. Journal of Computational Chemistry, 29(5):809–819, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Massova:2002:PKM

[MK02] Irina Massova and Peter A. Kollman. p K_{a}, MM, and QM studies of mechanisms of β-lactamases and penicillin-binding proteins: Acylation step. Journal of Computational Chemistry, 23(16):1559–1576, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Makowski:2006:EAE

[MKGA06] Marcin Makowski, Jacek Korchowiec, Feng Long Gu, and Yuriko Aoki. Efficiency and accuracy of the elongation method as applied to the electronic structures of large systems. Journal of Computational Chemistry, 27(13):1603–1619, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mancera:2004:LPD

[MKT04] Ricardo L. Mancera, Per Källblad, and Nikolay P. Todorov. Ligand–protein docking using a quantum stochastic tunneling optimization method. Journal of Computational Chemistry, 25(6):858–864, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Moskaleva:2000:UID

[ML00] L. V. Moskaleva and M. C. Lin. Unimolecular isomerization/ decomposition of cyclopentadienyl and related bimolecular reverse process: ab initio MO/statistical theory study. Journal of Computational Chemistry, 21(6):415–425, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ma:2000:MPI

[MLA00] Buyong Ma, Jenn-Huei Lii, and Norman L. Allinger. Molecular polarizabilities and induced dipole moments in molecular mechanics. Journal of Computational Chemistry, 21(10): 813–825, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mok:2001:ICS

[MLCD01] Daniel K. W. Mok, Edmond P. F. Lee, Foo-Tim Chau, and John M. Dyke. Ab initio calculations on the and states of AlNC and simulation of the AlNC– emission spectra. Journal of Computational Chemistry, 22(16):1896–1906, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Martineau:2004:BFD

[MLG04] Eric Martineau, Pierre-Jean L’Heureux, and John R. Gunn. Biased fragment distribution in MC simulation of protein folding. Journal of Computational Chemistry, 25(15):1895–1903, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Molina:2003:IED

[MLJ03] Pablo A. Molina, Hui Li, and Jan H. Jensen. Intraprotein electrostatics derived from first principles: Divide-and-conquer approaches for QM/MM calculations. Journal of Computational Chemistry, 24(16):1971–1979, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ma:2006:MCI

[MLL06] Jing Ma, Shuhua Li, and Wei Li. A multireference configuration interaction method based on the separated electron pair wave functions. Journal of Computational Chemistry, 27(1): 39–47, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Milgram:2008:SIK

[MLL08a] R. J. Milgram, Guanfeng Liu, and J. C. Latombe. On the structure of the inverse kinematics map of a fragment of protein backbone. Journal of Computational Chemistry, 29(1): 50–68, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Min:2008:CBS

[MLL^{+}08b] Seung Kyu Min, Eun Cheol Lee, Han Myoung Lee, Dong Young Kim, Dongwook Kim, and Kwang S. Kim. Complete basis set limit of ab initio binding energies and geometrical parameters for various typical types of complexes. Journal of Computational Chemistry, 29(8):1208–1221, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mayer:2000:CDC

[MM00] Bernd Mayer and Giancarlo Marconi. Circular dichroic constrained structure optimization of homoalanine peptides. Journal of Computational Chemistry, 21(4):270–281, March 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mishra:2002:ITS

[MM02] S. K. Mishra and P. C. Mishra. An ab initio theoretical study of electronic structure and properties of 2’-deoxyguanosine in gas phase and aqueous media. Journal of Computational Chemistry, 23(5):530–540, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Martinez:2003:POA

[MM03] José Mario Martínez and Leandro Martínez. Packing optimization for automated generation of complex system’s initial configurations for molecular dynamics and docking. Journal of Computational Chemistry, 24(7):819–825, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Macias:2005:CII

[MM05] Alba T. Macias and Alexander D. MacKerell Jr. CH/π interactions involving aromatic amino acids: Refinement of the CHARMM tryptophan force field. Journal of Computational Chemistry, 26(14):1452–1463, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Milani:2007:GAD

[MM07] G. Milani and F. Milani. Genetic algorithm for the determination of binodal curves in ternary systems polymer– liquid(1)–liquid(2) and polymer(1)–polymer(2)–solvent. Journal of Computational Chemistry, 28(13):2203–2215, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mallik:2002:DED

[MML02] Buddhadeb Mallik, Artem Masunov, and Themis Lazaridis. Distance and exposure dependent effective dielectric function.

Journal of Computational Chemistry, 23(11):1090–1099, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marechal:2006:DSR

[MML^{+}06] Jean-Didier Maréchal, Feliu Maseras, Agustí Lledós, Liliane Mouawad, and David Perahia. A DFT study on the relative affinity for oxygen of the α and β subunits of hemoglobin. Journal of Computational Chemistry, 27(12): 1446–1453, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Makowska:2005:TCH

[MMLC05] Joanna Makowska, Mariusz Makowski, Adam Liwo, and Lech Chmurzyński. Theoretical calculations of homoconjugation equilibrium constants in systems modeling acid–base interactions in side chains of biomolecules using the potential of mean force. Journal of Computational Chemistry, 26(3):235–242, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mayaan:2007:CFF

[MMMY07] Evelyn Mayaan, Adam Moser, Alexander D. MacKerell Jr., and Darrin M. York. CHARMM force field parameters for simulation of reactive intermediates in native and thio-substituted ribozymes. Journal of Computational Chemistry, 28(2):495– 507, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Moad:2007:NVD

[MMP^{+}07] Andrew J. Moad, Charles W. Moad, John M. Perry, Ronald D. Wampler, G. Scott Goeken, Nathan J. Begue, Tian Shen, Randy Heiland, and Garth J. Simpson. NLOPredict: Visualization and data analysis software for nonlinear optics. Journal of Computational Chemistry, 28(12):1996–2002, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mannfors:2001:PEM

[MMPK01] B. Mannfors, N. G. Mirkin, K. Palmo, and S. Krimm. A polarizable electrostatic model of the N-methylacetamide dimer. Journal of Computational Chemistry, 22(16):1933–1943, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Merino:2007:RAP

[MMRVH07] Gabriel Merino, Miguel A. Méndez-Rojas, Alberto Vela, and Thomas Heine. Recent advances in planar tetracoordinate carbon chemistry. Journal of Computational Chemistry, 28(1): 362–372, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mark:2002:SDL

[MN02] Pekka Mark and Lennart Nilsson. Structure and dynamics of liquid water with different long-range interaction truncation and temperature control methods in molecular dynamics simulations. Journal of Computational Chemistry, 23(13): 1211–1219, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Murga:2001:NAC

[MO01] Leonel F. Murga and Mary Jo Ondrechen. Numerical aspects of the calculation of second hyperpolarizabilities using the finite field method coupled with a simple Lanczos algorithm. Journal of Computational Chemistry, 22(4):468–474, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Miao:2009:MDO

[MO09] Yinglong Miao and Peter J. Ortoleva. Molecular dynamics/ order parameter extrapolation for bionanosystem simulations. Journal of Computational Chemistry, 30(3):423–437, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Maul:2007:DSU

[MOP^{+}07] R. Maul, F. Ortmann, M. Preuss, K. Hannewald, and F. Bechstedt. DFT studies using supercells and projector-augmented waves for structure, energetics, and dynamics of glycine, alanine, and cysteine. Journal of Computational Chemistry, 28 (11):1817–1833, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morita:2002:WPC

[Mor02] Akihiro Morita. Water polarizability in condensed phase: ab initio evaluation by cluster approach. Journal of Computational Chemistry, 23(15):1466–1471, November 30, 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Malcolm:2003:ADI

[MP03a] Nathaniel O. J. Malcolm and Paul L. A. Popelier. An algorithm to delineate and integrate topological basins in a threedimensional quantum mechanical density function. Journal of Computational Chemistry, 24(10):1276–1282, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Malcolm:2003:IAL

[MP03b] Nathaniel O. J. Malcolm and Paul L. A. Popelier. An improved algorithm to locate critical points in a 3D scalar field as implemented in the program MORPHY. Journal of Computational Chemistry, 24(4):437–442, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Murphy:2000:MQM

[MPF00] R. B. Murphy, D. M. Philipp, and R. A. Friesner. A mixed quantum mechanics/molecular mechanics (QM/MM) method for large-scale modeling of chemistry in protein environments. Journal of Computational Chemistry, 21(16):1442–1457, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marsal:2002:VTV

[MR02] Philippe Marsal and Michel Roche. Variational treatment of the vibrational Hamiltonian for NH_{3} and H_{2}NO. Journal of Computational Chemistry, 23(5):541–547, April 15, 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Maurer:2004:MAA

[MR04] Richard I. Maurer and Christopher A. Reynolds. A multilayered approach to approximating solute polarization. Journal of Computational Chemistry, 25(5):627–631, April 15, 2004.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marti:2009:HQC

[MR09] Konrad H. Marti and Markus Reiher. Haptic quantum chemistry. Journal of Computational Chemistry, 30(13):2010–2020, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meagher:2003:DPP

[MRC03] Kristin L. Meagher, Luke T. Redman, and Heather A. Carlson. Development of polyphosphate parameters for use with the AMBER force field. Journal of Computational Chemistry, 24 (9):1016–1025, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mutyala:2007:CRB

[MRS^{+}07] Ravichandra Mutyala, R. N. Reddy, M. Sumakanth, P. Reddanna, and M. Rami Reddy. Calculation of relative binding affinities of fructose 1,6-bisphosphatase mutants with adenosine monophosphate using free energy perturbation method. Journal of Computational Chemistry, 28(5):932–937, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mitra:2009:QAP

[MRS09] Indrani Mitra, Kunal Roy, and Achintya Saha. QSAR of antilipid peroxidative activity of substituted benzodioxoles using chemometric tools. Journal of Computational Chemistry, 30 (16):2712–2722, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mitchell:2000:MSP

[MS00] Anthony S. Mitchell and Mark A. Spackman. Molecular surfaces from the promolecule: a comparison with Hartree–Fock ab initio electron density surfaces. Journal of Computational Chemistry, 21(11):933–942, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morita:2001:INI

[MS01] Shoji Morita and Shogo Sakai. IMiCMO: a new integrated ab initio multicenter molecular orbitals method for molecular dynamics calculations in solvent cluster systems. Journal of Computational Chemistry, 22(10):1107–1112, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Massi:2003:SDA

[MS03] Francesca Massi and John E. Straub. Structural and dynamical analysis of the hydration of the Alzheimer’s β-amyloid peptide. Journal of Computational Chemistry, 24(2):143–153, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Motoki:2004:NRS

[MS04] Takao Motoki and Akinobu Shiga. New reaction simulator “LUMMOX” and its application for prediction of catalytic activities. Journal of Computational Chemistry, 25(1):106– 111, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meyer:2001:DFS

[MSBS01] Michael Meyer, Thomas Steinke, Maria Brandl, and Jürgen Sühnel. Density functional study of guanine and uracil quartets and of guanine quartet/metal ion complexes. Journal of Computational Chemistry, 22(1):109–124, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Marabotti:2008:EBP

[MSF^{+}08] Anna Marabotti, Francesca Spyrakis, Angelo Facchiano, Pietro Cozzini, Saverio Alberti, Glen E. Kellogg, and Andrea Mozzarelli. Energy-based prediction of amino acid-nucleotide base recognition. Journal of Computational Chemistry, 29 (12):1955–1969, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Matsuura:2006:AEA

[MSH^{+}06a] Azuma Matsuura, Hiroyuki Sato, Hirohiko Houjou, Shino Saito, Tomohiko Hayashi, and Minoru Sakurai. Accurate evaluation of the absorption maxima of retinal proteins based on a hybrid QM/MM method. Journal of Computational Chemistry, 27(14):1623–1630, November 15, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

McNamara:2006:CSP

[MSH^{+}06b] Jonathan P. McNamara, Mahesh Sundararajan, Ian H. Hillier, Jun Ge, Andrew Campbell, and Claudio Morgado. Can the semiempirical PM3 scheme describe iron-containing bioinorganic molecules? Journal of Computational Chemistry, 27 (12):1307–1323, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

McQuaid:2004:DVC

[MSR04] Michael J. McQuaid, Huai Sun, and David Rigby. Development and validation of COMPASS force field parameters for molecules with aliphatic azide chains. Journal of Computational Chemistry, 25(1):61–71, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Monticelli:2008:MSM

[MST^{+}08] Luca Monticelli, Eric J. Sorin, D. Peter Tieleman, Vijay S. Pande, and Giorgio Colombo. Molecular simulation of multistate peptide dynamics: a comparison between microsecond timescale sampling and multiple shorter trajectories. Journal of Computational Chemistry, 29(11):1740–1752, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

MacCallum:2003:CWC

[MT03] Justin L. MacCallum and D. Peter Tieleman. Calculation of the water–cyclohexane transfer free energies of neutral amino acid side-chain analogs using the OPLS all-atom force field. Journal of Computational Chemistry, 24(15):1930–1935, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mediavilla:2009:MHP

[MTB09] C. Mediavilla, J. Tortajada, and V. G. Baonza. Modeling high pressure reactivity in unsaturated systems: Application to dimethylacetylene. Journal of Computational Chemistry, 30 (3):415–422, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Michel:2004:PVG

[MTE04] Julien Michel, Richard D. Taylor, and Jonathan W. Essex. The parameterization and validation of generalized Born models using the pairwise descreening approximation. Journal of Computational Chemistry, 25(14):1760–1770, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Muegge:2001:ELV

[Mue01] Ingo Muegge. Effect of ligand volume correction on PMF scoring. Journal of Computational Chemistry, 22(4):418–425, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Muino:2005:OCR

[Mui05] Pedro L. Muiño. The OH^{•} + CH_{3}SH reaction: Support for an addition-elimination mechanism from ab initio calculations. Journal of Computational Chemistry, 26(6):612–618, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Manninen:2006:SGB

[MV06] Pekka Manninen and Juha Vaara. Systematic Gaussian basisset limit using completeness-optimized primitive sets. A case for magnetic properties. Journal of Computational Chemistry, 27(4):434–445, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meineke:2005:OOO

[MVL^{+}05] Matthew A. Meineke, Charles F. Vardeman II, Teng Lin, Christopher J. Fennell, and J. Daniel Gezelter. OOPSE: an object-oriented parallel simulation engine for molecular dynamics. Journal of Computational Chemistry, 26(3):252–271, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Medvedev:2006:ATD

[MVLG06] N. N. Medvedev, V. P. Voloshin, V. A. Luchnikov, and M. L. Gavrilova. An algorithm for three-dimensional Voronoi Snetwork. Journal of Computational Chemistry, 27(14):1676– 1692, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Momany:2000:CSC

[MW00] F. A. Momany and J. L. Willett. Computational studies on carbohydrates: I. Density functional ab initio geometry optimization on maltose conformations. Journal of Computational Chemistry, 21(13):1204–1219, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mazur:2009:ADT

[MW09] Grzegorz Mazur and Radosław Włodarczyk. Application of the dressed time-dependent density functional theory for the excited states of linear polyenes. Journal of Computational Chemistry, 30(5):811–817, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Meller:2002:MFG

[MWE02] Jaroslaw Meller, Michael Wagner, and Ron Elber. Maximum feasibility guideline in the design and analysis of protein folding potentials. Journal of Computational Chemistry, 23(1): 111–118, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ma:2008:SOA

[MWL^{+}08] Jian-Yi Ma, Jing-Bo Wang, Xiang-Yuan Li, Yao Huang, Quan Zhu, and Ke-Xiang Fu. A study on orientation and absorption spectrum of interfacial molecules by using continuum model. Journal of Computational Chemistry, 29(2):198–210, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morita:2008:PCSb

[MY08a] Masato Morita and Satoshi Yabushita. Photoionization cross sections of H and H_{2} with complex Gaussian-type basis functions optimized for the frequency-dependent polarizabilities. Journal of Computational Chemistry, 29(14):2471–2478, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Morita:2008:PCSa

[MY08b] Masato Morita and Satoshi Yabushita. Photoionization cross sections with optimized orbital exponents within the complex basis function method. Journal of Computational Chemistry, 29(14):2317–2329, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Martin:2005:CDW

[MZ05] F. Martin and H. Zipse. Charge distribution in the water molecule — a comparison of methods. Journal of Computational Chemistry, 26(1):97–105, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mor:2008:SPI

[MZL08] Amit Mor, Guy Ziv, and Yaakov Levy. Simulations of proteins with inhomogeneous degrees of freedom: the effect of thermostats. Journal of Computational Chemistry, 29(12): 1992–1998, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakamura:2006:TPA

[NA06] Hideaki Nakamura and Masaru Aniya. Thermodynamic properties of Ag_{2}OB_{2}O_{3} glasses by a modified scale-transformed energy space sampling Monte Carlo method. Journal of Computational Chemistry, 27(4):471–477, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakamura:2002:MCS

[Nak02] Hideaki Nakamura. Monte Carlo sampling algorithm for searching a scale-transformed energy space of polypeptides. Journal of Computational Chemistry, 23(4):511–516, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakagawa:2007:PMP

[Nak07] Setsuko Nakagawa. Polarizable model potential function for nucleic acid bases. Journal of Computational Chemistry, 28 (9):1538–1550, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nifosi:2007:VSS

[NAT07] Riccardo Nifosí, Pietro Amat, and Valentina Tozzini. Variation of spectral, structural, and vibrational properties within the intrinsically fluorescent proteins family: a density functional study. Journal of Computational Chemistry, 28(14): 2366–2377, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nair:2004:MDI

[NBJ04] Nisanth N. Nair, Thomas Bredow, and Karl Jug. Molecular dynamics implementation in MSINDO: Study of silicon clusters. Journal of Computational Chemistry, 25(10):1255–1263, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakata:2004:ETS

[NBTN04a] A. Nakata, T. Baba, H. Takahashi, and H. Nakai. Erratum: Theoretical study on the excited states of psoralen compounds bonded to a thymine residue. Journal of Computational Chemistry, 25(3):309, February 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [NBTN04b].

Nakata:2004:TSE

[NBTN04b] A. Nakata, T. Baba, H. Takahashi, and H. Nakai. Theoretical study on the excited states of psoralen compounds bonded to a thymine residue. Journal of Computational Chemistry, 25(2): 179–188, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [NBTN04a].

Nanias:2005:PSP

[NCO^{+}05] Marian Nanias, Maurizio Chinchio, Stanisław Ołdziej, Cezary Czaplewski, and Harold A. Scheraga. Protein structure prediction with the UNRES force-field using Replica-Exchange Monte Carlo-with-Minimization; Comparison with MCM, CSA, and CFMC. Journal of Computational Chemistry, 26 (14):1472–1486, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Neese:2003:IRI

[Nee03] Frank Neese. An improvement of the resolution of the identity approximation for the formation of the Coulomb matrix. Journal of Computational Chemistry, 24(14):1740–1747, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Netzloff:2004:FFD

[NG04] Heather M. Netzloff and Mark S. Gordon. Fast fragments: the development of a parallel effective fragment potential method. Journal of Computational Chemistry, 25(15):1926– 1936, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nemukhin:2003:FEF

[NGTB03] A. V. Nemukhin, B. L. Grigorenko, I. A. Topol, and S. K. Burt. Flexible effective fragment QM/MM method: Validation through the challenging tests. Journal of Computational Chemistry, 24(12):1410–1420, September 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nguyen:2005:CEI

[NHH05] Thanh Ha Nguyen, David E. Hibbs, and Siân T. Howard. Conformations, energies, and intramolecular hydrogen bonds in dicarboxylic acids: Implications for the design of synthetic dicarboxylic acid receptors. Journal of Computational Chemistry, 26(12):1233–1241, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakashima:2006:RBF

[NHN06] Hiroyuki Nakashima, Jun-Ya Hasegawa, and Hiroshi Nakatsuji. On the reversible O_{2} binding of the Fe–porphyrin complex. Journal of Computational Chemistry, 27(4):426–433, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nilsson:2009:ETL

[Nil09] Lennart Nilsson. Efficient table lookup without inverse square roots for calculation of pair wise atomic interactions in classical simulations. Journal of Computational Chemistry, 30(9): 1490–1498, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nantasenamat:2007:PGS

[NINAT^{+}07] Chanin Nantasenamat, Chartchalerm Isarankura-Na-Ayudhya, Natta Tansila, Thanakorn Naenna, and Virapong Prachayasittikul. Prediction of GFP spectral properties using artificial neural network. Journal of Computational Chemistry, 28(7): 1275–1289, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Naidoo:2001:WSA

[NK01] Kevin J. Naidoo and Michelle Kuttel. Water structure about the dimer and hexamer repeat units of amylose from molecular dynamics computer simulations. Journal of Computational Chemistry, 22(4):445–456, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakai:2006:EDA

[NK06] Hiromi Nakai and Yasuaki Kikuchi. Energy density analysis of cluster size dependence of surface–molecule interactions (II): Formate adsorption onto a Cu(111) surface. Journal of Computational Chemistry, 27(8):917–925, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakamura:2002:NMP

[NKIS02] Shugo Nakamura, Daisuke Kyono, Mitsunori Ikeguchi, and Kentaro Shimizu. New method for parallel computation of Hessian matrix of conformational energy function in internal coordinates. Journal of Computational Chemistry, 23(4):463– 469, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nikitin:2007:NSS

[NL07] Alexei M. Nikitin and Alexander P. Lyubartsev. New six-site acetonitrile model for simulations of liquid acetonitrile and its aqueous mixtures. Journal of Computational Chemistry, 28 (12):2020–2026, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Neihsial:2008:NHB

[NL08] Siamkhanthang Neihsial and Richard H. Duncan Lyngdoh. Novel H-bonded base dimers as repeat units for informationbearing self-associative duplexes: a B3LYP/6-31G* search. Journal of Computational Chemistry, 29(11):1788–1797, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Niu:2009:HPC

[NLL^{+}09] Bing Niu, Lin Lu, Liang Liu, Tian Hong Gu, Kai-Yan Feng, Wen-Cong Lu, and Yu-Dong Cai. HIV-1 protease cleavage site prediction based on amino acid property. Journal of Computational Chemistry, 30(1):33–39, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nobeli:2001:EKB

[NMAT01] Irene Nobeli, John B. O. Mitchell, Alexander Alex, and Janet M. Thornton. Evaluation of a knowledge-based potential of mean force for scoring docked protein–ligand complexes. Journal of Computational Chemistry, 22(7):673–688, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Editorship:2004:A

[Nor04] North American Editorship. Announcement. Journal of Computational Chemistry, 25(1):i, January 15, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Neugebauer:2004:VCL

[NR04] Johannes Neugebauer and Markus Reiher. Vibrational center– ligand couplings in transition metal complexes. Journal of Computational Chemistry, 25(4):587–597, March 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Neugebauer:2002:QCC

[NRKH02] Johannes Neugebauer, Markus Reiher, Carsten Kind, and Bernd A. Hess. Quantum chemical calculation of vibrational spectra of large molecules — Raman and IR spectra for Buckminsterfullerene. Journal of Computational Chemistry, 23(9): 895–910, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Norberg:2008:MRR

[NSB08] Daniel Norberg and Nessima Salhi-Benachenhou. McLafferty rearrangement of the radical cations of butanal and 3fluorobutanal: a theoretical investigation of the concerted and stepwise mechanisms. Journal of Computational Chemistry, 29(3):392–406, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Noro:2007:RCB

[NSO^{+}07] Takeshi Noro, Masahiro Sekiya, You Osanai, Toshikatsu Koga, and Hisashi Matsuyama. Relativistic correlating basis sets for actinide atoms from _{90}Th to _{103}Lr. Journal of Computational Chemistry, 28(16):2511–2516, December 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Novoselov:2002:CST

[NSU^{+}02] Konstantin P. Novoselov, Denis B. Shirabaikin, Stanislav Ya. Umanskii, Alexander S. Vladimirov, Airat Kh. Minushev, and Anatoli A. Korkin. CHIMERA: a software tool for reaction rate calculations and kinetics and thermodynamics analysis. Journal of Computational Chemistry, 23(14):1375–1389, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakagawa:2000:TSC

[NTH00] Yoshinao Nakagawa, Nobuo Tajima, and Kimihiko Hirao. A theoretical study of catalytic hydration reactions of ethylene. Journal of Computational Chemistry, 21(14):1292–1304, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakata:2009:MRS

[NTH09] Ayako Nakata, Takao Tsuneda, and Kimihiko Hirao. Modified regional self-interaction corrected time-dependent density functional theory for core excited-state calculations. Journal of Computational Chemistry, 30(16):2583–2593, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakano:2002:QDP

[NUH02] Haruyuki Nakano, Ryuma Uchiyama, and Kimihiko Hirao. Quasi-degenerate perturbation theory with general multiconfiguration self-consistent field reference functions. Journal of Computational Chemistry, 23(12):1166–1175, September 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nye:2007:EWT

[Nye07] Mary Jo Nye. Essay: Working tools for theoretical chemistry: Polanyi, Eyring, and debates over the “semiempirical method”. Journal of Computational Chemistry, 28(1):98–108, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakajima:2002:RES

[NYH02] Takahito Nakajima, Takeshi Yanai, and Kimihiko Hirao. Relativistic electronic structure theory. Journal of Computational Chemistry, 23(8):847–860, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakashima:2006:BFP

[NyHN06] Hiroyuki Nakashima, Jun ya Hasegawa, and Hiroshi Nakatsuji. On the O_{2} binding of Fe–porphyrin, Fe–porphycene, and Fe– corrphycene complexes. Journal of Computational Chemistry, 27(12):1363–1372, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Nakamura:2009:RKD

[NYK^{+}09] Takashi Nakamura, Azusa Yamaguchi, Hirotaka Kondo, Hirofumi Watanabe, Tatsuo Kurihara, Nobuyoshi Esaki, Shuichi Hirono, and Shigenori Tanaka. Roles of K151 and D180 in L-2-haloacid dehalogenase from Pseudomonas sp. YL: Analysis by molecular dynamics and ab initio fragment molecular orbital calculations. Journal of Computational Chemistry, 30 (16):2625–2634, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Narumi:2009:CPG

[NYTH09] Tetsu Narumi, Kenji Yasuoka, Makoto Taiji, and Siegfried Höfinger. Current performance gains from utilizing the GPU or the ASIC MDGRAPE-3 within an enhanced Poisson Boltzmann approach. Journal of Computational Chemistry, 30(14): 2351–2357, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oakes:2005:RSP

[OBBS05] Roma E. Oakes, Steven E. J. Bell, Zuzana Benkova, and Andrzej J. Sadlej. Reduced-size polarized basis sets for calculations of molecular electric properties. II. Simulation of the Raman spectra. Journal of Computational Chemistry, 26(2): 154–159, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Otte:2009:FFP

[OBT09] Nikolaj Otte, Marco Bocola, and Walter Thiel. Force-field parameters for the simulation of tetrahedral intermediates of serine hydrolases. Journal of Computational Chemistry, 30(1): 154–162, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Onufriev:2002:EBR

[OCB02] Alexey Onufriev, David A. Case, and Donald Bashford. Effective Born radii in the generalized Born approximation: the importance of being perfect. Journal of Computational Chemistry, 23(14):1297–1304, November 15, 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Orzechowski:2002:TCC

[OCP02] Marek Orzechowski, Piotr Cieplak, and Lucjan Piela. Theoretical calculation of the coiled-coil stability in water in the context of its possible use as a molecular rack. Journal of Computational Chemistry, 23(1):106–110, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oberhofer:2009:EEF

[OD09] Harald Oberhofer and Christoph Dellago. Efficient extraction of free energy profiles from nonequilibrium experiments. Journal of Computational Chemistry, 30(11):1726–1736, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Olson:2008:PPL

[OFB08] Mark A. Olson, Michael Feig, and Charles L. Brooks III. Prediction of protein loop conformations using multiscale modeling methods with physical energy scoring functions. Journal of Computational Chemistry, 29(5):820–831, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Omori:2009:LRT

[OFIK09] Satoshi Omori, Sotaro Fuchigami, Mitsunori Ikeguchi, and Akinori Kidera. Linear response theory in dihedral angle space for protein structural change upon ligand binding. Journal of Computational Chemistry, 30(16):2602–2608, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oakley:2005:LMP

[OGH05] Mark T. Oakley, Jonathan M. Garibaldi, and Jonathan D. Hirst. Lattice models of peptide aggregation: Evaluation of conformational search algorithms. Journal of Computational Chemistry, 26(15):1638–1646, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Otsuka:2002:TAE

[OKE^{+}02] Takao Otsuka, Seiji Koizumi, Kazunaka Endo, Hiroyuki Kawabe, and Delano P. Chong. Theoretical Auger electron spectra of polymers by density functional theory calculations using model dimers. Journal of Computational Chemistry, 23 (3):394–401, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ono:2002:CFF

[OKH^{+}02] Satoshi Ono, Masataka Kuroda, Junichi Higo, Nobuyuki Nakajima, and Haruki Nakamura. Calibration of force-field dependency in free energy landscapes of peptide conformations by quantum chemical calculations. Journal of Computational Chemistry, 23(4):470–476, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ozkan:2004:CSP

[OM04] S. Banu Ozkan and Hagai Meirovitch. Conformational search of peptides and proteins: Monte Carlo minimization with an adaptive bias method applied to the heptapeptide deltorphin. Journal of Computational Chemistry, 25(4):565–572, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

ODonohue:2000:PNT

[OML^{+}00] M. F. O’Donohue, E. Minasian, S. J. Leach, A. W. Burgess, and H. R. Treutlein. PEPCAT — a new tool for conformational analysis of peptides. Journal of Computational Chemistry, 21(6):446–461, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ode:2008:FFP

[OMNH08] Hirotaka Ode, Yuri Matsuo, Saburo Neya, and Tyuji Hoshino. Force field parameters for rotation around χ torsion axis in nucleic acids. Journal of Computational Chemistry, 29(15):

2531–2542, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Otsuka:2007:WTA

[ON07] Takao Otsuka and Hiromi Nakai. Wavelet transform analysis of ab initio molecular dynamics simulation: Application to core-excitation dynamics of BF_{3}. Journal of Computational Chemistry, 28(6):1137–1144, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ono:2000:PFE

[ONHN00] Satoshi Ono, Nobuyuki Nakajima, Junichi Higo, and Haruki Nakamura. Peptide free-energy profile is strongly dependent on the force field: Comparison of C96 and AMBER95. Journal of Computational Chemistry, 21(9):748–762, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ohta:2004:IBC

[OO04] Yasuhito Ohta and Koji Ohta. Interconversion behavior of the C H bond in the CH radical cation: ab initio molecular dynamics study. Journal of Computational Chemistry, 25(15): 1910–1919, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Okumura:2006:MME

[OO06] Hisashi Okumura and Yuko Okamoto. Multibaric–multithermal ensemble molecular dynamics simulations. Journal of Computational Chemistry, 27(3):379–395, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ozawa:2008:CHB

[OO08] Tomonaga Ozawa and Kosuke Okazaki. CH/π hydrogen bonds determine the selectivity of the Src homology 2 domain to tyrosine phosphotyrosyl peptides: an ab initio fragment molecular orbital study. Journal of Computational Chemistry, 29 (16):2656–2666, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ohtsuka:2001:ETT

[OON01] Yuhki Ohtsuka, Kazufumi Ohkawa, and Hiroshi Nakatsuji. Electron transfer in the c-type cytochrome subunit of the photosynthetic reaction center of Rhodopseudomonas viridis: ab initio theoretical study. Journal of Computational Chemistry, 22(5):521–527, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oostenbrink:2009:EFE

[Oos09] Chris Oostenbrink. Efficient free energy calculations on small molecule host-guest systems — a combined linear interaction energy/one-step perturbation approach. Journal of Computational Chemistry, 30(2):212–221, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Olano:2005:FCN

[OR05] L. Renee Olano and Steven W. Rick. Fluctuating charge normal modes: an algorithm for implementing molecular dynamics simulations with polarizable potentials. Journal of Computational Chemistry, 26(7):699–707, May 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ohno:2006:LSM

[OS06] Kazuki Ohno and Minoru Sakurai. Linear-scaling molecular orbital calculations for the pKa values of ionizable residues in proteins. Journal of Computational Chemistry, 27(7):906– 916, May 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oncak:2008:ESP

[OS08] Milan Ončák and Martin Srnec. Electronic structure and physical properties of M_{i} X_{i} clusters (M = B, Al; X = N, P; i = 1, 2, 3): ab initio study. Journal of Computational Chemistry, 29(2):233–246, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oliva:2006:CSL

[OSA06] Josep M. Oliva and Luis Serrano-Andrés. A computational study of the lowest singlet and triplet states of neutral and dianionic 1,2-substituted icosahedral and octahedral ocarboranes. Journal of Computational Chemistry, 27(4):524– 535, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Okur:2003:UPC

[OSHS03] Asim Okur, Bentley Strockbine, Viktor Hornak, and Carlos Simmerling. Using PC clusters to evaluate the transferability of molecular mechanics force fields for proteins. Journal of Computational Chemistry, 24(1):21–31, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

OBoyle:2008:CLP

[OTL08] Noel M. O’Boyle, Adam L. Tenderholt, and Karol M. Langner. cclib: a library for package-independent computational chemistry algorithms. Journal of Computational Chemistry, 29(5):839–845, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oostenbrink:2003:SSP

[OV03] Chris Oostenbrink and Wilfred F. Van Gunsteren. Singlestep perturbations to calculate free energy differences from unphysical reference states: Limits on size, flexibility, and character. Journal of Computational Chemistry, 24(14):1730– 1739, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oostenbrink:2004:BFF

[OVMV04] Chris Oostenbrink, Alessandra Villa, Alan E. Mark, and Wilfred F. Van Gunsteren. A biomolecular force field based on the free enthalpy of hydration and solvation: the GROMOS force-field parameter sets 53A5 and 53A6. Journal of Computational Chemistry, 25(13):1656–1676, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Owens:2005:SCE

[Owe05] Frank J. Owens. Stability of (C_{60})_{2} and epoxide dimers, (C_{60})_{2}O_{N}, and their anions. Journal of Computational Chemistry, 26(8):803–806, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Okumoto:2001:CSB

[OY01] Satoshi Okumoto and Shinichi Yamabe. A computational study of base-catalyzed reactions between isocyanates and epoxides affording 2-oxazolidones and isocyanurates. Journal of Computational Chemistry, 22(3):316–326, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Okumoto:2003:CSE

[OY03] Satoshi Okumoto and Shinichi Yamabe. Computational study of epoxy-amine reactions. Journal of Computational Chemistry, 24(2):244–253, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oda:2005:NAF

[OYH05] Akifumi Oda, Noriyuki Yamaotsu, and Shuichi Hirono. New AMBER force field parameters of heme iron for cytochrome P450s determined by quantum chemical calculations of simplified models. Journal of Computational Chemistry, 26(8): 818–826, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Oda:2009:ESA

[OYH09] Akifumi Oda, Noriyuki Yamaotsu, and Shuichi Hirono. Evaluation of the searching abilities of HBOP and HBSITE for binding pocket detection. Journal of Computational Chemistry, 30(16):2728–2737, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Okamoto:2009:CTC

[OYK^{+}09] Hajime Okamoto, Tetsuo Yamada, Shuichiro Kihara, Kazumasa Takechi, Hiroyuki Takagi, and Kyozaburo Takeda. Conformational transitions of cyclic D, L-peptides. Journal of Computational Chemistry, 30(6):962–973, April 30, 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Polo:2005:JSB

[PA05] Victor Polo and Juan Andrés. A joint study based on the electron localization function and catastrophe theory of the chameleonic and centauric models for the Cope rearrangement of 1,5-hexadiene and its cyano derivatives. Journal of Computational Chemistry, 26(14):1427–1437, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Petrella:2003:IMN

[PABK03] Robert J. Petrella, Ioan Andricioaei, Bernard R. Brooks, and Martin Karplus. An improved method for nonbonded list generation: Rapid determination of near-neighbor pairs. Journal of Computational Chemistry, 24(2):222–231, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pacios:2006:CSP

[Pac06] Luis F. Pacios. Computational study of the process of hydrogen bond breaking: the case of the formamide–formic acid complex. Journal of Computational Chemistry, 27(14):1641– 1649, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Panczyk:2007:CIG

[Pan07] Tomasz Panczyk. Collisions of ideal gas molecules with a rough/fractal surface. A computational study. Journal of Computational Chemistry, 28(3):681–688, February 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Polo:2007:NIB

[PAS07] Victor Polo, Juan Andres, and Bernard Silvi. New insights on the bridge carbon–carbon bond in propellanes: a theoretical study based on the analysis of the electron localization function. Journal of Computational Chemistry, 28(5):857–864, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Patil:2008:GCC

[PAS08] Sunil Patil, R. C. Aiyer, and K. C. Sharma. Globally convergent computation of chemical equilibrium composition. Journal of Computational Chemistry, 29(7):1032–1036, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Periole:2009:PFE

[PAT^{+}09] Xavier Periole, Lucy R. Allen, Kamil Tamiola, Alan E. Mark, and Emanuele Paci. Probing the free energy landscape of the FBP28WW domain using multiple techniques. Journal of Computational Chemistry, 30(7):1059–1068, May 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Price:2002:MPF

[PB02] Daniel J. Price and Charles L. Brooks III. Modern protein force fields behave comparably in molecular dynamics simulations. Journal of Computational Chemistry, 23(11):1045–1057, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Patel:2004:CFCa

[PB04] Sandeep Patel and Charles L. Brooks III. CHARMM fluctuating charge force field for proteins: I parameterization and application to bulk organic liquid simulations. Journal of Computational Chemistry, 25(1):1–16, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Price:2005:DCB

[PB05] Daniel J. Price and Charles L. Brooks III. Detailed considerations for a balanced and broadly applicable force field: a study of substituted benzenes modeled with OPLS-AA. Journal of Computational Chemistry, 26(14):1529–1541, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Page:2006:CMP

[PB06] Christopher S. Page and Paul A. Bates. Can MM-PBSA calculations predict the specificities of protein kinase inhibitors? Journal of Computational Chemistry, 27(16):1990–2007, December 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pendas:2007:CFR

[PBF07] A. Martín Pendás, M. A. Blanco, and E. Francisco. Chemical fragments in real space: Definitions, properties, and energetic decompositions. Journal of Computational Chemistry, 28(1): 161–184, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pendas:2009:SRR

[PBF09] A. Martín Pendás, M. A. Blanco, and E. Francisco. Steric repulsions, rotation barriers, and stereoelectronic effects: a real space perspective. Journal of Computational Chemistry, 30(1):98–109, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Phillips:2005:SMD

[PBW^{+}05] James C. Phillips, Rosemary Braun, Wei Wang, James Gumbart, Emad Tajkhorshid, Elizabeth Villa, Christophe Chipot, Robert D. Skeel, Laxmikant Kalé, and Klaus Schulten. Scalable molecular dynamics with NAMD. Journal of Computational Chemistry, 26(16):1781–1802, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Persson:2000:ICS

[PBZ00] Petter Persson, Johnny Bustad, and Michael C. Zerner. INDO calculations of small copper clusters and CO adsorbed on copper(100) surfaces. Journal of Computational Chemistry, 21 (14):1221–1228, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perczel:2000:TDD

[PC00] András Perczel and Attila G. Császár. Toward direct determination of conformations of protein building units from multidimensional NMR experiments I. A theoretical case study of For-Gly-NH_{2} and For-L-Ala-NH_{2}. Journal of Computational Chemistry, 21(10):882–900, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ponec:2005:EPC

[PC05] Robert Ponec and Joaquin Chaves. Electron pairing and chemical bonds. Electron fluctuation and pair localization in ELF domains. Journal of Computational Chemistry, 26(12):1205– 1213, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ponec:2007:EPC

[PC07] Robert Ponec and Joaquin Chaves. Electron pairing and chemical bonds: Pair localization in ELF domains from the analysis of domain averaged Fermi holes. Journal of Computational Chemistry, 28(1):109–116, January 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pinsky:2008:SOM

[PCA^{+}08] Mark Pinsky, David Casanova, Pere Alemany, Santiago Alvarez, David Avnir, Chaim Dryzun, Ziv Kizner, and Alexander Sterkin. Symmetry operation measures. Journal of Computational Chemistry, 29(2):190–197, January 30, 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pacios:2009:STB

[PCMG09] Luis F. Pacios, Víctor M. Campos, Irene Merino, and Luis Gómez. Structures and thermodynamics of biphenyl dihydrodiol stereoisomers and their metabolites in the enzymatic degradation of arene xenobiotics. Journal of Computational Chemistry, 30(15):2420–2432, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Palermo:2007:EAB

[PCO^{+}07a] Nicholas Y. Palermo, József Csontos, Michael C. Owen, R. F. Murphy, and Sándor Lovas. Erratum: Aromatic-backbone interactions in model α-helical peptides. Journal of Computational Chemistry, 28(15):2510, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [PCO^{+}07b].

Palermo:2007:ABI

[PCO^{+}07b] Nicholas Y. Palermo, József Csontos, Michael C. Owen, Richard F. Murphy, and Sándor Lovas. Aromatic-backbone interactions in model α-helical peptides. Journal of Computational Chemistry, 28(7):1208–1214, May 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [?].

Pelmenschikov:2004:CRR

[PCS04] Vladimir Pelmenschikov, Kyung-Bin Cho, and Per E. M. Siegbahn. Class I ribonucleotide reductase revisited: the effect of removing a proton on Glu441. Journal of Computational Chemistry, 25(3):311–321, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pinsky:2008:AMC

[PDC^{+}08] Mark Pinsky, Chaim Dryzun, David Casanova, Pere Alemany, and David Avnir. Analytical methods for calculating Continuous Symmetry Measures and the Chirality Measure. Journal of Computational Chemistry, 29(16):2712–2721, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perera:2002:PSS

[PDP02] Lalith Perera, Thomas A. Darden, and Lee G. Pedersen. Predicted solution structure of zymogen human coagulation FVII. Journal of Computational Chemistry, 23(1):35–47, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Poater:2001:PBL

[PDS01] Jordi Poater, Miquel Duran, and Miquel Solà. Parametrization of the Becke3-LYP hybrid functional for a series of small molecules using quantum molecular similarity techniques. Journal of Computational Chemistry, 22(14):1666– 1678, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Penotti:2006:OOC

[Pen06] Fabio E. Penotti. Orbital-orthogonality constraints and basisset optimization. Journal of Computational Chemistry, 27(6): 762–772, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pristovsek:2006:SAP

[PF06] Primož Pristovšek and Lorella Franzoni. Stereospecific assignments of protein NMR resonances based on the tertiary structure and 2D/3D NOE data. Journal of Computational Chemistry, 27(6):791–797, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pendas:2005:TEI

[PFB05] A. Martín Pendás, E. Francisco, and M. A. Blanco. Twoelectron integrations in the Quantum Theory of Atoms in Molecules with correlated wave functions. Journal of Computational Chemistry, 26(4):344–351, March 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perczel:2003:TDD

[PFC03] András Perczel, Anna K. Füzéry, and Attila G. Császár. Toward direct determination of conformations of protein building units from multidimensional NMR experiments. V. NMR chemical shielding analysis of N-formyl-serinamide, a model for polar side-chain containing peptides. Journal of Computational Chemistry, 24(10):1157–1171, July 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perczel:2003:PMX

[PFJ^{+}03] András Perczel, Ödön Farkas, Imre Jákli, Igor A. Topol, and Imre G. Csizmadia. Peptide models. XXXIII. Extrapolation of low-level Hartree–Fock data of peptide conformation to large basis set SCF, MP2, DFT, and CCSD(T) results. The Ramachandran surface of alanine dipeptide computed at various levels of theory. Journal of Computational Chemistry, 24(9): 1026–1042, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pereira:2004:MRR

[PFR04a] Susana Pereira, Pedro Alexandrino Fernandes, and Maria João Ramos. Mechanism for ribonucleotide reductase inactivation by the anticancer drug gemcitabine. Journal of Computational Chemistry, 25(10):1286–1294, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pereira:2004:TSR

[PFR04b] Susana Pereira, Pedro Alexandrino Fernandes, and Maria João Ramos. Theoretical study of ribonucleotide reductase mechanism-based inhibition by 2’-azido-2’-deoxyribonucleoside 5’-diphosphates. Journal of Computational Chemistry, 25(2): 227–237, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pacios:2001:III

[PG01] L. F. Pacios and P. C. Gómez. Intramolecular interactions and intramolecular hydrogen bonding in conformers of gaseous glycine. Journal of Computational Chemistry, 22(7):702–716, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Piacenza:2004:SQC

[PG04] M. Piacenza and S. Grimme. Systematic quantum chemical study of DNA-base tautomers. Journal of Computational Chemistry, 25(1):83–99, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perczel:2005:SSP

[PGC05] András Perczel, Zoltán Gáspári, and Imre G. Csizmadia. Structure and stability of β-pleated sheets. Journal of Computational Chemistry, 26(11):1155–1168, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pacios:2006:VAC

[PGG06] Luis F. Pacios, Pedro C. Gómez, and Oscar Gálvez. Variation of atomic charges on proton transfer in strong hydrogen bonds: the case of anionic and neutral imidazole–acetate complexes. Journal of Computational Chemistry, 27(14):1650– 1661, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pettersen:2004:UCV

[PGH^{+}04] Eric F. Pettersen, Thomas D. Goddard, Conrad C. Huang, Gregory S. Couch, Daniel M. Greenblatt, Elaine C. Meng, and Thomas E. Ferrin. UCSF Chimera — a visualization system for exploratory research and analysis. Journal of Computational Chemistry, 25(13):1605–1612, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Peiro-Garcia:2003:ISM

[PGNG03] Julio Peiró-García and Ignacio Nebot-Gil. Ab initio study of the mechanism of the atmospheric reaction: NO_{2} + O_{3} → NO_{3} + O_{2}. Journal of Computational Chemistry, 24(13): 1657–1663, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Peiro-Garcia:2003:TIS

[PGRRNG03] Julio Peiró-García, Víctor M. Ramírez-Ramírez, and Ignacio Nebot-Gil. A theoretical ab initio study on the H_{2}NO + O_{3} reaction. Journal of Computational Chemistry, 24(11): 1321–1328, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perczel:2004:SIC

[PHFC04] András Perczel, Péter Hudáky, Anna K. Füzéry, and Imre G. Csizmadia. Stability issues of covalently and noncovalently bonded peptide subunits. Journal of Computational Chemistry, 25(8):1084–1100, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Praprotnik:2008:NAA

[PHH^{+}08] Matej Praprotnik, Stanko Hočevar, Milan Hodošček, Matej Penca, and Dušanka Janežič. New all-atom force field for molecular dynamics simulation of an AlPO_{4}-34 molecular sieve. Journal of Computational Chemistry, 29(1):122–129, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Partay:2008:NMD

[PHJ^{+}08] Lívia B. Pártay, György Hantal, Pál Jedlovszky, Árpád Vincze, and George Horvai. A new method for determining the interfacial molecules and characterizing the surface roughness in computer simulations. Application to the liquid–vapor interface of water. Journal of Computational Chemistry, 29(6): 945–956, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Piacenza:2007:CQC

[PHKG07] M. Piacenza, I. Hyla-Kryspin, and S. Grimme. A comparative quantum chemical study of the ruthenium catalyzed olefin metathesis. Journal of Computational Chemistry, 28 (14):2275–2285, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Parsons:2005:PCT

[PHR^{+}05] Jerod Parsons, J. Bradley Holmes, J. Maurice Rojas, Jerry Tsai, and Charlie E. M. Strauss. Practical conversion from torsion space to Cartesian space for in silico protein synthesis. Journal of Computational Chemistry, 26(10):1063–1068, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pribil:2008:SDP

[PHRR08] Andreas B. Pribil, Thomas S. Hofer, Bernhard R. Randolf, and Bernd M. Rode. Structure and dynamics of phosphate ion in aqueous solution: an ab initio QMCF MD study. Journal of Computational Chemistry, 29(14):2330–2334, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pinak:2001:MDS

[Pin01] Miroslav Pinak. Molecular dynamics simulation of thymine glycol-lesioned DNA reveals specific hydration at the lesion. Journal of Computational Chemistry, 22(15):1723–1731, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pinak:2003:OLB

[Pin03] Miroslav Pinak. 8-oxoguanine lesioned B-DNA molecule complexed with repair enzyme hOGG1: a molecular dynamics study. Journal of Computational Chemistry, 24(7):898–907, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Porwal:2007:PSP

[PJB^{+}07] Gaurav Porwal, Swapnil Jain, S. Dhilly Babu, Deepak Singh, Hemant Nanavati, and Santosh Noronha. Protein structure prediction aided by geometrical and probabilistic constraints. Journal of Computational Chemistry, 28(12):1943– 1952, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Perez-Jimenez:2007:MMW

[PJPJdPRMI07] Ángel J. Pérez-Jiménez, José M. Pérez-Jordá, Ibério de P. R. Moreira, and Francesc Illas. Merging multiconfigurational wavefunctions and correlation functionals to predict magnetic coupling constants. Journal of Computational Chemistry, 28(16): 2559–2568, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Patra:2004:SCF

[PK04] Michael Patra and Mikko Karttunen. Systematic comparison of force fields for microscopic simulations of NaCl in aqueous solutions: Diffusion, free energy of hydration, and structural properties. Journal of Computational Chemistry, 25(5):678– 689, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Petrella:2005:EEF

[PK05] Robert J. Petrella and Martin Karplus. Electrostatic energies and forces computed without explicit interparticle interactions: a linear time complexity formulation. Journal of Computational Chemistry, 26(8):755–787, June 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ponec:2008:SBB

[PLC08] Robert Ponec, György Lendvay, and Joaquin Chaves. Structure and bonding in binuclear metal carbonyls from the analysis of domain averaged Fermi holes. I. Fe_{2} (CO)_{9} and Co_{2} (CO)_{8}. Journal of Computational Chemistry, 29(9):1387– 1398, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

PereiraDaSilva:2002:QBP

[PM02] Andre Elvas Pereira Da Silva and Geraldo Magela E. Silva. Quantum bits with polyacetylene. Journal of Computational Chemistry, 23(9):870–873, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Patel:2004:CFCb

[PMB04] Sandeep Patel, Alexander D. Mackerell Jr., and Charles L. Brooks III. CHARMM fluctuating charge force field for proteins: II. Protein/solvent properties from molecular dynamics simulations using a nonadditive electrostatic model. Journal of Computational Chemistry, 25(12):1504–1514, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pisani:2008:PLM

[PMC^{+}08] Cesare Pisani, Lorenzo Maschio, Silvia Casassa, Migen Halo, Martin Schütz, and Denis Usvyat. Periodic local MP2 method for the study of electronic correlation in crystals: Theory and preliminary applications. Journal of Computational Chemistry, 29(13):2113–2124, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pappalardo:2003:FEP

[PMGL03] Matteo Pappalardo, Danilo Milardi, Domenico M. Grasso, and Carmelo La Rosa. Free energy perturbation and molecular dynamics calculations of copper binding to azurin. Journal of Computational Chemistry, 24(6):779–785, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Politzer:2003:EPC

[PML03] Peter Politzer, Jane S. Murray, and Pat Lane. Electrostatic potentials and covalent radii. Journal of Computational Chemistry, 24(4):505–511, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prabhakar:2005:CSV

[PMM05] Rajeev Prabhakar, Keiji Morokuma, and Djamaladdin G. Musaev. A comparative study of various computational approaches in calculating the structure of pyridoxal 5’-phosphate (PLP)-dependent β-lyase protein. The importance of protein environment. Journal of Computational Chemistry, 26(5): 443–446, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prabhakar:2006:DSM

[PMM06] Rajeev Prabhakar, Keiji Morokuma, and Djamaladdin G. Musaev. A DFT study of the mechanism of Ni superoxide dismutase (NiSOD): Role of the active site cysteine-6 residue in the oxidative half-reaction. Journal of Computational Chemistry, 27(12):1438–1445, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Piquemal:2005:CSD

[PMPGP05] Jean-Philip Piquemal, Antonio Marquez, Olivier Parisel, and Claude Giessner-Prettre. A CSOV study of the difference between HF and DFT intermolecular interaction energy values: the importance of the charge transfer contribution. Journal of Computational Chemistry, 26(10):1052–1062, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Project:2008:PCP

[PNG08] Elad Project, Esther Nachliel, and Menachem Gutman. Parameterization of Ca^{+2}–protein interactions for molecular dynamics simulations. Journal of Computational Chemistry, 29 (7):1163–1169, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Page:2003:GES

[PO03] Christopher S. Page and Massimo Olivucci. Ground and excited state CASPT 2 geometry optimizations of small organic molecules. Journal of Computational Chemistry, 24(3):298– 309, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pogliani:2003:CGC

[Pog03] Lionello Pogliani. Complete graph conjecture for inner-core electrons: Homogeneous index case. Journal of Computational Chemistry, 24(9):1097–1109, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pogliani:2006:HPM

[Pog06] Lionello Pogliani. The hydrogen perturbation in molecular connectivity computations. Journal of Computational Chemistry, 27(7):868–882, May 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Price:2001:GPL

[POJ01] Melissa L. P. Price, Dennis Ostrovsky, and William L. Jorgensen. Gas-phase and liquid-state properties of esters, nitriles, and nitro compounds with the OPLS-AA force field. Journal of Computational Chemistry, 22(13):1340–1352, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pomelli:2004:TIG

[Pom04] Christian Silvio Pomelli. A tessellationless integration grid for the polarizable continuum model reaction field. Journal of Computational Chemistry, 25(12):1532–1541, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Palfi:2008:HSC

[PP08a] Villő K. Pálfi and András Perczel. How stable is a collagen triple helix? An ab initio study on various collagen and β-sheet forming sequences. Journal of Computational Chemistry, 29 (9):1374–1386, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pilme:2008:ABC

[PP08b] Julien Pilmé and Jean-Philip Piquemal. Advancing beyond charge analysis using the electronic localization function: Chemically intuitive distribution of electrostatic moments.

Journal of Computational Chemistry, 29(9):1440–1449, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pang:2001:ECP

[PPXP01] Yuan-Ping Pang, Emanuele Perola, Kun Xu, and Franklyn G. Prendergast. EUDOC: a computer program for identification of drug interaction sites in macromolecules and drug leads from chemical databases. Journal of Computational Chemistry, 22(15):1750–1771, November 30, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prabhu:2008:EII

[PPYS08] Ninad V. Prabhu, Manoranjan Panda, Qingyi Yang, and Kim A. Sharp. Explicit ion, implicit water solvation for molecular dynamics of nucleic acids and highly charged molecules. Journal of Computational Chemistry, 29(7):1113–1130, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prall:2001:VGT

[Pra01] Matthias Prall. VMD: a graphical tool for the modern chemists. Journal of Computational Chemistry, 22(1):132– 134, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Proschak:2008:SPL

[PRDS08] Ewgenij Proschak, Matthias Rupp, Swetlana Derksen, and Gisbert Schneider. Shapelets: Possibilities and limitations of shape-based virtual screening. Journal of Computational Chemistry, 29(1):108–114, January 15, 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pristovsek:2002:SSS

[PRJ02] Primož Pristovšek, Heinz Rüterjans, and Roman Jerala. Semiautomatic sequence-specific assignment of proteins based on the tertiary structure — the program st2nmr. Journal of Computational Chemistry, 23(3):335–340, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Park:2005:NSM

[PRKP05] Sanghyun Park, Randall J. Radmer, Teri E. Klein, and Vijay S. Pande. A new set of molecular mechanics parameters for hydroxyproline and its use in molecular dynamics simulations of collagen-like peptides. Journal of Computational Chemistry, 26(15):1612–1616, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Putz:2004:AHP

[PRS04] Mihai V. Putz, Nino Russo, and Emilia Sicilia. On the applicability of the HSAB principle through the use of improved computational schemes for chemical hardness evaluation. Journal of Computational Chemistry, 25(7):994–1003, May 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pitarch-Ruiz:2002:RCS

[PRSMM02] José Pitarch-Ruiz, José Sánchez-Marín, and Daniel Maynau. Reducing CAS-SDCI space. Using selected spaces in configuration interaction calculations in an efficient way. Journal of Computational Chemistry, 23(12):1157–1165, September 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pitarch-Ruiz:2003:VSC

[PRSMM03] José Pitarch-Ruiz, José Sánchez-Marín, and Daniel Maynau. Vertical spectrum of the C_{2}H system. An open shell (SC)^{2}– CAS–SDCI study. Journal of Computational Chemistry, 24 (5):609–617, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pitarch-Ruiz:2008:FCI

[PRSMV08] J. Pitarch-Ruiz, J. Sánchez-Marín, and A. M. Velasco. Full configuration interaction calculation of the low lying valence and Rydberg states of BeH. Journal of Computational Chemistry, 29(4):523–532, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Porta:2007:CMM

[PRT^{+}07] Josep M. Porta, Lluís Ros, Federico Thomas, Francesc Corcho, Josep Cantó, and Juan Jesús Pérez. Complete maps of molecular-loop conformational spaces. Journal of Computational Chemistry, 28(13):2170–2189, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Porta:2008:CMM

[PRT^{+}08] Josep M. Porta, Lluís Ros, Federico Thomas, Francesc Corcho, Josep Cantó, and Juan Jesús Pérez. Complete maps of molecular-loop conformational spaces. Journal of Computational Chemistry, 29(1):144–155, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prabhakar:2003:CMR

[PS03] Rajeev Prabhakar and Per E. M. Siegbahn. A comparison of the mechanism for the reductive half-reaction between pea seedling and other copper amine oxidases (CAOs). Journal of Computational Chemistry, 24(13):1599–1609, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Patra:2009:ANN

[PS09a] Jagdish C. Patra and Onkar Singh. Artificial neural networksbased approach to design ARIs using QSAR for diabetes mellitus. Journal of Computational Chemistry, 30(15):2494–2508, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Petridis:2009:MMF

[PS09b] Loukas Petridis and Jeremy C. Smith. A molecular mechanics force field for lignin. Journal of Computational Chemistry, 30 (3):457–467, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Paizs:2001:IBL

[PSC^{+}01] Béla Paizs, Pedro Salvador, Attila G. Császár, Miquel Duran, and Sándor Suhai. Intermolecular bond lengths: extrapolation to the basis set limit on uncorrected and BSSE-corrected potential energy hypersurfaces. Journal of Computational Chemistry, 22(2):196–207, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Poater:2009:MSP

[PSCD^{+}09] Albert Poater, Ana Gallegos Saliner, Ramon Carbó-Dorca, Jordi Poater, Miquel Solà, Luigi Cavallo, and Andrew P. Worth. Modeling the structure-property relationships of nanoneedles: a journey toward nanomedicine. Journal of Computational Chemistry, 30(2):275–284, January 30, 2009.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pierce:2000:CSM

[PSDM00] N. A. Pierce, J. A. Spriet, J. Desmet, and S. L. Mayo. Conformational splitting: a more powerful criterion for dead-end elimination. Journal of Computational Chemistry, 21(11): 999–1009, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Piacenza:2008:TEE

[PSF^{+}08] M. Piacenza, F. Della Sala, E. Fabiano, T. Maiolo, and G. Gigli. Torsional effects on excitation energies of thiophene derivatives induced by β-substituents: Comparison between time-dependent density functional theory and approximated coupled cluster approaches. Journal of Computational Chemistry, 29(3):451–457, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Poulain:2008:IPD

[PSHP08] P. Poulain, A. Saladin, B. Hartmann, and C. Prévost. Insights on protein-DNA recognition by coarse grain modelling. Journal of Computational Chemistry, 29(15):2582–2592, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pulay:2005:AEA

[PSMB05] Peter Pulay, Svein Saebo, Massimo Malagoli, and Jon Baker. Accuracy and efficiency of atomic basis set methods versus plane wave calculations with ultrasoft pseudopotentials for DNA base molecules. Journal of Computational Chemistry, 26 (6):599–605, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Preuss:2004:GES

[PSS^{+}04] M. Preuss, W. G. Schmidt, K. Seino, J. Furthmüller, and F. Bechstedt. Ground- and excited-state properties of DNA base molecules from plane-wave calculations using ultrasoft pseudopotentials. Journal of Computational Chemistry, 25(1): 112–122, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pomelli:2001:SAT

[PTC01] Christian Silvio Pomelli, Jacopo Tomasi, and Roberto Cammi. A Symmetry adapted tessellation of the GEPOL surface: applications to molecular properties in solution. Journal of Computational Chemistry, 22(12):1262–1272, September 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Piris:2009:IDO

[PU09] M. Piris and J. M. Ugalde. Iterative diagonalization for orbital optimization in natural orbital functional theory. Journal of Computational Chemistry, 30(13):2078–2086, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pullan:2005:UPB

[Pul05] Wayne Pullan. An unbiased population-based search for the geometry optimization of Lennard-Jones clusters: 2 ≤ N ≤ 372. Journal of Computational Chemistry, 26(9):899–906, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pernpointner:2003:PFC

[PV03] Markus Pernpointner and Lucas Visscher. Parallelization of four-component calculations. II. Symmetry-driven parallelization of the 4-Spinor CCSD algorithm. Journal of Computational Chemistry, 24(6):754–759, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pykavy:2007:SQC

[PV07] Mikhail Pykavy and Christoph Van Wüllen. A systematic quantum chemical investigation of the C H bond activation in methane by gas phase vanadium oxide cation VO^{+}. Journal of Computational Chemistry, 28(14):2252–2259, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pernpointner:2000:PFC

[PVdJB00] M. Pernpointner, L. Visscher, W. A. de Jong, and R. Broer. Parallelization of four-component calculations. I. Integral generation, SCF, and four-index transformation in the Dirac– Fock package MOLFDIR. Journal of Computational Chemistry, 21(13):1176–1186, October 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prall:2001:SEB

[PWFS01] Matthias Prall, Alexander Wittkopp, Andrey A. Fokin, and Peter R. Schreiner. Substituent effects on the Bergman cyclization of (Z)-1,5-hexadiyne-3-enes: a systematic computational study. Journal of Computational Chemistry, 22(13): 1605–1614, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Piquemal:2003:ILF

[PWHF^{+}03] Jean-Philip Piquemal, Ben Williams-Hubbard, Natalie Fey, Robert J. Deeth, Nohad Gresh, and Claude Giessner-Prettre. Inclusion of the ligand field contribution in a polarizable molecular mechanics: SIBFA-LF. Journal of Computational Chemistry, 24(16):1963–1970, December 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [PWHF^{+}04].

Piquemal:2004:EIL

[PWHF^{+}04] Jean-Philip Piquemal, Ben Williams-Hubbard, Natalie Fey, Robert J. Deeth, Nohad Gresh, and Claude Giessner-Prettre. Erratum: Inclusion of the ligand field contribution in a polarizable molecular mechanics: SIBFA-LF. Journal of Computational Chemistry, 25(2):308, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [PWHF^{+}03].

Ponec:2003:CSA

[PYCD03] Robert Ponec, Gleb Yuzhakov, and Ramon Carbó-Dorca. Chemical structures from the analysis of domain-averaged Fermi holes: Multiple metal metal bonding in transition metal compounds. Journal of Computational Chemistry, 24(15): 1829–1838, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pinsky:2003:SAE

[PYEA03] Mark Pinsky, Dina Yogev-Einot, and David Avnir. Statistical analysis of the estimation of distance measures. Journal of Computational Chemistry, 24(6):786–796, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ponec:2005:CSA

[PYS05] Robert Ponec, Gleb Yuzhakov, and Markku R. Sundberg. Chemical structures from the analysis of domainaveraged Fermi holes. Nature of the Mn Mn bond in bis(pentacarbonylmanganese). Journal of Computational Chemistry, 26(5):447–454, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Prabhu:2004:ITS

[PZS04] Ninad V. Prabhu, Peijuan Zhu, and Kim A. Sharp. Implementation and testing of stable, fast implicit solvation in molecular dynamics using the smooth-permittivity finite difference Poisson–Boltzmann method. Journal of Computational Chemistry, 25(16):2049–2064, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pascale:2004:CVF

[PZWG^{+}04] F. Pascale, C. M. Zicovich-Wilson, F. López Gejo, B. Civalleri, R. Orlando, and R. Dovesi. The calculation of the vibrational frequencies of crystalline compounds and its implementation in the CRYSTAL code. Journal of Computational Chemistry, 25(6):888–897, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qu:2005:ESS

[QB05] Yuhui Qu and Xiufang Bian. Electronic structure and stability of Al_{n}P_{n} (n = 2–4) clusters. Journal of Computational Chemistry, 26(3):226–234, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Quinet:2001:ATS

[QCK01] Olivier Quinet, Benoît Champagne, and Bernard Kirtman. Analytical TDHF second derivatives of dynamic electronic polarizability with respect to nuclear coordinates. Application to the dynamic ZPVA correction. Journal of Computational Chemistry, 22(16):1920–1932, December 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Quinet:2002:ATS

[QCK02] Olivier Quinet, Benoît Champagne, and Bernard Kirtman. Analytical TDHF second derivatives of dynamic electronic polarizability with respect to nuclear coordinates. Application to the dynamic ZPVA correction. Journal of Computational Chemistry, 23(15):1495–1496, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qiu:2009:USV

[QLHL09] Jian-Ding Qiu, San-Hua Luo, Jian-Hua Huang, and Ru-Ping Liang. Using support vector machines for prediction of protein structural classes based on discrete wavelet transform. Journal of Computational Chemistry, 30(8):1344–1350, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Queyroy:2009:NEE

[QNF09] Séverine Queyroy, Haruki Nakamura, and Ikuo Fukuda. Numerical examination of the extended phase-space volumepreserving integrator by the Nosé–Hoover molecular dynamics equations. Journal of Computational Chemistry, 30(12):1799– 1815, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qian:2001:NPO

[QSS01] Xiaoliang Qian, Daniel Strahs, and Tamar Schlick. A new program for optimizing periodic boundary models of solvated biomolecules (PBCAID). Journal of Computational Chemistry, 22(15):1843–1850, November 30, 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Queralt:2008:AMS

[QTdG^{+}08] Núria Queralt, David Taratiel, Coen de Graaf, Rosa Caballol, Renzo Cimiraglia, and Celestino Angeli. On the applicability of multireference second-order perturbation theory to study weak magnetic coupling in molecular complexes. Journal of Computational Chemistry, 29(6):994–1003, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Quapp:2001:CQR

[Qua01] Wolfgang Quapp. Comment “On the quadratic reaction path evaluated in a reduced potential energy surface model and the problem to locate transition states” [by J. M. Anglada, E. Besalú, J. M. Bofill, and R. Crehuet, J Comput Chem 2001, 22, 4, 387–406]. Journal of Computational Chemistry, 22(5):

537–540, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [ABBC01b, ABBC01b].

Quapp:2004:RPP

[Qua04] Wolfgang Quapp. Reaction pathways and projection operators: Application to string methods. Journal of Computational Chemistry, 25(10):1277–1285, July 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Quapp:2007:FTS

[Qua07] Wolfgang Quapp. Finding the transition state without initial guess: the growing string method for Newton trajectory to isomerization and enantiomerization reaction of alanine dipeptide and poly(15)alanine. Journal of Computational Chemistry, 28(11):1834–1847, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qu:2004:DFI

[QZL^{+}04] Zheng-Wang Qu, Hui Zhu, Ze-Sheng Li, Xing-Kang Zhang, and Qi-Yuan Zhang. Density functional investigation of reaction of borohydride cation BH with propylene. Journal of Computational Chemistry, 25(2):258–264, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qu:2003:DFI

[QZZZ03] Zheng-Wang Qu, Hui Zhu, Xing-Kang Zhang, and Qi-Yuan Zhang. Density functional investigation on electron-transfer catalysis of cycloreversion of cyclobutane: Radical anion mechanism. Journal of Computational Chemistry, 24(3):340– 344, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ramirez-Anguita:2009:FPDa

[RAGLL09a] Juan M. Ramírez-Anguita, Àngels González-Lafont, and José M. Lluch. Formation pathways of DMSO from DMS-OH in the presence of O_{2} and NO_{x}: a theoretical study. Journal of Computational Chemistry, 30(2):173–182, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ramirez-Anguita:2009:FPDb

[RAGLL09b] Juan M. Ramírez-Anguita, Àngels González-Lafont, and José M. Lluch. Formation pathways of DMSO_{2} in the addition channel of the OH-initiated DMS oxidation: a theoretical study. Journal of Computational Chemistry, 30(9):1477–1489, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rao:2000:MSH

[Rao00a] Sreedhara V. Rao. Mechanism of some 1:2 hydrogen transfer reactions through bond variation indices. Journal of Computational Chemistry, 21(14):1283–1291, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Raos:2000:IMF

[Rao00b] Nenad Raos. Iterative method for finding the low-energy conformations based on the concept of molecular volumes. Journal of Computational Chemistry, 21(15):1353–1360, November 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rapallo:2006:AUS

[Rap06] Arnaldo Rapallo. An algorithm for the uniform sampling of iso-energy surfaces and for the calculation of microcanonical averages. Journal of Computational Chemistry, 27(4):414– 425, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rankin:2001:DFT

[RB01] Kathryn N. Rankin and Russell J. Boyd. A density functional theory study of the dimers of HX (X = F, Cl, and Br). Journal of Computational Chemistry, 22(13):1590–1597, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ren:2004:MGL

[RC04] Yi Ren and San-Yan Chu. Modified Gaussian-2 level investigation of the identity ion-pair S_{N}2 reactions of lithium halide and methyl halide with inversion and retention mechanisms. Journal of Computational Chemistry, 25(4):461–471, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Repasky:2002:ISH

[RCJ02a] Matthew P. Repasky, Jayaraman Chandrasekhar, and William L. Jorgensen. Improved semiempirical heats of formation through the use of bond and group equivalents. Journal of Computational Chemistry, 23(4):498–510, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Repasky:2002:PPP

[RCJ02b] Matthew P. Repasky, Jayaraman Chandrasekhar, and William L. Jorgensen. PDDG/PM3 and PDDG/MNDO: Improved semiempirical methods. Journal of Computational Chemistry, 23(16):1601–1622, December 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rocha:2000:CIR

[RD00] Willian R. Rocha and Wagner B. De Almeida. Carbonyl insertion reaction into the Pt C bond in heterobimetallic Pt(SnCl_{3})(PH_{3})_{2}(CO)(CH_{3}) compound: Theoretical study. Journal of Computational Chemistry, 21(8):668–674, June 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rabone:2006:IPM

[RD06] J. A. L. Rabone and N. H. De Leeuw. Interatomic potential models for natural apatite crystals: Incorporating strontium and the lanthanides. Journal of Computational Chemistry, 27(2):253–266, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Roy:2008:DMI

[RDM^{+}08] Lindsay E. Roy, Tomasz Durakiewicz, Richard L. Martin, Juan E. Peralta, Gustavo E. Scuseria, Cliff G. Olson, John J. Joyce, and Ela Guziewicz. Dispersion in the Mott insulator UO_{2}: a comparison of photoemission spectroscopy and screened hybrid density functional theory. Journal of Computational Chemistry, 29(13):2288–2294, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rocha:2006:RRA

[RFSS06] Gerd B. Rocha, Ricardo O. Freire, Alfredo M. Simas, and James J. P. Stewart. RM1: a reparameterization of AM1 for H, C, N, O, P, S, F, Cl, Br, and I. Journal of Computational Chemistry, 27(10):1101–1111, July 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rajamani:2002:CQM

[RG02] Ramkumar Rajamani and Jiali Gao. Combined QM/MM study of the opsin shift in bacteriorhodopsin. Journal of Computational Chemistry, 23(1):96–105, January 15, 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rosso:2008:SDP

[RG08] Lula Rosso and Ian R. Gould. Structure and dynamics of phospholipid bilayers using recently developed general all-atom force fields. Journal of Computational Chemistry, 29(1):24– 37, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rosso:2008:ICS

[RGG08] Lula Rosso, Antony D. Gee, and Ian R. Gould. Ab initio computational study of positron emission tomography ligands interacting with lipid molecule for the prediction of nonspecific binding. Journal of Computational Chemistry, 29(14):2397– 2405, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Roux:2007:BPD

[RGP^{+}07] Celine Roux, Nohad Gresh, Lalith E. Perera, Jean-Philip Piquemal, and Laurent Salmon. Binding of 5-phospho-D-arabinonohydroxamate and 5-phospho- D-arabinonate inhibitors to zinc phosphomannose isomerase from Candida albicans studied by polarizable molecular mechanics and quantum mechanics. Journal of Computational Chemistry, 28(5): 938–957, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rohrig:2009:DHP

[RGZM09] Ute F. Röhrig, Aurélien Grosdidier, Vincent Zoete, and Olivier Michielin. Docking to heme proteins. Journal of Computational Chemistry, 30(14):2305–2315, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rurainski:2009:CLS

[RHL09] Alexander Rurainski, Andreas Hildebrandt, and Hans-Peter Lenhof. A consensus line search algorithm for molecular potential energy functions. Journal of Computational Chemistry, 30(9):1499–1509, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rathinavelan:2007:ETF

[RI07] Thenmalarchelvi Rathinavelan and Wonpil Im. Explicit treatment of force contribution from alignment tensor using overdetermined linear equations and its application in NMR structure determination. Journal of Computational Chemistry, 28 (11):1858–1864, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rathinavelan:2008:NSD

[RI08] Thenmalarchelvi Rathinavelan and Wonpil Im. A novel strategy to determine protein structures using exclusively residual dipolar coupling. Journal of Computational Chemistry, 29(10):1640–1649, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rulisek:2006:RMI

[RJLR06] Lubomír Rulíšek, Kasper P. Jensen, Kristoffer Lundgren, and Ulf Ryde. The reaction mechanism of iron and manganese superoxide dismutases studied by theoretical calculations. Journal of Computational Chemistry, 27(12):1398–1414, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Reveles:2004:GOD

[RK04] J. Ulises Reveles and Andreas M. Köster. Geometry optimization in density functional methods. Journal of Computational Chemistry, 25(9):1109–1116, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ruvinsky:2005:NFS

[RK05] A. M. Ruvinsky and A. V. Kozintsev. New and fast statisticalthermodynamic method for computation of protein-ligand binding entropy substantially improves docking accuracy. Journal of Computational Chemistry, 26(11):1089–1095, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rodriguez:2009:EGB

[RKA^{+}09] Juan I. Rodríguez, Andreas M. Köster, Paul W. Ayers, Ana Santos-Valle, Alberto Vela, and Gabriel Merino. An efficient grid-based scheme to compute QTAIM atomic properties without explicit calculation of zero-flux surfaces. Journal of Computational Chemistry, 30(7):1082–1092, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ryjacek:2003:NPC

[RKH03] Filip Ryjáček, Tomáš Kubař, and Pavel Hobza. New parameterization of the Cornell et al. empirical force field covering amino group nonplanarity in nucleic acid bases. Journal of Computational Chemistry, 24(15):1891–1902, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rassokhin:2001:NMM

[RLA01] Dmitrii N. Rassokhin, Victor S. Lobanov, and Dimitris K. Agrafiotis. Nonlinear mapping of massive data sets by fuzzy clustering and neural networks. Journal of Computational Chemistry, 22(4):373–386, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rivero:2009:PPW

[RLDI09] Pablo Rivero, Christoph Loschen, Ibério De P. R. Moreira, and Francesc Illas. Performance of plane-wave-based LDA + U and GGA + U approaches to describe magnetic coupling in molecular systems. Journal of Computational Chemistry, 30 (14):2316–2326, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rico:2004:EAC

[RLER04a] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Efficiency of the algorithms for the calculation of Slater molecular integrals in polyatomic molecules. Journal of Computational Chemistry, 25(16):1987–1994, December 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rico:2004:EPF

[RLER04b] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Electrostatic potentials and fields from density expansions of deformed atoms in molecules. Journal of Computational Chemistry, 25(11):1347–1354, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rico:2005:TSC

[RLER05] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Translation of STO charge distributions. Journal of Computational Chemistry, 26(8):846–855, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rico:2007:GBS

[RLER07] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Generation of basis sets with high degree of fulfillment of the Hellmann–Feynman theorem. Journal of Computational Chemistry, 28(4):748–758, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Robertson:2008:ELR

[RLP08] Alex Robertson, Edgar Luttmann, and Vijay S. Pande. Effects of long-range electrostatic forces on simulated protein folding kinetics. Journal of Computational Chemistry, 29(5):694–700, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rico:2004:AMR

[RLR^{+}04] J. Fernández Rico, R. López, G. Ramírez, I. Ema, and E. V. Ludeña. Analytical method for the representation of atomsin-molecules densities. Journal of Computational Chemistry, 25(11):1355–1363, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rico:2001:CBG

[RLRE01] J. Fernández Rico, R. López, G. Ramírez, and I. Ema. Correspondence between GTO and STO molecular basis sets. Journal of Computational Chemistry, 22(14):1655–1665, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rousseau:2000:AEC

[RM00] Emmanuelle Rousseau and Didier Mathieu. Atom equivalents for converting DFT energies calculated on molecular mechanics structures to formation enthalpies. Journal of Computational Chemistry, 21(5):367–379, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Raub:2007:QCI

[RM07] Stephan Raub and Christel M. Marian. Quantum chemical investigation of hydrogen-bond strengths and partition into donor and acceptor contributions. Journal of Computational Chemistry, 28(9):1503–1515, July 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rog:2003:MPP

[RMHK03] T. Róg, K. Murzyn, K. Hinsen, and G. R. Kneller. n Moldyn: a program package for a neutron scattering oriented analysis of molecular dynamics simulations. Journal of Computational Chemistry, 24(5):657–667, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Razumov:2001:MMC

[RMP01] Michael G. Razumov, Vladimir L. Melnikov, and Igor V. Pletnev. Molecular mechanics calculations of β-diketonate, aqua, and aqua-β-diketonate complexes of lanthanide ions using Gillespie–Kepert model. Journal of Computational Chemistry, 22(1):38–50, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rajamani:2003:IAU

[RNG03] Ramkumar Rajamani, Kevin J. Naidoo, and Jiali Gao. Implementation of an adaptive umbrella sampling method for the calculation of multidimensional potential of mean force of chemical reactions in solution. Journal of Computational Chemistry, 24(14):1775–1781, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rogalewicz:2000:INZ

[ROG00] Françoise Rogalewicz, Gilles Ohanessian, and Nohad Gresh. Interaction of neutral and zwitterionic glycine with Zn^{2+} in gas phase: ab initio and SIBFA molecular mechanics calculations. Journal of Computational Chemistry, 21(11):963–973, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ryde:2002:QCG

[RON02] Ulf Ryde, Lars Olsen, and Kristina Nilsson. Quantum chemical geometry optimizations in proteins using crystallographic raw data. Journal of Computational Chemistry, 23(11):1058– 1070, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ren:2002:CTI

[RP02] Pengyu Ren and Jay W. Ponder. Consistent treatment of inter- and intramolecular polarization in molecular mechanics calculations. Journal of Computational Chemistry, 23(16): 1497–1506, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Robertazzi:2004:HBS

[RP04] Arturo Robertazzi and James A. Platts. Hydrogen bonding, solvation, and hydrolysis of cisplatin: a theoretical study. Journal of Computational Chemistry, 25(8):1060–1067, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rafat:2007:AAP

[RP07a] M. Rafat and P. L. A. Popelier. Atom–atom partitioning of total (super)molecular energy: the hidden terms of classical force fields. Journal of Computational Chemistry, 28(1):292– 301, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rafat:2007:VIQ

[RP07b] M. Rafat and P. L. A. Popelier. Visualization and integration of quantum topological atoms by spatial discretization into finite elements. Journal of Computational Chemistry, 28(16): 2602–2617, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rafat:2007:LRB

[RP07c] Michel Rafat and Paul L. A. Popelier. Long range behavior of high-rank topological multipole moments. Journal of Computational Chemistry, 28(4):832–838, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rychkov:2007:JNA

[RP07d] Georgy Rychkov and Michael Petukhov. Joint neighbors approximation of macromolecular solvent accessible surface area. Journal of Computational Chemistry, 28(12):1974–1989, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Remmert:2009:EAC

[RP09] Sarah Remmert and Carol Parish. Energetic analyses of chair and boat conformations of maleimide substituted cyclohexane derivatives. Journal of Computational Chemistry, 30(6):992– 998, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Reith:2003:DEM

[RPMP03] Dirk Reith, Mathias Pütz, and Florian Müller-Plathe. Deriving effective mesoscale potentials from atomistic simulations. Journal of Computational Chemistry, 24(13):1624–1636, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rani:2007:ESB

[RPNJ07] Dandamudi Usha Rani, Dasari L. V. K. Prasad, John F. Nixon, and Eluvathingal D. Jemmis. Electronic structure and bonding studies on triple-decker sandwich complexes with a P_{6} middle ring. Journal of Computational Chemistry, 28(1): 310–319, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rungnim:2005:DFS

[RR05] Chompoonut Rungnim and Vithaya Ruangpornvisuti. A density functional study of propylene glycol conversion to propanal and propanone of various acid-catalyzed reaction models: a water-addition effect. Journal of Computational Chemistry, 26(15):1592–1599, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rodriguez-Ropero:2008:ICS

[RRCA08] Francisco Rodríguez-Ropero, Jordi Casanovas, and Carlos Alemán. Ab initio calculations on π-stacked thiophene dimer, trimer, and tetramer: Structure, interaction energy, cooperative effects, and intermolecular electronic parameters. Journal of Computational Chemistry, 29(1):69–78, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ruiz:2003:ACE

[RRFC^{+}03] Eliseo Ruiz, Antonio Rodríguez-Fortea, Joan Cano, Santiago Alvarez, and Pere Alemany. About the calculation of exchange coupling constants in polynuclear transition metal complexes. Journal of Computational Chemistry, 24(8):982– 989, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rassolov:2001:BST

[RRP^{+}01] Vitaly A. Rassolov, Mark A. Ratner, John A. Pople, Paul C. Redfern, and Larry A. Curtiss. 6-31G* basis set for third-row atoms. Journal of Computational Chemistry, 22(9):976–984, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rivalta:2006:MAC

[RRS06] Ivan Rivalta, Nino Russo, and Emilia Sicilia. Methane activation by chromium oxide cations in the gas phase: a theoretical study. Journal of Computational Chemistry, 27(2): 174–187, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rubensson:2007:HSM

[RRS07] Emanuel H. Rubensson, Elias Rudberg, and Paweł Sałek. A hierarchic sparse matrix data structure for large-scale Hartree– Fock/Kohn–Sham calculations. Journal of Computational Chemistry, 28(16):2531–2537, December 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rubensson:2009:TSM

[RRS09] Emanuel H. Rubensson, Elias Rudberg, and Paweł Sałek. Truncation of small matrix elements based on the Euclidean norm for blocked data structures. Journal of Computational Chemistry, 30(6):974–977, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rodriguez-Ropero:2008:MDC

[RRZA08] Francisco Rodríguez-Ropero, David Zanuy, and Carlos Alemán. Molecular dynamics of a calix[4]arene-containing polymer in dichloromethane solution: Ability of the solvent molecules to fill the cavity of the macrocycle. Journal of Computational Chemistry, 29(8):1233–1241, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rubensson:2005:SSM

[RS05] Emanuel H. Rubensson and Paweł Sałek. Systematic sparse matrix error control for linear scaling electronic structure calculations. Journal of Computational Chemistry, 26(15):1628– 1637, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ruedenberg:2007:EWD

[RS07a] Klaus Ruedenberg and Michael W. Schmidt. Erratum: Why does electron sharing lead to covalent bonding? A variational analysis. Journal of Computational Chemistry, 28(14):2389, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [RS07b].

Ruedenberg:2007:WDE

[RS07b] Klaus Ruedenberg and Michael W. Schmidt. Why does electron sharing lead to covalent bonding? A variational analysis. Journal of Computational Chemistry, 28(1):391–410, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [RS07a].

Rotkiewicz:2008:FPR

[RS08] Piotr Rotkiewicz and Jeffrey Skolnick. Fast procedure for reconstruction of full-atom protein models from reduced representations. Journal of Computational Chemistry, 29(9):1460– 1465, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Reddy:2007:IQM

[RSE07] M. Rami Reddy, U. C. Singh, and Mark D. Erion. Ab initio quantum mechanics-based free energy perturbation method for calculating relative solvation free energies. Journal of Computational Chemistry, 28(2):491–494, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rey-Stolle:2009:TBM

[RSER09] M. Fernanda Rey-Stolle, Marta Enciso, and Antonio Rey. Topology-based models and NMR structures in protein folding simulations. Journal of Computational Chemistry, 30(8): 1212–1219, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rocchia:2002:RGB

[RSN^{+}02] Walter Rocchia, Sundaram Sridharan, Anthony Nicholls, Emil Alexov, Alessandro Chiabrera, and Barry Honig. Rapid gridbased construction of the molecular surface and the use of induced surface charge to calculate reaction field energies: Applications to the molecular systems and geometric objects. Journal of Computational Chemistry, 23(1):128–137, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rankin:2003:THP

[RSP03] Kathryn N. Rankin, Traian Sulea, and Enrico O. Purisima. On the transferability of hydration-parametrized continuum electrostatics models to solvated binding calculations. Journal of Computational Chemistry, 24(8):954–962, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rupp:2009:DPH

[RSS09] Matthias Rupp, Petra Schneider, and Gisbert Schneider. Distance phenomena in high-dimensional chemical descriptor spaces: Consequences for similarity-based approaches. Journal of Computational Chemistry, 30(14):2285–2296, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rahal-Sekkal:2003:SED

[RSSKB03] Majda Rahal-Sekkal, Nezha Sekkal, Dirk C. Kleb, and Paul Bleckmann. Structures and energies of D-galactose and galabiose conformers as calculated by ab initio and semiempirical methods. Journal of Computational Chemistry, 24(7):806– 818, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Russo:2000:TDE

[RTG00] Nino Russo, Marirosa Toscano, and André Grand. Theoretical determination of electron affinity and ionization potential of DNA and RNA bases. Journal of Computational Chemistry, 21(14):1243–1250, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rudolph:2005:AECa

[Rud05a] Manfred Rudolph. Attaining exponential convergence for the flux error with second- and fourth-order accurate finitedifference equations. I. Presentation of the basic concept and application to a pure diffusion system. Journal of Computational Chemistry, 26(6):619–632, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rudolph:2005:AECb

[Rud05b] Manfred Rudolph. Attaining exponential convergence for the flux error with second- and fourth-order accurate finitedifference equations. II. Application to systems comprising first-order chemical reactions. Journal of Computational Chemistry, 26(6):633–641, April 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rudolph:2005:AECc

[Rud05c] Manfred Rudolph. Attaining exponential convergence for the flux error with second- and fourth-order accurate finitedifference equations. Part 3. Application to electrochemical systems comprising second-order chemical reactions. Journal of Computational Chemistry, 26(11):1193–1204, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Roy:2006:ATC

[RUPH06] Ram Kinkar Roy, V. Usha, Bhisma K. Patel, and Kimihiko Hirao. Acetalization and thioacetalization of cabonyl compounds: a case study based on global and local electrophilicity descriptors. Journal of Computational Chemistry, 27(6): 773–780, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ruvinsky:2007:RBE

[Ruv07] Anatoly M. Ruvinsky. Role of binding entropy in the refinement of protein–ligand docking predictions: Analysis based on the use of 11 scoring functions. Journal of Computational Chemistry, 28(8):1364–1372, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rosta:2009:ARC

[RWBH09] Edina Rosta, H. Lee Woodcock, Bernard R. Brooks, and Gerhard Hummer. Artificial reaction coordinate “tunneling” in free-energy calculations: the catalytic reaction of RNase H. Journal of Computational Chemistry, 30(11):1634–1641, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ren:2009:DEE

[RY09] Yi Ren and Hiroshi Yamataka. Does α-effect exist in E2 reactions? A G2(+) investigation. Journal of Computational Chemistry, 30(3):358–365, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Rutkowska-Zbik:2007:TDF

[RZWS07] Dorota Rutkowska-Zbik, Małgorzata Witko, and Grazyna Stochel. Theoretical density functional theory studies on interactions of small biologically active molecules with isolated heme group. Journal of Computational Chemistry, 28(4):825– 831, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

So:2007:VEE

[SA07] Remmick So and Saman Alavi. Vertical excitation energies for ribose and deoxyribose nucleosides. Journal of Computational Chemistry, 28(11):1776–1782, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Salvador:2006:GEI

[SAM06] P. Salvador, D. Asturiol, and I. Mayer. A general efficient implementation of the BSSE-free SCF and MP2 methods based on the chemical Hamiltonian approach. Journal of Computational Chemistry, 27(13):1505–1516, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sandak:2001:MFS

[San01] Bilha Sandak. Multiscale fast summation of long-range charge and dipolar interactions. Journal of Computational Chemistry, 22(7):717–731, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schuurman:2005:ILQ

[SAS05] Michael S. Schuurman, Wesley D. Allen, and Henry F. Schaefer III. The ab initio limit quartic force field of BH_{3}. Journal of Computational Chemistry, 26(11):1106–1112, August 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sayyed-Ahmad:2004:EST

[SATO04] Abdallah Sayyed-Ahmad, Kagan Tuncay, and Peter J. Ortoleva. Efficient solution technique for solving the Poisson– Boltzmann equation. Journal of Computational Chemistry, 25(8):1068–1074, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Saunders:2004:SSI

[Sau04] Martin Saunders. Stochastic search for isomers on a quantum mechanical surface. Journal of Computational Chemistry, 25 (5):621–626, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sommer:2001:MRI

[SB01] Ingolf Sommer and Richard Brimacombe. Methods for refining interactively established models of ribosomal RNA towards a physico-chemically plausible structure. Journal of Computational Chemistry, 22(4):407–417, March 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Swart:2008:QQR

[SB08] Marcel Swart and F. Matthias Bickelhaupt. QUILD: QUantum-regions interconnected by local descriptions. Journal of Computational Chemistry, 29(5):724–734, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Smith:2002:RQC

[SBB02] Grant D. Smith, Oleg Borodin, and Dmitry Bedrov. A revised quantum chemistry-based potential for poly(ethylene oxide) and its oligomers in aqueous solution. Journal of Computational Chemistry, 23(15):1480–1488, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sgrignani:2009:IDS

[SBG^{+}09a] Jacopo Sgrignani, Claudia Bonaccini, Giovanni Grazioso, Matteo Chioccioli, Andrea Cavalli, and Paola Gratteri. Insights into docking and scoring neuronal α4β2 nicotinic receptor agonists using molecular dynamics simulations and QM/ MM calculations. Journal of Computational Chemistry, 30 (15):2443–2454, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Simpson:2009:MSA

[SBG09b] Darren J. Simpson, Thomas Bredow, and Andrea R. Gerson. MSINDO study of acid promoted dissolution of planar MgO and NiO surfaces. Journal of Computational Chemistry, 30 (4):581–588, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Strassner:2002:MPF

[SBH02] Thomas Strassner, Markus Busold, and Wolfgang A. Herrmann. MM3 parametrization of four- and five-coordinated rhenium complexes by a genetic algorithm — which factors influence the optimization performance? Journal of Computational Chemistry, 23(2):282–290, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Santoro:2008:CTD

[SBI08] Fabrizio Santoro, Vincenzo Barone, and Roberto Improta. Can TD-DFT calculations accurately describe the excited states behavior of stacked nucleobases? The cytosine dimer as a test case. Journal of Computational Chemistry, 29(6): 957–964, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schluttig:2008:CMD

[SBJ08] Jakob Schluttig, Michael Bachmann, and Wolfhard Janke. Comparative molecular dynamics and Monte Carlo study of statistical properties for coarse-grained heteropolymers. Journal of Computational Chemistry, 29(15):2603–2612, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Seda:2005:SES

[ŠBL05] Josef Šeda, Jaroslav V. Burda, and Jerzy Leszczynski. Study of electronic spectra of free-base porphin and Mg-porphin: Comprehensive comparison of variety of ab initio, DFT, and semiempirical methods. Journal of Computational Chemistry, 26(3):294–303, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schaefer:2001:EAV

[SBLK01] Michael Schaefer, Christian Bartels, Fabrice Leclerc, and Martin Karplus. Effective atom volumes for implicit solvent models: comparison between Voronoi volumes and minimum fluctuation volumes. Journal of Computational Chemistry, 22(15): 1857–1879, November 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sherer:2001:QCC

[SC01] Edward C. Sherer and Christopher J. Cramer. Quantum chemical characterization of the cytosine: 2-Aminopurine base pair. Journal of Computational Chemistry, 22(11):1167–1179, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shao:2004:DLS

[SCC04] Xueguang Shao, Longjiu Cheng, and Wensheng Cai. A dynamic lattice searching method for fast optimization of Lennard-Jones clusters. Journal of Computational Chemistry, 25(14):1693–1698, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sousa:2009:CAP

[SCF^{+}09] Sérgio F. Sousa, Emanuela S. Carvalho, Diana M. Ferreira, Isabel S. Tavares, Pedro A. Fernandes, Maria João Ramos, and José A. N. F. Gomes. Comparative analysis of the performance of commonly available density functionals in the determination of geometrical parameters for zinc complexes. Journal of Computational Chemistry, 30(16):2752–2763, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Soto-Castro:2004:MVD

[SCG04] Delia Soto-Castro and Patricia Guadarrama. Macrocyclic vs. dendrimeric effect. A DFT study. Journal of Computational Chemistry, 25(10):1215–1226, July 30, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schuurmann:2000:PHL

[Sch00] Gerrit Schüürmann. Prediction of Henry’s law constant of benzene derivatives using quantum chemical continuumsolvation models. Journal of Computational Chemistry, 21(1): 17–34, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schlegel:2003:EPE

[Sch03] H. Bernhard Schlegel. Exploring potential energy surfaces for chemical reactions: an overview of some practical methods. Journal of Computational Chemistry, 24(12):1514–1527, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schmid:2004:CPS

[Sch04] Rochus Schmid. Car–Parrinello simulations with a real space method. Journal of Computational Chemistry, 25(6):799–812, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stashans:2008:ESC

[SCP08] Arvids Stashans, Gaston Chamba, and Henry Pinto. Electronic structure, chemical bonding, and geometry of pure and Sr-doped CaCO_{3}. Journal of Computational Chemistry, 29(3): 343–349, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Scemama:2007:MPD

[SCS07] Anthony Scemama, Michel Caffarel, and Andreas Savin. Maximum probability domains from Quantum Monte Carlo calculations. Journal of Computational Chemistry, 28(1):442–454, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Seeliger:2009:TGV

[SD09] Daniel Seeliger and Bert L. De Groot. tCONCOORD-GUI: Visually supported conformational sampling of bioactive molecules. Journal of Computational Chemistry, 30(7):1160– 1166, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sefcik:2002:DCE

[SDCG02] Jan Sefcik, Ersan Demiralp, Tahir Cagin, and William A. Goddard III. Dynamic Charge Equilibration-Morse stretch force field: Application to energetics of pure silica zeolites. Journal of Computational Chemistry, 23(16):1507–1514, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sun:2007:CSE

[SDL07] Yu Sun, Brian N. Dominy, and Robert A. Latour. Comparison of solvation-effect methods for the simulation of peptide interactions with a hydrophobic surface. Journal of Computational Chemistry, 28(11):1883–1892, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Salisburg:2009:RTP

[SDL^{+}09] Amanda M. Salisburg, Ashley L. Deline, Katrina W. Lexa, George C. Shields, and Karl N. Kirschner. Ramachandrantype plots for glycosidic linkages: Examples from molecular dynamic simulations using the Glycam06 force field. Journal of Computational Chemistry, 30(6):910–921, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Suarez:2002:MDS

[SDM02] Dimas Suárez, Natalia Díaz, and Kenneth M. Merz Jr. Molecular dynamics simulations of the dinuclear zinc-β-lactamase from Bacteroides fragilis complexed with imipenem. Journal of Computational Chemistry, 23(16):1587–1600, December 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schuler:2001:IGF

[SDvG01] Lukas D. Schuler, Xavier Daura, and Wilfred F. van Gunsteren. An improved GROMOS96 force field for aliphatic hydrocarbons in the condensed phase. Journal of Computational Chemistry, 22(11):1205–1218, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stepanenko:2007:GTS

[SE07] Svetlana Stepanenko and Bernd Engels. Gradient tabu search. Journal of Computational Chemistry, 28(2):601–611, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stepanenko:2008:NTS

[SE08] Svetlana Stepanenko and Bernd Engels. New tabu search based global optimization methods outline of algorithms and study of efficiency. Journal of Computational Chemistry, 29 (5):768–780, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schwobel:2009:MHB

[SEKS09] Johannes Schwöbel, Ralf-Uwe Ebert, Ralph Kühne, and Gerrit Schüürmann. Modeling the H bond donor strength of OH, NH, and CH sites by local molecular parameters. Journal of Computational Chemistry, 30(9):1454–1464, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Seno:2006:ATS

[Sen06] Yasunobu Seno. Allostery of the two-state model of hemoglobin studied by ECEPP energy minimization. Journal of Computational Chemistry, 27(6):701–710, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shen:2005:SMF

[SF05] Min-Yi Shen and Karl F. Freed. A simple method for faster nonbonded force evaluations. Journal of Computational Chemistry, 26(7):691–698, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sabzyan:2007:EFE

[SF07] Hassan Sabzyan and Davood Farmanzadeh. Electric field effects on the performance of a candidate multipole molecular switch: a quantum computational study. Journal of Computational Chemistry, 28(5):922–931, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Soler:2004:FTA

[SFC04] Philippe Soler, Franck Fuster, and Hilaire Chevreau. Fast topological analysis of 2D and 3D grids of data: Application to the atoms in molecule (AIM) and the electron localization function (ELF). Journal of Computational Chemistry, 25(15): 1920–1925, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sousa:2007:TSF

[SFR07] Sérgio Filipe Sousa, Pedro Alexandrino Fernandes, and Maria João Ramos. Theoretical studies on farnesyl transferase: Evidence for thioether product coordination to the active-site zinc sphere. Journal of Computational Chemistry, 28(7):1160–1168, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Smedarchina:2001:DPC

[SFRS01] Zorka Smedarchina, Antonio Fernández-Ramos, and Willem Siebrand. DOIT: a program to calculate thermal rate constants and mode-specific tunneling splittings directly from quantum-chemical calculations. Journal of Computational Chemistry, 22(7):787–801, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Steiger:2001:LSC

[SG01] Don Steiger and Rainer Glaser. Lattice sum calculations for 1/r^{p} interactions via multipole expansions and Euler summation. Journal of Computational Chemistry, 22(2):208–215, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schreiber:2007:SBA

[SG07a] Marko Schreiber and Leticia González. Structure and bonding of Ag(I)–DNA base complexes and Ag(I)–adenine–cytosine mispairs: an ab initio study. Journal of Computational Chemistry, 28(14):2299–2308, November 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Slipchenko:2007:EEE

[SG07b] Lyudmila V. Slipchenko and Mark S. Gordon. Electrostatic energy in the effective fragment potential method: Theory and application to benzene dimer. Journal of Computational Chemistry, 28(1):276–291, January 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Solomon:2006:MTB

[SGD06] Edward I. Solomon, Serge I. Gorelsky, and Abhishek Dey. Metal–thiolate bonds in bioinorganic chemistry. Journal of Computational Chemistry, 27(12):1415–1428, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sapre:2009:GCO

[SGPS09] Nitin S. Sapre, Swagata Gupta, Nilanjana Pancholi, and Neelima Sapre. A group center overlap based approach for “3D QSAR” studies on TIBO derivatives. Journal of Computational Chemistry, 30(6):922–933, April 30, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schulz-Gasch:2003:MIO

[SGS03] Tanja Schulz-Gasch and Martin Stahl. Mechanistic insights into oxidosqualene cyclizations through homology modeling. Journal of Computational Chemistry, 24(6):741–753, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Saue:2002:FCR

[SH02] Trond Saue and Trygve Helgaker. Four-component relativistic Kohn–Sham theory. Journal of Computational Chemistry, 23 (8):814–823, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Salek:2007:SCP

[SH07] Paweł Sałek and Andreas Hesselmann. A self-contained and portable density functional theory library for use in Ab Initio quantum chemistry programs. Journal of Computational Chemistry, 28(16):2569–2575, December 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shaik:2008:CGV

[SH08] Sason S. Shaik and Philippe C. Hiberty. A Chemist’s Guide to Valence Bond Theory. Wiley-Interscience, New York, NY, USA, 2008. ISBN 0-470-03735-0 (hardcover). xiv + 316 pp. LCCN QD469 .S48 2008. URL http://www.loc.

gov/catdir/enhancements/fy0745/2007040432-d.html;
http://www.loc.gov/catdir/enhancements/fy0801/2007040432-
t.html;   http://www.loc.gov/catdir/enhancements/fy0806/
2007040432-b.html.

Sandberg:2009:CAH

[SH09] Thomas Sandberg and Matti Hotokka. Conformational analysis of hydroxymatairesinol in aqueous solution with molecular dynamics. Journal of Computational Chemistry, 30(16):2666– 2673, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shalabi:2002:FAF

[Sha02] A. S. Shalabi. F_{A}(I):Au^{+} and F_{A}(II):Cu^{+} laser activity and photographic sensitization at the low coordinated surfaces of AgBr ab initio calculations. Journal of Computational Chemistry, 23(11):1104–1120, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shaw:2005:FSM

[Sha05] David E. Shaw. A fast, scalable method for the parallel evaluation of distance-limited pairwise particle interactions. Journal of Computational Chemistry, 26(13):1318–1328, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [?].

Shaik:2007:ELL

[Sha07] Sason Shaik. Essay: The Lewis legacy: the chemical bond — a territory and heartland of chemistry. Journal of Computational Chemistry, 28(1):51–61, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schreiber:2005:EEL

[SHBD05] Andreas Schreiber, Michael Humbert, Alexander Benz, and Ursula Dietrich. 3D-Epitope-Explorer (3DEX): Localization of conformational epitopes within three-dimensional structures of proteins. Journal of Computational Chemistry, 26(9): 879–887, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Song:2008:MMP

[SHD^{+}08] Kun Song, Viktor Hornak, Carlos De Los Santos, Arthur P. Grollman, and Carlos Simmerling. Molecular mechanics parameters for the FapydG DNA lesion. Journal of Computational Chemistry, 29(1):17–23, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Siegert:2007:GDP

[SHH07] Martin R. Siegert, Matthias Heuchel, and Dieter Hofmann. A generalized direct-particle-deletion scheme for the calculation of chemical potential and solubilities of small- and mediumsized molecules in amorphous polymers. Journal of Computational Chemistry, 28(5):877–889, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Soares:2005:INA

[SHK^{+}05] Thereza A. Soares, Philippe H. Hünenberger, Mika A. Kastenholz, Vincent Kräutler, Thomas Lenz, Roberto D. Lins, Chris Oostenbrink, and Wilfred F. van Gunsteren. An improved nucleic acid parameter set for the GROMOS force field. Journal of Computational Chemistry, 26(7):725–737, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shah:2004:PRP

[SHM04] Premal S. Shah, Geoffrey K. Hom, and Stephen L. Mayo. Preprocessing of rotamers for protein design calculations. Journal of Computational Chemistry, 25(14):1797–1800, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schwarzl:2005:NCS

[SHSF05] Sonja M. Schwarzl, Danzhi Huang, Jeremy C. Smith, and Stefan Fischer. Nonuniform charge scaling (NUCS): a practical approximation of solvent electrostatic screening in proteins. Journal of Computational Chemistry, 26(13):1359–1371, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Siegbahn:2001:MAM

[Sie01] P. E. M. Siegbahn. Modeling aspects of mechanisms for reactions catalyzed by metalloenzymes. Journal of Computational Chemistry, 22(14):1634–1645, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Simoes:2007:EBW

[Sim07] Ana Simões. Essay: In between worlds: G. N. Lewis, the shared pair bond and its multifarious contexts. Journal of Computational Chemistry, 28(1):62–72, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shen:2004:HPS

[SJJ^{+}04] Qi Shen, Jian-Hui Jiang, Chen-Xu Jiao, Wei-Qi Lin, Guo-Li Shen, and Ru-Qin Yu. Hybridized particle swarm algorithm for adaptive structure training of multilayer feed-forward neural network: QSAR studies of bioactivity of organic compounds. Journal of Computational Chemistry, 25(14):1726– 1735, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sun:2009:APE

[SJW09] Chang-Liang Sun, Xiao-Nan Jiang, and Chang-Sheng Wang. An analytic potential energy function for the amide–amide and amide–water intermolecular hydrogen bonds in peptides. Journal of Computational Chemistry, 30(15):2567–2575, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sun:2005:UIF

[SK05] H. Sun and P. W.-C. Kung. Urea: an ab initio and force field study of the gas and solid phases. Journal of Computational Chemistry, 26(2):169–174, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Seo:2008:NSC

[SK08] Dong-Kyun Seo and Sang-Hwan Kim. Nature of Stoner condition for metallic ferromagnetism. Journal of Computational Chemistry, 29(13):2172–2176, October 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sakuraba:2009:MMT

[SK09] Shun Sakuraba and Akio Kitao. Multiple Markov transition matrix method: Obtaining the stationary probability distribution from multiple simulations. Journal of Computational Chemistry, 30(12):1850–1858, September 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schweizer:2008:LSC

[SKDO08] Sabine Schweizer, Jörg Kussmann, Bernd Doser, and Christian Ochsenfeld. Linear-scaling Cholesky decomposition. Journal of Computational Chemistry, 29(6):1004–1010, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sayano:2000:MMC

[SKGS00] Kenji Sayano, Hidetoshi Kono, M. Michael Gromiha, and Akinori Sarai. Multicanonical Monte Carlo calculation of the freeenergy map of the base–amino acid interaction. Journal of Computational Chemistry, 21(11):954–962, August 2000. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Soderhjelm:2007:ADM

[SKK^{+}07] P. Söderhjelm, J. W. Krogh, G. Karlström, U. Ryde, and R. Lindh. Accuracy of distributed multipoles and polarizabilities: Comparison between the LoProp and MpProp models. Journal of Computational Chemistry, 28(6):1083–1090, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sanville:2007:IGB

[SKSH07] Edward Sanville, Steven D. Kenny, Roger Smith, and Graeme Henkelman. Improved grid-based algorithm for Bader charge allocation. Journal of Computational Chemistry, 28(5):899– 908, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shukla:2004:TIN

[SL04] M. K. Shukla and Jerzy Leszczynski. TDDFT investigation on nucleic acid bases: Comparison with experiments and standard approach. Journal of Computational Chemistry, 25(5):

768–778, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sun:2006:CIS

[SL06] Yu Sun and Robert A. Latour. Comparison of implicit solvent models for the simulation of protein–surface interactions. Journal of Computational Chemistry, 27(16):1908–1922, December 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sakharov:2009:FFI

[SL09] Dmitri V. Sakharov and Carmay Lim. Force fields including charge transfer and local polarization effects: Application to proteins containing multi/heavy metal ions. Journal of Computational Chemistry, 30(2):191–202, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shien:2009:ISC

[SLC^{+}09] Dray-Ming Shien, Tzong-Yi Lee, Wen-Chi Chang, Justin Bo-Kai Hsu, Jorng-Tzong Horng, Po-Chiang Hsu, Ting-Yuan Wang, and Hsien-Da Huang. Incorporating structural characteristics for identification of protein methylation sites. Journal of Computational Chemistry, 30(9):1532–1543, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Song:2009:ETS

[SLHW09] Xinli Song, Jicun Li, Hua Hou, and Baoshan Wang. Extensive theoretical studies of a new energetic material: Tetrazinotetrazine-tetraoxide (TTTO). Journal of Computational Chemistry, 30(12):1816–1820, September 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [JW12].

Sheng:2004:IDD

[SLL^{+}04a] Li Sheng, Ze-Sheng Li, Jing-Yao Liu, Jing-Fa Xiao, and Chia-Chung Sun. Ab initio direct dynamics studies on the reaction of H atom with CH_{3}CH_{2}Cl. Journal of Computational Chemistry, 25(1):72–82, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sheng:2004:TSR

[SLL^{+}04b] Li Sheng, Ze-Sheng Li, Jing-Yao Liu, Jing-Fa Xiao, and Chia-Chung Sun. Theoretical study on the rate constants for the C_{2}H_{5} + HBr → C_{2}H_{6} + Br reaction. Journal of Computational Chemistry, 25(3):423–428, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sunoj:2001:CID

[SLRC01] R. B. Sunoj, P. Lakshminarasimhan, V. Ramamurthy, and J. Chandrasekhar. Configuration interaction and density functional study of the influence of lithium cation complexation on vertical and adiabatic excitation energies of enones. Journal of Computational Chemistry, 22(13):1598–1604, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shinoda:2003:RBD

[SM03] Wataru Shinoda and Masuhiro Mikami. Rigid-body dynamics in the isothermal-isobaric ensemble: a test on the accuracy and computational efficiency. Journal of Computational Chemistry, 24(8):920–930, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shi:2006:DEP

[SM06] Chunsheng Shi and Masahiko Morinaga. Doping effects on proton incorporation and conduction in SrZrO_{3}. Journal of Computational Chemistry, 27(6):711–718, April 30, 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Samolyuk:2008:RBC

[SM08a] German D. Samolyuk and Gordon J. Miller. Relation between chemical bonding and exchange coupling approaches to the description of ordering in itinerant magnets. Journal of Computational Chemistry, 29(13):2177–2186, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shi:2008:IME

[SM08b] Wei Shi and Edward J. Maginn. Improvement in molecule exchange efficiency in Gibbs ensemble Monte Carlo: Development and implementation of the continuous fractional component move. Journal of Computational Chemistry, 29(15):

2520–2530, November 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sanchez:2000:SEM

[SMAdV00] M. L. Sánchez, M. E. Martín, M. A. Aguilar, and F. J. Olivares del Valle. Solvent effects by means of averaged solvent electrostatic potentials: Coupled method. Journal of Computational Chemistry, 21(9):705–715, July 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stoll:2002:REC

[SMD02] Hermann Stoll, Bernhard Metz, and Michael Dolg. Relativistic energy-consistent pseudopotentials — recent developments. Journal of Computational Chemistry, 23(8):767–778, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Song:2009:MIP

[SMG09] Yifan Song, Junjun Mao, and M. R. Gunner. MCCE2: Improving protein pK_{a} calculations with extensive side chain rotamer sampling. Journal of Computational Chemistry, 30(14): 2231–2247, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schlund:2008:CAA

[SMGE08] Sebastian Schlund, Robert Müller, Carsten Graßmann, and Bernd Engels. Conformational analysis of arginine in gas phase — a strategy for scanning the potential energy surface effectively. Journal of Computational Chemistry, 29(3): 407–415, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shukla:2000:ISE

[SMKM00] M. K. Shukla, S. K. Mishra, Anil Kumar, and P. C. Mishra. An ab initio study of excited states of guanine in the gas phase and aqueous media: Electronic transitions and mechanism of spectral oscillations. Journal of Computational Chemistry, 21 (10):826–846, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Spiegel:2008:PAB

[SMM^{+}08] Katrin Spiegel, Alessandra Magistrato, Patrick Maurer, Paolo Ruggerone, Ursula Rothlisberger, Paolo Carloni, Jan Reedijk, and Michael L. Klein. Parameterization of azole-bridged dinuclear platinum anticancer drugs via a QM/MM force matching procedure. Journal of Computational Chemistry, 29(1):38–49, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Suresh:2009:TAN

[SMV^{+}09] Cherumuttathu H. Suresh, Neetha Mohan, K. Periya Vijayalakshmi, Renjumon George, and Janice M. Mathew. Typical aromatic noncovalent interactions in proteins: a theoretical study using phenylalanine. Journal of Computational Chemistry, 30(9):1392–1404, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Song:2005:XPI

[SMZW05] Lingchun Song, Yirong Mo, Qianer Zhang, and Wei Wu. XMVB: a program for ab initio nonorthogonal valence bond computations. Journal of Computational Chemistry, 26(5): 514–521, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sarma:2000:LSY

[SN00] C. R. Sarma and A. V. V. Nampoothiri. A labeling scheme for young tableaux spanning representations of permutation group S (N). Journal of Computational Chemistry, 21(3): 185–190, February 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sinnecker:2006:QMC

[SN06] Sebastian Sinnecker and Frank Neese. QM/MM calculations with DFT for taking into account protein effects on the EPR and optical spectra of metalloproteins. Plastocyanin as a case study. Journal of Computational Chemistry, 27(12):1463– 1475, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sekiya:2006:RCB

[SNM^{+}06] Masahiro Sekiya, Takeshi Noro, Eisaku Miyoshi, You Osanai, and Toshikatsu Koga. Relativistic correlating basis sets for lanthanide atoms from Ce to Lu. Journal of Computational Chemistry, 27(4):463–470, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Saito:2007:NMD

[SO07] Minoru Saito and Isao Okazaki. A 45-ns molecular dynamics simulation of hemoglobin in water by vectorizing and parallelizing COSMOS90 on the earth simulator: Dynamics of tertiary and quaternary structures. Journal of Computational Chemistry, 28(6):1129–1136, April 30, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Saito:2009:FFP

[SO09] Minoru Saito and Isao Okazaki. Force-field parameters of the Ψ and Φ around glycosidic bonds to oxygen and sulfur atoms. Journal of Computational Chemistry, 30(16):2656–2665, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Saito:2005:FEC

[SOOF05] Minoru Saito, Isao Okazaki, Masayuki Oda, and Ikuo Fujii. A free energy calculation study of the effect of H→F substitution on binding affinity in ligand–antibody interactions. Journal of Computational Chemistry, 26(3):272–282, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sorin:2005:EFF

[SP05] Eric J. Sorin and Vijay S. Pande. Empirical force-field assessment: the interplay between backbone torsions and noncovalent term scaling. Journal of Computational Chemistry, 26 (7):682–690, May 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Salvador:2001:EBI

[SPDS01] Pedro Salvador, Béla Paizs, Miquel Duran, and Sándor Suhai. On the effect of the BSSE on intermolecular potential energy surfaces. Comparison of a priori and a posteriori BSSE correction schemes. Journal of Computational Chemistry, 22 (7):765–786, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stone:2007:AMM

[SPF^{+}07] John E. Stone, James C. Phillips, Peter L. Freddolino, David J. Hardy, Leonardo G. Trabuco, and Klaus Schulten. Accelerating molecular modeling applications with graphics processors. Journal of Computational Chemistry, 28(16): 2618–2640, December 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sapre:2008:CMT

[SPGS08] Nitin S. Sapre, Nilanjana Pancholi, Swagata Gupta, and Neelima Sapre. Computational modeling of tetrahydroimidazo- [4,5,1-jk][1,4]-benzodiazepinone derivatives: an atomistic drug design approach using Kier-Hall electrotopological state (Estate) indices. Journal of Computational Chemistry, 29(11): 1699–1706, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Scheraga:2002:EPB

[SPL^{+}02] Harold A. Scheraga, Jaroslaw Pillardy, Adam Liwo, Jooyoung Lee, Cezary Czaplewski, Daniel R. Ripoll, William J. Wedemeyer, and Yelena A. Arnautova. Evolution of physics-based methodology for exploring the conformational energy landscape of proteins. Journal of Computational Chemistry, 23 (1):28–34, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sirois:2003:DFS

[SPT^{+}03] S. Sirois, E. I. Proynov, J.-F. Truchon, C. M. Tsoukas, and D. R. Salahub. A density functional study of the hydrogenbond network within the HIV-1 protease catalytic site cleft. Journal of Computational Chemistry, 24(9):1110–1119, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Slepoy:2007:SGO

[SPT07] A. Slepoy, M. D. Peters, and A. P. Thompson. Searching for globally optimal functional forms for interatomic potentials using genetic programming with parallel tempering. Journal of Computational Chemistry, 28(15):2465–2471, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Soderhjelm:2009:CDC

[SR09] Pär Söderhjelm and Ulf Ryde. Conformational dependence of charges in protein simulations. Journal of Computational Chemistry, 30(5):750–760, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sigfridsson:2002:RPC

[SRB02] Emma Sigfridsson, Ulf Ryde, and Bruce L. Bush. Restrained point-charge models for disaccharides. Journal of Computational Chemistry, 23(3):351–364, February 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Swart:2006:PAM

[SRB06] Marcel Swart, Ernst Rösler, and F. Matthias Bickelhaupt. Proton affinities of maingroup-element hydrides and noble gases: Trends across the periodic table, structural effects, and DFT validation. Journal of Computational Chemistry, 27(13): 1486–1493, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Seok:2003:MMO

[SRCD03] Chaok Seok, J. B. Rosen, John D. Chodera, and Ken A. Dill. MOPED: Method for optimizing physical energy parameters using decoys. Journal of Computational Chemistry, 24(1):89– 97, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shi:2008:PSI

[SRE08] Huancong Shi, David C. Roettger, and Allan L. L. East. Photochemistry studied with ab initio orbital-correlation and state-correlation plots: Classic cyclobutene ring opening, and the reaction of N_{2} with photoexcited O_{2}. Journal of Computational Chemistry, 29(6):883–891, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Skowronek:2000:WDC

[SRK^{+}00] M. Skowronek, I. Roterman, L. Konieczny, B. Stopa, J. Rybarska, and B. Piekarska. Why do Congo Red, Evans Blue, and Trypan Blue differ in their complexation properties? Journal of Computational Chemistry, 21(8):656–667, June 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sateesh:2007:TDS

[SRS07] B. Sateesh, A. Srinivas Reddy, and G. Narahari Sastry. Towards design of the smallest planar tetracoordinate carbon and boron systems. Journal of Computational Chemistry, 28 (1):335–343, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sauer:2000:CQM

[SS00] Joachim Sauer and Marek Sierka. Combining quantum mechanics and interatomic potential functions in ab initio studies of extended systems. Journal of Computational Chemistry, 21(16):1470–1493, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Siegbahn:2005:QCM

[SS05] Per E. M. Siegbahn and Alexander F. Shestakov. Quantum chemical modeling of CO oxidation by the active site of molybdenum CO dehydrogenase. Journal of Computational Chemistry, 26(9):888–898, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Saravanan:2003:SMM

[SSB^{+}03] Chandra Saravanan, Yihan Shao, Roi Baer, Philip N. Ross, and Martin Head-Gordon. Sparse matrix multiplications for linear scaling electronic structure calculations in an atomcentered basis set using multiatom blocks. Journal of Computational Chemistry, 24(5):618–622, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Swart:2007:ELN

[SSB07] Marcel Swart, Miquel Solà, and F. Matthias Bickelhaupt. Energy landscapes of nucleophilic substitution reactions: a comparison of density functional theory and coupled cluster methods. Journal of Computational Chemistry, 28(9):1551–1560, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Suvire:2006:DFC

[SSBE06] Fernando D. Suvire, Luis N. Santagata, José A. Bombasaro, and Ricardo D. Enriz. Dynamics of flexible cycloalkanes. Ab initio and DFT study of the conformational energy hypersurface of cyclononane. Journal of Computational Chemistry, 27(2):188–202, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Smellie:2003:CAI

[SSHT03] Andrew Smellie, Robert Stanton, Randy Henne, and Steve Teig. Conformational analysis by intersection: CONAN. Journal of Computational Chemistry, 24(1):10–20, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shepard:2002:RCE

[SSL02] Ron Shepard, Isaiah Shavitt, and Hans Lischka. Reducing I/O costs for the eigenvalue procedure in large-scale configuration interaction calculations. Journal of Computational Chemistry, 23(11):1121–1125, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shinoda:2008:EFE

[SSM08] Keiko Shinoda, Wataru Shinoda, and Masuhiro Mikami. Efficient free energy calculation of water across lipid membranes. Journal of Computational Chemistry, 29(12):1912– 1918, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Song:2009:EAE

[SSMW09] Lingchun Song, Jinshuai Song, Yirong Mo, and Wei Wu. An efficient algorithm for energy gradients and orbital optimization in valence bond theory. Journal of Computational Chemistry, 30(3):399–406, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See comment [vLBBR12] and reply [WM12].

Sherrill:2009:ASF

[SSS^{+}09] C. David Sherrill, Bobby G. Sumpter, Mutasem O. Sinnokrot, Michael S. Marshall, Edward G. Hohenstein, Ross C. Walker, and Ian R. Gould. Assessment of standard force field models against high-quality ab initio potential curves for prototypes of π = π, CH/π, and SH/π interactions. Journal of Computational Chemistry, 30(14):2187–2193, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Su:2007:VBS

[SSW^{+}07] Peifeng Su, Lingchun Song, Wei Wu, Philippe C. Hiberty, and Sason Shaik. A valence bond study of the dioxygen molecule.

Journal of Computational Chemistry, 28(1):185–197, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stanger:2001:SIB

[ST01] Amnon Stanger and Elena Tkachenko. Strain induced bond localization in strained aromatic compounds with extended π systems. Journal of Computational Chemistry, 22(13):1377– 1386, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Szabelski:2004:KEM

[ST04] Paweł Szabelski and Julian Talbot. Kinetics and equilibrium of multicomponent adsorption on chiraly templated surfaces. Journal of Computational Chemistry, 25(14):1779–1786, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Studt:2006:TSM

[ST06] Felix Studt and Felix Tuczek. Theoretical, spectroscopic, and mechanistic studies on transition-metal dinitrogen complexes: Implications to reactivity and relevance to the nitrogenase problem. Journal of Computational Chemistry, 27(12):1278– 1291, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [Ano06a].

Stanton:2000:BRB

[Sta00] John F. Stanton. Book review: Reviews in Computational Chemistry, Volume 13. Edited by Kenny B. Lipkowitz and Donald B. Boyd, Wiley-VCH, 1999, 426 pp. $139.00. ISBN 0-471-33135-X. Journal of Computational Chemistry, 21(6): 505–507, April 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sun:2008:SSH

[STC^{+}08] Lili Sun, Shuwei Tang, Yingfei Chang, Zhanliang Wang, and Rongshun Wang. Searching for stable hept-C_{62}X_{2} (X = F, Cl, and Br): Structures and stabilities of heptagon-containing C_{62} halogenated derivatives. Journal of Computational Chemistry, 29(16):2631–2635, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Suarez:2008:POC

[STCJ08] María Suárez, Pablo Tortosa, Javier Carrera, and Alfonso Jaramillo. Pareto optimization in computational protein design with multiple objectives. Journal of Computational Chemistry, 29(16):2704–2711, December 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stern:2004:SAI

[Ste04] Harry A. Stern. Simple algorithm for isothermal–isobaric molecular dynamics. Journal of Computational Chemistry, 25 (5):749–761, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Skeel:2002:MGM

[STH02] Robert D. Skeel, Ismail Tezcan, and David J. Hardy. Multiple grid methods for classical molecular dynamics. Journal of Computational Chemistry, 23(6):673–684, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stortz:2005:CPM

[Sto05] Carlos A. Stortz. Comparative performance of MM3(92) and two TINKER^{T M} MM3 versions for the modeling of carbohydrates. Journal of Computational Chemistry, 26(5):471–483, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schwarzl:2002:CCL

[STSF02] Sonja M. Schwarzl, Thomas B. Tschopp, Jeremy C. Smith, and Stefan Fischer. Can the calculation of ligand binding free energies be improved with continuum solvent electrostatics and an ideal-gas entropy correction? Journal of Computational Chemistry, 23(12):1143–1149, September 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Santos:2006:QQS

[SURG06] Hugo R. R. Santos, Gregori Ujaque, Maria J. Ramos, and José A. N. F. Gomes. QM/QM study of the coverage effects on the adsorption of amino-cyclopentene at the Si(100) surface. Journal of Computational Chemistry, 27(15):1892–1897, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Swart:2001:CAD

[SvDS01] Marcel Swart, Piet Th. van Duijnen, and Jaap G. Snijders. A charge analysis derived from an atomic multipole expansion. Journal of Computational Chemistry, 22(1):79–88, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Storchi:2009:IUD

[SVT09] Loriano Storchi, Giuseppe Vitillaro, and Francesco Tarantelli. Implementation and use of a direct, partially integraldriven non-Dyson propagator method for molecular ionization. Journal of Computational Chemistry, 30(5):818–825, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Suresh:2008:RSW

[SVV^{+}08] Cherumuttathu H. Suresh, Aswathy Mary Vargheese, K. Periya Vijayalakshmi, Neetha Mohan, and Nobuaki Koga. Role of structural water molecule in HIV protease-inhibitor complexes: a QM/MM study. Journal of Computational Chemistry, 29(11):1840–1849, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Szilagyi:2006:ADF

[SW06] Robert K. Szilagyi and Mark A. Winslow. On the accuracy of density functional theory for iron–sulfur clusters. Journal of Computational Chemistry, 27(12):1385–1397, September 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schnupf:2008:DCS

[SWBM08] Udo Schnupf, Julious L. Willett, Wayne B. Bosma, and Frank A. Momany. DFT conformational studies of αmaltotriose. Journal of Computational Chemistry, 29(7):1103– 1112, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sims:2004:COI

[SWM04] Peter A. Sims, Chung F. Wong, and J. Andrew McCammon. Charge optimization of the interface between protein kinases and their ligands. Journal of Computational Chemistry, 25 (11):1416–1429, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sklenar:2006:UIC

[SWR06] Heinz Sklenar, Daniel Wüstner, and Remo Rohs. Using internal and collective variables in Monte Carlo simulations of nucleic acid structures: Chain breakage/closure algorithm and associated Jacobians. Journal of Computational Chemistry, 27 (3):309–315, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sims:2005:RCD

[SWV^{+}05] Peter A. Sims, Chung F. Wong, Danka Vuga, J. Andrew Mc-Cammon, and Bartholomew M. Sefton. Relative contributions of desolvation, inter- and intramolecular interactions to binding affinity in protein kinase systems. Journal of Computational Chemistry, 26(7):668–681, May 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Song:2004:PVB

[SWZS04] Lingchun Song, Wei Wu, Qianer Zhang, and Sason Shaik. A practical valence bond method: a configuration interaction method approach with perturbation theoretic facility. Journal of Computational Chemistry, 25(4):472–478, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shyu:2009:RBU

[SY09] Conrad Shyu and F. Marty Ytreberg. Reducing the bias and uncertainty of free energy estimates by using regression to fit thermodynamic integration data. Journal of Computational Chemistry, 30(14):2297–2304, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sherer:2003:FAM

[SYC03] Edward C. Sherer, Darrin M. York, and Christopher J. Cramer. Fast approximate methods for calculating nucleic acid base pair interaction energies. Journal of Computational Chemistry, 24(1):57–67, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shao:2008:DLS

[SYC08] Xueguang Shao, Xiaoli Yang, and Wensheng Cai. A dynamic lattice searching method with interior operation for unbiased optimization of large Lennard-Jones clusters. Journal of Computational Chemistry, 29(11):1772–1779, August 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shi:2003:ESF

[SYY^{+}03] Shenghua Shi, Lisa Yan, Yang Yang, Jodi Fisher-Shaulsky, and Tom Thacher. An extensible and systematic force field, ESFF, for molecular modeling of organic, inorganic, and organometallic systems. Journal of Computational Chemistry, 24(9):1059–1076, July 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Szabelski:2008:SSH

[Sza08] Paweł Szabelski. Spontaneous segregation on a hybrid chiral surface. Journal of Computational Chemistry, 29(10):1615– 1625, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schultz:2008:BAD

[SZT08] Nathan E. Schultz, Yan Zhao, and Donald G. Truhlar. Benchmarking approximate density functional theory for s/d excitation energies in 3d transition metal cations. Journal of Computational Chemistry, 29(2):185–189, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Suo:2005:PMB

[SZW^{+}05] Bing Suo, Gaohong Zhai, Yubin Wang, Zhenyi Wen, Xiangqian Hu, and Lemin Li. Parallelization of MRCI based on hole-particle symmetry. Journal of Computational Chemistry, 26(1):88–96, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tafipolsky:2007:IPM

[TAS07] Maxim Tafipolsky, Saeed Amirjalayer, and Rochus Schmid. Ab initio parametrized MM3 force field for the metal-organic framework MOF-5. Journal of Computational Chemistry, 28 (7):1169–1176, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Talarico:2000:ECI

[TBG00] Giovanni Talarico, Peter H. M. Budzelaar, and Anton W. Gal. Ethylene coordination, insertion, and chain transfer at a cationic aluminum center: a comparative study with Ab Initio correlated level and density functional methods. Journal of Computational Chemistry, 21(5):398–410, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tubert-Brohman:2004:EPP

[TBGRJ04] Ivan Tubert-Brohman, Cristiano Ruch Werneck Guimarães, Matthew P. Repasky, and William L. Jorgensen. Extension of the PDDG/PM3 and PDDG/MNDO semiempirical molecular orbital methods to the halogens. Journal of Computational Chemistry, 25(1):138–150, January 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tan:2009:AER

[TBSM09] Jeffrey S. Tan, Stephan X. M. Boerrigter, Raymond P. Scaringe, and Kenneth R. Morris. Application of errorranked singular value decomposition for the determination of potential-derived atomic-centered point charges. Journal of Computational Chemistry, 30(5):733–742, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Titmuss:2002:CLS

[TCR^{+}02] Stephen J. Titmuss, Peter L. Cummins, Alistair P. Rendell, Andrey A. Bliznyuk, and Jill E. Gready. Comparison of linearscaling semiempirical methods and combined quantum mechanical/molecular mechanical methods for enzymic reactions. II. An energy decomposition analysis. Journal of Computational Chemistry, 23(14):1314–1322, November 15, 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Turner:2003:SDP

[TCSM03] Joseph V. Turner, David J. Cutler, Ian Spence, and Desmond J. Maddalena. Selective descriptor pruning for QSAR/QSPR studies using artificial neural networks. Journal of Computational Chemistry, 24(7):891–897, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Thompson:2003:PCM

[TCT03] Jason D. Thompson, Christopher J. Cramer, and Donald G. Truhlar. Parameterization of charge model 3 for AM1, PM3, BLYP, and B3LYP. Journal of Computational Chemistry, 24 (11):1291–1304, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tokmachev:2006:EGF

[TD06] A. M. Tokmachev and R. Dronskowski. Electron group functions for the analysis of the electronic structures of molecules. Journal of Computational Chemistry, 27(3):296–308, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tchougreeff:2008:CSC

[TD08] Anderi L. Tchougréeff and Richard Dronskowski. A computational study of the crystal and electronic structure of the room temperature organometallic ferromagnet V(TCNE)_{2}. Journal of Computational Chemistry, 29(13):2220–2233, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tye:2006:CBC

[TDH06] Jesse W. Tye, Marcetta Y. Darensbourg, and Michael B. Hall. Correlation between computed gas-phase and experimentally determined solution-phase infrared spectra: Models of the iron–iron hydrogenase enzyme active site. Journal of Computational Chemistry, 27(12):1454–1462, September 2006. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tsipis:2007:GFE

[TDK07] Constantinos A. Tsipis, Ioannis G. Depastas, and Christos E. Kefalidis. Growth format, electronic architecture, magnetic, and optical properties of aromatic cyclo-Cu_{3}Au_{3} homotops. Journal of Computational Chemistry, 28(11):1893–1908, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Torras:2008:VAG

[TdMSD^{+}08] Juan Torras, Gustavo de M. Seabra, Erik Deumens, S. B. Trickey, and Adrian E. Roitberg. A versatile AMBER– Gaussian QM/MM interface through PUPIL. Journal of Computational Chemistry, 29(10):1564–1573, July 30, 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tatsumi:2004:HMM

[TFN04] Rie Tatsumi, Yoshifumi Fukunishi, and Haruki Nakamura. A hybrid method of molecular dynamics and harmonic dynamics for docking of flexible ligand to flexible receptor. Journal of Computational Chemistry, 25(16):1995–2005, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tiraboschi:2001:ICZ

[TFZRG01] Gilles Tiraboschi, Marie-Claude Fournié-Zaluski, Bernard-Pierre Roques, and Nohad Gresh. Intramolecular chelation of Zn^{2+} by α- and β-mercaptocarboxamides. A parallel ab initio and polarizable molecular mechanics investigation. Assessment of the role of multipole transferability. Journal of Computational Chemistry, 22(10):1038–1047, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tuffery:2005:IGA

[TGD05] Pierre Tuffery, Frédéric Guyon, and Philippe Derreumaux. Improved greedy algorithm for protein structure reconstruction. Journal of Computational Chemistry, 26(5):506–513, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tiraboschi:2000:PIM

[TGGP^{+}00] Gilles Tiraboschi, Nohad Gresh, Claude Giessner-Prettre, Lee G. Pedersen, and David W. Deerfield. Parallel ab initio and molecular mechanics investigation of polycoordinated Zn(II) complexes with model hard and soft ligands: Variations of binding energy and of its components with number and charges of ligands. Journal of Computational Chemistry, 21(12):1011–1039, September 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tejero:2007:PDS

[TGLL07] Ismael Tejero, Àngels González-Lafont, and José M. Lluch. A PM3/d specific reaction parameterization for iron atom in the hydrogen abstraction catalyzed by soybean lipoxygenase-1. Journal of Computational Chemistry, 28(6):997–1005, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tantillo:2002:TSD

[TH02] Dean J. Tantillo and K. N. Houk. Transition state docking: a probe for noncovalent catalysis in biological systems. Application to antibody-catalyzed ester hydrolysis. Journal of Computational Chemistry, 23(1):84–95, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Takahashi:2001:HQM

[THHN01] H. Takahashi, T. Hori, H. Hashimoto, and T. Nitta. A hybrid QM/MM method employing real space grids for QM water in the TIP4P water solvents. Journal of Computational Chemistry, 22(12):1252–1261, September 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tielens:2009:ERF

[Tie09] Frederik Tielens. Exploring the reactivity of framework vanadium, niobium, and tantalum sites in zeolitic materials using DFT reactivity descriptors. Journal of Computational Chemistry, 30(12):1946–1951, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Taylor:2003:FFL

[TJE03] Richard D. Taylor, Philip J. Jewsbury, and Jonathan W. Essex. FDS: Flexible ligand and receptor docking with a continuum solvent model and soft-core energy function. Journal of Computational Chemistry, 24(13):1637–1656, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Takahashi:2003:TSM

[TJM^{+}03] Osamu Takahashi, Masanori Joyabu, Masaki Mitani, Ko Saito, and Suehiro Iwata. Theoretical studies on the molecular dependence of bond dissociation after core excitations II: CH_{3}CO(CH_{2})_{n}CN, n = 0–3. Journal of Computational Chemistry, 24(11):1329–1335, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Togo:2008:FPL

[TK08] Atsushi Togo and Peter Kroll. First-principles lattice dynamics calculations of the phase boundary between β-Si_{3}N_{4} and γ-Si_{3}N_{4} at elevated temperatures and pressures. Journal of Computational Chemistry, 29(13):2255–2259, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tsuneda:2003:RSI

[TKH03] Takao Tsuneda, Muneaki Kamiya, and Kimihiko Hirao. Regional self-interaction correction of density functional theory. Journal of Computational Chemistry, 24(13):1592–1598, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tew:2007:ECM

[TKH07] David P. Tew, Wim Klopper, and Trygve Helgaker. Electron correlation: the many-body problem at the heart of chemistry. Journal of Computational Chemistry, 28(8):1307–1320, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tsuchimochi:2008:ASS

[TKN^{+}08] Takashi Tsuchimochi, Masato Kobayashi, Ayako Nakata, Yutaka Imamura, and Hiromi Nakai. Application of the Sakurai-Sugiura projection method to core-excited-state calculation by time-dependent density functional theory. Journal of Computational Chemistry, 29(14):2311–2316, November 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Takahashi:2001:TPV

[TKS^{+}01] Masae Takahashi, Mitsuo Kira, Kenkichi Sakamoto, Thomas Müller, and Yitzhak Apeloig. Theoretical prediction of vertical transition energies of diaminosilylenes and aminosubstituted disilenes. Journal of Computational Chemistry, 22(13): 1536–1541, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tappura:2000:NSC

[TLKT00] K. Tappura, M. Lahtela-Kakkonen, and O. Teleman. A new soft-core potential function for molecular dynamics applied to the prediction of protein loop conformations. Journal of Computational Chemistry, 21(5):388–397, April 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tielens:2000:QCC

[TLOG00] F. Tielens, W. Langenaeker, A. R. Ocakoglu, and P. Geerlings. Quantum chemical calculation of Henry constants of diatomic molecules in faujasite-type zeolites. Journal of Computational Chemistry, 21(11):909–922, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Torrent:2002:CSR

[TMBM02] Maricel Torrent, Djamaladdin G. Musaev, Harold Basch, and Keiji Morokuma. Computational studies of reaction mechanisms of methane monooxygenase and ribonucleotide reductase. Journal of Computational Chemistry, 23(1):59–76, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tuzun:2000:CIC

[TNS00] Robert E. Tuzun, Donald W. Noid, and Bobby G. Sumpter. Computation of internal coordinates, derivatives, and gradient expressions: torsion and improper torsion. Journal of Computational Chemistry, 21(7):553–561, May 2000. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Torii:2002:ICM

[Tor02] Hajime Torii. Intensity-carrying modes important for vibrational polarizabilities and hyperpolarizabilities of molecules: Derivation from the algebraic properties of formulas and applications. Journal of Computational Chemistry, 23(10):997– 1006, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Totrov:2004:AEG

[Tot04] Maxim Totrov. Accurate and efficient generalized Born model based on solvent accessibility: Derivation and application for Log P octanol/water prediction and flexible peptide docking. Journal of Computational Chemistry, 25(4):609–619, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Thormann:2001:MDF

[TP01a] Michael Thormann and Miquel Pons. Massive docking of flexible ligands using environmental niches in parallelized genetic algorithms. Journal of Computational Chemistry, 22 (16):1971–1982, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Toniolo:2001:ECF

[TP01b] A. Toniolo and M. Persico. Efficient calculation of Franck– Condon factors and vibronic couplings in polyatomics. Journal of Computational Chemistry, 22(9):968–975, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tsodikov:2002:NCP

[TRS02] Oleg V. Tsodikov, M. Thomas Record Jr., and Yuri V. Sergeev. Novel computer program for fast exact calculation of accessible and molecular surface areas and average surface curvature. Journal of Computational Chemistry, 23(6):600–609, April 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Truhlar:2007:EVB

[Tru07] Donald G. Truhlar. Essay: Valence bond theory for chemical dynamics. Journal of Computational Chemistry, 28(1):73–86, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tafipolsky:2005:CRP

[TS05] M. Tafipolsky and R. Schmid. Calculation of rotational partition functions by an efficient Monte Carlo importance sampling technique. Journal of Computational Chemistry, 26(15): 1579–1591, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Torrens:2001:NDI

[TSMNG01] Francisco Torrens, José Sánchez-Marín, and Ignacio Nebot-Gil. New dimension indices for the characterization of the solvent-accessible surface. Journal of Computational Chemistry, 22(5):477–487, April 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Torrent-Sucarrat:2007:QHK

[TSSGS07] Miquel Torrent-Sucarrat, Pedro Salvador, Paul Geerlings, and Miquel Solà. On the quality of the hardness kernel and the Fukui function to evaluate the global hardness. Journal of Computational Chemistry, 28(2):574–583, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Torrent-Sucarrat:2008:HKB

[TSSSG08] Miquel Torrent-Sucarrat, Pedro Salvador, Miquel Solà, and Paul Geerlings. The hardness kernel as the basis for global and local reactivity indices. Journal of Computational Chemistry, 29(7):1064–1072, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tokura:2008:DLS

[TST^{+}08] Seiken Tokura, Takeshi Sato, Takao Tsuneda, Takahito Nakajima, and Kimihiko Hirao. A dual-level state-specific timedependent density-functional theory. Journal of Computational Chemistry, 29(8):1187–1197, June 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tokmachev:2001:SIS

[TT01] A. M. Tokmachev and A. L. Tchougréeff. Semiempirical implementation of strictly localized geminals for analysis of molecular electronic structure. Journal of Computational Chemistry, 22(7):752–764, May 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tsipis:2002:MAD

[TT02] Athanassios C. Tsipis and Constantinos A. Tsipis. Mechanistic aspects of the dehydration and dehydrohalogenation of halo-hydroxyformaldoxime conformers. A quantum chemical model study. Journal of Computational Chemistry, 23(13): 1266–1280, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tokmachev:2005:TPS

[TT05] A. M. Tokmachev and A. L. Tchougréeff. Transferability of parameters of strictly local geminals’ wave function and possibility of sequential derivation of molecular mechanics. Journal of Computational Chemistry, 26(5):491–505, April 15, 2005.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tarasov:2008:NCM

[TT08] D. N. Tarasov and R. P. Tiger. Numerical computations of molecular reactions in associated systems caused by the formation of fractal structures. Journal of Computational Chemistry, 29(2):220–224, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tran:2001:CATa

[TTB01a] Tran Trung Tran, Herbert Treutlein, and Antony W. Burgess. Conformational analysis of thiopeptides: derivation of sp^{2} sulfur parameters for the CFF91 force field. Journal of Computational Chemistry, 22(10):1010–1025, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tran:2001:CATb

[TTB01b] Tran Trung Tran, Herbert Treutlein, and Antony W. Burgess. Conformational analysis of thiopeptides: (φ, ψ) maps of thiosubstituted dipeptides. Journal of Computational Chemistry, 22(10):1026–1037, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Toropov:2009:QME

[TTBM09] Andrey A. Toropov, Alla P. Toropova, Emilio Benfenati, and Alberto Manganaro. QSPR modeling of enthalpies of formation for organometallic compounds by SMART-based optimal descriptors. Journal of Computational Chemistry, 30(15): 2576–2582, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

teVelde:2001:CA

[tVBB^{+}01] G. te Velde, F. M. Bickelhaupt, E. J. Baerends, C. Fonseca Guerra, S. J. A. van Gisbergen, J. G. Snijders, and T. Ziegler. Chemistry with ADF. Journal of Computational Chemistry, 22(9):931–967, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tautermann:2003:EMA

[TVL^{+}03] Christofer S. Tautermann, Andreas F. Voegele, Thomas Loerting, Peter Kaps, and Klaus R. Liedl. Extended method for adiabatic mode reordering. Journal of Computational Chemistry, 24(3):386–395, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tatewaki:2003:GTF

[TW03] Hiroshi Tatewaki and Yoshihiro Watanabe. Gaussian-type function set without prolapse for the Dirac–Fock–Roothaan equation. Journal of Computational Chemistry, 24(15):1823– 1828, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tomlinson:2003:DIS

[TY03] Aimée Tomlinson and David Yaron. Direct INDO/SCI method for excited state calculations. Journal of Computational Chemistry, 24(14):1782–1788, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tamaoki:2005:STF

[TYN05] Mari Tamaoki, Yusuke Yamauchi, and Hiromi Nakai. Shorttime Fourier transform analysis of ab initio molecular dynamics simulation: Collision reaction between CN and C_{4}H_{6}. Journal of Computational Chemistry, 26(5):436–442, April 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Takashima:2002:NPA

[TYO^{+}02] Hajime Takashima, So Yamada, Shigeru Obara, Kunihiro Kitamura, Shinjiro Inabata, Nobuaki Miyakawa, Kazutoshi Tanabe, and Umpei Nagashima. A novel parallel algorithm for large-scale Fock matrix construction with small locally distributed memory architectures: RT parallel algorithm. Journal of Computational Chemistry, 23(14):1337–1346, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Tang:2001:ICS

[TZX01a] Pei Tang, Igor Zubryzcki, and Yan Xu. Ab initio calculation of structures and properties of halogenated general anesthetics: halothane and sevoflurane. Journal of Computational Chemistry, 22(4):436–444, March 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [TZX01b].

Tang:2001:EIC

[TZX01b] Pei Tang, Igor Zubrzycki, and Yan Xu. Erratum: Ab initio calculation of structures and properties of halogenated general anesthetics: halothane and sevoflurane. Journal of Computational Chemistry, 22(7):802, May 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [TZX01a].

Udier-Blagovic:2004:AFE

[UBDPJ04] Marina Udier-Blagović, Patricia Morales De Tirado, Shoshannah A. Pearlman, and William L. Jorgensen. Accuracy of free energies of hydration using CM1 and CM3 atomic charges. Journal of Computational Chemistry, 25(11):1322–1332, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Uchimaru:2003:IIR

[UCT^{+}03] Tadafumi Uchimaru, Asit K. Chandra, Seiji Tsuzuki, Masaaki Sugie, and Akira Sekiya. Ab initio investigation on the reaction path and rate for the gas-phase reaction of HO + H_{2}O ↔ H_{2}O + OH. Journal of Computational Chemistry, 24(13): 1538–1548, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Umeda:2009:PFM

[UIHN09] Hiroaki Umeda, Yuichi Inadomi, Hiroaki Honda, and Umpei Nagashima. Parallel Fock matrix construction program for molecular orbital calculation — specific computer with a hierarchical network. Journal of Computational Chemistry, 30(5): 826–831, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Umeda:2004:IPP

[UKN04] Hiroaki Umeda, Shiro Koseki, and Umpei Nagashima. Improvement of parallelization performance of GAMESS: Global sum and (semi-)direct integral calculation in multireference perturbation calculation. Journal of Computational Chemistry, 25(9):1175–1183, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Umeda:2001:PMP

[UKNS01] Hiroaki Umeda, Shiro Koseki, Umpei Nagashima, and Michael W. Schmidt. Parallelization of multireference perturbation calculations with GAMESS. Journal of Computational Chemistry, 22(12):1243–1251, September 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Upadhyay:2003:ISM

[UM03] D. M. Upadhyay and P. C. Mishra. An ab initio study of microsolvation of LiF in water: Structures and properties of LiF– W_{n}, n = 1–9 complexes. Journal of Computational Chemistry, 24(11):1336–1347, August 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ueno-Noto:2006:RTS

[UNHYT06] Kaori Ueno-Noto, Miki Hara-Yokoyama, and Keiko Takano. Recognition of tandem sialic acid residues by CD38: a theoretical study. Journal of Computational Chemistry, 27(1): 53–60, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Umanskii:2001:TKI

[UNM^{+}01] Stanislav Ya. Umanskii, Konstantin P. Novoselov, Airat Kh. Minushev, Magdalena Siodmiak, Gernot Frenking, and Anatoli A. Korkin. Thermodynamics and kinetics of initial gas phase reactions in chemical vapor deposition of titanium nitride. Theoretical study of TiCl_{4} ammonolysis. Journal of Computational Chemistry, 22(13):1366–1376, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ujaque:2003:CCT

[UTH^{+}03] Gregori Ujaque, Dean J. Tantillo, Yunfeng Hu, K. N. Houk, Kinya Hotta, and Donald Hilvert. Catalysis on the coastline: Theozyme, molecular dynamics, and free energy perturbation analysis of antibody 21D8 catalysis of the decarboxylation of 5-nitro-3-carboxybenzisoxazole. Journal of Computational Chemistry, 24(1):98–110, January 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Urata:2002:AII

[UTM^{+}02] Shingo Urata, Seiji Tsuzuki, Masuhiro Mikami, Akira Takada, Tadafumi Uchimaru, and Akira Sekiya. Analysis of the intermolecular interaction between CH_{3}OCH_{3}, CF_{3}OCH_{3}, CF_{3}OCF_{3}, and CH_{4}: High level ab initio calculations. Journal of Computational Chemistry, 23(15):1472–1479, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Urata:2004:AII

[UTT^{+}04] Shingo Urata, Seiji Tsuzuki, Akira Takada, Masuhiro Mikami, Tadafumi Uchimaru, and Akira Sekiya. Analysis of the intermolecular interactions between CH_{3}OCH_{3}, CF_{3}OCH_{3}, CF_{3}OCF_{3}, and CH_{2}F_{2}, CHF_{3}. Journal of Computational Chemistry, 25(3):447–459, February 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanEijck:2002:CSP

[Van02a] Bouke P. Van Eijck. Crystal structure predictions using five space groups with two independent molecules. The case of small organic acids. Journal of Computational Chemistry, 23 (4):456–462, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanWullen:2002:SDT

[Van02b] Christoph Van Wüllen. Spin densities in two-component relativistic density functional calculations: Noncollinear versus collinear approach. Journal of Computational Chemistry, 23 (8):779–785, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanMourik:2008:CAB

[Van08] Tanja Van Mourik. Comment on “Aromatic-Backbone Interactions in Model α-Helical Peptides” [Palermo et al., J. Comput Chem 2007, 28, 1208]. Journal of Computational Chemistry, 29(1):1–3, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [CPM03] and reply [CPML08b].

Varandas:2009:MPP

[Var09] A. J. C. Varandas. Møller–Plesset perturbation energies and distances for HeC_{20} extrapolated to the complete basis set limit. Journal of Computational Chemistry, 30(3):379–388, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vasilyev:2002:DED

[Vas02] Vladislav Vasilyev. Determination of the effective dielectric constant from the accurate solution of the Poisson equation. Journal of Computational Chemistry, 23(13):1254–1265, October 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanLenthe:2003:OST

[VB03] E. Van Lenthe and E. J. Baerends. Optimized Slater-type basis sets for the elements 1–118. Journal of Computational Chemistry, 24(9):1142–1156, July 15, 2003. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanDamme:2007:NCP

[VB07] Sofie Van Damme and Patrick Bultinck. A new computer program for QSAR-analysis: ARTE-QSAR. Journal of Computational Chemistry, 28(11):1924–1928, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanDamme:2009:CDP

[VB09] Sofie Van Damme and Patrick Bultinck. Conceptual DFT properties-based 3D QSAR: Analysis of inhibitors of the nicotine metabolizing CYP2A6 enzyme. Journal of Computational Chemistry, 30(12):1749–1757, September 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Villa:2000:EWM

[VBGL^{+}00] Jordi Villà, Jörg Bentzien, Àngels González-Lafont, José M. Lluch, Juan Bertran, and Arieh Warshel. Effective way of modeling chemical catalysis: Empirical valence bond picture of role of solvent and catalyst in alkylation reactions. Journal of Computational Chemistry, 21(8):607–625, June 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vila:2009:PDF

[VBS09] Jorge A. Vila, Héctor A. Baldoni, and Harold A. Scheraga. Performance of density functional models to reproduce observed ^{13}C^{α} chemical shifts of proteins in solution. Journal of Computational Chemistry, 30(6):884–892, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Volkov:2004:CEI

[VC04] Anatoliy Volkov and Philip Coppens. Calculation of electrostatic interaction energies in molecular dimers from atomic multipole moments obtained by different methods of electron density partitioning. Journal of Computational Chemistry, 25 (7):921–934, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanZele:2001:VNS

[VCM01] Christopher J. Van Zele, Mark A. Cunningham, and Marvin W. Makinen. Validation of nitroxyl spin-label force-field parameters through molecular dynamics simulations. Journal of Computational Chemistry, 22(10):1113–1123, July 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Venkatesan:2006:JTP

[VDM06] T. S. Venkatesan, K. Deepika, and S. Mahapatra. The Jahn– Teller and pseudo-Jahn–Teller effects in the anion photoelectron spectroscopy of B_{3} cluster. Journal of Computational Chemistry, 27(10):1093–1100, July 30, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanDroogenbroeck:2004:SMI

[vDSSvA04] Joris van Droogenbroeck, Ben Swerts, Lothar Schäfer, and Christian van Alsenoy. Solids modeled by ab initio crystal field methods, part 22: the Fock matrix transformed supermolecule model and the structure determination of D-erythronic acid-3,4-carbonate. Journal of Computational Chemistry, 25(16): 2065–2072, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanderVaart:2000:LSM

[vdVGDM00] Arjan van der Vaart, Valentin Gogonea, Steven L. Dixon, and Kenneth M. Merz, Jr. Linear scaling molecular orbital calculations of biological systems using the semiempirical divide and conquer method. Journal of Computational Chemistry, 21(16):1494–1504, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanEijck:2001:ICSb

[vE01] Bouke P. van Eijck. Ab initio crystal structure predictions for flexible hydrogen-bonded molecules. Part III. Effect of lattice vibrations. Journal of Computational Chemistry, 22(8):816– 826, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vanden-Eijnden:2009:SRT

[VE09] Eric Vanden-Eijnden. Some recent techniques for free energy calculations. Journal of Computational Chemistry, 30(11): 1737–1747, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanEijck:2001:ICSa

[vEMK01] Bouke P. van Eijck, Wijnand T. M. Mooij, and Jan Kroon. Ab initio crystal structure predictions for flexible hydrogenbonded molecules. Part II. Accurate energy minimization. Journal of Computational Chemistry, 22(8):805–815, June 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanZeist:2008:PSY

[VGB08] Willem-Jan Van Zeist, Célia Fonseca Guerra, and F. Matthias Bickelhaupt. PyFrag — streamlining your reaction path analysis. Journal of Computational Chemistry, 29(2):312–315, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vilar:2008:QMA

[VGDSU08] Santiago Vilar, Humberto González-Díaz, Lourdes Santana, and Eugenio Uriarte. QSAR model for alignment-free prediction of human breast cancer biomarkers based on electrostatic potentials of protein pseudofolding HP-lattice networks. Journal of Computational Chemistry, 29(16):2613–2622, December 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanGisbergen:2000:TEE

[vGGB00] S. J. A. van Gisbergen, C. Fonseca Guerra, and E. J. Baerends. Towards excitation energies and (hyper)polarizability calculations of large molecules. Application of parallelization and linear scaling techniques to time-dependent density functional response theory. Journal of Computational Chemistry, 21(16): 1511–1523, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Villar:2005:ALP

[VGGMM05] R. Villar, M. J. Gil, J. I. García, and V. Martínez-Merino. Are AM1 ligand-protein binding enthalpies good enough for use in the rational design of new drugs? Journal of Computational Chemistry, 26(13):1347–1358, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Verma:2007:AAN

[VGO^{+}07] Abhinav Verma, Srinivasa M. Gopal, Jung S. Oh, Kyu H. Lee, and Wolfgang Wenzel. All-atom de novo protein folding with a scalable evolutionary algorithm. Journal of Computational Chemistry, 28(16):2552–2558, December 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vchirawongkwin:2007:QMM

[VHRR07a] Viwat Vchirawongkwin, Thomas S. Hofer, Bernhard R. Randolf, and Bernd M. Rode. Quantum mechanical/molecular mechanical simulations of the Tl(III) ion in water. Journal of Computational Chemistry, 28(6):1057–1067, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vchirawongkwin:2007:TSS

[VHRR07b] Viwat Vchirawongkwin, Thomas S. Hofer, Bernhard R. Randolf, and Bernd M. Rode. Tl(I)-the strongest structurebreaking metal ion in water? A quantum mechanical/ molecular mechanical simulation study. Journal of Computational Chemistry, 28(6):1006–1016, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vemparala:2006:ICI

[VIP^{+}06] Satyavani Vemparala, Ivaylo Ivanov, Vojislava Pophristic, Katrin Spiegel, and Michael L. Klein. Ab initio calculations of intramolecular parameters for a class of arylamide polymers. Journal of Computational Chemistry, 27(6):693–700, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Visscher:2002:DEQ

[Vis02] Lucas Visscher. The Dirac equation in quantum chemistry: Strategies to overcome the current computational problems. Journal of Computational Chemistry, 23(8):759–766, June 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Varekova:2006:OPI

[VK06] R. Svobodová Vařeková and J. Koča. Optimized and parallelized implementation of the electronegativity equalization method and the atom-bond electronegativity equalization method. Journal of Computational Chemistry, 27(3):396–405, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Volkov:2009:BSD

[VKCK09] Anatoliy Volkov, Tibor Koritsanszky, Michal Chodkiewicz, and Harry F. King. On the basis-set dependence of local and integrated electron density properties: Application of a new computer program for quantum-chemical density analysis. Journal of Computational Chemistry, 30(9):1379–1391, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Valdes:2008:EIB

[VKP^{+}08] Haydée Valdés, Vojtĕch Klusák, Michal Piton̆ák, Otto Exner, Ivo Starý, Pavel Hobza, and Lubomír Rulíšek. Evaluation of the intramolecular basis set superposition error in the calculations of larger molecules: [n]helicenes and Phe-Gly-Phe tripeptide. Journal of Computational Chemistry, 29(6):861– 870, April 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vicens:2000:DFS

[VL00] Marie C. Vicens and Gustavo E. López. Density functional studies of cation–water complexes. Journal of Computational Chemistry, 21(1):63–68, January 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanLenthe:2012:LEE

[vLBBR12] J. H. van Lenthe, H. B. Broer-Braam, and Z. Rashid. Letters to the editor: On the efficiency of VBSCF algorithms, a comment on “An efficient algorithm for energy gradients and orbital optimization in valence bond theory”. Journal of Computational Chemistry, 33(8):911–913, March 30, 2012. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [SSMW09].

VanDerSpoel:2005:GFF

[VLH^{+}05] David Van Der Spoel, Erik Lindahl, Berk Hess, Gerrit Groenhof, Alan E. Mark, and Herman J. C. Berendsen. GROMACS: Fast, flexible, and free. Journal of Computational Chemistry, 26(16):1701–1718, December 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vreven:2000:AII

[VM00] Thom Vreven and Keiji Morokuma. On the application of the IMOMO (integrated molecular orbital + molecular orbital) method. Journal of Computational Chemistry, 21(16):1419– 1432, December 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Villa:2002:CFE

[VM02] Alessandra Villa and Alan E. Mark. Calculation of the free energy of solvation for neutral analogs of amino acid side chains. Journal of Computational Chemistry, 23(5):548–553, April 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vila:2007:AMI

[VM07] Antonio Vila and Ricardo A. Mosquera. Atoms in molecules interpretation of the anomeric effect in the O − C − O unit. Journal of Computational Chemistry, 28(9):1516–1530, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vetere:2003:CSQ

[VMA03] Valentina Vetere, Pascale Maldivi, and Carlo Adamo. Comparative studies of quasi-relativistic density functional methods for the description of lanthanide and actinide complexes. Journal of Computational Chemistry, 24(7):850–858, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vreven:2003:GOQ

[VMF^{+}03] Thom Vreven, Keiji Morokuma, Ödön Farkas, H. Bernhard Schlegel, and Michael J. Frisch. Geometry optimization with QM/MM, ONIOM, and other combined methods. I. Microiterations and constraints. Journal of Computational Chemistry, 24(6):760–769, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vasilyev:2002:FPE

[VP02] Vladislav Vasilyev and Enrico O. Purisima. A fast pairwise evaluation of molecular surface area. Journal of Computational Chemistry, 23(7):737–745, May 2002. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Voloshina:2008:CPC

[VP08] Elena Voloshina and Beate Paulus. Cohesive properties of CeN and LaN from first principles. Journal of Computational Chemistry, 29(13):2107–2112, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vitalis:2009:ANC

[VP09] Andreas Vitalis and Rohit V. Pappu. ABSINTH: a new continuum solvation model for simulations of polypeptides in aqueous solutions. Journal of Computational Chemistry, 30 (5):673–699, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Schleyer:1998:ECC

[vRS98] Paul von Ragué Schleyer, editor. Encyclopedia of Computational Chemistry. John Wiley, New York, NY, USA, 1998. ISBN 0-471-96588-X (hardback). xxix + 3429 (5 volumes) pp. LCCN QD39.3.E46 E53 1998. URL http://www.loc.gov/

catdir/description/wiley032/98037164.html.

Schleyer:2001:A

[vRS01] Paul von R. Schleyer. An appreciation. Journal of Computational Chemistry, 22(13):ix–xi, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Valdes:2002:IDS

[VS02] H. Valdés and J. A. Sordo. Ab initio and DFT studies on van der Waals trimers: The OCS · (CO_{2})_{2} complexes. Journal of Computational Chemistry, 23(4):444–455, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vijayalakshmi:2008:TSC

[VS08] K. Periya Vijayalakshmi and Cherumuttathu H. Suresh. Theoretical studies on the carcinogenicity of polycyclic aromatic hydrocarbons. Journal of Computational Chemistry, 29(11): 1808–1817, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vadali:2004:SFG

[VSK^{+}04] Ramkumar V. Vadali, Yan Shi, Sameer Kumar, Laxmikant V. Kale, Mark E. Tuckerman, and Glenn J. Martyna. Scalable fine-grained parallelization of plane-wave–based ab initio molecular dynamics for large supercomputers. Journal of Computational Chemistry, 25(16):2006–2022, December 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vaiana:2003:MMF

[VSW^{+}03] Andrea C. Vaiana, Andreas Schulz, Jürgen Wolfrum, Markus Sauer, and Jeremy C. Smith. Molecular mechanics force field parameterization of the fluorescent probe rhodamine 6G using automated frequency matching. Journal of Computational Chemistry, 24(5):632–639, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Voltz:2008:CGF

[VTT^{+}08] Karine Voltz, Joanna Trylska, Valentina Tozzini, Vandana Kurkal-Siebert, Jörg Langowski, and Jeremy Smith. Coarsegrained force field for the nucleosome from self-consistent multiscaling. Journal of Computational Chemistry, 29(9):1429– 1439, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VandeStreek:2002:IEC

[VVBV02] Jacco Van de Streek, Paul Verwer, Piet Bennema, and Elias Vlieg. On the irrelevance of electrostatics for the crystal structures and polymorphism of long even n-alkanes. Journal of Computational Chemistry, 23(3):365–370, February 2002. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vilar:2007:IDS

[VVS07] Santiago Vilar, M. Carmen Villaverde, and Fredy Sussman. Inhibitor docking screened by the modified SAFE p scoring function: Application to cyclic urea HIV-1 PR inhibitors. Journal of Computational Chemistry, 28(13):2216–2225, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vasilkoski:2000:GPF

[VW00] Zlatko Vasilkoski and David L. Weaver. A generator of protein folding kinetics states for the diffusion–collision model. Journal of Computational Chemistry, 21(11):923–932, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Valderrama:2003:EPA

[VW03] E. Valderrama and R. J. Wheatley. An environmental pseudopotential approach to molecular interactions: Implementation in MOLPRO. Journal of Computational Chemistry, 24 (16):2075–2082, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vasilkoski:2004:DCM

[VW04] Zlatko Vasilkoski and David L. Weaver. Diffusion–collision model algorithms for protein folding kinetics. Journal of Computational Chemistry, 25(8):1101–1107, June 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

vanWullen:2006:NIC

[vW06] Christoph van Wüllen. Numerical instabilities in the computation of pseudopotential matrix elements. Journal of Computational Chemistry, 27(2):135–141, January 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vyazovkin:2001:MII

[Vya01] Sergey Vyazovkin. Modification of the integral isoconversional method to account for variation in the activation energy. Journal of Computational Chemistry, 22(2):178–183, January 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vizcarra:2008:IPD

[VZM^{+}08] Christina L. Vizcarra, Naigong Zhang, Shannon A. Marshall, Ned S. Wingreen, Chen Zeng, and Stephen L. Mayo. An improved pairwise decomposable finite-difference Poisson– Boltzmann method for computational protein design. Journal of Computational Chemistry, 29(7):1153–1162, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

VanLenthe:2006:SSC

[VZVG06] J. H. Van Lenthe, R. Zwaans, H. J. J. Van Dam, and M. F. Guest. Starting SCF calculations by superposition of atomic densities. Journal of Computational Chemistry, 27(8):926– 932, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2009:PPT

[Wan09] Yang Wang. Perfect planar tetracoordinate carbon in neutral unsaturated hydrocarbon cages: a new strategy utilizing three-dimensional electron delocalization. Journal of Computational Chemistry, 30(13):2122–2126, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wagoner:2004:EJW

[WB04a] Jason Wagoner and Nathan A. Baker. Erratum: Jason Wagoner and Nathan A. Baker, “Solvation forces on biomolecular structures: A comparison of explicit solvent and Poisson– Boltzmann models,” Journal of Computational Chemistry (2004) 25(13) 1623–1629. Journal of Computational Chemistry, 25(14):1801, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [WB04b, WB05].

Wagoner:2004:SFB

[WB04b] Jason Wagoner and Nathan A. Baker. Solvation forces on biomolecular structures: a comparison of explicit solvent and Poisson–Boltzmann models. Journal of Computational Chemistry, 25(13):1623–1629, October 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See errata [WB04a, WB05].

Wagoner:2005:ESF

[WB05] Jason Wagoner and Nathan A. Baker. Erratum to “Solvation forces on biomolecular structures: A comparison of explicit solvent and Poisson–Boltzmann models,” by Jason Wagoner and Nathan A. Baker. Journal of Computational Chemistry, 26(8):863, June 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [WB04b, WB04a].

Waller:2007:VCG

[WB07] Mark P. Waller and Michael Bühl. Vibrational corrections to geometries of transition metal complexes from density functional theory. Journal of Computational Chemistry, 28(9): 1531–1537, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Whitehead:2003:TAE

[WBSR03] C. E. Whitehead, C. M. Breneman, N. Sukumar, and M. D. Ryan. Transferable atom equivalent multicentered multipole expansion method. Journal of Computational Chemistry, 24 (4):512–529, March 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Winget:2004:EFB

[WC04] Paul Winget and Timothy Clark. Enthalpies of formation from B3LYP calculations. Journal of Computational Chemistry, 25 (5):725–733, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2008:QMS

[WC08] Ruibo Wu and Zexing Cao. QM/MM study of catalytic methyl transfer by the N^{5}-glutamine SAM-dependent methyltransferase and its inhibition by the nitrogen analogue of coenzyme. Journal of Computational Chemistry, 29(3):350–357, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wan:2009:CSC

[WC09] Shunzhou Wan and Peter V. Coveney. A comparative study of the COX-1 and COX-2 isozymes bound to lipid membranes. Journal of Computational Chemistry, 30(7):1038–1050, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Walker:2008:IFA

[WCC08] Ross C. Walker, Michael F. Crowley, and David A. Case. The implementation of a fast and accurate QM/MM potential method in AMBER. Journal of Computational Chemistry, 29 (7):1019–1031, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wan:2004:LSM

[WCF04] Shunzhou Wan, Peter Coveney, and Darren R. Flower. Largescale molecular dynamics simulations of HLA-A*0201 complexed with a tumor-specific antigenic peptide: Can the α3 and β_{2}m domains be neglected? Journal of Computational Chemistry, 25(15):1803–1813, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Witek:2002:ISA

[WCFH02] Henryk A. Witek, Yoong-Kee Choe, James P. Finley, and Kimihiko Hirao. Intruder state avoidance multireference Møller–Plesset perturbation theory. Journal of Computational Chemistry, 23(10):957–965, July 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2009:RMP

[WCHW09] Meiyan Wang, Lin Cheng, Bo Hong, and Zhijian Wu. Reaction mechanism of palladium-catalyzed silastannation of allenes by density functional theory. Journal of Computational Chemistry, 30(9):1521–1531, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2000:HWD

[WCK00] Junmei Wang, Piotr Cieplak, and Peter A. Kollman. How well does a restrained electrostatic potential (RESP) model perform in calculating conformational energies of organic and biological molecules? Journal of Computational Chemistry, 21(12):1049–1074, September 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2005:TSN

[WCL05] Hui-Yin Wu, Ajay Chaudhari, and Shyi-Long Lee. Theoretical studies on nonlinear optical properties of formaldehyde oligomers by ab initio and density functional theory methods. Journal of Computational Chemistry, 26(15):1543–1564, November 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2009:OBC

[WCS09] Xia Wu, Wensheng Cai, and Xueguang Shao. Optimization of bimetallic Cu–Au and Ag–Au clusters by using a modified adaptive immune optimization algorithm. Journal of Computational Chemistry, 30(13):1992–2000, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2008:DFS

[WCW08] M. Y. Wang, L. Cheng, and Z. J. Wu. Density functional study on the reaction mechanism of palladium-catalyzed addition of cyanoboranes to alkynes. Journal of Computational Chemistry, 29(11):1825–1839, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2004:SEA

[WD04] Zhi-Xiang Wang and Yong Duan. Solvation effects on alanine dipeptide: a MP2/cc-pVTZ//MP2/6-31G** study of (Φ, Ψ) energy maps and conformers in the gas phase, ether, and water. Journal of Computational Chemistry, 25(14):1699–1716, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wolff:2008:FPM

[WD08] Holger Wolff and Richard Dronskowski. First-principles and molecular-dynamics study of structure and bonding in perovskite-type oxynitrides ABO_{2}N (A = Ca, Sr, Ba; B = Ta, Nb). Journal of Computational Chemistry, 29(13):2260–2267, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2006:TMS

[WDS06] Jian Wang, Yi-Hong Ding, and Chia-Chung Sun. Theoretical mechanistic study on the radical–molecule reactions of cyanomethylidyne with PH_{3}, H_{2}S, and HCl. Journal of Computational Chemistry, 27(14):1756–1764, November 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2006:GRM

[WDWS06] Jian Wang, Yi-Hong Ding, Gong-Bing Wu, and Chia-Chung Sun. Gaseous reaction mechanism of C_{2}F radical with water. Journal of Computational Chemistry, 27(3):363–367, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2002:DID

[WDX^{+}02] Yang Wu, Yi-Hong Ding, Jing-Fa Xiao, Ze-Sheng Li, Xu-Ri Huang, and Chia-Chung Sun. Direct ab initio dynamics study on the rate constants and kinetic isotope effect for the reactions of H atoms with GeD_{n} (CH_{3})_{4−n} (n = 1–4). Journal of Computational Chemistry, 23(14):1366–1374, November 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2006:TSS

[WDXS06] Qiang Wang, Yi-Hong Ding, Hong-Bin Xie, and Chia-Chung Sun. Theoretical study on structures and stabilities of [H,Ge,C,N]. Journal of Computational Chemistry, 27(4):505– 514, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2007:TSM

[WDZS07] Jian Wang, Yi-Hong Ding, Shao-Wen Zhang, and Chia-Chung Sun. Theoretical study on the methyl radical with chlorinated methyl radicals CH_{3−n}Cl_{n} (n = 1, 2, 3) and CCl_{2}. Journal of Computational Chemistry, 28(5):865–876, April 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Whitehead:2001:MDS

[WEE01] Lewis Whitehead, Colin M. Edge, and Jonathan W. Essex. Molecular dynamics simulation of the hydrocarbon region of a biomembrane using a reduced representation model. Journal of Computational Chemistry, 22(14):1622–1633, November 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Weigend:2008:HFE

[Wei08] Florian Weigend. Hartree–Fock exchange fitting basis sets for H to Rn. Journal of Computational Chemistry, 29(2):167– 175, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2001:APP

[WFHP01] Bing Wang, Ulrich Fleischer, James F. Hinton, and Peter Pulay. Accurate prediction of proton chemical shifts. I. Substituted aromatic hydrocarbons. Journal of Computational Chemistry, 22(16):1887–1895, December 2001. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wong:2008:TCM

[WFR08] Bryan M. Wong, Maria M. Fadri, and Sumathy Raman. Thermodynamic calculations for molecules with asymmetric internal rotors. II. Application to the 1,2-dihaloethanes. Journal of Computational Chemistry, 29(3):481–487, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Widjaja:2002:PCS

[WG02] Effendi Widjaja and Marc Garland. Pure component spectral reconstruction from mixture data using SVD, global entropy minimization, and simulated annealing. Numerical investigations of admissible objective functions using a synthetic 7species data set. Journal of Computational Chemistry, 23(9): 911–919, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wheatley:2008:TDC

[Whe08] Richard J. Wheatley. Time-dependent coupled-cluster calculations of polarizabilities and dispersion energy coefficients. Journal of Computational Chemistry, 29(3):445–450, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wittayanarakul:2008:APP

[WHF08] Kitiyaporn Wittayanarakul, Supot Hannongbua, and Michael Feig. Accurate prediction of protonation state as a prerequisite for reliable MM-PB(GB)SA binding free energy calculations of HIV-1 protease inhibitors. Journal of Computational Chemistry, 29(5):673–685, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Woodcock:2007:IQC

[WHG^{+}07] H. Lee Woodcock III, Milan Hodošček, Andrew T. B. Gilbert, Peter M. W. Gill, Henry F. Schaefer III, and Bernard R. Brooks. Interfacing Q-Chem and CHARMM to perform QM/ MM reaction path calculations. Journal of Computational Chemistry, 28(9):1485–1502, July 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2006:DDS

[WHH^{+}06] Li-Cheng Wu, Jorng-Tzong Horng, Shir-Ly Huang, Hsien-Da Huang, and Baw-Jhiune Liu. Detection of discriminative sequence motifs in proteins obtained from prokaryotes grown at various temperatures. Journal of Computational Chemistry, 27(6):798–808, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2002:APP

[WHP02] Bing Wang, James F. Hinton, and Peter Pulay. Accurate prediction of proton chemical shifts. II. Peptide analogues. Journal of Computational Chemistry, 23(4):492–497, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Whitnell:2008:CMV

[WHRG08] Robert M. Whitnell, Dow P. Hurst, Patricia H. Reggio, and Frank Guarnieri. Conformational memories with variable bond angles. Journal of Computational Chemistry, 29(5):741– 752, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wiberg:2004:BSE

[Wib04] Kenneth B. Wiberg. Basis set effects on calculated geometries: 6-311++G** vs. aug-cc-pVDZ. Journal of Computational Chemistry, 25(11):1342–1346, August 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Williams:2001:IIFa

[Wil01a] Donald E. Williams. Improved intermolecular force field for crystalline oxohydrocarbons including O − H⋯O hydrogen bonding. Journal of Computational Chemistry, 22(1):1–20, January 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Williams:2001:IIFb

[Wil01b] Donald E. Williams. Improved intermolecular force field for molecules containing H, C, N, and O atoms, with application to nucleoside and peptide crystals. Journal of Computational Chemistry, 22(11):1154–1166, August 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2000:VBC

[WJ00] Jian Wu and Yuansheng Jiang. The valence bond calculations for conjugated hydrocarbons having 24–28 π-electrons. Journal of Computational Chemistry, 21(10):856–869, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2008:TSH

[WJX^{+}08] Yong Wu, Lu Jin, Ying Xue, Dai Qian Xie, Chan Kyung Kim, Yong Guo, and Guo Sen Yan. Theoretical study on the hydrolysis mechanism of N, N-dimethyl–N’-(2-oxo-1, 2dihydro-pyrimidinyl)formamidine: Water-assisted mechanism and cluster-continuum model. Journal of Computational Chemistry, 29(8):1222–1232, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2001:APF

[WK01] Junmei Wang and Peter A. Kollman. Automatic parameterization of force field by systematic search and genetic algorithms. Journal of Computational Chemistry, 22(12):1219– 1228, September 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Watanabe:2001:CXR

[WKYU01] Noboru Watanabe, Yohei Kamata, Kota Yamauchi, and Yasuo Udagawa. Calculation of X-ray scattering intensities by means of the coupled cluster singles and doubles model. Journal of Computational Chemistry, 22(13):1315–1320, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wells:2000:CSE

[WL00] M. C. Wells and R. R. Lucchese. Computing the superexchange electronic factor of electron transfer theory using a grid-based numerical method. Journal of Computational Chemistry, 21(14):1262–1273, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2002:AEG

[WL02] Fan Wang and Lemin Li. Analytical energy gradient evaluation in relativistic and nonrelativistic density functional calculations. Journal of Computational Chemistry, 23(9):920–927, July 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2004:NEP

[WL04] Fan Wang and Lemin Li. Numerical examination of performance of some exchange-correlation functionals for molecules containing heavy elements. Journal of Computational Chemistry, 25(5):669–677, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Walsh:2009:MBW

[WL09a] T. R. Walsh and T. Liang. A multipole-based water potential with implicit polarization for biomolecular simulations. Journal of Computational Chemistry, 30(6):893–899, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2009:NBT

[WL09b] Jin Wang and Jin Huai Liu. Novel bi-transition metallic encapsulated naphthalene-like Si_{20} prismatic cage: a DFT investigation. Journal of Computational Chemistry, 30(7):1103–1110, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2001:TSP

[WLL01] Li Jie Wang, Se Li, and Qian Shu Li. Theoretical studies on a possible synthesis reaction pathway on N_{8} (C_{S}) clusters. Journal of Computational Chemistry, 22(13):1334–1339, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2003:DLD

[WLL^{+}03] Jia-Yan Wu, Jing-Yao Liu, Ze-Sheng Li, Xu-Ri Huang, and Chia-Chung Sun. Dual-level direct dynamics studies for the reactions of dimethyl ether with hydrogen atom and methyl radical. Journal of Computational Chemistry, 24(5):593–600, April 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2007:IDD

[WLL07a] Ying Wang, Jing-Yao Liu, and Ze-Sheng Li. Ab initio direct dynamics studies on the reactions of chlorine atom with CH_{3−n}F_{n}CH_{2}OH (n = 1–3). Journal of Computational Chemistry, 28(16):2517–2530, December 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2007:TSR

[WLL^{+}07b] Ying Wang, Jing-Yao Liu, Ze-Sheng Li, Li Wang, and Chia-Chung Sun. Theoretical study and rate constant calculation for reaction of CF_{3}CH_{2}OH with OH. Journal of Computational Chemistry, 28(4):802–810, March 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2004:DID

[WLLS04] Li Wang, Jing-Yao Liu, Ze-Sheng Li, and Chia-Chung Sun. Direct ab initio dynamics studies on the hydrogen-abstraction reactions of OH radicals with HOX (X = F, Cl, and Br). Journal of Computational Chemistry, 25(4):558–564, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2005:TSR

[WLLS05] Li Wang, Jing-Yao Liu, Ze-Sheng Li, and Chia-Chung Sun. Theoretical study and rate constant calculation for the reactions of SH (SD) with Cl_{2}, Br_{2}, and BrCl. Journal of Computational Chemistry, 26(2):184–193, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2005:PCS

[WLPF05] Zhigang Wang, Keyan Lian, Shoufu Pan, and Xianhong Fan. A path from I_{h} to C_{1} symmetry for C_{20} cage molecule. Journal of Computational Chemistry, 26(12):1279–1283, September 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2005:RBA

[WLX^{+}05] Gui-Chang Wang, Jun Li, Xiu-Fang Xu, Rui-Fang Li, and Junji Nakamura. The relationship between adsorption energies of methyl on metals and the metallic electronic properties: a first-principles DFT study. Journal of Computational Chemistry, 26(9):871–878, July 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wei:2007:TIE

[WLZ^{+}07] Zi-Zhang Wei, Bu-Tong Li, Hong-Xing Zhang, Chia-Chung Sun, and Ke-Li Han. A theoretical investigation of the excited states of OCLO radical, cation, and anion using the CASSCF/CASPT2 method. Journal of Computational Chemistry, 28(2):467–477, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wysokinski:2001:PDD

[WM01] Rafał Wysokiński and Danuta Michalska. The performance of different density functional methods in the calculation of molecular structures and vibrational spectra of platinum(II) antitumor drugs: cisplatin and carboplatin. Journal of Computational Chemistry, 22(9):901–912, July 15, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Witek:2004:SSV

[WM04] Henryk A. Witek and Keiji Morokuma. Systematic study of vibrational frequencies calculated with the self-consistent charge density functional tight-binding method. Journal of Computational Chemistry, 25(15):1858–1864, November 30, 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wassenaar:2006:EBS

[WM06] Tsjerk A. Wassenaar and Alan E. Mark. The effect of box shape on the dynamic properties of proteins simulated under periodic boundary conditions. Journal of Computational Chemistry, 27(3):316–325, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2012:LER

[WM12] Wei Wu and Yirong Mo. Letters to the editor: Reply to comment on the paper “An Efficient Algorithm for Energy Gradients and Orbital Optimization in Valence Bond Theory”. Journal of Computational Chemistry, 33(8):914–915, March 30, 2012. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Walker:2007:CBS

[WMGK07] Ross C. Walker, Ian P. Mercer, Ian R. Gould, and David R. Klug. Comparison of basis set effects and the performance of ab initio and DFT methods for probing equilibrium fluctuations. Journal of Computational Chemistry, 28(2):478–490, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2001:HRE

[WMRW^{+}01] F. Wang, H. Mackenzie-Ross, D. A. Winkler, I. E. McCarthy, L. Campbell, and M. J. Brunger. A high-resolution electron momentum spectroscopy and density functional theory study into the complete valence electronic structure of allene. Journal of Computational Chemistry, 22(13):1321–1333, October 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Whitten:2006:EFD

[WMS06] Andrew E. Whitten, Joshua J. McKinnon, and Mark A. Spackman. Electric field-derived point charges to mimic the electrostatics in molecular crystals. Journal of Computational Chemistry, 27(10):1063–1070, July 30, 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2003:CLD

[WMW03] Yi-Gui Wang, Cherif Matta, and Nick Henry Werstiuk. Comparison of localization and delocalization indices obtained with Hartree–Fock and conventional correlated methods: Effect of Coulomb correlation. Journal of Computational Chemistry, 24(14):1720–1729, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [WMW04].

Wang:2004:ECL

[WMW04] Yi-Gui Wang, Cherif F. Matta, and Nick Henry Werstiuk. Erratum: Comparison of localization and delocalization indices obtained with Hartree–Fock and conventional correlated methods: Effect of Coulomb correlation. Journal of Computational Chemistry, 25(2):309, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [WMW03].

Witek:2003:MPT

[WNH03] Henryk A. Witek, Haruyuki Nakano, and Kimihiko Hirao. Multireference perturbation theory with optimized partitioning. II. Applications to molecular systems. Journal of Computational Chemistry, 24(12):1390–1400, September 2003. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wladkowski:2003:HCP

[WOC^{+}03] Brian D. Wladkowski, Paul Ostazeski, Sarah Chenoweth, Steven J. Broadwater, and Morris Krauss. Hydrolysis of cyclic phosphates by ribonuclease A: a computational study using a simplified ab initio quantum model. Journal of Computational Chemistry, 24(14):1803–1811, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wong:2009:NIS

[Won09] Bryan M. Wong. Noncovalent interactions in supramolecular complexes: a study on corannulene and the double concave buckycatcher. Journal of Computational Chemistry, 30(1): 51–56, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Woods:2001:BRB

[Woo01] Robert J. Woods. Book review: Computational Chemistry. A Practical Guide for Applying Techniques to Real World Problems, By David Young. Journal of Computational Chemistry, 22(16):2040, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wouters:2000:UTD

[Wou00] Johan Wouters. Use of theoretical descriptors to characterize cation–π binding sites in (macro)molecules. Journal of Computational Chemistry, 21(10):847–855, July 30, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wannere:2007:ESO

[WPH^{+}07] Chaitanya S. Wannere, Ankan Paul, Rainer Herges, K. N. Houk, Henry F. Schaefer III, and Paul von Ragué Schleyer. The existence of secondary orbital interactions. Journal of Computational Chemistry, 28(1):344–361, January 15, 2007.

CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Witter:2002:CSD

[WPS02] Raiker Witter, Wolfram Prie, and Ulrich Sternberg. Chemical shift driven geometry optimization. Journal of Computational Chemistry, 23(2):298–305, January 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qu:2002:TSM

[wQZsLyZ02] Zheng wang Qu, Hui Zhu, Ze sheng Li, and Qi yuan Zhang. Theoretical study on the mechanism of the gas-phase reaction of diborane(3) anion with carbon disulfide. Journal of Computational Chemistry, 23(3):414–419, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wong:2007:TCM

[WR07] Bryan M. Wong and Sumathy Raman. Thermodynamic calculations for molecules with asymmetric internal rotors — application to 1,3-butadiene. Journal of Computational Chemistry, 28(4):759–766, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2003:DAG

[WRBV03] Guosheng Wu, Daniel H. Robertson, Charles L. Brooks III, and Michal Vieth. Detailed analysis of grid-based molecular docking: a case study of CDOCKER — a CHARMm-based MD docking algorithm. Journal of Computational Chemistry, 24(13):1549–1562, October 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Waller:2006:HDF

[WRP^{+}06] Mark P. Waller, Arturo Robertazzi, James A. Platts, David E. Hibbs, and Peter A. Williams. Hybrid density functional theory for π-stacking interactions: Application to benzenes, pyridines, and DNA bases. Journal of Computational Chemistry, 27(4):491–504, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Whitfield:2002:GSG

[WS02a] T. W. Whitfield and John E. Straub. Gravitational smoothing as a global optimization strategy. Journal of Computational Chemistry, 23(11):1100–1103, August 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wojciechowski:2002:DSL

[WS02b] Marek Wojciechowski and Jeffrey Skolnick. Docking of small ligands to low-resolution and theoretically predicted receptor structures. Journal of Computational Chemistry, 23(1):189– 197, January 15, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wada:2005:QCM

[WS05a] Mitsuhito Wada and Minoru Sakurai. A quantum chemical method for rapid optimization of protein structures. Journal of Computational Chemistry, 26(2):160–168, January 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wennmohs:2005:DMM

[WS05b] F. Wennmohs and M. Schindler. Development of a multipoint model for sulfur in proteins: a new parametrization scheme to reproduce high-level ab initio interaction energies. Journal of Computational Chemistry, 26(3):283–293, February 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wroblewska:2007:CPB

[WS07] Liliana Wroblewska and Jeffrey Skolnick. Can a physics-based, all-atom potential find a protein’s native structure among misfolded structures? I. Large scale AMBER benchmarking. Journal of Computational Chemistry, 28(12):2059–2066, September 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wright:2009:SCC

[WSC09] James S. Wright, Hooman Shadnia, and Leonid L. Chepelev. Stability of carbon-centered radicals: Effect of functional groups on the energetics of addition of molecular oxygen. Journal of Computational Chemistry, 30(7):1016–1026, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wei:2008:ATM

[WSM^{+}08] Y. Wei, T. Singer, H. Mayr, G. N. Sastry, and H. Zipse. Assessment of theoretical methods for the calculation of methyl cation affinities. Journal of Computational Chemistry, 29(2): 291–297, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wei:2009:PCT

[WSM^{+}09] Yin Wei, Bandaru Sateesh, Boris Maryasin, G. N. Sastry, and Hendrik Zipse. The performance of computational techniques in locating the charge separated intermediates in organocatalytic transformations. Journal of Computational Chemistry, 30(16):2617–2624, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Watanabe:2006:RGB

[WTKM06] Yoshihiro Watanabe, Hiroshi Tatewaki, Toshikatsu Koga, and Osamu Matsuoka. Relativistic Gaussian basis sets for molecular calculations: Fully optimized single-family exponent basis sets for H Hg. Journal of Computational Chemistry, 27(1): 48–52, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2006:ESM

[Wu06] Zhijian Wu. Electronic structures of 3d-metal mononitrides. Journal of Computational Chemistry, 27(3):267–276, February 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2003:PEM

[WW03] Yi-Gui Wang and Nick Henry Werstiuk. A practical and efficient method to calculate AIM localization and delocalization indices at post-HF levels of theory. Journal of Computational Chemistry, 24(3):379–385, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wiberg:2004:CES

[WW04] Kenneth B. Wiberg and Yi-Gui Wang. Conformational energies for 2-substituted butanes. Journal of Computational Chemistry, 25(9):1127–1132, July 15, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2004:DTG

[WWC^{+}04] Junmei Wang, Romain M. Wolf, James W. Caldwell, Peter A. Kollman, and David A. Case. Development and testing of a general AMBER force field. Journal of Computational Chemistry, 25(9):1157–1174, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [?].

Wang:2005:EJW

[WWC^{+}05] Junmei Wang, Romain M. Wolf, James W. Caldwell, Peter A. Kollman, and David A. Case. Erratum: Junmei Wang, Romain M. Wolf, James W. Caldwell, Peter A. Kollman, and David A. Case, “Development and testing of a general AM-BER force field” Journal of Computational Chemistry (2004) 25(9) 1157–1174. Journal of Computational Chemistry, 26(1): 114, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See [WWC^{+}04].

Wang:2009:SDE

[WWL^{+}09] Rong Wang, Jian-Sheng Wang, Gui-Rong Liu, Jongyoon Han, and Yu-Zong Chen. Simulation of DNA electrophoresis in systems of large number of solvent particles by coarse-grained hybrid molecular dynamics approach. Journal of Computational Chemistry, 30(4):505–513, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2007:ESC

[WWS07] Z. J. Wu, M. Y. Wang, and Z. M. Su. Electronic structures and chemical bonding in diatomic ScX to ZnX (X = S, Se, Te). Journal of Computational Chemistry, 28(3):703–714, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2008:DSC

[WWT08] Xueye Wang, Hengliang Wang, and Yuanqiang Tan. DFT study of the cryptand and benzocryptand and their complexes with alkali metal cations: Li^{+}, Na^{+}, K^{+}. Journal of Computational Chemistry, 29(9):1423–1428, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2009:APH

[WX09] Jianming Wu and Xin Xu. Accurate prediction of heats of formation by a combined method of B3LYP and neural network correction. Journal of Computational Chemistry, 30(9): 1424–1444, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2008:CSD

[WXK08] Yong Wu, Ying Xue, and Chan Kyung Kim. Computational studies on the dimers and the thermal dimerization of norbornadiene. Journal of Computational Chemistry, 29(8):1250– 1258, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wu:2003:ISP

[WXX03] Yong Wu, Daiqian Xie, and Ying Xue. Ab initio studies for the photodissociation mechanism of hydroxyacetone. Journal of Computational Chemistry, 24(8):931–938, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2009:DLD

[WyLG^{+}09] Li Wang, Jing yao Liu, Hong Gao, Su qin Wan, and Ze sheng Li. Dual-level direct dynamics studies for the reactions of OH radical with bromine-substituted ethanes. Journal of Computational Chemistry, 30(4):611–620, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2006:SBO

[WZW^{+}06] Zhi-Xiang Wang, Wei Zhang, Chun Wu, Hongxing Lei, Piotr Cieplak, and Yong Duan. Strike a balance: Optimization of backbone torsion parameters of AMBER polarizable force field for simulations of proteins and peptides. Journal of Computational Chemistry, 27(6):781–790, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic). See erratum [Ano06c].

Wu:2007:SSC

[WZXY07] Anan Wu, Ying Zhang, Xin Xu, and Yijing Yan. Systematic studies on the computation of nuclear magnetic resonance shielding constants and chemical shifts: the density functional models. Journal of Computational Chemistry, 28(15):2431– 2442, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wan:2004:QSA

[WZY04] Jian Wan, Li Zhang, and Guangfu Yang. Quantitative structure–activity relationships for phenyl triazolinones of protoporphyrinogen oxidase inhibitors: a density functional theory study. Journal of Computational Chemistry, 25(15):1827– 1832, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wang:2009:FPS

[WZZ^{+}09] Lu Wang, Jijun Zhao, Zhen Zhou, S. B. Zhang, and Zhongfang Chen. First-principles study of molecular hydrogen dissociation on doped Al_{12}X (X = B, Al, C, Si, P, Mg, and Ca) clusters. Journal of Computational Chemistry, 30(15):2509– 2514, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xia:2008:BSZ

[XB08] Sheng-Qing Xia and Svilen Bobev. Are Ba_{11}Cd_{6}Sb_{12} and Sr_{11}Cd_{6}Sb_{12} zintl phases or not? A density-functional theory study. Journal of Computational Chemistry, 29(13):2125– 2133, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2006:RRC

[XDS06a] Hong-Bin Xie, Yi-Hong Ding, and Chia-Chung Sun. Radical reaction C_{3}H + NO: a mechanistic study. Journal of Computational Chemistry, 27(5):641–660, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2006:YRM

[XDS06b] Hong-Bin Xie, Yi-Hong Ding, and Chia-Chung Sun. Ylidic radical–molecule reactions. Journal of Computational Chemistry, 27(5):545–551, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2006:TSI

[XFF06] Ju Xie, Dacheng Feng, and Shengyu Feng. Theoretical study on the isomeric structures and the stability of silylenoid (Tsi)Cl_{2}SiLi (Tsi = C(SiMe_{3})_{3}). Journal of Computational Chemistry, 27(8):933–940, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xu:2008:USS

[XK08] Jianxiao Xu and Holger Kleinke. Unusual Sb–Sb bonding in high temperature thermoelectric materials. Journal of Computational Chemistry, 29(13):2134–2143, October 2008. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xue:2005:DSM

[XKG^{+}05] Ying Xue, Chan Kyung Kim, Yong Guo, Dai Qian Xie, and Guo Sen Yan. DFT study and Monte Carlo simulation on proton transfers of 2-amino-2-oxazoline, 2-amino-2-thiazoline, and 2-amino-2-imidazoline in the gas phase and in water. Journal of Computational Chemistry, 26(10):994–1005, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xue:2003:TSG

[XKKL03] Ying Xue, Chun Ho Kang, Chan Kyung Kim, and Ikchoon Lee. Theoretical studies on the gas-phase pyrolysis of 2phenoxycarboxylic acids: an ONIOM approach. Journal of Computational Chemistry, 24(8):963–972, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2002:TSG

[XL02] Li Xie and Haiyan Liu. The treatment of solvation by a generalized Born model and a self-consistent charge-density functional theory-based tight-binding method. Journal of Computational Chemistry, 23(15):1404–1415, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xiao:2008:UGD

[XLC08] Xuan Xiao, Wei-Zhong Lin, and Kuo-Chen Chou. Using grey dynamic modeling and pseudo amino acid composition to predict protein structural classes. Journal of Computational Chemistry, 29(12):2018–2024, September 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xiao:2002:DID

[XLL^{+}02] Jing-Fa Xiao, Ze-Sheng Li, Jing-Yao Liu, Li Sheng, and Chia-Chung Sun. DFT and ab initio direct dynamics studies on the hydrogen abstraction reactions of chlorine atoms with CH_{4−n}F_{n} (n = 1–3). Journal of Computational Chemistry, 23(15):1456–1465, November 30, 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xu:2007:MOA

[XLT07] Zhitao Xu, Harry H. Luo, and D. Peter Tieleman. Modifying the OPLS-AA force field to improve hydration free energies for several amino acid side chains using new atomic charges and an off-plane charge model for aromatic residues. Journal of Computational Chemistry, 28(3):689–697, February 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xiong:2008:DSP

[XLZ08] Ying Xiong, Hai-Ting Lu, and Chang-Guo Zhan. Dynamic structures of phosphodiesterase-5 active site by combined molecular dynamics simulations and hybrid quantum mechanical/molecular mechanical calculations. Journal of Computational Chemistry, 29(8):1259–1267, June 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xu:2000:UFS

[XOW^{+}00] Y. Z. Xu, Q. Ouyang, J. G. Wu, J. A. Yorke, G. X. Xu, D. F. Xu, R. D. Soloway, and J. Q. Ren. Using fractal to solve the multiple minima problem in molecular mechanics calculation. Journal of Computational Chemistry, 21(12):1101–1108, September 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xue:2009:SOS

[XPW09] Li-Qin Xue, Xian-Yong Pang, and Gui-Chang Wang. Selective oxidation of styrene on an oxygen-adsorbed Au(111): a density functional theory study. Journal of Computational Chemistry, 30(3):438–446, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xiao:2006:UPA

[XSHC06] Xuan Xiao, Shi-Huang Shao, Zheng-De Huang, and Kuo-Chen Chou. Using pseudo amino acid composition to predict protein structural classes: Approached with complexity measure factor. Journal of Computational Chemistry, 27(4):478–482, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xiao:2009:GCC

[XWC09] Xuan Xiao, Pu Wang, and Kuo-Chen Chou. GPCR-CA: a cellular automaton image approach for predicting G-protein– coupled receptor functional classes. Journal of Computational Chemistry, 30(9):1414–1423, July 15, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xu:2009:NED

[XWL^{+}09] Hong-Liang Xu, Fang-Fang Wang, Zhi-Ru Li, Bing-Qiang Wang, Di Wu, Wei Chen, Guang-Tao Yu, Feng Long Gu, and Yuriko Aoki. The nitrogen edge-doped effect on the static first hyperpolarizability of the supershort single-walled carbon nanotube. Journal of Computational Chemistry, 30(7): 1128–1134, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2008:EEA

[XWXC08] Hujun Xie, Ruibo Wu, Fei Xia, and Zexing Cao. Effects of electron attachment on C_{5′}O_{5′} and C_{1′}N_{1} bond cleavages of pyrimidine nucleotides: a theoretical study. Journal of Computational Chemistry, 29(12):2025–2032, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xi:2006:DGQ

[XYN^{+}06] Zhen Xi, Zhihong Yu, Congwei Niu, Shurong Ban, and Guangfu Yang. Development of a general quantum-chemical descriptor for steric effects: Density functional theory based QSAR study of herbicidal sulfonylurea analogues. Journal of Computational Chemistry, 27(13):1571–1576, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2004:HBS

[XZ04] Daiqian Xie and Jun Zeng. Hydrogen bonding and solvatochromatic shift of the lowest^{1} (n, π^{∗}) excitation of stetrazine in its hydrated clusters and dilute solutions. Journal of Computational Chemistry, 25(12):1487–1495, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xie:2005:EEG

[XZ05] Daiqian Xie and Jun Zeng. Electronic excitations of green fluorescent proteins: Protonation states of chromophore model compound in solutions. Journal of Computational Chemistry, 26(14):1487–1496, November 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Xiang:2004:FQM

[XZZ04] Yun Xiang, Da W. Zhang, and John Z. H. Zhang. Fully quantum mechanical energy optimization for protein–ligand structure. Journal of Computational Chemistry, 25(12):1431–1437, September 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yasar:2002:EMS

[YAÇ^{+}02] Fatih Yaşar, Handan Arkin, Tarik Çelik, Bernd A. Berg, and Hagai Meirovitch. Efficiency of the multicanonical simulation method as applied to peptides of increasing size: the heptapeptide deltorphin. Journal of Computational Chemistry, 23(12):1127–1134, September 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2004:DEG

[Yan04] Jinn-Moon Yang. Development and evaluation of a generic evolutionary method for protein–ligand docking. Journal of Computational Chemistry, 25(6):843–857, April 30, 2004. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yasuda:2008:TEI

[Yas08] Koji Yasuda. Two-electron integral evaluation on the graphics processor unit. Journal of Computational Chemistry, 29(3): 334–342, February 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yasar:2000:MPC

[YÇBM00] Fatih Yaşar, Tarik Çelik, Bernd A. Berg, and Hagai Meirovitch. Multicanonical procedure for continuum peptide models. Journal of Computational Chemistry, 21(14): 1251–1261, November 15, 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Chiu:2008:SEF

[yCkHmY08] Yi yuan Chiu, Jenn kang Hwang, and Jinn moon Yang. Soft energy function and generic evolutionary method for discriminating native from nonnative protein conformations. Journal of Computational Chemistry, 29(9):1364–1373, July 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2007:DLS

[YCS07] Xiaoli Yang, Wensheng Cai, and Xueguang Shao. A dynamic lattice searching method with constructed core for optimization of large Lennard-Jones clusters. Journal of Computational Chemistry, 28(8):1427–1433, June 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2009:PAC

[YCW^{+}09] Xue-Gang Yang, Duan Chen, Min Wang, Ying Xue, and Yu-Zong Chen. Prediction of antibacterial compounds by machine learning approaches. Journal of Computational Chemistry, 30 (8):1202–1211, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yokojima:2003:DMF

[YCXY03] Satoshi Yokojima, Guanhua Chen, Ruixue Xu, and Yijing Yan. A dynamic mean field theory for dissipative interacting many-electron systems: Markovian formalism and its implementation. Journal of Computational Chemistry, 24(16): 2083–2092, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2006:MSC

[YDWS06] Li-Ming Yang, Yi-Hong Ding, Qiang Wang, and Chia-Chung Sun. Monosilicon-substituted cyanoacetylene: a computational study. Journal of Computational Chemistry, 27(5):578– 595, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2005:TSE

[YFR05] Li Yang, Ji-Kang Feng, and Ai-Min Ren. Theoretical studies on the electronic and optical properties of two new alternating fluorene/carbazole copolymers. Journal of Computational Chemistry, 26(10):969–979, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yanover:2007:DEE

[YFS07] Chen Yanover, Menachem Fromer, and Julia M. Shifman. Dead-end elimination for multistate protein design. Journal of Computational Chemistry, 28(13):2122–2129, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yu:2006:MDS

[YGLvG06] Haibo Yu, Daan P. Geerke, Haiyan Liu, and Wilfred F. van Gunsteren. Molecular dynamics simulations of liquid methanol and methanol–water mixtures with polarizable models. Journal of Computational Chemistry, 27(13):1494–1504, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2005:MAO

[YGZZ05] Zhong-Zhi Yang, Li-Dong Gong, Dong-Xia Zhao, and Ming-Bo Zhang. Method and algorithm of obtaining the molecular intrinsic characteristic contours (MICCs) of organic molecules. Journal of Computational Chemistry, 26(1):35–47, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yoshida:2006:NMD

[YH06] Norio Yoshida and Fumio Hirata. A new method to determine electrostatic potential around a macromolecule in solution from molecular wave functions. Journal of Computational Chemistry, 27(4):453–462, March 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yoshizawa:2007:CFD

[YH07] Terutaka Yoshizawa and Masahiko Hada. Calculations of frequency-dependent molecular magnetizabilities with quasirelativistic time-dependent generalized unrestricted Hartree– Fock method. Journal of Computational Chemistry, 28(4): 740–747, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yoshizawa:2009:REC

[YH09] Terutaka Yoshizawa and Masahiko Hada. Relativistic and electron-correlation effects on magnetizabilities investigated by the Douglas-Kroll-Hess method and the second-order Møller–Plesset perturbation theory. Journal of Computational Chemistry, 30(15):2550–2566, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yu:2006:TSS

[YHD^{+}06] Guang-Tao Yu, Xu-Ri Huang, Yi-Hong Ding, Chia-Chung Sun, and Au-Chin Tang. Theoretical study on the [Si, C, N, O] potential energy surface. Journal of Computational Chemistry, 27(6):749–761, April 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yamanaka:2003:TSS

[YIN03] Masahiro Yamanaka, Akiko Inagaki, and Eiichi Nakamura. Theoretical studies on structures and reactivities of organocuprate(I) and organocopper(III) species. Journal of Computational Chemistry, 24(12):1401–1409, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yu:2006:WRD

[YJF06] Zhiyun Yu, Matthew P. Jacobson, and Richard A. Friesner. What role do surfaces play in GB models? A new-generation of surface-generalized Born model based on a novel Gaussian surface for biomolecules. Journal of Computational Chemistry, 27(1):72–89, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2000:FLD

[YK00] Jinn-Moon Yang and Cheng-Yan Kao. Flexible ligand docking using a robust evolutionary algorithm. Journal of Computational Chemistry, 21(11):988–998, August 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yim:2008:AMA

[YK08] Wai-Leung Yim and Thorsten Klüner. Atoms-in-molecules analysis for planewave DFT calculations — a numerical approach on a successively interpolated charge density grid. Journal of Computational Chemistry, 29(8):1306–1315, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yuzlenko:2009:MMA

[YKK09] Olga Yuzlenko and Katarzyna Kieć-Kononowicz. Molecular modeling of A_{1} and A_{2A} adenosine receptors: Comparison of rhodopsin- and β_{2}-adrenergic-based homology models through the docking studies. Journal of Computational Chemistry, 30 (1):14–32, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2006:GAP

[YL06] Yuedong Yang and Haiyan Liu. Genetic algorithms for protein conformation sampling and optimization in a discrete backbone dihedral angle space. Journal of Computational Chemistry, 27(13):1593–1602, October 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2009:SMN

[YL09] Pei-Kun Yang and Carmay Lim. Strategies to model the near-solute solvent molecular density/polarization. Journal of Computational Chemistry, 30(5):700–709, April 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yu:2009:CSR

[YLL^{+}09] Wenbo Yu, Lei Liang, Zijing Lin, Sanliang Ling, Maciej Haranczyk, and Maciej Gutowski. Comparison of some representative density functional theory and wave function theory methods for the studies of amino acids. Journal of Computational Chemistry, 30(4):589–600, March 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2008:TSR

[YLW^{+}08] Lei Yang, Jing-Yao Liu, Li Wang, Hong-Qing He, Ying Wang, and Ze-Sheng Li. Theoretical study of the reactions CF_{3}CH_{2}OCHF_{2} + OH/Cl and its product radicals and parent ether (CH_{3}CH_{2}OCH_{3}) with OH. Journal of Computational Chemistry, 29(4):550–561, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2009:TSR

[YLWL09] Lei Yang, Jing-Yao Liu, Su-Qin Wan, and Ze-Sheng Li. Theoretical studies of the reactions of CF_{3}CHCLOCHF_{2}/ CF_{3}CHFOCHF_{2} with OH radical and Cl atom and their product radicals with OH. Journal of Computational Chemistry, 30(4):565–580, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2008:DDS

[YLZ08] Jing Yang, Qian Shu Li, and Shaowen Zhang. Direct dynamics study on the reaction of acetaldehyde with ozone. Journal of Computational Chemistry, 29(2):247–255, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yuan:2004:SSV

[YMT04] Zheng Yuan, John S. Mattick, and Rohan D. Teasdale. SVM^{T M}: Support vector machines to predict transmembrane segments. Journal of Computational Chemistry, 25(5):632– 636, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yao:2005:CGR

[YNW05] Yu-Hua Yao, Xu-Ying Nan, and Tian-Ming Wang. A class of 2D graphical representations of RNA secondary structures and the analysis of similarity based on them. Journal of Computational Chemistry, 26(13):1339–1346, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2008:QCS

[YNZ^{+}08] Wei Yang, Hugh Nymeyer, Huan-Xiang Zhou, Bernd Berg, and Rafael Brüschweiler. Quantitative computer simulations of biomolecules: a snapshot. Journal of Computational Chemistry, 29(4):668–672, March 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yakovenko:2008:KAC

[YOB^{+}08] Olexander Yakovenko, Alexander A. Oliferenko, Volodymyr G. Bdzhola, Vladimir A. Palyulin, and Nikolai S. Zefirov. Kirchhoff atomic charges fitted to multipole moments: Implementation for a virtual screening system. Journal of Computational Chemistry, 29(8):1332–1343, June 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yoshioki:2002:DPW

[Yos02] Shuzo Yoshioki. Dynamics of a protein and water molecules surrounding the protein: Hydrogen-bonding between vibrating water molecules and a fluctuating protein. Journal of Computational Chemistry, 23(3):402–413, February 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Young:2011:CCP

[You11] David C. Young. Computational Chemistry: a Practical Guide for Applying Techniques to Real-World Problems. John Wiley, New York, NY, USA, 2011. ISBN 0-470-88008-2. xxiv + 381 pp. LCCN QD39.3.M3.

Yoon:2009:MMM

[YPNE09] Gwonchan Yoon, Hyeong-Jin Park, Sungsoo Na, and Kilho Eom. Mesoscopic model for mechanical characterization of biological protein materials. Journal of Computational Chemistry, 30(6):873–880, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2009:TSG

[YQQH09] Hua-Qing Yang, Song Qin, Song Qin, and Chang-Wei Hu. Theoretical study on the gas-phase reaction mechanism between nickel monoxide and methane for syngas production. Journal of Computational Chemistry, 30(6):847–863, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ye:2000:DAC

[YS00] Yuan-Jie Ye and Ling-Ling Shen. DFT approach to calculate electronic transfer through a segment of DNA double helix. Journal of Computational Chemistry, 21(12):1109–1117, September 2000. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yoshida:2003:MMM

[YSA^{+}03] Takashi Yoshida, Kazuhisa Sakakibara, Masatoshi Asami, Kuo-Hsiang Chen, Jenn-Huei Lii, and Norman L. Allinger.

Molecular mechanics (MM3) calculations on lithium amide compounds. Journal of Computational Chemistry, 24(3):319– 327, February 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yan:2009:EED

[YSJ09] Liuming Yan, Changle Shao, and Xiaobo Ji. Evaluation of electroosmotic drag coefficient of water in hydrated sodium perfluorosulfonate electrolyte polymer. Journal of Computational Chemistry, 30(9):1361–1370, July 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yamabe:2003:CSI

[YT03] Shinichi Yamabe and Noriko Tsuchida. A computational study of interactions between acetic acid and water molecules. Journal of Computational Chemistry, 24(8):939–947, June 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yamabe:2004:CSR

[YT04] Shinichi Yamabe and Noriko Tsuchida. A computational study of the role of hydrogen bonds in S_{N}1 and E1 reactions. Journal of Computational Chemistry, 25(4):598–608, March 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yanagisawa:2001:IUD

[YTH01] Susumu Yanagisawa, Takao Tsuneda, and Kimihiko Hirao. Investigation of the use of density functionals in second- and third-row transition metal dimer calculations. Journal of Computational Chemistry, 22(16):1995–2009, December 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yuki:2007:IIM

[YTH^{+}07] Hitomi Yuki, Yoshikazu Tanaka, Masayuki Hata, Hidenori Ishikawa, Saburo Neya, and Tyuji Hoshino. Implementation of π-π interactions in molecular dynamics simulation. Journal of Computational Chemistry, 28(6):1091–1099, April 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yamabe:2007:TSR

[YTY07] Shinichi Yamabe, Noriko Tsuchida, and Shoko Yamazaki. Theoretical study of the role of solvent H_{2}O in neopentyl and pinacol rearrangements. Journal of Computational Chemistry, 28(9):1561–1571, July 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2004:MWI

[YWH04] Feng Yang, Zhen-Dong Wang, and Yun-Ping Huang. Modification of the Wiener index 4. Journal of Computational Chemistry, 25(6):881–887, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2003:NTI

[YWHZ03] Feng Yang, Zhen-Dong Wang, Yun-Ping Huang, and Hai-Liang Zhu. Novel topological index F based on incidence matrix. Journal of Computational Chemistry, 24(14):1812– 1820, November 15, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2007:FEP

[YXC^{+}07] Sheng-Yong Yang, Ming-Li Xiang, Li-Juan Chen, Guo-Bin Xie, Bing Shi, Yu-Quan Wei, and Tom Ziegler. Free energy profiles for monomer capture in Grubbs- and SHOP-type olefin polymerization catalysts: a constraint ab initio molecular dynamics study. Journal of Computational Chemistry, 28(2): 513–518, January 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yu:2009:ECS

[YXL^{+}09] Wenbo Yu, Xuee Xu, Hongbao Li, Rui Pang, Kun Fang, and Zijing Lin. Extensive conformational searches of 13 representative dipeptides and an efficient method for dipeptide structure determinations based on amino acid conformers. Journal of Computational Chemistry, 30(13):2105–2121, October 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yan:2004:TSS

[YXZ^{+}04] Weizhong Yan, Ying Xue, Hua Zhu, Jun Zeng, and Daiqian Xie. A theoretical study of solvent effects on tautomerism and electronic absorption spectra of 3-hydroxy-2-mercaptopyridine and 2,3-dihydroxypyridine. Journal of Computational Chemistry, 25(15):1833–1839, November 30, 2004. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yu:2007:PRI

[YYW07] Xinliang Yu, Bing Yi, and Xueye Wang. Prediction of refractive index of vinyl polymers by using density functional theory. Journal of Computational Chemistry, 28(14):2336– 2341, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ytreberg:2004:EUN

[YZ04] F. Marty Ytreberg and Daniel M. Zuckerman. Efficient use of nonequilibrium measurement to estimate free energy differences for molecular systems. Journal of Computational Chemistry, 25(14):1749–1759, November 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:2006:SPC

[YZ06] Zhong-Zhi Yang and Qiang Zhang. Study of peptide conformation in terms of the ABEEM/MM method. Journal of Computational Chemistry, 27(1):1–10, January 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2007:SOS

[ZA07] Fangqiang Zhu and Dimitris K. Agrafiotis. Self-organizing superimposition algorithm for conformational sampling. Journal of Computational Chemistry, 28(7):1234–1239, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zanuy:2003:TIP

[ZALMG03] David Zanuy, Carlos Alemán, Manuel Laso, and Sebastián Muñoz-Guerra. Thermally induced phase transition in helical comblike poly(β-peptide)s: an atomistic simulation. Journal of Computational Chemistry, 24(6):770–778, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zheng:2007:CSS

[ZAT07] Jingjing Zheng, Ahmet Altun, and Walter Thiel. Common system setup for the entire catalytic cycle of cytochrome P450_{cam} in quantum mechanical/molecular mechanical studies. Journal of Computational Chemistry, 28(13):2147–2158, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zubarev:2007:CAC

[ZB07] Dmitry Yu. Zubarev and Alexander I. Boldyrev. Comprehensive analysis of chemical bonding in boron clusters. Journal of Computational Chemistry, 28(1):251–268, January 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2003:CIS

[ZBS03] Qing Zhang, Daniel A. Beard, and Tamar Schlick. Constructing irregular surfaces to enclose macromolecular complexes for mesoscale modeling using the discrete surface charge optimization (DISCO) algorithm. Journal of Computational Chemistry, 24(16):2063–2074, December 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhan:2009:CPN

[ZCL09] Lixin Zhan, Jeff Z. Y. Chen, and Wing-Ki Liu. Comparison of predicted native structures of Met-enkephalin based on various accessible-surface-area solvent models. Journal of Computational Chemistry, 30(7):1051–1058, May 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zarzycki:2004:SPA

[ZCS04] P. Zarzycki, R. Charmas, and P. Szabelski. Study of proton adsorption at heterogeneous oxide/electrolyte interface. Prediction of the surface potential using Monte Carlo simulations and 1-pK approach. Journal of Computational Chemistry, 25 (5):704–711, April 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2003:MCF

[ZCZ03] D. W. Zhang, X. H. Chen, and J. Z. H. Zhang. Molecular caps for full quantum mechanical computation of peptide–water interaction energy. Journal of Computational Chemistry, 24(15): 1846–1852, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhan:2005:FPS

[ZDS^{+}05] Chang-Guo Zhan, Shi-Xian Deng, Jaime G. Skiba, Beth A. Hayes, Sarah M. Tschampel, George C. Shields, and Donald W. Landry. First-principle studies of intermolecular and intramolecular catalysis of protonated cocaine. Journal of Computational Chemistry, 26(10):980–986, July 30, 2005. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zerara:2008:PCP

[Zer08] Mohamed Zerara. pyVib, a computer program for the analysis of infrared and Raman optical activity. Journal of Computational Chemistry, 29(2):306–311, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2005:CMT

[ZFL^{+}05] Quan Zhu, Ke-Xiang Fu, Xiang-Yuan Li, Zhen Gong, and Jian-Yi Ma. Continuous medium theory for nonequilibrium solvation: III. Solvation shift by monopole approximation and multipole expansion in spherical cavity. Journal of Computational Chemistry, 26(4):399–409, March 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2008:HAB

[ZFW08] Y. C. Zhou, Michael Feig, and G. W. Wei. Highly accurate biomolecular electrostatics in continuum dielectric environments. Journal of Computational Chemistry, 29(1):87–97, January 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2001:SMP

[ZGFL01] Linda Yu Zhang, Emilio Gallicchio, Richard A. Friesner, and Ronald M. Levy. Solvent models for protein–ligand binding: Comparison of implicit solvent Poisson and surface generalized Born models with explicit solvent simulations. Journal of Computational Chemistry, 22(6):591–607, April 30, 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2006:IPE

[ZGXX06] Hua Zhu, Yong Guo, Ying Xue, and Daiqian Xie. Ab initio potential energy surface and predicted microwave spectra for Ar − OCS dimer and structures of Ar_{n} − OCS (n = 2–14) clusters. Journal of Computational Chemistry, 27(9):1045– 1053, July 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2007:CIS

[ZGZX07] Yuanqiang Zhu, Yong Guo, Lidong Zhang, and Daiqian Xie. A computational investigation on the sequential rearrangement mechanism of 2-allyl-2,4,5-hexatrienaldehyde involving [1,5]-hydrogen migration and 8π-electrocyclization. Journal of Computational Chemistry, 28(13):2164–2169, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhao:2008:TDD

[ZH08] Guang-Jiu Zhao and Ke-Li Han. Time-dependent density functional theory study on hydrogen-bonded intramolecular charge-transfer excited state of 4-dimethylamino-benzonitrile in methanol. Journal of Computational Chemistry, 29(12): 2010–2017, September 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zou:2009:HPQ

[ZHH09] Dongsheng Zou, Zhongshi He, and Jingyuan He. β-hairpin prediction with quadratic discriminant analysis using diversity measure. Journal of Computational Chemistry, 30(14):2277– 2284, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhao:2009:FPI

[ZHMW09] Erjun Zhao, Bo Hong, Jian Meng, and Zhijian Wu. First principles investigation on the ultra-incompressible and hard TaN. Journal of Computational Chemistry, 30(14):2358–2363, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2006:RCP

[Zho06] Shiqi Zhou. Rapidly convergent procedure to solve the density profile equation in the classical density functional theory. Journal of Computational Chemistry, 27(8):941–947, June 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zeng:2007:EAH

[ZJM^{+}07] T. Zeng, Z. Jamshidi, H. Mori, E. Miyoshi, and M. Klobukowski. Electron affinities of heavier phosphoryl and thiophosphoryl halides APX_{3} (A = O, S and X = Br, I). Journal of Computational Chemistry, 28(12):2027–2033, September 2007. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhao:2007:CDQ

[ZKZ^{+}07] Yan-Li Zhao, Wei Kan, Hua Zhong, Hai-Tao Yu, and Hong-Gang Fu. Combined DFT, QCISD(T), and G2 mechanism investigation for the reactions of carbon monophosphide CP with unsaturated hydrocarbons allene CH_{2}CCH_{2} and methylacetylene CH_{3}CCH. Journal of Computational Chemistry, 28 (7):1221–1233, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zou:2005:ETS

[ZL05] Wenli Zou and Wenjian Liu. Extensive theoretical studies on the low-lying electronic states of indium monochloride cation, InCl^{+}. Journal of Computational Chemistry, 26(1):106–113, January 15, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zou:2007:TSL

[ZL07] Wenli Zou and Wenjian Liu. Theoretical study on the lowlying electronic states of NiH and NiAt. Journal of Computational Chemistry, 28(14):2286–2298, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2009:NME

[ZL09a] Xiaolei Zhu and Luhua Lai. A novel method for enzyme design. Journal of Computational Chemistry, 30(2):256–267, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zou:2009:CIC

[ZL09b] Wenli Zou and Wenjian Liu. Comprehensive ab initio calculation and simulation on the low-lying electronic states of TlX (X = F, Cl, Br, I, and At). Journal of Computational Chemistry, 30(4):524–539, March 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zheng:2009:STC

[ZLD09] Anmin Zheng, Shang-Bin Liu, and Feng Deng. ^{13}C shielding tensors of crystalline amino acids and peptides: Theoretical predictions based on periodic structure models. Journal of Computational Chemistry, 30(2):222–235, January 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zacharias:2003:PSM

[ZLJS03] Carlos Renato Zacharias, Maurício Ruv Lemes, Arnaldo Dal Pino Júnior, and David Santo Orcero. Predicting structural models for silicon clusters. Journal of Computational Chemistry, 24(7):869–875, May 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2004:TSR

[ZLLS04a] Jia-Xu Zhang, Jing-Yao Liu, Ze-Sheng Li, and Chia-Chung Sun. Theoretical study on reaction mechanism of the fluoromethylene radical with nitrogen dioxide. Journal of Computational Chemistry, 25(15):1888–1894, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2004:TSM

[ZLLS04b] Jia-Xu Zhang, Jing-Yao Liu, Ze-Sheng Li, and Chia-Chung Sun. Theoretical study on the mechanism of the ^{1}CHCl + NO_{2} reactions. Journal of Computational Chemistry, 25(9): 1184–1190, July 15, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2005:TSR

[ZLLS05] Jia-Xu Zhang, Jing-Yao Liu, Ze-Sheng Li, and Chia-Chung Sun. Theoretical study on the reaction mechanism of the methyl radical with nitrogen oxides. Journal of Computational Chemistry, 26(8):807–817, June 2005. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2006:TMS

[ZLLS06a] Jia-Xu Zhang, Ze-Sheng Li, Jing-Yao Liu, and Chia-Chung Sun. Theoretical mechanistic study on the radical-molecule reaction of CHCl_{2} /CCl_{3} with NO_{2}. Journal of Computational Chemistry, 27(5):661–671, April 15, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2006:TSM

[ZLLS06b] Jia-Xu Zhang, Ze-Sheng Li, Jing-Yao Liu, and Chia-Chung Sun. Theoretical study on the mechanism of the CH_{2}F + NO_{2} reaction. Journal of Computational Chemistry, 27(7):894–905, May 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2007:NID

[ZLY07] Xizhen Zhang, Jiawei Luo, and Li Yang. New invariant of DNA sequence based on 3DD-curves and its application on phylogeny. Journal of Computational Chemistry, 28(14):2342– 2346, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zyubin:2003:PTD

[ZM03] A. S. Zyubin and A. M. Mebel. Performance of timedependent density functional and Green functions methods for calculations of excitation energies in radicals and for Rydberg electronic states. Journal of Computational Chemistry, 24(6): 692–700, April 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zoete:2006:IIS

[ZM06] Vincent Zoete and Markus Meuwly. Importance of individual side chains for the stability of a protein fold: Computational alanine scanning of the insulin monomer. Journal of Computational Chemistry, 27(15):1843–1857, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zyubin:2009:QCM

[ZMH^{+}09] A. S. Zyubin, A. M. Mebel, M. Hayashi, H. C. Chang, and S. H. Lin. Quantum chemical modeling of photoadsorption properties of the nitrogen-vacancy point defect in diamond. Journal of Computational Chemistry, 30(1):119–131, January 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2009:AFE

[ZMZ09] Xin Zhang, Artem B. Mamonov, and Daniel M. Zuckerman. Absolute free energies estimated by combining precalculated molecular fragment libraries. Journal of Computational Chemistry, 30(11):1680–1691, August 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2007:NSE

[ZNLL07] Congyi Zhou, Changming Nie, Shan Li, and Zhonghai Li. A novel semi-empirical topological descriptor Nt and the application to study on QSPR/QSAR. Journal of Computational Chemistry, 28(15):2413–2423, November 30, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2006:CSG

[ZOJ^{+}06] Yu Zhou, Chris Oostenbrink, Aldo Jongejan, Wilfred F. Van Gunsteren, Wilfred R. Hagen, Simon W. De Leeuw, and Jaap A. Jongejan. Computational study of ground-state chiral induction in small peptides: Comparison of the relative stability of selected amino acid dimers and oligomers in homochiral and heterochiral combinations. Journal of Computational Chemistry, 27(7):857–867, May 2006. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zagrovic:2003:SVD

[ZP03] Bojan Zagrovic and Vijay Pande. Solvent viscosity dependence of the folding rate of a small protein: Distributed computing study. Journal of Computational Chemistry, 24(12): 1432–1436, September 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zanasi:2007:CDN

[ZPL07] R. Zanasi, S. Pelloni, and P. Lazzeretti. Chiral discrimination via nuclear magnetic shielding polarisabilities from NMR spectroscopy: Theoretical study of (R_{a})-1,3-dimethylallene, (2 R)-2-methyloxirane, and (2 R)- N-methyloxaziridine. Journal of Computational Chemistry, 28(13):2159–2163, October 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2004:SCA

[ZS04] Yang Zhang and Jeffrey Skolnick. SPICKER: a clustering approach to identify near-native protein folds. Journal of Computational Chemistry, 25(6):865–871, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zarzycki:2005:RSH

[ZSC05] Piotr Zarzycki, Paweł Szabelski, and Robert Charmas. Role of the surface heterogeneity in adsorption of hydrogen ions on metal oxides: Theory and simulations. Journal of Computational Chemistry, 26(10):1079–1088, July 30, 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zamora:2008:RIC

[ZSE08] Miguel A. Zamora, Fernando D. Suvire, and Ricardo D. Enriz. Ring inversion in 1,4,7 cyclononatriene and analogues: ab initio and DFT calculations and topological analysis. Journal of Computational Chemistry, 29(2):280–290, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zwier:2007:HMD

[ZSK07] Matthew C. Zwier, Justin M. Shorb, and Brent P. Krueger. Hybrid molecular dynamics-quantum mechanics simulations of solute spectral properties in the condensed phase: Evaluation of simulation parameters. Journal of Computational Chemistry, 28(9):1572–1581, July 15, 2007. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zotti:2008:ABF

[ZTP^{+}08] L. A. Zotti, G. Teobaldi, K. Palotás, W. Ji, H.-J. Gao, and W. A. Hofer. Adsorption of benzene, fluorobenzene and metadi-fluorobenzene on Cu(110): a computational study. Journal of Computational Chemistry, 29(10):1589–1595, July 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2009:DNI

[ZTS09] Peng Zhou, Feifei Tian, and Zhicai Shang. 2D depiction of nonbonding interactions for protein complexes. Journal of Computational Chemistry, 30(6):940–951, April 30, 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2009:EIM

[ZW09] Yan Zhang and Chang-Sheng Wang. Estimation on the intramolecular 10-membered ring N H⋯O C hydrogen-bonding energies in glycine and alanine peptides. Journal of Computational Chemistry, 30(8):1251–1260, June 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2009:CAC

[ZWB09] Zhigang Zhou, Yanli Wang, and Stephen H. Bryant. Computational analysis of the cathepsin B inhibitors activities through LR-MMPBSA binding affinity calculation based on docked complex. Journal of Computational Chemistry, 30(14):2165– 2175, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2009:FFC

[ZWC^{+}09] Chenggang Zhou, Jinping Wu, Liang Chen, Yang Wang, Hansong Cheng, and Robert C. Forrey. Force field for copper clusters and nanoparticles. Journal of Computational Chemistry, 30(14):2255–2266, November 15, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2005:DLD

[ZWL^{+}05] Hui Zhang, Jia-Yan Wu, Ze-Sheng Li, Jing-Yao Liu, Li Sheng, and Chia-Chung Sun. Dual-level direct dynamics studies on the reaction Cl + CHBr_{2}Cl. Journal of Computational Chemistry, 26(13):1421–1426, October 2005. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhong:2008:TSP

[ZWP08] Yang Zhong, G. Lee Warren, and Sandeep Patel. Thermodynamic and structural properties of methanol–water solutions using nonadditive interaction models. Journal of Computational Chemistry, 29(7):1142–1152, May 2008. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zicovich-Wilson:2004:CVF

[ZWPR^{+}04] C. M. Zicovich-Wilson, F. Pascale, C. Roetti, V. R. Saunders, R. Orlando, and R. Dovesi. Calculation of the vibration frequencies of α-quartz: the effect of Hamiltonian and basis set. Journal of Computational Chemistry, 25(15):1873– 1881, November 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zheng:2002:SSE

[ZWS^{+}02] Kangcheng Zheng, Juping Wang, Yong Shen, Wenlie Peng, and Fengcun Yun. Studies on 4, 7-di-substitution effects of one ligand in [Ru(Phen)_{3}]^{2} with DFT method. Journal of Computational Chemistry, 23(4):436–443, March 2002. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2009:MDS

[ZWS^{+}09] Xin Zhang, Ruibo Wu, Lingchun Song, Yuchun Lin, Menghai Lin, Zexing Cao, Wei Wu, and Yirong Mo. Molecular dynamics simulations of the detoxification of paraoxon catalyzed by phosphotriesterase. Journal of Computational Chemistry, 30 (15):2388–2401, November 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zicovich-Wilson:2008:ISI

[ZWTP^{+}08] C. M. Zicovich-Wilson, F. J. Torres, F. Pascale, L. Valenzano, R. Orlando, and R. Dovesi. Ab initio simulation of the IR spectra of pyrope, grossular, and andradite. Journal of Computational Chemistry, 29(13):2268–2278, October 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zheng:2009:DSC

[ZWY^{+}09] Xiaoyan Zheng, Xueye Wang, Shanfeng Yi, Nuanqing Wang, and Yueming Peng. DFT study of the carbon- and nitrogenpivot lariat crown ethers and their complexes with alkali metal cations: Na^{+}, K^{+}. Journal of Computational Chemistry, 30 (16):2674–2683, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zheng:2009:ITS

[ZWZ09] Han Zheng, Shenglong Wang, and Yingkai Zhang. Increasing the time step with mass scaling in Born–Oppenheimer ab initio QM/MM molecular dynamics simulations. Journal of Computational Chemistry, 30(16):2706–2711, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zeng:2004:HBS

[ZX04] Jun Zeng and Daiqian Xie. Hydrogen bonding and solvent effects on the lowest ^{1}(n, π^{∗}) excitations of triazines in water. Journal of Computational Chemistry, 25(6):813–822, April 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2008:ISE

[ZX08] Weihua Zhu and Heming Xiao. Ab initio study of electronic structure and optical properties of heavy-metal azides: TlN_{3}, AgN_{3}, and CuN_{3}. Journal of Computational Chemistry, 29(2): 176–184, January 30, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2009:SPH

[ZX09] Hua Zhu and Daiqian Xie. N_{2}O in small para-hydrogen clusters: Structures and energetics. Journal of Computational Chemistry, 30(6):841–846, April 30, 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhao:2004:ACT

[ZXL^{+}04] Yi Zhao, Wenguo Xu, Qianshu Li, Yaoming Xie, and Henry F. Schaefer III. The arsenic clusters As_{n} (n = 1–5) and their anions: Structures, thermochemistry, and electron affinities. Journal of Computational Chemistry, 25(7):907–920, May 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2003:TSS

[ZXY03] Hua Zhu, Daiqian Xie, and Guosen Yan. Theoretical studies for structures and energetics of Rg_{n}N_{2}O (Rg = He, Ne, Ar) clusters. Journal of Computational Chemistry, 24(15):1839– 1845, November 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2008:SIG

[ZXY08] Ying Zhang, Xin Xu, and Yijing Yan. Systematic investigation on the geometric dependence of the calculated nuclear magnetic shielding constants. Journal of Computational Chemistry, 29(11):1798–1807, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2009:FPC

[ZXYF09] Guo-Xu Zhang, Ying Xie, Hai-Tao Yu, and Hong-Gang Fu. First-principles calculations of the stability and electronic properties of the PbTiO_{3} (110) polar surface. Journal of Computational Chemistry, 30(12):1785–1798, September 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2001:CRE

[ZY01] Yong Zhang and Xiao-Zeng You. Charge redistribution effect on the properties of charge transfer complexes H_{n}R·XY and H_{n}R·X_{2} (X, Y = F, Cl, Br, I; R = O, S, N, P). Journal of Computational Chemistry, 22(3):327–338, February 2001. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2008:REE

[ZZ08] Naigong Zhang and Chen Zeng. Reference energy extremal optimization: a stochastic search algorithm applied to computational protein design. Journal of Computational Chemistry, 29(11):1762–1771, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2004:DLI

[ZZL04] Yue Zhang, Shaowen Zhang, and Qian Shu Li. A dual-level ab initio and hybrid density functional theory dynamics study on the unimolecular decomposition reaction C_{2}H_{5}O → CH_{2}O + CH_{3}. Journal of Computational Chemistry, 25(2):218–226, January 30, 2004. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2008:TSO

[ZZL^{+}08] Hui Zhang, Gui-Ling Zhang, Jing-Yao Liu, Bo Liu, Xiao-Yang Yu, and Ze-Sheng Li. Theoretical study on the OH + CH_{3}NHC(O)OCH_{3} reaction. Journal of Computational Chemistry, 29(7):1170–1176, May 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2009:TSR

[ZZL^{+}09] Hui Zhang, Gui-Ling Zhang, Jing-Yan Liu, Miao Sun, Bo Liu, and Ze-Sheng Li. Theoretical study on the reaction of SiH(CH_{3})_{3} with SiH_{3} radical. Journal of Computational Chemistry, 30(2):236–242, January 30, 2009. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2007:SEG

[ZZS^{+}07] Guiling Zhang, Hui Zhang, Miao Sun, Yanhong Liu, Xiaohong Pang, Xiaoyang Yu, Bo Liu, and Zesheng Li. Substitution effect on the geometry and electronic structure of the ferrocene. Journal of Computational Chemistry, 28(14):2260– 2274, November 15, 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2009:GSB

[ZZTS09] Peng Zhou, Jianwei Zou, Feifei Tian, and Zhicai Shang. Geometric similarity between protein–RNA interfaces. Journal of Computational Chemistry, 30(16):2738–2751, December 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhou:2008:ICS

[ZZvRSC08] Zhen Zhou, Jijun Zhao, Paul von Ragué Schleyer, and Zhongfang Chen. Insertion of C_{50} into single-walled carbon nanotubes: Selectivity in interwall spacing and C_{50} isomers. Journal of Computational Chemistry, 29(5):781–787, April 15, 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2007:TSB

[ZZW^{+}07] Hui Zhang, Gui-Ling Zhang, Li Wang, Bo Liu, Xiao-Yang Yu, and Ze-Sheng Li. Theoretical study on the Br + CH_{3}SCH_{3} reaction. Journal of Computational Chemistry, 28(7):1153– 1159, May 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2009:TUV

[ZZW09] Hao Zhang, Yun-Hong Zhang, and Feng Wang. Theoretical understanding on the v_{1} -SO band perturbed by the formation of magnesium sulfate ion pairs. Journal of Computational Chemistry, 30(3):493–503, February 2009. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2007:EPW

[ZZY07] Tian Zhang, Haoping Zheng, and Shu Yan. Equivalent potential of water molecules for electronic structure of glutamic acid. Journal of Computational Chemistry, 28(11):1848–1857, August 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2008:EPW

[ZZY08] Tian Zhang, Haoping Zheng, and Shu Yan. The equivalent potential of water for electronic structure of aspartic acid. Journal of Computational Chemistry, 29(11):1780–1787, August 2008. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhang:2006:ISI

[ZZZ^{+}06] Jian Zhang, Mingyu Zhang, Yuanyuan Zhao, Baoguo Chen, and Chia-Chung Sun. Ab initio studies on isomers of macropolyhedral borane ions [B_{20}H_{18}]^{n} (n = 0, −2, −4). Journal of Computational Chemistry, 27(15):1817–1821, November 30, 2006. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).